A Complete Bibliography of Publications in the Journal of Mathematical Biology

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

07 May 2024 Version 1.31

Title word cross-reference

(n, 2) [2872]. (τ, α, ρ) [32]. 1 [1159]. 2 [570, 3026, 2660, 3031, 70]. 2: 1 [890]. 2n = 40 [2815]. 3 [2779, 3010, 1707, 1368, 2948]. ^{+} [3192, 1836]. ^{2+} [1413]. _{2} [3396]. _{V} [3477]. A [1340]. Aβ [3158]. B [1340]. B_{2} [3236]. β [2197, 3532, 1176, 2692, 2767, 1849]. d [2657]. F [3320]. F_{st} [1921]. G · U [1402]. i [3559]. k [2294]. K^{+} [1495]. κ [2358]. L^{∞} [2509]. λ [246]. λ = α_{1} + (α_{2} + α_{3}λ)e^{−λ} [2489]. m [3425]. M^{5} [1609]. N [3531, 933, 840, 482, 418, 481, 480, 581]. n: m [315]. ω [246]. p [3615, 2813]. q [3361]. R^{3} [419]. R^{4} [419]. R_{0} [2060, 784, 2942, 2395, 2127, 1430]. R_{5} [2745]. T [2126, 2557]. G [3480]. × [1307]. Daphnia [2662]. X [2713, 833]. X_{4} [2745].

-actin [3320]. -amyloid [2197]. -ary [3361]. -catenin [2692, 2767, 1849]. -cells [1176]. -component [570]. -D [1159]. -dimensional [840, 70, 1368]. -hydroxypropionic [2779]. -level [933]. -linked [833]. -mer [2294]. -mers [2294]. -order [3425]. -parameter [3026]. -periodic [2557]. -person [482].

-stable [2657]. -state [3031]. -structures [2660]. -th [3559, 2813]. -trees [2713]. -type [418, 481, 480, 581]. -values [3615].

/CD8 [3192].

1 [3073, 3581, 1393, 2005, 2187, 2560, 2007]. 1.1/ [3477]. 111 [2760]. 19 [3405, 3223, 3359, 3296, 3242, 3430, 3243, 3109, 3448, 3489, 3466].

2 [3476, 3507]. 2D [2343, 2417].

60th [2485]. 65th [2992, 2109].

Abdurahman [1834]. aberrant [3585, 3387]. Abrupt [2644, 1372]. absence [2460]. Absolute [2262, 3369, 3607, 3110]. absorbing [68]. absorption [2943, 1417, 143, 360]. abundance [3347, 3565, 2603, 1949, 3559, 3082]. abundancy [3343]. abundant [2802, 3559]. Accelerating [2734]. acceleration [2183, 2117]. access [939, 3246]. Accessibility [3431]. according [2645, 251, 3006]. account [3014]. accounting [1577]. accretion [990]. accumulation [3216, 1024]. accumulations [122]. accuracy [2906, 2715]. Accurate [336, 1870]. acetone [1983]. achieve [3061, 2965]. acid [2779]. Acidithiobacillus [2293]. acidity [2254]. acidosis [1694]. acquired [743]. AcrAB [3150]. AcrAB-TolC [3150]. acrosomal [387, 582]. across [3590, 1006, 253, 3054, 2175]. act [2424]. Actin [928, 2978, 1776, 1308, 2106, 1984, 3130, 3320]. actin-filament [1308]. action [151, 2043, 1339, 2052, 2096, 2346, 834, 2894, 2227]. Activation [927, 947, 975, 3131, 2043, 771, 3170, 2981, 2064, 2341, 1860]. activator [102, 1028]. activator-inhibitor [102, 1028]. activator-inhibitor/immobilizer [1028]. active [567, 3267, 2560]. activities [3161]. activity [772, 903, 1176, 1826, 2967, 3279, 125, 3064, 792, 1860, 1149]. activity-dependent [1149]. acto [2641]. acto-myosin [2641]. actomyosin [2271, 3320]. acuity [3116]. acute [2412, 2968, 1870]. acyclic [3558]. Adaptation [3551, 3202, 1194, 2253, 2255, 860, 1053, 676, 2322, 385, 616, 2474, 386]. Adaptive [1253, 2308, 1633, 2025, 1726, 2835, 1883, 2495, 3024, 2988, 2389, 2510, 2439, 2497, 1437, 1851, 1673, 2168, 2159, 2153, 832, 2124, 2375, 2972, 3423, 2498, 3570, 1547, 2166, 2642, 1970, 1095, 3248, 1982]. adaptively [2635]. added [2039]. addicts [1347]. additional [1812, 1078, 1889, 1866, 3541]. additive [1018, 1078, 117, 809, 597, 1894, 270]. additivity [1253]. ADE [3470]. adherent [2240]. adhesion [1791, 2724, 922, 2971, 2573, 1780, 2647, 2223, 2564, 2248, 3620]. adhesion-mediated [922]. Adiabatic [2250]. adipocyte [3587]. Adjusting [3537]. adjuvant [1447]. administration [1073]. admixed [3112]. adrenal [1972]. adult [826, 1606, 2488]. adults [2131]. advance [656, 619]. advance-delay-differential [619]. Advanced [32]. advancing [3539]. advantage [934]. advantageous [317, 474, 2982]. advantages [1546]. Advection [1180, 3250, 1750, 493, 3372, 1187, 2484, 2070, 2910, 3299, 3377]. advection-diffusion [3372]. advective [2320, 3310, 3117]. adversity [2887]. Aedes [2750, 2770, 2067]. aegypti [2750, 2770]. Aerodynamics [585]. aeruginosa [1541]. affect [1552]. affected [2640, 2295]. affecting [764]. Affine [1821, 2324, 1567, 1562, 2163]. affinity [11, 297]. Africa [3476]. African [1758]. after [1406, 3316, 2389, 3198]. Age [349, 1632, 432, 1361, 3163, 3467, 708, 971, 1653, 1737, 564, 547, 1851, 733, 2470, 774, 1211, 150, 174, 129, 362, 503, 982, 1896, 2536, 200, 343, 543, 338, 1734, 1049, 1918, 2456, 3406, 1639, 3164, 3549, 327, 2697, 1204, 1338, 2388, 786, 2814, 2685, 3617, 1948, 1017, 3194, 3529, 3257, 820, 3337, 693, 4, 229, 311, 3345, 632, 2057, 2705, 383, 1392, 1809, 805, 641, 963, 3278, 573, 1451, 273, 994, 2143, 3270, 236, 1614, 234, 888, 2982, 53, 736, 1876, 2048, 483, 680, 3535, 50, 1084, 1314, 1463, 3621, 2571, 2755, 1290, 1555]. age-and-cyclin-structured [1737]. age-class-structured [2982]. age-dependence [820]. Age-dependent [349, 564, 547, 1851, 774, 200, 543, 3406, 3549, 2388, 1948, 3337, 693, 2057, 383, 641, 2143, 236, 234, 53, 736, 1876, 1290]. age-distributions [994]. age-of-infection [3467]. age-specific [273]. Age-structured [1632, 1361, 708, 971, 1653, 1211, 503, 982, 1896, 2536, 1049, 1918, 1639, 3164, 327, 1204, 1338, 786, 2814, 3617, 3194, 3529, 3257, 229, 311, 3345, 632, 2705, 1809, 805, 3278, 1451, 3270, 1614, 2048, 680, 1084, 1314, 1463, 3621, 2755, 1555]. age/stage [3465]. ageing [2947]. agent [3226, 2764, 3498, 3341, 3599, 1792]. agent-based [3226, 2764, 3341, 1792]. Agents [1554]. ages [3584, 2576]. aggregate [2501]. aggregate-level [2501]. aggregated [708, 1260]. aggregates [3212, 2815, 2621, 2513, 3266, 1573]. aggregating [1177]. Aggregation [370, 1200, 2197, 1096, 1791, 1840, 2602, 2425, 705, 3344, 2303, 1399, 714, 2394, 2676, 2857, 1419, 927, 1487, 1141, 928]. aggregations [2041]. Aggregative [493, 819]. aging [60, 1031]. agr [1879]. agreement [2601]. agreement-based [2601]. aid [2818]. aided [486]. AIDS [743, 777, 894, 751, 843, 1652]. Aims [1219]. Akt [2874]. al [735]. Alan [2992]. alanine [2779]. alba [3]. Albert [3097]. albopictus [2067]. aléatoire [2556, 2656]. algae [2136, 2380, 167, 122, 2746]. Algal [2522, 499, 1096, 3528, 314, 3210]. algebra [3033, 672, 154, 1476, 2883, 1536, 627, 5, 49]. Algebraic [2437, 1355, 2974, 1709, 548, 1402, 3517, 2332, 3110, 3461, 1536, 3314, 3427]. algebraic-delay [3427]. algebras [232, 896, 959, 107, 69, 144]. algorithm [1560, 3576, 1498, 1635]. Algorithmic [2609, 3510]. Algorithms [17, 2747, 739, 884]. alignment [1702, 1458, 2859, 1358]. alignments [3419, 3615]. all-atom [1708]. all-leader [3226]. Allee [2986, 2643, 2154, 3565, 3504, 3530, 1945, 2325, 2996, 3375, 3606, 3272, 1578, 2444, 1326, 1922, 2135, 2910]. Allee-type [2325]. Allele [415, 384, 1345, 886, 2804, 1715, 842, 24, 117, 252, 1377, 1835, 704, 2659, 166, 2718, 1992, 2173, 2258]. alleles [900, 604, 1018, 332, 390, 657, 36, 651, 250, 1962, 637, 1021, 251, 375, 605, 296, 326, 1854, 1385, 1207, 934, 22]. allelic [857, 284]. allelopathy [1197, 3490]. Allen [2727]. allocation [1571, 3161, 3384, 646, 2612, 2608, 3460, 3071, 3124, 2698, 3481, 1152, 1332]. Allometric [1300, 456, 1416]. allospecies [139]. allow [980, 3248]. allowing [647]. allows [2857]. almost [3400, 515, 1155, 3022, 3555]. along [2255, 2346, 3531, 1910, 2599, 2532]. alterations [3402, 2275]. altered [3552, 3470]. Alternative [2949, 2628, 1294, 452, 2247, 3219]. alternatives [3090]. altruism [2259]. always [2204]. Alzheimer [2197, 3158, 3532, 2317, 2844]. Amari [2958]. ambient [2295]. Amendment [3193]. Amoeboid [100, 1657, 2523]. among [1253, 2457, 1895, 2770, 1172, 439, 1212, 1043, 744, 847, 1430]. amongst [2122]. amount [154]. amphibian [1685]. amplification [2328, 2348, 1459, 3542, 2868]. amplification-driven [3542]. amplifiers [3186, 3486]. amplitude [344, 870, 367]. amyloid [2197]. anaerobic [2878]. analyse [1652]. analyses [2380, 1028]. Analysing [1336, 1561, 1453]. Analysis [3552, 1641, 238, 1008, 1517, 2323, 434, 776, 3239, 990, 661, 1085, 1339, 352, 2588, 2797, 1184, 1421, 3056, 3191, 334, 1002, 3164, 2000, 397, 39, 3079, 3141, 229, 311, 2082, 893, 1109, 2546, 1592, 228, 411, 2994, 1141, 1558, 65, 1521, 1267, 2481, 3354, 742, 2107, 1523, 3143, 314, 3333, 321, 2488, 3507, 2274, 2765, 1356, 1149, 1096, 1742, 3211, 1913, 3612, 1953, 3581, 1880, 3468, 2264, 2655, 2683, 2387, 772, 567, 248, 1167, 1537, 3425, 1739, 1415, 465, 1303, 2515, 3351, 886, 900, 3032, 1653, 412, 604, 1499, 1198, 2591, 151, 2701, 802, 1700, 1027, 1731]. analysis [690, 1343, 3562, 2779, 3077, 2760, 1976, 2695, 1138, 1301, 2843, 486, 2194, 2529, 2914, 1857, 1810, 3544, 1284, 1690, 1824, 2926, 2134, 1067, 771, 1741, 780, 1399, 2928, 1090, 3165, 79, 849, 1903, 3453, 55, 1189, 781, 3615, 2317, 2844, 3075, 331, 644, 1494, 983, 1289, 1984, 1952, 3296, 840, 1474, 1210, 1254, 3228, 2400, 1832, 3617, 2162, 2566, 3194, 178, 487, 2929, 1723, 1983, 2506, 2925, 1139, 949, 1950, 2983, 1859, 2690, 87, 1262, 1981, 2029, 2092, 921, 1573, 287, 3404, 1033, 1629, 1882, 2362, 2174, 1476, 3140, 2800, 1665, 1666, 1849, 1387, 1526, 866]. analysis [2386, 3117, 1819, 31, 1850, 3316, 1137, 2462, 2098, 2326, 1259, 244, 3556, 1233, 2196, 3270, 3186, 1243, 1668, 1388, 2424, 3173, 1566, 2195, 3272, 3579, 822, 51, 2257, 2901, 2886, 1035, 2428, 1635, 2499, 2892, 2048, 296, 3087, 795, 326, 1188, 2922, 1506, 3535, 63, 2523, 981, 3451, 2037, 294, 1314, 3198, 3150, 2774, 3221, 3377, 2702, 2216, 2824, 1823, 864, 1246, 1247]. Analytic [1739, 3128, 400]. Analytical [3494, 1135, 3504, 1231, 152, 1423, 2272, 2928, 2276, 3608, 179, 2369, 1894, 1890, 2782]. analytically [1626]. analyzer [388]. Analyzing [1788, 1905, 1674, 2369, 1970, 2555]. ancestor [1331]. Ancestral [3455, 3284, 1595, 706, 2261, 2906, 1910, 1442, 1340, 3473]. Ancestries [2472]. ancestry [1637, 2799]. and/or [3268]. androdioecy [1410]. anemia [328]. aneurysm [1367]. aneurysms [2213]. angioedema [2476]. Angiogenesis [1799, 1772, 1027, 1363, 1382, 1278, 1466, 3140, 2829, 1754, 3288, 2355, 2072, 2824]. angiotensin [3443]. Anguilliform [947, 1620]. animal [890, 1933, 2529, 1857, 1810, 2014, 999, 261, 811, 2532, 1481, 3456]. animals [1300, 1358, 217, 2532]. Anisotropic [1630, 1628, 3155, 1933, 1223, 2912, 771, 677]. anisotropy [2520, 878]. Announcement [77, 219, 291, 353, 363, 438, 454, 521, 544, 584, 635, 674, 689, 695, 710, 726, 791, 844]. Announcements [378, 484, 577, 762]. annual [1398, 1571]. Anomalous [3600, 1669, 1842]. answer [2763]. ant [2875, 2514, 1475]. antagonism [1579]. antennal [479]. antennules [1319]. Anti [1005, 1541, 3116, 2642, 2560, 3057, 2824]. anti-angiogenesis [2824]. Anti-phase [1005, 3057]. anti-predator [2642]. anti-quorum [1541]. anti-retroviral [2560]. anti-vascular [3116]. antibiotic [1541, 2701, 1830, 1896, 2551, 2259]. antibiotic- [1541]. antibiotic-resistant [2259]. antibiotics [2575]. antibody [2412, 263, 2335, 3054]. anticancer [1473, 2856]. Antigen [3229, 322]. antigenic [1806, 2091]. antigens [1768]. antimalarial [1941]. antimalarial-drug [1941]. Antimicrobial [1554, 3561]. Antiparallel [2657]. antiretroviral [3605, 3422, 3084, 2061, 1836, 2179]. antitoxin [3180]. Antivin [2198]. Antivin/ [2198]. antiviral [2229]. ants [1014, 528]. any [1064]. aortic [2570]. APC [2692, 2767]. apoptosis [3221]. Aposematic [3329]. app [3243]. apparent [1148]. appearance [440, 751, 124]. appears [1226]. apple [149]. applicability [2715]. Application [3319, 55, 3167, 222, 431, 853, 1497, 960, 1164, 3032, 2687, 1285, 75, 40, 3201, 2828, 1636, 3432, 2557, 1447, 2766, 30, 2476, 2394, 134, 1632, 3617, 2605, 3142, 3528, 1579, 2213, 1652, 2758, 1904, 3279, 135, 3388, 2982, 3174, 1894, 1581, 2824, 1495, 1334, 2344, 305]. Applications [1341, 1892, 2043, 2058, 3294, 1683, 2303, 3392, 1763, 451, 2739, 1821, 1377, 2082, 2917, 2570, 2081, 1257, 1801, 2379, 1760, 1889, 1866, 1635, 3187, 2765, 3577, 2565, 3421, 2413]. applied [420, 2033, 2953, 2148, 1264, 2358]. applying [2515]. Approach [343, 1554, 1593, 2987, 3212, 1637, 1917, 3083, 1674, 2105, 2314, 3223, 2770, 3239, 1505, 2433, 409, 3544, 3076, 3464, 3603, 1777, 2352, 3593, 2736, 2806, 479, 1327, 308, 441, 3011, 3439, 452, 181, 1171, 905, 2019, 2475, 3217, 2903, 1831, 2678, 2721, 1835, 1699, 765, 3243, 2282, 2383, 2740, 1675, 3511, 2160, 1102, 3312, 2330, 2148, 3263, 3341, 135, 2994, 2990, 3546, 27, 3092, 2769, 1435, 3314, 1894, 1890, 3364, 3620, 289]. Approaches [978, 2590, 3497, 1712]. Approximate [320, 490, 2022, 1870, 2285, 3599]. Approximating [1292, 2184, 1593, 1680]. Approximation [2359, 2284, 579, 419, 2798, 1364, 3083, 1445, 2843, 3319, 1143, 1691, 3091, 30, 3407, 3075, 1238, 2982, 1150]. Approximations [1941, 223, 1417, 1135, 2252, 2681, 2574, 719, 2649, 790, 865, 3065, 1835, 3413, 1583, 2030, 588, 764]. aquaporin [1771]. aquatic [3597, 2031, 2867, 2517, 2746]. Arabidopsis [3247]. arbitrary [2987, 2881, 1991, 913, 154, 2375, 2468, 155]. arboviral [2723]. architecture [2387, 2520, 1777, 2605, 1280]. area [3591]. areas [2016]. arguments [1058]. Arid [1523, 2797, 3101, 2531, 3366, 3505]. arises [40]. arising [139, 2409, 698, 908, 3569, 1094, 1284, 2962, 2622, 2952, 1691, 39, 2238, 814, 2763]. arithmetic [3419]. arm [1486, 3032]. army [1014]. Arnol’d [1459]. array [1388]. arrays [803]. arrival [3464]. arrythmias [307]. arterial [2885]. arteries [1195, 3440, 443, 2451]. arteriosus [2201]. arthritis [3056]. artificial [82, 1681]. ary [3361]. asexual [1024, 3571]. aspect [364]. aspects [903, 2190, 1447, 2663, 1832, 2272]. aspergillosis [3489]. assay [3094, 1629, 1869, 135, 1771]. assembly [3370, 2394, 1756, 981]. assess [3287]. assessed [2950]. Assessing [2794, 3488, 2998, 3381, 3100]. Assessment [3224, 2637]. Assigning [2331]. Assignments [3139]. associated [352, 1142, 660, 3489]. Association [1558, 968]. associations [578, 606]. associative [3434]. Assortative [399, 82, 334, 333, 497]. assuming [62]. assumptions [1434, 1490]. astrocyte [2591]. asymmetric [3069, 3251, 3616, 3185, 283, 1583, 3082]. asymmetries [479, 3531]. asymmetry [2120, 2327]. asymptomatic [2784, 3467]. asymptomatics [3242]. Asymptotic [740, 817, 350, 1895, 3523, 1119, 1646, 967, 3069, 1728, 473, 2557, 511, 696, 2852, 873, 3529, 2800, 259, 3583, 2334, 1558, 703, 1591, 207, 1489, 3265, 3338, 3225, 2278, 3267, 2934, 771, 126, 472, 3543, 1587, 452, 2260, 748, 464, 2960, 945, 224, 43, 581, 2040, 583, 597, 194, 539, 392, 993, 3150]. asymptotical [1647]. asymptotically [447, 970, 885, 3019]. Asymptotics [1265, 2877]. atherosclerosis [2448, 2051, 2276]. atherosclerotic [3586, 3096]. atmosphere [2522]. atom [1708]. atria [2454]. attachment [2240]. attainability [829]. attempt [2494]. attracting [1596]. attraction [256, 1644, 3246]. attractive [448, 2460]. Attractivity [1931, 1206, 3315]. attractor [1346, 1412, 1411]. attractors [356, 3361, 1186]. attribute [1864]. attrition [1622, 1043]. auditory [770]. Augmentative [2643]. aureus [1879, 1501, 2913]. autocatalysis [3342]. Autocatalytic [1190, 1855, 973, 877]. autoimmune [2980]. autologous [1406]. Automata [1025]. automaton [2306, 1954, 2689]. automorphism [1930]. autonomous [1008, 3527, 1122, 1882, 2576, 885, 16]. autophagy [3266]. autoreactivity [2980]. autoregressive [479]. autoregulated [2810]. autoregulation [2695, 3472]. autosomal [835, 838, 845, 910, 923]. auxin [3357, 2065, 2194, 2243, 773]. AV [307]. availability [2131]. avascular [1354, 1198, 1584, 1240, 1306]. Average [3343, 479, 2940, 3480, 2543, 3397]. averaged [3393]. Averaging [245, 812]. avian [2552, 2553, 2158]. avian-only [2552, 2553]. Avner [3418]. avoid [462, 560]. awareness [2742, 3497]. axial [1458, 754, 1333]. Axin [2692]. axis [1972]. axisymmetric [2062]. axon [442, 1721, 1525, 2346, 12, 566, 115, 47, 792]. axonal [3572]. Ayala [96].

B [2536, 1865, 322, 403, 3420, 2358]. B-cell [3420]. Bacaër [1834]. Back [3619]. backbone [1708]. Backward [1815, 2561, 3462, 2187, 1392, 2390, 284, 1613, 2582, 1242]. Backwards [1133]. bacteria [1427, 2371, 950, 1656, 469, 2027, 1705, 2259, 3150]. bacteria-plasmid [950]. Bacterial [2384, 1221, 330, 990, 1275, 3517, 3135, 2287, 192, 2332, 1586, 237, 2645, 1529, 404, 2568, 2899, 1188, 602, 3524, 1403, 2338, 2562, 2876]. bacteriocin [1656]. bacterium [1501]. balance [319, 3236, 428, 955, 2591, 2212, 1427, 2954, 781, 21, 1518, 2024, 2069, 305, 2702]. balanced [3037, 2541]. balancing [280, 65]. baleen [75]. band [1146, 872, 469]. band-type [872]. banded [2797]. bank [2930, 3172, 2278, 2068]. banks [3200]. Barabási [3097]. bargaining [2806]. barley [1495]. baroreflex [1738, 1123]. baroreflex-feedback [1123]. barrier [1006, 3111, 68]. Barton [2851]. based [3300, 2875, 3213, 1013, 1917, 1397, 2133, 2601, 2128, 1875, 1483, 1795, 3224, 1673, 3338, 3005, 1988, 3562, 2954, 3226, 2716, 2025, 2194, 1979, 2414, 3359, 1327, 2645, 3512, 3588, 295, 3079, 487, 2855, 2981, 1831, 2351, 2764, 3498, 1261, 3437, 1588, 2947, 1954, 2121, 2178, 3471, 3312, 3341, 1309, 2873, 1692, 2786, 1916, 2186, 1079, 2502, 2816, 2054, 2901, 648, 2545, 2077, 3071, 2773, 3456, 3599, 2188, 1373, 3220, 3434, 3290, 1792]. baseline [1767]. Basener [3619]. Basic [3030, 3417, 828, 1939, 2208, 1167, 2203, 784, 2042, 1543, 2050, 2942, 3529, 2155, 1291, 1693, 2125, 822, 1828, 1430]. basin [1644]. basins [1471, 3246]. basis [3215, 568, 1787, 360, 1137, 27, 3197]. bass [853]. batch [662]. Batesian [2929]. Bayesian [1670, 3201, 1072, 2200, 2764, 2678, 3599]. be [1545, 3551, 2731, 2516, 2364, 2477, 2521, 2127, 1836, 2382, 1054]. beam [1798]. bearing [1273, 983, 1041, 1033]. beating [859]. Becker [3587]. become [3003]. becoming [1604]. bed [2773]. Beddington [2661, 3621]. beds [2870]. bee [3238, 3502, 861, 3070]. bees [3070]. behavior [626, 740, 817, 470, 1599, 1164, 1895, 1215, 3523, 3251, 590, 1129, 2433, 3328, 3269, 473, 563, 2934, 3180, 3368, 2604, 959, 999, 1587, 1297, 3529, 3573, 693, 1365, 569, 650, 945, 2839, 2995, 775, 1487, 3185, 2309, 2642, 3583, 889, 2107, 1987, 1927, 3299, 207, 400]. behavioral [3204, 617, 385, 616, 3173, 957]. behaviors [1180, 3021]. behaviour [2869, 3300, 350, 2420, 2430, 2725, 1119, 967, 774, 3338, 1898, 3332, 517, 2579, 962, 1569, 1734, 1535, 126, 566, 2757, 3217, 1122, 587, 1681, 136, 1292, 2020, 2171, 259, 1058, 1655]. behavioural [2587, 3427]. Bellman [2940]. Belousov [410, 124]. Benasque [1702]. bending [3320]. Bendixson [1816, 885]. beneficial [2806]. benefits [1642]. bent [1600]. benthic [2746]. berlese [55]. Bernoulli [160]. Bernstein [896]. Best [2908, 3045, 2754, 3166, 3139, 3023, 1642, 2379]. bet [3162]. bet-hedging [3162]. beta [2779, 3015]. beta-alanine [2779]. beta-splitting [3015]. better [2385, 3508]. between [2324, 1455, 1273, 2805, 3380, 2959, 2440, 3072, 3611, 947, 1995, 428, 2590, 3005, 36, 2059, 342, 3127, 2621, 2170, 1817, 3015, 237, 2806, 2835, 333, 2933, 253, 310, 2371, 2018, 2463, 21, 669, 1041, 3192, 459, 659, 787, 3561, 2455, 277, 1556, 1656, 3614, 963, 2191, 159, 3341, 1992, 1825, 3271, 66, 2587, 658, 989, 2185, 2403, 1277, 1994, 987, 1516, 942, 3398, 2704, 3335, 2167, 1152, 1332]. between-host [2805, 3561]. Beyond [2468, 3245, 2821, 3473]. bi [3413]. bi-molecular [3413]. BIAcore [1470, 1683]. biallelic [439, 497, 781]. bias [3455, 2016, 3143, 2774]. Biased [223, 522, 1223, 3603, 1901, 2167]. bidegree [2872]. Bidirectional [2937]. bidirectionally [3554]. bidomain [1716]. biennial [687, 737]. biennials [1378]. Bifunctional [3607]. bifurcating [3061]. Bifurcation [2723, 2998, 313, 2701, 2448, 2202, 140, 1399, 949, 2055, 3404, 1387, 269, 3272, 3451, 3377, 864, 274, 2683, 1169, 2497, 1349, 2226, 2423, 1129, 890, 2670, 352, 1888, 2662, 486, 2194, 1284, 681, 734, 79, 1815, 781, 911, 840, 1299, 2929, 3286, 2029, 3528, 1392, 2390, 2669, 292, 125, 1743, 1613, 460, 2769, 3579, 1897, 425, 2749, 1512, 1413, 2132, 3621, 1557, 3221, 2582, 1242, 287]. Bifurcations [1088, 429, 477, 636, 3035, 3479, 1806, 2154, 3358, 1094, 1133, 345, 2694, 1401, 3374, 3462, 518, 1074, 3613, 3500, 1007, 1127, 2561, 2119]. bilayers [2669]. bile [908]. bilipid [3501]. bimatrix [1059, 668]. binary [2033, 2660, 2713, 1948, 3104, 3313, 3231]. binding [1044, 2703, 263, 2616, 2519, 1805, 1229, 2872, 3198]. binocular [212]. binomial [587, 3242]. bio [2236]. bio-polymerization [2236]. biochemical [274, 123, 3351, 2475, 2322, 2275, 3463, 112, 1665, 1666, 56, 2568, 65, 1958, 824, 825, 1355, 951, 19, 2358]. biochemistry [2338]. biocontrol [2643]. bioconvection [1179]. Biodiversity [3018, 1763, 2189, 2816]. Bioeconomic [2424]. bioenergetics [611]. biofilm [2508, 1407, 2740, 2417]. biofilms [2138, 1403]. biogas [2878]. biological [2324, 2337, 1913, 1953, 500, 2060, 1986, 2147, 1700, 933, 916, 1124, 2504, 3527, 1217, 2041, 1562, 1899, 2, 320, 3548, 2524, 145, 177, 308, 345, 3255, 1856, 3002, 2739, 393, 368, 2400, 89, 459, 230, 3307, 887, 2240, 2597, 673, 1954, 2344, 691, 3493, 1692, 3500, 2210, 3174, 1889, 1866, 17, 747, 9]. biologically [1373]. biologically-based [1373]. Biology [1774, 485, 1696, 1621, 2150, 2114, 3201, 2779, 2825, 3569, 3033, 168, 1658, 2372, 3520, 2145, 1760, 1631, 1784, 76, 1594, 2992]. biomarkers [2818]. biomass [3597]. Biomathematics [485]. biomechanical [3154]. biomedicine [2152]. Biomembranes [1534]. biomolecular [1818, 1252]. biomorphoelastic [3444]. bioproduction [3428]. bioreactor [2508]. biosensor [2413]. biosensors [2735]. biparental [3301]. bipedal [1601]. biphasic [1427, 244, 3599]. bird [2697, 2230, 1600]. birds [2552, 2553, 3226, 2699, 2265]. Birth [2739, 3486, 490, 2281, 2656, 254, 1209, 1369, 2058, 3435, 731, 303, 2947, 647, 699, 738, 3186, 1328, 638, 3598]. birth-and-death [2281]. birth-death [2058, 2739, 3435, 731, 2947]. Birth/birth [2739]. Birth/birth-death [2739]. birthday [2485, 2992, 2109]. bisexual [713]. Bistability [1827, 2672, 1940, 1989, 3255, 2592, 2966, 2349, 2442]. bistable [2075, 1587, 2538, 2758, 3271, 1647, 1695, 3118, 1655]. bivalent [263]. bivalents [2815]. black [3498]. black- [3498]. blastula [636, 3010]. blebbing [2765]. blobs [3395]. Block [3540, 307, 2944, 180]. Block-pulse [3540]. blocking [2004]. blocks [1036, 1332]. blood [1160, 1073, 1724, 1575, 1605, 1006, 86, 3013, 1839, 2213, 717, 941, 1846, 233, 1967, 2492, 2216]. blood-retina [1006]. blood-stage [1967]. blooms [991, 1390, 3528]. blow [2962, 1342, 1011, 3098]. blow-up [2962, 1342, 1011, 3098]. blue [122, 1212]. blue-green [122]. BMP [3010]. BMP-mediated [3010]. boat [712]. body [182, 947, 760, 1391, 1082, 391]. Bogdanov [3500]. Boltzmann [1700, 1585, 2939, 1529, 1701, 1112]. Bombay [2003]. bond [3365]. Bone [2858, 3229, 2322, 1148, 827, 682, 3230, 985]. Boolean [97, 3087]. booster [1559, 2985]. boosting [2458, 2838, 3452]. boring [1647]. borne [1598, 3302, 1543, 3593, 1639, 2953, 1436, 2257, 2283, 3146, 3555, 3417]. botflies [1274]. both [3467, 2552, 2553, 3135, 67, 2939, 2559]. bottlenecks [1509]. bottom [1777]. bottom-up [1777]. bound [2407, 3515, 1762, 892]. boundaries [287, 2034, 3410]. boundary [2959, 592, 2151, 2737, 2542, 40, 2641, 3362, 2663, 3374, 2634, 3217, 956, 2062, 2699, 3404, 3430, 2213, 941, 582, 3514, 376, 2392, 2093, 2932, 1909, 2624]. boundary-layer [2392]. boundary-value [2093]. bounded [1231, 2044]. Boundedness [2952, 165, 727]. Bounding [2796, 1522]. Bounds [2748, 2714, 3252, 3407, 3559]. bovine [1002, 2326]. box [3498, 1456]. brain [2542, 1839, 1802, 1153]. brainworms [1274]. branched [1974]. Branching [468, 2721, 366, 1637, 841, 3200, 2128, 2978, 3293, 3455, 3251, 1797, 2547, 1429, 317, 474, 2238, 2388, 1948, 1961, 1448, 2764, 1034, 2868, 1770, 2344, 694, 3409, 641, 3518, 1651, 163, 1353, 876, 2626, 68, 1515, 1602, 935, 1481, 612]. Branching-diffusion [468]. branching-extinction [1429]. branching-process [841, 3409, 641]. Brauer [3520]. breakage [2461, 1309]. breaking [890, 3374]. breaks [2172]. breast [1422, 1893, 3381]. breath [1983, 360]. breathing [1268, 3233]. breeders [3227]. breeding [1528, 2112]. bridge [504, 2461, 1213, 2587]. Bridging [3005]. brief [2041]. broadened [2715]. brood [1453]. Brownian [2104, 3458, 2446, 1259, 2272, 3473]. BRSV [1815]. Brucellosis [1274, 1002]. bruit [1367]. BSE [1430]. bubbles [2073]. buckle [2604]. buckling [2869]. budding [935]. Budget [1261, 1457, 2891]. Budgeted [1991]. budgets [568, 1787]. budworm [1189, 198]. buffered [1587, 2397, 1647, 1695]. buffers [1914, 1647, 2204]. building [3399, 1036, 1332]. building-blocks [1036]. bulk [431, 1556, 3271]. bumble [3238]. bundles [2271]. burden [3326, 3476]. burn [3214, 3444]. burns [2654]. burst [2640, 2627]. bursting [823, 996, 1008, 911, 269, 671, 1080, 1844]. bursty [3025]. Butterflie [1505]. Butterfly [2081]. by-products [1608]. Bye [3526]. bystander [2666].

C [2122, 3337, 1064]. Ca [1413]. cable [2151, 1842]. cadherins [2647]. Cahn [2727, 2669]. calcic [3089]. Calcium [1914, 3552, 796, 146, 2909, 2146, 115, 1876, 1716, 2204]. calcium-induced [146]. calcium-regulated [796]. calculate [2691]. Calculating [1540, 741, 2418, 1430]. Calculation [678, 409, 2939, 1355, 2734]. Calculations [1948]. calculus [1547]. Calibration [2891, 3599, 3405, 2200, 2195]. Call [354]. Calvin [164]. CaMKII [2142]. cAMP [1573]. Can [2144, 3111, 3511, 2964, 1836, 2471, 1545, 2136, 3590, 1725, 2731, 3591, 3003, 3254, 2516, 1552, 401, 3513, 568, 1787, 2672, 2521, 2095, 2587, 3473, 1054, 2965, 3490]. Canadian [1663]. canalicular [908]. canard [2994, 3063]. Canard-induced [3063]. Canards [1553, 2182]. cancellous [1148]. cancer [3095, 1953, 2108, 2097, 530, 2144, 2139, 2507, 2557, 3362, 3453, 2516, 1210, 3491, 2254, 3367, 3568, 3381, 3288, 3585, 3514, 3174, 1804, 3012, 1792]. cancer-induced [3288]. cancerous [723, 1799]. Canham [2235]. canine [239]. Cannibalism [712, 1546]. Cannibalistic [3541, 623]. Cannings [3249, 3046]. Canonical [1809, 2498, 1015, 260]. canopy [1667]. cap [3096]. capacity [649, 3085, 3099, 3193, 2316, 3573]. capillaries [1388, 655, 233]. capillary [1278, 2216]. capsid [2394]. capsule [2177, 1761]. capture [2845, 1319]. Capturing [2459, 3454]. carbon [3241, 281, 2712, 297, 2418, 3210]. carcinogenesis [366, 1477]. carcinogenic [294]. carcinoma [930, 1422]. Cardiac [3279, 1679, 1745, 2454, 771, 345, 307, 2894, 2981, 659, 2160, 2967, 878, 1409, 498, 1077]. cardiorespiratory [1738]. Cardiovascular [2950, 1497, 1755, 1539, 924]. cardiovascular-respiratory [1497]. Carey [2415]. cargo [2629, 3572]. caricature [698, 779]. Carlo [3223, 2747, 3498]. carnivore [432]. carriage [2784]. Carrier [524, 181, 3452]. carrier-facilitated [181]. carrier-state [3452]. carrying [649, 2674, 3085, 3099, 3193, 2491, 2634, 2316]. Cartesian [3431]. cartilage [2686]. cascade [1989, 2998, 1176]. cascades [2880, 2286, 1908, 2064, 1075, 1782]. case [3095, 3274, 1880, 943, 1397, 3249, 1103, 2622, 2496, 1412, 2732, 1411, 1910, 1065, 2941, 2642, 1307, 2586, 1881, 1685]. cases [751, 2826]. catalysed [1434, 1490]. catalysis [305]. catalysts [2624]. catalytic [279]. catalyzed [265, 2293]. catastrophe [1133, 60, 699, 738]. catastrophes [437, 1401, 647]. Catastrophic [1557, 2497]. category [3604]. catenin [2692, 2767, 1849]. Caterpillars [3104]. catheterised [1840]. cations [46]. Cattaneo [1380, 1395, 1396]. cattle [2295, 1430]. Cauchy [3312]. causal [2150, 1852]. cause [552, 3591, 3527, 3216]. caused [3477, 1941]. cavities [2939]. Cayley [2081]. CD4 [3192, 3079, 1836]. CD8 [3192, 2766, 2048]. celebration [3520]. Cell [1774, 2377, 1427, 1877, 2222, 1183, 1612, 478, 533, 572, 238, 3380, 1953, 1791, 2409, 740, 817, 2617, 2387, 1167, 2888, 1044, 1415, 1473, 2136, 1221, 1960, 1674, 1294, 1354, 1737, 1166, 975, 3265, 2724, 2978, 1795, 3005, 1472, 3131, 1605, 2433, 2621, 2408, 1393, 3567, 705, 2401, 922, 1223, 476, 2489, 2703, 2194, 2005, 2520, 3113, 1690, 2971, 3067, 2663, 2456, 2106, 2928, 3009, 2766, 263, 1901, 2613, 1975, 2018, 2463, 2002, 3295, 554, 2038, 3192, 2286, 1431, 2676, 86, 2074, 3491, 2342, 3013, 2506, 1780, 3286, 464, 2240, 1843, 1978, 2207, 1630, 893, 2062, 3260, 3006, 1878]. cell [1838, 1480, 370, 1181, 750, 2916, 2647, 2282, 2341, 1781, 2024, 2069, 3185, 1280, 2118, 2689, 2011, 2863, 1406, 116, 1800, 44, 43, 1312, 2980, 1233, 2196, 2768, 3298, 3381, 3271, 2223, 3514, 3537, 1079, 985, 3057, 2564, 2054, 600, 3386, 2426, 892, 2048, 703, 539, 3246, 1783, 2248, 3153, 1258, 2368, 3420, 575, 2338, 1153, 3387, 3620, 631]. cell-based [2054]. cell-bound [892]. cell-bulk [3271]. Cell-cell [1612, 1354, 2724, 3067, 3537, 3620]. cell-haptotaxis [750]. cell-induced [3298]. cell-to-cell [1791, 1393, 2768, 3420]. cells [723, 1322, 1946, 570, 1775, 1073, 81, 3358, 1745, 1176, 701, 1817, 803, 3105, 3176, 1657, 2340, 3464, 2205, 3603, 1421, 3548, 1777, 3181, 3170, 871, 852, 625, 2909, 1723, 1893, 3399, 659, 1842, 1681, 3147, 1358, 773, 3367, 318, 424, 505, 872, 2184, 641, 3148, 436, 269, 909, 2560, 1836, 3174, 1876, 1844, 1516, 795, 2765, 1784, 2792, 1768]. cellular [2306, 2139, 1852, 2419, 1777, 2463, 70, 1183, 997, 1954, 1102, 2689, 2612, 2608, 116, 1257, 3514, 1782, 20]. cementation [827]. center [448, 2625, 2646]. Central [133, 1601, 930, 1281, 1282, 3238, 1741, 1564, 1552, 1579, 1904]. central-place [3238]. centrally [2619, 159]. centre [2440]. centrosymmetric [3454]. cerebral [1351, 1538, 2213]. cerebrospinal [1839]. cerevisiae [589]. certain [1641, 626, 240, 7]. cervical [3568]. Cevolutionary [449]. CFSE [1629, 1838, 2327]. Chain [2975, 1196, 29, 1168, 2303, 953, 992, 545, 1831, 587, 3242, 2546, 1699, 880, 2671, 394, 3012, 2214]. chain-binomial [587]. chains [3025, 365, 3392, 1561, 2778, 69, 966, 1122, 1365, 2329, 1229, 2646]. challenge [2651]. challenges [1733]. chamber [1275]. Change [1377, 3107, 3250, 2801, 1297, 3416, 3513, 324, 2827, 1451, 3173, 2816, 109, 957, 3299]. changes [1044, 2318, 3601, 3046, 2849, 107, 1372, 1521, 3354, 792]. changing [316, 1052, 3339, 3412]. channel [114, 147, 1789, 536, 2417, 1495]. channel-clogging [2417]. channels [2902, 1861]. Chaos [2914, 966, 2158, 1279, 419, 1320, 1087, 734, 2856, 1090, 345, 787, 1365, 2073, 3122, 1663, 1201, 2290, 574, 3579, 2229, 2442, 124, 1100, 1314]. Chaotic [2707, 891]. chaperone [2513]. chaperone-mediated [2513]. Chapman [46]. Characean [872]. character [399, 512, 764]. characterisations [2873]. characteristic [352, 2489, 32]. characteristics [3514, 2816, 989]. Characterization [3574, 3139, 1336, 2752, 487, 2395, 3461]. characterizations [1900, 2949]. characterize [3047]. characterized [3544, 601]. Characterizing [1503, 2125, 3480]. characters [2789, 90, 745, 597]. charged [46]. charges [235]. chase [3007]. chase-and-run [3007]. checkerboard [3537]. checkpoint [2836]. Chemical [2301, 1900, 960, 1827, 306, 1626, 2759, 2566, 33, 88, 645, 1859, 3461, 11, 1665, 1666, 2785, 557, 2054, 3174, 1150, 3342, 3245, 528]. chemosensitive [1380, 1395, 1396, 1493, 2002]. chemosensory [643]. Chemostat [968, 1273, 1362, 769, 313, 423, 593, 2226, 2681, 1376, 1563, 1816, 919, 1162, 1608, 983, 1041, 2729, 1033, 3132, 2031, 3360, 2397, 1062, 1099, 2027, 892, 1095, 1185, 1502, 839, 3490]. chemostat-like [2226]. chemostat-type [1162]. Chemotactic [1083, 3212, 1427, 1723, 3498, 856, 469, 1257, 1188]. Chemotaxis [1676, 1118, 223, 626, 2628, 1892, 1947, 2724, 1275, 3567, 1542, 3603, 755, 1778, 1529, 2690, 1664, 3142, 3404, 1906, 1011, 2964, 2269, 1264, 1258, 2132, 2338, 2562, 2384]. Chemotaxis-induced [1676]. chemotherapeutic [3514]. chemotherapy [723, 1447, 1210, 1110, 2010, 1406, 2503, 1500, 985]. cherries [3225]. Cheyne [1755]. chiasma [506, 1142]. chickens [403]. Chikungunya [2067]. child [2273]. childhood [2879, 1829]. China [3617]. choice [787, 1987, 1864]. cholera [3167, 2441, 2435, 3232, 3299, 3554]. chondrocyte [1975]. choosing [2074]. chosen [310]. ChR2 [2728]. ChR2-controlled [2728]. chromaffin [3148]. chromatic [386]. chromatid [155]. chromatin [1309]. chromium [2938]. chromosomal [209, 1064]. chromosome [1926, 2000, 169, 522, 155, 2274]. chromosome- [155]. chromosomes [2219, 2645]. chronic [1946, 2677, 2412, 2291, 2822, 1527, 2143]. cicadas [1236]. Ciliary [859]. circadian [1709, 32, 3216, 1409, 648, 3247]. Circle [860]. Circle-map [860]. circuits [2854, 1852, 56, 507, 1969, 2370]. Circular [2391, 2783]. circulating [677, 985]. circulation [1538, 3524]. circulatory [492]. circumferential [748]. cirratum [131]. cities [2440]. Clades [2398]. clamp [1617, 1456]. clans [2398]. clashes [1708]. class [1641, 2218, 330, 2580, 1378, 2509, 1562, 95, 2896, 2487, 1686, 2781, 1921, 2234, 2900, 453, 56, 2618, 2397, 2982, 275, 3118]. class-structured [2781]. Classes [3307, 2952, 1036, 963]. Classical [1917, 2222, 308, 225, 620]. Classification [3285, 260, 2193, 2308, 778, 915, 1713, 2634, 3207]. classify [55]. classifying [2645]. cleaning [2295]. clear [2842]. cleavage [636]. climate [2737, 3107, 3250, 2801, 2827, 3312, 2773]. climate-based [3312, 2773]. climax [986]. clines [3274, 139, 768, 283]. clinical [751]. Clock [1565, 2717, 1885]. clocks [1709, 393]. clogging [2417]. clonal [2563, 3265, 2968]. clonotypes [2617]. Closed [2446, 1534, 666, 2241, 1389, 3090]. Clostridium [2364, 2709]. closure [1832, 1935, 2030, 2072, 3451]. closures [2945, 3519, 3498]. clot [2480]. clumped [1141]. Cluster [2946, 2512, 2202, 3038, 2214]. cluster-similarity [3038]. clustered [2945, 1267]. clustering [2123, 3325, 2377, 2529, 2270, 2919, 714, 2468]. clusters [2651, 2783]. clutch [628]. Co [2955, 1351, 1767, 2857, 3524, 3146]. co-circulation [3524]. Co-divergence [2955]. co-evolution [1767]. co-feeding [3146]. Co-incidence [1351]. co-stability [2857]. coagulation [1151]. Coalescence [1311, 2849, 1904]. Coalescent [1418, 1942, 3395, 3600, 3293, 3240, 741, 3294, 2848, 2609, 937, 804, 1904, 2379, 2369]. coalescents [1753, 3046]. coalgebras [2251, 2598]. Coarse [2387, 3008, 2683, 2077]. Coarse-grained [2387, 2683]. Coarse-graining [3008]. cochlear [1233]. cocirculating [1113]. cockroaches [3344]. code [3325, 1536]. codes [2391]. codifying [2494]. Coding [2339]. codon [2645, 3090]. coefficient [678, 3059, 1007, 2358, 137]. coefficients [2577, 1169, 790, 865, 2466, 3078, 764, 289]. Coevolution [465, 1056, 1692, 540]. Coevolutionary [1277, 3331]. Coexistence [801, 299, 423, 2226, 3582, 1375, 1608, 1544, 663, 1428, 880, 602, 1408, 980, 1725, 3565, 1097, 1731, 2668, 1378, 855, 3227, 1488, 2371, 1491, 2277, 3439, 3175, 3370, 526, 545, 2857, 3506, 813, 3050, 2034, 925, 3490]. coexisting [106]. coffee [3601]. cognitive [3602, 3456]. cographs [2100, 2660]. Coherent [2621, 2683, 1931]. cohesive [1146]. cohort [853, 1116, 3489]. coin [2716]. coinfection [3531]. cointegrated [2682]. Coleochaete [2368]. coli [238, 2244, 2295]. collage [2147]. collagen [1901, 2210]. collapse [1575, 1083, 3070, 1200]. collaterally [1817, 3248]. collecting [413]. collection [3418]. collections [1064]. collective [3269, 3603, 3009, 2689, 3570, 2493, 2514]. Colless [3061, 3385]. collisions [1753]. collocation [409]. colon [3288]. colonic [2692, 2767]. colonies [1221, 3199, 3161, 861, 937, 3570]. colonisation [2531]. colony [1551, 462, 1134, 3161, 3070, 2158]. Color [185, 290, 10, 386]. colorectal [2139, 3174]. colored [2603]. Colorimetry [1520]. colours [1520]. colours- [1520]. column [802, 2522, 3210]. combating [3086, 2552, 2553]. combination [2822]. Combinatorial [1912, 3017, 2303, 3615, 2383, 359]. Combinatorics [2225, 2009, 2744]. combine [2846]. combined [3326, 3209]. combining [82, 2072]. Comments [108]. common [701, 1022, 1066, 1331]. communicating [2184]. communication [1612, 3537, 728]. communities [926, 1026, 714, 3175, 752, 3203, 35, 2809, 3595, 1973, 1128]. community [2457, 396, 2478, 1530, 1412, 3616, 3512, 3370, 1411, 2929, 1756, 1692, 2374, 1890]. community-dynamical [1412, 1411]. commuting [3580]. compact [2848, 3169]. compactness [2132]. comparable [716]. Comparative [3140, 1558, 1237, 1700, 3165, 3608]. comparing [3480]. comparison [1336, 3095, 793, 1795, 3223, 662, 3542, 3364, 2398]. compartment [351, 1724, 3544, 1762, 3259, 2786]. Compartmental [294, 3178, 3030, 3414, 725, 2178, 2576]. compartmentalized [2963]. Compatibility [1595, 1070]. compatible [90, 3313]. Compensation [739, 1709]. competent [1516]. Competing [1640, 616, 839, 3020, 980, 423, 905, 545, 662, 594, 2006, 591]. Competition [2371, 2018, 461, 279, 1041, 629, 718, 2031, 66, 376, 728, 289, 1273, 3043, 1050, 139, 356, 1576, 1537, 1917, 1216, 2356, 3044, 926, 1499, 593, 727, 2817, 3004, 266, 607, 826, 2153, 3582, 746, 1105, 919, 329, 1172, 96, 1998, 2463, 1746, 2277, 1117, 1063, 221, 295, 983, 1993, 1329, 665, 3217, 325, 2316, 2861, 1357, 1682, 2772, 3297, 1033, 2597, 1065, 1199, 813, 2036, 2245, 1667, 3545, 2522, 963, 3448, 2410, 3280, 288, 742, 1566, 2867, 513, 1333, 52, 1062, 2648, 892, 1095, 2865, 3126, 1185, 3080]. competition-diffusion [3004]. Competitive [1408, 1173, 1874, 1105, 1098, 3227, 1610, 2269, 3497, 2228, 938, 1091, 730, 1378, 855, 986, 93, 3040, 1965, 2491, 2634, 1920, 215, 1805, 1516, 932, 3082]. competitors [752]. complementary [80, 738]. Complete [574, 3139, 23, 1713]. completely [310]. Complex [3004, 1122, 1307, 850, 36, 2412, 3277, 3063, 1777, 1584, 2479, 2738, 3365, 2330, 1137, 1744, 1077, 2571, 400]. complexes [2421, 3058, 2594]. complexity [3416]. compliance [2622]. complicated [661, 1163, 2862]. component [570, 3145, 3157, 2174, 611, 388, 2444, 1712]. components [1852, 459]. Composite [1916, 1707, 1619]. composition [452, 2468]. compositions [62]. compounds [1983]. compressive [3403]. compromise [697, 735, 9]. computable [2739]. computation [2263, 784, 2126, 2818, 2802, 2213, 336, 3279, 3482, 3599]. Computational [1774, 1696, 2363, 1629, 1698, 2382, 2523, 1953, 3264, 1977, 2139, 2297, 3011, 2475, 2074, 2282, 1937, 2451, 1774]. computations [1096, 1739, 1782]. Computer [146, 828, 990, 486, 308, 1803]. computer-aided [486]. Computing [1136, 1753, 3127, 2524, 1743, 3234]. concentration [2563, 3369, 3607, 1388]. concentrations [1938, 1850, 824, 825]. concept [1052, 1412, 1411, 1036, 2110, 536]. concepts [3141]. condition [290, 3592, 1746, 68]. condition-dependent [1746]. Conditional [2979, 1108, 1340, 1750, 3521, 2301, 2242, 2974]. Conditioned [99, 1874]. Conditions [657, 30, 2759, 2987, 592, 442, 2585, 1968, 240, 1533, 1925, 3519, 2275, 941, 278, 1230, 2127, 376, 2087, 2396, 3123, 3442, 3424, 386]. condom [1652]. conductance [1511, 115, 1149]. conduction [619, 307, 132, 180, 37, 125]. conductivities [878]. conductivity [1233, 3321]. cone [3552]. Conference [485]. conferring [1113]. configuration [3519]. configurations [86, 178, 2698]. confined [3187]. conflict [3335]. conflicts [914]. Conformal [2413]. conformational [2025]. congestive [1497]. conjecture [3043, 1235, 201, 2763]. conjugate [1210]. connected [3171]. Connecting [2449]. connection [2170]. connections [2578, 2616]. connectivity [2970, 2483, 3069, 1819, 1710]. consanguineous [1817]. Consequences [1957, 1329, 1469, 2253, 1318, 2247, 1248, 963, 3136, 1328, 1496, 1463]. conservation [1125, 1889, 1866, 3398]. conserving [3105, 3176, 3597]. consideration [106, 158]. considered [514, 550, 551]. consist [3329]. Consistency [1255, 1619]. Consistent [1522, 3297, 1847, 1693, 1258, 1661]. consortium [3506]. conspecific [1453]. constancy [185, 386]. Constant [268, 186, 299, 1752, 1190, 325, 401, 2073, 709, 16, 3621, 2833]. Constant-rate [268, 186]. constantly [212]. constants [1470, 1359]. constitutive [1860]. Constrained [2831, 3618]. Constraint [1534, 209, 349]. constraints [3380, 2771, 999, 1458, 821, 1905, 2443, 2822, 3130, 2545, 3019, 3233]. constructed [1561, 3186]. Construction [2375, 821, 1560, 610, 2124]. constructs [2191]. consumer [1026, 3384, 2897, 3522, 3614, 1075, 799, 2214]. consumer-one [799]. consumer-resource [3384, 2897, 3522, 1075]. consumers [2865]. consuming [2712]. consumption [78]. Contact [3409, 803, 920, 2823, 2155, 1583, 2389, 1258, 981, 2938]. contact-mediated [803]. contacts [2185, 2403]. contain [3243]. containing [135]. containment [3612]. contaminated [3231]. contamination [3070]. content [2789]. contest [1063]. context [1507, 2489, 2334]. context-free [2334]. continual [549]. continuation [3105, 3176]. Continuous [3073, 1752, 3223, 374, 1656, 2650, 1468, 490, 1648, 2469, 2619, 1394, 896, 959, 3543, 640, 696, 763, 1204, 1848, 366, 2781, 700, 430, 1714, 2611, 121, 1479, 2474, 2626, 282, 3476, 447, 910, 923, 19, 1828]. continuous-space [696, 2626]. Continuous-time [3223, 2619, 3543, 1714, 121]. continuously [3483]. Continuum [2038, 1780, 2740, 2011, 947, 604, 1795, 3005, 1824, 3260, 3140, 436, 3341, 411, 2355, 1385]. continuum-based [3341]. continuum-of-alleles [604]. contractile [2271, 2244]. contraction [504, 3214, 3444, 2654, 3298, 2223, 3320, 2072, 2248]. contractions [105, 3341]. contractive [2169]. contrast [1677, 2786]. contrast-enhanced [2786]. contributes [2800]. Contributions [290, 3168, 2127]. Control [3178, 1929, 646, 322, 19, 2723, 3189, 494, 103, 274, 2151, 1246, 1247, 1497, 2890, 2581, 1960, 2039, 2750, 2311, 3222, 3032, 564, 556, 2858, 2770, 1126, 2348, 1376, 2536, 3323, 2094, 1979, 3267, 92, 1539, 1877, 131, 479, 145, 242, 2373, 1928, 3438, 3557, 840, 3079, 3229, 3195, 1110, 2903, 3428, 488, 881, 3371, 112, 2953, 2024, 2069, 2794, 2028, 2160, 2390, 1798, 56, 278, 2166, 2378, 2990, 3288, 3585, 15, 2927, 2186, 3460, 3579, 2185, 2403, 1889, 1866, 2709, 2426, 1876, 2551, 20, 2227, 2985, 1077, 1784, 3538, 2876, 3421, 3080, 2582, 1430, 1081]. controlled [1649, 120, 2173, 2258, 1373, 2728]. controlling [2345, 3585]. controls [2463, 736, 2755]. Convected [1754]. convection [579, 519, 1964, 1597]. convenient [2597]. Convergence [1478, 330, 2684, 2296, 746, 62, 497, 768, 836, 709, 1178, 258, 72, 658, 1155, 885, 1854, 2606, 2665, 608, 739, 2496, 3031, 377, 670, 809, 366, 1442, 87, 413, 1444, 3263, 1333, 588, 335, 340, 907, 1808]. convergent [3588, 1307]. converging [884]. conversion [2907, 522]. convolution [1677]. cooperation [1957, 2979, 3560, 279, 785, 3343, 3495, 2586, 979]. Cooperative [2519, 3124, 1336, 1874, 2214, 1356, 1669]. Cooperativity [3110, 2648, 1336]. coordinated [2619]. coordinates [1900, 2715]. coordination [2223]. copy [2022]. core [3404]. corneal [2884]. coronavirus [3326]. corpus [2326]. Correction [3102, 3176, 3184, 3024, 3412, 3442, 3387]. Correlated [2978, 100, 1076, 2040]. Correlation [3015, 2123, 1343, 1817, 2455, 2077]. correlations [2945, 3527, 329, 2835, 2956]. Corrigendum [3166]. cortex [1351, 1120, 893]. cortical [3477, 1897, 2077, 2792]. corticotropin [1579, 1409]. corticotropin-releasing [1579, 1409]. corvids [2267]. Cost [3560, 2430, 2725]. Costs [1642]. cotyledons [3]. count [3079, 506]. Counter [2864, 3411]. Counter-propagating [2864]. countercurrent [413]. Counting [2304, 3115, 3041, 654, 3327, 1714]. country [3348]. coupled [2943, 570, 123, 3094, 3137, 3569, 315, 563, 535, 812, 3125, 2270, 444, 114, 147, 495, 2322, 1005, 659, 2745, 436, 3049, 379, 1267, 2688, 648, 2077, 1159, 2248, 1860, 2746, 3427]. coupling [342, 1745, 3081, 3616, 2290, 2224, 2892, 3610]. couplings [2284]. course [391]. cousins [2813]. CoV [3476, 3507]. CoV-2 [3476, 3507]. Covalent [1558, 3607]. covariance [333, 2275, 98]. Coverage [2188, 2816, 2965]. coverage-based [2816]. covers [2042, 2713]. COVID [3405, 3223, 3359, 3296, 3242, 3243, 3448, 3489, 3466, 3430, 3109]. COVID-19 [3405, 3223, 3359, 3296, 3242, 3243, 3448, 3466, 3430, 3109]. COVID-19-associated [3489]. cows [2861]. craniofacial [1521]. crawling [1472, 2342]. created [2369]. crest [3603, 3009]. Crick [1814]. crisis [3570]. criteria [2518, 704, 3510, 1241]. Criterion [1679, 103, 222, 1816, 445, 2827, 1181, 1268, 3290]. Critical [1486, 357, 2788, 621, 2547, 2623, 1811, 3375, 199, 2620, 3610, 3451]. critiques [2656]. CRNs [3526]. crocodilian [1765]. Crohn [3206]. crop [3238, 3601]. Cross [2114, 850, 1610, 2362, 1503, 1113, 1417, 3237, 733, 2580, 504, 2366, 3525, 263, 956, 3448, 2899]. Cross-diffusion [1610, 3237, 2366, 956, 2899]. Cross-diffusion-driven [2362]. cross-immunity [1113, 733, 2580, 3525, 3448]. cross-linking [263]. Cross-membrane [2114]. cross-section [1417]. crossing [108, 2167]. crossings [965]. Crossover [1468, 2261, 1871]. Crosstalk [2874]. Crow [910, 923]. crowding [417, 462, 560, 136]. crustacean [1740, 792]. crypt [2692, 2767]. crypts [2139]. crystal [1840]. crystallization [3269]. crystallographic [2019]. CSF [1500]. CSTR [1190]. CTL [2687, 2045, 2179]. cubic [706, 751]. cue [690]. cues [3006, 1257]. culling [2311, 2552, 2553, 1639, 1911]. culture [656, 1294, 402, 192, 662, 2729, 1079, 1153]. cultures [990, 1975]. cumulant [3453]. cumulants [858]. Cumulative [3432]. current [996, 3393, 241]. Curvature [1350, 947, 834]. curve [2422, 1979, 1531, 2217, 1994]. curved [212, 941]. curves [293, 2393, 2501, 1743, 3290]. cusp [2924]. cusp-shaped [2924]. cut [2437, 3154]. cutaneous [1363, 1382, 2072]. cycle [238, 2888, 1415, 1196, 2377, 975, 313, 2293, 1877, 2456, 293, 495, 2161, 2506, 2984, 3244, 164, 3516, 3312, 1867, 2095, 1635, 703, 539, 572, 1735, 575, 3290]. Cycles [623, 2324, 1576, 2169, 265, 28, 1429, 2461, 1037, 860, 272, 337, 2491, 1684, 3035, 921, 1663, 782, 199, 2618, 3623, 2209, 2023, 9, 3248, 1140, 3490]. Cyclic [3361, 1065, 230, 555, 639, 275, 2023]. cyclical [1694]. cyclin [1737, 975]. Cycling [339, 3415, 684, 1744]. cyclostat [167]. cylinder [748]. cylinder-like [748]. cylindrical [3440]. cytokine [3399]. cytokines [2686]. cytomechanics [1047]. cytometry [1044, 1166]. cyton [1732]. cytoplasm [2365]. cytoplasmic [3148]. cytoskeletal [1974, 981]. cytotoxic [505].

D [3010, 1707, 2948, 1159]. damage [3592, 3432, 871, 909]. Dangerous [2616]. dans [419, 2556, 2656, 3027]. Daphnia [1888, 692]. Daphnias [2121]. Darcy [3321]. Darwinian [2255, 3251, 1053, 3014]. Darwinism [2289]. Data [3487, 2434, 2943, 1336, 3178, 1913, 853, 3294, 2094, 2926, 1069, 3215, 2017, 2275, 3498, 2595, 1838, 3509, 3448, 2539, 1916, 2837, 2587, 3092, 2901, 1705]. Data-driven [2434]. datasets [3509]. Dating [2943]. day [3064]. day-to-day [3064]. DeAngelis [2661, 3621]. death [490, 1244, 2281, 2656, 1473, 254, 2590, 2058, 3406, 1467, 2739, 3435, 731, 303, 2947, 647, 699, 738, 3186, 3486]. death-Birth [3486, 3186]. debris [1166]. decay [347, 875, 943, 3066, 3297, 2599]. deceleration [2470]. deciding [3291]. decision [1977, 2532, 1864]. decision-making [1977]. decisions [2725, 1987, 1203, 1864]. decline [3073]. declining [946]. Decomposable [1615, 1747]. decomposing [3051]. Decomposition [216, 3183, 2729, 1635, 1703]. decompositions [1742]. Decoupled [2872, 2103]. decouples [3309]. decrease [1424]. decreases [1668]. deep [1179, 3421]. deer [2291]. defence [2113]. defense [3570]. deficient [1409]. defined [3515]. Defining [2337, 2743]. definite [1333]. definition [2060, 784, 1834, 2126, 33]. definitions [2648]. deforestation [3250]. deformation [2885, 2564, 1032]. deforming [1902, 3279]. degenerate [3044, 2899, 1000, 1107, 1106, 76]. degradation [968, 975, 1173, 2191, 3280]. degrade [2270]. degrade-and-fire [2270]. degrading [3556]. degree [2123, 1919, 2089]. degrees [2623]. Delay [2432, 2442, 2985, 350, 1507, 2312, 1246, 1247, 1497, 1486, 708, 619, 972, 1209, 1369, 1393, 3252, 1672, 1284, 320, 1539, 1765, 2928, 1467, 2076, 3181, 721, 1039, 1952, 435, 2029, 344, 3297, 2632, 1304, 180, 1123, 3448, 2350, 2927, 1760, 912, 83, 1295, 447, 2898, 275, 1841, 1414, 3209, 3420, 1084, 1185, 3283, 3427, 3334]. delay-difference [708]. delay-differential [1765]. delay-diffusion [344]. Delayed [1186, 2227, 3030, 1486, 769, 3083, 75, 3582, 471, 919, 1162, 2241, 3609, 245, 997, 2229, 3232, 2749, 3219, 1502]. delayed-logistic [245]. delayed-recruitment [75, 471]. Delays [509, 206, 715, 1004, 556, 2638, 1085, 74, 94, 140, 3422, 1738, 3268, 3525, 1114, 1321, 3549, 1220, 488, 546, 1666, 3049, 3556, 1241, 1287, 2119]. Delbrück [1276]. deleterious [250]. deletion [3517]. delimitation [2609]. delineates [3568]. delivery [3275, 2249, 2007]. deme [2253]. demes [1442]. Demographic [3254, 2360, 2361, 1742, 3088, 2180, 3415, 2812, 603, 2814, 3513, 2678, 1255]. demographically [1194]. demographics [883, 3021]. Demography [483, 618, 1923, 2860, 451, 2033, 282, 1435, 907]. Demyelination [2667]. dendrites [2142]. dendritic [567, 3089, 2000, 2224, 1687, 1149]. dendro [2224]. dendro-dendritic [2224]. dengue [1182, 1375, 1098, 2335, 3470, 3488, 3054]. densities [3348]. Density [929, 1719, 1248, 1328, 194, 1237, 894, 2336, 1963, 470, 2128, 312, 3025, 3591, 2032, 2164, 2588, 1766, 883, 1172, 208, 686, 999, 85, 640, 830, 2732, 1297, 1318, 3435, 2339, 709, 1148, 2599, 116, 1511, 205, 1079, 2235, 540, 2502, 912, 888, 1145, 50, 3126, 3331]. density- [205]. Density-dependent [929, 1719, 1248, 1328, 194, 312, 3025, 1766, 883, 1172, 208, 686, 999, 830, 2732, 1297, 3435, 709, 1145, 3126, 3331]. dependence [1880, 417, 338, 1467, 1762, 751, 820, 2494, 912]. dependency [802, 3]. Dependent [591, 3058, 2312, 1963, 470, 1303, 2497, 312, 564, 547, 1851, 657, 2226, 2252, 774, 3276, 3025, 2714, 1061, 1425, 798, 3446, 200, 543, 1766, 883, 349, 1172, 3268, 2928, 3406, 208, 686, 999, 85, 3549, 1746, 830, 2732, 1297, 1289, 1310, 3435, 2388, 1948, 3141, 2929, 1139, 3337, 693, 1457, 1293, 2781, 2895, 1042, 917, 2057, 709, 929, 1719, 2916, 383, 285, 2165, 641, 1248, 1819, 1511, 43, 205, 3114, 2143, 1916, 1743, 236, 1566, 1881, 540, 234, 888, 53, 736, 1145, 1328, 1512, 1496, 1876, 3408, 194, 2782, 50, 3126, 149, 819, 1149, 1305, 3488, 372, 3331, 1207]. dependent [3150, 1290, 988]. depending [3210]. depletion [1125]. depolarization [3477]. deposition [1840, 2133, 1749]. depth [690, 241, 1811, 2035]. Derivation [1493, 1723, 3260, 1937, 2568, 2899, 3072, 2724, 1056, 2235, 2216]. derivatives [3230]. derived [1415, 3258]. Deriving [1439]. dermatitis [2938]. descent [2799, 511]. describe [1146]. describing [817, 3265, 479, 2844, 997, 2749]. description [3347, 1252, 2282, 859, 1137]. descriptors [2617]. Design [1958, 1929, 1458, 2383]. destabilisation [1068]. Destabilization [3122]. destruction [2996, 2800, 400]. desynchronization [3074]. detailed [2338]. detect [2504]. Detecting [3372, 2337, 2803, 3354]. detection [1351, 3504, 3098, 321, 2565]. determinacy [1357, 1356]. Determination [1359, 3471, 2239, 1944, 2613, 835]. determine [2180, 3094, 2889]. determined [3394, 1550]. determines [2933, 340]. Determining [2151, 2535, 3059, 1732, 1837, 3006, 3231, 3156, 3329]. Deterministic [304, 208, 1374, 1501, 3207, 1748, 2799, 2449, 900, 3223, 2590, 1138, 1301, 2528, 1286, 1945, 2353, 765, 1330, 418, 481, 480, 581, 53, 703, 572, 1855, 1859]. deterministic/probabilistic [572]. Deux [3027]. Developing [2706, 2337]. Development [2213, 1257, 1373, 1551, 1875, 1302, 1101, 2139, 2051, 1134, 568, 1787, 2740, 744, 1885, 1660, 1403]. developmental [1449, 1373]. deviations [1768]. devices [1782]. di-cographs [2660]. diagnostics [1313]. diagrams [681, 26, 3360]. dialects [2230]. Diallelic [630, 1927, 838, 833, 845, 910, 923]. diameter [3388]. diapause [2915, 2116]. dichotomy [3344, 3014]. dichroism [389]. Dictyostelium [705, 1573]. Diekmann [2109]. dielectric [2939]. diet [1245]. dietary [3237]. difference [708, 3611, 473, 2952, 663, 382, 445, 2286, 670, 87, 2772, 3375, 3606, 1206, 617, 889, 51, 3219, 942, 1751]. differences [346, 25]. differencing [1653]. different [3072, 1375, 3348, 509, 3229, 2816, 989, 2167]. Differential [904, 3238, 1988, 1508, 2016, 1611, 10, 470, 3574, 1915, 3265, 2147, 619, 3562, 3496, 3000, 1284, 2432, 1765, 365, 2291, 2928, 2766, 30, 497, 2613, 1903, 670, 3491, 1645, 3465, 1878, 997, 2632, 1476, 1722, 2947, 1665, 866, 1617, 1969, 2370, 2378, 2927, 305, 658, 889, 52, 2898, 2305, 885, 1414, 16, 2774, 3427]. differentially [2640, 2802]. Differentiated [2433, 1946]. differentiation [2253, 1221, 701, 2005, 2766, 3439, 322, 3013, 287, 2426, 1373, 1174, 1153]. difficile [2364, 2709]. diffuse [2200, 1712]. diffusing [2328, 2316]. Diffusion [719, 1622, 448, 1610, 1835, 1843, 1593, 3050, 2982, 862, 2943, 2197, 1361, 570, 3291, 223, 3158, 2606, 1791, 592, 2780, 1892, 3235, 2890, 2549, 1169, 2115, 3445, 1902, 2930, 3172, 3044, 3155, 3237, 2882, 2991, 556, 3250, 1439, 3426, 330, 967, 2441, 2435, 2226, 3576, 1750, 1275, 171, 309, 2817, 3004, 2465, 1445, 1324, 3201, 200, 1154, 2509, 3499, 2296, 524, 409, 2051, 3484, 286, 3125, 3066, 2366, 3145, 3091, 3042, 349, 2473, 3372, 667, 1068, 1045, 79, 3483, 666, 1180, 2359, 41, 3011, 141, 2897, 3522, 1729, 181, 429, 477, 836, 395, 468, 3157, 3317, 2074, 3194, 1873]. diffusion [2260, 325, 757, 748, 3286, 956, 693, 3068, 136, 3119, 3609, 344, 287, 880, 1932, 945, 977, 709, 2995, 1882, 2362, 2615, 2597, 2254, 1556, 1580, 172, 215, 813, 2953, 1060, 453, 157, 1046, 1093, 297, 2693, 2666, 1511, 2484, 135, 2899, 2840, 2070, 2768, 2539, 72, 179, 1038, 1000, 1107, 1007, 1578, 341, 432, 600, 1927, 1368, 3366, 987, 2961, 296, 326, 1655, 1990, 3442, 3424, 810, 3535, 2910, 3093, 3126, 3299, 358, 1736, 3198, 993, 1106, 2167, 1661, 1032, 1964, 1597, 3315, 2066]. Diffusion-driven [862]. diffusion-ODE [2615]. diffusions [1728, 3057]. Diffusive [1581, 819, 3043, 442, 2510, 2252, 2296, 416, 2661, 1379, 3363, 401, 458, 1065, 2184, 259, 3564, 3280, 3333, 3209, 3282]. diffusivities [862, 1202]. diffusivity [830, 3449]. digestion [2878]. digraphs [1635]. dilemma [2586, 815]. diluted [3342]. dilution [2508]. dim [3552]. dim-light [3552]. Dimension [527, 1505, 1612, 1662, 1927]. dimension-reduction [1612]. dimensional [2324, 579, 1791, 1246, 1247, 1136, 3155, 1968, 1225, 1432, 1437, 2924, 1472, 1575, 1728, 265, 3102, 3001, 3214, 3444, 246, 1816, 2557, 2297, 667, 666, 852, 836, 840, 2634, 1821, 2894, 745, 70, 3391, 1978, 2207, 2627, 1192, 3122, 2137, 2550, 3185, 598, 1441, 1389, 941, 153, 1388, 1028, 1368, 3386, 2920, 2537, 1077, 1014, 819]. Dimensionality [3234]. dimensions [2681, 93, 3374, 859]. dimers [1229]. dimorphism [2544]. dioecious [190, 1046, 838, 845]. dioxide [3241, 281, 2522]. diploid [1088, 2804, 1018, 2465, 1293, 1770, 1418, 2659, 2472, 1277]. diploids [2168]. diplophase [67]. dipolar [1135]. dipole [420]. Direct [69, 2245, 2867, 1979, 2124, 1265, 1663, 2891]. directed [2133, 3558, 3553]. direction [1664, 1022, 1066, 3118]. directional [1035]. directly [1944]. Dirichlet [376]. dis [1546]. disambiguation [3435]. disappearances [1279]. disassembly [3267]. disassortative [3389, 3303]. disasters [729]. discoideum [705, 1573]. discontinues [419]. discontinuous [1648, 419]. discordance [3398]. Discrepancies [2959]. Discrete [3240, 1567, 2456, 867, 430, 1142, 2280, 1468, 1615, 3208, 490, 801, 1322, 3073, 1786, 1362, 708, 2789, 1492, 1711, 1943, 2469, 1175, 954, 1596, 95, 549, 855, 986, 1049, 1766, 3091, 2744, 3543, 696, 763, 276, 2977, 3355, 3462, 377, 665, 2487, 2759, 1448, 1433, 785, 1365, 2073, 1177, 3122, 3260, 945, 2740, 559, 2650, 1312, 3049, 2521, 2883, 3476, 1062, 1747, 3623, 1323, 1695, 1100, 1255, 1449, 1084, 1314, 1463, 3577, 934, 2675, 189, 1871]. discrete- [3543]. Discrete-time [867, 3208, 1786, 549, 1766, 696, 2977, 2487, 1433, 3122, 2883, 3623, 2675]. discrete-time-continuous-time [3476]. discretization [2510, 2555]. discretized [3105, 3176, 3165]. Discussion [17, 1834]. Disease [1766, 883, 2558, 2374, 1591, 2548, 2197, 2228, 2784, 1559, 3326, 2951, 3158, 3468, 203, 248, 3566, 991, 3532, 2402, 434, 776, 2638, 3223, 3624, 1370, 3302, 918, 3601, 2067, 3353, 1182, 516, 1773, 3593, 2291, 242, 2085, 3348, 2317, 2844, 2276, 3438, 1713, 374, 2903, 3258, 587, 2616, 1428, 2429, 1919, 3503, 2089, 224, 3262, 3622, 2178, 1313, 3206, 1527, 3279, 2671, 2501, 2652, 2927, 3390, 2185, 2403, 3476, 2749, 207, 3538, 3563, 2561, 3146, 3555, 2066, 3417, 1111]. disease-free [2951]. Disease-induced [1766, 2228, 3566]. diseased [3029]. diseases [2723, 3584, 3167, 2805, 1598, 1364, 2311, 788, 2879, 330, 1097, 784, 831, 675, 1133, 2841, 3432, 1829, 1639, 3521, 214, 331, 2345, 757, 1868, 917, 882, 2386, 2583, 767, 2980, 1662, 1481, 2630, 1519, 2808, 2283, 1140, 3273, 988, 1081]. disequilibria [2530]. disequilibrium [347, 1999, 902]. disjoint [1064]. disk [2177, 1761]. disk-shaped [2177, 1761]. disorientation [389]. Dispersal [1050, 2206, 2467, 2727, 3340, 1685, 431, 552, 2071, 2993, 3322, 3426, 1750, 2911, 309, 2159, 3591, 1154, 3530, 2015, 2797, 3400, 338, 3304, 3543, 1352, 1746, 398, 448, 1640, 1315, 1426, 1040, 2316, 2242, 2320, 3437, 1719, 691, 1353, 2754, 2481, 1295, 2288, 3408, 1287, 3082, 819]. Dispersal-induced [3340]. disperse [462, 560]. dispersed [2270, 944, 1368]. dispersing [697, 217]. Dispersion [763, 1661, 2518, 698, 2360, 2361, 512]. displacement [139]. disposition [2264, 2655, 2925, 391]. Dissection [671, 2885, 1498, 2570, 951]. dissipation [3330, 1334]. dissipative [440]. dissociation [2104]. distal [2249]. distance [3292, 2791, 2059, 2015, 2934, 3060, 3588, 3441, 1419, 747, 1994, 3508]. distance-based [3588]. distances [2535, 2601, 1887, 3127, 3397]. distancing [2143]. distensibility [1337]. distinct [277]. distinguish [2849]. distinguishable [251]. Distinguishing [3215]. Distributed [1838, 2151, 793, 715, 3222, 1026, 2591, 2621, 3422, 3406, 1130, 3452, 404, 863, 573, 3049, 2997, 3152, 194, 3208, 3420, 1997, 1185]. Distribution [2934, 3068, 1942, 1276, 2123, 3347, 1135, 2255, 2882, 2991, 2924, 2483, 2321, 450, 476, 2853, 1308, 317, 808, 437, 554, 751, 2697, 2939, 514, 110, 1021, 2260, 2161, 1337, 1645, 2325, 4, 1921, 1148, 3506, 277, 1015, 941, 1205, 2334, 2525, 2710, 583, 217, 314, 847, 1368, 597, 3187, 539, 512]. distributional [468]. Distributions [2279, 2464, 3234, 3058, 3485, 2022, 3083, 1048, 402, 3225, 241, 746, 3008, 2603, 520, 511, 1151, 589, 995, 13, 852, 3182, 3183, 2975, 2685, 1949, 2284, 111, 3545, 2650, 2251, 994, 2785, 483, 391, 492, 372, 1557]. disturbance [932]. Disturbances [980, 552, 1216, 549]. disturbed [2788]. distylous [2904]. divalent [46]. divergence [2955, 595]. diversity [2796, 1991, 3611, 3237, 1995, 3530, 1296, 3304, 3141, 364, 2082, 3327, 3536, 2196, 2410, 932, 3017, 3398]. diversity-dependent [3141]. divided [3348]. division [2409, 740, 2149, 2456, 554, 1183, 2062, 2327, 3185, 572, 2368, 2543]. division-linked [2543]. DNA [1560, 3454, 2219, 1397, 209, 3592, 2032, 389, 1561, 871, 1334, 1524, 1131, 158, 1309, 1259, 2710, 909, 2172, 2163, 1438, 2188, 2565]. DNA-drug [389]. Do [1146, 2841, 2204, 1571, 2845]. Does [1316, 2731, 2203, 705, 1157, 1424, 2821, 2531]. dogs [3373]. Doi [3245]. Dolores [2736]. domain [1649, 1231, 1321, 1842, 1978, 813, 2620]. domain-shape [813]. domains [1324, 1114, 748, 1882, 453, 1990]. domestic [3373]. dominance [412, 1715, 1962, 869, 1987]. dominant [2861, 53]. done [2521]. doomed [3571]. Döring [3587]. dormancy [3200, 1790, 2399, 3098]. dorso [1634]. dorso-ventral [1634]. Dose [917, 2407, 2593, 2393, 2836, 1406, 909]. Dose-dependent [917]. dose-response [2393]. doses [1265]. dosing [2505, 1837, 2822, 2990]. Double [2039, 463, 3454, 2437, 154, 2172]. double-helix [3454]. double-strand [2172]. doubling [1324, 1176, 345, 3462]. doubly [2899]. down [2567, 2964, 2204]. downwards [1600]. doxorubicin [3381]. DPD [2480]. drainage [2542]. drift [1895, 1728, 1835, 1664, 1202]. drive [3530, 3254, 3416, 3513]. driven [1794, 1971, 2737, 2439, 3237, 3358, 3004, 3458, 345, 3575, 3487, 1729, 2629, 2260, 2316, 3542, 2362, 2239, 3207, 870, 1930, 862, 3479, 2434, 3281]. driving [1818]. Droop [984, 1099]. dropping [2905]. Drosophila [2946, 1634]. Drug [1260, 2300, 2264, 2655, 3275, 1473, 1073, 2122, 1296, 2806, 2925, 1265, 303, 2061, 1941, 1803, 3209, 391, 1860, 2007, 3248, 389]. Drug-resistant [1260]. drugs [2266, 3514, 1836]. drying [1462]. dryland [2527, 3400]. drylands [2830]. DTI [2414]. DTI-based [2414]. Dual [565, 3125, 3497]. dual-strain [3497]. duality [2530]. duals [2722]. duct [413]. ductal [1422]. ductus [2201]. due [406, 140, 501, 1630, 3070, 682, 1628]. Dulac [1816]. d’un [419]. duplicate [226]. Duplicates [1593]. duplication [3041, 672]. duplication-loss [3041]. duplication-loss-transfer [3041]. duplications [2521]. duration [2933, 2894, 1020]. during [3300, 3131, 3530, 3179, 3084, 3216, 1634, 1240, 3595, 827, 3623, 2950]. durotaxis [2916]. dusky [646]. duty [3244]. dwell [2975]. Dynamic [373, 1773, 1975, 261, 1457, 1261, 882, 3230, 2233, 1504, 3071, 2702, 1023, 1269, 2441, 2435, 3504, 3515, 1394, 3459, 1952, 1329, 3369, 1677, 1122, 1787, 2721, 2395, 893, 1878, 1248, 2612, 2608, 963, 2499, 1328, 1496, 1047, 3149, 2891]. Dynamical [470, 1535, 650, 2280, 3556, 2503, 889, 1463, 3485, 2890, 903, 2321, 2262, 1530, 1056, 1412, 1367, 1411, 3416, 2925, 262, 1518, 1387, 1280, 2883, 614, 829, 3342, 799, 3126, 3524]. dynamically [2427]. Dynamics [1398, 3171, 1689, 2308, 1432, 1633, 2832, 1943, 2580, 3151, 3161, 126, 1321, 2268, 3170, 398, 2897, 3522, 1227, 1325, 2192, 3204, 2207, 1868, 2960, 430, 2490, 3049, 379, 1468, 1615, 1627, 2129, 3146, 2723, 2267, 2637, 1946, 502, 1545, 823, 3274, 2421, 2300, 1113, 3468, 3020, 3163, 578, 606, 2237, 2683, 2799, 248, 2312, 2707, 2510, 1960, 3547, 2439, 1320, 2091, 1674, 2497, 2128, 1944, 947, 2377, 564, 2452, 2687, 3394, 1648, 858, 547, 346, 2978, 1966, 3103, 1009, 1119, 1437, 1851, 818, 743, 2701, 2448, 3502, 3624, 3036, 2138, 75, 3131, 2232, 2168, 661, 2923, 1260]. dynamics [2465, 571, 503, 982, 2828, 2641, 3567, 2513, 1370, 3277, 3422, 2838, 3190, 3135, 3601, 1147, 2153, 1817, 2199, 1726, 1222, 2801, 3586, 1484, 3008, 2432, 471, 323, 2557, 3359, 3362, 722, 2352, 2297, 3268, 2106, 1346, 1488, 1327, 1531, 2613, 281, 85, 2335, 2085, 3549, 860, 1938, 2002, 510, 3295, 2431, 1865, 1059, 1584, 1297, 1310, 3435, 1338, 1572, 2331, 395, 1965, 1872, 3317, 665, 1713, 2634, 3306, 1017, 1848, 2929, 1334, 1789, 1861, 2894, 1213, 1092, 1883, 2124, 2375, 2495, 3024, 2187, 1469, 2453, 2972, 785, 820, 2459, 956, 1391, 3337, 693, 2700, 887, 3258, 3531, 2010]. dynamics [2781, 2616, 2827, 1341, 1682, 2029, 2429, 2809, 2845, 3609, 119, 2699, 1365, 1839, 2057, 2627, 3259, 2061, 3310, 1838, 2327, 950, 475, 2538, 1423, 2745, 3561, 2999, 1055, 3595, 3132, 3133, 2947, 2498, 1579, 3587, 1924, 2851, 884, 1313, 2921, 2184, 1526, 2354, 2118, 3356, 2330, 1436, 1730, 2836, 56, 116, 2493, 482, 3278, 3613, 1547, 2166, 3448, 3568, 973, 2642, 2290, 2718, 3050, 2405, 273, 1259, 3021, 1935, 2397, 2671, 1825, 3486, 2468, 1163, 2862, 2141, 1577, 2514, 2834, 3280, 1958, 1970, 744, 3197, 2349, 1381, 1574, 3064, 2769, 3579, 2867, 1583, 3232, 1886, 1747, 53, 1019, 1409, 2426]. dynamics [3366, 736, 910, 923, 935, 1145, 1095, 1876, 3408, 1516, 2680, 3208, 2961, 499, 361, 907, 1744, 1890, 3019, 2922, 957, 1519, 3246, 1710, 2179, 2985, 2865, 3604, 1077, 3028, 2950, 1106, 1185, 1305, 2283, 3331, 3554, 2307, 2774, 2561, 3080, 3273, 3427, 3555, 2066, 2720, 400, 1081, 2533]. dysrhythmias [345].

earliest [2981]. Early [3109, 1403, 2306, 3158, 2602, 975, 2448, 530, 1260, 1422, 751, 1034, 1770, 1313, 3098, 3354]. early-season [1260]. echinoids [1335]. eco [2439, 2478, 3261, 3595]. eco-epidemiological [3261]. eco-evolutionary [2439, 2478, 3595]. Ecological [2180, 2817, 1533, 867, 2443, 1532, 3468, 2931, 3107, 2206, 1173, 1056, 3323, 3336, 2511, 2111, 3175, 3182, 752, 295, 688, 1883, 1055, 2776, 2125, 1128, 1279]. ecology [1439, 2232, 2058, 2952, 2086, 1299, 2063, 1858, 2917, 814, 3371, 2110, 2494, 3169, 614, 1635, 1581]. economy [2418]. ecosystem [313, 1211, 3481, 149, 2746]. ecosystems [3382, 440, 3250, 106, 3597, 87, 1663, 267, 2639, 3016]. Edge [2155, 1274, 1143, 3284, 2995, 2599, 3595, 2178, 3486, 2167]. edge-based [2178]. edge-length [3284]. edges [2905, 2839, 525]. Editorial [1218, 1051]. Effect [2329, 813, 443, 1554, 1461, 1462, 78, 2517, 2571, 2755, 1559, 2951, 2986, 3275, 466, 462, 2090, 3044, 2430, 2725, 2407, 240, 2154, 3565, 332, 390, 371, 1898, 2181, 1464, 3567, 3504, 3353, 3116, 2593, 549, 846, 1877, 1837, 263, 3543, 2757, 1132, 752, 3557, 1297, 1374, 374, 1265, 2861, 3375, 3606, 2073, 1580, 2081, 1999, 195, 2410, 2834, 1941, 594, 3272, 1578, 314, 2001, 1803, 2749, 3333, 113, 2948, 1446, 1287, 2444, 3541, 957, 1519, 1475, 2272, 1922, 2569, 2910, 3538, 1253]. Effective [1938, 1891, 1919, 2089, 1202, 3563, 1494, 2313, 2416]. effectively [439]. Effectiveness [3243, 2039]. Effects [796, 926, 3426, 1715, 241, 1933, 441, 509, 1040, 1684, 2610, 385, 467, 2836, 941, 3613, 614, 912, 1490, 1197, 3398, 2306, 939, 1757, 777, 2643, 2594, 2677, 1350, 2091, 818, 1156, 1492, 1097, 171, 1129, 2159, 493, 2633, 3422, 2032, 831, 3530, 1217, 1683, 3161, 365, 1045, 3085, 3305, 3618, 2266, 1729, 625, 1945, 1914, 2996, 136, 2029, 247, 2061, 1882, 488, 3622, 1652, 3356, 58, 2666, 116, 717, 2350, 3298, 3381, 2481, 2867, 513, 1125, 3386, 2087, 1326, 2135, 819, 1031]. Efficiency [3349]. Efficient [2263, 1701, 2803, 3092]. efflux [3150]. effort [2073, 199, 1613]. egg [623, 568, 1787, 2540]. egg-larval [623]. eggs [568, 2183, 2117]. Eigen [624]. Eigenanalysis [284]. eigenmodes [1351]. Eigensolutions [2886]. eigenvalue [2456, 2416, 2754, 2680, 512]. Eikonal [2981, 834]. eikonal-curvature [834]. either [2817]. elastic [2869, 3252, 1620, 1389]. elasticity [915]. electric [3089, 1745, 682, 872, 2967, 3514, 2537]. electrical [3402, 1176, 3388, 2224]. electro [3279]. electro-mechanical [3279]. electrocardiology [420]. electrodiffusion [2843, 1842]. electrodiffusive [3211]. electrolyte [2024, 2069, 3187]. electroneutral [3187]. Electropermeabilization [2222]. electrophysiology [2454, 2160]. electroporated [3514]. electroporation [3501]. electrostatic [2939]. element [1986, 3089, 2843, 595, 410, 1770, 3279, 1754, 3514, 2072]. Elementary [2665, 2427, 2973, 3533]. elements [841, 694]. elicited [1579]. elimination [2096, 505, 2600]. elinguis [1286]. elliptic [814, 2964]. elongation [2065, 1630, 387, 582]. Embeddability [3454, 3026]. embedded [1852]. embedding [3213]. embeddings [2596]. embeds [2156, 2935]. embryo [3010, 2233]. Embryonic [1787, 975, 2909, 2506, 928]. EMDUniFrac [2802]. emergence [2255, 3448, 979]. emerging [3584, 2457, 1693]. emphasis [1983]. Empirical [246]. enable [3453]. enables [3400, 3537, 2698]. encephalitis [1543]. encircled [2604]. encode [3218, 2273]. Encoding [1717]. encodings [2044]. end [2599, 3320, 1111]. endangered [2424]. endemic [214, 374, 2960, 843, 1020]. endemicity [433]. endocrine [1884]. endosome [1501]. endotacticity [2475]. endothelial [3116, 1690]. endowed [3380]. endurance [766]. enemies [2643, 2672]. enemy [3541]. enemy-pest [3541]. Energetics [1825, 938, 2120]. energies [2577]. Energy [1261, 2891, 1336, 2336, 1091, 3372, 3384, 2939, 568, 1787, 1457, 754, 1701, 2235, 2343]. engineered [1568, 2191]. engulfing [3537]. enhanced [1677, 1392, 1579, 2786]. enhancement [525]. enrichment [1013, 1460, 1790, 432]. ensemble [1700, 1701]. ensembles [3558]. Ensuring [2750]. Entanglement [1515, 1602]. enterica [1971]. Entrainment [2567, 1934, 601, 1741, 345, 495]. Entropy [907, 1455, 850, 2854, 2645, 3559, 1935, 3071, 340]. entropy-based [2645]. entry [2160]. Enumeration [2848, 3313, 2193]. environment [3551, 3167, 523, 2176, 1764, 2208, 2281, 2231, 2411, 2556, 2656, 3027, 1216, 1517, 1752, 371, 2911, 174, 266, 3277, 1223, 2528, 338, 1347, 3202, 696, 243, 398, 848, 1315, 3340, 757, 35, 1868, 277, 3121, 227, 718, 3545, 3589, 3339, 3412, 166, 2378, 687, 737, 213, 599, 2867, 3232, 3526, 3366, 979, 296, 3022, 3282, 3505, 2283, 2755, 2963, 3417]. environmental [2180, 2105, 1571, 2603, 1037, 948, 3305, 3416, 2124, 941, 1197, 2374, 2396]. environmentally [1533]. environments [3020, 301, 2549, 3200, 3155, 3103, 332, 390, 2071, 316, 304, 2797, 3101, 2102, 2406, 1671, 3085, 2732, 2852, 3040, 2050, 2126, 2942, 2357, 3099, 3193, 1645, 54, 2360, 2361, 2320, 3310, 3117, 2754, 1668, 2319, 1936, 1886, 1481, 361, 483, 618, 1625, 1510]. environnement [2556, 2656, 3027]. Enzyme [716, 265, 3094, 3319, 3607, 161, 1850, 3556, 305, 394, 1434, 1490]. enzyme-catalysed [1434, 1490]. enzyme-catalyzed [265]. enzyme-substrate [716]. enzymes [892]. enzymic [716]. Ephaptic [1745]. Ephedra [585]. Epidemic [2945, 3445, 406, 1589, 2352, 1973, 2140, 2635, 1408, 2228, 3178, 1023, 1923, 1366, 2176, 777, 2237, 1598, 1764, 2003, 2411, 2556, 1959, 2123, 2905, 3379, 2707, 2890, 1004, 3171, 2777, 3574, 2422, 1320, 1915, 1670, 2787, 3523, 967, 3069, 3562, 2741, 1208, 1085, 1209, 1369, 517, 352, 3414, 3190, 2914, 2296, 2528, 3179, 1067, 3359, 1734, 1766, 3525, 3165, 1815, 257, 2359, 3616, 3512, 920, 1039, 3557, 2753, 3462, 1508, 786, 1632, 2814, 2623, 3306, 2162, 2742, 3194, 3529, 3195, 3519, 3378, 949, 2610, 3573, 3242, 1868, 2960, 2895, 3365, 3465, 3602, 3259, 3437, 2826, 1392, 2900, 3352, 2165]. epidemic [2941, 3356, 1330, 3109, 3564, 3114, 418, 481, 480, 581, 3270, 348, 555, 3173, 1071, 460, 531, 3064, 3062, 1614, 1930, 426, 425, 3333, 2257, 275, 3623, 1841, 2030, 2285, 101, 3100, 3118, 1519, 392, 3535, 358, 2704, 2471, 2571, 3577, 3421, 2833, 189, 1242, 1907]. Epidemics [288, 3167, 1269, 1398, 1758, 3264, 2457, 2673, 2726, 466, 240, 3527, 2579, 3432, 2919, 2302, 2649, 2268, 2533, 1928, 2047, 2353, 2721, 433, 3243, 1214, 1730, 2500, 1651, 870, 1583, 2733, 83, 2499, 3299, 1020, 1111]. épidémie [2556]. Epidemiologic [202, 1220]. Epidemiological [2440, 733, 675, 3596, 3468, 351, 2154, 3224, 3406, 1090, 1535, 3054, 721, 816, 1952, 2850, 3228, 879, 569, 650, 1226, 3261, 1400, 2921, 3156, 2350, 2026, 574, 1748, 2709, 3497, 3554, 1828]. epidemiology [1559, 2458, 3302, 3287, 2917, 1074, 3511, 2125]. epidermal [822, 928]. epidermis [1918, 2584]. epileptic [3477, 3492]. Epistasis [711, 3394, 3447, 1178]. epistatic [913]. epithelial [2909, 1893, 231, 2054]. epithelium [2884]. epitope [2687, 2832]. EPO [2990]. equal [2631, 3283]. Equality [2436, 2415]. Equation [2301, 1809, 1534, 3160, 740, 868, 1757, 2799, 1507, 417, 470, 2356, 1674, 3238, 1427, 619, 3328, 493, 352, 2489, 2149, 3048, 473, 2134, 3067, 2771, 365, 734, 2928, 2766, 1256, 2076, 64, 613, 830, 768, 214, 2727, 562, 670, 395, 3059, 1626, 2685, 3491, 834, 435, 731, 2981, 2903, 1842, 1187, 245, 1226, 1878, 997, 2632, 132, 249, 2121, 2498, 172, 1665, 1387, 782, 866, 2689, 284, 2568, 2669, 2785, 1461, 51, 83, 2749, 201, 2898, 113, 3219, 2305, 1112, 361, 3245, 1414, 16, 63, 993, 1964, 631]. equation-free [1674]. Equations [1611, 1641, 1618, 356, 772, 350, 431, 2312, 769, 971, 904, 1653, 1875, 240, 3265, 1646, 2147, 3562, 3496, 972, 607, 3000, 2843, 416, 1284, 1657, 2404, 1816, 2952, 1765, 3540, 30, 1068, 12, 2613, 1180, 41, 663, 1769, 382, 445, 670, 216, 1299, 899, 325, 3519, 821, 1645, 87, 2827, 3375, 3606, 1192, 977, 1719, 617, 2999, 1722, 215, 336, 1553, 2664, 453, 2298, 2299, 259, 984, 2776, 1511, 135, 973, 1617, 3114, 37, 72, 2927, 1000, 1107, 2688, 305, 889, 47, 367, 52, 998, 885, 1326, 2135, 76, 2338, 541]. equilibration [2854]. Equilibria [1548, 561, 622, 913, 2234, 910, 923, 2951, 2959, 2799, 558, 1646, 3565, 2441, 2435, 503, 2588, 2853, 225, 843, 267, 1251, 22]. Equilibrium [930, 3496, 1528, 1293, 632, 273, 847, 1534, 1503, 2798, 3144, 529, 330, 343, 1147, 263, 1815, 7, 2313, 86, 178, 252, 325, 401, 2960, 709, 1444, 2599, 11, 3510, 72, 658, 1333, 447, 113, 970]. Equivalence [1124, 2189, 1036, 2544]. equivalent [2634, 507]. eradicate [2841, 1613]. eradicated [1836]. Eradicating [1639]. eradication [2402, 3076, 2000, 2558, 2560, 2274]. ergodic [954, 2]. ergodicity [3573, 361]. Erik [1924]. ERK [2043, 3151, 2966]. Erlang [3222]. Erlang-distributed [3222]. Erratum [91, 188, 218, 414, 422, 446, 455, 534, 576, 2108, 2441, 2553, 2662, 197, 2533, 2183, 642, 2024, 664, 2612, 2101, 2370, 2403, 1889, 2135, 2173, 128]. error [60, 1509, 942]. erythropoiesis [2340, 2131, 2388, 2107]. escape [2365, 2479, 1501, 815, 408]. Escherichia [238, 2244, 2295]. ESMTB [1594, 1621]. ESMTB-European [1621]. esophageal [3585]. ESS [976, 2771, 869, 704, 829, 3019]. essential [1576]. essentially [1489]. establish [2477]. establishing [2770]. Establishment [857, 222, 3052, 2486, 1271]. estimate [264, 3441, 2539]. estimates [2509, 3622, 160, 17, 3610, 319]. Estimating [2457, 2090, 2149, 3129, 549, 2837, 934, 552, 271]. Estimation [853, 2649, 1660, 2167, 891, 2606, 431, 793, 1724, 2807, 3066, 2294, 2473, 3284, 1729, 2609, 2339, 757, 2678, 673, 1954, 2148, 1693, 17, 3198, 1820, 1495]. estimations [2630]. estimator [3483, 3031, 1847, 3508]. estimators [1067, 2850, 1212, 1619]. Euclidean [2596]. euglycaemic [1456]. euglycemic [1617]. eukaryotic [2106, 2792]. Eulerian [999]. European [1594, 1621]. eutrophication [2736]. evaluating [3264, 32]. Evaluation [2726, 3010, 199, 17, 3480]. evasive [1569]. event [3224, 2976, 1121]. event-labeled [2976]. event-triggered [3224]. Events [1522, 2959, 3047, 2860, 2762, 2759, 887, 2956, 1438]. eventual [2788]. eventually [3099, 3193]. Evidence [512, 1663]. Evolution [3380, 3155, 3237, 1633, 1851, 2993, 1750, 2015, 1746, 871, 2242, 1910, 2320, 2036, 1659, 2586, 652, 2046, 222, 1489, 841, 1818, 466, 1767, 2255, 1957, 1985, 3474, 412, 3265, 527, 2711, 1673, 3624, 2203, 1445, 2058, 3515, 1154, 3134, 3037, 1052, 1348, 2541, 2332, 1533, 2461, 1352, 1925, 1053, 3588, 2238, 1315, 1426, 368, 1684, 1509, 3141, 3597, 1883, 668, 2360, 2361, 1423, 1057, 11, 3495, 2110, 1580, 1714, 694, 1060, 2217, 2598, 2669, 2883, 1616, 2381, 2095, 624, 2997, 1941, 1404, 2112, 499, 2630, 2492]. Evolutionarily [1425, 2099, 2995, 1725, 2308, 1235, 1437, 851, 2496, 2242, 542]. Evolutionary [502, 3020, 1522, 2497, 2071, 571, 1061, 3277, 1394, 1352, 1059, 1572, 1961, 2793, 3259, 775, 1055, 2611, 2354, 3162, 2493, 482, 1353, 2718, 944, 2007, 3335, 2218, 2180, 440, 1136, 3547, 2439, 1944, 3041, 2911, 2923, 2478, 2670, 1429, 3558, 2588, 2853, 1052, 264, 2406, 2736, 2762, 2479, 3065, 2063, 374, 2316, 814, 2247, 3561, 3595, 3133, 2494, 2851, 3205, 3493, 3050, 2446, 3546, 2696, 3329, 2871, 1438, 829, 1251, 3019, 2220, 2847, 3508, 2398, 2877]. evolutive [2947]. evolvability [3480]. evolving [1986, 262, 1518, 3123, 1990]. Ewens [491, 1854]. Exact [2022, 2212, 3407, 3615, 1861, 2353, 2738, 1360, 1420, 1930, 3480, 1985, 3519, 2802, 2285, 1385, 483]. Exactly [2532]. examination [1013]. examined [1348]. examining [2306]. example [706, 357, 1888, 2662, 1355]. Examples [594, 1913, 3411, 867, 1400]. excessive [2522]. exchange [1556]. exchangeable [565, 1654, 1418]. Excitability [1211, 235, 792, 2182, 2111, 33]. Excitable [33, 88, 1225, 1985, 286, 2574, 667, 276, 677, 1201, 1170, 1149, 2935]. Excitation [196, 59, 677]. excitations [175]. excitatory [3063, 1005]. excited [114, 147]. Exciting [1320]. Exclusion [3203, 1408, 2228, 730, 1378, 2419, 1105, 1098, 855, 986, 2371, 3040, 2829, 2269]. exhibiting [1291, 3220]. exhibits [2918]. Existence [724, 503, 702, 2766, 1379, 337, 175, 1785, 407, 249, 2599, 2669, 1000, 447, 16, 1964, 1048, 265, 1925, 1342, 110, 3391, 1104, 927, 3279, 267, 20, 1990, 1557]. existence-uniqueness [3279]. exit [2641]. exotic [2817]. expanding [3403]. Expansion [1945, 2783, 3530, 3595, 3320, 794]. expansions [43]. expectancy [2039]. expectations [2974]. Expected [2059, 1237, 2588, 3536, 1732]. experiment [735]. Experimental [1, 2807, 2400, 1509, 1389]. experimentally [1540]. experimentation [1691]. experiments [2150, 256, 1893, 3154, 321, 3451, 294]. explain [636, 3206, 27]. explained [1888, 2662]. explaining [2868]. explanatory [1826]. explicit [1276, 3302, 2914, 107, 3307, 1443, 2221, 1038, 1268, 3577]. exploration [2698]. Exploring [3468, 2438, 3553, 3244, 3579, 2545]. Exponential [2854, 1416, 3573, 153, 2446]. exposed [1473]. exposure [1724]. exposures [2350]. expression [2409, 1044, 2022, 2056, 2640, 3083, 2810, 2695, 3464, 1671, 3192, 1678, 2157, 3423, 1771, 2462, 1886, 2782, 3246, 1258, 3472, 2250]. expressions [2974, 3480]. extant [1340]. extended [2094, 1815, 1238, 2116]. Extending [2185, 2403, 3558]. extension [2308, 2216]. extensions [1821]. extent [3129]. Exterior [1724]. External [704, 2931, 1173, 1339, 365, 349, 881, 1798, 2776, 1202]. externally [1971]. extinct [3099, 3193]. Extinction [2787, 1168, 1049, 1998, 3546, 621, 2904, 1316, 2959, 1269, 1923, 2176, 2231, 2673, 3415, 2931, 3044, 1907, 1274, 1874, 2860, 1429, 2636, 640, 3202, 2486, 3175, 873, 1374, 2487, 2759, 1945, 2819, 2996, 262, 3571, 3536, 699, 2788, 2644, 958, 68, 946, 2295, 2910, 2554, 1111]. extinction/fixation [2860]. extra [2529]. extracellular [2520, 1843, 1800, 44, 1801, 878, 892, 3321]. extracisternal [1233]. Extracting [2689]. Extrapolating [2847]. extremal [3047]. extreme [3068]. Eye [1545, 525, 2604, 2035].

fabric [1148]. Facets [2541]. Facilitated [297, 524, 181]. facilitates [3320]. facilitation [802]. facing [3107]. Factor [2195, 3116, 55, 1579, 1409, 3096]. factor-regulated [3096]. factorization [2017]. factors [3291, 2180, 417, 3262, 2054, 2006]. failed [3513]. Failure [1830, 1497, 2347, 1413]. Fair [2189, 2436]. families [1455, 1476, 1616]. family [3041, 595]. famine [602]. far [11]. far-from-equilibrium [11]. farmers [3145]. farms [2558]. Farris [3294]. Fast [3449, 1560, 3027, 2945, 2115, 1721, 2465, 3344, 1914, 1556, 2462, 3057, 1695, 541]. fast-or-slow [3344]. fast-slow [2462]. fat [3216]. fatal [788, 1133, 2841]. fatalities [3179]. fate [2617, 3537]. favor [1533]. favors [2128]. fear [2569]. Feasibility [3382]. feast [602]. feature [1991, 3536, 2077]. features [1700, 3397]. fecundity [686, 270]. Feedback [1376, 1709, 2264, 2655, 2387, 2888, 1486, 558, 2750, 2640, 3083, 2377, 1186, 1852, 337, 2234, 112, 1123, 3288, 2249, 3366, 1876, 20, 2537]. feedbacks [2439, 3254, 2776]. feeding [3146]. Female [53, 1097, 646, 1987, 1203]. femoral [827]. fencing [2368]. Fenichel [3319]. Fennoscandia [3479]. ferrooxidans [2293]. fertilities [1345]. fertility [346, 25, 538, 2127, 457, 2982, 1744]. fertilizing [1451]. fever [1098]. fibers [180, 3388]. fibres [590, 389]. Fibrillation [498]. fibrin [2480]. fibroblasts [1893]. fibroblasts/myofibroblasts [1893]. fibronectin [2686]. fibrous [2177, 2476]. fidelity [2014]. field [2665, 2902, 3358, 933, 2404, 2134, 496, 79, 2573, 3295, 2975, 3501, 2958, 2605, 2956, 2282, 2256, 2468, 2141, 2733, 1191, 2428, 2499, 3482, 864, 2119]. fields [796, 705, 1549, 175, 682, 1912]. fight [2430, 2725]. fighting [2430, 2725, 2453]. fighting-predators-prey [2453]. filament [1308, 1984, 2599]. filamentary [1531]. filaments [1825]. Filippov [2553, 3086, 2552, 3422]. filled [1389]. filling [3212, 1791]. film [1221]. films [389]. filopodia [1956]. filters [84]. Final [3616, 3441, 2176, 481, 583].

Finding [1072, 2859, 2973, 2379, 1707]. findings [2570]. fine [3537]. fine-tuning [3537]. Finetti [26]. fingerprints [2912]. Finite [2633, 3102, 3001, 1011, 2040, 1253, 1986, 3089, 2711, 2681, 741, 2843, 3560, 250, 1951, 280, 1321, 2394, 3195, 324, 410, 1580, 2730, 3205, 2354, 3279, 2871, 2072]. finite-element [2072]. Finite-size [2633]. finite-time [3195]. fire [1225, 2169, 2156, 2164, 2578, 2270, 1482, 1441, 2349, 2935, 2639]. firefly [832]. firing [2134, 2956, 1267]. First [1611, 747, 2641, 3060, 1877, 3068, 3475, 1722, 1850, 2785, 441]. first-order [1850, 2785]. fish [1788]. Fisher [1757, 2115, 2882, 2991, 2252, 2714, 537, 2684, 3048, 1348, 719, 3125, 3091, 2447, 1256, 1582, 830, 768, 1572, 3059, 384, 3546, 1000, 1155, 993]. fishery [939, 312, 529, 229, 311]. fishes [1759]. fishing [3371, 3613]. fission [450, 1550]. fit [2173, 2258]. Fitch [2820, 2949]. Fitness [3134, 1338, 1253, 2581, 2090, 2255, 3394, 2743, 3447, 2946, 1424, 1317, 2803, 929, 1423, 1518, 2498, 2821, 2559, 1550, 1191, 1912, 1058, 2544, 3324]. fitness-returns [2498]. fitnesses [1228, 768, 1042, 3583, 889]. fittest [1253]. Fitting [3550, 2422, 2372]. FitzHugh [1618, 12, 899, 336, 125, 367, 541, 47, 2935]. Fixation [3547, 1018, 2469, 2812, 3081, 3133, 3324, 2877, 99, 1715, 2714, 2860, 2313, 1654, 3583, 2982]. fixed [2218, 2601, 1596, 3179, 2085, 1433, 2161, 3209]. fixing [2517]. FKPP [3308]. FKPP-type [3308]. flagella [1586]. flagellar [3267]. flagellates [849]. flashing [3433, 2272]. flattening [3429]. flexible [2421, 2975]. flight [3349, 3220]. flights [2063]. flocking [3226]. Floquet [2456]. Flow [1997, 2253, 579, 431, 1044, 1221, 1166, 1575, 1464, 2236, 702, 780, 2161, 1337, 2958, 1568, 2213, 2550, 598, 717, 941, 2795, 1846, 2786, 2249, 2201, 2392, 2564, 1581, 2480, 795, 3289, 1783, 1749, 1661]. Flow-distributed [1997]. flows [1135, 1202]. flu [3577]. Fluctuating [3108, 1158, 3020, 523, 301, 3200, 1957, 3103, 461, 221, 277, 213, 2319, 1936, 979, 2396]. Fluctuating-rate [3108]. fluctuation [1880, 2811, 1197, 1255]. Fluctuations [985, 1237, 939, 3177, 2812, 603, 3305, 1132, 1984, 2984, 1692]. fluid [2614, 2138, 1839, 1389, 2795, 2249, 2130, 2195, 519, 3289]. fluid-filled [1389]. fluid-mechanical [519]. fluktuierender [523]. fluorescence [1779, 3198]. fluttering [3309]. Flux [2315, 2591, 1343, 2243, 2427, 116, 2973, 2892, 1258, 3533, 2562, 2702]. fly [2946]. Focal [981]. focus [3345]. Foerster [3359]. Fokker [2549, 1977, 3288, 631]. Folding [2589, 1702, 2747, 2461, 2334, 2343]. foldings [2477]. foliage [111, 149]. follicle [3131]. follicles [2861]. Follicular [2861]. follow [2414]. following [2412, 962, 2884, 3536, 985, 792]. food [1196, 1968, 1168, 365, 1321, 2757, 2778, 2727, 545, 966, 1122, 1469, 1365, 2546, 880, 2329, 3360, 341, 1889, 1866, 3541]. food-chain [1168, 2546]. food-limited [1321]. food-web [1968]. foods [1461, 1462]. footing [2404]. footprinting [1560]. foragers [2318, 3238, 962]. foraging [1215, 3349, 1010, 3070, 735]. force [938, 1091, 1343, 3008, 2382]. Forced [302, 2707, 2914, 2952, 518, 1201, 1934, 2026, 531, 1099]. forces [3260, 3386, 2792]. forcing [365, 2984, 984, 2209]. foreign [1768]. forest [3250, 956, 2349]. forestry [1125]. Foreword [2145]. form [3058, 803, 478, 2446, 2769, 260]. Formal [1741]. formalism [2881]. Formation [872, 1523, 1032, 1986, 1818, 2549, 355, 2133, 3594, 2425, 2202, 1285, 908, 2412, 3004, 1001, 1101, 2527, 45, 964, 2376, 2194, 803, 1810, 2243, 102, 79, 93, 1975, 713, 940, 2013, 152, 3011, 3487, 1634, 2629, 468, 3157, 861, 1407, 184, 677, 1873, 161, 3286, 1956, 1278, 893, 506, 287, 1480, 1142, 2570, 1280, 2011, 2913, 1312, 2460, 994, 2233, 1038, 1504, 2512, 1028, 2480, 3096, 1964, 3434, 1290, 3016]. formations [2663]. forming [3123]. forms [2516, 662]. formula [3249, 107, 491, 1377, 1854]. formulas [2212, 952, 1360, 1420]. Formulation [294, 2628, 1875, 1988, 1138, 1301, 2473, 110, 2958]. formulations [120]. Forward [2012, 3462, 3528, 2064, 1708]. Forward-backward [3462]. Foundations [2289, 3369, 1687]. founder [369, 298]. founders [3078]. four [504, 3285, 108, 3104, 1112, 207]. four-node [3285]. four-state [504]. Fourier [1124]. FR3D [1707]. Fractal [1505, 1579, 1409]. fraction [2407, 1477, 3514]. Fractional [1842, 1264, 3458, 2754]. fractionated [2407]. fragmentation [2590, 2513, 2149, 2989, 1151, 1668]. fragmented [1517, 2775, 1640]. fragments [2686, 2592, 2008]. Frame [1534]. framework [2337, 3264, 2777, 1453, 1010, 2017, 1883, 3578, 3570, 2521, 605, 3288, 3585, 2369, 2555, 2901, 1100, 2639]. Fred [3520]. Free [2737, 2663, 1273, 2951, 1674, 2071, 3322, 3039, 2321, 3362, 2939, 983, 1041, 3374, 3217, 2260, 1645, 2325, 2699, 3404, 3430, 1033, 3106, 2334, 3410, 2235, 2635, 2932, 1909, 1860]. free-energy [2235]. French [419, 2556, 2656, 3027]. frequencies [535, 497, 415, 3483, 384]. Frequency [250, 2781, 2494, 1397, 470, 2640, 2497, 657, 2252, 1427, 2714, 1605, 890, 1061, 1425, 2294, 107, 3256, 230, 324, 1293, 1377, 1365, 2284, 277, 3244, 3583, 205, 1992, 1566, 1881, 109, 1844, 1207]. Frequency-dependent [2781, 2497, 2252, 2714, 1061, 1425, 1293, 205, 1566, 1881, 1207]. frequency-independent [3583]. frequency-locking [890]. frequent [1992]. Freter [1407]. friction [3252]. Friedman [3418]. Frobenius [1341]. Front [3457, 1256, 3157, 2758, 1108]. fronts [1361, 3004, 41, 3355, 927, 1647]. Frost [583]. fruit [819]. FTC [802]. FtsZ [2244]. Fuji [1253]. Fuji-type [1253]. full [550]. fully [114, 147, 1387]. Function [1547, 2166, 894, 1688, 2134, 1644, 2613, 1645, 754, 11, 42, 3174, 2702]. Function-valued [1547, 2166, 1688]. functional [740, 868, 3072, 1585, 1129, 2113, 2661, 514, 1564, 162, 1981, 1878, 617, 550, 551, 2691, 3613, 1743, 2769, 3579, 16, 3621]. functional-integral [740]. functionalized [2669]. functionals [3349, 2685, 2446]. functions [2987, 1029, 1215, 2314, 1208, 2094, 2751, 520, 3384, 1785, 929, 2691, 2397, 1894, 183, 1660, 2544, 1751]. fundamental [2761, 924]. fundamentals [3619]. fungal [1101, 3601, 3191, 1200]. funnels [2924]. Further [369]. fusion [2461, 1550]. futile [2618, 1735]. fuzzy [8].

G [2266, 1102, 1500, 1860]. G-CSF [1500]. G-protein [1102, 1860]. G-Quadruplex [2266]. G2 [2836]. G2-checkpoint [2836]. Gain [1160, 3174]. Gain-induced [1160]. gaits [1281, 1282]. Galerkin [410]. Galton [1360, 1420]. game [2180, 1833, 3373, 2856, 1903, 1572, 2450, 3371, 1656, 482, 3050, 3136, 2696, 1251, 2220, 2007, 3149]. game-theoretic [3373]. game-theoretical [3149]. Games [2143, 502, 2180, 2333, 2452, 2923, 571, 2099, 2588, 2853, 1394, 1951, 3368, 2771, 1059, 2793, 668, 3343, 430, 2466, 1043, 1199, 3311, 2611, 1025, 482, 944, 829, 3019, 2847, 652, 2877]. gametic [627]. gametophytic [970, 1019]. gamma [2137]. gap [3005, 1511, 2587]. gap-junction [1511]. gapped [2294]. gas [1983, 360, 1954, 1089, 1405]. gases [655]. gastric [3211]. GATA [2005]. gated [1192]. gatherers [3145]. Gause [1029, 1845, 789]. Gause-type [1029, 789]. Gauss [2641]. Gaussian [2255, 2681, 389, 3008, 2939, 3356, 492]. gel [3298]. Gene [656, 2640, 2907, 1593, 1536, 1886, 2253, 1942, 2409, 875, 943, 1044, 2022, 2056, 3083, 2438, 3041, 3025, 2810, 369, 3330, 2695, 2848, 2957, 3135, 1929, 317, 474, 3464, 1671, 2820, 3045, 3254, 2976, 298, 873, 1678, 2331, 3513, 3097, 2157, 3128, 324, 3108, 277, 522, 2913, 3263, 2290, 2462, 2783, 2521, 3196, 639, 2886, 2428, 2006, 2974, 109, 2389, 1885, 2782, 3246, 3472, 3398, 3150, 2250]. Gene-culture [656]. gene-type [3097]. gene/species [2438]. Genealogical [758, 3160, 1857, 654, 1024, 1036, 804, 898, 897, 638, 1386]. Genealogical-tree [758]. genealogies [2238, 1680, 1272]. Genealogy [1939, 1174, 651, 1950]. General [751, 2785, 3424, 1534, 1709, 3612, 1960, 1996, 2882, 2991, 1208, 3389, 2058, 1138, 1301, 2153, 317, 2751, 3081, 702, 333, 1347, 3175, 1584, 526, 1949, 1377, 2917, 111, 2254, 3578, 2330, 384, 3518, 2397, 2369, 1266, 3057, 2418, 2769, 2163, 1461, 2901, 1685, 799, 189, 3442]. generalised [3293, 1763]. Generalist [1128, 2192, 3310, 2672, 3479]. Generalization [2415, 579, 2188]. Generalizations [2975]. Generalized [46, 532, 1230, 2871, 2077, 1029, 608, 707, 715, 3025, 807, 309, 1223, 964, 2856, 1180, 2482, 1623, 1574, 267, 703, 492, 1864]. Generalizing [2857]. generate [2531]. generated [2937, 1343, 3215, 1474, 159]. generating [2987, 976, 3461, 478, 2427]. generation [490, 1294, 1492, 1826, 345, 1367, 1640, 4, 1480, 2940, 2125, 2172, 2543, 1084]. generations [1345, 62, 1504]. generator [342, 1741]. generators [1281, 1282, 1552, 1601]. generic [2483, 2064]. genes [1311, 2717, 271, 2849, 430, 630, 898, 1994]. genesis [1070]. Genetic [1548, 232, 253, 26, 501, 1054, 339, 3274, 2881, 706, 2387, 2796, 3419, 466, 2515, 3325, 428, 604, 955, 556, 2854, 82, 36, 304, 1567, 3294, 1147, 1348, 1296, 3304, 896, 959, 62, 107, 154, 21, 452, 69, 144, 337, 745, 1921, 2082, 2803, 488, 546, 2251, 2598, 2539, 1916, 1536, 594, 658, 2901, 1058, 14, 512, 1855]. Genetical [1157]. genetically [433, 258]. genetics [724, 40, 357, 2058, 565, 2684, 38, 1053, 637, 39, 377, 836, 3141, 1858, 1835, 3511, 148, 258, 375, 876, 1970, 889, 2763, 680, 2173, 2258, 1054]. genetics-epidemiology [3511]. genic [415]. genome [1455, 3517, 271, 170, 2332, 3060, 1910, 1205, 3314]. genomes [2287, 1616]. genomic [3292, 2920, 1744]. genomics [1585, 2081]. genotype [2479, 285, 597]. genotype-dependent [285]. genotype-phenotype [2479]. genotypes [156]. Genotypic [1003, 346, 333, 497, 1017]. genus [2193, 55]. Geographical [1249, 138, 3303, 3348, 501]. geographically [157, 220, 804, 898, 897, 1131]. Geometric [1980, 1818, 1562, 1856, 2506, 2925, 2626, 3173, 1812, 3228, 2462, 2669, 289]. geometrical [2994]. Geometrically [1534]. geometries [2540]. Geometry [1822, 3447, 183, 3199, 3089, 3402, 1988, 45, 3138, 666, 2803, 10, 3533]. German [523, 61]. germinal [2625]. germination [2835]. Getting [2792]. giant [2346]. Gierer [152]. Gillespie [2747]. Gilpin [96]. Ginglymostoma [131]. give [386]. given [271, 2623, 3559, 1340]. gives [3246]. glass [1863]. glide [1600]. glioblastoma [3300, 3581, 2505, 3129, 2948]. Glioma [2414, 3474, 3575, 1729, 3168]. gliomas [2473, 3207]. Global [1946, 3581, 1537, 715, 1004, 886, 2687, 346, 2638, 774, 3036, 2423, 3077, 1563, 2278, 3267, 2751, 71, 472, 3549, 193, 141, 127, 1162, 983, 1289, 1952, 2753, 395, 2400, 3617, 809, 518, 2260, 3257, 789, 1139, 3337, 1682, 1981, 3609, 679, 3465, 3310, 945, 1033, 997, 1663, 224, 3132, 2900, 2386, 2618, 2397, 620, 3280, 1075, 2749, 539, 1990, 392, 2865, 1185, 1305, 2066, 103, 3468, 1960, 2402, 604, 2226, 2701, 300, 338, 93, 3615, 1206, 1181, 383, 3021, 1825, 1333, 1127, 2132, 3397, 3315]. Globally [1596, 464, 1307, 447, 970, 267, 3187]. gluconeogenesis [278]. Glucose [1176, 1224, 3216, 3556]. Glucose-induced [1176]. glucose-insulin [3556]. glycans [2304]. glycemia [2526]. glycemia-structured [2526]. glycocalyx [1843]. glycolysis [3120]. glycolytic [1127]. GO [3615]. goals [2511]. Goldbeter [2880, 2506, 1908]. Golden [2842]. Gompertz [195]. goods [1951, 2611]. Goodwin [840]. Gordon [485]. got [1781]. Gouy [46]. governance [3557]. governed [79, 663, 445]. governing [142, 1416, 1730, 3505]. Gower [2634]. grade [3474, 2473]. gradient [579, 1136, 2255, 1275, 1549, 953, 2958, 1954, 2384]. gradient-based [1954]. gradient-sensing [2384]. Gradients [668, 409, 2347, 2498, 3197, 2562]. gradostat [629, 728, 847, 580]. gradual [2255]. gradually [469]. Grafen [2333]. grained [2683, 2387]. graining [3008, 2077]. gram [3150]. gram-negative [3150]. grammar [2334]. granulocyte [29, 559]. granulocytopoiesis [239]. Granulopoiesis [1500, 1406, 1479]. Graph [1742, 1900, 1665, 1666, 2324, 3325, 2052, 3558, 2623, 1910, 1442, 2921, 1930, 1703, 1828, 2877]. graph-automorphism [1930]. Graph-theoretic [1900, 1665, 1666, 1828]. Graphical [2600, 1917]. Graphs [3139, 3532, 3171, 3048, 2919, 2908, 3045, 3023, 3370, 2081, 3186, 2095, 3166, 3431, 1930]. grass [2639]. gravitactic [2136]. gravitropism [1630]. gray [3498]. gray-box [3498]. grazer [3241, 1981]. grazing [755, 120]. Greater [2715, 2128]. green [122]. Gregarious [1569]. Grey [1456]. Grey-box [1456]. grid [3358]. GRO [2595]. GRO-seq [2595]. Group [2287, 2113, 1550, 3213, 2777, 1957, 3328, 1445, 1857, 1810, 2751, 1814, 2534, 3184, 1821, 1835, 3441, 3364, 2833]. group-based [3213]. group-structured [3328, 1835]. Group-theoretic [2287]. grouping [2529, 1561]. groups [3379, 1933, 1899, 2871]. grouse [1436]. growing [2869, 1795, 3005, 1324, 1882, 2054, 2548, 631, 1571]. grows [1442]. Growth [450, 3515, 1531, 490, 1096, 2306, 1450, 817, 1362, 523, 706, 1649, 2673, 1013, 853, 417, 2585, 1354, 634, 240, 346, 1198, 1009, 2590, 915, 1285, 3293, 1302, 1275, 530, 2138, 150, 174, 309, 129, 316, 1384, 1797, 1339, 94, 140, 1571, 45, 3332, 1530, 3504, 831, 2177, 3135, 2149, 3116, 1636, 2102, 3440, 1422, 156, 1184, 1877, 167, 237, 1918, 437, 3085, 800, 64, 554, 1130, 751, 2200, 2732, 1865, 1584, 1729, 873, 2508, 2712, 1329, 456, 1240, 3340, 745, 662, 3391, 3308, 821, 1568, 2240, 2781, 54, 1785, 2062, 3297, 2915, 3404, 1793, 383, 754, 505]. growth [2719, 2217, 3121, 2612, 2608, 2191, 3589, 984, 2788, 1973, 1628, 411, 153, 1801, 2326, 1200, 3098, 1521, 1761, 236, 1306, 234, 1070, 1115, 51, 1062, 1099, 113, 3219, 2948, 580, 892, 3482, 2305, 2961, 2093, 499, 572, 3433, 749, 1687, 1414, 16, 63, 2910, 3599, 3096, 1371, 1383, 1909, 1373, 2517, 1502, 372, 2876, 2746, 3210, 3387, 2824, 3321, 1152]. growth-fragmentation [2149]. growth-induced [1302]. growth-related [1521]. guidance [1258]. guide [3359, 1778]. guided [3486]. Gyllenberg [2485, 2215]. gynodioecy [1410]. gyrotactic [1179].

H [3396]. H7 [2295]. Habitat [2989, 2318, 1173, 2775, 2076, 1132, 2712, 2996, 3606, 2839, 1330, 3278, 1668, 1577, 3280, 1371, 1383, 2167]. habitats [3590, 1274, 2839, 2675]. Hadeler [2679, 2825, 1604]. haemodynamics [2363]. hair [1233]. Haldane [3249]. Half [906, 2646]. half-center [2646]. Half-sib [906]. Hamiltonian [1059, 1213, 2783]. hammerhead [1359]. Hamming [747]. hamper [3254]. handling [551]. Hans [3002]. haploid [1088, 1228, 1092, 1277]. haplophase [67]. haplotype [3483]. haptotaxis [922, 750]. hard [1072]. Hardy [978, 895, 178, 809]. hares [1436]. harmful [2806, 3528]. harmonic [65, 457, 305]. harmony [785]. Harris [2940]. harvest [1463]. harvested [176, 3613]. Harvesting [686, 2936, 3075, 1510, 1158, 1234, 186, 299, 1752, 661, 327, 2852, 3305, 2977, 1017, 2546, 2073, 3122, 1304, 805, 866, 3500, 599, 236, 1446, 3541, 3621]. Hastings [2992]. having [506, 1255]. hazard [1457]. HBV [1982]. head [1727, 2950]. head-up [2950]. healing [1363, 1382, 822, 928, 2072, 3451]. health [2879, 2794, 1031]. healthy [1737]. heart [1497, 444, 2148]. heartbeat [1579]. heat [459, 1804]. hedging [3162]. Heisenberg [1814]. Hele [2719]. Helfrich [2235]. Helical [2136, 2336, 1586]. helix [3454]. Helly [2132]. helminth [761, 1012]. help [2144, 2839]. hematological [1773]. hematopoiesis [3177]. hematopoietic [1946, 1723]. hemodialysis [3494, 2990]. hemolytic [45, 135]. hepatitis [2122, 2536, 1865, 3337]. hepatocyte [1865]. hepatocytes [1805]. her [568]. herbivore [2986, 586, 1325, 3350, 2057, 3579]. herbivore-plant-pollinator [1325]. herbivores [1010]. Herd [3318, 3566, 3253, 755, 761]. hereditary [171, 2049, 247, 195]. hermaphrodite [2450]. hermaphrodites [439]. hermaphroditic [2381]. Herzel [1738]. hESCs [3235]. Heteroclinic [3415, 1512, 1196, 2491, 1743, 2023]. heterogeneities [2181, 1640]. heterogeneity [3445, 818, 1156, 2467, 3414, 3190, 2735, 263, 2111, 3428, 3423, 1676, 2837, 3143, 3333, 3386, 1519, 2792, 3248]. Heterogeneous [2673, 1554, 931, 3291, 3551, 2893, 901, 2549, 1674, 3155, 2142, 2911, 2790, 590, 3007, 3277, 784, 3253, 2102, 2366, 2268, 2533, 3085, 3202, 2732, 3438, 1315, 2050, 2126, 2942, 2047, 3204, 843, 3367, 718, 2754, 2378, 2539, 1504, 3232, 2512, 3497, 1705, 2630, 3153, 1710, 1625, 2774, 3417]. heterozygosity [3112]. heterozygous [310]. hexagonal [2011]. hexagonal-cell [2011]. hidden [1831]. hierarchic [2785]. Hierarchical [1063, 3474, 982, 3512]. hierarchically [1119]. hierarchies [2178]. hierarchy [916, 1950, 2710, 2285]. High [2128, 1605, 1968, 1359, 398, 1365, 1142, 3185, 1406, 3388, 909, 27, 3433]. high-dimensional [1968, 3185]. high-frequency [1365]. high/low [3433]. high/low-flashing [3433]. Higher [2385, 1720, 3058, 222, 1596, 1064, 3413]. higher-molecular [3413]. higher-order [1720]. highly [1476, 2560]. hiking [1272]. Hill [2577]. Hilliard [2669]. hillslope [2531]. hindrance [1683]. Hindrances [2758]. histogram [1838]. historical [2801]. histories [3041, 2848, 3558, 2498, 2095, 1031]. History [1522, 3304, 3384, 3622, 1438, 1386]. hitch [1272]. hitch-hiking [1272]. hitchhiking [578, 606, 1121, 1941]. hitting [2811]. HIV [3073, 777, 1758, 1507, 2687, 2412, 3605, 1393, 1454, 3084, 3170, 3459, 1297, 3079, 1110, 2187, 1448, 1642, 2082, 843, 2745, 1652, 2490, 2768, 2560, 1836, 957, 3420, 2007]. HIV-1 [3073, 1393, 2187, 2560, 2007]. HIV/AIDS [777, 843, 1652]. HIV/CTL [2687]. Hive [3199]. Hodgkin [772, 2902, 1553]. hole [3594]. Holliday [2710]. Holling [514, 1981, 550, 551, 3613, 3579]. home [1739, 3580]. Homeostasis [3383, 2607, 3034, 3285, 3188, 3290, 3387]. Homeostatic [2627]. homing [1549]. Homoclinic [1196, 1087, 1334]. homogeneity [2515, 748]. homogeneous [143, 1632, 1254, 3228, 175, 2360, 2705, 2680]. Homogenization [2291, 2774, 2184]. homology [2]. Homophilic [2664]. homosexually [1713]. homosexually-transmitted [1713]. Homozygosity [553, 1992]. honey [3502, 861, 3070]. honeybee [3199]. honeycomb [610]. honor [2992, 3418]. honour [2109, 3002]. Hopf [274, 1349, 2423, 352, 1094, 734, 840, 1299, 3528, 292, 125, 2749, 3621]. horizontal [3135, 3196]. hormone [2305]. horseshoe [2658]. hospital [3573]. hospitals [2701]. Host [2668, 1544, 2823, 1574, 1503, 3405, 2805, 3468, 901, 465, 2128, 1186, 2638, 1731, 2203, 1955, 3135, 2841, 1098, 2352, 1172, 3054, 510, 1232, 1374, 2887, 2187, 3609, 433, 2745, 3561, 3595, 974, 863, 1676, 2768, 348, 958, 1504, 2093, 2554, 2046, 3331, 3427]. host-one [2668]. host-parasite [901, 465, 1232]. Host-parasitoid [1574, 1186, 974, 1676, 1504]. Host-Pathogen [1544, 1503, 2203, 3331]. host-symbiont [3595]. hosts [2004, 1767, 2385]. Household [1959, 2457, 2726, 3566, 1907, 2089, 1973]. households [2457, 2500, 1973]. HP [2658]. HTLV [2045]. HTLV-I [2045]. Human [1458, 3503, 1500, 648, 1415, 1473, 182, 1246, 1247, 29, 3032, 972, 1738, 2692, 2767, 1182, 3303, 3229, 761, 2345, 435, 611, 766, 2390, 1479, 783, 941, 2967, 3279, 3568, 3173, 3376, 3064, 1894, 1268, 3299, 2046]. humans [1497, 2412, 2700, 3471, 1627]. hump [3591, 986]. hump-shaped [3591]. hungry [514, 550, 551]. hunter [3145]. hunter-gatherers [3145]. Huxley [772, 2902, 1553]. Hybrid [3142, 1624, 2671, 3476, 1985, 1003, 2401, 3560, 3159, 3346, 3128, 3263]. Hybridisation [1522]. hydra [2073, 2001]. hydraulic [2691, 1233]. Hydrodynamic [1542, 2542]. hydrodynamics [2628]. hydrogen [2306]. hydrolysis [265]. hydroxypropionic [2779]. hyper [2836]. hyper-radiosensitivity [2836]. Hyperbolic [2041, 2425, 2741, 1810, 1899, 1493, 3067]. hyperbolicity [2658]. hypercycle [613, 89]. hypercycles [732]. Hyperelastic [2476, 3403]. hyperexcitability [3477]. hyperinsulinemic [1617]. hypertension [3443]. hypnozoite [3578]. hypothalamic [1972]. hypothalamic-pituitary-adrenal [1972]. hypothesis [3349, 60, 837, 1756, 2502]. hypoxia [1694]. Hysteresis [161, 659, 2627, 3528, 2290]. hysteresis-type [2290].

icosahedral [2019, 1821]. Ideal [3322, 2071, 2321, 2260, 1645, 2325]. Idealized [3477, 2570, 3509]. identical [2937]. Identifiability [3292, 1166, 2175, 3316, 3397, 3395, 2211, 3026, 3562, 2215, 1476, 3270, 2088]. Identification [1852, 2275, 3576, 908, 2408, 517, 2516, 1838, 2826, 2174, 1937, 2802, 164, 2941, 278, 104, 3196]. Identifying [1942, 2856, 2017, 2826, 3486]. identities [2706]. identity [253]. IgG [255]. IgM [255]. II [1253, 3551, 2655, 943, 1247, 1650, 1282, 1807, 390, 1943, 1988, 1301, 2507, 2622, 3180, 1412, 441, 2566, 324, 2361, 2595, 424, 551, 1666, 1881, 2195, 305, 3579, 1462, 49]. III [319, 174, 514, 255]. ill [1436]. illness [2933]. illumination [290]. illustrating [1390]. IM [2263]. image [849, 1729]. image-driven [1729]. images [2760]. imaging [1677]. imitation [2965, 3149]. immersed [2062, 2213]. immigration [2951, 1734, 2388, 1949, 638, 3598]. immobilized [161]. immobilizer [1028]. Immune [998, 2458, 2832, 3394, 1454, 3422, 2838, 320, 3191, 2352, 3550, 2887, 1161, 2010, 2690, 318, 424, 1526, 3098, 2229, 1732, 1516, 2048, 2179, 3420, 1982]. immune-mediated [3098]. immunity [1503, 1113, 2004, 3566, 2458, 733, 2580, 1898, 3253, 3544, 3525, 1535, 331, 761, 3452, 3258, 3318, 3448, 1841, 2554, 2307, 988]. immunization [3488]. Immuno [2458, 1398, 3406, 3054, 2350, 3554]. immuno-epidemiological [3406, 3054, 2350, 3554]. Immuno-epidemiology [2458]. immuno-selection [1398]. immunocompetent [3489]. immunodeficiency [743]. immunological [403, 3471]. immunology [73, 118, 255]. immunosuppressive [2229]. immunotherapy [3095, 2557, 1161]. Impact [2860, 3601, 3470, 3545, 3390, 3334, 3326, 3566, 2039, 3089, 1274, 3624, 1817, 3303, 1509, 2345, 3513, 3573, 3258, 2705, 2794, 2829, 2749, 3333, 1994, 3289, 2808, 3488]. Impacts [3502, 3250, 3542, 1571]. impedance [1690]. Imperfect [3500, 2247, 2720]. implant [2177, 2098]. implants [1761]. Implications [3438, 3098, 2349, 2440, 2219, 3238, 3605, 1837, 2390]. Implicit [3302, 2393, 1038, 3577]. Implicit-explicit [1038]. imply [2821]. importance [2419, 3466]. important [2823]. impossibility [1491]. impracticalities [3090]. imprecise [2545]. imprinting [1744]. improve [1662]. Improved [1195, 1640]. improvements [2609]. improving [1216]. impulse [1852]. impulses [3030]. Impulsive [1911, 2766, 2061, 1496, 2755, 2639]. In-111 [2760]. in-host [3405]. in-phase [3057]. in/on [748]. inactivation [169]. inbred [3078]. Inbreeding [260, 548, 706, 707, 678, 683, 778, 1377, 228]. Incentives [3296, 3560]. incidence [1351, 2741, 2751, 1535, 721, 816, 1952, 2124, 569, 650, 1869, 3518, 3339, 3412, 1693, 988]. includes [1393]. including [1497, 1985, 2104, 2857]. inclusions [1350, 2947]. Inclusive [1058, 1550]. incompatibilities [3274]. incompatibility [2904, 1996, 900, 970, 1019]. incompatible [602]. incompetence [2793]. incomplete [1833]. incompressible [3332]. Inconsistencies [764]. Incorporating [3241, 2054, 1791, 2726, 841, 915, 2552, 2553, 3414, 2703, 320, 1321, 2317, 3370, 2975, 3298, 2179]. Incorrect [3139]. increase [657, 704, 3511, 2471]. increased [1316, 2039]. increases [1664, 3623]. increasing [343]. incremental [2885]. incubation [894, 2780, 743, 1932, 3315]. indefinite [2754]. independence [1756]. independent [3576, 1645, 3006, 3583, 2310]. independently [729]. indeterminacy [2738]. index [3236, 2954, 3061, 2436, 2189, 2313, 2493]. indicator [2163]. indicators [3224, 2626]. indices [3611, 2212, 3385, 3262, 3017]. Indirect [3196, 2447, 2867]. Indiscriminate [508]. Individual [2420, 1875, 1795, 2839, 3220, 2799, 1009, 1673, 3005, 962, 3359, 714, 999, 2855, 2455, 2947, 2110, 2121, 1419, 3318, 2583, 1692, 1748]. Individual-based [1875, 1795, 1673, 3005, 3359, 2947, 1692]. individuals [3291, 3347, 2422, 3204, 1487, 3450]. induce [1725]. Induced [682, 1160, 2228, 3212, 3566, 1772, 1302, 3592, 3591, 146, 1176, 3530, 3063, 1766, 1531, 2592, 3317, 1240, 3340, 328, 2690, 3122, 2672, 1466, 2666, 1676, 44, 3396, 3403, 3298, 3271, 3288, 1521, 2172, 1566, 2229, 2442, 3623, 1502, 3220]. Inducing [2290]. industrial [2558]. inequality [2545]. inevitable [2486]. infected [1023, 1589, 2552, 2553, 3170, 2385, 3608, 2560, 1836, 3080]. Infection [3054, 3623, 3467, 1959, 2457, 2673, 2770, 2202, 1208, 1898, 2181, 3605, 1260, 138, 543, 2199, 1824, 2674, 1759, 3191, 1440, 3303, 756, 751, 1865, 3452, 2045, 404, 3258, 2395, 1292, 1291, 3261, 2745, 2455, 1392, 1869, 2490, 863, 1141, 1693, 2560, 2229, 1870, 1967, 3535, 3420, 2554, 2046, 2571, 1982, 1430]. infection- [1260]. infection-derived [3258]. Infection-induced [3623]. infections [3467, 2323, 3478, 2412, 2790, 2335, 3181, 2085, 1012, 761, 753, 731, 1868, 3609, 1141, 1465, 2143, 1613, 2186, 3460, 1727, 3524]. Infectious [1591, 3584, 3167, 2805, 248, 2311, 3222, 784, 831, 3353, 2841, 3432, 3521, 214, 3438, 2345, 374, 587, 2616, 917, 882, 2178, 1313, 2386, 2583, 2652, 3390, 2185, 2403, 2749, 2246, 1481, 2630, 2808, 3563, 1111, 1081]. infective [2085]. infectives [2951, 3500]. infectivity [3574, 1915, 1347, 3470, 3465, 2501]. infector [2246]. infector-probabilities [2246]. INFER [3508]. Inference [3356, 641, 806, 3472, 1670, 3338, 2, 2359, 2764, 3449, 2175, 2572, 2706]. inferences [1916]. inferior [2055]. Inferring [2762, 3435, 2521, 946, 2543, 1386, 3359, 3205]. infestations [1189]. Infinite [460, 1136, 1437, 2681, 3330, 62, 439, 23, 745, 1842, 1518, 3205, 605, 1854]. infinite-dimensional [1136, 745]. infinitely [1753, 3523, 651, 758, 1491, 2586]. infinitely-many [651]. infinitely-many-site [758]. Infinitesimal [3034, 3285, 3188, 3290]. infinity [1127]. inflammation [626, 760, 404]. inflammatory [2276]. Inflection [2182]. Influence [3107, 2441, 2435, 569, 3514, 729, 2115, 1302, 45, 2520, 1296, 156, 2524, 3378, 364, 3262, 1927, 1804, 3610, 1203, 2007]. influences [2053, 878]. influenza [3326, 1113, 1398, 1939, 3086, 2552, 2580, 2941, 2158, 2862, 2823, 2553]. Information [2150, 204, 211, 3138, 1494, 2789, 2791, 1833, 2483, 2849, 1564, 645]. informed [3453]. inheritance [1346, 520]. inherited [575]. inhibiting [3320]. inhibition [475, 3087, 3490]. inhibitions [1033]. inhibitor [1273, 102, 1041]. inhibitor/immobilizer [1028]. inhibitors [1339]. inhomogeneity [501]. inhomogeneous [2815, 1670, 1006, 2361, 160, 19]. initial [2987, 1923, 592, 552, 3044, 657, 716, 704, 2570, 443, 2093]. Initial/ [2093]. initiating [1278]. initiation [2507, 3477]. initio [1703]. injecting [2122]. injection [2122]. injectivity [3383]. inorganic [2712, 3210]. Input [210, 796, 2631, 3034, 3285, 1490, 3188, 839]. Input-output [210, 3034, 3285, 3188]. input/output [796]. inputs [2040]. insect [552, 2067, 3161, 397, 1132, 1204, 3570, 744]. insecticidal [3076]. insecticide [3511]. insecticides [2227]. insects [3003, 448, 3309, 628, 2488]. insertion [2090]. insight [1953, 400]. insights [3211, 2402, 2819, 3122, 3595, 3443]. inspired [1511]. instabilities [1545, 2399, 2079, 2101, 1266, 1089, 3442, 3424, 819]. Instability [2888, 325, 2615, 972, 3336, 449, 1114, 2690, 344, 2362, 862, 191]. instantaneous [919, 3368, 925]. institutional [3560]. instructing [80]. insulin [3144, 3556]. insulin-degrading [3556]. insurance [2493]. intact [278]. Integral [2373, 214, 2987, 740, 868, 240, 489, 352, 1636, 1785, 1226, 407, 1230, 83, 612, 3245]. Integrals [254, 1825]. Integrate [2169, 1225, 2156, 2164, 2578, 1482, 1441, 2935]. integrate-and-fire [1225, 2156, 2164, 2578, 1482, 1441, 2935]. integrated [1496, 2087]. integration [3234, 209, 2080, 2338]. integrative [1733]. Integrator [210]. integro [3265, 2772, 3375, 3606, 2378, 1751]. integro-difference [2772, 3375, 3606, 1751]. integro-differential [3265, 2378]. Integrodifference [2357, 2827, 2298, 2299, 3493, 1326, 2135, 3322, 3101, 3540, 2903, 1719, 2999]. integrodifferential [965]. intensities [24]. inter [2535, 2791, 1552]. inter-network [1552]. inter-taxa [2535, 2791]. interacting [3274, 3401, 462, 560, 904, 1156, 1439, 622, 1270, 701, 95, 3372, 2936, 526, 2845, 3449, 1487, 2319, 1071, 98, 806, 2209]. Interaction [556, 1209, 1369, 1610, 2577, 2986, 1221, 947, 3394, 690, 1129, 3004, 1001, 798, 2052, 2005, 2912, 93, 3384, 1232, 3110, 1893, 1161, 303, 2009, 2137, 172, 215, 1526, 3614, 2693, 1801, 1825, 432, 1994, 499, 819, 3331]. Interactions [3072, 159, 1336, 1532, 3485, 465, 1350, 1354, 2452, 1513, 1650, 36, 2668, 171, 1933, 74, 2520, 3105, 3176, 812, 722, 2297, 71, 3453, 193, 623, 2844, 1401, 1965, 1005, 2803, 2057, 1419, 1137, 3579, 1277, 799, 1783, 2569, 2639, 3016]. interactive [609, 3427]. intercellular [1449]. interchange [3127, 2482]. Interdependency [2285]. interface [2683, 2200, 2392, 1712]. interfacial [905]. interference [3469, 3326, 2594, 202, 295, 108, 1064, 1142, 2405, 513]. interferential [3393]. interlocking [467]. interlocus [578, 606]. intermediate [1533, 824, 825]. intermediate-virulence [1533]. Intermediates [2624]. Intermingled [1471]. intermingling [1926]. intermittent [2112, 2305]. Internal [2612, 2608, 2931, 2363, 2588, 2853, 281, 2712, 1682, 2309, 2776, 3386]. Internal-Stores [2612]. internalization [3514]. Internodons [1036]. interplay [3192, 2191]. interpret [2400]. interpretation [1234, 2060, 1483, 2595, 3479]. interpreting [389]. interspecies [1667]. Interspecific [1172]. interstitial [1846]. interval [1579]. intervality [1968]. intervention [2879, 1208, 3353, 3512, 3195]. interventions [3224, 2345, 3287, 3602]. intestinal [2139]. intestine [795]. intimidatory [3282]. intra [1063]. intra-specific [1063]. Intracellular [2286, 1525, 2463, 2029, 44, 3271, 878, 2918, 3209, 3220]. intrachromosomal [522]. intracranial [1367, 1391]. Intraguild [2911, 3317, 3416, 2192, 2442]. intrahost [2082]. intrarenal [3443]. Intraspecific [432, 1499, 2410, 1566]. intratumoral [2837]. intravenous [1224]. intrinsic [316, 3085]. Introducing [1850, 3230]. Introduction [203, 1697, 3418, 2943, 2750]. introgression [2389]. intuitive [3411]. invadability [2496]. invader [3103, 2099, 1488]. invading [1117]. invariance [2515, 290, 565, 1249]. Invariant [1789, 374]. invariants [2631, 1814, 2534, 3184, 1515, 1602, 2706]. Invasibility [727]. Invasion [1673, 2775, 1346, 2851, 3117, 2811, 2943, 2075, 1953, 1794, 3539, 2590, 2817, 1797, 1143, 3129, 2463, 3575, 3370, 3491, 1262, 1263, 2772, 3111, 2254, 2758, 2821, 2034, 958, 513, 1804, 1145, 2374, 2548, 2774]. invasions [2347, 2672, 3493, 1381, 1326, 2135]. Invasive [3217, 2000, 3345, 2410, 2288, 2274]. Inverse [1299, 2981, 2152, 2146, 2147, 3569, 1691, 2473, 2145, 1894, 492]. inversion [3517, 3201, 2287]. invertebrate [514, 550, 551]. investigate [403]. Investigating [1363, 1382, 3089, 1822, 1852]. Investigation [3014, 1794, 1565, 2520, 1146, 1899, 3221]. investigations [580]. investment [529, 2725]. involved [2453]. involving [2725, 1064]. ion [1789, 1861, 1495]. ionic [1112]. ionizing [909]. ions [2306]. IPM [1496]. iron [2293, 381, 2131]. Irradiance [3, 27]. irradiated [871]. irradiation [3592, 909]. irregular [1335]. Irreversible [2116, 529]. ischemic [2967, 3279]. Ising [3097]. island [621, 2930, 1442, 2020, 953, 1205]. islands [324]. islet [1884]. islets [1511]. isochores [1675]. Isochrons [18]. isolated [1434]. isolation [1757, 3222, 3243, 3500, 2927]. isometry [2871]. isostable [2715]. isotonic [579]. isotope [1460]. Issue [1774, 1697, 2992, 2012, 3002]. Issues [1696]. iterated [3311, 3136]. iteration [810]. iterative [1307]. IV [36, 439, 361].

Jacobian [2275]. jamming [3573]. jellyfish [3244]. Jiangsu [3617]. JMB [1696]. join [2437]. Joining [3051]. Joint [2029, 1758, 3032, 3225, 2926, 1644, 597]. jointly [3530]. Jones [1189]. journal [1219, 2992]. jugular [2363]. jump [3172, 2724, 1540, 836, 2916, 2899, 2250]. jump-diffusion [3172]. jumps [2164, 1130, 2819, 599]. junction [1511]. junctional [1745]. junctions [2710]. Juvenile [826, 1606]. juvenile-adult [1606].

Karl [2679, 2825, 1604]. Karl-Peter [2825]. Karlin [14]. keep [1571, 2887]. Keller [1947, 1483, 3067, 1083, 1342, 3391, 2964]. Kermack [2003, 2704]. kernel [3622, 1475]. kernels [3100]. key [1072, 2463]. keystone [1245]. kidney [780, 162, 2764, 2249]. killer [318, 424]. killing [510]. Kimura [604, 3026, 251]. kin [1746, 1835, 1043]. kinases [2114]. kind [271, 636]. Kinematics [1115, 2217, 1708]. Kinetic [3190, 1542, 1134, 3006, 244, 2779, 3094, 2041, 2692, 1359, 3053, 105, 2568, 2980, 3247]. Kinetics [2394, 1556, 3552, 3494, 740, 182, 2717, 2747, 3592, 45, 3319, 2096, 716, 441, 403, 559, 1850, 3316, 3271, 951, 2343, 394, 600, 1997]. Kingman [2379]. Kirkpatrick [2851]. Kleptoparasitic [1633, 1444]. kleptoparasitism [1833, 2757]. Knot [2078]. knowledge [2762, 2545, 3456]. knowledge-based [3456]. Kolmogorov [1757, 1433, 1611, 1722, 284, 1155]. Koshland [2880, 1908]. KPP [2115, 3048, 3059, 1000]. Krebs [735]. Kries [386]. krill [3613]. Kupffer [2550]. Kuramoto [3492].

L [1064]. L. [3]. L1 [3581, 3423]. label [1838]. label-structured [1838]. labeled [2976, 1714]. labeling [2543]. labelled [1570, 2327]. labor [3506]. lack [108]. lactate [2306]. lactating [319]. lag [681, 64, 2819, 2327, 2749, 3333]. lagoon [1532]. lagopus [1436, 1436]. Lagrangian [1902, 1988, 3593, 3246]. lags [1167, 1552, 113]. Lake [3452, 2736]. lamellipod [1308, 533]. lamellipodial [2008, 2592]. lamellipodium [2599]. lamprey [342]. landscape [1253, 3394, 2232, 1640, 1423, 1518]. Landscapes [1523, 2893, 2581, 3447, 2946, 1640, 3311, 2298, 2299, 2539, 2343, 1191, 1912, 2774]. Langerhans [1511]. Langevin [2689]. language [1387, 1251]. lanternfly [3345]. Laplacian [2752]. Large [2870, 1947, 3177, 1067, 1459, 693, 3382, 3401, 2440, 2469, 3338, 3496, 280, 3284, 1564, 1265, 2121, 2175, 297, 2988, 3572, 2558, 2696, 2813, 1927, 638, 2562, 1768, 2877]. Large-sample [1067, 3338]. large-scale [2175]. large-time [1927]. larvae [3283]. larval [853, 623]. ‘Lassoing’ [2042]. last [1121]. lasting [3076]. latency [2784, 2463, 3549, 1868, 3465, 2554, 2307]. latency-age [3465]. latency-age/stage-structured [3465]. Latent [3100, 3478, 2085, 2895, 917, 1330, 3509, 2246, 2565]. latently [2560, 1836]. lateral [2401, 3087]. Lattice [1529, 904, 1143, 1954, 375, 958, 1089]. Lattice-Boltzmann [1529]. lattice-gas [1954]. lattices [3105, 1312, 3176]. Law [3321, 2875, 2549, 895, 809]. layer [420, 2392, 2132]. layers [2454]. LDL [1805]. LDL/VLDL [1805]. lead [182]. leader [3226]. leading [2599, 3595]. leaf [1460, 3601, 2969, 281, 2418, 1495]. leaf-reconstructible [2969]. leak [2024, 2069]. Leaky [210, 2156, 2578, 2935]. learned [2516]. Learning [3498, 2988, 2077, 3421]. least [3576]. leaves [3104]. Lecar [1175, 2156]. left [3055]. Lefty [2198]. Lefty-Nodal [2198]. leg [1458]. legacy [2825, 3520]. leishmaniasis [2582]. Length [2161, 1308, 3267, 3179, 2266, 3284, 1984, 1205, 3028, 575]. lent [3027]. Leonard [3415]. Lepus [1436]. Leslie [1966, 1813, 451, 1165, 2634, 1686, 3035, 912]. Leslie/Gower [2634]. less [2173, 2258]. lessons [1702, 2209]. lethality [3262]. leukemia [1946, 2677, 2822]. leukemias [2968]. leukocyte [3142, 643]. leukocytes [1047]. Level [3037, 2575, 3115, 3032, 933, 1168, 2044, 2651, 3215, 3054, 965, 1584, 3553, 2501, 1583, 1748, 2273]. Level-1 [3037, 2651, 3215, 3553]. level-2 [2273]. level-set [1584]. levels [1599, 2022, 2725, 1513, 1650, 365, 509]. Levins [1478]. Lévy [3349, 2063, 3220, 3281]. Lévy-driven [3281]. Lévy-flight [3220]. LGT [2708, 2708]. Liapunov [183, 191]. lice [1759, 1727]. Lie [2367]. Liénard [782]. Life [1244, 3304, 2039, 405, 1037, 860, 3384, 3229, 2498, 3312, 687, 737, 2095, 1635, 1784, 712]. life-cycles [1037, 860]. life-history [3384]. lifetime [110]. lifetimes [1949]. Lifshitz [3587]. ligand [2297, 263, 1307, 2564]. ligand-protein [1307]. ligands [1044, 370]. light [3552, 2356, 256, 143, 122, 3439, 499, 3433, 2517]. like [2226, 1222, 3180, 491, 748, 104, 1137]. likelihood [3031, 2174, 1916, 3100, 1619]. likelihood-based [1916]. likelihoods [1753]. likely [2859]. Limit [1862, 3025, 647, 699, 50, 2324, 1576, 3172, 2575, 3237, 313, 2169, 3151, 265, 3332, 3113, 2456, 272, 2727, 911, 1564, 3157, 293, 495, 2819, 133, 2781, 921, 2984, 3587, 782, 220, 1249, 897, 1904, 2899, 3290, 1140, 3490]. Limitations [2092]. limited [3584, 1321, 2463, 1928, 3462, 2245, 2794]. Limiting [136, 2018, 2538, 2020, 2171]. Limits [1366, 2810, 3352, 1542, 93, 1213, 3053, 974, 550, 2221, 2011, 2696, 185]. limulus [59]. line [2115, 2401, 511]. line-of-descent [511]. lineage [2351, 522]. lineage-based [2351]. linéaires [2656]. Linear [2975, 1798, 236, 1241, 500, 1489, 2281, 2656, 1537, 1179, 2631, 1548, 40, 1324, 200, 389, 1672, 3319, 333, 1762, 2477, 662, 1357, 2802, 1330, 2990, 2600, 854, 1000, 1107, 3090, 234, 3526, 2545, 16, 1356, 3233, 1138]. linearity [3526]. lines [2799, 1415, 1473, 3381]. link [2842, 1995]. linkage [347, 1288, 2530, 23, 1494, 252, 1042, 1999, 5, 49, 902]. Linked [1625, 310, 6, 1680, 1340, 833, 2543, 3554]. Linking [2014, 3561, 2232, 263]. Links [2324, 2590]. Linnaean [1950]. lipid [3586, 1712]. Lipschitzian [2684]. liquefaciens [1221]. liquid [1188]. literature [3287]. lived [3379, 3055]. liver [2786]. lives [3055]. living [602, 1784]. LNP [3471]. load [955, 3446, 2997]. load-dependent [3446]. loads [901]. Local [2798, 3265, 722, 106, 1254, 1552, 1127, 2253, 2518, 2724, 1888, 2662, 3137, 3113, 2366, 2085, 860, 3453, 175, 459, 1868, 632, 3006, 3053, 1193, 2217, 2693, 1707, 1089, 3397, 3620, 2119]. Localised [2527]. locality [916]. localization [2629, 1210, 2981, 2184]. Localized [2134, 3042, 404]. localizing [2943]. locally [2225, 1935, 3187]. located [3320]. location [2286, 506, 111]. locations [811]. loci [1895, 332, 390, 36, 807, 3389, 232, 253, 310, 790, 865, 669, 1416, 1680, 1916, 627, 764, 1625]. lockdown [3296]. Locked [3355]. Locking [860, 890, 315, 1482, 177, 345, 393, 2224]. locomotion [1281, 1282, 342, 890, 3244, 1601, 2434]. locomotor [131]. locus [2253, 347, 578, 606, 1228, 1345, 711, 1061, 1425, 719, 808, 208, 497, 857, 842, 24, 117, 225, 252, 487, 1442, 1835, 538, 1580, 121, 1659, 6, 1940, 166, 2718, 1881, 1340, 970, 1035, 838, 833, 845, 910, 923, 1744, 1207, 270]. locusts [1146]. logarithmic [237]. logDet [3292]. logical [2094, 1950]. logistic [350, 2787, 774, 3293, 3484, 1865, 3435, 3363, 3391, 245, 632, 383, 3121, 2811, 599, 2964, 201, 523]. Logistisches [523]. Long [2677, 3251, 3328, 3400, 3076, 959, 1057, 3185, 743, 1898, 1605, 1348, 3459, 1292, 598, 278, 3354, 870, 1744, 1054]. Long-lasting [3076]. long-period [1744]. Long-range [3400]. Long-term [2677, 1057, 1898, 1348, 3459, 1292, 278, 3354, 1054]. Long-time [3251, 3328, 959]. longevity [1338, 2418]. longitude [3374]. loop [1742, 112, 3074]. loops [103, 274, 1852, 2880, 2353, 2234, 620, 2127]. loss [1991, 3041, 3309, 3536, 1850, 3174, 902]. losses [2521]. Lotka [3382, 3043, 2510, 715, 1215, 3565, 1874, 3496, 607, 416, 2366, 734, 515, 141, 2778, 663, 1769, 3203, 134, 1993, 216, 3065, 653, 401, 458, 1206, 2329, 259, 72, 620, 1089, 1333, 2116, 361, 3281, 810, 3126, 3324]. Louping [1436]. Low [3392, 1365, 2920, 2022, 3032, 3474, 2473, 2627, 2836, 1701, 3388, 909, 3476, 2523, 1077]. Low- [1365]. low-copy [2022]. Low-dimensional [2920, 2627, 1077]. low-dose [2836]. low-energy [1701]. low-flashing [3433]. low-level [3032]. Low-rank [3392]. lower [1840, 365, 2933, 68, 3321]. LPA [2694]. lumped [2591, 1561]. lumping [1930]. lung [2376, 1034, 3154]. lung-slice [3154]. Luria [1276]. luteum [2326]. Lux [3271]. Lyapunov [1029, 1215, 1644, 1645, 3263]. Lyme [2066]. lymphocyte [1817, 322, 403, 2327]. lymphocytes [2625]. lymphoid [2490]. lynx [1663].

m [1307]. MacArthur [3368, 2855, 2769]. Macdonald [3022]. machines [2736]. Macro [3289]. Macro-scale [3289]. macromolecular [80, 142]. macromolecules [747]. macroparasite [1465]. macroparasites [1172]. macroparasitic [1141, 1140]. Macrophage [1440, 3586]. macrophages [3084, 3221]. Macroscopic [2683, 2338, 1252, 1609, 1906, 1435]. magic [1961]. magna [692]. magnetic [1677, 682]. magneto [3275]. magneto-responsive [3275]. magnitude [3334]. mainland [2020]. mainland-island [2020]. Maintaining [1740]. Maintenance [961, 781, 3211, 604, 1296, 439, 2884, 2230, 1807]. major [3354, 1020, 1111]. majorization [2796]. makes [2286]. making [1977, 2532]. malaccae [832]. malaria [2300, 2004, 2780, 1806, 2801, 3076, 2989, 1069, 1932, 3511, 2390, 3143, 1967, 2227, 2773, 3604, 3315]. male [1987, 1203, 3080]. males [439]. malignant [2130, 2195]. Malthusian [1451]. mammal [644]. mammalian [151]. mammals [3216]. man [757, 15, 321]. man-environment [757]. management [3302, 2373, 2511, 2903, 1889, 1866, 1496, 2087, 2889]. mange [3618]. manifold [978, 1690, 1789]. manifolds [357, 666, 1931, 3123, 22]. mantid [514, 550, 551]. Mantis [1319]. many [1753, 3523, 651, 3544, 1114, 758, 1491, 2459]. many-compartment [3544]. MAP [1072, 973, 860, 1192, 2137]. MAPK [1989, 2205]. mapped [2163]. mapping [2894, 1878, 1416, 3456, 2413]. Mappings [685]. Maps [1615, 2169, 2491, 2981, 1747, 2077, 1159]. Marginal [2318, 2251, 2785]. marine [241, 1976, 2016, 76]. mark [552]. mark-release [552]. marker [2439]. Markov [3213, 1963, 29, 3293, 3025, 2641, 2367, 2303, 3392, 3215, 1561, 514, 69, 2534, 3184, 1831, 1699, 550, 551, 1714, 952, 2309, 163, 2671, 394, 2706, 3012, 2250, 1855]. Markov-chain [3012]. Markovian [2528, 3464, 2033, 377, 1789, 1861, 2353, 1675, 2733, 2030]. marriage [920]. marriage- [920]. marrow [985]. marsh [3016]. Mass [2563, 1947, 2043, 2052, 2096, 1729, 1015, 2948, 528]. mass-action [2052, 2096]. massive [2270]. mast [641]. Master [2301, 1626, 2785, 3245]. masting [3530]. Match [3139, 2908, 3045, 3023, 3166]. matched [43]. matching [2848, 1699]. mate [1987, 1864]. material [684]. maternal [2720]. Mathematical [2637, 1594, 1953, 2686, 3425, 465, 610, 1733, 975, 3250, 1275, 2412, 1638, 1224, 1621, 2992, 2005, 2067, 2199, 586, 3116, 2801, 2297, 780, 2106, 2086, 122, 2080, 2592, 1879, 2364, 3194, 1590, 645, 2690, 1568, 1278, 2327, 2916, 1466, 3262, 2570, 61, 2625, 2344, 2024, 2069, 3559, 199, 1805, 2326, 2210, 2172, 1388, 2201, 822, 1804, 580, 892, 2048, 1506, 1687, 2938, 3028, 1631, 1344, 1405, 2492, 1534, 1597, 2244, 2881, 3612, 2264, 2655, 1565, 2458, 1415, 2289, 938, 1073, 2842, 3532, 2255, 2105, 2402, 256, 1239, 2701, 760, 2448, 1027, 2717, 402, 342, 590, 1339, 1393]. mathematical [2513, 28, 3266, 2401, 2825, 3569, 3076, 1010, 381, 1184, 3056, 3191, 1447, 1765, 2584, 145, 345, 2002, 3348, 1189, 2317, 181, 644, 1120, 2372, 337, 368, 1299, 1832, 1240, 761, 1538, 924, 1848, 1893, 3399, 1983, 322, 3307, 2654, 3520, 3014, 1391, 2045, 3287, 2010, 2857, 1908, 3010, 1839, 2667, 413, 2103, 559, 2691, 2198, 1849, 3511, 1667, 1406, 3608, 1801, 1259, 3572, 2079, 2101, 1867, 3298, 1802, 1283, 15, 633, 2503, 1550, 2769, 2107, 1306, 3443, 3174, 239, 2698, 2248, 3153, 528, 3534, 2746, 3479, 3387, 2824, 504, 3216]. Mathematics [1781, 2660, 1658, 1784]. Mathematische [61]. Mating [130, 886, 906, 82, 2756, 334, 333, 497, 399, 923, 1203, 3080]. matrices [3454, 1966, 3026, 451, 624, 1706, 2871, 512, 3480]. Matrix [2534, 3184, 3069, 396, 2670, 2828, 2099, 2520, 2771, 2524, 1901, 1165, 2017, 1686, 3035, 1843, 632, 1443, 1800, 1801, 2125, 912, 3429, 987, 3019, 3321].

Mats [2485]. matter [2414, 2729]. maturation [1209, 1369, 3582, 2928, 1467, 940, 2968, 1181, 1479, 573, 1660, 3283]. maturing [126]. Maximal [3363, 1235, 2284]. maximally [2957]. maximising [1454]. Maximization [1881, 2876]. maximize [2203]. Maximizing [3121]. Maximum [2049, 1847, 888, 1489, 3565, 1072, 2906, 3031, 2174, 3327, 2498, 1935, 3071, 2037]. may [3326, 552, 636, 2839, 2382, 3415]. Mayer [1506]. McKendrick [2003, 3359, 1165, 2704]. McKendrick-von [3359]. MCM [2301]. MCMC [1704]. MDSCs [3581]. meals [3216]. Mean [2232, 2733, 2428, 2665, 1417, 2902, 241, 3060, 2365, 2975, 2157, 1121, 3583, 2576, 2468, 457, 3526, 747, 2499, 1844, 3472]. Mean-field [2733, 2665, 2468, 2499]. means [2750]. measure [2433, 3592, 2706]. measure-valued [3592]. measurement [3469, 1470]. measurements [1724, 389]. Measures [2218, 2796, 3373, 1822, 2189, 2345, 3243, 3480]. mechanical [1334, 1914, 3260, 421, 1016, 231, 2863, 387, 3279, 615, 3396, 1200, 519]. mechanism [3469, 1649, 1343, 890, 2542, 1837, 1367, 1634, 1034, 3206, 1123, 3510, 1694, 712]. Mechanisms [1068, 1542, 2297, 3435, 2605, 3542, 2248]. mechanistic [1739, 2136, 2232, 2086, 1903, 3092]. Mechano [2537, 960, 2322]. mechano-biochemical [2322]. mechano-chemical [960]. Mechano-electric [2537]. mechanobiology [2382]. mechanochemical [2909]. media [2628, 2574, 276, 677, 2786, 2749, 3333, 1323]. medial [3440]. mediate [1128]. mediated [2264, 2655, 1350, 593, 2513, 922, 803, 905, 2925, 3010, 2647, 3098, 2249, 2564, 1047]. medical [3462]. medium [2356, 1302, 2366, 2184, 3403, 3187, 1188]. medullary [1597]. Meinhardt [152]. meioses [1494]. meiotic [2815]. melanoblast [2983]. Melnikov [1090]. meltdown [3571]. members [595]. membership [2630]. membrane [567, 2114, 1350, 265, 759, 3501, 1372, 1556, 717, 3386, 2396, 137, 2792]. membrane-mediated [1350]. membranes [46, 1985, 507, 1712]. MEMo [2973]. memoriam [2679]. memories [660]. Memory [3622, 2642, 3238, 2894, 2864, 770, 3152, 3282, 3434, 3450]. MEMS [3458]. men [3273]. Mendelian [2804, 2168, 787, 2659]. menopause [3475]. Menten [1289, 3510, 600]. Menten-type [1289]. mer [2294]. merged [3336]. merger [3046]. mers [2294]. mesenchymal [2002, 1609, 421, 1016]. mesoscale [2863]. Mesoscopic [2355, 1586, 1609]. message [2704]. meta [2831, 3538]. meta-population [2831, 3538]. metabolic [558, 2211, 2779, 373, 1608, 840, 2183, 2117, 112, 2315, 2427, 3071, 3124, 2698]. metabolism [2591, 381, 3216, 3132, 2612, 2608, 1355]. metabolite [116, 2876]. metacommunities [2996]. metagenomic [2017, 2188]. Metaheuristics [3095]. metalloporphyrin [1060]. metaphor [419]. métaphore [419]. Metapopulation [2996, 2004, 1478, 2449, 2420, 2798, 995, 1040, 1931, 2020, 1082, 1577, 2481, 2512, 3507, 1555]. metapopulations [3304, 1317, 1352, 1353, 2559]. metastabilty [3246]. Metastable [2309]. metastases [2516]. metastasis [2858, 3229, 3230]. metastatic [2170, 2516, 3098]. Method [2301, 2151, 1986, 3072, 1902, 3089, 2147, 915, 40, 741, 3105, 3176, 451, 1584, 2074, 2124, 1372, 2605, 2062, 2803, 1954, 2213, 1214, 3156, 1754, 1307, 2418, 2480, 2782, 2358, 1828]. methodological [1137]. methodology [1507]. Methods [2196, 3095, 793, 1653, 976, 2854, 3338, 1124, 32, 3392, 2928, 245, 2818, 1698, 410, 1665, 1666, 2891, 2040, 1038, 1264]. metric [3038, 2802, 2871]. metrics [2748, 2873]. metronomic [2503]. MHC [3192]. MHC-I [3192]. mice [1579, 1409]. Michaelis [1289, 3510, 600]. Micro [2502, 1179, 1531, 1938]. Micro-object [2502]. micro-organisms [1179, 1531]. micro-swimmers [1938]. microalgae [3433]. microalgal [1417]. microarray [2926]. microbe [1332]. Microbial [1588, 1554, 1830, 1122, 4, 1682, 3506, 2612, 2608, 3360, 1062, 892, 2698, 2889, 1152]. microbiome [2946]. microcirculation [717]. Microenvironment [1794, 3399, 3168]. microfluidic [2508]. microorganisms [2018, 580, 1202]. microparasitic [1140]. Microphysical [2235]. microscopic [1906]. microscopy [1779]. microstructure [3289]. microswimmer [2765]. microtissues [2807]. microtube [3575]. microtubule [2978, 2445, 2629, 1964]. microtubules [2937]. microvasculature [2130, 2195]. Migration [324, 190, 109, 3160, 724, 1807, 1808, 406, 3007, 2401, 1223, 2703, 2419, 3113, 1231, 38, 3009, 1901, 39, 1318, 3491, 1780, 3014, 1104, 1194, 3053, 3142, 929, 2916, 187, 285, 220, 1046, 1093, 1249, 1659, 937, 2689, 897, 1904, 1800, 283, 2223, 282, 2763, 1371, 1383]. migration-recombination [3160]. migratory [1146]. milk [319]. millipede [479]. Milnor [1814]. mimicry [2929]. minded [2965]. mini [2765]. mini-robots [2765]. Minimal [1455, 1345, 2728, 1231, 2541, 2506, 1972]. minimally [2868, 2503]. Minimization [824, 825, 850, 2831]. minimize [3179]. minimizes [3059]. Minimizing [328]. Minimum [2713, 453, 3095, 3565, 3061, 3037, 3372, 1790, 11, 2973, 1613, 1438]. mirabilis [1551, 1134, 1826]. misfolded [2348, 2857]. misrepair [625]. mistletoes [2265]. mite [1129, 149, 819]. mites [1607]. mitigating [3488]. mitochondrial [2032, 720]. mitosis [2029, 463, 19]. mixed [1989, 2203, 449, 242, 1144, 1640, 2712, 3356, 3210]. mixing [3379, 733, 543, 3303, 1347, 3438, 1106]. Mixture [3083, 1797, 1568, 2130, 2195, 155]. mixtures [2631]. MMPs [2686]. mobility [2440, 3264, 3503]. modality [2785]. Mode [1324, 3130, 498, 1823]. Mode-doubling [1324]. Model [1913, 1721, 1589, 1590, 1978, 2429, 1588, 2178, 3381, 1693, 2381, 210, 1238, 1188, 1591, 3016, 1495, 490, 1709, 2943, 2324, 723, 828, 1336, 2197, 1096, 1408, 2228, 930, 1322, 1273, 1946, 2970, 2253, 1545, 2784, 3405, 2306, 1361, 238, 3043, 1786, 2805, 2875, 2300, 2263, 3158, 1880, 1791, 1029, 3213, 3163, 419, 817, 1362, 2004, 2665, 1366, 2986, 578, 606, 2264, 2655, 996, 2151, 1551, 2003, 2411, 2556, 567, 1576, 1008, 2780, 3467, 351, 1565, 2905, 3379, 3347, 1013, 793, 901, 1478, 2707, 1415, 841, 1246, 1247, 1497, 1364, 2878, 440, 2136, 2890, 938, 610, 1599, 442]. model [2510, 1767, 1960, 1517, 1004, 1073, 1303, 3532, 2255, 1947, 1996, 1753, 3200, 1806, 3402, 1895, 2575, 2998, 3083, 3222, 2614, 1228, 1345, 2128, 2526, 3144, 1294, 2804, 947, 29, 3032, 1354, 1737, 698, 1907, 1833, 724, 2402, 2687, 1166, 1985, 3474, 634, 2882, 604, 3523, 256, 434, 776, 2563, 346, 1198, 1483, 2154, 1009, 1119, 1239, 1349, 3565, 332, 3373, 2380, 3238, 2212, 1977, 2226, 2701, 3393, 760, 802, 1711, 1943, 2448, 2252, 1027, 1087, 1175, 1750, 1724, 1988, 3069, 3086, 908, 690, 1168, 3036, 1472, 2552, 2553, 1575, 711, 807, 2714, 2580, 2412, 2668, 2138]. model [75, 3517, 3131, 309, 342, 2168, 590, 1129, 504, 1085, 3592, 2756, 3301, 3389, 2122, 2321, 2478, 2454, 1324, 1425, 3226, 1797, 3201, 1001, 2716, 1339, 1393, 362, 1606, 2619, 2779, 3077, 1101, 3241, 2236, 2513, 2527, 3252, 1896, 28, 3266, 1955, 3446, 2536, 1530, 918, 3504, 2401, 517, 922, 1223, 352, 2489, 2843, 2156, 3601, 1154, 1094, 1147, 2153, 3029, 537, 651, 389, 2684, 2194, 2529, 3458, 2914, 1738, 2578, 1857, 2912, 2528, 1308, 1810, 3214, 3444, 2797, 3101, 3400, 3586, 1286, 2051, 3484, 3344, 2692, 1484, 2414, 286, 832, 1824, 317, 474, 1348, 1134]. model [2530, 1182, 2971, 719, 3267, 192, 1231, 1098, 2622, 1899, 1067, 516, 92, 3066, 38, 2557, 320, 1539, 2636, 1734, 381, 771, 1826, 1184, 1421, 1525, 3056, 3191, 3362, 1586, 759, 167, 3368, 3067, 3145, 3015, 2131, 1363, 1382, 3042, 1765, 1918, 780, 1069, 2456, 913, 1399, 131, 1829, 2473, 1390, 3406, 1644, 333, 3009, 177, 1090, 3525, 3165, 3457, 1002, 2447, 1321, 1467, 2076, 1815, 439, 758, 566, 257, 999, 2573, 755, 345, 3054, 3181, 2085, 3164, 3170, 3549, 3459, 2000, 800, 995, 2002, 122, 2694, 152, 3295, 3616, 2516, 554, 696, 3348, 397, 1117, 2215]. model [1189, 781, 2939, 2661, 2897, 514, 2317, 2844, 1132, 1204, 1083, 721, 1039, 1865, 2276, 1076, 1529, 3618, 3031, 1729, 3588, 3355, 1564, 637, 39, 1342, 2372, 134, 221, 295, 983, 1310, 1993, 1379, 3462, 2238, 337, 1116, 786, 1632, 2814, 1474, 2676, 3079, 368, 468, 114, 147, 1040, 625, 3501, 3228, 2534, 3184, 86, 861, 3317, 3416, 1713, 2477, 3617, 80, 89, 1240, 1407, 2162, 1538, 162, 2192, 3194, 3529, 3597, 3363, 2909, 924, 33, 24, 117, 225, 252, 809, 2929, 1677, 545, 3452, 2894, 251, 3217, 3350, 1271, 1893, 2342, 1983, 745, 1883, 3519, 322, 2322, 2241, 3013, 1448, 662]. model [2161, 2729, 3391, 3204, 2819, 2654, 2506, 3257, 3308, 753, 731, 165, 3097, 2925, 2855, 2958, 1780, 1790, 3286, 956, 3014, 1391, 1831, 2351, 3337, 3573, 659, 2045, 108, 2678, 404, 3309, 2010, 3477, 2857, 1293, 1681, 1442, 1654, 136, 229, 311, 2082, 2207, 1109, 1868, 1908, 1981, 2064, 2068, 2772, 2885, 2895, 2956, 3010, 3108, 3119, 3365, 3609, 538, 1261, 843, 879, 2699, 1839, 632, 2546, 2627, 2073, 1074, 1177, 2595, 2667, 2968, 3259, 413, 1932, 3310, 3053, 3404, 3437, 1292, 709, 814, 929, 1033, 1148, 1878, 1838, 2089, 3528, 3070, 3580, 546, 881, 750, 779, 2647, 2864, 224, 1392, 1937, 559, 2254]. model [1620, 2171, 2256, 2947, 3441, 318, 424, 550, 551, 2121, 3587, 2800, 2198, 180, 1580, 2851, 1214, 1849, 2953, 1193, 927, 2921, 121, 694, 2884, 1060, 611, 2719, 1652, 1280, 1623, 3409, 2941, 421, 866, 856, 157, 1046, 1093, 1527, 3312, 641, 2659, 1656, 1667, 1015, 2390, 1016, 3614, 385, 616, 2031, 284, 630, 1904, 384, 1330, 2158, 2966, 231, 2271, 3148, 388, 2863, 2666, 2148, 1406, 1479, 387, 533, 436, 1257, 941, 6, 1884, 3278, 3339, 3412, 3613, 3608, 3448, 2484, 3341, 582, 615, 2568, 2795, 269, 411, 2994, 1309, 2718, 205, 2334, 195, 2098, 2326, 994, 3410]. model [1141, 2840, 2350, 3021, 3572, 2980, 2079, 2101, 1200, 3403, 1867, 2768, 3298, 3271, 1082, 125, 671, 348, 1163, 2862, 1243, 1802, 2378, 574, 2990, 1283, 1577, 605, 15, 1616, 2927, 2514, 3354, 633, 3280, 2644, 1089, 282, 2503, 3173, 742, 457, 1574, 2130, 2195, 460, 557, 658, 2871, 3272, 3064, 2769, 2107, 2201, 822, 1306, 3143, 1614, 2229, 2442, 3232, 115, 1333, 2512, 233, 426, 425, 519, 639, 888, 1062, 1099, 1870, 3174, 2257, 806, 2886, 2116, 239, 1019, 1035, 1028, 2709, 2269, 1727, 3219, 648, 3623, 2428, 1732, 2948, 3366, 747, 2499, 910, 923, 987, 1876, 2305]. model [1841, 2209, 2048, 2088, 101, 1716, 2961, 2532, 588, 1413, 2396, 643, 1047, 1481, 3541, 14, 703, 124, 572, 1264, 3481, 1744, 749, 799, 1655, 3246, 2072, 3012, 2248, 2537, 1414, 1972, 3153, 3364, 392, 3535, 63, 1660, 2179, 2265, 2295, 2658, 2765, 2773, 2808, 2865, 2910, 2932, 2866, 3022, 3209, 3299, 3282, 3420, 3524, 1854, 1014, 3096, 358, 575, 2214, 2173, 2258, 2358, 528, 1373, 294, 1314, 2417, 1185, 819, 3534, 1820, 1153, 2935, 839, 2571, 2755, 3554, 591, 2639, 3324, 3283, 2307, 2250, 3508, 3233, 3492, 1597, 2244, 2746, 3146, 3149, 3210, 3377, 3315, 3387, 3436, 3555, 2066]. model [2833, 2216, 2824, 3334, 2675, 3490, 22, 864, 1555, 1242]. Model-consistent [1693]. model-specific [3213]. Model-supported [1913]. modèle [419, 2556]. modeled [894, 313, 2924, 3204, 998, 1841, 1855]. modèles [3027]. Modeling [1096, 1758, 552, 991, 2356, 1926, 2741, 2170, 1837, 2461, 3084, 2084, 1678, 1509, 2222, 1161, 2983, 1956, 1262, 2057, 1358, 3423, 3516, 1554, 2967, 1617, 2786, 2501, 1761, 2223, 1079, 2288, 3187, 1783, 2259, 1982, 3080, 3073, 1953, 960, 3445, 3177, 2105, 2858, 3250, 2441, 2435, 3605, 972, 2641, 2199, 2801, 2767, 3076, 1284, 2962, 3359, 3603, 2584, 3303, 2476, 2613, 3216, 3487, 2592, 2038, 3374, 2388, 1873, 1723, 3399, 3168, 2690, 1278, 1466, 2740, 2570, 2625, 2178, 3511, 2344, 3570, 3230, 199, 3589, 3279, 3207, 2446, 3388, 1825, 2210, 1071, 1295, 1927, 3497, 3482, 1909, 2492, 1964]. modeling [3248, 1792, 3427, 2582]. modelled [1643]. Modelling [1559, 1541, 1167, 1840, 1460, 1473, 3089, 1513, 1650, 915, 1126, 3624, 3007, 1607, 2674, 1422, 2302, 2346, 3009, 3575, 2266, 3192, 2453, 303, 2240, 2915, 3053, 2745, 766, 1123, 1800, 1771, 2997, 1500, 1125, 2480, 957, 1712, 2569, 1860, 3538, 3620, 3291, 1532, 626, 2150, 203, 248, 2686, 2065, 2056, 1453, 240, 1483, 330, 1772, 3358, 2621, 2408, 1224, 2005, 2419, 102, 1447, 1777, 2736, 2297, 2106, 1901, 3439, 1879, 2364, 720, 3491, 1501, 757, 3287, 3242, 1568, 1793, 2916, 2341, 3622, 1487, 1798, 1805, 1624, 1801, 2560, 1550, 83, 3443, 2355, 17, 1804, 935, 2551, 1456]. modelling [2938, 3028, 1344, 1405, 3466, 2639, 2434]. Models [788, 1281, 1282, 1548, 1776, 1755, 803, 1222, 846, 919, 714, 713, 2757, 1544, 364, 433, 488, 691, 573, 1159, 2337, 891, 3596, 3178, 1720, 3469, 1050, 2218, 3088, 3167, 223, 1023, 500, 592, 3213, 2176, 347, 1088, 1489, 1637, 2060, 3027, 3030, 1959, 3425, 431, 1507, 2602, 2737, 2888, 2449, 2420, 301, 3415, 2945, 3235, 1091, 470, 2931, 708, 2211, 2777, 3574, 2422, 2323, 3249, 1915, 312, 1215, 1670, 1752, 3120, 1432, 1875, 2425, 412, 406, 556, 547, 975, 2724, 1795, 2638, 3223, 2591, 2590, 1439, 2993, 3322, 967, 2147, 82, 1492, 2631, 733, 774]. models [2202, 3005, 3276, 3041, 2717, 727, 1275, 3562, 1208, 1898, 396, 2681, 2232, 661, 1209, 1369, 304, 1567, 1061, 2159, 94, 140, 1813, 2670, 2828, 1370, 1376, 2139, 3332, 3302, 2182, 1270, 784, 1138, 1301, 1888, 2662, 3000, 3414, 543, 675, 3190, 486, 1380, 1395, 1396, 1542, 2703, 565, 2164, 2578, 1726, 2507, 1133, 2278, 3113, 962, 3116, 2041, 2287, 3105, 3176, 1596, 3544, 473, 746, 1143, 1636, 3008, 2445, 2751, 1562, 681, 3033, 1105, 2367, 1493, 95, 471, 549, 323, 3359, 1766, 1759, 3046, 3548, 883, 3593, 365, 329, 1482, 3392, 2086, 1763, 520, 1078, 3372, 145]. models [953, 96, 1535, 896, 959, 1459, 310, 3543, 2896, 85, 3550, 2511, 860, 2359, 948, 7, 25, 261, 940, 1762, 2111, 3407, 120, 603, 640, 1903, 3453, 3011, 763, 272, 276, 2200, 1162, 3522, 920, 448, 1063, 1165, 1232, 214, 331, 816, 1220, 2977, 3438, 1609, 1778, 1120, 3557, 1952, 2753, 2975, 445, 526, 1508, 377, 1431, 1299, 3157, 2357, 1227, 1325, 2634, 2487, 761, 2742, 169, 1945, 178, 487, 809, 374, 435, 1452, 1686, 3035, 645, 2903, 3128, 949, 2242, 1949, 2764, 2016, 1842, 324, 1372, 1457, 1377, 1835, 2616, 1921, 1357, 3345, 700, 1428, 2960, 2845, 725, 247]. models [344, 3297, 1919, 569, 650, 3147, 3465, 3602, 1226, 3122, 3260, 3461, 3463, 3006, 3142, 1291, 1906, 1443, 2327, 2995, 2826, 2705, 775, 1142, 1663, 1476, 1400, 1304, 3140, 2567, 1931, 2900, 3352, 882, 2570, 1954, 505, 2221, 884, 1665, 1714, 1418, 3205, 2891, 2024, 2069, 3185, 598, 2175, 2165, 478, 130, 3356, 2689, 2612, 2608, 199, 258, 375, 1940, 3263, 3510, 3509, 2788, 166, 3564, 3316, 2829, 2831, 153, 3493, 2462, 767, 3154, 1935, 1011, 2026, 2671, 2620, 3270, 2728, 1692, 555, 2140, 2040, 213, 925, 3500, 2481, 1613, 2186, 2172, 744, 614, 2555, 2587, 1241, 3092]. models [1279, 531, 889, 1748, 1155, 2733, 1578, 3090, 2001, 3476, 912, 1930, 2054, 447, 2763, 1191, 275, 1127, 936, 969, 3497, 109, 892, 1095, 1446, 1496, 2087, 3408, 1058, 1170, 2030, 2285, 1287, 3314, 2444, 692, 2077, 3100, 296, 3118, 539, 3289, 1890, 2782, 1100, 1687, 680, 3507, 942, 50, 2889, 3093, 3456, 3599, 1255, 1385, 1449, 1356, 1371, 1383, 1751, 1054, 207, 1084, 1463, 1174, 2704, 2950, 3247, 2562, 3220, 2554, 1031, 3577, 3563, 3421, 1290, 1140, 2561, 988, 2398, 270, 1828, 2119, 2650]. moderate [1938]. moderately [3249]. modern [3370]. modes [2206, 2328, 1962, 779, 2427, 2973, 3533]. modification [2096, 1901, 3607, 1042, 244, 3492]. modifications [2629]. modified [3097, 163]. modular [2017]. modulated [1971]. Modulation [2006, 1482, 1102]. modules [2315]. moisture [1461, 1462]. molar [1359]. Mold [1590, 1661]. Molecular [115, 2150, 1528, 2703, 3008, 3033, 308, 1498, 2538, 3413, 1760, 2734]. Molecule [1319]. molecules [89, 2872]. Moment [1611, 160, 2562, 1722, 1935, 2901, 2030]. moment-based [2901]. Moment-flux [2562]. Moments [858, 391, 99, 904, 741, 2301, 110, 747]. monodomain [2454]. monoecious [910, 923]. monogonont [1239, 1349]. monolayer [1690]. monolayers [2011]. monomolecular [1626, 3245]. monophasic [3599]. Monostability [3255]. Monotone [1361, 1762, 3021, 323, 929, 810]. monotonic [727]. Monotonically [1615, 1747]. Monotonicity [2581, 1900, 2456, 1230]. monoxide [297]. Monte [3223, 2747, 3498]. Moran [1895, 2714, 3301, 2633, 2507, 2530, 3081, 3091, 3583, 1854, 2877]. Moran-type [2714]. Morone [853]. morphoelastic [3214]. morphogen [1649, 3066]. morphogen-controlled [1649]. morphogenesis [355, 698, 152, 2764, 1034, 779, 172, 421, 478, 1016, 231, 615, 179, 1079, 519]. morphogenetic [496, 79]. morphogens [1514]. morphological [3235, 1832, 2795]. morphology [1047, 1475, 3321]. Morris [1175, 2156]. mort [2656]. mortality [2228, 2470, 1766, 686, 2678, 3331, 1031]. mosaics [48]. mosquito [2637, 2737, 2674, 2915, 2390, 3021, 2883, 2227, 3080, 3436, 3417, 3334]. mosquito-borne [3417]. mosquitoes [2267, 1607, 2076, 2699, 3283, 2720]. most [2859, 3559, 1331, 1992]. motif [2284]. motifs [2334, 1707]. motile [3147]. motility [2408, 922, 2291, 2463, 2038, 2282, 1781, 533, 819, 2774]. motion [2619, 3458, 3344, 666, 1609, 1431, 2983, 2502, 2035, 3473]. motivated [1, 3601, 3242]. motor [3189, 3032, 2629, 792, 2272]. motor-driven [2629]. motors [2937, 1776, 3320]. mountain [1436]. mouse [2815, 1675]. move [2532]. Movement [1590, 1545, 2737, 29, 2232, 1380, 1395, 1396, 2041, 1493, 1788, 2014, 100, 1723, 2016, 1358, 2995, 3390, 2555, 2587, 3152, 2374, 643, 3456, 2167, 3220, 3450]. movements [1486, 525]. moving [81, 479, 2819, 2839, 582]. moving-optimum [2819]. MPF [975]. MRI [3300]. mRNA [3234, 2022, 3471, 3316]. mRNA-LNP-based [3471]. MS2 [170]. MSEIR [1632]. Mt. [1253]. MTT [2365]. mucus [3211]. MUL [2976]. MUL-trees [2976]. mule [2291]. Muller [942]. Multi [2777, 171, 2779, 3332, 2767, 672, 1864, 1503, 2228, 3468, 1366, 2323, 3177, 2128, 2452, 2687, 2832, 2202, 2139, 1955, 3499, 2588, 3484, 2751, 2134, 2622, 1816, 3432, 1777, 3372, 714, 3054, 1570, 1254, 1948, 2566, 117, 3343, 1859, 1835, 1480, 121, 284, 375, 1360, 1420, 1676, 166, 2786, 513, 2209, 1712, 3427]. Multi-algebra [672]. multi-allele [117, 1835, 166]. multi-alleles [375]. multi-allelic [284]. Multi-attribute [1864]. multi-cell [1480]. multi-cellular [2139, 1777]. multi-compartment [2786]. multi-component [1712]. multi-dimensional [1816]. multi-epitope [2687, 2832]. multi-group [2777, 2751]. multi-host [3468]. multi-labelled [1570]. multi-locus [121, 166]. multi-oral [2622]. multi-parameter [2134]. Multi-patch [2777, 1366, 3484, 1254]. multi-player [2452, 2588, 3343]. Multi-scale [2767, 3177]. Multi-species [171, 3332, 3499, 714, 1676, 513]. multi-stability [3372, 3427]. multi-stage [1955]. multi-strain [1503, 2228, 2323, 2202, 3054]. multi-strategy [2588]. multi-time-scale [2566, 1859]. Multi-timescale [2779]. multi-type [2128, 3432, 1948, 1360, 1420]. multi-year [2209]. multiallele [225, 487, 1566, 1881]. multiallelic [1659]. multicell [1450, 1302]. multicomponent [2669]. multidimensional [474, 874, 384]. multiforme [3300]. Multilevel [2722, 2923, 3328]. Multilocus [1807, 1808, 1092, 835, 1499, 961, 895, 178, 809, 1057, 1178, 630, 1381, 1566]. multilocus-multiallele [1566]. multimeric [2464]. multimodel [1794]. multinomial [3546]. Multiparameter [681]. Multiparametric [2154, 1074]. Multiphase [1801, 1483, 1568, 1681, 3096]. multiplayer [2466]. Multiple [1576, 1812, 812, 272, 1779, 2772, 1699, 1441, 531, 426, 2937, 1753, 1895, 1186, 3373, 2993, 1097, 711, 807, 2670, 1370, 1563, 3046, 3525, 1327, 2511, 3085, 3522, 2375, 3013, 3531, 3477, 2956, 3108, 2667, 2968, 2519, 112, 1849, 3356, 58, 1257, 2967, 3279, 2572, 2350, 3556, 639, 326, 1735, 2865, 1736, 1820, 3146, 22, 36]. multiple-loop [112]. multiple-merger [3046]. multiple-replication [1820]. Multiple-spike [1441]. multiple-timescale [3477]. multiplex [2753]. multiplicative [807, 1423]. multiplicative-fitness [1423]. multiplicatively [3090]. multiplicatively-closed [3090]. multiplicity [3203]. Multiscale [2805, 2065, 2056, 3002, 3168, 1793, 2257, 1792, 1794, 2433, 2414, 3337, 1624, 1935, 3620]. multisite [3607, 2310]. multisolute [702, 780]. Multispecies [341, 3600, 3294, 2609, 35, 2948]. Multistability [1665]. Multistage [1477, 366]. multistate [230]. Multistationarity [2043, 2880]. multistep [2987]. multistrain [1662]. multitype [943, 474, 641, 876, 583, 2442]. multivalent [2412, 370]. Multivariate [3256, 2496]. Mum [1571]. murine [2760]. Mus [319]. muscle [947, 590, 504, 1644, 878, 78]. muscle-skeletal [1644]. muscular [92, 31, 58, 105]. musculus [319]. mussel [2870, 2380]. mussel-algae [2380]. mustard [3]. mutant [875, 943, 3052, 634, 1715, 317, 474, 873, 874, 1377, 1654, 704, 1340, 2982]. mutants [1253]. Mutation [1288, 1637, 955, 877, 2799, 2581, 2882, 2991, 3265, 1437, 1646, 2722, 2633, 3135, 253, 781, 562, 1509, 251, 1949, 324, 1835, 3542, 1518, 898, 384, 553, 1940, 1999, 109, 747, 833, 845, 2088, 2021, 1820]. Mutational [2547, 3571]. mutations [2761, 1895, 1024, 2187, 3477, 2369, 3174, 902]. Mutual [855, 2496, 1419]. Mutualism [3595, 2865]. mutualism-competition [2865]. mutualistic [968, 2265]. Mutually [168, 535]. Mycobacteriumtuberculosis [1440]. myelin [3402]. myelinated [442, 566, 3388]. myeloid [1946, 2677, 2822]. myocardium [834, 2537]. myofibroblasts [1893]. myosin [2641, 3320, 2592]. myotonia [146]. mystery [3476].

n [93, 1307]. NaCl [2249]. Nagumo [1618, 899, 249, 336, 1511, 125, 1107, 367, 2935, 541]. Nagylaki [910, 923]. Naidis [1286]. naissance [2656]. naive [2617]. nanophysiology [3187]. narrow [150, 174, 129, 2365]. Nash [449, 1251]. Natural [427, 2963, 2881, 3454, 2643, 2761, 2750, 82, 1087, 1425, 2672, 318, 424, 314, 3604]. Nature [1991, 342, 3513, 3206]. navigation [2875, 1549]. Near [2623, 1244, 2496, 448, 1811, 2023, 2965]. Near-critical [2623]. near-optimal [2965]. nearest [3127, 2482]. nearest-neighbor [2482]. Necessary [2585, 1925, 3519, 2127, 386]. necrosis [930, 1480, 827]. necrosis- [930]. necrotic [3404]. need [2385]. needle [1347]. needle-sharing [1347]. needles [1347]. Negative [329, 1709, 2377, 1816, 235, 2364, 2017, 3119, 1064, 112, 3381, 1969, 2370, 2763, 20, 3150]. negatively [1524]. neglected [1054]. Negligible [551, 1850]. neighbor [2482]. neighbour [3127]. nematode [1472]. Neolithic [3145]. nephron [780]. Nernst [2843]. nerve [772, 619, 12, 566, 1842, 132, 249, 269, 37, 125, 47]. Nested [1595, 2687, 2352, 3452]. net [1298, 1265, 1261, 3320]. nets [739, 2094, 3076, 173, 2773, 2708]. network [3292, 3395, 3401, 2683, 2387, 2123, 2673, 2905, 1364, 2211, 1225, 2687, 2832, 3394, 1711, 1943, 3383, 2779, 1929, 3063, 1562, 2748, 3407, 3615, 3512, 3228, 1552, 2859, 3097, 2721, 2395, 2429, 1919, 1192, 2089, 775, 1937, 2490, 2427, 1441, 2290, 3049, 2428, 2499, 2918, 3150]. Networks [2970, 1548, 2280, 2959, 3058, 3600, 1900, 2726, 2945, 3361, 2515, 3547, 2535, 2791, 2902, 3115, 2452, 3239, 2708, 3358, 2169, 933, 1822, 3025, 2790, 3039, 2043, 1567, 3330, 3137, 1827, 2052, 3323, 3048, 3037, 1672, 3336, 2969, 2096, 2044, 2651, 2303, 2607, 3034, 2268, 3051, 3215, 2757, 2976, 3183, 2753, 3285, 1570, 2589, 2482, 2479, 2331, 3306, 2759, 3607, 2566, 2353, 3519, 2155, 3307, 1036, 2738, 2275, 3128, 1859, 2809, 3365, 1905, 3104, 3553, 3461, 3463, 1863, 2282, 112, 3495, 1924, 1665, 1666, 2315, 2386, 2330, 1819, 3263, 973, 2873, 1974, 2070, 2079, 2101, 2785, 2468, 1267, 2600, 2292]. networks [2626, 824, 825, 3218, 1583, 2733, 3174, 877, 998, 1150, 2635, 3497, 2974, 1170, 2285, 2374, 3342, 2922, 2220, 1710, 3188, 3397, 2571, 1855, 3149, 2273, 2533]. Neumann [2754]. Neural [2141, 2151, 2683, 609, 3358, 2169, 1711, 1943, 2483, 933, 739, 1298, 812, 2404, 2134, 3603, 3009, 175, 2958, 2956, 775, 2567, 2256, 1441, 3049, 1170, 3610, 1030, 3074, 2119]. neuro [31, 58]. neuro-muscular [31, 58]. neuroglobin [2525]. neuromuscular [159]. Neuron [210, 996, 567, 1721, 1432, 2591, 3393, 2156, 759, 1271, 388, 380, 2728, 2040, 2935]. neuronal [2970, 1008, 903, 2182, 2578, 173, 1819, 2146, 671]. neurones [2443]. neurons [3401, 2902, 1225, 1711, 1943, 2164, 1564, 2055, 2845, 2028, 2224, 98]. neurophysiological [419]. neurophysiologique [419]. neuroscience [1977, 1638]. Neurotransmitter [160, 1721]. neutral [1880, 2312, 2439, 651, 250, 3304, 637, 251, 1949, 3531, 284, 375, 630, 898, 1680, 2410, 296, 1890, 1625]. neutral-alleles [296]. neutrality [428]. Neutrally [37]. neutrophil [15]. next [2125]. next-generation [2125]. NF [2358]. NF- [2358]. Niche [3439, 3107, 1968]. nickel [2938]. Nile [2267, 2674, 2699, 3410, 3021, 2932]. NIMFA [3114]. nitric [3440, 1597]. nitrogen [2517]. nitrogen-fixing [2517]. NMF [2926]. No [1212, 408, 1875, 2580, 767]. No-escape [408]. Nodal [1887, 2198]. node [3402, 3034, 3285]. Noise [3358, 3330, 1971, 2594, 1320, 2640, 2206, 1943, 2810, 3319, 2603, 2463, 2157, 2538, 1025, 614, 2006, 3220]. Noise-driven [3358, 1971]. Noisy [2578, 3047, 2902, 3516, 3509]. Non [1179, 916, 200, 1816, 2049, 2366, 1558, 854, 234, 2267, 1253, 2518, 2497, 2724, 1874, 2252, 40, 3094, 3137, 1147, 1852, 2684, 2296, 3008, 3055, 143, 722, 2085, 2215, 377, 2017, 3195, 325, 401, 1104, 1868, 3006, 3053, 929, 1882, 2567, 1869, 1193, 2175, 2693, 1436, 699, 1850, 3156, 97, 418, 481, 480, 581, 3169, 2990, 909, 1000, 1107, 2733, 1434, 3187, 1475, 16, 2889, 1749, 2950, 2517, 3620]. non-additivity [1253]. non-autonomous [1882]. non-Boolean [97]. non-catastrophic [2497]. non-causal [1852]. non-compact [3169]. non-constant [325, 401, 16]. non-corvids [2267]. non-diffusive [2252, 2296]. non-equilibrium [1147]. non-existence [1104]. non-extinction [1874, 699]. non-Gaussian [3008]. Non-hereditary [2049]. non-homogeneous [143]. non-identifiability [2215]. non-isolated [1434]. Non-linear [1179, 200, 854, 234, 40, 2990, 1000, 1107, 16]. non-Lipschitzian [2684]. non-local [2518, 2724, 3137, 722, 2085, 1868, 3006, 3053, 1193, 2693, 3620]. Non-locality [916]. non-locally [3187]. non-Markovian [377, 2733]. non-monotone [929]. non-negative [2017]. non-negligible [1850]. non-nitrogen-fixing [2517]. non-parallel [1749]. non-parametric [2175, 3156]. Non-periodicity [1816]. non-pharmaceutical [3195]. non-progressors [1869]. non-proliferating [909]. non-pulsatile [2950]. non-reducible [418, 481, 480, 581]. non-smooth [2567, 2889]. non-standard [3094]. non-stationary [3055]. non-uniform [1475]. non-viraemic [1436]. nonautonomous [2228, 3515, 3614, 2576, 1287, 749]. noncoding [1698, 1705]. nonconstant [918]. noncooperative [1049]. nondensely [3515]. nonepistatic [178]. nonexistence [724]. nonexpanding [3361]. Nonhomogeneous [13]. Nonlinear [412, 2380, 1384, 1797, 3484, 2894, 34, 615, 2795, 507, 1207, 1545, 3212, 740, 817, 868, 596, 971, 1198, 1009, 915, 619, 3131, 1209, 1369, 489, 954, 352, 1810, 1691, 1535, 93, 800, 41, 1367, 1132, 721, 816, 444, 1952, 429, 477, 395, 518, 1686, 3035, 693, 1263, 3119, 1187, 569, 650, 407, 1793, 2611, 531, 912, 1028, 2048, 499, 84, 993, 2702, 864, 1301]. nonlinear-diffusion [993]. nonlinearly [98]. Nonlocal [3113, 2830, 3291, 2563, 3105, 3176, 2797, 1899, 3362, 3372, 1045, 1256, 1321, 2896, 3194, 3609, 2597, 2484, 3410, 3232, 2288, 3456, 3016]. nonmonotone [1785, 3339, 3412, 889, 1751]. nonnegative [3088]. nonnegative-valued [3088]. nonpacemaker [659]. nonpharmaceutical [3224, 3602]. nonphotochemical [2594]. nonrandom [512]. nonsteady [655]. nontraditional [334]. Normal [2700, 2791, 3277, 3104, 1799, 15, 2130, 2195]. normalized [2422]. norms [745]. Northern [2736]. nosocomial [3524]. Note [1600, 960, 48, 1924, 738, 3062, 811, 1019, 935, 492]. nothing [494]. notions [2846]. Novel [3564, 3469, 1505, 2857, 1937, 1954, 2691, 3153]. NP [1072]. NP-hard [1072]. NPZ [2241]. Npzd [3528]. Nuclear [1410, 2605]. nucleation [2602]. nucleotide [170, 1355]. nucleotides [2059, 1561]. nucleus [2365]. Number [1522, 494, 621, 1939, 2003, 2208, 1345, 2083, 2203, 2853, 271, 3179, 1543, 310, 1494, 2394, 3159, 3346, 2050, 2126, 2942, 3529, 225, 251, 2155, 324, 1442, 1781, 1131, 1973, 1121, 2831, 1693, 2125, 2140, 2813, 2185, 2403, 1340, 1438, 1735, 2523, 1828]. numbers [3284, 2157, 2917, 3345, 3156, 985, 3093]. Numerical [723, 1417, 8, 619, 1343, 2194, 1284, 2613, 2276, 827, 1846, 936, 969, 1385, 3347, 2602, 3072, 1285, 741, 409, 2574, 771, 2663, 2584, 2106, 2928, 3075, 2038, 3217, 1372, 410, 3587, 884, 1214, 1761, 655, 600, 3480, 1823]. nurse [131]. nutrient [769, 919, 461, 120, 1162, 992, 221, 1227, 2241, 2899, 1283, 925, 1095, 839, 1152, 1332]. nutrient-phytoplankton-zooplankton [1227]. nutrient-taxis [2899]. nutrients [1563, 3210]. nystagmus [1545].

O [3396]. O157 [2295]. oasis [1244]. obeying [382]. object [2502]. objective [1894, 2702]. objects [143, 1779, 673]. oblate [477]. Oblique [420]. observational [1509]. observed [2731, 1540, 3483, 1651, 1340]. obtain [2375]. obtained [2077]. occasion [2485, 2109]. Occupancy [3276, 2232, 2171]. occurrence [2994, 2639]. occurring [729, 2921, 394]. ODE [1364, 3550, 2215, 2975, 2615, 3271]. ODEs [1643, 2184]. ODEs-homogenization [2184]. Odo [2109]. oesophagus [3608]. off [2277]. offs [2375, 3324]. Ohta [251]. old [2697]. oldest [842]. olfaction [1691]. olfactory [1319]. oligodendrocytes [1294]. oligomers [3158]. Oligomorphic [1970]. olive [2055]. omnivory [365]. One [246, 24, 380, 166, 1791, 1345, 1225, 1432, 313, 2328, 1724, 1472, 2668, 1728, 3214, 264, 2557, 986, 852, 1310, 836, 225, 487, 688, 3391, 1835, 1192, 3122, 2137, 180, 598, 1441, 1389, 2290, 1881, 1028, 799, 2537, 2865, 1014, 819]. One- [166, 598]. one-compartment [1724]. One-dimensional [246, 1791, 1225, 1432, 1472, 3214, 2557, 852, 836, 3391, 1192, 3122, 2137, 1441, 1389, 1028, 2537, 1014, 819]. one-hump [986]. One-locus [24, 1345, 225, 487, 1835, 1881]. one-prey [1310]. only [2552, 2553]. onset [2122, 1370, 1278, 3354]. onto [860]. Ontogenetic [2120]. ontogenic [3601]. oocytes [1989]. Open [3456, 2965, 1534, 939, 2744, 1908, 2064, 2320, 3310, 11, 3117, 1712, 3074]. open-loop [3074]. Open-minded [2965]. Operating [3360]. operations [2482]. operator [489, 2685, 3204, 407, 3169]. operators [2598]. Operon [3268, 1879]. opioid [2890]. Opposition [628]. optical [2735, 2413]. Optimal [3584, 1158, 2890, 2108, 2097, 312, 529, 2407, 2879, 2770, 1208, 1454, 3411, 2536, 2579, 1928, 3195, 1110, 2903, 2010, 3428, 3506, 2953, 805, 2160, 73, 118, 255, 1974, 2990, 599, 3460, 2709, 2698, 3074, 2876, 3421, 1152, 1332, 2723, 3095, 1960, 3325, 2311, 3222, 1215, 2318, 2858, 316, 1571, 92, 1010, 2225, 1535, 1458, 327, 1144, 3305, 3557, 3079, 3059, 3229, 487, 2345, 1104, 3371, 697, 2028, 2239, 2166, 2652, 3288, 3585, 3092, 2418, 1889, 1866, 2116, 2007, 2965, 3233, 2582]. optimality [3251, 1365, 824, 825]. optimisation [3239, 3560, 1925]. optimising [3327]. Optimization [2505, 1091, 2348, 2822, 1709, 3189, 1980, 2878, 2614, 3134, 3110, 775, 3367, 1954, 3311, 2427, 2754, 1692, 1894]. Optimizing [1688]. optimum [2819]. option [1234]. oral [2622]. orbital [2239]. orbits [1196, 1094, 1686, 2055, 1863, 2023]. orchard [3218]. orchards [1081]. order [1720, 3058, 222, 3425, 3330, 1596, 441, 790, 865, 3038, 87, 787, 1064, 1142, 3206, 3230, 1850, 2785, 408, 1987]. orders [2803, 2783]. ordinary [3000, 670, 1665, 866, 305, 2305]. ore [2657]. organic [1983, 2729]. organisation [1777, 1821]. organised [2425]. organism [278]. organisms [1273, 1179, 1531, 1117, 1041, 1416, 2802, 1381]. organization [2445, 2340, 1899, 496, 1190, 512, 1030]. organized [2041, 3308]. organizing [2401, 1823]. organs [2199]. orientation [1135, 261, 636, 1974]. orientational [1399]. oriented [1417, 1975]. Origin [3389, 405, 2443]. Ornstein [2606]. ortho [1630]. ortho-gravitropism [1630]. Orthogonal [2926, 148]. Orthology [2100, 3139]. Oscillating [2682, 2499, 168, 1511, 2290]. oscillation [1351, 1929, 19, 1885]. Oscillations [684, 1298, 2742, 3013, 3286, 881, 112, 2966, 639, 1449, 1160, 570, 1963, 302, 1605, 140, 3527, 1827, 173, 3063, 192, 145, 911, 161, 458, 2045, 2092, 1665, 3516, 2390, 31, 58, 159, 620, 1089, 408, 2229, 1127, 2396, 2918, 19, 1502, 1666]. oscillator [596, 345, 518, 292, 951, 498, 3074]. oscillators [123, 2206, 342, 315, 563, 535, 812, 2270, 177, 168, 345, 444, 293, 495, 1005, 1934, 379, 65, 1267, 648, 1159, 2646]. Oscillatory [545, 977, 1370, 286, 2632, 3622, 2898, 1170, 20, 1030, 1661]. osmotic [579, 3298]. osseointegration [2098]. osteoarthritic [2686]. osteolytic [3230]. other [1962, 1984, 745, 1921, 1043]. Othmer [3002]. our [527, 3233]. Outbreak [2162, 1662, 2295, 3577, 1111]. outbreaks [1320, 3042, 531, 870, 1020]. outcome [1401]. outcomes [1944, 3451]. outer [1233]. output [796, 3034, 3285, 210, 3188]. outward [1495]. outward-rectifier [1495]. ovarian [3131]. ovary [1126]. Overcompensatory [1084]. overflow [3132]. overlap [2196]. overlapping [1926, 62]. overshooting [1486]. oviposition [2076]. Ovulation [494, 1126]. oxide [3440, 1597]. oxygen [2306, 1538, 3309, 297, 2525, 854, 1388, 600, 78, 1344]. oxygenation [233].

p62 [3266]. p62-ubiquitin [3266]. Pace [3189]. pacemaker [659]. Pacemakers [705, 601]. paclitaxel [1473]. page [2083]. pain [3388]. Pair [940, 2013, 1583, 1585, 1143, 713, 994, 1290]. pair-edge [1143]. pair-formation [994, 1290]. Pair-level [1583]. pairing [1048, 1814]. pairs [1402]. pairwise [2945, 2030, 2704]. pancreas [1884]. pancreatic [3144, 1176]. pandemic [1939, 3296]. paper [1924, 1656]. papers [354]. papillomavirus [3303, 3568]. Parabolic [1080, 1008, 1947, 2425, 3458, 2899, 671, 2964, 76]. parabolic-elliptic [2964]. parabolic-parabolic [1947]. paradigm [3009, 2158, 2821]. paradox [2716, 1790, 432]. parallax [2035]. parallel [2937, 803, 1749]. paralog [1616]. Paramecium [235, 196]. Parameter [891, 431, 2408, 3066, 1114, 2215, 757, 1954, 164, 2941, 3198, 3581, 2606, 2151, 2060, 793, 1507, 2601, 3026, 802, 3576, 1822, 908, 1690, 2134, 1067, 2473, 2694, 1729, 2029, 1838, 2826, 2148, 17, 2985, 1495]. Parameterization [3092]. parameterized [2503]. parameters [552, 1540, 3094, 32, 3550, 1552, 2691, 2891, 1451, 806, 1619]. Parametric [1303, 1592, 2516, 3243, 2984, 2175, 3156]. parametrized [2868]. parasite [901, 465, 2128, 1533, 1232, 1163, 2997, 2046]. parasite-host [2128]. Parasites [405, 1260, 510, 1374]. parasitic [242, 756, 753, 731, 1292, 1291, 863]. parasitism [1453, 3502, 1607]. parasitoid [1186, 2668, 974, 697, 1676, 1504, 1574]. Pareto [2238]. parity [2329]. parsimonious [3558, 2859]. Parsimony [2049, 3429, 2906, 2037]. Part [1534, 3274, 1513, 1650, 550, 551, 1461, 1462]. Partial [23, 1113, 2004, 2613, 3038, 1377, 2803, 997, 3441, 3510, 1230, 1243, 2305, 1841]. Partially [1651, 1926, 3393, 1680]. particle [2628, 1902, 904, 1439, 3449, 1487, 3341, 1071, 2502]. particles [1135, 1525, 2394, 1368]. particular [2586]. partition [2314]. Partitioning [2530, 1261, 1667]. partitions [2946, 2846]. partnership [2721, 2395, 952]. partnerships [1023, 3459, 1448]. passage [3060, 3068, 3475]. passages [911]. passing [2704]. passive [432]. past [732]. Patch [1757, 621, 1786, 1366, 2420, 2777, 3069, 2159, 3484, 2406, 2636, 3303, 3543, 3557, 1254, 1945, 3529, 1645, 2960, 3375, 3437, 2171, 3082]. patch-selection [2406]. Patch-size [1757]. patches [1975, 3257, 1295, 1911, 3146]. patchy [2071, 338, 1868, 3261, 2298, 2299, 3206, 3121, 3022, 2283, 2755]. patent [2201]. Path [2987, 3234, 2940, 2783, 3245]. Pathogen [1544, 958, 1503, 2105, 730, 1274, 1097, 1731, 2203, 1955, 374, 2429, 1526, 2350, 2374, 3331]. pathogen-immune [1526]. pathogens [1408, 1767, 2039, 1725, 2497, 3452, 2124, 2459, 1428]. pathological [2700]. pathologies [1755]. paths [1455, 2421]. pathway [2190, 2205, 1849, 2568]. pathways [373, 58, 3585, 1958, 2343]. patient [3300, 3489]. patients [3489, 2990, 2560]. Pattern [592, 355, 81, 964, 2243, 93, 1312, 1986, 2510, 2549, 3594, 1235, 1281, 1282, 2425, 1285, 3004, 1001, 1101, 2527, 2547, 2376, 2194, 803, 102, 1741, 79, 152, 3011, 3487, 2629, 861, 1552, 184, 1873, 161, 893, 287, 1699, 478, 2011, 2913, 615, 1601, 2460, 2233, 1038, 1504, 1028, 321, 3123, 3016]. patterning [3235, 1634, 3010, 198, 2198]. Patterns [1235, 831, 9, 1786, 1775, 123, 1167, 1739, 2870, 976, 2380, 1505, 1862, 3576, 990, 1324, 2504, 3504, 1857, 2797, 3101, 3400, 170, 246, 226, 276, 3002, 1474, 468, 2017, 677, 2667, 750, 2864, 2615, 172, 872, 1418, 3185, 453, 1016, 1267, 3537, 2531, 1897, 1188, 914, 2866, 3282, 3505, 1449, 1032]. Payoff [3311]. payoffs [2180, 3343]. PC12 [2205]. PCR [2731]. PD [3581, 3581, 3423, 3230]. PD-L1 [3423]. PD-L1-PD-1 [3581]. PDE [3328, 3569, 1165, 1778, 3309, 3345, 3270, 3271, 2480, 3481, 3456]. PDEs [3359, 3498]. PDEs/closures [3498]. PDGF [3207]. PDGF-driven [3207]. peaks [3447]. Pedigree [3301]. pedigrees [133, 2088, 3078]. pelagic [2746]. pelagic-benthic [2746]. Peliti [3245]. pelvic [413]. penalization [2213]. people [3580]. peptides [2856]. perceived [2738]. perception [527, 690, 57, 10, 3456, 185]. Perceptive [3450]. percolation [2815, 3365, 3109, 3431]. Perfect [785, 3313]. perfectly [1537]. Performance [598, 2030]. perfused [459]. perfusion [2786]. peri [2098]. peri-implant [2098]. period [894, 2780, 3222, 1605, 1176, 345, 2085, 940, 3462, 2895, 917, 1932, 1330, 870, 1744, 3282, 3283, 3315, 3334, 1020]. period-doubling [1176]. Periodic [1273, 1532, 596, 607, 2828, 1094, 257, 820, 112, 1214, 984, 2396, 2023, 2565, 2784, 3167, 419, 1764, 2060, 2231, 2411, 3027, 350, 3030, 59, 1728, 140, 266, 1606, 517, 649, 192, 516, 2557, 1826, 167, 2766, 515, 1037, 1132, 768, 911, 337, 1315, 1116, 3059, 1509, 3340, 949, 659, 2045, 2895, 1863, 132, 249, 546, 1065, 3278, 2484, 1668, 540, 1155, 367, 83, 426, 201, 2898, 1490, 1268, 392, 16, 3022, 1371, 1383, 1751, 839, 3146, 3555, 289, 174]. Periodical [1236, 3003]. periodically [3020, 301, 1517, 316, 2775, 345, 518, 2239, 213, 1241, 1099]. periodicities [2565]. Periodicity [1672, 1544, 1967, 868, 2504, 1816, 917, 1969, 2370]. périodique [3027]. périodiques [419]. periods [2784, 743, 1482]. peripherally [159]. periphery [2617]. perivascular [2542]. Permanence [3370, 688, 1433, 1452, 1206, 1769, 653, 2475, 2776, 1128]. Permanent [526, 3134]. permeability [536, 2396, 137]. permeable [287, 598]. permissive [1861]. Perron [1341]. persist [3590, 2839]. Persistence [621, 2893, 3401, 2931, 2518, 2790, 515, 603, 243, 965, 382, 2487, 87, 1664, 2026, 2319, 925, 1936, 1578, 2027, 2910, 2228, 1517, 3044, 2105, 1156, 2232, 3389, 1530, 3029, 3048, 1105, 2952, 2989, 365, 3084, 3202, 2486, 1171, 2357, 2053, 2475, 2827, 3606, 1400, 1436, 2070, 1668, 3390, 2292, 2626, 1245, 2488, 3281, 2675, 2624]. Persistent [276, 753, 2390, 994, 1290, 1788, 3181, 327, 3182]. person [482]. personality [2519]. personalized [3288]. perspective [3383, 2050, 1055, 2494, 952, 1054]. pertaining [2890]. Perturbation [1691, 3495, 1035, 3094, 1856, 2925, 2610, 1573, 2818, 2884, 135, 3173]. perturbations [2890, 971, 1002, 2738]. perturbative [3544, 2092]. perturbed [772, 649, 1643, 2158, 1241]. pertussis [2186]. pest [2039, 3345, 2672, 2385, 2378, 1889, 1866, 1496, 2087, 3541, 2889]. pesticide [3070]. pesticides [2087]. Peter [2679, 2825]. Petri [2094]. phages [2027]. phagocytosis [1073]. pharmaceutical [3195]. pharmacodynamic [1456]. pharmacokinetic [2622, 1456]. pharmacokinetic/pharmacodynamic [1456]. pharmacokinetics [2818, 1803, 492]. pharmacological [1265]. pharmacology [2114, 359]. Phase [2606, 1979, 1482, 345, 3295, 393, 3501, 344, 2137, 3482, 1008, 1354, 1239, 1349, 177, 2573, 2697, 3256, 1552, 293, 495, 1005, 2605, 2868, 2282, 1740, 3516, 2271, 2682, 1934, 65, 3403, 2224, 3057, 1870, 2715, 9, 2646, 315]. Phase-field [3501, 2605]. phase-locking [2224, 315]. phase-type [3256]. phase-waves [2646]. phaseless [18]. phases [237, 3064]. phasic [1740]. phenomena [176, 2226, 237, 1815, 3002, 659, 854, 1107]. phenomenological [2647]. phenomenon [3479]. Phenotype [1853, 3020, 2479, 3259]. phenotype-structured [3020, 3259]. phenotypes [66, 1207]. Phenotypic [2819, 1157, 2846, 1017, 2474, 2381, 2644, 512, 1054]. phenotyping [2691]. pheromone [2133]. phosphate [3494]. phospholipid [46]. photobleaching [3198]. photomorphogenesis [27]. photophobic [122]. photoreceptors [3552]. photosynthesis [164]. photosynthetic [2594]. Phototactic [256]. phototrophic [2138]. phototropism [1628]. phyllotaxis [2842]. Phylogenetic [2631, 1570, 2596, 1595, 1720, 3236, 2535, 2601, 2789, 3611, 3115, 1995, 1887, 1862, 3225, 3127, 2954, 2042, 3037, 2969, 2934, 2748, 2436, 3015, 2044, 2651, 3138, 3385, 3215, 1717, 1718, 2189, 2762, 3038, 2906, 2846, 2589, 2482, 2713, 3159, 2534, 3184, 1212, 3307, 3327, 3536, 3218, 3018, 1515, 1602, 2706, 2037, 3017, 3398, 3397, 2273]. phylogenies [2175]. Phylogeny [1950, 2609, 3313]. Physarum [2614]. physical [3053, 2689, 2863]. Physico [557]. Physico-chemical [557]. Physiological [3211, 2219, 1284, 1832, 391, 206]. physiologically [1009, 1888, 2662, 3102, 3001, 3000, 1726, 2116, 692, 1660]. physiology [1733, 104]. phytochrome [3, 27]. Phytophthora [2084]. Phytoplankton [1811, 1362, 991, 2752, 241, 1390, 2371, 3439, 395, 1227, 984, 2484, 925, 1099, 2517, 1557]. Piecewise [1548, 2324, 2169, 1567, 1324, 1672, 1562, 662, 1645, 3233, 1855]. piecewise-deterministic [1855]. piecewise-independent [1645]. Piecewise-linear [1548, 1672, 662]. pigeon [1549]. pinning [2775, 3457, 2198, 2851]. pinpointed [1322]. Pioneer [986]. pioneer-climax [986]. pitchforks [3225]. pituitary [1972]. PK [3230]. PK/PD [3230]. place [3238]. plague [2003]. planarian [3197]. Planck [2549, 1977, 2843, 3288, 631]. plane [636, 665, 685, 1368]. plankton [1211, 2716, 120, 1162, 1197, 1581]. planktonic [926, 2508]. planned [1486]. planning [1234]. Plant [1667, 2904, 2986, 2065, 2356, 1530, 2278, 586, 2445, 516, 439, 852, 1325, 456, 3350, 1630, 2068, 2057, 773, 754, 2691, 1280, 1093, 2863, 1628, 1283, 3579, 2288, 2129]. plant-herbivore [2986, 586, 3350, 2057, 3579]. plant-pollinator-robber [2129]. plant-root [2065]. plant-scale [2691]. plant-seed [2278]. Planta [3396]. plants [1050, 1571, 1300, 2277]. plaque [45, 135]. plaques [3586, 3374, 2492]. plasmid [1273, 209, 983, 1041, 1033, 950]. plasmid-bearing [1273, 983, 1041, 1033]. plasmid-free [1273, 983, 1041, 1033]. plasmids [2886, 602]. Plasmodium [3578]. Plastic [3384]. plasticity [3380, 2474, 3610]. plateau [2868]. Platelet [1749, 927, 2480]. platelet-fibrin [2480]. platelets [2700]. player [2452, 2588, 3343, 3136]. players [851]. plea [1872]. ploidy [3346]. plus [3320]. plus-end-located [3320]. PNP [2924]. Poincaré [134, 885]. Poincaré-transform [134]. point [3160, 3264, 1052, 3066, 1877, 2511, 506, 135, 44, 1935, 125, 3019]. point-source [3066]. point-stimulation [125]. points [1596, 911, 1433, 2561]. pointwise [2621]. Poisson [3058, 2843, 2939, 2238, 366, 1112]. Pol [2595]. polar [226, 773]. polarity [1271, 3286, 3185, 3057]. polarization [3006]. policies [3224, 3538]. policy [312, 529, 3086]. poliomyelitis [2402]. pollen [2961]. pollination [3238]. pollinator [1325, 2129]. polluted [848]. pollution [1125]. Poly [1500]. Poly-chemotherapy [1500]. Pólya [491]. Pólya-like [491]. polycephalum [2614]. polyclonality [1329]. polycystic [1126]. polygenic [858, 781, 790, 865, 597, 764]. polygons [1926]. polyhedra [652]. polymer [2477]. polymer-zeta [2477]. polymerase [1831]. polymerization [2236, 1776, 1222, 981]. polymers [1222, 1984, 3130]. polymorphic [252]. Polymorphism [187, 285, 222, 1228, 1345, 1807, 332, 390, 3389, 39, 190]. polymorphisms [2406, 1078, 106, 857, 2495, 3024, 2369]. polynomials [2872]. polyploid [669]. Polyploidy [155, 154, 627]. polytope [2421, 2541]. polytopes [3037]. polytopic [3263]. Pontryagin [3095, 564, 2498]. pool [701, 535]. poor [2622]. poorly [2712, 3210]. Population [729, 2164, 3368, 948, 510, 1130, 1017, 1858, 3297, 963, 767, 2070, 213, 2292, 1095, 3408, 361, 2922, 3082, 988, 270, 490, 2637, 1096, 1408, 2218, 3088, 2904, 1237, 3380, 2300, 1541, 532, 608, 1551, 2060, 3027, 1311, 1892, 431, 2888, 2312, 2458, 1599, 3590, 3107, 708, 3112, 1674, 2526, 312, 971, 3539, 1944, 1752, 1653, 1737, 564, 724, 428, 1648, 240, 406, 547, 776, 3265, 346, 1966, 1009, 1119, 1349, 818, 2993, 967, 82, 1492, 774, 3502, 2181, 75, 3131, 2232, 309, 40, 357, 1209, 1369, 571, 2159, 503, 2828, 369, 1896, 28, 2547, 918, 200]. population [1270, 831, 1138, 1301, 1888, 2489, 2662, 701, 565, 537, 651, 3530, 2684, 2296, 1817, 1726, 1596, 1286, 473, 1143, 1636, 474, 1052, 1394, 1182, 2102, 250, 2432, 2674, 471, 549, 323, 38, 2952, 156, 3046, 722, 1765, 1157, 349, 1172, 472, 1321, 62, 310, 437, 2613, 3543, 2849, 85, 2373, 860, 800, 64, 2757, 603, 640, 1053, 696, 763, 3348, 2732, 1132, 1204, 1401, 1039, 21, 298, 452, 2977, 3355, 637, 874, 39, 337, 1116, 3435, 1338, 377, 1891, 836, 1872, 2053, 2487, 3141, 3363, 1848, 2345, 2819, 3257, 821, 820, 693, 4, 1293, 1681, 1442, 1835]. population [1921, 1341, 2068, 3111, 247, 344, 843, 2073, 2968, 1932, 2915, 3428, 1838, 1518, 2745, 383, 1770, 2385, 1869, 882, 2121, 884, 1418, 3205, 2921, 805, 3121, 227, 1046, 1093, 2118, 1248, 199, 148, 258, 375, 804, 1131, 553, 863, 3589, 2500, 647, 1884, 3278, 2788, 2642, 2831, 205, 273, 2840, 1935, 2415, 2620, 3130, 2883, 599, 3354, 2472, 3390, 2696, 236, 2626, 744, 1241, 1279, 889, 2813, 1125, 234, 51, 888, 1295, 2982, 946, 2763, 53, 113, 3219, 3623, 936, 969, 736, 838, 845, 910, 923, 1328, 1446, 1911, 692, 2680, 3208, 14, 340, 749, 602, 1159, 1519]. population [1100, 914, 680, 16, 63, 50, 2910, 3022, 3093, 1255, 2173, 2258, 1084, 1463, 1174, 1106, 1386, 1619, 3538, 372, 2755, 427, 3563, 2548, 3377, 3315, 3436, 3334, 2675, 2877, 794]. population-controlled [2173, 2258]. population-size [62]. Populations [3099, 1554, 621, 2893, 1158, 2421, 1269, 3020, 3163, 2387, 470, 2518, 1164, 462, 560, 2356, 2750, 850, 1086, 1875, 1018, 1807, 1808, 1795, 2770, 2469, 2711, 1673, 3005, 2790, 75, 150, 174, 129, 316, 3007, 3328, 2465, 3591, 561, 622, 982, 2828, 2812, 741, 1830, 3277, 343, 784, 3102, 3001, 3000, 202, 3560, 3253, 2270, 3055, 95, 1951, 1643, 280, 106, 126, 439, 253, 2463, 327, 1412, 13, 713, 756, 441, 2486, 2852, 2936, 3075, 23, 1640, 1024, 644, 2313, 2416, 2508, 1374, 2126, 468, 1411, 3193, 2047, 184, 1686, 2316, 1858, 464, 1183, 1835]. populations [2781, 364, 700, 632, 1177, 1194, 433, 3260, 1444, 2839, 2455, 2597, 2567, 3367, 1809, 1485, 2344, 2730, 3205, 1487, 157, 190, 220, 1249, 2354, 2390, 898, 897, 3571, 3162, 2988, 573, 1451, 944, 2196, 2319, 2381, 2997, 958, 3329, 935, 1145, 638, 2209, 194, 3208, 3610, 1685, 2630, 575, 1751, 3074, 931, 631]. porous [2628, 2356, 1568, 2786]. poroviscoelastic [2342]. posedness [2321, 2164, 692, 3535]. position [3291, 1706]. Positional [287]. positioning [2244]. Positive [814, 1611, 2387, 350, 3447, 497, 2279, 1064, 2234, 1226, 1722, 3312]. possibility [2504]. Possible [2594, 890, 1642]. Post [2869, 2365, 2943, 2987, 3214, 3444, 2096, 2629]. Post-buckling [2869]. post-burn [3214, 3444]. post-introduction [2943]. post-transcription [2987]. Post-transcriptional [2365]. post-translation [2987]. post-translational [2096, 2629]. postantibiotic [1837]. posteriori [1072]. potassium [996]. Potential [44, 46, 2617, 1818, 151, 1745, 759, 2346, 834, 2894, 2570, 2794, 43, 2460]. potentially [1549]. potentials [420, 11, 1419]. pour [419, 2556]. power [1976, 720, 766, 3431, 2565]. power-law [1976]. powerful [2706]. powers [451]. Pozón [2736]. practical [1171, 1642, 2148]. practice [1856]. practices [2295]. PRC [601]. pre [3477]. pre-epileptic [3477]. precaution [3624]. Precipitation [3505]. precise [3449]. precision [1690, 3154]. precision-cut [3154]. predation [2911, 3530, 1607, 397, 3317, 3416, 2192, 364, 3613, 508, 633, 2442, 3479]. predation-induced [3530]. Predator [593, 3591, 74, 156, 1318, 2405, 828, 3469, 1980, 301, 3072, 1303, 302, 1086, 1215, 186, 176, 268, 299, 3394, 1513, 1650, 313, 300, 1168, 1129, 2478, 362, 2262, 3077, 798, 3504, 1094, 486, 3029, 649, 681, 846, 1643, 365, 2113, 1045, 1467, 756, 2431, 193, 243, 1845, 272, 2661, 1289, 1379, 905, 3317, 1684, 2192, 184, 1213, 688, 165, 789, 1139, 1790, 2399, 136, 2082, 119, 921, 679, 3310, 1055, 1304, 782, 1592, 866, 3117, 385, 3613, 2642, 2994, 685, 1743, 3354, 508, 3329, 408, 3272, 2001, 432, 1889, 1866, 1512, 2209, 2444, 862, 1922, 2569, 2866, 3282, 1314]. predator [149, 819, 3621, 1305]. Predator-induced [3591]. Predator-mediated [593]. predator-one [313]. predator-prey [3621]. predators [3485, 423, 752, 1310, 545, 2453, 1790, 2399, 1128, 839]. predict [2291, 1148]. predictability [531, 1662]. predicted [1545, 413]. Predicting [2277, 1704, 1803, 1069, 2148]. Prediction [1643, 783, 1705, 3405, 81, 2906, 3352, 1675, 3516, 1703]. predictions [646]. Predictive [2736, 2990]. Preface [1796, 1603]. preference [3389, 1348]. preferences [2756, 3521]. preferential [3291]. pregnancy [3282]. prehension [1894]. Preisach [3204]. Prelude [352]. prepare [2183, 2117]. prepatterns [676]. presence [1503, 1237, 82, 3544, 1024, 821, 2827, 3428, 2031, 283, 2288, 2246, 902]. present [3536]. Preservation [1593, 560]. preserves [706]. press [2738]. pressure [1160, 1892, 2339, 1391, 2451, 2007]. pressures [82]. presynaptic [1721]. prevalence [3600, 3503, 3437, 2455, 1869, 3518]. prevalence-based [3437]. Prevention [2583, 2143]. preventive [2905, 3373]. prey [828, 3469, 1980, 301, 3072, 1303, 302, 1086, 1215, 186, 176, 268, 299, 3394, 1513, 1650, 313, 300, 1129, 2478, 3591, 74, 362, 2262, 3077, 798, 3504, 1094, 486, 3029, 649, 1569, 681, 846, 1643, 1045, 1467, 756, 2431, 193, 243, 1845, 272, 2661, 1289, 1310, 1318, 1379, 1684, 184, 1213, 688, 2453, 165, 789, 1139, 1790, 2399, 136, 364, 2082, 119, 921, 679, 3310, 1055, 1304, 782, 1592, 866, 3117, 385, 3613, 2994, 2405, 685, 1743, 3354, 508, 3329, 408, 3272, 2001, 1245, 432, 1889, 1866, 2116, 1512, 2209, 2444, 862, 1922, 2569, 2866, 3282, 1314, 149, 819, 3621]. prey [1305, 839]. prey-predator [1980, 302, 1086, 3394, 2478, 3504, 1094, 184, 1790, 2399, 136, 685, 3329, 3282, 1314]. preys [3485]. preytaxis [819]. primary [2170, 2335, 2048, 1281]. prime [1520]. primer [3562]. primitive [979]. Primitivity [335]. principal [2174, 388, 2754]. principle [3095, 1489, 730, 564, 1925, 3040, 2498, 824, 825, 3071]. principled [2372]. principles [1913, 142, 1863, 1881]. prion [2513, 2844, 2857]. prions [2317]. priori [2509]. prisoner [815, 2586]. Probabilistic [1857, 548, 2799, 1271, 703, 539, 572]. probabilistic/deterministic [703]. Probabilities [1593, 1018, 1585, 2058, 2812, 3081, 758, 2739, 2331, 2738, 3133, 467, 2811, 876, 2246, 2877]. Probability [2464, 2037, 1455, 875, 943, 2231, 3547, 1715, 2714, 317, 474, 1072, 873, 874, 1654, 505, 1798, 2502, 2982, 2288, 2167, 2554, 902, 1111]. probable [2957]. problem [2904, 939, 3213, 2108, 2097, 3325, 1989, 2469, 908, 40, 2962, 1691, 2663, 2473, 1729, 2365, 3374, 1299, 820, 693, 3430, 2826, 2719, 3312, 2783, 2657, 1894, 499, 2368, 359, 1909]. problem-in [3430]. problems [793, 564, 2147, 2408, 3269, 3569, 3515, 271, 702, 2744, 181, 2145, 2152, 3371, 813, 2754, 2831, 2146, 1038, 2680, 2093, 3456]. procedure [3414, 3113, 2165, 1355]. process [2665, 841, 3200, 904, 3293, 3251, 2633, 343, 2287, 317, 474, 192, 3125, 3081, 3091, 1998, 2461, 514, 2388, 836, 1572, 1448, 731, 3378, 2721, 2868, 1770, 1809, 2940, 550, 551, 952, 694, 3409, 641, 804, 898, 897, 387, 647, 699, 738, 3518, 2811, 163, 3583, 2899, 1935, 1760, 68, 1368, 1145, 638, 588, 612, 1864, 3466, 372, 591, 1031, 2250, 3598, 2877]. processes [1096, 2987, 3468, 2281, 2656, 3264, 729, 1963, 2890, 2815, 2128, 254, 3223, 2470, 3455, 1540, 2058, 2641, 2236, 1056, 2703, 565, 3215, 1561, 1151, 2359, 1412, 2739, 3435, 377, 836, 3065, 1411, 1948, 924, 230, 303, 1262, 506, 1064, 1181, 2344, 2309, 1360, 1420, 2682, 1651, 2829, 876, 3493, 3546, 583, 2587, 935, 20, 1820, 1855]. Processing [84, 1564]. processive [3151]. processus [2656]. produce [3490]. producer [3241, 1981]. producer-grazer [3241, 1981]. producing [1656]. product [3058, 2800]. product-form [3058]. production [319, 2878, 2779, 2364, 3013, 1261, 3506, 1015, 15, 2551, 2876, 400]. productivity [2329]. products [1608, 69]. profile [3346, 2017]. profiles [1044, 2879, 3069, 1811, 2960, 783, 3564]. progeny [317]. program [242]. programs [3497]. Progression [1392, 3392, 3014, 1498, 3423, 3568, 3012, 3221]. progressive [1527]. progressors [1869]. projection [527, 2483, 1636]. prolate [429]. proliferating [464, 909, 575]. Proliferation [1880, 1167, 3014, 1629, 2426, 3153, 1660, 1820]. proliferation/migration [3014]. promoter [2150]. promotes [2989, 3439]. promoting [3560]. proneural [2505]. proof [2665]. proofs [3335]. propagating [2864, 1716]. Propagation [2142, 3499, 3278, 3118, 3158, 3264, 2115, 1989, 771, 2346, 3457, 566, 1120, 834, 70, 443, 682, 2758, 34, 581, 1413, 1323, 2204, 1014]. propagators [2782]. Properties [1118, 828, 930, 1322, 1237, 2930, 299, 2687, 1808, 2638, 2321, 2042, 3499, 1562, 3055, 2205, 2456, 1763, 280, 696, 2364, 809, 464, 1372, 2900, 2940, 2024, 2069, 2280, 3169, 624, 2503, 1355, 2871, 1340, 811, 3289, 3017]. property [506, 3057]. Proportion [2189, 2436]. Proportional [2977]. Proportionate [543, 733]. proportions [1003]. prostate [2557, 2305]. prostheses [827]. Protected [1228, 2406, 2016]. Protection [2661, 2583]. protective [2345]. protein [1336, 2594, 2022, 2104, 2939, 2157, 3506, 1102, 1307, 244, 3198, 1860]. proteins [2336, 3532, 2348, 2844, 2078, 1804]. Proteus [1551, 1134, 1826]. protocell [2585, 557]. protocol [2348]. protocols [3095]. proton [1556]. Protoplasmic [1590]. prototypes [168]. protracted [3592, 2351]. protrusion [2663, 2223]. provide [2136, 2849, 3057]. provided [1889, 1866]. provides [3607]. proving [2994]. PSA [1599]. pseudo [1850]. pseudoknot [2944]. pseudoknots [2083, 1704, 1703]. Pseudomonas [1541]. pseudopalisade [3487]. pseudopalisades [3168]. pseudospectral [1653]. PSSMTS [1706]. psychophysical [42]. Pteroptyx [832]. PU.1 [2005]. PU.1-GATA-1 [2005]. Public [1951, 2879, 2611, 2794, 2583]. Pulled [3513]. pulmonary [2451, 3489, 655, 233, 1405]. pulsatile [596, 2950]. pulsating [2296]. Pulsations [2542]. Pulse [1175, 1752, 2402, 3137, 812, 3540, 992, 3514, 541]. pulse-coupled [3137]. pulses [1225, 380, 1328]. pump [2024, 2069, 3150]. pump-leak [2024, 2069]. pumping [1389]. purine [2367]. purine/pyrimidine [2367]. Purkinje [180]. pursues [798]. pursuit [2035]. pushed [3043, 3513]. pyloric [1740]. pyrimidine [2367]. pyrite [2293]. pyrosequencing [2279].

qPCR [2868]. quadratic [1798]. quadruped [1281, 1282]. Quadruplex [2266]. Qualitative [1164, 1769, 119, 921, 2197, 2321, 3332, 90, 441, 2331, 3350, 1139, 56, 3092, 51]. quality [1987, 1203]. quantification [2591, 2200, 3526, 3012]. quantify [2816]. Quantifying [2409, 3611, 2906, 2685, 1430]. Quantitative [3091, 840, 2892, 2300, 2219, 955, 1499, 1052, 520, 1078, 745, 1831, 1416, 3561, 2691, 1580, 2221, 859, 399, 1970, 512, 3198]. quantities [1012]. quarantine [2580, 3466]. quarantines [3179]. quartet [3338, 2706]. quartet-based [3338]. quartets [2954, 1717]. Quasi [995, 2313, 3531, 3148, 1150, 1434, 1269, 2176, 3144, 2465, 3319, 3510, 2669, 1238, 1490]. quasi-bilayers [2669]. quasi-equilibrium [3144]. Quasi-neutral [3531]. quasi-stationarity [1269, 2176, 2465]. Quasi-stationary [995]. Quasi-steady [1434, 3510, 1490]. Quasi-steady-state [3148, 1150, 3319, 1238]. quasispecies [2644, 2871]. Queen [2431]. quenching [2594]. quest [3076]. questions [2088]. queues [3330]. Quiescence [2111, 1588, 1694, 800]. Quorum [3271, 1541, 1501, 1612, 1403]. Quorum-sensing [3271]. quota [2018].

R [3042, 1214, 2671, 425]. rôle [492]. rabies [3373]. radially [1342].

Radiation [1309, 2505, 3592, 625, 3367, 2666, 1802, 2172]. radiation-induced [3592, 2666, 2172]. radiosensitivity [2836]. radiotherapy [2108, 2097, 2407, 1798]. raffles [1199]. raids [2514]. rainbow [1213]. rainfall [2637]. Raising [451]. Random [1966, 24, 366, 247, 2984, 1229, 1191, 932, 1253, 2970, 223, 2176, 2208, 2281, 2556, 2656, 2726, 3425, 2890, 1926, 2978, 1700, 371, 3225, 2619, 922, 2588, 2853, 2528, 535, 2919, 3603, 437, 696, 2852, 3305, 21, 1076, 2623, 2155, 3343, 3573, 1835, 54, 2429, 2809, 48, 1924, 3589, 166, 1309, 687, 737, 742, 3526, 923, 1710]. random-walk [1309]. Randomization [3353]. randomly [1417, 3200, 310, 1130, 35, 2788, 1886, 2272]. randomness [2409, 1812]. Range [2221, 2737, 1739, 3400, 252, 3595, 2851, 2719, 3537]. ranges [1934]. rank [3392, 3429]. ranked [3127, 2957]. Ranvier [3402]. Rapid [1444, 1703, 1739, 1684, 3536]. rapide [3027]. Rapidly [884]. rare [2762, 1021, 1377]. rarefaction [2816]. rat [162, 2249]. ratchet [1343, 2104, 942]. ratcheting [2104]. ratchets [2272]. rate [2606, 608, 2457, 2581, 3199, 186, 268, 299, 1009, 1437, 3026, 2470, 3036, 2741, 1208, 3077, 2149, 1470, 2134, 1069, 2933, 1467, 3483, 2613, 3085, 920, 721, 3031, 3435, 1104, 1263, 3108, 2165, 553, 2148, 1973, 3376, 233, 728, 2709, 1328, 78, 340, 3433, 2167]. Rates [347, 875, 943, 1670, 2882, 2991, 3426, 2226, 316, 1606, 1154, 3406, 253, 1762, 751, 398, 873, 1315, 1457, 569, 650, 2916, 2175, 1248, 573, 2446, 2539, 909, 744, 2021, 1820]. rates-across-sites [2175]. ratio [1303, 2842, 1445, 784, 1359, 1289, 1310, 3192, 487, 2929, 1139, 1743, 1512, 3100, 3472, 1305, 3417, 1430]. ratio-dependent [1303, 1289, 1310, 2929, 1139, 1743, 1512, 1305]. rational [1234, 2579]. ratios [3030, 716, 1291]. re [1348, 2160]. re-entry [2160]. re-examined [1348]. reachability [1157]. React [2311]. reacting [34]. Reaction [3235, 2051, 2693, 2280, 1355, 2220, 2943, 1361, 2959, 3058, 3158, 1900, 592, 2780, 1892, 1902, 3250, 1439, 3426, 330, 967, 2441, 2435, 3576, 1750, 3383, 1324, 265, 493, 1154, 1827, 2052, 2684, 2509, 3499, 2296, 409, 286, 2096, 716, 3145, 2303, 2473, 667, 1068, 1045, 666, 3051, 3011, 2897, 3522, 3183, 1729, 429, 477, 3157, 1626, 2475, 2759, 2566, 2074, 33, 88, 1873, 745, 2275, 757, 748, 3286, 1859, 1843, 3119, 3609, 1187, 287, 1932, 3461, 3463, 977, 2995, 1882, 2362, 410, 2615, 2254, 2800, 2953, 453, 2666, 3510, 2484, 135, 2785, 2768, 179, 1038, 2600, 1578]. reaction [3174, 1927, 877, 1150, 2961, 124, 3342, 1655, 3245, 1990, 3442, 3424, 2910, 3093, 3299, 1736, 1661, 1964, 3315, 2066]. reaction-advection-diffusion [3299]. Reaction-diffusion [3235, 3426, 3522, 3286, 2995, 2961, 3315]. reaction-diffusion-advection [3250]. reactions [1217, 1683, 2887, 1831, 11, 3413, 244, 1434, 1490, 3198]. reactive [3534]. real [3088, 2673, 2731, 849, 1663]. real-time [2673]. real-valued [3088]. realistic [417, 161, 1330]. realize [1157]. reanalysis [2318]. rearrangement [3517, 3060, 3314]. rebound [2264, 2655]. Recall [660, 770]. Receptor [2735, 960, 2114, 2297, 2676, 3229, 303, 2196, 1409, 1047, 1860]. receptor-ligand [2297]. receptor-mediated [1047]. receptors [1044, 370, 1860]. recessive [2804, 1962, 2659]. Reciprocal [3023]. recognition [321]. recognize [1768]. recombinants [463]. Recombination [669, 3160, 1288, 2261, 1895, 2530, 3483, 1910, 1042, 1940, 1438, 1035, 2088, 1871]. recombining [2472]. reconciliation [3045, 2907]. reconciliations [2438]. Reconciling [2976]. reconstitution [1406]. reconstruct [3104, 3473]. reconstructed [2351]. reconstructible [2969]. Reconstructing [2820, 2088, 2651, 1905]. Reconstruction [2708, 2094, 3284]. recordings [1495]. Recovering [2791]. recovery [2150, 2804, 3406, 2884, 2165, 767, 3604, 358, 3198]. recovery-relapse [358]. recruitment [3199, 75, 343, 471, 1539, 509, 1751, 528, 1084, 2833]. rectifier [1495]. rectum [1597]. recurrence [2823, 3542]. recurrences [408]. recurrent [991, 887, 475, 1730, 1999, 1707, 531]. recursion [1920, 1669]. recursions [1785]. recycling [319, 769, 919, 1162, 1227, 2241, 925]. red [1073, 86, 1436, 2431]. reduce [1708]. Reduced [3154, 3144, 2845, 1192, 2385]. reduces [2137]. reducible [418, 481, 480, 581, 583]. Reducing [1216]. Reduction [3246, 3234, 527, 3000, 154, 2855, 1553, 1612, 3148, 97, 2600, 1238, 1662, 2305, 2358, 2715, 2250]. reductions [3319, 1789, 1861]. redundancy [3475]. Reed [583]. reentry [1679]. reference [2496, 2186]. refinement [264]. Refining [1470, 1992]. reflected [2542]. reflection [1052]. reflections [2451]. reflex [58]. reflexes [435]. Reflux [1464]. refractoriness [2256]. refractory [1298, 1482]. regard [2139]. regeneration [3197, 864]. regime [2022, 297, 3354, 602, 2935]. regimens [1837, 1642]. regimes [319, 1556]. region [498]. regional [2402, 905]. Regions [2880, 186, 176, 3348, 3618, 1064, 408, 3321]. regression [1875, 2216]. Regular [683, 548, 706, 707, 778, 228, 3186]. Regularity [1335]. regularly [2316]. Regulated [54, 2150, 796, 1935, 892, 3096]. Regulation [701, 337, 2863, 817, 1971, 3151, 2692, 2767, 3440, 3548, 2511, 3216, 2266, 2365, 2286, 924, 1192, 2966, 2913, 2148, 205, 1958, 194, 427]. regulator [2364]. regulators [882]. Regulatory [1548, 1689, 1822, 3025, 2423, 1567, 3567, 3330, 1929, 2331, 3128, 1937, 3263, 2290, 3556, 2428, 2974, 2918, 3150]. rehabilitation [2979]. reinfection [3318, 2186]. reinforce [3581]. relapse [3164, 358]. related [223, 377, 836, 3261, 1521, 1340, 3538, 270]. Relatedness [3078]. Relating [1869]. relation [698, 2820, 2949, 1661]. Relations [459, 3357, 2100, 2660, 456, 989]. relationship [1992, 210, 987, 3019]. Relationships [744, 457, 3591, 264, 334, 1416, 508, 2704]. Relative [3437]. relatives [333]. Relaxation [609, 168, 899, 1005, 951, 1461, 1462]. relaxed [2348]. release [552, 3440, 397, 160, 1413, 3283, 3436, 3334]. releases [3080]. releasing [1579, 1409, 3334]. relevant [2427]. relief [3388]. Remarks [1429, 489, 3018]. remodeling [3230]. remodelling [1799]. removal [3594, 1670, 1208, 2155]. removals [516, 3486]. removed [1023, 2422]. Renal [413, 702, 780, 1372, 1597]. renewal [731, 2900, 588]. renewing [3265]. renin [3443]. renin-angiotensin [3443]. reorganization [2605]. reorientation [1540]. repair [209, 3592, 625]. repair-misrepair [625]. repeated [1025, 2586]. Repetitive [125, 772]. replacement [403]. replanting [1081]. replicas [2025]. replicates [1421]. replicating [3181]. replication [1190, 997, 1181, 158, 747, 1516, 1820]. replication-competent [1516]. replicator [1641, 3163, 2923, 3328, 3134, 1424, 2771, 653, 2664, 973, 998, 3019, 2023]. reply [3619]. reporting [2302, 2649]. represent [1372]. representable [2660]. Representation [178, 2421, 1276, 1402, 265, 3102, 3001, 2856, 2353, 2103, 2920, 3534]. representations [2621, 3364]. represented [369, 298]. Representing [3558, 16]. repressilator [1623]. repression [556, 1678, 488, 546, 881, 639]. reproduce [510]. reproducing [142]. reproduction [1939, 2003, 2208, 3030, 1239, 1571, 784, 2636, 2050, 2126, 2942, 3529, 2155, 2917, 1291, 1973, 3156, 2831, 1693, 2125, 2140, 2185, 2403, 912, 2444, 3093, 3335, 3417, 1828, 1430]. reproductive [3455, 150, 174, 129, 2203, 3504, 1834, 1543, 1582, 646, 1017, 3345, 1451]. repulsion [3067, 2719]. required [2364, 1613]. requirements [3294, 2080]. Research [485, 1698]. Reserve [1991]. reserves [1173]. reservoir [3578, 1836]. resetting [1979, 2137]. reshuffling [2963]. Resident [3103, 1488, 1944]. Resident-invader [3103, 1488]. Residual [1480, 1070, 2087]. Resistance [2575, 1216, 2701, 1896, 3452, 3542, 3561, 3511, 1941, 2046]. resistant [1260, 2259, 3524]. resolution [2379]. resolving [2806]. Resonance [1764, 1741, 1677, 1213, 1201, 269]. Resource [1588, 3481, 801, 1013, 1917, 1026, 1105, 2018, 2463, 3384, 2897, 3522, 295, 2036, 2245, 3545, 3614, 1075, 1404, 3408, 979, 3071, 3124, 2698, 799, 2865]. Resource-based [1588, 1013, 1917]. resource-consumer [1026]. resource-dependent [3408]. resource-limited [2463]. resources [1576, 1537, 1928, 3462, 1682, 2794, 2608, 1125, 3126, 2612]. respect [2859]. respiration [972, 1755, 78]. respirationssystems [61]. respiratory [1246, 1247, 1497, 192, 2933, 3287, 3309, 61, 3376]. respond [2318]. Response [992, 2632, 2898, 1959, 1537, 1136, 3072, 3402, 2832, 59, 1129, 1454, 1955, 1852, 1834, 1979, 3191, 1440, 2352, 2297, 2113, 2661, 514, 1318, 293, 3204, 2045, 2010, 550, 551, 2393, 3570, 1257, 3613, 984, 2397, 1802, 1743, 2769, 3579, 2229, 1516, 2048, 2179, 3420, 3621, 1502, 1982, 3290]. responses [3552, 3089, 3393, 3422, 2593, 320, 2806, 1981, 27, 1732]. responsive [3275]. responsiveness [1102]. rest [3177, 3376, 3019]. resting [3216]. restricted [2540]. restriction [1877]. result [868, 565, 1608, 879, 2916, 2385]. resulting [2532]. Results [707, 1, 715, 3351, 916, 300, 369, 3081, 471, 2636, 1078, 3385, 452, 786, 2400, 2245, 179, 885]. reticulation [1905]. retina [59, 1006]. Retinal [2525]. retreat [3090, 2173, 2258]. retrieval [3434]. retroviral [2560]. return [3580]. return-to-home [3580]. returns [2498]. revealing [3221]. reveals [2364, 3207]. reversal [1851]. reversals [1173]. reverse [2542, 2672]. Reversible [2474]. Reversing [2075]. review [2139, 2041, 2801, 1773, 3287, 2740]. revised [2614]. Revision [1677]. revisited [2333, 2304, 1235, 1888, 2662, 1459, 1199, 1080]. Revisiting [3236, 2918]. reward [1951]. rewiring [2162]. Reynolds [2523]. rheological [3252]. rheumatoid [3056]. rhythm [1740]. Rhythmic [1590, 2137]. rhythms [32, 2092, 648, 3247]. ribozyme [1359]. ribozyme-substrate [1359]. Rich [1865, 1310, 3554, 2561]. richness [3062, 2816]. Riemannian [1690]. right [494]. rigid [3159]. rigidity [57, 1579]. Rigorous [1968, 2168, 3113, 2404, 1036]. Ring [1250, 1679, 3049, 2244]. rings [563, 977]. risk [3262, 1477, 2548]. risks [591]. river [3048, 2357, 2070, 2292, 2922]. riverine [2000]. RNA [2193, 1702, 2314, 1402, 1700, 1697, 2059, 2083, 2747, 3039, 2225, 2477, 2009, 2653, 2944, 2595, 1698, 1704, 467, 1701, 2334, 1707, 2343, 1708, 1703]. RNABC [1708]. RNAs [1705]. road [3107]. robber [2129]. robots [2765]. Robust [2228, 2294, 2511, 1756, 2776, 2646, 2136, 3351, 3594, 1214, 3585]. Robustness [1144, 2587, 1822, 3369, 3607, 3311, 2555, 2224]. rock [1656]. rod [3552, 2869, 1222, 1620]. rod-like [1222]. rodent [28]. rodents [319]. roguing [1081]. Role [2105, 3399, 413, 2637, 1450, 996, 3264, 991, 2594, 3532, 1989, 1354, 975, 1173, 743, 2340, 3033, 1390, 3084, 398, 2317, 1658, 2124, 2187, 318, 424, 1436, 1082, 2525, 3130, 519, 728, 2227, 2259, 3604, 2792, 2554, 3221]. roles [1199]. root [2065, 2691, 1015, 411, 2392]. rooted [1942, 1887, 2954, 3015, 3407, 3038]. rooted-tree [3407]. Rosenzweig [3368, 2855, 2769]. Ross [3022]. Rotating [1618, 2574, 306]. rotavirus [1614]. rotifers [1239, 1349]. routes [1355]. RSV [2933]. RTK [3585]. RTKs [2114]. rules [1826, 2689, 2077]. run [3007, 1664, 2568]. run-and-tumble [2568]. runner [3189]. running [938, 1091]. rust [3601].

s [2165, 1766, 3042, 587, 1214, 2671, 425]. S-I-R [3042, 1214, 2671, 425]. S-I-S [1766, 587]. S. [14]. saccadic [1545]. Saccharomyces [589]. saccular [1367]. Sackin [3385]. Saharan [3476]. salamanders [646]. Salmonella [1971]. salmonid [2344]. same [2716, 2921]. Sample [2940, 2882, 2991, 3338, 1724, 2547, 1067, 1131, 2813]. sampling [3041, 2025, 808, 2697, 21, 491, 637, 1704, 1360, 1420, 1701, 1854, 3598]. Sanford [3619]. SARS [3476, 3507]. satellite [2440]. satiation [514]. saturated [2495, 3024, 1433, 3339, 3412, 3604, 3436]. saturating [920]. savanna [2349, 2639]. saxatilis [853]. scaffold [1681, 2191]. scalar [2574]. Scale [2515, 2222, 2870, 3177, 3039, 3319, 2767, 2566, 1859, 2691, 3244, 2921, 2175, 2635, 3289, 3577]. scale-free [3039, 2635]. scaled [1619]. scales [1812, 3054, 3561, 2994, 2587]. Scaling [3357, 951, 3565, 1300, 3161, 3092]. scar [589]. scenarios [3546]. schedule [2985]. schedules [2505, 912]. Scheduling [3179, 2407, 1447]. scheme [3347, 2602, 3512, 3231]. schemes [2697]. schistosome [1296, 2046]. schistosomes [130]. Schnakenberg [1474, 1997]. schools [2500, 1973]. scissors [1656]. scleral [2604]. sclerosis [2667]. SCN1A} [3477]. scope [1219, 3534]. scoring [1706]. scoticus [1436]. scramble [1063]. screening [3470]. SDE [3088]. Sea [1759, 2233]. Sea-lice [1759]. seabird [2828, 2158]. Search [2540, 3077, 170, 3180, 3231, 1987, 1203, 1864, 2891]. searchers [1987]. searching [1180]. searchlight [837]. seashells [1873]. season [1260]. Seasonal [2237, 2053, 1730, 2209, 939, 417, 2828, 2579, 3227, 3305, 1993, 3410, 236, 1446, 1510, 3436]. seasonality [1598, 1939, 2003, 2780, 3502, 3101, 2866, 3299]. seasonally [2707, 2914, 221, 2158, 2026, 531, 870]. seasons [860]. Second [1774, 790, 865, 3330, 636, 87, 408, 1987, 747]. Second-order [790, 865, 1987]. secondary [2314, 1402, 1700, 2059, 2083, 2225, 2335, 2477, 1704, 467, 1701, 2389, 1282]. secretion [3144]. secretory [3148]. section [1417]. sections [673]. sedentary [462]. Seed [3530, 338, 2930, 3200, 3172, 2278, 3400, 2835, 2068]. seeding [3075]. Segel [1947, 1483, 3067, 1083, 1342, 3391, 2964]. segmental [342]. segmentation [2717, 1885]. segregating [1131, 1121, 1340]. Segregation [227, 1106, 560, 1995, 905, 787, 3602, 215, 155]. Seidel [1738]. Seidel-Herzel [1738]. SEIQR [1829]. SEIR [3407, 949, 460]. SEIRD [3441]. SEIRS [1085, 2895, 2941, 2307]. seizures [3492]. selected [1625]. Selecting [1022, 1066]. Selection [1991, 36, 439, 25, 838, 833, 845, 2881, 1398, 347, 1637, 2799, 2761, 841, 1136, 1688, 2333, 3249, 724, 955, 1499, 1807, 1808, 2563, 346, 1646, 657, 82, 1087, 807, 2923, 3328, 1445, 1061, 1425, 741, 2722, 537, 2406, 38, 913, 961, 280, 1078, 2447, 2076, 208, 415, 7, 67, 781, 790, 865, 1024, 562, 39, 2238, 377, 26, 1965, 80, 89, 142, 1848, 24, 178, 225, 252, 809, 1092, 2861, 1654, 1835, 2781, 538, 2968, 709, 929, 1518, 628, 2466, 3133, 2498, 2625, 2730, 121, 187, 285, 190, 1046, 1093, 1178, 1659, 6, 1940, 166, 615]. selection [1451, 2718, 205, 3381, 1243, 605, 2696, 2688, 1566, 1881, 989, 1245, 2982, 1927, 2763, 1035, 910, 923, 326, 764, 2847, 1054, 2007, 427, 1207, 2963, 931, 270, 22, 1495]. selection-balance [955]. Selective [1962, 3388, 817, 428, 1910, 1680, 1999, 934]. selectively [1421]. Self [1530, 1692, 1030, 2904, 1947, 1996, 900, 2425, 3265, 2401, 2041, 2445, 1899, 1227, 142, 2345, 3308, 3461, 1451, 314, 1258, 1661]. self-consistent [1258, 1661]. self-generating [3461]. self-incompatibility [2904, 1996, 900]. Self-optimization [1692]. self-organised [2425]. Self-organization [1030, 2445, 1899]. self-organized [2041, 3308]. self-organizing [2401]. self-protective [2345]. self-renewing [3265]. self-reproducing [142]. self-shading [1227, 314]. self-similar [1947]. Self-thinning [1530]. Selfing [144, 497, 1243]. selfish [3311]. selfreplication [279]. selfreproductive [89]. semelparous [1813, 3003, 1017, 1686, 3035, 1485, 3208]. Semi [1709, 1523, 1175, 2797, 3101, 276, 3553, 952, 2531, 3505]. Semi-algebraic [1709]. semi-arid [2797, 3101, 2531, 3505]. semi-directed [3553]. semi-discrete [1175, 276]. semi-Markov [952]. Semigroup [3463, 817, 3461, 3312]. Semigroups [1611, 1722]. semilinear [1990]. Senescence [1896, 3550]. senescent [3399]. sensing [1541, 1501, 3006, 3053, 1612, 3271, 1403, 2384]. sensitive [2843, 3524, 3335, 3248]. Sensitivity [248, 3402, 2843, 2769, 2337, 3211, 500, 3468, 592, 1507, 120, 2400, 3345, 3244, 2195, 3221]. sensory [999]. Separable [547, 680, 3414]. Separation [3172, 1524]. seq [2595]. Sequence [1611, 1809, 1088, 2090, 170, 670, 2339, 1722]. sequences [3292, 1397, 1700, 1124, 310, 1561, 3284, 3106, 1131, 2163, 1438, 2920, 2565]. sequencing [2161, 3509, 1205, 2188]. Sequential [976, 1203, 1987, 1864, 3598]. serial [2508]. series [891, 3047, 2914, 2094, 479, 1838, 1409]. serological [1069]. serostatus [3488]. serostatus-dependent [3488]. serotype [3470]. serotypes [1375]. serovar [1971]. Serratia [1221]. server [3330]. Set [2630, 2722, 2511, 1584, 699, 1557]. Set-membership [2630]. set-point [2511]. set-valued [2722]. sets [8, 1425, 18, 3327, 48, 2973, 2587, 542]. setting [2369]. settling [1685]. several [332, 390, 2212, 3569, 232, 669, 688, 3257, 273, 627]. severity [3054, 3512, 3062]. Sex [1445, 1652, 1851, 2674, 2291, 23, 487, 835, 1042, 6, 457, 888, 2680, 3208, 5, 49, 3273]. sex-age-structured [888]. sex-dependent [1042]. sex-linked [6]. sex-ratio [487]. sex-reversal [1851]. Sex-structured [1652, 2674]. sexes [658]. sexual [2333, 1239, 1349, 3277, 1348, 1105, 2636, 1157, 1448, 2834]. sexually [1364, 1097, 675, 1428, 3460, 3273]. sexually-transmitted [1428]. shading [926, 1227, 314]. shadow [3457]. shallow [2746]. Shannon [3559]. Shao [3236]. shape [3468, 2879, 2752, 3105, 3176, 3304, 745, 57, 813, 2792, 3290]. shaped [2924, 3591, 2177, 2477, 1761, 601]. shapes [1417, 3199, 3447, 3385, 3386, 2877]. shaping [3248]. Shapley [2436, 3015, 1718, 2189, 3018]. shared [1731, 3615]. sharing [1347]. shark [131]. sharp [1000]. Shaw [2719]. Shear [3440, 1135, 241, 1938, 2213, 2564]. sheath [3402]. Sheldon [2716]. shellings [2042]. shield [3399]. shift [65, 3354]. shifting [3606]. shifts [3516, 1557]. shock [1804]. short [732, 3379, 2719]. short-lived [3379]. short-range [2719]. shortest [2791, 2783]. should [2318, 2477]. show [2731]. shrimp [1319]. shrinkage [3482]. shrinking [3212]. shuffles [1700]. SI [2741, 2395, 3259]. sib [906]. siblings [2813]. sides [2716]. sigmoidal [1245]. sign [2504]. Signal [1989, 70, 2114, 1978, 2207, 2562]. Signaling [2282, 3197, 1782, 2333, 2190, 2286, 1908, 2064, 3396, 2146, 2079, 2101, 3585, 1079, 2338]. signalling [2297, 2909, 1573, 2198, 1849, 3329, 2054, 1449]. signals [2279, 3509, 84]. significance [520, 235, 3307]. significant [3471]. signs [3354]. silico [1784]. similar [1947, 3103, 271, 1488]. similarity [2796, 3038]. Simple [2333, 962, 1170, 1322, 339, 3405, 103, 2255, 1689, 1575, 28, 1530, 1147, 3048, 1133, 409, 286, 2574, 167, 1363, 1382, 1390, 177, 1090, 1407, 2909, 1821, 435, 2073, 880, 180, 125, 1163, 2862, 574, 179, 1958, 457, 1279, 1368, 648, 3563, 2561, 1242]. simplex [2634]. simplices [2491]. simplicial [3364]. simplification [351, 3230]. Simplified [388, 861]. simplifies [1850]. simulate [2628, 685, 3492]. simulated [2387, 3223, 1351]. Simulating [3300, 2815, 1584, 2912, 3348, 2062]. Simulation [2008, 1272, 2948, 723, 828, 1953, 1285, 990, 1384, 146, 308, 2038, 33, 2654, 3498, 1793, 827, 3587, 61, 2118, 2967, 1754, 1846, 1803, 2492, 2216]. simulations [3211, 1275, 1797, 771, 2584, 2106, 2362, 2223, 1712]. Simultaneous [849, 2725, 2962, 1181, 1258]. Sinapis [3]. since [2122, 1121]. Single [2261, 2712, 1686, 970, 1871, 801, 1361, 417, 2136, 3032, 590, 1061, 1425, 94, 140, 2467, 705, 2153, 846, 1777, 780, 1644, 208, 714, 2018, 3085, 398, 3522, 1861, 1501, 1452, 3470, 1654, 2062, 945, 2632, 360, 2689, 2484, 3329, 1844, 3408, 1287, 1744, 16, 1207, 3377, 270]. single- [3408]. single-cell [2136]. Single-class [1686]. Single-crossover [2261, 1871]. single-joint [3032, 1644]. Single-locus [970, 1061, 1425, 208, 1744]. single-serotype [3470]. single-species [2153, 846, 1452, 945, 1287]. Singular [1573, 419, 2308, 698, 3237, 3094, 2496, 1002, 1856, 2727, 1299, 3157, 2925, 2818, 135, 3173]. singularities [2607, 2375]. Singularity [2544, 2622]. singularly [772]. singulière [419]. sink [3590, 2594, 44]. sinking [1811]. SIQ [2927]. SIR [2228, 2665, 2237, 2457, 2726, 2905, 2707, 1004, 3222, 3036, 3446, 2914, 3499, 1734, 2919, 2896, 3407, 3616, 2623, 3195, 2353, 3204, 949, 3573, 2809, 2826, 3352, 1924, 2921, 2158, 2500, 2650, 1841, 1710, 2704, 2833]. SIR-models [2650]. SIRC [3525]. SIRS [3228, 569, 3339, 3412, 2186]. SIS [1786, 1923, 1366, 2411, 2556, 2575, 2787, 1907, 3069, 2790, 2528, 3165, 1039, 1220, 3306, 2162, 3194, 3529, 2610, 2960, 2031, 3564, 3064, 3497, 2571, 1242]. SIS-household-epidemic [1907]. SIT [2903]. Site [3365, 567, 2815, 2593, 2014, 758, 2516, 1189, 3256, 2616, 628]. Site-bond [3365]. site-specific [2516]. site-to-site [2593]. sites [1753, 2104, 2981, 2103, 2175, 1131, 1121, 1340]. situ [1422]. Size [676, 1408, 2218, 1362, 2176, 1757, 853, 729, 2640, 312, 1957, 2407, 428, 1648, 776, 3565, 402, 3251, 2465, 2716, 1976, 798, 2633, 1530, 450, 476, 537, 3105, 3176, 1286, 317, 3161, 3046, 2835, 1998, 62, 439, 3543, 2849, 589, 852, 3616, 554, 1039, 21, 452, 2313, 2416, 1116, 1891, 3597, 2047, 1858, 1457, 3375, 628, 2385, 3441, 3587, 3614, 553, 647, 2559, 481, 2620, 1082, 3390, 2696, 583, 1550, 2816, 1062, 3623, 539, 2471, 2639, 3563, 988]. size- [2559, 2816]. size-dependent [1457]. size-specific [312]. size-spectrum [2716, 3597]. size-structured [1362, 853, 3251, 1530, 1286, 1116, 3614, 1062, 2639]. sizes [882, 2675]. skeletal [1644]. skewness [329]. skin [960, 1001, 2118, 783, 941, 3341, 3012]. sleep [1867, 648]. sleep/wake [1867]. slice [3154]. sliding [2937]. slightly [317, 474]. Slime [1590, 1661]. slip [717]. Slow [2465, 3027, 1154, 3319, 3344, 911, 1556, 598, 2462, 2964, 3057, 2204]. Slow-fast [2465]. Slowest [3475]. slowing [469]. slowly [262]. Slyozov [3587]. SMAD [2190]. Small [971, 1437, 870, 367, 1253, 319, 1135, 2452, 2882, 2991, 1938, 644, 3374, 1334, 2819, 1705, 795, 2021, 2540, 3324]. small-jumps [2819]. smaller [2440]. smooth [2628, 2134, 2939, 2567, 2889, 2734]. smoothed [3341]. snail [2046]. snail-schistosome-human [2046]. snapshots [1165]. SNR [2163]. social [3445, 3190, 3161, 1419, 3570, 2143]. social-insect [3570]. socially [3178]. Society [1594, 1621]. sodium [114, 147, 115]. soft [2476, 3099, 3193]. soil [2392]. soil-root [2392]. sojourn [2284]. sojourn-times [2284]. Sokal [3236]. solid [1198, 915, 1285, 1797, 57, 3014, 1846]. solitary [1195, 336]. Soliton [1655]. solute [2130, 2195]. Solution [617, 1461, 1112, 490, 930, 3494, 2888, 2108, 2097, 529, 619, 1195, 320, 830, 761, 899, 401, 1226, 1423, 1556, 3114, 2785, 2392, 1870, 201, 936, 969, 392, 16, 993]. Solutions [419, 76, 1273, 1618, 1118, 1532, 350, 1599, 1947, 3120, 724, 1472, 489, 2509, 2914, 416, 1146, 2574, 449, 1231, 1691, 702, 2766, 1256, 257, 1587, 3407, 1342, 1379, 337, 1116, 395, 1254, 1334, 325, 1005, 753, 820, 3128, 748, 659, 1842, 2068, 3119, 1187, 407, 814, 1906, 132, 249, 546, 1214, 927, 259, 2994, 1108, 273, 418, 480, 37, 540, 83, 20, 1385, 2214, 541, 1290, 2833]. solutivory [992]. solvable [89, 2532]. solvation [1988]. Solving [3160, 2147, 1626, 3245, 810, 181, 2685, 3233]. somatic [3131]. Some [930, 82, 300, 489, 1078, 816, 142, 1400, 179, 989, 274, 1913, 706, 1167, 1397, 299, 412, 2470, 661, 473, 3055, 2557, 2697, 2850, 688, 645, 87, 1476, 550, 952, 228, 2811, 3564, 876, 1011, 614, 3018, 2088, 3281]. somitogenesis [1565]. song [2230]. Sorption [1462]. sorting [308]. source [3066, 135, 44, 2964]. sources [1812, 3530]. sous [2656]. sous-critiques [2656]. South [1758]. southern [2291]. soybeans [1462]. space [1088, 1739, 1246, 1247, 2724, 2438, 1822, 212, 2748, 2461, 696, 763, 3355, 70, 3106, 718, 3356, 1015, 3278, 2840, 1935, 1233, 2626, 1927, 372]. space-dependent [372]. space-jump [2724]. space-time [1935]. Spaces [2482, 1611, 1968, 2759, 3553, 1722, 1853, 2871]. Spain [2736]. sparse [3382, 2926]. sparsely [3401]. sparsification [2477]. Spatial [1786, 123, 1492, 1351, 2190, 2376, 2507, 3521, 3453, 3011, 2629, 184, 2699, 198, 750, 3371, 172, 215, 1089, 3152, 217, 3366, 914, 2881, 3158, 2356, 406, 556, 2328, 818, 1156, 2993, 330, 1540, 171, 3592, 2467, 2139, 1222, 2340, 329, 3254, 2018, 13, 1117, 276, 3439, 768, 3513, 820, 2242, 693, 2347, 3602, 3070, 488, 224, 2198, 453, 2650, 3564, 481, 2233, 2964, 2349, 3143, 3333, 1782, 2428, 101, 3100, 3012, 3535, 3282, 358, 84, 794]. spatial-temporal [768]. Spatially [1026, 2361, 1443, 3291, 1541, 1169, 2591, 1673, 2911, 2507, 3105, 3176, 2102, 3085, 3453, 398, 2732, 1640, 3618, 3059, 1254, 1177, 2221, 863, 2754, 944, 3232, 194, 3208, 3124, 19, 1449, 1751, 2360]. spatially-constrained [3618]. Spatially-explicit [1443]. spatially-structured [1541, 3124]. Spatio [1285, 1016, 3274, 1985, 1542, 1777, 458, 750, 1676, 1504, 1583, 1890]. Spatio-temporal [1285, 1016, 3274, 1985, 1542, 1777, 458, 750, 1676, 1504, 1583, 1890]. Spatiotemporal [3317, 2961, 2265, 3282, 2837]. Special [1696, 2992, 1697, 2496, 2012, 3002, 1983, 757, 3418, 2186]. specialist [2192, 1128, 3479]. speciation [2756, 2351, 1970]. Species [1517, 2603, 1544, 2846, 1949, 1668, 1361, 1942, 3292, 3395, 356, 1576, 417, 980, 968, 2328, 1156, 1274, 2226, 2438, 3338, 1731, 396, 3496, 171, 2817, 571, 94, 140, 266, 2467, 3332, 1270, 2848, 2153, 3499, 3582, 746, 2841, 1231, 1049, 846, 3045, 479, 3372, 71, 1998, 714, 3085, 2000, 2907, 127, 398, 3522, 3439, 2976, 1471, 2712, 848, 382, 445, 526, 905, 2609, 1017, 3141, 3363, 2450, 688, 401, 458, 966, 1433, 1452, 1036, 165, 2857, 1357, 2772, 119, 1847, 945, 1065, 2036, 866, 3545, 616, 1676, 2484, 273, 2410, 3500, 3354, 2600, 2424, 3329, 540, 594, 3062]. species [2867, 513, 2816, 1245, 52, 847, 806, 797, 2269, 2288, 1277, 3408, 1287, 1112, 16, 2274, 3398]. Species-abundance [2603]. species-dependent [2226]. specific [3213, 312, 520, 2516, 2431, 1063, 2994, 273, 1706]. specificity [1037]. specimens [3231]. SPECT [2760]. spectra [609, 1976, 250]. Spectral [500, 3169, 3186, 3486, 624, 1762, 2871, 2886]. spectrum [2716, 3439, 3256, 3597, 2944, 2565]. Speed [1143, 1316, 3539, 489, 3059, 1357, 3375, 3006, 581, 2964, 194, 2932, 2358]. speeds [2159, 2053, 2347, 2772, 1920, 1511, 1371, 1383, 1669, 1751]. speedup [2477]. sperm [1472, 1199, 387]. spermatocytes [2815]. spermatozoa [2540]. sphere [3594]. spherical [930, 1285, 2807, 2199, 44, 43, 600]. spheroid [1450]. spheroids [1302, 429, 477, 1480]. SPI1 [1971]. spike [1441]. spikes [1605, 3391, 1997]. spiking [1432, 1711, 1943, 2028, 2728]. Spiky [2132]. spin [2077]. spinal [342]. spine [1149]. spines [3089]. spins [3097]. spiny [2142]. Spiracular [3309]. spiral [3147]. split [1991]. splitting [3015]. Spontaneous [570, 3173]. sporophytic [2904, 1996]. Spots [3147, 2870, 1736]. spotted [3345]. Spread [1263, 3606, 2893, 3425, 1364, 788, 406, 3223, 330, 1898, 2181, 150, 174, 129, 2122, 1393, 138, 516, 2291, 3254, 3521, 3348, 1117, 2084, 2053, 2742, 3513, 1448, 3430, 224, 1770, 1514, 2178, 2794, 2386, 863, 3109, 481, 2349, 2749, 101, 194, 2532, 1481, 2488, 2808, 1356, 3488]. spreaders [3409]. Spreading [1357, 1920, 2330, 2932, 1751, 3590, 3171, 3295, 3059, 3157, 2053, 3477, 2347, 3365, 2699, 533, 2964, 2635, 2248, 1371, 1383, 1669, 2720]. spruce [1189, 198]. squared [11]. squares [3576]. squid [115]. stabilisation [2262, 1068]. stabilising [2266]. Stability [139, 1246, 1247, 769, 708, 1806, 1915, 186, 176, 1653, 1548, 802, 972, 3252, 3214, 3444, 3336, 306, 95, 471, 2928, 96, 554, 2844, 331, 1474, 2929, 35, 1882, 132, 546, 1526, 1479, 2098, 348, 555, 1243, 275, 2214, 372, 541, 1555, 828, 103, 3581, 1960, 558, 715, 1004, 796, 3265, 1646, 2380, 2071, 2441, 2435, 2202, 300, 690, 851, 1061, 241, 739, 503, 2262, 1976, 1563, 373, 476, 1888, 2662, 2278, 1484, 246, 1052, 1394, 2406, 3267, 2751, 2952, 2766, 3372, 71, 472, 1815, 755, 64, 193, 141, 127, 509, 965, 1162, 2753, 216, 786, 395, 1254]. stability [252, 487, 435, 2260, 175, 3391, 2316, 3257, 165, 1645, 789, 1183, 2857, 1293, 2207, 704, 679, 632, 3465, 2234, 407, 997, 1181, 249, 224, 383, 2900, 974, 2611, 3205, 927, 2239, 2795, 1108, 2405, 2618, 1692, 620, 3329, 1241, 1075, 513, 2442, 912, 113, 1028, 2048, 1647, 183, 191, 335, 703, 539, 1100, 1268, 392, 3535, 63, 207, 3335, 1964, 3577, 3427]. Stability-instability [3336]. stabilization [373, 588]. Stabilized [1592]. stabilizing [1499, 371, 961, 1078, 398, 752]. Stable [145, 458, 2045, 1969, 2370, 540, 1744, 222, 532, 608, 1725, 1235, 1679, 1437, 1048, 3004, 851, 1425, 450, 2099, 1636, 2496, 106, 3203, 39, 225, 401, 2242, 2995, 2984, 336, 3185, 482, 37, 2657, 282, 531, 426, 447, 970, 267, 542, 3019, 514]. stacked [3274]. Stage [1037, 2893, 3158, 2878, 1164, 1648, 1875, 3426, 3276, 3605, 1813, 1955, 1467, 1815, 1459, 1452, 2678, 2999, 2559, 1967, 1145, 1446].

Stage-specificity [1037]. stage-structured [2893, 1164, 3276, 1459, 1452, 2678, 3465, 2999, 2559, 1145, 1446]. Staged [3459]. stages [1923, 3229, 3013, 2968, 2488]. standard [1576, 3094, 519, 764]. standing [579]. Staphylococcus [1879, 1501, 2913]. starvation [3237, 2260, 1627]. State [2312, 2850, 1496, 2888, 1739, 1246, 1247, 1497, 2878, 2998, 724, 2924, 1528, 504, 1976, 3102, 3001, 3319, 1824, 2584, 3268, 2928, 30, 1151, 64, 554, 2844, 2906, 3031, 3284, 181, 114, 147, 2759, 252, 293, 1861, 3452, 2275, 1444, 11, 3356, 258, 3148, 3510, 72, 1238, 113, 1150, 1434, 1490, 3473, 981]. State-dependent [2312, 1496, 2928]. state-space [1246, 1247, 3356]. states [3072, 850, 330, 3069, 373, 3048, 2134, 323, 1157, 3372, 1114, 7, 2939, 3108, 277, 2567, 2309, 2572, 3271, 639, 2922, 1735, 2132, 1964, 2624, 3507]. statespace [1476]. Static [86, 1237, 373, 3414, 3369, 3147]. stationarity [1269, 2176, 2465]. Stationary [3182, 3183, 277, 1736, 2882, 3504, 3400, 746, 3055, 237, 437, 995, 64, 395, 659, 893, 11, 2415, 372]. Statistical [793, 720, 1044, 2359, 887, 811]. statistically [2706]. Statistics [2653, 1397, 1599, 1540, 3330, 1921, 765, 160]. status [2458, 2838, 1327, 3170]. status-based [1327]. STDP [3434]. STDP-based [3434]. Steady [2924, 1824, 323, 852, 2888, 1739, 2878, 1135, 724, 3069, 1976, 373, 3319, 2584, 30, 1151, 554, 2844, 181, 293, 1448, 2275, 3148, 3510, 3271, 1238, 639, 1150, 1434, 1490, 2922, 1735, 2132, 981, 1964, 3321, 2624]. Steady-state [2924, 1824, 1739, 724, 30, 1151, 554, 181]. steady-states [3271]. stealing [2757]. Steiner [2037]. stem [1322, 1946, 2005, 1723, 3367, 1406, 2426, 1149]. stenotic [2363]. stents [1464]. step [716, 1210, 2555, 1264]. stepping [258]. Steps [859, 1823, 2602]. Steric [1683, 1708]. sterile [2067, 397, 3283, 3334]. sterilization [2385]. sterilizing [2039]. Stevens [2188]. stiff [1825, 305]. stiffness [2916, 2863]. stimulated [759, 3148]. stimulation [596, 3393, 3388, 125, 878, 2560]. stimulus [2632, 1819, 2898]. stimulus-dependent [1819]. Stochastic [3596, 1720, 3167, 1269, 2176, 2617, 2986, 2439, 903, 2328, 1772, 3131, 2478, 1270, 3323, 2529, 3319, 1636, 2445, 2102, 2404, 3548, 1671, 1763, 1901, 67, 2778, 2732, 1232, 2388, 1431, 3306, 169, 3491, 2956, 1291, 2341, 505, 2118, 2500, 3568, 3413, 2560, 797, 2426, 2551, 643, 1047, 191, 1020, 1111, 490, 723, 828, 1336, 1237, 1923, 2683, 2411, 2556, 2905, 2449, 1599, 2931, 1996, 2575, 2056, 3083, 2902, 1294, 1670, 2787, 1907, 1166, 1985, 3474, 634, 3223, 2590, 1874, 2711, 3025, 727, 2741, 1898, 2465, 2756, 2641, 3294, 2695, 922, 1056, 3458, 1738, 3586]. stochastic [473, 746, 3008, 2971, 2406, 1067, 3055, 323, 3464, 759, 131, 3165, 2268, 2533, 2461, 999, 995, 2936, 3040, 3075, 3175, 3183, 1564, 670, 1678, 1872, 625, 1945, 2074, 1873, 1501, 2157, 2958, 3128, 2610, 324, 3309, 2082, 1262, 1263, 725, 2546, 3261, 2984, 2567, 2171, 2921, 2884, 2941, 1527, 2330, 1651, 2811, 3207, 1617, 153, 2462, 3050, 2334, 195, 2620, 1668, 2040, 1577, 599, 1760, 1987, 1727, 2428, 109, 747, 2499, 483, 618, 612, 2782, 1100, 3246, 3281, 50, 2295, 3420, 3524, 294, 2704, 1820, 1153, 2935, 591, 1031, 3324, 2250, 2434, 2566, 324]. stochastically [1158, 2387, 1891]. stochasticity [1316, 3415, 948, 2360, 2361, 2545, 1255]. stochastique [2556]. stocking [268, 299, 3305, 866]. Stoechiometric [3342]. Stoichiometric [2174, 3241, 2729, 1981]. Stoichiometry [3416]. Stokes [1755]. Stokeslet [1581]. stomatal [379, 292]. stone [258]. stop [3111]. storage [2712, 1682, 770]. stored [2807]. Stores [2612, 2608]. Storing [98]. strain [1503, 2228, 2323, 796, 2202, 2580, 1327, 3054, 2857, 754, 3390, 3143, 3497]. strains [1113, 1097, 1370, 55, 2459, 3531, 2501, 2652, 3524]. strand [2534, 3184, 108, 2172, 1524]. strands [209]. Strange [356, 974]. strategic [1061, 2979]. strategies [1980, 2726, 2308, 3222, 1235, 3103, 1437, 2071, 3478, 851, 1571, 3302, 2579, 2406, 3081, 1488, 2496, 1491, 1144, 3453, 1640, 2242, 2995, 2239, 1800, 482, 1353, 2754, 73, 118, 255, 2034, 2586, 3136, 1496, 2877]. Strategy [1077, 2311, 2099, 2588, 2597, 2466, 687, 737, 1987, 2116, 2274, 1864, 2007, 3436, 1081]. stream [3464, 3545, 2488]. streams [2053]. streetcar [1053]. strength [913]. stress [1450, 1302, 2807, 3440, 2388, 1240, 1480, 2863, 278, 3396, 3403, 1070]. stress-induced [3396]. stressed [3589]. stresses [2213, 443]. stretch [435]. stretching [3154]. striated [1120]. strict [1378]. Stripe [1523, 1634]. striped [853]. stripes [2870, 3147]. stroboscopic [1165]. Strong [851, 3527, 3281, 3249, 1807, 1808, 2154, 2817, 961, 3375, 3606, 2245, 220, 1249, 897, 1904, 1922, 2910]. strong-migration [220, 1249, 897]. stronger [2657]. strongly [2496]. Structural [2930, 3562, 2205, 120, 1863, 1476, 3270, 2337, 351, 1164, 2524, 2019, 2691, 1707]. Structure [3106, 2466, 2218, 2881, 3454, 2726, 3566, 2888, 440, 3325, 2430, 2725, 2226, 733, 1933, 362, 2032, 2052, 2445, 1734, 2018, 2849, 2757, 1024, 216, 2823, 2534, 3184, 820, 3286, 1293, 1921, 3261, 2740, 1309, 2840, 1536, 513, 432, 1912, 638, 483, 3188, 1174, 3538]. Structured [1648, 2703, 3597, 1295, 1361, 3178, 2893, 3380, 2421, 1541, 3020, 3163, 1362, 1551, 853, 2458, 1478, 2420, 708, 1164, 2575, 2526, 971, 1653, 1737, 1875, 1009, 1119, 1349, 1673, 3276, 3251, 1211, 3328, 503, 561, 622, 982, 2828, 1896, 2536, 1530, 343, 1138, 1301, 1888, 2662, 3102, 3001, 3000, 1726, 2507, 3586, 1286, 2674, 1049, 1918, 2456, 1467, 1639, 1459, 3054, 3164, 327, 800, 1317, 1352, 1412, 1204, 1640, 2416, 1116, 1338, 786, 1632, 2126, 2814, 2487, 3617, 3194, 3529, 1452, 3257, 2678, 1835, 2781, 229, 311, 3345, 700, 632, 3465, 2968, 3259, 1838, 2705, 2999]. structured [1476, 1809, 1652, 805, 157, 220, 2118, 3614, 804, 898, 897, 1131, 1904, 3162, 3278, 1353, 2559, 573, 1451, 2840, 3270, 2319, 958, 1614, 888, 1062, 2982, 936, 969, 1145, 1446, 1911, 2048, 692, 3124, 680, 1660, 3022, 1255, 1084, 1314, 1463, 3621, 2755, 2639, 1555]. structured-population [1551]. structures [2193, 2683, 1648, 2314, 906, 1812, 1402, 1700, 2059, 2083, 2225, 2660, 3557, 2477, 2653, 2944, 1704, 467, 1701, 1707, 1706]. studies [1397, 2067, 2795]. study [3095, 2617, 2458, 2452, 2168, 504, 2177, 2692, 3303, 2297, 2276, 2365, 3350, 2124, 1908, 410, 2941, 105, 3608, 2642, 2350, 2451, 3388, 3489, 1761, 655, 600, 2763, 1456, 1685, 2398]. Studying [2165, 2764, 233]. sub [1535, 1869, 3476]. sub-optimal [1535]. sub-population [1869]. Sub-Saharan [3476]. subcritical [2656, 2787]. subcutaneous [783, 941]. subdivided [1311, 1807, 1808, 310, 1442, 1249, 931]. subgraph [2468]. Subharmonic [1201, 425, 460]. subharmonics [426]. subject [2643, 437, 2277, 1117]. subjected [909]. submitted [723]. Submultiplicative [390]. subpopulation [1174]. subspace [3050]. substitutable [1537]. Substitution [1809, 3065, 2534, 3184, 2821]. Substrate [968, 3490, 3252, 716, 1359]. substrates [1850]. substructure [2313]. subtilin [2551]. subtype [1409]. subtype-2 [1409]. subunit [2577]. Success [2347, 2218]. successful [2750]. succession [1993, 3410, 3436]. sudden [1634]. suffice [705]. sufficient [2585, 2131, 2849, 1925, 3519, 3104, 2127, 386]. suggested [824, 825]. suicide [2497, 1352]. sum [502, 2127]. summary [3338]. sunfleck [111]. sunlight [1667]. super [2510, 1824, 2939, 1392, 3409]. super-diffusive [2510]. super-Gaussian [2939]. super-infection [1824, 1392]. super-spreaders [3409]. supercoiled [1524]. supercritical [699, 738]. superdiffusive [2532]. Superinfection [3578]. Superinfections [1725]. Superposition [779]. supplementary [1019]. supply [3584, 720, 1538, 979]. Support [1500, 2736, 1329, 3243, 2873]. supported [1913]. supports [2424]. suppression [2627, 3283, 3436]. Surface [2217, 2564, 1044, 1986, 1217, 1683, 263, 235, 2676, 370, 1556, 717, 3197, 1115, 2734]. surface-volume [1217]. surfaces [1986, 1818, 1902, 1285, 2240]. surrounding [1302, 3403]. surveillance [318, 424, 3448]. surveillance- [424]. survey [2219]. Survival [2636, 700, 2659, 876, 875, 943, 853, 369, 317, 474, 298, 874, 848, 445, 765, 1230, 3136, 3514, 3335]. Surviving [2479, 3496, 1998]. susceptibilities [2455]. susceptibility [3574, 1915, 1097, 3353, 1508, 2047]. Susceptible [1023, 1589, 2552, 2553, 2296, 3259, 3450]. Susceptible-infected [1589]. Susceptible-infected-removed [1023]. suspension [1179]. suspensions [1938]. sustainability [2329]. sustainable [888]. Sustained [1963, 2229, 1929]. swallowing [3608]. swarm [1551, 1134, 1826, 2864, 1193]. swarm-colony [1551]. swarming [1221, 999]. swarms [1146]. sweep [1910, 1680, 1999]. sweeps [1962]. swelling [2476, 1771]. swimmers [1938]. swimming [2136, 1179, 1620, 2523, 2792, 1202]. Swine [3577]. switch [1709, 1971, 1689, 2621, 3180, 2538, 255, 1137, 2442]. switch-like [3180, 1137]. switch-type [1709]. switched [3036]. switches [3255, 3416]. switching [2150, 3547, 2619, 2633, 1672, 1671, 3130, 3136, 2808]. symbiont [3595]. symbiosis [1469]. symbolic [2100, 2660, 3509]. Symmetric [1585, 2153, 1503, 3213, 332, 711, 3031, 1342, 2534, 3184, 665, 538, 1387, 937, 970]. Symmetries [2425, 2483, 2323, 2391]. Symmetry [890, 3374, 2091, 2120, 2367, 459]. Symmetry-breaking [890, 3374]. sympatho [1579]. sympatho-vagal [1579]. symptomatic [3467]. symptoms [751]. synapses [1897]. synaptic [2077, 3610]. synaptically [1267]. synchronisation [1037]. synchronised [1504]. Synchronization [2902, 535, 19, 2627, 1485, 1967, 3492]. Synchronized [1885, 168]. synchronous [1351]. synchrony [2206, 832, 2055]. syncytial [3287]. syndrome [743, 1126]. synergistic [2861]. synonymous [2645]. synthesis [2161]. synthetic [3454, 2779, 3506, 1937, 2918]. syntrophy [3506]. System [1633, 1592, 278, 1545, 2830, 1118, 626, 706, 2237, 3275, 2458, 1246, 1247, 1497, 715, 684, 1086, 886, 796, 1186, 3237, 1026, 3265, 3250, 2328, 3426, 330, 2441, 2435, 3358, 3393, 300, 2423, 3496, 3004, 3499, 2296, 1738, 2366, 1539, 986, 1045, 3550, 1587, 2000, 1845, 965, 1059, 1471, 444, 1289, 848, 840, 924, 495, 1873, 1723, 325, 401, 458, 459, 2453, 2855, 1139, 757, 1790, 2610, 3531, 2690, 921, 679, 413, 814, 1878, 3261, 2984, 2691, 3587, 61, 813, 1487, 598, 866, 3117, 2031, 1389, 104, 2899, 3556, 179, 2883, 3, 1071, 513, 3443, 3333, 1927, 806]. system [1889, 1866, 498, 1512, 2048, 1112, 862, 1647, 1695, 1922, 2023, 3126, 2358, 1030, 2517, 819, 76, 3621, 1305, 2046, 2755, 3150, 3427]. systematic [3414, 3180, 673]. Systems [1784, 1503, 828, 339, 2464, 823, 2075, 548, 707, 1892, 793, 3485, 2114, 558, 1902, 3351, 3044, 1991, 683, 778, 3032, 609, 186, 176, 268, 299, 2377, 1439, 967, 1874, 3576, 916, 36, 2817, 2203, 954, 2504, 622, 2262, 2722, 2139, 964, 3569, 3134, 2509, 586, 409, 649, 812, 2432, 102, 3180, 855, 1049, 846, 1777, 2524, 145, 953, 30, 667, 472, 515, 1068, 1114, 168, 3453, 2431, 243, 141, 127, 2660, 3182, 3002, 663, 3203, 1564, 429, 477, 636, 2850, 1632, 1626, 653, 1254, 2475, 80, 142, 2566, 2074, 33, 88, 1334, 161, 688, 230]. systems [1433, 645, 165, 789, 748, 2399, 262, 1859, 119, 1920, 3449, 287, 945, 1206, 950, 1882, 2362, 2615, 1055, 112, 974, 11, 782, 228, 1178, 31, 58, 2239, 2682, 34, 691, 1201, 1676, 1137, 1307, 2576, 1969, 2370, 685, 1743, 2310, 2964, 1266, 3329, 376, 1504, 1355, 432, 797, 260, 1028, 932, 3124, 3123, 3245, 1990, 3281, 3442, 3424, 810, 2129, 2132, 3082, 1736, 1669, 1631, 1661, 3290, 1823, 206, 289, 1332]. systems-theoretic [3032].

T [2617, 3567, 2766, 3192, 2341, 2980, 2196, 1836, 2048, 1768]. T-cell [3567, 2766, 3192, 2341, 2196]. T-cells [1768]. taboo [3106]. taboo-free [3106]. tachycardia [498]. tactic [1180]. Tajima [2379]. Takens [3500]. takes [3003]. Target [3093, 2264, 2655, 246, 2925, 2917]. target-mediated [2264, 2655]. targeted [3389, 2760]. targeting [2185, 2403]. targets [3581]. task [3161]. Tau [3532]. Taxa [1595, 1720, 2535, 2791]. Taxis [1657, 3113, 1431, 2899]. taxon [3559]. taxonomic [90, 1072]. taxonomy [8]. TB [1758]. TcdC [2364]. technique [2067, 2645, 3230, 65]. techniques [431, 552, 2092, 3514, 2816, 3421]. tectal [802]. tell [568, 1787, 2095]. telomerase [2266]. telomere [2266, 3028]. Temperature [3345, 149, 1709, 2637, 783, 941, 2295, 819]. Temperature-dependent [149, 819]. templates [1560]. Temporal [173, 496, 3367, 3274, 2421, 3547, 1985, 1285, 1542, 1777, 768, 24, 458, 750, 1016, 1676, 2587, 1504, 1583, 1890]. Temporally [326, 537, 2357, 3257, 1007]. temporary [1841, 2307]. tendon [1079]. tensor [3392, 1148]. term [2677, 1898, 2684, 1348, 3459, 1292, 1057, 278, 3354, 1054]. terminal [1721, 2059]. termination [2160]. terms [1044, 1717, 3615]. terrace [1826].

Territorial [2460, 2452]. territoriality [1903]. territories [1109]. test [1224, 2786]. testing [3296, 3466]. testosterone [1484]. tests [246, 1756, 360, 3100]. tether [1259]. tetraploidy [232]. tetravalent [3470]. th [3559, 2813]. thaliana [3247]. theatre [2180, 1883]. their [1397, 1548, 915, 3000, 3037, 3385, 126, 2359, 2516, 2739, 1811, 3307, 3542, 1142, 1248, 1994, 1328, 1496, 194, 2227]. Theme [1774]. Theorem [2318, 2761, 954, 2670, 1564, 1520, 653, 1904, 2415, 2132, 2188, 189]. Theorems [1145, 2575, 1862, 3025, 2144, 1549, 133, 647, 699, 50, 1255]. theoretic [1900, 3032, 3373, 2287, 1665, 1666, 3364, 2007, 1828]. Theoretical [1594, 485, 1621, 2177, 1210, 360, 795, 3433, 3248, 2421, 1833, 2663, 1903, 1984, 1148, 3371, 1102, 1850, 233, 1631, 3149]. theories [1348, 97, 2188]. Theory [59, 1138, 263, 2508, 97, 1591, 1534, 3600, 1, 532, 1044, 2289, 290, 915, 933, 619, 402, 2433, 3094, 1056, 1301, 1888, 2662, 586, 1636, 1052, 2, 143, 2744, 1582, 345, 1053, 1856, 3175, 3256, 3370, 646, 637, 1848, 2450, 1677, 4, 1568, 2917, 1341, 617, 111, 3495, 2078, 3312, 148, 375, 1389, 2811, 2166, 158, 3396, 2141, 2379, 2696, 3173, 2130, 2195, 2769, 2220, 2544, 652, 3577, 2130]. therapeutic [3095, 3581]. therapeutically [328]. therapies [3550, 3288]. therapy [3073, 3605, 3422, 3084, 3229, 3423, 3367, 2822, 3230, 1802, 2560, 2305, 2179, 3248, 2824]. thereon [2123]. thermal [827]. Thermodynamic [11]. Thermodynamics [1558]. theta [2137]. theta-gamma [2137]. thin [1221]. thin-film [1221]. thinning [1530]. Thompson [1574]. Three [1650, 3605, 1813, 1549, 667, 2550, 3551, 578, 606, 1576, 1247, 2924, 1168, 396, 3081, 156, 3145, 2297, 3034, 1815, 2849, 3374, 526, 3157, 2634, 1821, 401, 458, 966, 165, 1978, 2207, 119, 3104, 1291, 1065, 859, 611, 3360, 2586, 3354, 797, 3386]. three-component [3145, 3157]. Three-dimensional [667, 2550, 1247, 2924, 2297, 2634, 1821, 1978, 2207, 3386]. three-level [1168]. three-locus [606]. three-node [3034]. three-species [3354]. Three-stage [3605, 1815]. three-tiered [3360]. Threshold [442, 1734, 2085, 1012, 786, 3573, 587, 2061, 3232, 2087, 2307, 3555, 1598, 1764, 2945, 2182, 3515, 2578, 566, 2977, 2365, 848, 70, 3378, 879, 2239, 3114, 2350, 1716, 189]. Thresholds [138, 1465, 2723, 1013, 3086, 3253, 3438, 1730, 2788, 942, 2889]. thrombus [2570, 1909]. Thucydides [3296]. thymocyte [1660]. Thyone [387]. tick [2432, 1543, 3312, 1436, 3146, 3427]. tick-borne [1543, 3146]. tick-host [3427]. ticking [2887]. ticks [2887]. tiered [3360]. tight [1288, 2629]. Tilapia [3452]. tilt [2950]. Time [3058, 94, 192, 2028, 3114, 1468, 1889, 1866, 2782, 3577, 3150, 99, 891, 801, 1786, 1923, 894, 1311, 3030, 1167, 2673, 1004, 3047, 1670, 1907, 2879, 2638, 3223, 1711, 1943, 3576, 3276, 2731, 2790, 3251, 1585, 2232, 3328, 2122, 1567, 1393, 74, 2641, 2619, 2094, 681, 95, 549, 3060, 1766, 2771, 2744, 2835, 479, 3525, 3165, 896, 959, 3543, 3549, 64, 849, 3216, 696, 763, 768, 2977, 2365, 3355, 2975, 1315, 377, 2942, 2487, 2566, 1848, 3340, 3195, 1433, 867, 693, 4, 1859, 366, 2895, 3609, 3122, 430, 3430, 1838, 2327, 488, 2802, 1065, 551, 3622, 1714, 2921, 121, 3185]. time [2165, 1015, 1123, 3518, 3278, 3448, 43, 2994, 1935, 1011, 2728, 2883, 3476, 447, 1927, 113, 1409, 3623, 3208, 588, 862, 3019, 1255, 1854, 391, 934, 2675, 1828, 189, 1871]. time-delay [1123, 3448]. time-delayed [3030, 3609]. Time-dependent [3058, 3114, 2782, 3276, 2895, 2165, 43]. time-discretized [3165]. time-for [3430]. time-heterogeneous [2942]. time-inhomogeneous [1670]. time-lag [64, 2327]. time-lags [1167]. Time-periodic [192, 768, 1315, 3340]. time-profiles [2879]. time-series [2094, 1838]. time-space [3278]. time-trees [2744]. time-varying [894, 3518, 862, 2675]. times [2409, 3485, 2787, 3276, 741, 2633, 1998, 2849, 751, 3068, 3475, 2284, 1904, 2811, 3583, 2576, 1340, 2246, 492]. timescale [2779, 3477, 1849]. timescales [3172, 2993, 3471]. timidus [1436]. Timing [1192, 3475, 1905]. tip [1200, 2961]. tipping [2998]. tissue [626, 1953, 3275, 431, 1679, 760, 402, 1001, 3332, 2177, 2507, 2520, 771, 1877, 1777, 2476, 3491, 459, 404, 1568, 2490, 2191, 2967, 3279, 1761, 2210, 854, 1804, 2093, 1077, 1344, 1153, 3387]. tissues [1737, 852, 1799, 783, 941, 3514, 2130, 2195, 2054, 3289, 1258]. TMDD [2800]. TolC [3150]. tolerance [1224, 2277, 403]. tolerance-competition [2277]. tongues [1459]. tool [772, 2764]. tools [1801]. Top [365, 1592, 1168]. top-predator [1168]. Topological [2193, 158, 2658, 3425, 209, 2760, 2032, 2339, 2653]. topologies [3292, 2957, 2205]. Topology [2572, 2324, 2219, 3419, 2955, 1822, 2957]. Torulopsis [55]. total [271, 3179, 3543, 3363, 1265, 3121, 1850]. township [1758]. toxicant [846, 3589]. toxicant-stressed [3589]. toxicants [441]. toxicity [2057]. toxicokinetic [1724]. toxicology [1373]. toxin [3180, 2364]. toxin-antitoxin [3180]. toxins [2867, 2808]. tracer [1677]. traces [2657]. tracheal [2476]. tracing [3409]. tracker [2502]. tracking [2573, 849, 1779, 2502]. tract [1840]. tractability [2601]. tractable [1330]. Traction [1775, 478]. tracts [2414, 2933]. trade [2277, 2375, 3324]. trade-off [2277]. trade-offs [2375, 3324]. tradeoff [2806]. traffic [2236, 380]. Trail [2133, 962, 1475, 1014]. trail-following [962]. trails [962]. Trait [1809, 1136, 2255, 2439, 1499, 2716, 1078, 3065, 1961, 1416, 1580, 2036]. trait-based [2716]. trait-substitution [3065]. traits [3380, 1688, 955, 2670, 3277, 3304, 2496, 2453, 3531, 1057, 2691, 2221]. trajectories [2270, 270]. transcendental [63]. transcription [2987, 3548, 3268, 2595, 2006]. transcriptional [2365, 1844]. transduction [960, 2114, 1978, 2207]. transectional [748]. transectional/circumferential [748]. transfection [3316]. transfer [3041, 3135, 2934, 2933, 459, 3196, 2886]. transfers [2438, 2521]. transform [3294, 134]. transformation [148]. transformations [2391, 2163]. Transformed [1893]. transgenic [1381]. Transient [626, 1006, 293, 1313, 1381, 1502, 777, 996, 280, 114, 147, 2045, 3186, 3486, 2229]. transients [3354]. Transit [2576, 391, 492]. Transition [2058, 226, 237, 2324, 2606, 2998, 176, 2991, 3336, 3145, 734, 2874, 1762, 2739, 3031, 2814, 987, 2132]. Transitions [3057, 2019, 1811, 1714, 3271, 2644]. Translating [999]. translation [2987, 3268, 3316]. translational [2096, 2629]. translocated [1685]. translocation [2142, 2104, 2874]. Transmission [2267, 2283, 2723, 2784, 3596, 2004, 2497, 2105, 776, 1274, 3036, 2741, 3446, 2536, 918, 543, 2801, 883, 1069, 3459, 1297, 2124, 3470, 843, 3511, 2165, 1436, 2768, 2834, 3143, 2709, 957, 3507, 2295, 2582, 2720, 2119]. transmission-blocking [2004]. transmitted [1364, 434, 1097, 675, 1533, 1713, 1428, 3460, 3273]. transplant [985]. transplantation [1406]. Transport [1217, 2065, 431, 2136, 1721, 2614, 2542, 1006, 1223, 2194, 1252, 1525, 1677, 3378, 3261, 773, 3148, 3572, 854, 2688, 2130, 2195, 655, 115, 1461, 1462, 1581, 1344, 1405]. transport-related [3261]. transportation [2386]. transposable [841, 595, 1770, 694]. trap [256]. traps [122]. travel [3538]. travel-related [3538]. Traveling [1225, 1472, 3137, 3145, 3308, 2068, 1187, 1906, 856, 1695, 2833, 613, 830, 1379, 1785, 1920, 37, 1647, 3118, 1371, 1383, 993]. traveling-wave [830]. Travelling [1001, 1101, 416, 1256, 41, 869, 1681, 3119, 2256, 1107, 1361, 357, 489, 138, 2684, 2296, 1146, 755, 867, 2827, 407, 132, 249, 927, 469, 1000, 541]. Travelling-wave [1681]. treat [1103, 2144]. Treatment [3212, 3478, 2300, 1541, 548, 2677, 2858, 1454, 1830, 1296, 3164, 3549, 3459, 152, 3453, 3618, 1297, 1329, 1642, 3339, 3412, 3098, 2652, 1521, 957, 3604, 2007, 2471]. Treatment-induced [3212, 1521]. treatments [3405, 3300, 3209]. Tree [2873, 2710, 1942, 3395, 2606, 2955, 3338, 3293, 2957, 2042, 758, 3407, 2907, 2033, 2660, 2906, 2859, 2351, 1847, 2369, 1706, 2706, 149, 2639, 1703, 2273]. Tree-based [2873]. tree-child [2273]. tree-grass [2639]. Tree-hierarchy [2710]. tree-representable [2660]. Trees [1595, 2970, 1720, 1942, 2261, 1887, 1862, 3225, 3127, 3240, 2954, 3061, 2848, 2049, 2436, 3015, 3138, 2744, 2820, 3045, 3385, 654, 1717, 1718, 2762, 2976, 3038, 1024, 2846, 1570, 2589, 2713, 3159, 1337, 2596, 3313, 2873, 2572, 3018, 2037, 3398, 1386, 819]. tri [1196, 1040, 2546, 1740, 1082, 2708]. tri-LGT-nets [2708]. tri-phasic [1740]. tri-trophic [1196, 1040, 2546, 1082]. trials [160]. tribute [2485]. trichome [1152]. trichotomy [885]. Trick [2975]. trifurca [585]. triggered [3224]. Trinets [3218, 2273]. tripartite [3336]. triple [3381]. triple-negative [3381]. triplet [2713]. triplets [2651]. tripling [1324]. Trojan [2000, 2274]. trophic [1196, 2998, 1968, 1513, 1650, 3239, 365, 509, 1040, 3350, 2546, 1082, 3354]. True [1590, 1661]. TruEst [3508]. truncated [1518]. truncations [912]. Trunk [962]. trypanosomes [2032]. tube [1389, 2961]. tube-system [1389]. tuberculosis [3478, 1103, 3164, 3549, 3617, 3221]. Tubular [2249, 682]. tubule [598]. tubules [1372]. tubuloglomerular [2249]. tumble [1664, 2568]. tumbling [1427]. Tumor [3612, 3589, 3321, 1775, 1384, 1797, 2962, 1421, 3392, 3550, 800, 2200, 1584, 1329, 1893, 3399, 1161, 3168, 3014, 2010, 3542, 1498, 3404, 1466, 505, 2719, 3608, 2795, 1846, 1802, 2644, 2948, 1516, 2305, 2093, 1414, 3387, 2824]. tumor- [1161]. tumor-immune [2010]. tumor-induced [1466]. tumor-infected [3608]. tumoral [1737]. tumors [1415, 1339, 2760, 2170, 1184, 1803]. tumour [2306, 3212, 1473, 1354, 1198, 1772, 1027, 1285, 1240, 3423, 1480, 1793, 3367, 1798, 1801, 3403, 1306, 3289, 3599]. tumour-induced [1772]. tumours [915, 3192]. tunable [2123]. tuning [3537]. turbulence [1390]. Turing [2510, 1169, 3594, 2380, 3576, 964, 1076, 636, 676, 1474, 2399, 2615, 3206, 2079, 2101, 1266, 1007, 1028, 3123, 3442, 3424, 2866]. turning [3226, 1788, 1475, 2561]. turnover [2210]. tutorial [3562]. tutorial-based [3562]. Two [579, 2159, 266, 2641, 3385, 2694, 1220, 1595, 153, 3403, 2324, 801, 2253, 570, 2937, 347, 1311, 1246, 2878, 2930, 1228, 1345, 886, 3155, 1354, 1513, 968, 313, 3426, 1731, 3086, 1575, 36, 2580, 2668, 1085, 2454, 1124, 571, 851, 2716, 1929, 3444, 3582, 746, 90, 719, 808, 716, 3303, 2297, 479, 71, 1998, 857, 310, 666, 3543, 2573, 3616, 2189, 127, 1471, 3557, 1310, 848, 3159, 2388, 905, 1210, 1945, 24, 117, 252, 495, 2894, 3350, 1433, 1645, 2275, 1293, 1442, 1357, 1682, 2029, 2772, 538, 1064, 277, 3441, 598, 866, 3516, 937, 2271, 941, 1940, 2795, 2718]. two [2034, 1692, 348, 2501, 2652, 3354, 3136, 3390, 1388, 2369, 457, 540, 3064, 3143, 1870, 1295, 1035, 2269, 498, 648, 3497, 1994, 3408, 1911, 2680, 3208, 3082, 2167, 1207, 934, 2833, 2398, 2877, 3027]. two-allele [1345, 886, 24, 252, 2718]. Two-boundary [2641]. two-community [3616]. two-cycle [3516]. two-deme [2253]. Two-dimensional [579, 153, 2324, 1246, 3155, 3444, 2297, 666, 2894, 598]. two-fluid [2795]. two-gene [1929]. two-group [3441, 2833]. two-locus [2253, 347, 1228, 719, 808, 857, 117, 252, 1442, 1940, 2718, 1035]. two-oscillator [498]. Two-parameter [2694, 2029]. Two-patch [2159, 3303, 3543, 3557, 1945, 3082]. Two-phase [3403, 1354, 2271, 1870]. two-player [3136]. two-population [1293]. two-predators [1310]. two-sex [457, 2680, 3208]. two-species [3582, 746, 1998, 1357, 2772, 866, 2269, 3408]. two-stage [2878, 3426]. two-step [716, 1210]. two-strain [2580, 3390, 3143]. two-thresholds [3086]. two-trophic [3350, 3354]. two-type [2388]. type [1709, 1253, 1029, 2003, 2998, 2128, 1175, 2714, 739, 3432, 3180, 1845, 1162, 3256, 663, 1289, 2388, 2126, 1948, 3308, 3097, 2325, 789, 1981, 3587, 872, 1360, 1420, 3613, 2290, 418, 481, 480, 581, 3579, 2185, 2403, 267, 3220]. type-II [3579]. type-reproduction [2126]. types [801, 626, 851, 645, 2247, 2185, 2403, 3246, 2167]. Typhimurium [1971]. typhoid [2794]. Typical [2270]. tyrosine [2114].

ubiquitin [3266]. Uhlenbeck [2606]. ultrametrics [2100, 2660]. ultrasensitive [2593]. Ultrasensitivity [2310, 1989, 1850]. Umwelt [523]. Unbounded [3120, 2134]. unbranched [373]. uncertain [3178, 3351, 2743, 243]. uncertainties [3263]. Uncertainty [2591, 3012, 2979, 2200, 2140, 1864]. Uncovering [1897]. under-reporting [2302, 2649]. underdominant [3274]. Understanding [911, 2450, 1505, 2605, 3561, 2078]. Unequal [1468, 740]. unexpected [1320]. unfolding [2589]. unidirectional [1509]. unidirectionally [3049]. unidirectionally-coupled [3049]. Unified [2314]. Uniform [3029, 1245, 2228, 1926, 1114, 1659, 258, 1475, 2658]. UniFrac [2802]. Unifying [3518, 1917, 1835, 1100]. uninvadable [850]. Uniqueness [1647, 1140, 2422, 702, 843, 1226, 782, 3279, 480, 1000]. United [3507]. universal [1322, 1560, 2377, 3447, 2516]. Universality [3247]. universally [788]. unlabelled [559]. Unlinked [209]. unpredictable [1228, 1345, 3416]. unprotected [2034]. unranked [2957]. Unravelling [1169]. unreported [2826]. Unrestricted [305]. unrooted [1942]. unstable [1194, 2107]. unstructured [1411, 3614]. unveiling [3509]. updating [3186, 3486]. upon [2181, 3618, 569]. upper [2407, 2933]. Upscaling [3341]. uptake [2760, 281, 3309, 1805, 1283, 600, 3366, 1095, 1152, 1332]. uptake-diffusion [3366]. upwards [2136]. urban [2440]. urchin [2233]. urea [413]. ureteric [1464]. urinary [1840]. urine [1464]. urn [637]. urns [491]. usage [2645]. Use [55, 3287, 148, 1739, 2489, 403, 1652, 1404, 340, 2773, 1768]. used [2400]. useful [1549]. user [1778]. users [2122]. Using [3047, 2504, 3094, 2539, 1662, 2889, 3292, 3468, 2796, 1169, 1073, 2452, 2854, 2147, 915, 1724, 2747, 2433, 2348, 1445, 1797, 2094, 1765, 2649, 2736, 2856, 2294, 1002, 1561, 3407, 1584, 2388, 3141, 1568, 3461, 1838, 410, 1476, 2282, 3352, 3205, 2891, 2427, 2239, 1479, 43, 2795, 2990, 3514, 3092, 2343, 1930, 1803, 1732, 2948, 2706, 935, 1705, 321, 1159, 3599, 2715, 1619]. Utah [2291]. utricular [2483].

V [45, 601]. V-shaped [601]. Vaccination [3253, 2186, 3584, 1559, 3326, 2726, 2402, 2879, 3411, 2579, 2919, 3303, 3462, 3204, 3470, 3365, 2953, 1250, 2239, 2583, 1614, 2709, 2985, 2965]. Vaccinations [1589]. Vaccine [2705, 3584, 3460]. vaccines [2247, 2383, 3471]. vacuum [2939]. vagal [1579]. validate [1826]. validated [3448, 1355]. validation [1164, 504, 2200, 1937, 1389, 2195]. validity [3510]. valley [1912]. Value [2318, 2039, 40, 3061, 1834, 2436, 1582, 1718, 2189, 3018, 2093]. valued [3088, 1688, 3592, 2722, 1547, 2166]. values [3044, 3015, 3615, 1451, 3451]. Valveless [1389]. Variability [2622, 3088, 2409, 706, 1571, 2593, 961, 1017, 501, 1987, 1732, 2246]. Variable [2096, 2608, 1408, 1752, 2725, 332, 390, 3077, 1182, 1347, 1039, 1227, 3257, 433, 3230, 553, 1141, 744, 2555, 296, 361, 2612]. variable-environment [296]. variable-infectivity [1347]. Variable-Internal-Stores [2608]. variable-yield [1227]. variables [2931]. variance [2673, 2212, 1424, 2313, 2455, 1121, 2141, 3062, 3472]. variance-to-mean [3472]. variant [3448]. variants [3523]. Variation [1401, 1806, 2091, 604, 2018, 781, 24, 3318, 594, 1844, 512]. variational [1637]. Variations [1702, 174, 1037, 3244, 1547]. varied [3048]. variety [1417]. Various [2206, 319, 3411, 2728, 1174]. vary [253]. varying [894, 1959, 1169, 290, 428, 776, 304, 2465, 2159, 537, 398, 452, 1891, 2357, 35, 2390, 3518, 1007, 1886, 862, 1481, 326, 2675]. vas [1597]. Vascular [1329, 3116, 2971, 1584, 459, 1337, 3404, 1793, 1799]. vascularized [930, 1803]. vasoconstricted [3514]. Vector [3593, 1598, 2638, 3302, 1098, 2736, 479, 1639, 3521, 2903, 1932, 224, 2953, 3143, 2257, 1912, 3022, 2283, 3315, 3555]. vector-bias [3143]. Vector-borne [3593, 1598, 3302, 1639, 2953, 2257, 2283, 3555]. vector-host [2638, 1098]. Vectorial [3269]. Vegetation [1523, 2830, 2527, 2797, 3101, 3400, 2531, 3366, 3505]. veins [2363, 1575]. velocities [1202]. velocity [1540, 1728, 1811, 1664, 717, 2899, 794]. velocity-jump [2899]. venereal [203, 207]. ventral [1634]. ventricular [151]. versa [1631]. versatile [1364]. version [2657]. versions [3165]. versus [1752, 2425, 826, 3302, 962, 3161, 855, 3202, 1063, 1338, 2487, 1642, 3602, 1194, 430, 3536, 3388, 1054]. vertebrate [1885]. Vertex [3351]. vertical [314, 2886, 1557]. vertically [434]. very [2817]. vesicle [2573, 3482, 2550]. vesicles [3148, 3572]. Vessel [1337]. vessels [2213, 847, 2492]. vestibular [2080]. via [3234, 2151, 2799, 3223, 2858, 2779, 3330, 1639, 3615, 3183, 3370, 3229, 2491, 2634, 3217, 1448, 2339, 2855, 3014, 2284, 2217, 2386, 2501, 2927, 1712, 2132, 3456, 2543, 2950, 3221, 1703]. viabilities [390]. viability [711, 807, 1061, 913, 117, 487, 809, 821, 538, 658, 2982]. Viable [1086, 736]. vice [1631]. view [3160, 2799, 2469, 1052, 2332, 2103, 3245, 3598]. viewpoint [2418]. villi [795]. villous [137]. viraemic [1436]. Viral [3422, 2179, 2805, 991, 3394, 330, 3446, 2199, 2751, 2933, 3181, 1865, 2394, 2029, 3609, 1836, 1870]. virtual [3489]. virulence [2128, 730, 1533, 2997]. Virus [3394, 2335, 2267, 3326, 2832, 2412, 202, 1375, 2674, 1543, 1421, 3337, 3287, 2699, 2061, 1878, 2953, 1436, 2031, 3448, 3410, 3021, 2834, 1516, 2932, 3209, 3488, 3452]. virus-antibody [2412]. virus-cell [1878]. Virus-immune [3394]. viruses [2019, 1821]. visceral [2582]. visco [3252]. visco-elastic [3252]. viscoelastic [796, 1461, 1462, 2216]. viscoelasticity [3275]. viscosity [1938]. viscous [3332, 2981, 2271]. Visual [57, 2035, 3116, 212, 2080, 84]. visual-vestibular [2080]. visualization [1521]. vital [1606, 1248]. vitro [1956]. vivax [3578]. vivo [2346, 1761, 1079]. VLDL [1805]. volatile [1983]. voltage [2924, 1192, 3187]. voltage-gated [1192]. Volterra [3382, 3043, 356, 2510, 715, 1215, 3565, 1874, 3496, 607, 352, 1094, 416, 2366, 734, 472, 515, 141, 2778, 663, 1769, 3203, 134, 1993, 216, 3065, 653, 401, 458, 165, 1206, 2329, 259, 72, 620, 1089, 1333, 447, 2116, 267, 3281, 810, 3126, 3324]. Volume [673, 3212, 1791, 1960, 2419, 1217, 2213, 2024, 2069, 2040]. volume-filling [3212]. volumes [2516]. volumetric [1070]. Voluntary [1589, 2879, 3373, 92]. Volz [1924]. vs [1216, 722].

Wachstum [523]. wait [2311]. wake [1867]. walk [223, 3603, 100, 1076, 1309]. walker [1788]. Walks [366, 1253, 2978, 2972]. wall [2971, 1280, 2863, 137]. Waning [2838, 2458, 1898, 3544, 3452]. war [1043]. warming [2737]. warning [3354]. warping [1585]. was [1331, 2246]. wastewater [3448]. wasting [2291]. Water [319, 1460, 1607, 1826, 2939, 3309, 2522, 3210]. waterborne [2429]. waters [314]. Watson [1814, 1360, 1420]. Wave [1120, 3375, 2198, 1511, 418, 1413, 1323, 1618, 1472, 489, 2296, 416, 3457, 755, 830, 1379, 899, 1681, 2068, 3119, 1187, 407, 1906, 132, 249, 2864, 443, 682, 336, 34, 2451, 480, 37, 1000, 1107, 1647, 3442, 3424, 2833]. wave-pinning [3457]. waveforms [3396]. Wavefront [1565, 771, 3043]. Wavenumber [2688]. Waves [286, 656, 3357, 1225, 2142, 1175, 2542, 357, 1001, 1101, 138, 3137, 2684, 1195, 2574, 306, 3145, 667, 666, 613, 869, 1914, 867, 3308, 2827, 1785, 1920, 3147, 2256, 2851, 927, 856, 2160, 1441, 367, 1716, 1695, 2204, 3118, 1506, 1371, 1383, 993, 2646]. Wavespeed [1892]. way [1888, 2662, 1826, 1518, 180, 380]. Wazewski [40]. Weak [377, 670, 1807, 1808, 1874, 2817, 2781, 3133, 1178, 2718, 605, 2696, 1578, 2847, 361]. Weakly [1810, 1028, 1198, 315]. web [1968, 1469, 3360]. Webb [2215]. Weber [2875]. webs [341]. weekly [2407]. weighted [1717, 1635, 2635]. Weinberg [978, 895, 178, 809]. Well [2321, 692, 3535, 2203, 2164, 2845]. well-mixed [2203]. Well-posedness [2321, 692, 3535, 2164]. West [2267, 2674, 2699, 3410, 3021, 2932]. whale [75, 3613]. whaling [3613]. where [1331]. Which [842, 2311, 40, 510, 1226]. white [2414, 3013]. Who [2246, 798, 3273]. whole [1391]. whole-body [1391]. whom [798]. Wiener [1299, 104]. Wiener-like [104]. wild [2770, 2685, 3283]. wildlife [2311]. windy [1244]. wing [1505]. winged [2488]. Wings [1505, 1600]. winner [3003]. Within [3135, 2805, 2457, 2807, 253, 2187, 3609, 2745, 3561, 963, 2768, 2093, 2554]. within- [3561]. Within-host [3135, 2805, 2187, 3609, 2768, 2554]. within-household [2457]. without [331, 2353, 2538, 2384, 988, 2264]. WKB [3083]. wMelPop [2770]. Wnt [1849]. Wnt/ [1849]. Wolbachia [2750, 2770, 2674, 2758, 3080, 3436, 2720]. Wolbachia-infected [3080]. wolf [1109]. wombat [3618]. woody [2057]. Words [1397]. work [3161]. workers [3580]. workplaces [1973]. world [2743]. worm [1286]. wound [1363, 1382, 1832, 822, 928, 3451]. wounds [2072]. Wright [2882, 2991, 2252, 2714, 537, 719, 3125, 3091, 2447, 613, 1572, 384, 3546].

X [169]. xenology [2820, 2660, 2949]. Xenopus [1989].

Y-Chromosome [2274]. Year [1378, 2209]. yeast [589, 1937, 935, 1876]. yeasts [55]. yield [1227, 888]. yielding [901]. Yolky [2183, 2117]. young [633]. Yule [2212, 1763, 2572].

Zahavi [1348]. Zealand [1685]. zealots [2354]. zebrafish [2717, 2401, 3010, 2434]. zero [502, 265, 1850]. zero-dimensional [265]. zero-order [1850]. zero-sum [502]. zeta [2477]. Zhabotinskii [410]. Zika [3470, 2953, 2834]. Zimbabwe [1652]. ZIP [2423]. zone [2065, 2661, 3111]. zones [1003, 2967]. Zoonotic [2582]. zooplankton [1227, 925]. zooplankton-phytoplankton-nutrient [925]. zygotic [5, 49].

References

Anonymous:1974:ERM

[1] Anonymous. Experimental results motivated by a theory. Journal of Mathematical Biology, 1(1):1–2, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02339483.

Foias:1974:BHI

[2] Prof. Dr. C. Foiaş. A biological homology inference from ergodic theory. Journal of Mathematical Biology, 1(1):3–7, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02339484.

Schafer:1974:IDP

[3] Dr. E. Schäfer and Prof. Dr. H. Mohr. Irradiance dependency of the phytochrome system in cotyledons of mustard (Sinapis alba L.). Journal of Mathematical Biology, 1(1):9–15, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF02339485.

Lebowitz:1974:TAG

[4] Dr. J. L. Lebowitz and Dr. S. I. Rubinow. A theory for the age and generation time distribution of a microbial population. Journal of Mathematical Biology, 1(1):17–36, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF02339486.

Worz-Busekros:1974:ZAS

[5] Dr. Angelika Wörz-Busekros. The zygotic algebra for sex linkage. Journal of Mathematical Biology, 1(1):37–46, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF02339487.

Palm:1974:SMS

[6] G. Palm. On the selection model for a sex-linked locus. Journal of Mathematical Biology, 1(1):47–50, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF02339488.

Hadeler:1974:ESC

[7] Prof. Dr. K. P. Hadeler. On the equilibrium states in certain selection models. Journal of Mathematical Biology, 1(1):51–56, ???? 1974. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF02339489.

Bezdek:1974:NTF

[8] Dr. J. C. Bezdek. Numerical taxonomy with fuzzy sets. Journal of Mathematical Biology, 1(1):57–71, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF02339490.

Winfree:1974:PPC

[9] Dr. A. T. Winfree. Patterns of phase compromise in biological cycles. Journal of Mathematical Biology, 1(1):73–93, ???? 1974. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF02339491.

Resnikoff:1974:DGC

[10] Dr. H. L. Resnikoff. Differential geometry and color perception. Journal of Mathematical Biology, 1(2):97–131, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275798.

Mel:1974:TPE

[11] H. C. Mel and D. A. Ewald. Thermodynamic potentials and evolution towards the stationary state in open systems of far-from-equilibrium chemical reactions: The affinity squared minimum function. Journal of Mathematical Biology, 1(2):133–151, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275799.

Green:1974:FNA

[12] Dr. M. W. Green and Dr. B. D. Sleeman. On FitzHugh’s nerve axon equations. Journal of Mathematical Biology, 1(2):153–163, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275800.

Hadeler:1974:NSD

[13] Prof. Dr. K. P. Hadeler, U. an der Heiden, and F. Rothe. Nonhomogeneous spatial distributions of populations. Journal of Mathematical Biology, 1(2):165–176, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275801.

Tuffner-Denker:1974:PGM

[14] I. Tuffner-Denker. On a population genetic model by S. Karlin. Journal of Mathematical Biology, 1(2):177–185, ???? 1974. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275802.

Rubinow:1975:MMN

[15] S. I. Rubinow and J. L. Lebowitz. A mathematical model of neutrophil production and control in normal man. Journal of Mathematical Biology, 1(3):187–225, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF01273744.

Walther:1975:ENC

[16] Hans-Otto Walther. Existence of a non-constant periodic solution of a non-linear autonomous functional differential equation representing the growth of a single species population. Journal of Mathematical Biology, 1(3):227–240, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF01273745.

Swartz:1975:DPE

[17] Dr. J. Swartz and Prof. H. Bremermann. Discussion of parameter estimation in biological modelling: Algorithms for estimation and evaluation of the estimates. Journal of Mathematical Biology, 1(3):241–257, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF01273746.

Guckenheimer:1975:IPS

[18] Prof. J. Guckenheimer. Isochrons and phaseless sets. Journal of Mathematical Biology, 1(3):259–273, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF01273747.

Tyson:1975:CMC

[19] Dr. J. Tyson and S. Kauffman. Control of mitosis by a continuous biochemical oscillation: Synchronization; spatially inhomogeneous oscillations. Journal of Mathematical Biology, 1(4):289–310, ???? 1975. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00279848.

Tyson:1975:EOS

[20] Dr. J. J. Tyson. On the existence of oscillatory solutions in negative feedback cellular control processes. Journal of Mathematical Biology, 1(4):311–315, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00279849.

Heyde:1975:GBB

[21] Dr. C. C. Heyde and Dr. E. Seneta. The genetic balance between random sampling and random population size. Journal of Mathematical Biology, 1(4):317–320, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00279850.

anderHeiden:1975:MES

[22] Dr. U. an der Heiden. On manifolds of equilibria in the selection model for multiple alleles. Journal of Mathematical Biology, 1(4):321–330, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00279851.

Heuch:1975:PCS

[23] I. Heuch. Partial and complete sex linkage in infinite populations. Journal of Mathematical Biology, 1(4):331–343, ???? 1975. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00279852.

Karlin:1975:RTV

[24] Dr. S. Karlin and Dr. U. Liberman. Random temporal variation in selection intensities: One-locus two-allele model. Journal of Mathematical Biology, 2(1):1–17, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276012.

Hadeler:1975:SMF

[25] K. P. Hadeler and U. Liberman. Selection models with fertility differences. Journal of Mathematical Biology, 2(1):19–32, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276013.

Ineichen:1975:GSF

[26] Prof. Robert Ineichen and Prof. Eduard Batschelet. Genetic selection and de Finetti diagrams. Journal of Mathematical Biology, 2(1):33–39, ????

1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276014.

Schafer:1975:NAE

[27] Dr. E. Schäfer. A new approach to explain the “high irradiance responses” of photomorphogenesis on the basis of phytochrome. Journal of Mathematical Biology, 2(1):41–56, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276015.

Dekker:1975:SMM

[28] Dr. H. Dekker. A simple mathematical model of rodent population cycles. Journal of Mathematical Biology, 2(1):57–67, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276016.

Boyarsky:1975:MCM

[29] Dr. A. Boyarsky. A Markov chain model for human granulocyte movement. Journal of Mathematical Biology, 2(1):69–78, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276017.

Gobber:1975:CAS

[30] Dr. F. Göbber and Prof. Dr. F. F. Seelig. Conditions for the application of the steady-state approximation to systems of differential equations. Journal of Mathematical Biology, 2(1):79–86, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276018.

Oguztoreli:1975:AON

[31] Prof. M. N. Öguztöreli and Prof. R. B. Stein. An analysis of oscillations in neuro-muscular systems. Journal of Mathematical Biology, 2 (2):87–105, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275922.

Dorrscheidt:1975:AME

[32] Dr. G. J. Dörrscheidt and Prof. Dr. L. Beck. Advanced methods for evaluating characteristic parameters (τ, α, ρ) of circadian rhythms. Journal of Mathematical Biology, 2(2):107–121, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275923.

Karfunkel:1975:ECR

[33] Dr. H. R. Karfunkel and F. F. Seelig. Excitable chemical reaction systems I. Definition of excitability and simulation of model systems. Journal of Mathematical Biology, 2(2):123–132, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275924.

Othmer:1975:NWP

[34] Dr. H. G. Othmer. Nonlinear wave propagation in reacting systems. Journal of Mathematical Biology, 2(2):133–163, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275925.

Ladde:1975:SMC

[35] Prof. G. S. Ladde and Prof. D. D. Siljak. Stability of multispecies communities in randomly varying environment. Journal of Mathematical Biology, 2(2):165–178, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275926.

Christiansen:1975:SCG

[36] F. B. Christiansen and M. W. Feldman. Selection in complex genetic systems IV. Multiple alleles and interactions between two loci. Journal of Mathematical Biology, 2(2):179–204, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275927.

Rinzel:1975:NST

[37] J. Rinzel. Neutrally stable traveling wave solutions of nerve conduction equations. Journal of Mathematical Biology, 2(3):205–217, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00277150.

Fleming:1975:SMM

[38] W. H. Fleming. A selection–migration model in population genetics. Journal of Mathematical Biology, 2(3):219–233, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00277151.

Hoppensteadt:1975:ASP

[39] F. C. Hoppensteadt. Analysis of a stable polymorphism arising in a selection–migration model in population genetics. Journal of Mathematical Biology, 2(3):235–240, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00277152.

Conley:1975:AWM

[40] C. Conley. An application of Wazewski’s method to a non-linear boundary value problem which arises in population genetics. Journal of Mathematical Biology, 2(3):241–249, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00277153.

Hadeler:1975:TFN

[41] K. P. Hadeler and F. Rothe. Travelling fronts in nonlinear diffusion equations. Journal of Mathematical Biology, 2(3):251–263, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00277154.

Resnikoff:1975:PF

[42] H. L. Resnikoff. On the psychophysical function. Journal of Mathematical Biology, 2(3):265–276, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00277155.

Peskoff:1975:TDP

[43] A. Peskoff and R. S. Eisenberg. The time-dependent potential in a spherical cell using matched asymptotic expansions. Journal of Mathematical Biology, 2(3):277–300, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00277156.

Peskoff:1975:PIS

[44] Dr. A. Peskoff and Mr. D. M. Ramirez. Potential induced in a spherical cell by an intracellular point source and an extracellular point sink. Journal of Mathematical Biology, 2(4):301–316, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00817388.

DeLisi:1975:KHP

[45] Dr. Ch. DeLisi. The kinetics of hemolytic plaque formation V. The influence of geometry on plaque growth. Journal of Mathematical Biology, 2(4):317–331, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00817389.

Abraham-Shrauner:1975:GGC

[46] Dr. Barbara Abraham-Shrauner. Generalized Gouy–Chapman potential of charged phospholipid membranes with divalent cations. Journal of Mathematical Biology, 2(4):333–339, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00817390.

Sleeman:1975:FNA

[47] Dr. B. D. Sleeman. Fitzhugh’s nerve axon equations. Journal of Mathematical Biology, 2(4):341–349, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00817391.

May:1975:NRS

[48] Dr. R. M. May. A note on random sets mosaics. Journal of Mathematical Biology, 2(4):351–357, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00817392.

Worz-Busekros:1975:ZAS

[49] Dr. Angelika Wörz-Busekros. The zygotic algebra for sex linkage II. Journal of Mathematical Biology, 2(4):359–371, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00817393.

Wang:1975:LTA

[50] Dr. F. J. S. Wang. Limit theorems for age and density dependent stochastic population models. Journal of Mathematical Biology, 2(4): 373–400, ???? 1975. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00817394.

Smale:1976:QAD

[51] Prof. S. Smale and Prof. R. F. Williams. The qualitative analysis of a difference equation of population growth. Journal of Mathematical Biology, 3(1):1–4, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00307853.

Smale:1976:DES

[52] Professor S. Smale. On the differential equations of species in competition. Journal of Mathematical Biology, 3(1):5–7, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00307854.

Sowunmi:1976:FDA

[53] Dr. C. O. A. Sowunmi. Female dominant age-dependent deterministic population dynamics. Journal of Mathematical Biology, 3(1):9–17, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00307855.

Levikson:1976:RGR

[54] Dr. B. Levikson. Regulated growth in random environments. Journal of Mathematical Biology, 3(1):19–26, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00307856.

Harman:1976:UFA

[55] H. H. Harman and A. Kocková-Kratochvílová. Use of factor analysis to classify strains of yeasts: Application to genus Torulopsis berlese. Journal of Mathematical Biology, 3(1):27–52, ???? 1976. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00307857.

Othmer:1976:QDC

[56] Dr. H. G. Othmer. The qualitative dynamics of a class of biochemical control circuits. Journal of Mathematical Biology, 3(1):53–78, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00307858.

Koenderink:1976:VPR

[57] Dr. J. J. Koenderink and A. J. van Doorn. Visual perception of rigidity of solid shape. Journal of Mathematical Biology, 3(1):79–85, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00307859.

Oguztoreli:1976:EMR

[58] Professor M. N. Oguztöreli and Dr. R. B. Stein. The effects of multiple reflex pathways on the oscillations in neuro-muscular systems. Journal of Mathematical Biology, 3(1):87–101, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00307860.

Coleman:1976:TRL

[59] B. D. Coleman and G. H. Renninger. Theory of the response of the limulus retina to periodic excitation. Journal of Mathematical Biology, 3(2):103–119, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276201.

Goel:1976:ECH

[60] Dr. Narendra S. Goel and Dr. Martynas Ycas. The error catastrophe hypothesis and aging. Journal of Mathematical Biology, 3(2):121–147, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276202.

Middendorf:1976:MSR

[61] Dipl.-Math. Dr. T. Middendorf and Prof. Dr. Dr. H. H. Loeschcke. Mathematische simulation des respirationssystems. (German) [mathematical simulation of the respiratory system]. Journal of Mathematical Biology, 3(2):149–177, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276203.

Gregorius:1976:CGC

[62] Dr. H.-R. Gregorius. Convergence of genetic compositions assuming infinite population-size and overlapping generations. Journal of Mathematical Biology, 3(2):179–186, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276204.

Walther:1976:TES

[63] Dr. H.-O. Walther. On a transcendental equation in the stability analysis of a population growth model. Journal of Mathematical Biology, 3(2):187–195, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276205.

Hadeler:1976:SSS

[64] Prof. Dr. K. P. Hadeler. On the stability of the stationary state of a population growth equation with time-lag. Journal of Mathematical Biology, 3(2):197–201, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276206.

Rapp:1976:ABP

[65] Dr. P. Rapp. Analysis of biochemical phase shift oscillators by a harmonic balancing technique. Journal of Mathematical Biology, 3(3–4): 203–224, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275057.

Rocklin:1976:CBP

[66] Dr. S. Rocklin and Dr. G. Oster. Competition between phenotypes. Journal of Mathematical Biology, 3(3–4):225–261, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275058.

Hartl:1976:SSB

[67] Dr. D. L. Hartl. Stochastic selection in both haplophase and diplophase. Journal of Mathematical Biology, 3(3–4):263–269, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275059.

Schuh:1976:CEB

[68] Dr. H.-J. Schuh. A condition for the extinction of a branching process with an absorbing lower barrier. Journal of Mathematical Biology, 3 (3–4):271–287, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275060.

Holgate:1976:DPG

[69] P. Holgate. Direct products of genetic algebras and Markov chains. Journal of Mathematical Biology, 3(3–4):289–295, ???? 1976. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275061.

Kobuchi:1976:SPD

[70] Dr. Y. Kobuchi. Signal propagation in 2-dimensional threshold cellular space. Journal of Mathematical Biology, 3(3–4):297–312, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275062.

Goh:1976:GST

[71] Dr. B. S. Goh. Global stability in two species interactions. Journal of Mathematical Biology, 3(3–4):313–318, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275063.

Rothe:1976:CES

[72] Dr. F. Rothe. Convergence to the equilibrium state in the Volterra– Lotka diffusion equations. Journal of Mathematical Biology, 3(3–4): 319–324, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275064.

Perelson:1976:OSI

[73] Alan S. Perelson, Majdedin Mirmirani, and George F. Oster. Optimal strategies in immunology. Journal of Mathematical Biology, 3(3– 4):325–367, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275065.

Cushing:1976:PPI

[74] Dr. J. M. Cushing. Predator prey interactions with time delays. Journal of Mathematical Biology, 3(3–4):369–380, ???? 1976. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275066.

Clark:1976:DRM

[75] Dr. Colin W. Clark. A delayed-recruitment model of population dynamics, with an application to baleen whale populations. Journal of Mathematical Biology, 3(3–4):381–391, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275067.

Worz-Busekros:1976:SDS

[76] Dr. Angelika Wörz-Busekros. Solutions to a degenerate system of parabolic equations from marine biology. Journal of Mathematical Biology, 3(3–4):393–406, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275068.

Anonymous:1976:A

[77] Anonymous. Announcement. Journal of Mathematical Biology, 3(3–4): 407, ???? 1976. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00275069.

Taylor:1977:ERO

[78] B. A. Taylor and Dr. J. D. Murray. Effect of the rate of oxygen consumption on muscle respiration. Journal of Mathematical Biology, 4 (1):1–20, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276348.

Granero:1977:BAP

[79] M. I. Granero, Dr. A. Porati, and D. Zanacca. A bifurcation analysis of pattern formation in a diffusion governed morphogenetic field. Journal of Mathematical Biology, 4(1):21–27, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276349.

Jones:1977:MMS

[80] Dr. B. L. Jones. A model for macromolecular selection in complementary instructing systems. Journal of Mathematical Biology, 4(1):29–34, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276350.

Boyarsky:1977:PPM

[81] Dr. A. Boyarsky. Pattern prediction for moving cells. Journal of Mathematical Biology, 4(1):35–47, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276351.

Carmelli:1977:SPG

[82] Dr. Dorit Carmelli. Some population genetic models combining artificial and natural selection pressures in the presence of assortative mating. Journal of Mathematical Biology, 4(1):49–67, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276352.

Smith:1977:PSD

[83] Dr. H. L. Smith. On periodic solutions of a delay integral equation modelling epidemics. Journal of Mathematical Biology, 4(1):69–80, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276353.

Wevelsiep:1977:PVS

[84] Dr. K. Wevelsiep. Processing visual signals with nonlinear spatial filters. Journal of Mathematical Biology, 4(1):81–99, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276354.

Guckenheimer:1977:DDD

[85] Professor J. Guckenheimer, Professor G. Oster, and Prof. A. Ipaktchi. The dynamics of density dependent population models. Journal of Mathematical Biology, 4(2):101–147, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275980.

Jenkins:1977:SEC

[86] Dr. J. T. Jenkins. Static equilibrium configurations of a model red blood cell. Journal of Mathematical Biology, 4(2):149–169, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275981.

Levine:1977:PCE

[87] Dr. S. H. Levine, Dr F. M. Scudo, and D. J. Plunkett. Persistence and convergence of ecosystems: An analysis of some second order difference equations. Journal of Mathematical Biology, 4(2):171–182, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275982.

Karfunkel:1977:ECR

[88] Dr. H. R. Karfunkel and C. Kahlert. Excitable chemical reaction systems. Journal of Mathematical Biology, 4(2):183–185, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275983.

Jones:1977:SSH

[89] Dr. B. L. Jones. A solvable selfreproductive hypercycle model for the selection of biological molecules. Journal of Mathematical Biology, 4 (2):187–193, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275984.

Estabrook:1977:WTQ

[90] Prof. G. F. Estabrook and Dr. F. R. McMorris. When are two qualitative taxonomic characters compatible? Journal of Mathematical Biology, 4 (2):195–200, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275985.

Anonymous:1977:E

[91] Anonymous. Erratum. Journal of Mathematical Biology, 4(2):201, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00275986.pdf.

FitzHugh:1977:MOV

[92] Dr. R. FitzHugh. A model of optimal voluntary muscular control. Journal of Mathematical Biology, 4(3):203–236, ???? 1977. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00280973.

Grossberg:1977:PFG

[93] Dr. S. Grossberg. Pattern formation by the global limits of a nonlinear competitive interaction in n dimensions. Journal of Mathematical Biology, 4(3):237–256, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00280974.

Cushing:1977:TDS

[94] Dr. J. M. Cushing. Time delays in single species growth models. Journal of Mathematical Biology, 4(3):257–264, ???? 1977. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00280975.

Fisher:1977:SCD

[95] M. E. Fisher and Dr. B. S. Goh. Stability in a class of discrete time models of interacting populations. Journal of Mathematical Biology, 4 (3):265–274, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00280976.

Goh:1977:SGA

[96] Dr. B. S. Goh and T. T. Agnew. Stability in Gilpin and Ayala’s models of competition. Journal of Mathematical Biology, 4(3):275–279, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00280977.

Primas:1977:TRN

[97] Prof. H. Primas. Theory reduction and non-Boolean theories. Journal of Mathematical Biology, 4(3):281–301, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00280978.

Sejnowski:1977:SCN

[98] Dr. T. J. Sejnowski. Storing covariance with nonlinearly interacting neurons. Journal of Mathematical Biology, 4(4):303–321, ???? 1977. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275079.

Aase:1977:CMT

[99] Dr. K. Aase. Conditioned moments of time to fixation. Journal of Mathematical Biology, 4(4):323–326, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275080.

Hall:1977:AMC

[100] Dr. R. L. Hall. Amoeboid movement as a correlated walk. Journal of Mathematical Biology, 4(4):327–335, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275081.

Thieme:1977:MSS

[101] Dr. H. R. Thieme. A model for the spatial spread of an epidemic. Journal of Mathematical Biology, 4(4):337–351, ???? 1977. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275082.

Fife:1977:MPF

[102] P. C. Fife. On modelling pattern formation by activator-inhibitor systems. Journal of Mathematical Biology, 4(4):353–362, ???? 1977. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275083.

Allwright:1977:GSC

[103] Dr. D. J. Allwright. A global stability criterion for simple control loops. Journal of Mathematical Biology, 4(4):363–373, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275084.

Palm:1977:WLS

[104] G. Palm and T. Poggio. Wiener-like system identification in physiology. Journal of Mathematical Biology, 4(4):375–381, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275085.

Oguztoreli:1977:KSM

[105] M. N. Oguztöreli and R. B. Stein. A kinetic study of muscular contractions. Journal of Mathematical Biology, 5(1):1–31, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275803.

Ginzburg:1977:LCP

[106] Lev R. Ginzburg. Local consideration of polymorphisms for populations coexisting in stable ecosystems. Journal of Mathematical Biology, 5(1):33–41, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275804.

Heuch:1977:EFF

[107] Ivar Heuch. An explicit formula for frequency changes in genetic algebras. Journal of Mathematical Biology, 5(1):43–53, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275805.

Lange:1977:CLI

[108] K. Lange and N. Risch. Comments on lack of interference in the four strand model of crossing over. Journal of Mathematical Biology, 5 (1):55–59, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275806.

Sved:1977:MMS

[109] J. A. Sved and B. D. H. Latter. Migration and mutation in stochastic models of gene frequency change. Journal of Mathematical Biology, 5(1):61–73, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275807.

Hsieh:1977:FLD

[110] J. J. Hsieh. A formulation of the lifetime distribution and the existence of its moments. Journal of Mathematical Biology, 5(1):75–86, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275808.

Mann:1977:STG

[111] J. E. Mann and G. L. Curry. A sunfleck theory for general foliage location distributions. Journal of Mathematical Biology, 5(1):87–97, ???? 1977. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275809.

Mees:1978:PMS

[112] A. I. Mees and P. E. Rapp. Periodic metabolic systems: Oscillations in multiple-loop negative feedback biochemical control networks. Journal of Mathematical Biology, 5(2):99–114, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275893.

Stech:1978:ETL

[113] Harlan W. Stech. The effect of time lags on the stability of the equilibrium state of a population growth equation. Journal of Mathematical Biology, 5(2):115–120, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275894.

Jakobsson:1978:FCTa

[114] Eric Jakobsson. A fully coupled transient excited state model for the sodium channel. Journal of Mathematical Biology, 5(2):121–142, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275895.

Siep:1978:MMS

[115] Elke Siep. Molecular model for sodium conductance and calcium transport in the squid axon. Journal of Mathematical Biology, 5(2): 143–168, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275896.

Othmer:1978:ECD

[116] Hans G. Othmer and John A. Aldridge. The effects of cell density and metabolite flux on cellular dynamics. Journal of Mathematical Biology, 5(2):169–200, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275897.

Karlin:1978:TLM

[117] Samuel Karlin and Uri Liberman. The two-locus multi-allele additive viability model. Journal of Mathematical Biology, 5(3):201–211, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276119.

Perelson:1978:OSI

[118] Alan S. Perelson, Majdedin Mirmirani, and George F. Oster. Optimal strategies in immunology. Journal of Mathematical Biology, 5(3):

213–256, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276120.

Lin:1978:QDT

[119] Juan Lin and Peter B. Kahn. Qualitative dynamics of three species predator–prey systems. Journal of Mathematical Biology, 5(3):257– 268, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276121.

Hallam:1978:SSG

[120] Thomas G. Hallam. Structural sensitivity of grazing formulations in nutrient controlled plankton models. Journal of Mathematical Biology, 5(3):269–280, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276122.

Moody:1978:MLC

[121] Michael Moody. A multi-locus continuous-time selection model. Journal of Mathematical Biology, 5(3):281–291, ???? 1978. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276123.

Hader:1978:MMP

[122] Donat-P. Häder and Uhland Burkart. Mathematical model for photophobic accumulations of blue-green algae in light traps. Journal of Mathematical Biology, 5(3):293–304, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276124.

Ashkenazi:1978:SPC

[123] M. Ashkenazi and H. G. Othmer. Spatial patterns in coupled biochemical oscillators. Journal of Mathematical Biology, 5(4):305–350, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276105.

Tyson:1978:ACM

[124] John J. Tyson. On the appearance of chaos in a model of the Belousov reaction. Journal of Mathematical Biology, 5(4):351–362, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276106.

Rinzel:1978:RAH

[125] John Rinzel. Repetitive activity and Hopf bifurcation under pointstimulation for a simple FitzHugh–Nagumo nerve conduction model. Journal of Mathematical Biology, 5(4):363–382, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276107.

Gopalsamy:1978:DMP

[126] K. Gopalsamy. Dynamics of maturing populations and their asymptotic behaviour. Journal of Mathematical Biology, 5(4):383–398, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276108.

Hastings:1978:GST

[127] Alan Hastings. Global stability of two species systems. Journal of Mathematical Biology, 5(4):399–403, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276109.

Worz-Busekros:1978:E

[128] Angelika Wörz-Busekros. Erratum. Journal of Mathematical Biology, 5 (4):405, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00276110.pdf.

Coleman:1978:GPN

[129] Bernard D. Coleman. On the growth of populations with narrow spread in reproductive age. Journal of Mathematical Biology, 6(1):1–19, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF02478513.

Naasell:1978:MMS

[130] Ingemar Nåsell. Mating models for schistosomes. Journal of Mathematical Biology, 6(1):21–35, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF02478514.

Gerald:1978:SLC

[131] K. B. Gerald, J. H. Matis, and H. Kleerekoper. A stochastic locomotor control model for the nurse shark, Ginglymostoma cirratum. Journal of Mathematical Biology, 6(1):37–48, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF02478515.

Maginu:1978:SPT

[132] Kenjiro Maginu. Stability of periodic travelling wave solutions of a nerve conduction equation. Journal of Mathematical Biology, 6(1):49–57, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF02478516.

Lange:1978:CLT

[133] Kenneth Lange. Central limit theorems of pedigrees. Journal of Mathematical Biology, 6(1):59–66, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF02478517.

Hsu:1978:APT

[134] S. B. Hsu. The application of the Poincaré-transform to the Lotka– Volterra model. Journal of Mathematical Biology, 6(1):67–73, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF02478518.

Perelson:1978:SPA

[135] Alan S. Perelson and Lee A. Segel. A singular perturbation approach to diffusion reaction equations containing a point source, with application to the hemolytic plaque assay. Journal of Mathematical Biology, 6(1):75–85, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02478519.

Leung:1978:LBP

[136] Anthony Leung. Limiting behaviour for a prey-predator model with diffusion and crowding effects. Journal of Mathematical Biology, 6 (1):87–93, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02478520.

Winne:1978:PCW

[137] D. Winne. The permeability coefficient of the wall of a villous membrane. Journal of Mathematical Biology, 6(1):95–108, ???? 1978. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF02478521.

Diekmann:1978:TTW

[138] O. Diekmann. Thresholds and travelling waves for the geographical spread of infection. Journal of Mathematical Biology, 6(2):109–130, ????

1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF02450783.

Anderson:1978:SDC

[139] R. F. V. Anderson. Stability of displacement clines arising in allospecies competition. Journal of Mathematical Biology, 6(2):131–144, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF02450784.

Cushing:1978:BPO

[140] J. M. Cushing. Bifurcation of periodic oscillations due to delays in single species growth models. Journal of Mathematical Biology, 6(2): 145–161, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02450785.

Hastings:1978:GSL

[141] Alan Hastings. Global stability in Lotka–Volterra systems with diffusion. Journal of Mathematical Biology, 6(2):163–168, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02450786.

Jones:1978:SPG

[142] B. L. Jones. Some principles governing selection in self-reproducing macromolecular systems. Journal of Mathematical Biology, 6(2):169– 175, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02450787.

Fukshansky:1978:TLA

[143] Leonid Fukshansky. On the theory of light absorption in nonhomogeneous objects. Journal of Mathematical Biology, 6(2):177– 196, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02450788.

Holgate:1978:SGA

[144] P. Holgate. Selfing in genetic algebras. Journal of Mathematical Biology, 6(2):197–206, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02450789.

Glass:1978:SOM

[145] Leon Glass and Joel S. Pasternack. Stable oscillations in mathematical models of biological control systems. Journal of Mathematical Biology, 6(3):207–223, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02547797.

DaRonch:1978:CSC

[146] A. Da Ronch, H. Radu, and V. Caldesi-Valeri. Computer simulation of calcium-induced myotonia. Journal of Mathematical Biology, 6(3): 225–234, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02547798.

Jakobsson:1978:FCTb

[147] Eric Jakobsson. A fully coupled transient excited state model for the sodium channel. Journal of Mathematical Biology, 6(3):235–248, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF02547799.

Notohara:1978:UOT

[148] Morihiro Notohara, Kazushige Ishii, and Hirotsugu Matsuda. Use of orthogonal transformation in population genetics theory. Journal of Mathematical Biology, 6(3):249–263, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF02547800.

Wollkind:1978:TDP

[149] D. J. Wollkind and J. A. Logan. Temperature-dependent predator–prey mite ecosystem on apple tree foliage. Journal of Mathematical Biology, 6(3):265–283, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02547801.

Coffman:1978:GPN

[150] Charles V. Coffman and Bernard D. Coleman. On the growth of populations with narrow spread in reproductive age. Journal of Mathematical Biology, 6(3):285–303, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF02547802.

Carpenter:1978:AMV

[151] Gail A. Carpenter and Victor R. Knapp. An analysis of the mammalian ventricular action potential. Journal of Mathematical Biology, 6(4):305–316, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02462996.

Haken:1978:ATP

[152] H. Haken and H. Olbrich. Analytical treatment of pattern formation in the Gierer–Meinhardt model of morphogenesis. Journal of Mathematical Biology, 6(4):317–331, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF02462997.

Polansky:1978:TDS

[153] Paul J. Polansky. Two-dimensional stochastic exponential growth models. Journal of Mathematical Biology, 6(4):333–342, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02462998.

Heuch:1978:GAP

[154] Ivar Heuch. The genetic algebra for polyploidy with an arbitrary amount of double reduction. Journal of Mathematical Biology, 6(4):343–352, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF02462999.

Worz-Busekros:1978:PAM

[155] Angelika Wörz-Busekros. Polyploidy with an arbitrary mixture of chromosome- and chromatid segregation. Journal of Mathematical Biology, 6(4):353–365, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF02463000.

Freedman:1978:PIG

[156] H. I. Freedman and Paul Waltman. Predator influence on the growth of a population with three genotypes. Journal of Mathematical Biology, 6(4):367–374, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02463001.

Nagylaki:1978:DMG

[157] Thomas Nagylaki. A diffusion model for geographically structured populations. Journal of Mathematical Biology, 6(4):375–382, ???? 1978. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF02463002.

Pohl:1978:TCT

[158] William F. Pohl and George W. Roberts. Topological consideration in the theory of replication of DNA. Journal of Mathematical Biology, 6 (4):383–402, ???? 1978. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02463003.

Oguztoreli:1979:IBC

[159] M. N. Oguztöreli and R. B. Stein. Interactions between centrally and peripherally generated neuromuscular oscillations. Journal of Mathematical Biology, 7(1):1–29, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276411.

Perkel:1979:NRS

[160] Donald H. Perkel and Marcus W. Feldman. Neurotransmitter release statistics: Moment estimates for inhomogeneous Bernoulli trials. Journal of Mathematical Biology, 7(1):31–40, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00276412.

Kernevez:1979:HOP

[161] J. P. Kernevez, G. Joly, M. C. Duban, B. Bunow, and D. Thomas. Hysteresis, oscillations, and pattern formation in realistic immobilized enzyme systems. Journal of Mathematical Biology, 7(1):41–56, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276413.

Kainer:1979:FMR

[162] R. Kainer. A functional model of the rat kidney. Journal of Mathematical Biology, 7(1):57–94, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276414.

Parthasarathy:1979:MMB

[163] P. R. Parthasarathy. On a modified Markov branching process. Journal of Mathematical Biology, 7(1):95–97, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276415.

Milstein:1979:PIC

[164] Jaime Milstein and Hans J. Bremermann. Parameter identification of the Calvin photosynthesis cycle. Journal of Mathematical Biology, 7 (2):99–116, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276924.

Krikorian:1979:VMT

[165] Nishan Krikorian. The Volterra model for three species predator–prey systems: Boundedness and stability. Journal of Mathematical Biology, 7(2):117–132, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276925.

Peil:1979:OML

[166] H. Peil. One- and multi-locus multi-allele selection models in a random environment. Journal of Mathematical Biology, 7(2):133–148, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276926.

Frisch:1979:SMP

[167] H. L. Frisch and I. J. Gotham. A simple model for periodic cyclostat growth of algae. Journal of Mathematical Biology, 7(2):149–169, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276927.

Grasman:1979:MSR

[168] J. Grasman and M. J. W. Jansen. Mutually synchronized relaxation oscillators as prototypes of oscillating systems in biology. Journal of Mathematical Biology, 7(2):171–197, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276928.

Kane:1979:SMX

[169] Victor E. Kane. Stochastic models for X chromosome inactivation. Journal of Mathematical Biology, 7(3):199–218, ???? 1979. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275724.

Erickson:1979:SPN

[170] John W. Erickson and Gary G. Altman. A search for patterns in the nucleotide sequence of the MS2 genome. Journal of Mathematical Biology, 7(3):219–230, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275725.

Cohen:1979:MSI

[171] Donald S. Cohen and S. Rosenblat. Multi-species interactions with hereditary effects and spatial diffusion. Journal of Mathematical Biology, 7 (3):231–241, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275726.

Mimura:1979:SPI

[172] Masayasu Mimura and Yasumasa Nishiura. Spatial patterns for an interaction–diffusion equation in morphogenesis. Journal of Mathematical Biology, 7(3):243–263, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275727.

Ermentrout:1979:TON

[173] G. B. Ermentrout and J. D. Cowan. Temporal oscillations in neuronal nets. Journal of Mathematical Biology, 7(3):265–280, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275728.

Coffman:1979:GPN

[174] Charles V. Coffman and Bernard D. Coleman. On the growth of populations with narrow spread in reproductive age: III. Periodic variations in the environment. Journal of Mathematical Biology, 7(3):281–301, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275729.

Kishimoto:1979:ESL

[175] K. Kishimoto and S. Amari. Existence and stability of local excitations in homogeneous neural fields. Journal of Mathematical Biology, 7(4):303–318, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275151.

Brauer:1979:SRT

[176] F. Brauer and A. C. Soudack. Stability regions and transition phenomena for harvested predator–prey systems. Journal of Mathematical Biology, 7(4):319–337, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275152.

Glass:1979:SMP

[177] Leon Glass and Michael C. Mackey. A simple model for phase locking of biological oscillators. Journal of Mathematical Biology, 7(4): 339–352, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275153.

Karlin:1979:RNS

[178] Samuel Karlin and Uri Liberman. Representation of nonepistatic selection models and analysis of multilocus Hardy–Weinberg equilibrium configurations. Journal of Mathematical Biology, 7(4):353–374, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275154.

Rothe:1979:SAR

[179] Franz Rothe. Some analytical results about a simple reaction–diffusion system for morphogenesis. Journal of Mathematical Biology, 7(4): 375–384, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275155.

Miller:1979:SMD

[180] Robert N. Miller. A simple model of delay, block and one way conduction in Purkinje fibers. Journal of Mathematical Biology, 7(4):385–398, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275156.

Hoofd:1979:NMA

[181] L. Hoofd and F. Kreuzer. A new mathematical approach for solving carrier-facilitated steady-state diffusion problems. Journal of Mathematical Biology, 8(1):1–13, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00280581.

Batschelet:1979:KLH

[182] E. Batschelet, L. Brand, and A. Steiner. On the kinetics of lead in the human body. Journal of Mathematical Biology, 8(1):15–23, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00280582.

Tuljapurkar:1979:LFG

[183] S. D. Tuljapurkar and J. S. Semura. Liapunov functions: Geometry and stability. Journal of Mathematical Biology, 8(1):25–32, ???? 1979.

CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00280583.

Kawasaki:1979:SPF

[184] Kohkichi Kawasaki and Ei Teramoto. Spatial pattern formation of preypredator populations. Journal of Mathematical Biology, 8(1):33–46, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00280584.

West:1979:CPL

[185] G. West. Color perception and the limits of color constancy. Journal of Mathematical Biology, 8(1):47–53, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00280585.

Brauer:1979:SRP

[186] F. Brauer and A. C. Soudack. Stability regions in predator–prey systems with constant-rate prey harvesting. Journal of Mathematical Biology, 8(1):55–71, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00280586.

Moody:1979:PMS

[187] Michael Moody. Polymorphism with migration and selection. Journal of Mathematical Biology, 8(1):73–109, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00280587.

Anonymous:1979:E

[188] Anonymous. Erratum. Journal of Mathematical Biology, 8(1):111, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00280588.pdf.

deHoog:1979:TTG

[189] F. de Hoog, J. Gani, and D. J. Gates. A threshold theorem for the general epidemic in discrete time. Journal of Mathematical Biology, 8 (2):113–121, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00279715.

Nagylaki:1979:MSP

[190] Thomas Nagylaki. Migration–selection polymorphism in dioecious populations. Journal of Mathematical Biology, 8(2):123–131, ???? 1979. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00279716.

Tuljapurkar:1979:SIL

[191] S. D. Tuljapurkar and J. S. Semura. Stochastic instability and Liapunov stability. Journal of Mathematical Biology, 8(2):133–145, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00279717.

Fairen:1979:TPO

[192] V. Fairén and M. G. Velarde. Time-periodic oscillations in a model for the respiratory process of a bacterial culture. Journal of Mathematical Biology, 8(2):147–157, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00279718.

Harrison:1979:GSP

[193] Gary W. Harrison. Global stability of predator–prey interactions. Journal of Mathematical Biology, 8(2):159–171, ???? 1979. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00279719.

Thieme:1979:DDR

[194] Horst R. Thieme. Density-dependent regulation of spatially distributed populations and their asymptotic speed of spread. Journal of Mathematical Biology, 8(2):173–187, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00279720.

Prajneshu:1979:SGM

[195] Prajneshu. A stochastic Gompertz model with hereditary effect. Journal of Mathematical Biology, 8(2):189–196, ???? 1979. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00279721.

Hook:1979:EP

[196] C. Hook and E. Hildebrand. Excitation of paramecium. Journal of Mathematical Biology, 8(2):197–214, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00279722.

Frisch:1979:E

[197] H. L. Frisch and I. J. Gotham. Erratum. Journal of Mathematical Biology, 8(2):215, ???? 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/BF00279723.pdf.

Ludwig:1979:SPS

[198] D. Ludwig, D. G. Aronson, and H. F. Weinberger. Spatial patterning of the spruce budworm. Journal of Mathematical Biology, 8(3): 217–258, October 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276310.

Nichols:1979:MMP

[199] James D. Nichols, Jay B. Hestbeck, and Walt Conley. Mathematical models and population cycles: A critical evaluation of a recent modeling effort. Journal of Mathematical Biology, 8(3):259–263, October 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276311.

DiBlasio:1979:NLA

[200] Gabriella Di Blasio. Non-linear age-dependent population diffusion. Journal of Mathematical Biology, 8(3):265–284, October 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276312.

Sonneveld:1979:CAP

[201] P. Sonneveld and J. van Kan. On a conjecture about the periodic solution of the logistic equation. Journal of Mathematical Biology, 8(3): 285–289, October 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276313.

Dietz:1979:EIV

[202] K. Dietz. Epidemiologic interference of virus populations. Journal of Mathematical Biology, 8(3):291–300, October 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00276314.

Bailey:1979:IMV

[203] Norman T. J. Bailey. Introduction to the modelling of venereal disease. Journal of Mathematical Biology, 8(3):301–322, October 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276315.

Anonymous:1979:I

[204] Anonymous. Information. Journal of Mathematical Biology, 8(4):323, December 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00275829.

Poulsen:1979:MPR

[205] Ebbe Thue Poulsen. A model for population regulation with densityand frequency-dependent selection. Journal of Mathematical Biology, 8 (4):325–343, December 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275830.

anderHeiden:1979:DPS

[206] U. an der Heiden. Delays in physiological systems. Journal of Mathematical Biology, 8(4):345–364, December 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275831.

Wichmann:1979:ABS

[207] H. E. Wichmann. Asymptotic behavior and stability in four models of venereal disease. Journal of Mathematical Biology, 8(4):365– 373, December 1979. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275832.

Gregorius:1979:DSL

[208] Hans-Rolf Gregorius. Deterministic single-locus density-dependent selection. Journal of Mathematical Biology, 8(4):375–391, December 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275833.

Bremermann:1979:UST

[209] Hans J. Bremermann. Unlinked strands as a topological constraint on chromosomal DNA, plasmid integration, and DNA repair. Journal of Mathematical Biology, 8(4):393–401, December 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275834.

Scharstein:1979:IOR

[210] Hans Scharstein. Input-output relationship of the leaky–integrator neuron model. Journal of Mathematical Biology, 8(4):403–420, December 1979. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275835.

Anonymous:1980:I

[211] Anonymous. Information. Journal of Mathematical Biology, 9(1):1, March 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00276031.

Eschenburg:1980:BVS

[212] J.-H. Eschenburg. Is binocular visual space constantly curved? Journal of Mathematical Biology, 9(1):3–22, March 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00276032.

Rosenblat:1980:PMP

[213] S. Rosenblat. Population models in a periodically fluctuating environment. Journal of Mathematical Biology, 9(1):23–36, March 1980. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276033.

Hethcote:1980:IEM

[214] Herbert W. Hethcote and David W. Tudor. Integral equation models for endemic infectious diseases. Journal of Mathematical Biology, 9(1):37–47, March 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276034.

Mimura:1980:SSC

[215] Masayasu Mimura and Kohkichi Kawasaki. Spatial segregation in competitive interaction–diffusion equations. Journal of Mathematical Biology, 9(1):49–64, March 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276035.

Ikeda:1980:LVE

[216] M. Ikeda and D. D. Siljak. Lotka–Volterra equations: Decomposition, stability, and structure. Journal of Mathematical Biology, 9 (1):65–83, March 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276036.

Shigesada:1980:SDD

[217] Nanako Shigesada. Spatial distribution of dispersing animals. Journal of Mathematical Biology, 9(1):85–96, March 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276037.

Anonymous:1980:E

[218] Anonymous. Erratum. Journal of Mathematical Biology, 9(1):97, March 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00276038.pdf.

Anonymous:1980:A

[219] Anonymous. Announcement. Journal of Mathematical Biology, 9(1): 99, March 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00276039.

Nagylaki:1980:SML

[220] Thomas Nagylaki. The strong-migration limit in geographically structured populations. Journal of Mathematical Biology, 9(2):101–114, April 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275916.

Hsu:1980:CMS

[221] S. B. Hsu. A competition model for a seasonally fluctuating nutrient. Journal of Mathematical Biology, 9(2):115–132, April 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275917.

Aoki:1980:CES

[222] Kenichi Aoki. A criterion for the establishment of a stable polymorphism of higher order with an application to the evolution of polymorphism. Journal of Mathematical Biology, 9(2):133–146, April 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275918.

Alt:1980:BRW

[223] Wolgang Alt. Biased random walk models for chemotaxis and related diffusion approximations. Journal of Mathematical Biology, 9 (2):147–177, April 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275919.

Marcati:1980:GAS

[224] Pierangelo Marcati and M. Assunta Pozio. Global asymptotic stability for a vector disease model with spatial spread. Journal of Mathematical Biology, 9(2):179–187, April 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275920.

Karlin:1980:NSE

[225] Samuel Karlin. The number of stable equilibria for the classical onelocus multiallele selection model. Journal of Mathematical Biology, 9 (2):189–192, April 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275921.

Erneux:1980:TPD

[226] T. Erneux and J. Hiernaux. Transition from polar to duplicate patterns. Journal of Mathematical Biology, 9(3):193–211, May 1980. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276025.

Nagasawa:1980:SPE

[227] Masao Nagasawa. Segregation of a population in an environment. Journal of Mathematical Biology, 9(3):213–235, May 1980. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276026.

Nagylaki:1980:ASR

[228] Thomas Nagylaki. Analysis of some regular systems of inbreeding. Journal of Mathematical Biology, 9(3):237–244, May 1980. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276027.

Levin:1980:AAS

[229] Simon A. Levin and C. Phillip Goodyear. Analysis of an age-structured fishery model. Journal of Mathematical Biology, 9(3):245–274, May 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276028.

Kohler:1980:FCP

[230] Hans-Helmut Kohler and Eva Vollmerhaus. The frequency of cyclic processes in biological multistate systems. Journal of Mathematical Biology, 9(3):275–290, May 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276029.

Odell:1980:MME

[231] G. Odell, G. Oster, B. Burnside, and P. Alberch. A mechanical model for epithelial morphogenesis. Journal of Mathematical Biology, 9(3): 291–295, May 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276030.

Fortini:1980:GAT

[232] Peter Fortini and Richard Barakat. Genetic algebras for tetraploidy with several loci. Journal of Mathematical Biology, 9(4):297–304, June 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276495.

Singh:1980:TMS

[233] M. P. Singh, K. Khetarpal, and Mathili Sharan. A theoretical model for studying the rate of oxygenation of blood in pulmonary capillaries. Journal of Mathematical Biology, 9(4):305–330, June 1980. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276496.

Sinestrari:1980:NLA

[234] Eugenio Sinestrari. Non-linear age-dependent population growth. Journal of Mathematical Biology, 9(4):331–345, June 1980. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276497.

Hook:1980:EPS

[235] C. Hook and E. Hildebrand. Excitability of paramecium and the significance of negative surface charges. Journal of Mathematical Biology, 9(4):347–360, June 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276498.

Sanchez:1980:LAD

[236] David A. Sánchez. Linear age-dependent population growth with seasonal harvesting. Journal of Mathematical Biology, 9(4):361–368, June 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276499.

Funakoshi:1980:TPB

[237] Hiroumi Funakoshi and Akio Yamada. Transition phenomena in bacterial growth between logarithmic and stationary phases. Journal of Mathematical Biology, 9(4):369–387, June 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276500.

Alberghina:1980:ACC

[238] Lilia Alberghina and Luigi Mariani. Analysis of a cell cycle model for Escherichia coli. Journal of Mathematical Biology, 9(4):389–398, June 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276501.

Steinbach:1980:MMC

[239] K. H. Steinbach, H. Raffler, G. Pabst, and T. M. Fliedner. A mathematical model of canine granulocytopoiesis. Journal of Mathematical Biology, 10(1):1–12, August 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276392.

Busenberg:1980:EIC

[240] Stavros Busenberg and Kenneth L. Cooke. The effect of integral conditions in certain equations modelling epidemics and population growth. Journal of Mathematical Biology, 10(1):13–32, August 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276393.

Criminale:1980:EMC

[241] William O. Criminale. Effects of mean current and shear on stability of depth distributions of marine phytoplankton. Journal of Mathematical Biology, 10(1):33–51, August 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276394.

Gonzalez-Guzman:1980:MPP

[242] Jorge Gonzalez-Guzman. A mixed program for parasitic disease control. Journal of Mathematical Biology, 10(1):53–64, August 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276395.

Harrison:1980:PPP

[243] Gary W. Harrison. Persistence of predator–prey systems in an uncertain environment. Journal of Mathematical Biology, 10(1):65–77, August 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276396.

Rakitzis:1980:KAB

[244] Emmanuel T. Rakitzis. Kinetic analysis of biphasic protein modification reactions. Journal of Mathematical Biology, 10(1):79–87, August 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276397.

Lin:1980:AMD

[245] Juan Lin and Peter B. Kahn. Averaging methods in the delayed-logistic equation. Journal of Mathematical Biology, 10(1):89–96, August 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276398.

Ermentrout:1980:ODT

[246] G. Bard Ermentrout and John Rinzel. One-dimensional λ–ω target patterns: Empirical stability tests. Journal of Mathematical Biology, 10 (1):97–100, August 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276399.

Lin:1980:REP

[247] Juan Lin and Peter B. Kahn. Random effects in population models with hereditary effects. Journal of Mathematical Biology, 10(2): 101–112, October 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275836.

Bailey:1980:SAM

[248] Norman T. J. Bailey and J. Duppenthaler. Sensitivity analysis in the modelling of infectious disease dynamics. Journal of Mathematical Biology, 10(2):113–131, October 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275837.

Maginu:1980:ESP

[249] Kenjiro Maginu. Existence and stability of periodic travelling wave solutions to Nagumo’s nerve equation. Journal of Mathematical Biology, 10 (2):133–153, October 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275838.

Ewens:1980:FSN

[250] W. J. Ewens and W.-H. Li. Frequency spectra of neutral and deleterious alleles in a finite population. Journal of Mathematical Biology, 10(2): 155–166, October 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275839.

Kesten:1980:NDA

[251] Harry Kesten. The number of distinguishable alleles according to the Ohta–Kimura model of neutral mutation. Journal of Mathematical Biology, 10(2):167–187, October 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275840.

Karlin:1980:RSP

[252] Samuel Karlin. The range of stability of a polymorphic linkage equilibrium state in a two-locus two-allele selection model. Journal of Mathematical Biology, 10(2):189–194, October 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275841.

Griffiths:1980:GIB

[253] R. C. Griffiths. Genetic identity between populations when mutation rates vary within and across loci. Journal of Mathematical Biology, 10 (2):195–204, October 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275842.

Buhler:1980:IBD

[254] W. J. Bühler and B. Mellein. Integrals of birth and death processes. Journal of Mathematical Biology, 10(2):205–207, October 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275843.

Perelson:1980:OSI

[255] Alan S. Perelson, Byron Goldstein, and Sol Rocklin. Optimal strategies in immunology III. The IgM–IgG switch. Journal of Mathematical Biology, 10(3):209–256, November 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276984.

Burkart:1980:PAL

[256] Uhland Burkart and Donat-P. Häder. Phototactic attraction in light trap experiments: A mathematical model. Journal of Mathematical Biology, 10(3):257–269, November 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276985.

Gripenberg:1980:PSE

[257] Gustaf Gripenberg. Periodic solutions of an epidemic model. Journal of Mathematical Biology, 10(3):271–280, November 1980. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276986.

Notohara:1980:CGU

[258] Morihiro Notohara and Tokuzo Shiga. Convergence to genetically uniform state in stepping stone models of population genetics. Journal of Mathematical Biology, 10(3):281–294, November 1980. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276987.

Namba:1980:ABS

[259] Toshiyuki Namba. Asymptotic behaviour of solutions of the diffusive Lotka–Volterra equations. Journal of Mathematical Biology, 10(3):295– 303, November 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276988.

Stark:1980:ISC

[260] Alan E. Stark. Inbreeding systems: Classification by a canonical form. Journal of Mathematical Biology, 10(3):305, November 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/BF00276989.

Hadeler:1980:DMA

[261] K. P. Hadeler, P. de Mottoni, and K. Schumacher. Dynamic models for animal orientation. Journal of Mathematical Biology, 10(4): 307–332, December 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276092.

Lasota:1980:ESE

[262] A. Lasota and Michael C. Mackey. The extinction of slowly evolving dynamical systems. Journal of Mathematical Biology, 10(4):333– 345, December 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276093.

Goldsten:1980:TEB

[263] Byron Goldsten and Carla Wofsy. Theory of equilibrium binding of a bivalent ligand to cell surface antibody: The effect of antibody heterogeneity on cross-linking. Journal of Mathematical Biology, 10(4): 347–366, December 1980. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276094.

Estabrook:1980:WOE

[264] G. F. Estabrook and F. R. McMorris. When is one estimate of evolutionary relationships a refinement of another? Journal of Mathematical Biology, 10(4):367–373, December 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276095.

Dai:1980:ELC

[265] Lo-Sheng Dai. On the existence of limit cycles for the enzyme-catalyzed hydrolysis reaction in a zero-dimensional representation of a membrane. Journal of Mathematical Biology, 10(4):375–384, December 1980. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276096.

Cushing:1980:TSC

[266] J. M. Cushing. Two species competition in a periodic environment. Journal of Mathematical Biology, 10(4):385–400, December 1980. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276097.

Takeuchi:1980:EGS

[267] Yasuhiro Takeuchi and Norihiko Adachi. The existence of globally stable equilibria of ecosystems of the generalized Volterra type. Journal of Mathematical Biology, 10(4):401–415, December 1980. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276098.

Brauer:1981:CRS

[268] F. Brauer and A. C. Soudack. Constant-rate stocking of predator–prey systems. Journal of Mathematical Biology, 11(1):1–14, January 1981.

CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275820.

Plant:1981:BRM

[269] Richard E. Plant. Bifurcation and resonance in a model for bursting nerve cells. Journal of Mathematical Biology, 11(1):15–32, January 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275821.

Ziehe:1981:PTS

[270] Martin Ziehe. Population trajectories for single locus additive fecundity selection and related selection models. Journal of Mathematical Biology, 11(1):33–43, January 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275822.

Engels:1981:ETN

[271] William R. Engels. On estimating the total number of genes of a given kind in the genome and similar problems. Journal of Mathematical Biology, 11(1):45–50, January 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275823.

Hastings:1981:MLC

[272] Alan Hastings. Multiple limit cycles in predator–prey models. Journal of Mathematical Biology, 11(1):51–63, January 1981. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275824.

Pruss:1981:ESA

[273] Jan Prüß. Equilibrium solutions of age-specific population dynamics of several species. Journal of Mathematical Biology, 11(1):65–84, January 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275825.

Allwright:1981:HBS

[274] D. J. Allwright. The Hopf bifurcation in some biochemical control loops. Journal of Mathematical Biology, 11(1):85–93, January 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275826.

Stech:1981:SCC

[275] Harlan Stech and Michael Williams. Stability in a class of cyclic epidemic models with delay. Journal of Mathematical Biology, 11(1): 95–103, January 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275827.

Hastings:1981:PSP

[276] Stuart Hastings. Persistent spatial patterns for semi-discrete models of excitable media. Journal of Mathematical Biology, 11(1):105– 117, January 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275828.

Matsuda:1981:SGF

[277] H. Matsuda and K. Ishii. Stationary gene frequency distribution in the environment fluctuating between two distinct states. Journal of Mathematical Biology, 11(2):119–141, February 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275437.

Paletta:1981:SIC

[278] B. Paletta, W. Estelberger, and G. Porod. System identification of the control of gluconeogenesis in the intact organism under conditions of long-term stress. Journal of Mathematical Biology, 11(2): 143–153, February 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275438.

Hofbauer:1981:CCC

[279] J. Hofbauer, P. Schuster, and K. Sigmund. Competition and cooperation in catalytic selfreplication. Journal of Mathematical Biology, 11 (2):155–168, February 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275439.

Gillespie:1981:TPB

[280] John H. Gillespie. The transient properties of balancing selection in large finite populations. Journal of Mathematical Biology, 11(2):169– 180, February 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275440.

Gross:1981:DIL

[281] Louis J. Gross. On the dynamics of internal leaf carbon dioxide uptake. Journal of Mathematical Biology, 11(2):181–191, February 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275441.

Sawyer:1981:CMM

[282] Stanley Sawyer and Joseph Felsenstein. A continuous migration model with stable demography. Journal of Mathematical Biology, 11(2):193– 205, February 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275442.

Pauwelussen:1981:CPA

[283] J. P. Pauwelussen and L. A. Peletier. Clines in the presence of asymmetric migration. Journal of Mathematical Biology, 11(2):207–233, February 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275443.

Notohara:1981:EKB

[284] Morihiro Notohara. Eigenanalysis for the Kolmogorov backward equation for the neutral multi-allelic model. Journal of Mathematical Biology, 11 (3):235–244, March 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276895.

Moody:1981:PSG

[285] Michael Moody. Polymorphism with selection and genotype-dependent migration. Journal of Mathematical Biology, 11(3):245–267, March 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276896.

Ermentrout:1981:WSE

[286] G. Bard Ermentrout and John Rinzel. Waves in a simple, excitable or oscillatory, reaction–diffusion model. Journal of Mathematical Biology, 11(3):269–294, March 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276897.

Livshits:1981:PDP

[287] M. A. Livshits, G. T. Gurija, B. N. Belintsev, and M. V. Volkenstein. Positional differentiation as pattern formation in reaction–diffusion systems with permeable boundaries. Bifurcation analysis. Journal of Mathematical Biology, 11(3):295–310, March 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276898.

Saunders:1981:EC

[288] I. W. Saunders. Epidemics in competition. Journal of Mathematical Biology, 11(3):311–318, March 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276899.

deMottoni:1981:CSP

[289] P. de Mottoni and A. Schiaffino. Competition systems with periodic coefficients: A geometric approach. Journal of Mathematical Biology, 11 (3):319–335, March 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276900.

Brill:1981:CTI

[290] Michael Brill and Gerhard West. Contributions to the theory of invariance of color under the condition of varying illumination. Journal of Mathematical Biology, 11(3):337–350, March 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00276901.

Anonymous:1981:A

[291] Anonymous. Announcement. Journal of Mathematical Biology, 11(3): 350, March 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00276902.

Rand:1982:HBS

[292] R. H. Rand, S. K. Upadhyaya, J. R. Cooke, and D. W. Storti. Hopf bifurcation in a stomatal oscillator. Journal of Mathematical Biology, 12(1):1–11, May 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275199.

Kawato:1982:TSS

[293] M. Kawato. Transient and steady state phase response curves of limit cycle oscillators. Journal of Mathematical Biology, 12(1):13–30, May 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275200.

Whitmore:1982:CAC

[294] Roy W. Whitmore and J. H. Matis. Compartmental analysis of carcinogenic experiments: Formulation of a stochastic model. Journal of Mathematical Biology, 12(1):31–43, May 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275201.

Hsu:1982:RBE

[295] Sze-Bi Hsu. On a resource based ecological competition model with interference. Journal of Mathematical Biology, 12(1):45–52, May 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275202.

Tier:1982:ANA

[296] Charles Tier. An analysis of neutral-alleles and variable-environment diffusion models. Journal of Mathematical Biology, 12(1):53–71, May 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275203.

Nedelman:1982:FDO

[297] J. Nedelman and S. I. Rubinow. Facilitated diffusion of oxygen and carbon monoxide in the large affinity regime. Journal of Mathematical Biology, 12(1):73–90, May 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275204.

Heyde:1982:SGR

[298] C. C. Heyde. On the survival of a gene represented in a founder population. Journal of Mathematical Biology, 12(1):91–99, May 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275205.

Brauer:1982:CPS

[299] F. Brauer and A. C. Soudack. Coexistence properties of some predator– prey systems under constant rate harvesting and stocking. Journal of Mathematical Biology, 12(1):101–114, May 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275206.

Cheng:1982:SRG

[300] Kuo-Shung Cheng, Sze-Bi Hsu, and Song-Sun Lin. Some results on global stability of a predator–prey system. Journal of Mathematical Biology, 12(1):115–126, May 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275207.

Bardi:1982:PPM

[301] Martino Bardi. predator–prey models in periodically fluctuating environments. Journal of Mathematical Biology, 12(1):127–140, May 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275208.

Blom:1981:FPP

[302] J. G. Blom, R. de Bruin, J. Grasman, and J. G. Verwer. Forced prey-predator oscillations. Journal of Mathematical Biology, 12(2): 141–152, June 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276125.

Lam:1981:MDR

[303] H. S. Lam and D. G. Lampard. Modelling of drug receptor interaction with birth and death processes. Journal of Mathematical Biology, 12 (2):153–172, June 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276126.

Cornette:1981:DGM

[304] James L. Cornette. Deterministic genetic models in varying environments. Journal of Mathematical Biology, 12(2):173–186, June 1981. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276127.

Seelig:1981:UHB

[305] Friedrich Franz Seelig. Unrestricted harmonic balance II. Application to stiff ordinary differential equations in enzyme catalysis. Journal of Mathematical Biology, 12(2):187–198, June 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276128.

Erneux:1981:SRC

[306] T. Erneux. Stability of rotating chemical waves. Journal of Mathematical Biology, 12(2):199–214, June 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276129.

Keener:1981:CAA

[307] James P. Keener. On cardiac arrythmias: AV conduction block. Journal of Mathematical Biology, 12(2):215–225, June 1981. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276130.

Greenspan:1981:CMA

[308] Donald Greenspan. A classical molecular approach to computer simulation of biological sorting. Journal of Mathematical Biology, 12(2): 227–235, June 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276131.

Cohen:1981:GDM

[309] Donald S. Cohen and James D. Murray. A generalized diffusion model for growth and dispersal in a population. Journal of Mathematical Biology, 12(2):237–249, June 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276132.

Griffiths:1981:NHL

[310] R. C. Griffiths. The number of heterozygous loci between two randomly chosen completely linked sequences of loci in two subdivided population models. Journal of Mathematical Biology, 12(2):251–261, June 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276133.

Levin:1981:AAS

[311] Simon A. Levin and C. Phillip Goodyear. Analysis of an age-structured fishery model. Journal of Mathematical Biology, 12(2):263, June 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/BF00276134.

Botsford:1981:OFP

[312] Louis W. Botsford. Optimal fishery policy for size-specific, densitydependent population models. Journal of Mathematical Biology, 12(3): 265–293, August 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276917.

Butler:1981:BLC

[313] G. J. Butler and Paul Waltman. Bifurcation from a limit cycle in a two predator-one prey ecosystem modeled on a chemostat. Journal of Mathematical Biology, 12(3):295–310, August 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276918.

Shigesada:1981:ASS

[314] Nanako Shigesada and Akira Okubo. Analysis of the self-shading effect on algal vertical distribution in natural waters. Journal of Mathematical Biology, 12(3):311–326, August 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276919.

Ermentrout:1981:PLW

[315] G. Bard Ermentrout. n: m Phase-locking of weakly coupled oscillators. Journal of Mathematical Biology, 12(3):327–342, August 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276920.

Coleman:1981:OIG

[316] Bernard D. Coleman. On optimal intrinsic growth rates for populations in periodically changing environments. Journal of Mathematical Biology, 12(3):343–354, August 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276921.

Eshel:1981:SPS

[317] Ilan Eshel. On the survival probability of a slightly advantageous mutant gene with a general distribution of progeny size — a branching process model. Journal of Mathematical Biology, 12(3):355–362, August 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276922.

Merrill:1981:MRN

[318] Stephen J. Merrill. A model of the role of natural killer cells in immune surveillance — I. Journal of Mathematical Biology, 12(3): 363–373, August 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276923.

Baverstock:1981:WBS

[319] P. R. Baverstock and S. Elhay. Water balance of small lactating rodents — III. Estimates of milk production and water recycling in lactating Mus musculus under various water regimes. Journal of Mathematical Biology, 13(1):1–22, October 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276863.

Fowler:1981:ASM

[320] A. C. Fowler. Approximate solution of a model of biological immune responses incorporating delay. Journal of Mathematical Biology, 13 (1):23–45, October 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276864.

Turke:1981:APR

[321] Bernhard Türke. Analysis of pattern recognition by man using detection experiments. Journal of Mathematical Biology, 13(1):47–65, October 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276865.

Klein:1981:MMB

[322] P. Klein, J. Sterzl, and J. Dolezal. A mathematical model of B lymphocyte differentiation: Control by antigen. Journal of Mathematical Biology, 13(1):67–86, October 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276866.

Flaam:1981:SSP

[323] Sjur D. Flåm. Steady states in population models with monotone, stochastic dynamics. Journal of Mathematical Biology, 13(1): 87–93, October 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276867.

Latter:1981:MMS

[324] B. D. H. Latter and J. A. Sved. Migration and mutation in stochastic models of gene frequency change. II. Stochastic migration with a finite number of islands. Journal of Mathematical Biology, 13(1): 95–104, October 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276868.

Kishimoto:1981:INC

[325] Kazuo Kishimoto. Instability of non-constant equilibrium solutions of a system of competition–diffusion equations. Journal of Mathematical Biology, 13(1):105–114, October 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276869.

Turelli:1981:TVS

[326] Michael Turelli. Temporally varying selection on multiple alleles: A diffusion analysis. Journal of Mathematical Biology, 13(1):115–129, October 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276870.

Gurtin:1981:OHP

[327] Morton E. Gurtin and Lea F. Murphy. On the optimal harvesting of persistent age-structured populations. Journal of Mathematical Biology, 13 (2):131–148, December 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275209.

Lasota:1981:MTI

[328] A. Lasota, Michael C. Mackey, and Maria Wazewska-Czyzewska. Minimizing therapeutically induced anemia. Journal of Mathematical Biology, 13(2):149–158, December 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275210.

Gates:1981:NSN

[329] David J. Gates and Mark Westcott. Negative skewness and negative correlations in spatial competition models. Journal of Mathematical Biology, 13(2):159–171, December 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275211.

Capasso:1981:CES

[330] V. Capasso and L. Maddalena. Convergence to equilibrium states for a reaction–diffusion system modelling the spatial spread of a class of bacterial and viral diseases. Journal of Mathematical Biology, 13(2): 173–184, December 1981. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275212.

Hethcote:1981:SAM

[331] Herbert W. Hethcote, Harlan W. Stech, and P. van den Driessche. Stability analysis for models of diseases without immunity. Journal of Mathematical Biology, 13(2):185–198, December 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275213.

Campbell:1981:EVE

[332] R. B. Campbell. The effect of variable environments on polymorphism at loci with several alleles I. A symmetric model. Journal of Mathematical Biology, 13(2):199–208, December 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275214.

Gimelfarb:1981:GLM

[333] A. Gimelfarb. A general linear model for the genotypic covariance between relatives under assortative mating. Journal of Mathematical Biology, 13(2):209–226, December 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275215.

Gimelfarb:1981:ANR

[334] A. Gimelfarb. Analysis of “nontraditional” relationships under assortative mating. Journal of Mathematical Biology, 13(2):227–240, December 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275216.

Tuljapurkar:1981:PCS

[335] Shripad D. Tuljapurkar. Primitivity and convergence to stability. Journal of Mathematical Biology, 13(2):241–246, December 1981. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275217.

Miura:1982:ACS

[336] Robert M. Miura. Accurate computation of the stable solitary wave for the FitzHugh–Nagumo equations. Journal of Mathematical Biology, 13 (3):247–269, January 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276063.

Hunt:1982:RPC

[337] Fern Hunt. Regulation of population cycles by genetic feedback: Existence of periodic solutions of a mathematical model. Journal of Mathematical Biology, 13(3):271–282, January 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276064.

Elderkin:1982:SDP

[338] Richard H. Elderkin. Seed dispersal in a patchy environment with global age dependence. Journal of Mathematical Biology, 13(3):283– 303, January 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276065.

Akin:1982:CSG

[339] Ethan Akin. Cycling in simple genetic systems. Journal of Mathematical Biology, 13(3):305–324, January 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276066.

Tuljapurkar:1982:WUP

[340] Shripad D. Tuljapurkar. Why use population entropy? it determines the rate of convergence. Journal of Mathematical Biology, 13 (3):325–337, January 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276067.

Shukla:1982:MFW

[341] V. P. Shukla and J. B. Shukla. Multispecies food webs with diffusion. Journal of Mathematical Biology, 13(3):339–344, January 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276068.

Cohen:1982:NCB

[342] Avis H. Cohen, Philip J. Holmes, and Richard H. Rand. The nature of the coupling between segmental oscillators of the lamprey spinal generator for locomotion: A mathematical model. Journal of Mathematical Biology, 13(3):345–369, January 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276069.

DiBlasio:1982:AEA

[343] Gabriella Di Blasio, Mimmo Iannelli, and Eugenio Sinestrari. Approach to equilibrium in age structured populations with an increasing recruitment process. Journal of Mathematical Biology, 13(3):371–382, January 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276070.

Lin:1982:PAI

[344] Juan Lin and Peter B. Kahn. Phase and amplitude instability in delaydiffusion population models. Journal of Mathematical Biology, 13(3): 383–393, January 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276071.

Guevara:1982:PLP

[345] Michael R. Guevara and Leon Glass. Phase locking, period doubling bifurcations and chaos in a mathematical model of a periodically driven oscillator: A theory for the entrainment of biological oscillators and the generation of cardiac dysrhythmias. Journal of Mathematical Biology, 14(1):1–23, March 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02154750.

Butler:1982:GDS

[346] G. J. Butler, H. I. Freedman, and Paul Waltman. Global dynamics of a selection model for the growth of a population with genotypic fertility differences. Journal of Mathematical Biology, 14(1):25–35, March 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF02154751.

Asmussen:1982:RDL

[347] Marjorie A. Asmussen and Michael T. Clegg. Rates of decay of linkage disequilibrium under two-locus models of selection. Journal of Mathematical Biology, 14(1):37–70, March 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF02154752.

Roberts:1982:STH

[348] M. G. Roberts. Stability in a two host epidemic model. Journal of Mathematical Biology, 14(1):71–75, March 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF02154753.

Garroni:1982:ADP

[349] Maria Giovanna Garroni and Michel Langlais. Age-dependent population diffusion with external constraint. Journal of Mathematical Biology, 14 (1):77–94, March 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02154754.

Badii:1982:ABP

[350] M. Badii and A. Schiaffino. Asymptotic behaviour of positive solutions of periodic delay logistic equations. Journal of Mathematical Biology, 14(1):95–100, March 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02154755.

Bailey:1982:SSE

[351] Norman T. J. Bailey. The structural simplification of an epidemiological compartment model. Journal of Mathematical Biology, 14(1): 101–116, March 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02154756.

Diekmann:1982:PHB

[352] O. Diekmann and R. Montijn. Prelude to Hopf bifurcation in an epidemic model: Analysis of a characteristic equation associated with a nonlinear Volterra integral equation. Journal of Mathematical Biology, 14(1):117–127, March 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02154757.

Anonymous:1982:Aa

[353] Anonymous. Announcement. Journal of Mathematical Biology, 14(1): 129, March 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF02154758.

Anonymous:1982:CP

[354] Anonymous. Call for papers. Journal of Mathematical Biology, 14(1): 131, March 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF02154759.

Berding:1982:PFM

[355] C. Berding and H. Haken. Pattern formation in morphogenesis. Journal of Mathematical Biology, 14(2):133–151, ???? 1982. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF01832840.

Arneodo:1982:SAV

[356] A. Arneodo, P. Coullet, J. Peyraud, and C. Tresser. Strange attractors in Volterra equations for species in competition. Journal of Mathematical Biology, 14(2):153–157, ???? 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF01832841.

Conley:1982:CMT

[357] C. Conley and P. Fife. Critical manifolds, travelling waves, and an example from population genetics. Journal of Mathematical Biology, 14 (2):159–176, ???? 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF01832842.

Webb:1982:RRE

[358] G. F. Webb. A recovery-relapse epidemic model with spatial diffusion. Journal of Mathematical Biology, 14(2):177–194, ???? 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01832843.

Weiss:1982:CPP

[359] George H. Weiss and John Rice. A combinatorial problem in pharmacology. Journal of Mathematical Biology, 14(2):195–201, ???? 1982. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF01832844.

Martonen:1982:TBS

[360] T. B. Martonen and A. F. Wilson. Theoretical basis of single breath gas absorption tests. Journal of Mathematical Biology, 14(2):203–220, ???? 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF01832845.

Tuljapurkar:1982:PDV

[361] Shripad D. Tuljapurkar. Population dynamics in variable environments. IV. Weak ergodicity in the Lotka equation. Journal of Mathematical Biology, 14(2):221–230, ???? 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF01832846.

Cushing:1982:PPM

[362] J. M. Cushing and M. Saleem. A predator prey model with age structure. Journal of Mathematical Biology, 14(2):231–250, ???? 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01832847.

Anonymous:1982:Ab

[363] Anonymous. Announcement. Journal of Mathematical Biology, 14(2): 251, ???? 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF01832848.

Levin:1982:MIP

[364] Simon A. Levin and Lee A. Segel. Models of the influence of predation on aspect diversity in prey populations. Journal of Mathematical Biology, 14(3):253–284, July 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275393.

Gard:1982:TPP

[365] Thomas C. Gard. Top predator persistence in differential equation models of food chains: The effects of omnivory and external forcing of lower trophic levels. Journal of Mathematical Biology, 14(3):285–299, July 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275394.

Lenz:1982:PCC

[366] Norbert Lenz. Poisson convergence on continuous time branching random walks and multistage carcinogenesis. Journal of Mathematical Biology, 14(3):301–307, July 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275395.

Sleeman:1982:SAP

[367] B. D. Sleeman. Small amplitude periodic waves for the FitzHugh– Nagumo equations. Journal of Mathematical Biology, 14(3):309– 325, July 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275396.

Ishii:1982:MMB

[368] K. Ishii, H. Matsuda, and N. Ogita. A mathematical model of biological evolution. Journal of Mathematical Biology, 14(3):327–353, July 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275397.

Daley:1982:FRS

[369] D. J. Daley, Peter Hall, and C. C. Heyde. Further results on the survival of a gene represented in a founder population. Journal of Mathematical Biology, 14(3):355–363, July 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275398.

Macken:1982:ACS

[370] Catherine A. Macken and Alan S. Perelson. Aggregation of cell surface receptors by multivalent ligands. Journal of Mathematical Biology, 14 (3):365–370, July 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275399.

Chesson:1982:SER

[371] Peter L. Chesson. The stabilizing effect of a random environment. Journal of Mathematical Biology, 15(1):1–36, September 1982. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275786.

Yanagida:1982:SSD

[372] Eiji Yanagida. Stability of stationary distributions in a space-dependent population growth process. Journal of Mathematical Biology, 15(1): 37–50, September 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275787.

Dibrov:1982:DSS

[373] B. F. Dibrov, A. M. Zhabotinsky, and B. N. Kholodenko. Dynamic stability of steady states and static stabilization in unbranched metabolic pathways. Journal of Mathematical Biology, 15(1):51–63, September 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275788.

Kemper:1982:EEE

[374] John T. Kemper. The evolutionary effect of endemic infectious disease: Continuous models for an invariant pathogen. Journal of Mathematical Biology, 15(1):65–77, September 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275789.

Notohara:1982:LMN

[375] Morihiro Notohara. The lattice models of neutral multi-alleles in population genetics theory. Journal of Mathematical Biology, 15(1):79– 92, September 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275790.

Schiaffino:1982:CSD

[376] A. Schiaffino and A. Tesei. Competition systems with Dirichlet boundary conditions. Journal of Mathematical Biology, 15(1):93–105, September 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275791.

Iizuka:1982:WCD

[377] Masaru Iizuka and Hirotsugu Matsuda. Weak convergence of discrete time non-Markovian processes related to selection models in population genetics. Journal of Mathematical Biology, 15(1):107–127, September 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275792.

Anonymous:1982:Ac

[378] Anonymous. Announcements. Journal of Mathematical Biology, 15(1): 129–130, September 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00275793.

Rand:1982:DCS

[379] R. H. Rand, D. W. Storti, S. K. Upadhyaya, and J. R. Cooke. Dynamics of coupled stomatal oscillators. Journal of Mathematical Biology, 15 (2):131–149, October 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275070.

Pauwelussen:1982:OWT

[380] Joop Pauwelussen. One way traffic of pulses in a neuron. Journal of Mathematical Biology, 15(2):151–171, October 1982. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275071.

Franzone:1982:MMI

[381] P. Colli Franzone, A. Paganuzzi, and M. Stefanelli. A mathematical model of iron metabolism. Journal of Mathematical Biology, 15(2): 173–201, October 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275072.

Hutson:1982:PSO

[382] V. Hutson and W. Moran. Persistence of species obeying difference equations. Journal of Mathematical Biology, 15(2):203–213, October 1982.

CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275073.

Marcati:1982:GSL

[383] P. Marcati. On the global stability of the logistic age-dependent population growth. Journal of Mathematical Biology, 15(2):215–226, October 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275074.

OBrien:1982:AFM

[384] Paul O’Brien. Allele frequencies in a multidimensional Wright–Fisher model with general mutation. Journal of Mathematical Biology, 15(2): 227–237, October 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275075.

Noonburg:1982:EBA

[385] V. W. Noonburg. Effects of behavioral adaptation on a predator–prey model. Journal of Mathematical Biology, 15(2):239–247, October 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275076.

West:1982:NSC

[386] Gerhard West and Michael H. Brill. Necessary and sufficient conditions for von kries chromatic adaptation to give color constancy. Journal of Mathematical Biology, 15(2):249–258, October 1982. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275077.

Oster:1982:MME

[387] George F. Oster, Alan S. Perelson, and Lewis G. Tilney. A mechanical model for elongation of the acrosomal process in thyone sperm. Journal of Mathematical Biology, 15(2):259–265, October 1982. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275078.

Oja:1982:SNM

[388] Erkki Oja. Simplified neuron model as a principal component analyzer. Journal of Mathematical Biology, 15(3):267–273, November 1982. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275687.

Dougherty:1982:GDM

[389] Geoffrey Dougherty. A Gaussian disorientation model for interpreting linear dichroism measurements of DNA-drug fibres and films. Journal of Mathematical Biology, 15(3):275–291, November 1982. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275688.

Campbell:1982:EVE

[390] R. B. Campbell. The effect of variable environments on polymorphism at loci with several alleles II. Submultiplicative viabilities. Journal of Mathematical Biology, 15(3):293–303, November 1982. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275689.

Weiss:1982:MPT

[391] M. Weiss. Moments of physiological transit time distributions and the time course of drug disposition in the body. Journal of Mathematical Biology, 15(3):305–318, November 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275690.

Volz:1982:GAS

[392] Rudolf Volz. Global asymptotic stability of a periodic solution to an epidemic model. Journal of Mathematical Biology, 15(3):319–338, November 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275691.

Hoppensteadt:1982:PLB

[393] F. C. Hoppensteadt and J. P. Keener. Phase locking of biological clocks. Journal of Mathematical Biology, 15(3):339–349, November 1982. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275692.

Shapiro:1982:MCO

[394] Louis W. Shapiro and Doron Zeilberger. A Markov chain occurring in enzyme kinetics. Journal of Mathematical Biology, 15(3):351– 357, November 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275693.

Ishii:1982:GSS

[395] Hitoshi Ishii and Izumi Takagi. Global stability of stationary solutions to a nonlinear diffusion equation in phytoplankton dynamics. Journal of Mathematical Biology, 16(1):1–24, December 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275157.

Clark:1982:CMT

[396] C. E. Clark and Thomas G. Hallam. The community matrix in three species community models. Journal of Mathematical Biology, 16(1): 25–31, December 1982. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275158.

Harrison:1982:ASI

[397] G. W. Harrison, Hugh J. Barclay, and P. van den Driessche. Analysis of a sterile insect release model with predation. Journal of Mathematical Biology, 16(1):33–48, December 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275159.

Hastings:1982:DSS

[398] Alan Hastings. Dynamics of a single species in a spatially varying environment: The stabilizing role of high dispersal rates. Journal of Mathematical Biology, 16(1):49–55, December 1982. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275160.

Nagylaki:1982:AMQ

[399] Thomas Nagylaki. Assortative mating for a quantitative character. Journal of Mathematical Biology, 16(1):57–74, December 1982. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275161.

anderHeiden:1982:DPD

[400] Uwe an der Heiden and Michael C. Mackey. The dynamics of production and destruction: Analytic insight into complex behavior. Journal of Mathematical Biology, 16(1):75–101, December 1982. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275162.

Kishimoto:1982:DLV

[401] Kazuo Kishimoto. The diffusive Lotka–Volterra system with three species can have a stable non-constant equilibrium solution. Journal of Mathematical Biology, 16(1):103–112, December 1982. CODEN JMBLAJ.

ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275163.

Chipot:1983:MTS

[402] Michel Chipot and Leah Edelstein. A mathematical theory of size distributions in tissue culture. Journal of Mathematical Biology, 16(2): 115–130, January 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276052.

Klein:1983:UIT

[403] P. Klein, T. Hraba, and J. Dolezal. The use of immunological tolerance to investigate B lymphocyte replacement kinetics in chickens. Journal of Mathematical Biology, 16(2):131–140, January 1983. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276053.

Lauffenburger:1983:LBI

[404] Douglas A. Lauffenburger and Clinton R. Kennedy. Localized bacterial infection in a distributed model for tissue inflammation. Journal of Mathematical Biology, 16(2):141–163, January 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276054.

Bremermann:1983:POL

[405] Hans J. Bremermann. Parasites at the origin of life. Journal of Mathematical Biology, 16(2):165–180, January 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276055.

Busenberg:1983:EMS

[406] Stavros N. Busenberg and Curtis C. Travis. Epidemic models with spatial spread due to population migration. Journal of Mathematical Biology, 16 (2):181–198, January 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276056.

Lui:1983:EST

[407] Roger Lui. Existence and stability of travelling wave solutions of a nonlinear integral operator. Journal of Mathematical Biology, 16(3): 199–220, February 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276502.

Schumacher:1983:NER

[408] Konrad Schumacher. No-escape regions and oscillations in second order predator–prey recurrences. Journal of Mathematical Biology, 16(3): 221–231, February 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276503.

Eilbeck:1983:CAN

[409] J. C. Eilbeck. A collocation approach to the numerical calculation of simple gradients in reaction–diffusion systems. Journal of Mathematical Biology, 16(3):233–249, February 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276504.

Manoranjan:1983:NSB

[410] V. S. Manoranjan and A. R. Mitchell. A numerical study of the Belousov– Zhabotinskii reaction using Galerkin finite element methods. Journal of Mathematical Biology, 16(3):251–260, February 1983. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276505.

Plant:1983:ACM

[411] Richard E. Plant. Analysis of a continuum model for root growth. Journal of Mathematical Biology, 16(3):261–268, February 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276506.

Burger:1983:NAS

[412] Reinhard Bürger. Nonlinear analysis of some models for the evolution of dominance. Journal of Mathematical Biology, 16(3):269–280, February 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276507.

Lory:1983:RCS

[413] Peter Lory, Albert Gilg, and Michael Horster. Renal countercurrent system: Role of collecting duct convergence and pelvic urea predicted from a mathematical model. Journal of Mathematical Biology, 16(3): 281–304, February 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276508.

Anonymous:1983:Ea

[414] Anonymous. Erratum. Journal of Mathematical Biology, 16(3):305, February 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00276509.pdf.

Griffiths:1983:AFG

[415] R. C. Griffiths. Allele frequencies with genic selection. Journal of Mathematical Biology, 17(1):1–10, May 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276111.

Dunbar:1983:TWS

[416] Steven R. Dunbar. Travelling wave solutions of diffusive Lotka–Volterra equations. Journal of Mathematical Biology, 17(1):11–32, May 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276112.

Bardi:1983:EGS

[417] Martino Bardi. An equation of growth of a single species with realistic dependence on crowding and seasonal factors. Journal of Mathematical Biology, 17(1):33–43, May 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276113.

Radcliffe:1983:WSD

[418] J. Radcliffe and L. Rass. Wave solutions for the deterministic nonreducible n-type epidemic. Journal of Mathematical Biology, 17(1): 45–66, May 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276114.

Argemi:1983:SPD

[419] José Argémi and Bruno Rossetto. Solutions périodiques discontinues pour l’approximation singulière d’un modèle neurophysiologique dans R^{4} — une métaphore dans R^{3} avec chaos. (French) [Discontinuous periodic solutions for the singular approximation of a neurophysiological model in R^{4} — a metaphor in R^{3} with chaos]. Journal of Mathematical Biology, 17(1):67–92, May 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276115.

Colli-Franzone:1983:ODL

[420] P. Colli-Franzone, L. Guerri, and C. Viganotti. Oblique dipole layer potentials applied to electrocardiology. Journal of Mathematical Biology, 17(1):93–124, May 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276116.

Murray:1983:MMM

[421] J. D. Murray, G. F. Oster, and A. K. Harris. A mechanical model for mesenchymal morphogenesis. Journal of Mathematical Biology, 17 (1):125–129, May 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276117.

Anonymous:1983:Eb

[422] Anonymous. Erratum. Journal of Mathematical Biology, 17(1): 131, May 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00276118.pdf.

Butler:1983:CCP

[423] G. J. Butler, S. B. Hsu, and P. Waltman. Coexistence of competing predators in a chemostat. Journal of Mathematical Biology, 17 (2):133–151, June 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00305755.

Merrill:1983:MRN

[424] Stephen J. Merrill. A model of the role of natural killer cells in immune surveillance- II. Journal of Mathematical Biology, 17(2):153–162, June 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00305756.

Smith:1983:SBR

[425] H. L. Smith. Subharmonic bifurcation in an S-I-R epidemic model. Journal of Mathematical Biology, 17(2):163–177, June 1983. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00305757.

Smith:1983:MSS

[426] H. L. Smith. Multiple stable subharmonics for a periodic epidemic model. Journal of Mathematical Biology, 17(2):179–190, June 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00305758.

Yellin:1983:NSU

[427] Joel Yellin. Natural selection under population regulation. Journal of Mathematical Biology, 17(2):191–216, June 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00305759.

Buckley:1983:GBB

[428] M. J. Buckley and E. Seneta. The genetic balance between varying population size and selective neutrality. Journal of Mathematical Biology, 17(2):217–222, June 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00305760.

Hunding:1983:BNR

[429] A. Hunding. Bifurcations of nonlinear reaction–diffusion systems in prolate spheroids. Journal of Mathematical Biology, 17(2):223–239, June 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00305761.

Losert:1983:DGG

[430] V. Losert and E. Akin. Dynamics of games and genes: Discrete versus continuous time. Journal of Mathematical Biology, 17(2):241–251, June 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00305762.

Banks:1983:PET

[431] H. T. Banks and P. Kareiva. Parameter estimation techniques for transport equations with application to population dispersal and tissue bulk flow models. Journal of Mathematical Biology, 17(3):253–273, July 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276516.

Smth:1983:ASP

[432] J. L. Smth and D. J. Wollkind. Age structure in predator–prey systems: Intraspecific carnivore interaction, passive diffusion, and the paradox of enrichment. Journal of Mathematical Biology, 17(3):275–288, July 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276517.

Longini:1983:MEE

[433] Ira M. Longini, Jr. Models of epidemics and endemicity in genetically variable host populations. Journal of Mathematical Biology, 17 (3):289–304, July 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276518.

Busenberg:1983:AMV

[434] S. Busenberg, K. L. Cooke, and M. A. Pozio. Analysis of a model of a vertically transmitted disease. Journal of Mathematical Biology, 17 (3):305–329, July 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276519.

Kilmer:1983:SDE

[435] W. Kilmer, W. Kroll, and R. Pelosi. On the stability of delay equation models of simple human stretch reflexes. Journal of Mathematical Biology, 17(3):331–349, July 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276520.

Othmer:1983:CMC

[436] H. G. Othmer. A continuum model for coupled cells. Journal of Mathematical Biology, 17(3):351–369, July 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276521.

Gripenberg:1983:SDG

[437] G. Gripenberg. A stationary distribution for the growth of a population subject to random catastrophes. Journal of Mathematical Biology, 17 (3):371–379, July 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276522.

Anonymous:1983:Aa

[438] Anonymous. Announcement. Journal of Mathematical Biology, 17 (3):381–382, July 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00276523.

Gregorius:1983:SPP

[439] H.-R. Gregorius, M. D. Ross, and E. Gillet. Selection in plant populations of effectively infinite size: IV. The maintenance of males among hermaphrodites for a biallelic model. Journal of Mathematical Biology, 18(1):1–12, August 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275906.

Bazykin:1983:MEA

[440] A. D. Bazykin, A. I. Khibnik, and E. A. Aponina. A model of evolutionary appearance of dissipative structure in ecosystems. Journal of Mathematical Biology, 18(1):13–23, August 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275907.

Hallam:1983:ETP

[441] T. G. Hallam, C. E. Clark, and G. S. Jordan. Effects of toxicants on populations: A qualitative approach II. First order kinetics. Journal of Mathematical Biology, 18(1):25–37, August 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275908.

Bell:1983:TCD

[442] J. Bell and C. Cosner. Threshold conditions for a diffusive model of a myelinated axon. Journal of Mathematical Biology, 18(1):39–52, August 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275909.

Misra:1983:EIS

[443] J. C. Misra and K. Roy Choudhury (Sen). Effect of initial stresses on the wave propagation in arteries. Journal of Mathematical Biology, 18 (1):53–67, August 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275910.

Honerkamp:1983:HSC

[444] J. Honerkamp. The heart as a system of coupled nonlinear oscillators. Journal of Mathematical Biology, 18(1):69–88, August 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275911.

Hutson:1983:CSS

[445] V. Hutson, W. Moran, and G. T. Vickers. On a criterion for survival of species in models governed by difference equations. Journal of Mathematical Biology, 18(1):89–90, August 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00275912.

Anonymous:1983:Ec

[446] Anonymous. Erratum. Journal of Mathematical Biology, 18(1):91– 92, August 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00275913.pdf.

Solimano:1983:EGA

[447] F. Solimano and E. Beretta. Existence of a globally asymptotically stable equilibrium in Volterra models with continuous time delay. Journal of Mathematical Biology, 18(2):93–102, November 1983. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00280659.

Helland:1983:DMD

[448] Inge S. Helland. Diffusion models for the dispersal of insects near an attractive center. Journal of Mathematical Biology, 18(2):103– 122, November 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00280660.

Eshel:1983:CIM

[449] I. Eshel and E. Akin. Cevolutionary instability of mixed Nash solutions. Journal of Mathematical Biology, 18(2):123–133, November 1983. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00280661.

Diekmann:1983:GFS

[450] O. Diekmann, H. A. Lauwerier, T. Aldenberg, and J. A. J. Metz. Growth, fission and the stable size distribution. Journal of Mathematical Biology, 18(2):135–148, November 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00280662.

Hansen:1983:RLM

[451] Poul. E. Hansen. Raising Leslie matrices to powers: a method and applications to demography. Journal of Mathematical Biology, 18(2): 149–161, November 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00280663.

Heyde:1983:AAA

[452] C. C. Heyde. An alternative approach to asymptotic results on genetic composition when the population size is varying. Journal of Mathematical Biology, 18(2):163–168, November 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00280664.

Murray:1983:MDS

[453] J. D. Murray and R. P. Sperb. Minimum domains for spatial patterns in a class of reaction diffusion equations. Journal of Mathematical Biology, 18 (2):169–184, November 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00280665.

Anonymous:1983:Ab

[454] Anonymous. Announcement. Journal of Mathematical Biology, 18(2): 185, November 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00280666.

Anonymous:1983:Ed

[455] Anonymous. Erratum. Journal of Mathematical Biology, 18(2):187, November 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00280667.pdf.

Jean:1983:ARP

[456] R. V. Jean. Allometric relations in plant growth. Journal of Mathematical Biology, 18(3):189–200, December 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276086.

Schoen:1983:RSH

[457] R. Schoen. Relationships in a simple harmonic mean two-sex fertility model. Journal of Mathematical Biology, 18(3):201–211, December 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276087.

Kishimoto:1983:SST

[458] K. Kishimoto, M. Mimura, and K. Yoshida. Stable spatio-temporal oscillations of diffusive Lotka–Volterra system with three or more species. Journal of Mathematical Biology, 18(3):213–221, December 1983. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276088.

Klinger:1983:RBH

[459] H. G. Klinger. Relations between heat transfer in perfused biological tissue and the local symmetry components of the vascular system. Journal of Mathematical Biology, 18(3):223–231, December 1983. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276089.

Schwartz:1983:ISB

[460] Ira B. Schwartz and H. L. Smith. Infinite subharmonic bifurcation in an SEIR epidemic model. Journal of Mathematical Biology, 18(3): 233–253, December 1983. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276090.

Hale:1983:CFN

[461] J. K. Hale and A. S. Somolinos. Competition for fluctuating nutrient. Journal of Mathematical Biology, 18(3):255–280, December 1983. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276091.

Bertsch:1984:IPD

[462] M. Bertsch, M. E. Gurtin, D. Hilhorst, and L. A. Peletier. On interacting populations that disperse to avoid crowding: The effect of a sedentary colony. Journal of Mathematical Biology, 19(1):1–12, January 1984. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275928.

Nagylaki:1984:DRM

[463] T. Nagylaki. Double recombinants in mitosis. Journal of Mathematical Biology, 19(1):13–42, January 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275929.

Lasota:1984:GAP

[464] A. Lasota and M. C. Mackey. Globally asymptotic properties of proliferating cell populations. Journal of Mathematical Biology, 19(1): 43–62, January 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275930.

Beck:1984:CMA

[465] K. Beck. Coevolution: Mathematical analysis of host-parasite interactions. Journal of Mathematical Biology, 19(1):63–77, January 1984. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275931.

Beck:1984:EEG

[466] K. Beck, J. P. Keener, and P. Ricciardi. The effect of epidemics on genetic evolution. Journal of Mathematical Biology, 19(1):79–94, January 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275932.

Nussinov:1984:ESS

[467] Ruth Nussinov. Effects of secondary structures in RNA on interlocking probabilities. Journal of Mathematical Biology, 19(1):95–107, January 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275933.

Iwasa:1984:BDM

[468] Y. Iwasa and E. Teramoto. Branching-diffusion model for the formation of distributional patterns in populations. Journal of Mathematical Biology, 19(1):109–124, January 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275934.

Novick-Cohen:1984:GST

[469] A. Novick-Cohen and L. A. Segel. A gradually slowing travelling band of chemotactic bacteria. Journal of Mathematical Biology, 19(1): 125–132, January 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275935.

Belgrade:1984:DBD

[470] J. F. Belgrade and G. Namkoong. Dynamical behavior of differential equation models of frequency and density dependent populations. Journal of Mathematical Biology, 19(1):133–146, January 1984. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275936.

Fisher:1984:SRD

[471] M. E. Fisher and B. S. Goh. Stability results for delayed-recruitment models in population dynamics. Journal of Mathematical Biology, 19 (1):147–156, January 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275937.

Gopalsamy:1984:GAS

[472] K. Gopalsamy. Global asymptotic stability in Volterra’s population systems. Journal of Mathematical Biology, 19(2):157–168, April 1984. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00277744.

Ellner:1984:ABS

[473] S. Ellner. Asymptotic behavior of some stochastic difference equation population models. Journal of Mathematical Biology, 19(2): 169–200, April 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277745.

Eshel:1984:SPS

[474] I. Eshel. On the survival probability of a slightly advantageous mutant gene in a multitype population: A multidimensional branching process model. Journal of Mathematical Biology, 19(2):201–209, April 1984. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00277746.

Mackey:1984:DRI

[475] M. C. Mackey and U. an der Heiden. The dynamics of recurrent inhibition. Journal of Mathematical Biology, 19(2):211–225, April 1984. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00277747.

Diekmann:1984:SCS

[476] O. Diekmann, H. J. A. M. Heijmans, and H. R. Thieme. On the stability of the cell size distribution. Journal of Mathematical Biology, 19 (2):227–248, April 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277748.

Hunding:1984:BNR

[477] Axel Hunding. Bifurcations of nonlinear reaction–diffusion systems in oblate spheroids. Journal of Mathematical Biology, 19(3):249–263, July 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00277098.

Murray:1984:CTM

[478] J. D. Murray and G. F. Oster. Cell traction models for generating pattern and form in morphogenesis. Journal of Mathematical Biology, 19 (3):265–279, July 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277099.

Giszter:1984:VAM

[479] S. F. Giszter, S. G. Koreisha, and R. F. Franklin. A vector autoregressive moving average time series approach for describing asymmetries of antennal control of two millipede species. Journal of Mathematical Biology, 19(3):281–302, July 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277100.

Radcliffe:1984:UWS

[480] J. Radcliffe and L. Rass. The uniqueness of wave solutions for the deterministic non-reducible n-type epidemic. Journal of Mathematical Biology, 19(3):303–308, July 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00277101.

Radcliffe:1984:SSF

[481] J. Radcliffe and L. Rass. The spatial spread and final size of the deterministic non-reducible n-type epidemic. Journal of Mathematical Biology, 19(3):309–327, July 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277102.

Palm:1984:ESS

[482] Günther Palm. Evolutionary stable strategies and game dynamics for n-person games. Journal of Mathematical Biology, 19(3):329–334, July 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00277103.

Tuljapurkar:1984:DSE

[483] S. Tuljapurkar. Demography in stochastic environments. I. Exact distributions of age structure. Journal of Mathematical Biology, 19(3): 335–350, July 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277104.

Anonymous:1984:Aa

[484] Anonymous. Announcements. Journal of Mathematical Biology, 19(3): 351, July 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00277105.

Anonymous:1984:GRC

[485] Anonymous. Gordon research conference on theoretical biology and biomathematics. Journal of Mathematical Biology, 19(3):353–354, July 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/BF00277106.

Doedel:1984:CAB

[486] Eusebius Doedel. The computer-aided bifurcation analysis of predator– prey models. Journal of Mathematical Biology, 20(1):1–14, August 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275858.

Karlin:1984:OSR

[487] Samuel Karlin and Sabin Lessard. On the optimal sex-ratio: A stability analysis based on a characterization for one-locus multiallele viability models. Journal of Mathematical Biology, 20(1):15–38, August 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275859.

Mahaffy:1984:MGC

[488] J. M. Mahaffy and C. V. Pao. Models of genetic control by repression with time delays and spatial effects. Journal of Mathematical Biology, 20(1):39–57, August 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275860.

Creegan:1984:SRA

[489] Paul Creegan and Roger Lui. Some remarks about the wave speed and travelling wave solutions of a nonlinear integral operator. Journal of Mathematical Biology, 20(1):59–68, August 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275861.

Aagaard-Hansen:1984:SDG

[490] Helle Aagaard-Hansen and G. F. Yeo. A stochastic discrete generation birth, continuous death population growth model and its approximate solution. Journal of Mathematical Biology, 20(1):69–90, August 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275862.

Hoppe:1984:PLU

[491] Fred M. Hoppe. Pólya-like urns and the Ewens’ sampling formula. Journal of Mathematical Biology, 20(1):91–94, August 1984. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275863.

Weiss:1984:NRG

[492] M. Weiss. A note on the rôle of generalized inverse Gaussian distributions of circulatory transit times in pharmacokinetics. Journal of Mathematical Biology, 20(1):95–102, August 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275864.

DalPasso:1984:AER

[493] R. Dal Passo and P. de Mottoni. Aggregative effects for a reaction– advection equation. Journal of Mathematical Biology, 20(1):103– 112, August 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275865.

Akin:1984:OCR

[494] Ethan Akin and H. Michael Lacker. Ovulation control: The right number or nothing. Journal of Mathematical Biology, 20(2):113–132, September 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00285341.

Keith:1984:PES

[495] W. L. Keith and R. H. Rand. 1:1 and 2:1 phase entrainment in a system of two coupled limit cycle oscillators. Journal of Mathematical Biology, 20(2):133–152, September 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00285342.

Granero-Porati:1984:TOM

[496] M. I. Granero-Porati and A. Porati. Temporal organization in a morphogenetic field. Journal of Mathematical Biology, 20(2):153–157, September 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00285343.

Gregorius:1984:CGF

[497] Hans-Rolf Gregorius. Convergence of genotypic frequencies for differential selfing and positive assortative mating at a biallelic locus. Journal of Mathematical Biology, 20(2):159–169, September 1984. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00285344.

Strittmatter:1984:FCR

[498] W. Strittmatter and J. Honerkamp. Fibrillation of a cardiac region and the tachycardia mode of a two-oscillator system. Journal of Mathematical Biology, 20(2):171–184, September 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00285345.

Totaro:1984:AGD

[499] Silvia Totaro. Algal growth dynamics under light interaction: A nonlinear evolution problem. Journal of Mathematical Biology, 20(2):185– 201, September 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00285346.

Anderson:1984:SSL

[500] David H. Anderson. Spectral sensitivity in linear biological models. Journal of Mathematical Biology, 20(2):203–221, September 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00285347.

Keller:1984:GVD

[501] Joseph B. Keller. Genetic variability due to geographical inhomogeneity. Journal of Mathematical Biology, 20(2):223–230, September 1984. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00285348.

Akin:1984:EDZ

[502] Ethan Akin and Viktor Losert. Evolutionary dynamics of zero-sum games. Journal of Mathematical Biology, 20(3):231–258, October 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275987.

Cushing:1984:ESE

[503] J. M. Cushing. Existence and stability of equilibria in age-structured population dynamics. Journal of Mathematical Biology, 20(3):259– 276, October 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275988.

Comincioli:1984:FSC

[504] V. Comincioli, A. Torelli, C. Poggesi, and C. Reggiani. A four-state cross bridge model for muscle contraction. Mathematical study and validation. Journal of Mathematical Biology, 20(3):277–304, October 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275989.

Merrill:1984:SMT

[505] Stephen J. Merrill. Stochastic models of tumor growth and the probability of elimination by cytotoxic cells. Journal of Mathematical Biology, 20 (3):305–320, October 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275990.

Liberman:1984:CFP

[506] Uri Liberman. On chiasma formation point processes having the count location property. Journal of Mathematical Biology, 21(1):1–10, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275218.

Pickard:1984:NEC

[507] William F. Pickard. Nonlinear equivalent circuits for membranes. Journal of Mathematical Biology, 21(1):11–23, December 1984. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275219.

Saleem:1984:PPR

[508] M. Saleem. predator–prey relationships: Indiscriminate predation. Journal of Mathematical Biology, 21(1):25–34, December 1984. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275220.

Hastings:1984:DRD

[509] Alan Hastings. Delays in recruitment at different trophic levels: Effects on stability. Journal of Mathematical Biology, 21(1):35–44, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275221.

Hadeler:1984:PDK

[510] K. P. Hadeler and K. Dietz. Population dynamics of killing parasites which reproduce in the host. Journal of Mathematical Biology, 21(1): 45–65, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275222.

Griffiths:1984:ALD

[511] R. C. Griffiths. Asymptotic line-of-descent distributions. Journal of Mathematical Biology, 21(1):67–75, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275223.

Wagner:1984:EDG

[512] G. P. Wagner. On the eigenvalue distribution of genetic and phenotypic dispersion matrices: Evidence for a nonrandom organization of quantitative character variation. Journal of Mathematical Biology, 21 (1):77–95, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275224.

Shigesada:1984:EIC

[513] Nanako Shigesada, Kohkichi Kawasaki, and Ei Teramoto. The effects of interference competition on stability, structure and invasion of a multi-species system. Journal of Mathematical Biology, 21(2): 97–113, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277664.

Heijmans:1984:HHM

[514] H. J. A. M. Heijmans. Holling’s “hungry mantid” model for the invertebrate functional response considered as a Markov process. III. Stable satiation distribution. Journal of Mathematical Biology, 21(2):115– 143, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277665.

Gopalsamy:1984:PPA

[515] K. Gopalsamy. Persistence in periodic and almost periodic Lotka– Volterra systems. Journal of Mathematical Biology, 21(2):145–148, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277666.

Fishman:1984:MSP

[516] Svetlana Fishman and Ruth Marcus. A model for spread of plant disease with periodic removals. Journal of Mathematical Biology, 21(2):

149–158, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277667.

DiLena:1984:IPB

[517] Giovanni Di Lena and Gabriella Serio. The identification of the periodic behaviour in an epidemic model. Journal of Mathematical Biology, 21(2): 159–174, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277668.

Keener:1984:GBP

[518] James P. Keener and Leon Glass. Global bifurcations of a periodically forced nonlinear oscillator. Journal of Mathematical Biology, 21(2): 175–190, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277669.

Smith:1984:RCS

[519] D. A. Smith. The role of convection in the standard fluid-mechanical model for morphogenesis. Journal of Mathematical Biology, 21(2):191– 203, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277670.

Gimelfarb:1984:SSD

[520] A. Gimelfarb. The significance of specific distributions and functions in models of quantitative inheritance. Journal of Mathematical Biology, 21 (2):205–211, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277671.

Anonymous:1984:Ab

[521] Anonymous. Announcement. Journal of Mathematical Biology, 21(2): 213, December 1984. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00277672.

Nagylaki:1985:BIG

[522] Thomas Nagylaki. Biased intrachromosomal gene conversion in a chromosome lineage. Journal of Mathematical Biology, 21(3):215–235, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276223.

Arrigoni:1985:LWF

[523] M. Arrigoni and A. Steiner. Logistisches Wachstum in fluktuierender Umwelt. (German) [Logistic growth in a fluctuating environment]. Journal of Mathematical Biology, 21(3):237–241, May 1985. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276224.

Ebel:1985:CFD

[524] Witta Ebel. Carrier facilitated diffusion. Journal of Mathematical Biology, 21(3):243–271, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276225.

Woods:1985:EME

[525] Steven D. Woods, Richard H. Rand, H. David Block, and Donald C. Lewis. Eye movements and the enhancement of edges. Journal of Mathematical Biology, 21(3):273–283, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276226.

Hutson:1985:PCG

[526] V. Hutson and R. Law. Permanent coexistence in general models of three interacting species. Journal of Mathematical Biology, 21(3):285–298, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276227.

Campbell:1985:DRP

[527] R. B. Campbell. Dimension reduction projection and our perception of evolution. Journal of Mathematical Biology, 21(3):299–306, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276228.

Weyer:1985:MMC

[528] J. Weyer. A mathematical model for chemical mass recruitment of ants. Journal of Mathematical Biology, 21(3):307–315, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276229.

Botsford:1985:OFP

[529] Louis W. Botsford and Thomas C. Wainwright. Optimal fishery policy: An equilibrium solution with irreversible investment. Journal of Mathematical Biology, 21(3):317–327, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276230.

Chover:1985:EGC

[530] Joshua Chover and James H. King. The early growth of cancer. Journal of Mathematical Biology, 21(3):329–346, May 1985. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276231.

Schwartz:1985:MSR

[531] Ira B. Schwartz. Multiple stable recurrent outbreaks and predictability in seasonally forced nonlinear epidemic models. Journal of Mathematical Biology, 21(3):347–361, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276232.

Artzrouni:1985:GSP

[532] Marc Artzrouni. Generalized stable population theory. Journal of Mathematical Biology, 21(3):363–381, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276233.

Oster:1985:CSM

[533] George F. Oster and Alan S. Perelson. Cell spreading and motility: A model lamellipod. Journal of Mathematical Biology, 21(3):383–388, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276234.

Anonymous:1985:E

[534] Anonymous. Erratum. Journal of Mathematical Biology, 21(3): 389, May 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00276235.pdf.

Ermentrout:1985:SPM

[535] G. Bard Ermentrout. Synchronization in a pool of mutually coupled oscillators with random frequencies. Journal of Mathematical Biology, 22(1):1–9, June 1985. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276542.

Pickard:1985:CCP

[536] William F. Pickard. The concept of channel permeability. Journal of Mathematical Biology, 22(1):11–19, June 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276543.

Donnelly:1985:WFM

[537] Peter Donnelly and Neville Weber. The Wright–Fisher model with temporally varying selection and population size. Journal of Mathematical Biology, 22(1):21–29, June 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276544.

Liberman:1985:STL

[538] Uri Liberman and Marcus W. Feldman. A symmetric two locus model with viability and fertility selection. Journal of Mathematical Biology, 22(1):31–60, June 1985. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276545.

Tyson:1985:GAS

[539] John J. Tyson and Kenneth B. Hannsgen. Global asymptotic stability of the size distribution in probabilistic models of the cell cycle. Journal of Mathematical Biology, 22(1):61–68, June 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276546.

Selgrade:1985:SPS

[540] James F. Selgrade and Gene Namkoong. Stable periodic solutions for two species, density dependent coevolution. Journal of Mathematical Biology, 22(1):69–80, June 1985. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276547.

Yanagida:1985:SFT

[541] Eiji Yanagida. Stability of fast travelling pulse solutions of the FitzHugh– Nagumo equations. Journal of Mathematical Biology, 22(1):81–104, June 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276548.

Thomas:1985:ESS

[542] Bernhard Thomas. On evolutionarily stable sets. Journal of Mathematical Biology, 22(1):105–115, June 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276549.

Dietz:1985:PMM

[543] Klaus Dietz and Dieter Schenzle. Proportionate mixing models for agedependent infection transmission. Journal of Mathematical Biology, 22 (1):117–120, June 1985. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276550.

Anonymous:1985:A

[544] Anonymous. Announcement. Journal of Mathematical Biology, 22(1): 121, June 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00276551.

Keener:1985:OCF

[545] James P. Keener. Oscillatory coexistence in a food chain model with competing predators. Journal of Mathematical Biology, 22(2): 123–135, August 1985. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275711.

Mahaffy:1985:SPS

[546] J. M. Mahaffy. Stability of periodic solutions for a model of genetic repression with delays. Journal of Mathematical Biology, 22(2): 137–144, August 1985. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275712.

Busenberg:1985:SMA

[547] S. Busenberg and M. Iannelli. Separable models in age-dependent population dynamics. Journal of Mathematical Biology, 22(2):145–173, August 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275713.

Arzberger:1985:PAT

[548] Peter Arzberger. A probabilistic and algebraic treatment of regular inbreeding systems. Journal of Mathematical Biology, 22(2):175– 197, August 1985. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275714.

Fisher:1985:EEC

[549] M. E. Fisher and W. J. Grantham. Estimating the effect of continual disturbances on discrete-time population models. Journal of Mathematical Biology, 22(2):199–207, August 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275715.

Metz:1985:HHMa

[550] J. A. J. Metz and F. H. D. van Batenburg. Holling’s “hungry mantid” model for the invertebrate functional response considered as a Markov process. Part I: The full model and some of its limits. Journal of Mathematical Biology, 22(2):209–238, August 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275716.

Metz:1985:HHMb

[551] J. A. J. Metz and F. H. D. van Batenburg. Holling’s “hungry mantid” model for the invertebrate functional response considered as a Markov process. Part II: Negligible handling time. Journal of Mathematical Biology, 22(2):239–257, August 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275717.

Banks:1985:MID

[552] H. T. Banks, P. M. Kareiva, and P. K. Lamm. Modeling insect dispersal and estimating parameters when mark-release techniques may cause initial disturbances. Journal of Mathematical Biology, 22(3):259–277, ???? 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276485.

OBrien:1985:HPV

[553] P. O’Brien. Homozygosity in a population of variable size and mutation rate. Journal of Mathematical Biology, 22(3):279–291, ???? 1985. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276486.

Hannsgen:1985:SSS

[554] Kenneth B. Hannsgen and John J. Tyson. Stability of the steady-state size distribution in a model of cell growth and division. Journal of Mathematical Biology, 22(3):293–301, ???? 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276487.

Roberts:1985:SCE

[555] M. G. Roberts. Stability in cyclic epidemic models. Journal of Mathematical Biology, 22(3):303–311, ???? 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276488.

Busenberg:1985:ISD

[556] Stavros Busenberg and Joseph Mahaffy. Interaction of spatial diffusion and delays in models of genetic control by repression. Journal of Mathematical Biology, 22(3):313–333, ???? 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276489.

Schwegler:1985:PCM

[557] H. Schwegler, K. Tarumi, and B. Gerstmann. Physico-chemical model of a protocell. Journal of Mathematical Biology, 22(3):335–348, ???? 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276490.

Berding:1985:SEM

[558] Christoph Berding and Georg Haubs. On the stability of equilibria in metabolic feedback systems. Journal of Mathematical Biology, 22 (3):349–352, ???? 1985. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276491.

McFarland:1985:DMM

[559] Blanche L. McFarland and H. Robert van der Vaart. A discrete mathematical model of unlabelled granulocyte kinetics. Journal of Mathematical Biology, 22(3):353–367, ???? 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276492.

Bertsch:1985:IPD

[560] M. Bertsch, M. E. Gurtin, D. Hilhorst, and L. A. Peletier. On interacting populations that disperse to avoid crowding: preservation of segregation. Journal of Mathematical Biology, 23(1):1–13, December 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276555.

Cushing:1985:ESP

[561] J. M. Cushing. Equilibria in structured populations. Journal of Mathematical Biology, 23(1):15–39, December 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276556.

Hofbauer:1985:SME

[562] Josef Hofbauer. The selection mutation equation. Journal of Mathematical Biology, 23(1):41–53, December 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276557.

Ermentrout:1985:BRC

[563] G. B. Ermentrout. The behavior of rings of coupled oscillators. Journal of Mathematical Biology, 23(1):55–74, December 1985. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276558.

Brokate:1985:PPC

[564] Martin Brokate. Pontryagin’s principle for control problems in agedependent population dynamics. Journal of Mathematical Biology, 23 (1):75–101, December 1985. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276559.

Donnelly:1985:DPI

[565] Peter Donnelly. Dual processes and an invariance result for exchangeable models in population genetics. Journal of Mathematical Biology, 23(1): 103–118, December 1985. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276560.

Grindrod:1985:MMN

[566] P. Grindrod and B. D. Sleeman. A model of a myelinated nerve axon: threshold behaviour and propagation. Journal of Mathematical Biology, 23(1):119–135, December 1985. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276561.

Baer:1986:ADN

[567] Steven M. Baer and Charles Tier. An analysis of a dendritic neuron model with an active membrane site. Journal of Mathematical Biology, 23 (2):137–161, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276954.

Kooijman:1986:WCT

[568] S. A. L. M. Kooijman. What the hen can tell about her eggs: egg development on the basis of energy budgets. Journal of Mathematical Biology, 23(2):163–185, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276955.

Liu:1986:INI

[569] Wei min Liu, Simon A. Levin, and Yoh Iwasa. Influence of nonlinear incidence rates upon the behavior of SIRS epidemiological models. Journal of Mathematical Biology, 23(2):187–204, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276956.

Alexander:1986:SOT

[570] J. C. Alexander. Spontaneous oscillations in two 2-component cells coupled by diffusion. Journal of Mathematical Biology, 23(2):205– 219, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276957.

Cressman:1986:EGT

[571] R. Cressman, A. T. Dash, and E. Akin. Evolutionary games and two species population dynamics. Journal of Mathematical Biology, 23(2): 221–230, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276958.

Tyson:1986:CGD

[572] John J. Tyson and Kenneth B. Hannsgen. Cell growth and division: a deterministic/probabilistic model of the cell cycle. Journal of Mathematical Biology, 23(2):231–246, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276959.

Plant:1986:MAS

[573] Richard E. Plant and L. T. Wilson. Models for age structured populations with distributed maturation rates. Journal of Mathematical Biology, 23 (2):247–262, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276960.

Rogers:1986:CCS

[574] Thomas D. Rogers, Zhuo-Cheng Yang, and Lee-Wah Yip. Complete chaos in a simple epidemiological model. Journal of Mathematical Biology, 23(2):263–268, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276961.

Webb:1986:MPC

[575] G. F. Webb. A model of proliferating cell populations with inherited cycle length. Journal of Mathematical Biology, 23(2):269–282, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276962.

Anonymous:1986:E

[576] Anonymous. Erratum. Journal of Mathematical Biology, 23(2):283, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00276963.pdf.

Anonymous:1986:Aa

[577] Anonymous. Announcements. Journal of Mathematical Biology, 23 (2):284, February 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00276964.

Asmussen:1986:DIAa

[578] Marjorie A. Asmussen. The dynamics of interlocus associations in the three locus hitchhiking model. Journal of Mathematical Biology, 23 (3):285–304, April 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275250.

Andrietti:1986:TDS

[579] F. Andrietti. Two-dimensional standing gradient osmotic flow: a generalization of the “isotonic convection approximation”. Journal of Mathematical Biology, 23(3):305–318, April 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275251.

Tang:1986:MIG

[580] Betty Tang. Mathematical investigations of growth of microorganisms in the gradostat. Journal of Mathematical Biology, 23(3):319–339, April 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275252.

Radcliffe:1986:ASP

[581] J. Radcliffe and L. Rass. The asymptotic speed of propagation of the deterministic non-reducible n-type epidemic. Journal of Mathematical Biology, 23(3):341–359, April 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275253.

Perelson:1986:MBM

[582] Alan S. Perelson and Evangelos A. Coutsias. A moving boundary model of acrosomal elongation. Journal of Mathematical Biology, 23 (3):361–379, April 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275254.

Scalia-Tomba:1986:AFS

[583] Gianpaolo Scalia-Tomba. The asymptotic final size distribution of reducible multitype Reed–Frost processes. Journal of Mathematical Biology, 23(3):381–392, April 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275255.

Anonymous:1986:Ab

[584] Anonymous. Announcement. Journal of Mathematical Biology, 23(3): 393, April 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00275256.

Niklas:1986:AET

[585] Karl J. Niklas and Vincent Kerchner. Aerodynamics of ephedra trifurca. Journal of Mathematical Biology, 24(1):1–24, May 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275718.

Edelstein-Keshet:1986:MTP

[586] Leah Edelstein-Keshet. Mathematical theory for plant-herbivore systems. Journal of Mathematical Biology, 24(1):25–58, May 1986. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275719.

Lefevre:1986:TBC

[587] Claude Lefèvre. Threshold behaviour for a chain-binomial S-I-S infectious disease. Journal of Mathematical Biology, 24(1):59–70, May 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275720.

Timischl:1986:PSR

[588] Werner Timischl. On the process of stabilization in the renewal model: approximations for the time to convergence. Journal of Mathematical Biology, 24(1):71–79, May 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275721.

Gyllenberg:1986:SSD

[589] Mats Gyllenberg. The size and scar distributions of the yeast saccharomyces cerevisiae. Journal of Mathematical Biology, 24(1):81–101, May 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275722.

Colli:1986:MMH

[590] Pierluigi Colli. A mathematical model of heterogeneous behavior of single muscle fibres. Journal of Mathematical Biology, 24(1):103–118, May 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275723.

Yashin:1986:DCR

[591] Anatoli I. Yashin, Kenneth G. Manton, and Eric Stallard. Dependent competing risks: a stochastic process model. Journal of Mathematical Biology, 24(2):119–140, July 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275995.

Arcuri:1986:PSB

[592] P. Arcuri and J. D. Murray. Pattern sensitivity to boundary and initial conditions in reaction–diffusion models. Journal of Mathematical Biology, 24(2):141–165, July 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275996.

Butler:1986:PMC

[593] G. J. Butler and G. S. K. Wolkowicz. Predator-mediated competition in the chemostat. Journal of Mathematical Biology, 24(2):167–191, July 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275997.

Selgrade:1986:EEG

[594] James F. Selgrade and Gene Namkoong. Examples of the effect of genetic variation on competing species. Journal of Mathematical Biology, 24 (2):193–206, July 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275998.

Hudson:1986:DMT

[595] Richard R. Hudson and Norman L. Kaplan. On the divergence of members of a transposable element family. Journal of Mathematical Biology, 24(2):207–215, July 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275999.

Belair:1986:PPS

[596] Jacques Bélair. Periodic pulsatile stimulation of a nonlinear oscillator. Journal of Mathematical Biology, 24(2):217–232, July 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276000.

Tallis:1986:JAD

[597] G. M. Tallis. On the joint asymptotic distribution of additive genotype for polygenic characters. Journal of Mathematical Biology, 24(3): 233–235, August 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275635.

Morrish:1986:POT

[598] Kathleen Morrish. Performance of one- and two-dimensional models for a slow flow system in a long, permeable tubule. Journal of Mathematical Biology, 24(3):237–258, August 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275636.

Ryan:1986:OHL

[599] Dennis Ryan and Floyd B. Hanson. Optimal harvesting of a logistic population in an environment with stochastic jumps. Journal of Mathematical Biology, 24(3):259–277, August 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275637.

Sod:1986:NSO

[600] Gary A. Sod. A numerical study of oxygen diffusion in a spherical cell with the Michaelis–Menten oxygen uptake kinetics. Journal of Mathematical Biology, 24(3):279–289, August 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275638.

TorrasiGenis:1986:EPC

[601] Carme Torras i Genís. Entrainment in pacemakers characterized by a Vshaped PRC. Journal of Mathematical Biology, 24(3):291–312, August 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275639.

VanderHoeven:1986:CIP

[602] Nelly Van der Hoeven. Coexistence of incompatible plasmids in a bacterial population living under a feast and famine regime. Journal of Mathematical Biology, 24(3):313–325, August 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275640.

Hallam:1986:PPM

[603] Thomas G. Hallam and Ma Zhien. Persistence in population models with demographic fluctuations. Journal of Mathematical Biology, 24 (3):327–339, August 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275641.

Burger:1986:MGV

[604] Reinhard Bürger. On the maintenance of genetic variation: global analysis of Kimura’s continuum-of-alleles model. Journal of Mathematical Biology, 24(3):341–351, August 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275642.

Rothman:1986:MWS

[605] E. D. Rothman and N. C. Weber. A model of weak selection in the infinite alleles framework. Journal of Mathematical Biology, 24(3): 353–360, August 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275643.

Asmussen:1986:DIAb

[606] Marjorie A. Asmussen. The dynamics of interlocus associations in the three-locus hitchhiking model. Journal of Mathematical Biology, 24 (4):361–380, ???? 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF01236887.

Cushing:1986:PLV

[607] J. M. Cushing. Periodic Lotka–Volterra competition equations. Journal of Mathematical Biology, 24(4):381–403, ???? 1986. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF01236888.

Artzrouni:1986:RCG

[608] Marc Artzrouni. The rate of convergence of a generalized stable population. Journal of Mathematical Biology, 24(4):405–422, ???? 1986. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF01236889.

Brannan:1986:RSI

[609] James R. Brannan. Relaxation spectra of interactive neural systems. Journal of Mathematical Biology, 24(4):423–436, ???? 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01236890.

Belic:1986:MMH

[610] M. R. Belić, V. Skarka, J. L. Deneubourg, and M. Lax. Mathematical model of honeycomb construction. Journal of Mathematical Biology, 24 (4):437–449, ???? 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF01236891.

Morton:1986:TCM

[611] R. Hugh Morton. A three component model of human bioenergetics. Journal of Mathematical Biology, 24(4):451–466, ???? 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01236892.

Udayabaskaran:1986:SIB

[612] S. Udayabaskaran and G. Sudalaiyandi. On a stochastic integral of a branching process. Journal of Mathematical Biology, 24(4):467–472, ???? 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF01236893.

Hadeler:1986:HTW

[613] K. P. Hadeler. The hypercycle, traveling waves, and Wright’s equation. Journal of Mathematical Biology, 24(5):473–477, December 1986. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275680.

Schaffer:1986:ENS

[614] W. M. Schaffer, S. Ellner, and M. Kot. Effects of noise on some dynamical models in ecology. Journal of Mathematical Biology, 24(5): 479–523, December 1986. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275681.

Perelson:1986:NPS

[615] A. S. Perelson, P. K. Maini, J. D. Murray, J. M. Hyman, and G. F. Oster. Nonlinear pattern selection in a mechanical model for morphogenesis. Journal of Mathematical Biology, 24(5):525–541, December 1986. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275682.

Noonburg:1986:CSM

[616] V. W. Noonburg. Competing species model with behavioral adaptation. Journal of Mathematical Biology, 24(5):543–555, December 1986. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275683.

Mangel:1986:SFD

[617] Marc Mangel. Solution of functional difference equations from behavioral theory. Journal of Mathematical Biology, 24(5):557–567, December 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275684.

Tuljapurkar:1986:DSE

[618] Shripad Tuljapurkar. Demography in stochastic environments. Journal of Mathematical Biology, 24(5):569–581, December 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275685.

Chi:1986:NSN

[619] Henjin Chi, Jonathan Bell, and Brian Hassard. Numerical solution of a nonlinear advance-delay-differential equation from nerve conduction theory. Journal of Mathematical Biology, 24(5):583–601, December 1986. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275686.

Roy:1987:GSO

[620] A. B. Roy and F. Solimano. Global stability and oscillations in classical Lotka–Volterra loops. Journal of Mathematical Biology, 24(6):

603–616, February 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275505.

Allen:1987:PEC

[621] Linda J. S. Allen. Persistence, extinction, and critical patch number for island populations. Journal of Mathematical Biology, 24(6): 617–625, February 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275506.

Cushing:1987:ESI

[622] J. M. Cushing. Equilibria in systems of interacting structured populations. Journal of Mathematical Biology, 24(6):627–649, February 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275507.

Hastings:1987:CCE

[623] Alan Hastings. Cycles in cannibalistic egg-larval interactions. Journal of Mathematical Biology, 24(6):651–666, February 1987. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275508.

Rumschitzki:1987:SPE

[624] David S. Rumschitzki. Spectral properties of eigen evolution matrices. Journal of Mathematical Biology, 24(6):667–680, February 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275509.

Janssen:1987:SRM

[625] Imke Janssen. A stochastic repair-misrepair model for the effects of radiation on cells. Journal of Mathematical Biology, 24(6):681–689, February 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275510.

Alt:1987:TBC

[626] Wolfgang Alt and Douglas A. Lauffenburger. Transient behavior of a chemotaxis system modelling certain types of tissue inflammation. Journal of Mathematical Biology, 24(6):691–722, February 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275511.

Singh:1987:GAP

[627] Mithilesh Kumar Singh. The gametic algebra for polyploidy with several loci. Journal of Mathematical Biology, 24(6):723–727, February 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275512.

Mangel:1987:OSS

[628] Marc Mangel. Opposition site selection and clutch size in insects. Journal of Mathematical Biology, 25(1):1–22, April 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275885.

Jager:1987:CG

[629] W. Jäger, Joseph W.-H. So, Betty Tang, and Paul Waltman. Competition in the gradostat. Journal of Mathematical Biology, 25(1): 23–42, April 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275886.

Notohara:1987:DMM

[630] Morihiro Notohara. Diallelic multilocus model of neutral genes. Journal of Mathematical Biology, 25(1):43–60, April 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275887.

vanderMee:1987:FPE

[631] C. V. M. van der Mee and P. F. Zweifel. A Fokker–Planck equation for growing cell populations. Journal of Mathematical Biology, 25(1):61–72, April 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275888.

Liu:1987:ELS

[632] Laifu Liu and J. E. Cohen. Equilibrium and local stability in a logistic matrix model for age-structured populations. Journal of Mathematical Biology, 25(1):73–88, April 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275889.

Saleem:1987:MMY

[633] M. Saleem, S. U. Siddiqui, and V. Gupta. A mathematical model with young predation. Journal of Mathematical Biology, 25(1):89–101, April 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275890.

Buhler:1987:SMM

[634] W. J. Bühler. A stochastic model of mutant growth. Journal of Mathematical Biology, 25(1):103–106, April 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275891.

Anonymous:1987:Aa

[635] Anonymous. Announcement. Journal of Mathematical Biology, 25(1): 107, April 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00275892.

Hunding:1987:BTS

[636] Axel Hunding. Bifurcations in Turing systems of the second kind may explain blastula cleavage plane orientation. Journal of Mathematical Biology, 25(2):109–121, June 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276385.

Hoppe:1987:STN

[637] Fred M. Hoppe. The sampling theory of neutral alleles and an urn model in population genetics. Journal of Mathematical Biology, 25 (2):123–159, June 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276386.

Tavare:1987:BPI

[638] Simon Tavaré. The birth process with immigration, and the genealogical structure of large populations. Journal of Mathematical Biology, 25(2):161–168, June 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276387.

Smith:1987:OMS

[639] Hal Smith. Oscillations and multiple steady states in a cyclic gene model with repression. Journal of Mathematical Biology, 25(2):169–190, June 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276388.

Hallam:1987:DEC

[640] Thomas G. Hallam and Ma Zhien. On density and extinction in continuous population models. Journal of Mathematical Biology, 25(2): 191–201, June 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276389.

Nedelman:1987:IAD

[641] Jerry Nedelman, Heather Downs, and Pamela Pharr. Inference for an age-dependent, multitype branching-process model of mast cells. Journal of Mathematical Biology, 25(2):203–226, June 1987. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276390.

Mangel:1987:E

[642] Marc Mangel. Erratum. Journal of Mathematical Biology, 25(2): 227, June 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/BF00276391.pdf.

Tranquillo:1987:SML

[643] R. T. Tranquillo and D. A. Lauffenburger. Stochastic model of leukocyte chemosensory movement. Journal of Mathematical Biology, 25 (3):229–262, July 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276435.

Hoppensteadt:1987:MAS

[644] F. C. Hoppensteadt and L. Murphy. A mathematical analysis of small mammal populations. Journal of Mathematical Biology, 25(3): 263–274, July 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276436.

Korner:1987:MMS

[645] B. Körner and J. Weyer. Mathematical models for some types of chemical information systems. Journal of Mathematical Biology, 25 (3):275–288, July 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276437.

Hom:1987:CTP

[646] Carole L. Hom. Control theory predictions of reproductive allocation in female dusky salamanders. Journal of Mathematical Biology, 25 (3):289–306, July 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276438.

Pakes:1987:LTP

[647] Anthony G. Pakes. Limit theorems for the population size of a birth and death process allowing catastrophes. Journal of Mathematical Biology, 25(3):307–325, July 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276439.

Strogatz:1987:HSC

[648] Steven H. Strogatz. Human sleep and circadian rhythms: a simple model based on two coupled oscillators. Journal of Mathematical Biology, 25 (3):327–347, July 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276440.

El-Owaidy:1987:PPS

[649] H. El-Owaidy, M. A. Allam, and A. A. Ammar. predator–prey systems with a perturbed periodic carrying capacity. Journal of Mathematical Biology, 25(3):349–358, July 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276441.

Liu:1987:DBE

[650] Wei min Liu, Herbert W. Hethcote, and Simon A. Levin. Dynamical behavior of epidemiological models with nonlinear incidence rates. Journal of Mathematical Biology, 25(4):359–380, September 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00277162.

Donnelly:1987:PGI

[651] Peter Donnelly and Simon Tavaré. The population genealogy of the infinitely-many neutral alleles model. Journal of Mathematical Biology, 25(4):381–391, September 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00277163.

Williams:1987:EGT

[652] H. P. Williams. Evolution, games theory and polyhedra. Journal of Mathematical Biology, 25(4):393–409, September 1987. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00277164.

Jansen:1987:PTR

[653] Wolfgang Jansen. A permanence theorem for replicator and Lotka– Volterra systems. Journal of Mathematical Biology, 25(4):411–422, September 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277165.

Griffiths:1987:CGT

[654] R. C. Griffiths. Counting genealogical trees. Journal of Mathematical Biology, 25(4):423–431, September 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00277166.

Sharan:1987:NSN

[655] Maithili Sharan, A. Aminataei, and M. P. Singh. A numerical study of the nonsteady transport of gases in the pulmonary capillaries. Journal of Mathematical Biology, 25(4):433–452, September 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00277167.

Aoki:1987:GCW

[656] Kenichi Aoki. Gene-culture waves of advance. Journal of Mathematical Biology, 25(5):453–464, November 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276192.

Cannings:1987:CII

[657] C. Cannings and R. Law. Conditions for the initial increase of new alleles with frequency dependent selection. Journal of Mathematical Biology, 25 (5):465–476, November 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276193.

Selgrade:1987:CEG

[658] James F. Selgrade and Martin Ziehe. Convergence to equilibrium in a genetic model with differential viability between the sexes. Journal of Mathematical Biology, 25(5):477–490, November 1987. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276194.

Landau:1987:HPB

[659] Mayer Landau, Paco Lorente, Jacques Henry, and Stephane Canu. Hysteresis phenomena between periodic and stationary solutions in a model of pacemaker and nonpacemaker coupled cardiac cells. Journal of Mathematical Biology, 25(5):491–509, November 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00276195.

Nebenzahl:1987:RAM

[660] I. Nebenzahl. Recall of associated memories. Journal of Mathematical Biology, 25(5):511–519, November 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276196.

Cooke:1987:ACD

[661] Kenneth L. Cooke and Helena E. Nusse. Analysis of the complicated dynamics of some harvesting models. Journal of Mathematical Biology, 25 (5):521–542, November 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276197.

Kolker:1987:PLG

[662] Yuri Kolker. A piecewise-linear growth model: comparison with competing forms in batch culture. Journal of Mathematical Biology, 25(5): 543–551, November 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276198.

Hofbauer:1987:CSG

[663] J. Hofbauer, V. Hutson, and W. Jansen. Coexistence for systems governed by difference equations of Lotka–Volterra type. Journal of Mathematical Biology, 25(5):553–570, November 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00276199.

Nedelman:1987:E

[664] Jerry Nedelman, Heather Downs, and Pamela Pharr. Erratum. Journal of Mathematical Biology, 25(5):571, November 1987. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/BF00276200.pdf.

Jiang:1987:DDS

[665] Hong Jiang and Thomas D. Rogers. The discrete dynamics of symmetric competition in the plane. Journal of Mathematical Biology, 25(6): 573–596, December 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275495.

Grindrod:1987:GMR

[666] Peter Grindrod and Jagannathan Gomatam. The geometry and motion of reaction–diffusion waves on closed two-dimensional manifolds. Journal of Mathematical Biology, 25(6):597–610, December 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275496.

Gomatam:1987:TDW

[667] Jagannathan Gomatam and Peter Grindrod. Three-dimensional waves in excitable reaction–diffusion systems. Journal of Mathematical Biology, 25 (6):611–622, December 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275497.

Koth:1987:GEB

[668] Maria Koth and Karl Sigmund. Gradients for the evolution of bimatrix games. Journal of Mathematical Biology, 25(6):623–635, December 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275498.

Holgate:1987:RBS

[669] Philip Holgate and Silvia D. A. Machado. Recombination between several polyploid loci. Journal of Mathematical Biology, 25(6):637– 642, December 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275499.

Iizuka:1987:WCS

[670] Masaru Iizuka. Weak convergence of a sequence of stochastic difference equations to a stochastic ordinary differential equation. Journal of Mathematical Biology, 25(6):643–652, December 1987. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00275500.

Rinzel:1987:DMN

[671] John Rinzel and Young Seek Lee. Dissection of a model for neuronal parabolic bursting. Journal of Mathematical Biology, 25(6):653– 675, December 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275501.

Gonshor:1987:MAD

[672] Harry Gonshor. Multi-algebra duplication. Journal of Mathematical Biology, 25(6):677–683, December 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00275502.

Mattfeldt:1987:VEB

[673] Torsten Mattfeldt. Volume estimation of biological objects by systematic sections. Journal of Mathematical Biology, 25(6):685–695, December 1987. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275503.

Anonymous:1987:Ab

[674] Anonymous. Announcement. Journal of Mathematical Biology, 25(6): 697, December 1987. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00275504.

Dietz:1988:EMS

[675] K. Dietz and K. P. Hadeler. Epidemiological models for sexually transmitted diseases. Journal of Mathematical Biology, 26(1):1–25, February 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00280169.

Hunding:1988:SAT

[676] Axel Hunding and Preben Graae Sørensen. Size adaptation of Turing prepatterns. Journal of Mathematical Biology, 26(1):27–39, February 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00280170.

Keener:1988:FCP

[677] James P. Keener. On the formation of circulating patterns of excitation in anisotropic excitable media. Journal of Mathematical Biology, 26 (1):41–56, February 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00280171.

Boucher:1988:CIC

[678] Wayne Boucher. Calculation of the inbreeding coefficient. Journal of Mathematical Biology, 26(1):57–64, February 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00280172.

Liou:1988:GSP

[679] Lii-Perng Liou and Kuo-Shung Cheng. Global stability of a predator– prey system. Journal of Mathematical Biology, 26(1):65–71, February 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00280173.

Vlad:1988:SMA

[680] Marcel Ovidiu Vlad. Separable models for age-structured population genetics. Journal of Mathematical Biology, 26(1):73–92, February 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00280174.

Farkas:1988:MBD

[681] A. Farkas, M. Farkas, and G. Szabó. Multiparameter bifurcation diagrams in predator–prey models with time lag. Journal of Mathematical Biology, 26(1):93–103, February 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00280175.

Misra:1988:IEM

[682] J. C. Misra, G. C. Bera, and S. Samanta. Induced electric and magnetic fields due to wave propagation in a tubular bone. Journal of Mathematical Biology, 26(1):105–120, February 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00280176.

Boucher:1988:RSI

[683] Wayne Boucher and Thomas Nagylaki. Regular systems of inbreeding. Journal of Mathematical Biology, 26(2):121–142, April 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00277729.

Beretta:1988:OSM

[684] E. Beretta, G. I. Bischi, and F. Solimano. Oscillations in a system with material cycling. Journal of Mathematical Biology, 26(2):143–167, April 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00277730.

Rotenberg:1988:MPS

[685] Manuel Rotenberg. Mappings of the plane that simulate prey-predator systems. Journal of Mathematical Biology, 26(2):169–191, April 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00277731.

Grey:1988:HUD

[686] D. R. Grey. Harvesting under density-dependent mortality and fecundity. Journal of Mathematical Biology, 26(2):193–197, April 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00277732.

Roerdink:1988:BLS

[687] J. B. T. M. Roerdink. The biennial life strategy in a random environment. Journal of Mathematical Biology, 26(2):199–215, April 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00277733.

Kirlinger:1988:PSE

[688] Gabriela Kirlinger. Permanence of some ecological systems with several predator and one prey species. Journal of Mathematical Biology, 26 (2):217–232, April 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277734.

Anonymous:1988:Aa

[689] Anonymous. Announcement. Journal of Mathematical Biology, 26(2): 233, April 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00277735.

Chipalkatti:1988:CIM

[690] Renu Chipalkatti and Michael A. Arbib. The cue interaction model of depth perception: a stability analysis. Journal of Mathematical Biology, 26(3):235–262, June 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277391.

Othmer:1988:MDB

[691] H. G. Othmer, S. R. Dunbar, and W. Alt. Models of dispersal in biological systems. Journal of Mathematical Biology, 26(3):263–298, June 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00277392.

Thieme:1988:WPP

[692] H. R. Thieme. Well-posedness of physiologically structured population models for Daphnia magna. Journal of Mathematical Biology, 26 (3):299–317, June 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00277393.

Langlais:1988:LTB

[693] Michel Langlais. Large time behavior in a nonlinear age-dependent population dynamics problem with spatial diffusion. Journal of Mathematical Biology, 26(3):319–346, June 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00277394.

Moody:1988:BPM

[694] Michael E. Moody. A branching process model for the evolution of transposable elements. Journal of Mathematical Biology, 26(3):347–357, June 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00277395.

Anonymous:1988:Ab

[695] Anonymous. Announcement. Journal of Mathematical Biology, 26(3): 359, June 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00277396.

Hardin:1988:APC

[696] D. P. Hardin, P. Takác, and G. F. Webb. Asymptotic properties of a continuous-space discrete-time population model in a random environment. Journal of Mathematical Biology, 26(4):361–374, August 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276367.

Mesterton-Gibbons:1988:OCD

[697] M. Mesterton-Gibbons. On the optimal compromise for a dispersing parasitoid. Journal of Mathematical Biology, 26(4):375–385, August 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276368.

Britton:1988:SDR

[698] N. F. Britton. A singular dispersion relation arising in a caricature of a model for morphogenesis. Journal of Mathematical Biology, 26(4):

387–403, August 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276369.

Pakes:1988:SBD

[699] Anthony G. Pakes. The supercritical birth, death and catastrophe process: Limit theorems on the set of non-extinction. Journal of Mathematical Biology, 26(4):405–420, August 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276370.

Li:1988:SCS

[700] Jia Li and T. G. Hallam. Survival in continuous structured populations models. Journal of Mathematical Biology, 26(4):421–433, August 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276371.

Doi:1988:RDP

[701] S. Doi and S. Sato. Regulation of differentiation in a population of cells interacting through a common pool. Journal of Mathematical Biology, 26 (4):435–454, August 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276372.

Garner:1988:EUS

[702] J. B. Garner and R. B. Kellogg. Existence and uniqueness of solutions in general multisolute renal flow problems. Journal of Mathematical Biology, 26(4):455–464, August 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276373.

Tyrcha:1988:ASG

[703] Joanna Tyrcha. Asymptotic stability in a generalized probabilistic/ deterministic model of the cell cycle. Journal of Mathematical Biology, 26(4):465–475, August 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276374.

Liberman:1988:ESE

[704] U. Liberman. External stability and ESS: criteria for initial increase of new mutant allele. Journal of Mathematical Biology, 26(4): 477–485, August 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276375.

DeYoung:1988:PAF

[705] G. DeYoung, P. B. Monk, and H. G. Othmer. Pacemakers in aggregation fields of Dictyostelium discoideum: does a single cell suffice? Journal of Mathematical Biology, 26(5):487–517, October 1988. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276057.

Arzberger:1988:ERI

[706] Peter Arzberger. An example of a regular inbreeding system with cubic ancestral growth that preserves some genetic variability. Journal of Mathematical Biology, 26(5):519–533, October 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276058.

Arzberger:1988:RGR

[707] Peter Arzberger. Results for generalized regular inbreeding systems. Journal of Mathematical Biology, 26(5):535–550, October 1988. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276059.

Bergh:1988:SDA

[708] M. O. Bergh and W. M. Getz. Stability of discrete age-structured and aggregated delay-difference population models. Journal of Mathematical Biology, 26(5):551–581, October 1988. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276060.

Lui:1988:CCE

[709] Roger Lui. Convergence to constant equilibrium for a density-dependent selection model with diffusion. Journal of Mathematical Biology, 26(5): 583–592, October 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276061.

Anonymous:1988:Ac

[710] Anonymous. Announcement. Journal of Mathematical Biology, 26 (5):593, October 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00276062.

Christiansen:1988:EML

[711] F. B. Christiansen. Epistasis in the multiple locus symmetric viability model. Journal of Mathematical Biology, 26(6):595–618, December 1988.

CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276143.

vandenBosch:1988:CLB

[712] F. van den Bosch, A. M. de Roos, and W. Gabriel. Cannibalism as a life boat mechanism. Journal of Mathematical Biology, 26(6):619– 633, December 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276144.

Hadeler:1988:MPF

[713] K. P. Hadeler, R. Waldstätter, and A. Wörz-Busekros. Models for pair formation in bisexual populations. Journal of Mathematical Biology, 26 (6):635–649, December 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276145.

Grindrod:1988:MIA

[714] Peter Grindrod. Models of individual aggregation or clustering in single and multi-species communities. Journal of Mathematical Biology, 26(6): 651–660, December 1988. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276146.

Beretta:1988:GSR

[715] E. Beretta, V. Capasso, and F. Rinaldi. Global stability results for a generalized Lotka–Volterra system with distributed delays. Journal of Mathematical Biology, 26(6):661–688, December 1988. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276147.

Frenzen:1988:EKT

[716] C. L. Frenzen and P. K. Maini. Enzyme kinetics for a two-step enzymic reaction with comparable initial enzyme-substrate ratios. Journal of Mathematical Biology, 26(6):689–703, December 1988. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276148.

Pal:1988:ESV

[717] Dulal Pal, N. Rudraiah, and Rathna Devanathan. The effects of slip velocity at a membrane surface on blood flow in the microcirculation. Journal of Mathematical Biology, 26(6):705–712, December 1988. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276149.

Namba:1989:CSH

[718] T. Namba. Competition for space in a heterogeneous environment. Journal of Mathematical Biology, 27(1):1–16, February 1989. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276077.

Ethier:1989:DAT

[719] S. N. Ethier and Thomas Nagylaki. Diffusion approximations of the two-locus Wright–Fisher model. Journal of Mathematical Biology, 27 (1):17–28, February 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276078.

James:1989:SMM

[720] A. T. James, J. T. Wiskich, and R. A. J. Conyers. Statistical modelling of mitochondrial power supply. Journal of Mathematical Biology, 27 (1):29–48, February 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276079.

Hethcote:1989:EMD

[721] H. W. Hethcote, M. A. Lewis, and P. van den Driessche. An epidemiological model with a delay and a nonlinear incidence rate. Journal of Mathematical Biology, 27(1):49–64, February 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00276080.

Furter:1989:LVN

[722] J. Furter and M. Grinfeld. Local vs. non-local interactions in population dynamics. Journal of Mathematical Biology, 27(1):65–80, February 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276081.

Abundo:1989:NSS

[723] M. Abundo and C. Rossi. Numerical simulation of a stochastic model for cancerous cells submitted to chemotherapy. Journal of Mathematical Biology, 27(1):81–90, February 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276082.

Brown:1989:ENS

[724] K. J. Brown, S. S. Lin, and A. Tertikas. Existence and nonexistence of steady-state solutions for a selection–migration model in population genetics. Journal of Mathematical Biology, 27(1):91–104, February 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276083.

Limic:1989:SCM

[725] N. Limić. On stochastic compartmental models. Journal of Mathematical Biology, 27(1):105–113, February 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00276084.

Anonymous:1989:Aa

[726] Anonymous. Announcement. Journal of Mathematical Biology, 27(1): 115–116, February 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00276085.

Chesson:1989:ISB

[727] P. L. Chesson and S. Ellner. Invasibility and stochastic boundedness in monotonic competition models. Journal of Mathematical Biology, 27 (2):117–138, April 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276099.

Smith:1989:CGR

[728] Hal Smith and Betty Tang. Competition in the gradostat: the role of the communication rate. Journal of Mathematical Biology, 27(2): 139–165, April 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276100.

Bartoszynski:1989:PPU

[729] R. Bartoszynski, W. J. Bühler, Wenyaw Chan, and D. K. Pearl. Population processes under the influence of disasters occurring independently of population size. Journal of Mathematical Biology, 27(2):167–178, April 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00276101.

Bremermann:1989:CEP

[730] H. J. Bremermann and H. R. Thieme. A competitive exclusion principle for pathogen virulence. Journal of Mathematical Biology, 27(2): 179–190, April 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276102.

Kretzschmar:1989:REB

[731] Mirjam Kretzschmar. A renewal equation with a birth-death process as a model for parasitic infections. Journal of Mathematical Biology, 27(2):191–221, April 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276103.

Aulbach:1989:PSH

[732] B. Aulbach and D. Flockerzi. The past in short hypercycles. Journal of Mathematical Biology, 27(2):223–231, April 1989. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00276104.

Castillo-Chavez:1989:EMA

[733] C. Castillo-Chavez, H. W. Hethcote, V. Andreasen, S. A. Levin, and W. M. Liu. Epidemiological models with age structure, proportionate mixing, and cross-immunity. Journal of Mathematical Biology, 27 (3):233–258, May 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275810.

Gardini:1989:HBT

[734] L. Gardini, R. Lupini, and M. G. Messia. Hopf bifurcation and transition to chaos in Lotka–Volterra equation. Journal of Mathematical Biology, 27(3):259–272, May 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275811.

Mesterton-Gibbons:1989:CFE

[735] M. Mesterton-Gibbons. On compromise in foraging and an experiment by Krebs et al. (1977). Journal of Mathematical Biology, 27(3):273–296, May 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00275812.

Swart:1989:VCA

[736] J. H. Swart. Viable controls in age-dependent population dynamics. Journal of Mathematical Biology, 27(3):297–308, May 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275813.

Roerdink:1989:BLS

[737] J. B. T. M. Roerdink. The biennial life strategy in a random environment. Journal of Mathematical Biology, 27(3):309–319, May 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00275814.

Pakes:1989:CNS

[738] A. G. Pakes. A complementary note on the supercritical birth, death and catastrophe process. Journal of Mathematical Biology, 27(3): 321–325, May 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275815.

Cromme:1989:CTA

[739] L. J. Cromme and I. E. Dammasch. Compensation type algorithms for neural nets: stability and convergence. Journal of Mathematical Biology, 27(3):327–340, May 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00275816.

Arino:1989:ABN

[740] O. Arino and M. Kimmel. Asymptotic behavior of a nonlinear functionalintegral equation of cell kinetics with unequal division. Journal of Mathematical Biology, 27(3):341–354, May 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00275817.

Darden:1989:NMC

[741] T. Darden, N. L. Kaplan, and R. R. Hudson. A numerical method for calculating moments of coalescent times in finite populations with selection. Journal of Mathematical Biology, 27(3):355–368, May 1989. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00275818.

Scheib:1989:AMR

[742] Artur Scheib. Analysis of a model for random competition. Journal of Mathematical Biology, 27(3):369–372, May 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00275819.

Castillo-Chavez:1989:RLI

[743] C. Castillo-Chavez, K. Cooke, W. Huang, and S. A. Levin. On the role of long incubation periods in the dynamics of acquired immunodeficiency syndrome (AIDS). Journal of Mathematical Biology, 27(4): 373–398, August 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00290636.

Schaalje:1989:RAR

[744] G. B. Schaalje and H. R. van der Vaart. Relationships among recent models for insect population dynamics with variable rates of development. Journal of Mathematical Biology, 27(4):399–428, August 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00290637.

Kirkpatrick:1989:QGM

[745] Mark Kirkpatrick and Nancy Heckman. A quantitative genetic model for growth, shape, reaction norms, and other infinite-dimensional characters. Journal of Mathematical Biology, 27(4):429–450, August 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00290638.

Ellner:1989:CSD

[746] S. Ellner. Convergence to stationary distributions in two-species stochastic competition models. Journal of Mathematical Biology, 27(4):451– 462, August 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00290639.

Swetina:1989:FSM

[747] J. Swetina. First and second moments and the mean Hamming distance in a stochastic replication–mutation model for biological macromolecules. Journal of Mathematical Biology, 27(4):463–483, August 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00290640.

Kurata:1989:ATC

[748] Koji Kurata, Kazuo Kisimoto, and Eiji Yanagida. The asymptotic transectional/circumferential homogeneity of the solutions of reaction– diffusion systems in/on cylinder-like domains. Journal of Mathematical Biology, 27(5):485–490, September 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00288429.

Vance:1989:NMP

[749] R. R. Vance and E. A. Coddington. A nonautonomous model of population growth. Journal of Mathematical Biology, 27(5):491–506, September 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00288430.

Maini:1989:SST

[750] P. K. Maini. Spatial and spatio-temporal patterns in a cell-haptotaxis model. Journal of Mathematical Biology, 27(5):507–522, September 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00288431.

Harrison:1989:CGA

[751] Michael J. Harrison. The cubic growth of AIDS cases: General dependence on early infection rates and distribution of times to appearance of clinical symptoms. Journal of Mathematical Biology, 27(5): 523–535, September 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00288432.

Hofbauer:1989:SEP

[752] J. Hofbauer and K. Sigmund. On the stabilizing effect of predators and competitors on ecological communities. Journal of Mathematical Biology, 27(5):537–548, September 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00288433.

Kretzschmar:1989:PSM

[753] Mirjam Kretzschmar. Persistent solutions in a model for parasitic infections. Journal of Mathematical Biology, 27(5):549–573, September 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00288434.

McCoy:1989:SEF

[754] E. L. McCoy. The strain energy function in axial plant growth. Journal of Mathematical Biology, 27(5):575–594, September 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00288435.

Gueron:1989:MHG

[755] Shay Gueron and Nadav Liron. A model of herd grazing as a travelling wave, chemotaxis and stability. Journal of Mathematical Biology, 27(5): 595–608, September 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00288436.

Hadeler:1989:PPP

[756] K. P. Hadeler and H. I. Freedman. predator–prey populations with parasitic infection. Journal of Mathematical Biology, 27(6):609–631, November 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276947.

Kunisch:1989:PES

[757] K. Kunisch and H. Scheich. Parameter estimation in a special reaction– diffusion system modelling man-environment diseases. Journal of Mathematical Biology, 27(6):633–665, November 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00276948.

Griffiths:1989:GTP

[758] R. C. Griffiths. Genealogical-tree probabilities in the infinitely-many-site model. Journal of Mathematical Biology, 27(6):667–680, November 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00276949.

Frigessi:1989:SMM

[759] A. Frigessi and F. den Hollander. A stochastic model for the membrane potential of a stimulated neuron. Journal of Mathematical Biology, 27(6): 681–692, November 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276950.

Cerrai:1989:MMB

[760] P. Cerrai. On a mathematical model for body tissue inflammation. Journal of Mathematical Biology, 27(6):693–706, November 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00276951.

Jose:1989:SMM

[761] Marco V. José. On the solution of mathematical models of herd immunity in human helminth infections. Journal of Mathematical Biology, 27(6): 707–715, November 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00276952.

Anonymous:1989:Ab

[762] Anonymous. Announcements. Journal of Mathematical Biology, 27 (6):717, November 1989. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00276953.

Hardin:1990:DPM

[763] D. P. Hardin, P. Takác, and G. F. Webb. Dispersion population models discrete in time and continuous in space. Journal of Mathematical Biology, 28(1):1–20, January 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00171515.

Walsh:1990:ISA

[764] J. B. Walsh. Inconsistencies in standard approximations for selection coefficients at loci affecting a polygenic character. Journal of Mathematical Biology, 28(1):21–31, January 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00171516.

Mackey:1990:DAS

[765] M. C. Mackey and J. G. Milton. A deterministic approach to survival statistics. Journal of Mathematical Biology, 28(1):33–48, January 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00171517.

Morton:1990:MHP

[766] R. H. Morton. Modelling human power and endurance. Journal of Mathematical Biology, 28(1):49–64, January 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00171518.

Pugliese:1990:PMD

[767] A. Pugliese. Population models for diseases with no recovery. Journal of Mathematical Biology, 28(1):65–82, January 1990. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00171519.

Hess:1990:CST

[768] P. Hess and H. Weinberger. Convergence to spatial-temporal clines in the Fisher equation with time-periodic fitnesses. Journal of Mathematical Biology, 28(1):83–98, January 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00171520.

Beretta:1990:SCE

[769] E. Beretta, G. I. Bischi, and F. Solimano. Stability in chemostat equations with delayed nutrient recycling. Journal of Mathematical Biology, 28(1):99–111, January 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00171521.

Nebenzahl:1990:SRA

[770] I. Nebenzahl and Y. Albeck. The storage and recall of auditory memory. Journal of Mathematical Biology, 28(1):113–119, January 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00171522.

Franzone:1990:WPA

[771] P. Colli Franzone, L. Guerri, and S. Rovida. Wavefront propagation in an activation model of the anisotropic cardiac tissue: asymptotic analysis and numerical simulations. Journal of Mathematical Biology, 28 (2):121–176, February 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163143.

Awiszus:1990:SPH

[772] F. Awiszus, J. Dehnhardt, and T. Funke. The singularly perturbed Hodgkin–Huxley equations as a tool for the analysis of repetitive nerve activity. Journal of Mathematical Biology, 28(2):177–195, February 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00163144.

Martin:1990:PAT

[773] M. H. Martin, M. H. M. Goldsmith, and T. H. Goldsmith. On polar auxin transport in plant cells. Journal of Mathematical Biology, 28(2): 197–223, February 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163145.

Chan:1990:GBA

[774] W. L. Chan and B. Z. Guo. Global behaviour of age-dependent logistic population models. Journal of Mathematical Biology, 28(2):225– 235, February 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163146.

Mangel:1990:EON

[775] Marc Mangel. Evolutionary optimization and neural network models of behavior. Journal of Mathematical Biology, 28(3):237–256, April 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00178775.

Busenberg:1990:ADT

[776] S. Busenberg and P. van den Driessche. Analysis of a disease transmission model in a population with varying size. Journal of Mathematical Biology, 28(3):257–270, April 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00178776.

Artzrouni:1990:TEH

[777] M. Artzrouni. On transient effects in the HIV/AIDS epidemic. Journal of Mathematical Biology, 28(3):271–291, April 1990. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00178777.

Boucher:1990:CRS

[778] W. Boucher and C. W. Cotterman. On the classification of regular systems of inbreeding. Journal of Mathematical Biology, 28(3): 293–305, April 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00178778.

Maini:1990:SMC

[779] P. K. Maini. Superposition of modes in a caricature of a model for morphogenesis. Journal of Mathematical Biology, 28(3):307–315, April 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00178779.

Garner:1990:MAM

[780] J. B. Garner. Mathematical analysis of multisolute renal flow in a single nephron model of the kidney. Journal of Mathematical Biology, 28 (3):317–327, April 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00178780.

Hastings:1990:MPV

[781] A. Hastings. Maintenance of polygenic variation through mutation– selection balance: bifurcation analysis of a biallelic model. Journal of Mathematical Biology, 28(3):329–340, April 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00178781.

Moreira:1990:LEU

[782] Helmar Nunes Moreira. On Liénard’s equation and the uniqueness of limit cycles in predator–prey systems. Journal of Mathematical Biology, 28(3):341–354, April 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00178782.

Pal:1990:PTP

[783] D. S. Pal and S. Pal. Prediction of temperature profiles in the human skin and subcutaneous tissues. Journal of Mathematical Biology, 28(3):355–364, April 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00178783.

Diekmann:1990:DCB

[784] O. Diekmann, J. A. P. Heesterbeek, and J. A. J. Metz. On the definition and the computation of the basic reproduction ratio R_{0} in models for infectious diseases in heterogeneous populations. Journal of Mathematical Biology, 28(4):365–382, June 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00178324.

Krawcewicz:1990:PHD

[785] W. Krawcewicz and T. D. Rogers. Perfect harmony: the discrete dynamics of cooperation. Journal of Mathematical Biology, 28(4):383–410, June 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00178325.

Inaba:1990:TSR

[786] H. Inaba. Threshold and stability results for an age-structured epidemic model. Journal of Mathematical Biology, 28(4):411–434, June 1990. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00178326.

Liberman:1990:MSC

[787] Uri Liberman. Mendelian segregation: a choice between “order” and “chaos”. Journal of Mathematical Biology, 28(4):435–449, June 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00178327.

Brauer:1990:MSU

[788] F. Brauer. Models for the spread of universally fatal diseases. Journal of Mathematical Biology, 28(4):451–462, June 1990. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00178328.

Kuang:1990:GSG

[789] Y. Kuang. Global stability of Gause-type predator–prey systems. Journal of Mathematical Biology, 28(4):463–474, June 1990. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00178329.

Hastings:1990:SOA

[790] Alan Hastings. Second-order approximations for selection coefficients at polygenic loci. Journal of Mathematical Biology, 28(4):475–483, June 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00178330.

Anonymous:1990:A

[791] Anonymous. Announcement. Journal of Mathematical Biology, 28(4): 485, June 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00178331.

Stockbridge:1990:ECC

[792] N. Stockbridge and N. Yamoah. Excitability changes in the crustacean motor axon following activity. Journal of Mathematical Biology, 28 (5):487–499, ???? 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00164160.

Banks:1990:SMM

[793] H. T. Banks and B. G. Fitzpatrick. Statistical methods for model comparison in parameter estimation problems for distributed systems. Journal of Mathematical Biology, 28(5):501–527, ???? 1990. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00164161.

vandenBosch:1990:VSP

[794] F. van den Bosch, J. A. J. Metz, and O. Diekmann. The velocity of spatial population expansion. Journal of Mathematical Biology, 28 (5):529–565, ???? 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00164162.

Totafurno:1990:TAF

[795] J. Totafurno, M. Bjerknes, M. J. Allen, and H. Cheng. Theoretical analysis of the flow of cells over villi of the small intestine. Journal of Mathematical Biology, 28(5):567–584, ???? 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00164163.

Briere:1990:ECI

[796] C. Brière and B. C. Goodwin. Effects of calcium input/output on the stability of a system for calcium-regulated viscoelastic strain fields. Journal of Mathematical Biology, 28(5):585–593, ???? 1990. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00164164.

Sridhara:1990:STS

[797] Rajeshwari Sridhara and Ray Watson. Stochastic three species systems. Journal of Mathematical Biology, 28(5):595–607, ???? 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00164165.

DeRoos:1990:SDP

[798] A. M. De Roos, J. A. J. Metz, E. Evers, and A. Leipoldt. A size dependent predator–prey interaction: who pursues whom? Journal of Mathematical Biology, 28(6):609–643, September 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160229.

Vance:1990:GDM

[799] R. R. Vance. A general dynamical model of one consumer-one resource interactions. Journal of Mathematical Biology, 28(6):645–669, September 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00160230.

Gyllenberg:1990:NSP

[800] M. Gyllenberg and G. F. Webb. A nonlinear structured population model of tumor growth with quiescence. Journal of Mathematical Biology, 28(6): 671–694, September 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160231.

Adler:1990:CTT

[801] F. R. Adler. Coexistence of two types on a single resource in discrete time. Journal of Mathematical Biology, 28(6):695–713, September 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00160232.

Cervantes-Perez:1990:SPD

[802] F. Cervantes-Pérez and M. A. Arbib. Stability and parameter dependency analysis of a facilitation tectal column (FTC) model. Journal of Mathematical Biology, 29(1):1–32, October 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00173907.

Edelstein-Keshet:1990:MCM

[803] Leah Edelstein-Keshet and G. Bard Ermentrout. Models for contactmediated pattern formation: cells that form parallel arrays. Journal of Mathematical Biology, 29(1):33–58, October 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00173908.

Notohara:1990:CGP

[804] M. Notohara. The coalescent and the genealogical process in geographically structured population. Journal of Mathematical Biology, 29(1): 59–75, October 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00173909.

Murphy:1990:OHA

[805] L. F. Murphy and S. J. Smith. Optimal harvesting of an age-structured population. Journal of Mathematical Biology, 29(1):77–90, October 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00173910.

Sridhara:1990:ISP

[806] R. Sridhara. Inference on system parameters in a model for interacting species. Journal of Mathematical Biology, 29(1):91–97, October 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00173911.

Christiansen:1990:GMM

[807] F. B. Christiansen. The generalized multiplicative model for viability selection at multiple loci. Journal of Mathematical Biology, 29(2):

99–129, December 1990. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168174.

Ethier:1990:TLS

[808] S. N. Ethier and R. C. Griffiths. On the two-locus sampling distribution. Journal of Mathematical Biology, 29(2):131–159, December 1990. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00168175.

Karlin:1990:GCP

[809] S. Karlin and U. Liberman. Global convergence properties in multilocus viability selection models: the additive model and the Hardy–Weinberg law. Journal of Mathematical Biology, 29(2):161–176, December 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00168176.

Voller:1990:SVL

[810] R. L. Voller. Solving Volterra–Lotka systems with diffusion by monotone iteration. Journal of Mathematical Biology, 29(2):177–187, December 1990. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00168177.

Solow:1990:NSP

[811] A. R. Solow. A note on the statistical properties of animal locations. Journal of Mathematical Biology, 29(2):189–193, December 1990. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00168178.

Ermentrout:1991:MPI

[812] G. B. Ermentrout and N. Kopell. Multiple pulse interactions and averaging in systems of coupled neural oscillators. Journal of Mathematical Biology, 29(3):195–217, January 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160535.

Mimura:1991:EDS

[813] M. Mimura, S.-I. Ei, and Q. Fang. Effect of domain-shape on coexistence problems in a competition–diffusion system. Journal of Mathematical Biology, 29(3):219–237, January 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160536.

Lui:1991:PSE

[814] R. Lui. Positive solutions of an elliptic system arising from a model in evolutionary ecology. Journal of Mathematical Biology, 29(3):239– 250, January 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160537.

Mesterton-Gibbons:1991:EPD

[815] M. Mesterton-Gibbons. An escape from ‘the prisoner’s dilemma‘. Journal of Mathematical Biology, 29(3):251–269, January 1991. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00160538.

Hethcote:1991:SEM

[816] H. W. Hethcote and P. van den Driessche. Some epidemiological models with nonlinear incidence. Journal of Mathematical Biology, 29(3): 271–287, January 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160539.

Arino:1991:ABN

[817] O. Arino and M. Kimmel. Asymptotic behavior of nonlinear semigroup describing a model of selective cell growth regulation. Journal of Mathematical Biology, 29(4):289–314, February 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00167154.

Cantrell:1991:ESH

[818] R. S. Cantrell and C. Cosner. The effects of spatial heterogeneity in population dynamics. Journal of Mathematical Biology, 29(4):315– 338, February 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00167155.

Wollkind:1991:DIO

[819] D. J. Wollkind, J. B. Collings, and M. C. B. Barba. Diffusive instabilities in a one-dimensional temperature-dependent model system for a mite predator–prey interaction on fruit trees: dispersal motility and aggregative preytaxis effects. Journal of Mathematical Biology, 29(4): 339–362, February 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00167156.

Kubo:1991:PSP

[820] M. Kubo and M. Langlais. Periodic solutions for a population dynamics problem with age-dependence and spatial structure. Journal of Mathematical Biology, 29(4):363–378, February 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00167157.

Krivan:1991:CPG

[821] V. Krivan. Construction of population growth equations in the presence of viability constraints. Journal of Mathematical Biology, 29(4): 379–387, February 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00167158.

Sherratt:1991:MAB

[822] J. A. Sherratt and J. D. Murray. Mathematical analysis of a basic model for epidermal wound healing. Journal of Mathematical Biology, 29(5):389–404, April 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160468.

Alexander:1991:DBS

[823] J. C. Alexander and Da-Yong Cai. On the dynamics of bursting systems. Journal of Mathematical Biology, 29(5):405–423, April 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00160469.

Schuster:1991:MICa

[824] S. Schuster and R. Heinrich. Minimization of intermediate concentrations as a suggested optimality principle for biochemical networks. Journal of Mathematical Biology, 29(5):425–442, April 1991. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00160470.

Schuster:1991:MICb

[825] S. Schuster, R. Schuster, and R. Heinrich. Minimization of intermediate concentrations as a suggested optimality principle for biochemical networks. Journal of Mathematical Biology, 29(5):443–455, April 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00160471.

Cushing:1991:JVA

[826] J. M. Cushing and Jia Li. Juvenile versus adult competition. Journal of Mathematical Biology, 29(5):457–473, April 1991. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00160472.

Mazzullot:1991:NST

[827] S. Mazzullot, M. Paolini, and C. Verdi. Numerical simulation of thermal bone necrosis during cementation of femoral prostheses. Journal of Mathematical Biology, 29(5):475–494, April 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00160473.

Abundo:1991:SMP

[828] M. Abundo. A stochastic model for predator–prey systems: basic properties, stability and computer simulation. Journal of Mathematical Biology, 29(6):495–511, June 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00164048.

Takada:1991:DAE

[829] T. Takada and J. Kigami. The dynamical attainability of ESS in evolutionary games. Journal of Mathematical Biology, 29(6):513–529, June 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00164049.

Hayes:1991:NTW

[830] C. K. Hayes. A new traveling-wave solution of Fisher’s equation with density-dependent diffusivity. Journal of Mathematical Biology, 29(6): 531–537, June 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00164050.

Diekmann:1991:PEI

[831] O. Diekmann and M. Kretzschmar. Patterns in the effects of infectious diseases on population growth. Journal of Mathematical Biology, 29(6):539–570, June 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00164051.

Ermentrout:1991:AMS

[832] B. Ermentrout. An adaptive model for synchrony in the firefly Pteroptyx malaccae. Journal of Mathematical Biology, 29(6):571–585, June 1991.

CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00164052.

Szucs:1991:SMDa

[833] Joseph M. Szucs. Selection and mutation at a diallelic X-linked locus. Journal of Mathematical Biology, 29(7):587–627, July 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00163915.

Keener:1991:ECE

[834] James P. Keener. An eikonal-curvature equation for action potential propagation in myocardium. Journal of Mathematical Biology, 29 (7):629–651, July 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163916.

Liberman:1991:MAS

[835] U. Liberman. Multilocus autosomal sex determination. Journal of Mathematical Biology, 29(7):653–669, July 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00163917.

Iizukah:1991:COD

[836] Masaru Iizukah and Yukio Ogura. Convergence of one-dimensional diffusion processes to a jump process related to population genetics. Journal of Mathematical Biology, 29(7):671–687, July 1991. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00163918.

Hoppensteadt:1991:SH

[837] F. C. Hoppensteadt. The searchlight hypothesis. Journal of Mathematical Biology, 29(7):689–691, July 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00163919.

Szucs:1991:SDA

[838] Joseph M. Szucs. Selection at a diallelic autosomal locus in a dioecious population. Journal of Mathematical Biology, 29(8):693–713, August 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00160187.

Yang:1991:CPP

[839] Feng Yang and H. I. Freedman. Competing predators for a prey in a chemostat model with periodic nutrient input. Journal of Mathematical Biology, 29(8):715–732, August 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160188.

Invernizzi:1991:QAH

[840] Sergio Invernizzi and Giulia Treu. Quantitative analysis of the Hopf bifurcation in the Goodwin n-dimensional metabolic control system. Journal of Mathematical Biology, 29(8):733–742, August 1991. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00160189.

Basten:1991:BPM

[841] Christopher J. Basten and Michael E. Moody. A branching-process model for the evolution of transposable elements incorporating selection. Journal of Mathematical Biology, 29(8):743–761, August 1991. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00160190.

Griffiths:1991:WLO

[842] R. C. Griffiths. Which locus has the oldest allele? Journal of Mathematical Biology, 29(8):763–777, August 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00160191.

Lin:1991:UEE

[843] Xiaodong Lin. On the uniqueness of endemic equilibria of an HIV/AIDS transmission model for a heterogeneous population. Journal of Mathematical Biology, 29(8):779–790, August 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00160192.

Anonymous:1991:A

[844] Anonymous. Announcement. Journal of Mathematical Biology, 29(8): 791, August 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/BF00160193.

Szucs:1991:SMDb

[845] Joseph M. Szucs. Selection–mutation at a diallelic autosomal locus in a dioecious population. Journal of Mathematical Biology, 30(1): 1–14, October 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168003.

Freedman:1991:MET

[846] H. I. Freedman and J. B. Shukla. Models for the effect of toxicant in single-species and predator–prey systems. Journal of Mathematical Biology, 30(1):15–30, October 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00168004.

Smith:1991:EDS

[847] H. L. Smith. Equilibrium distribution of species among vessels of a gradostat. Journal of Mathematical Biology, 30(1):31–48, October 1991. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00168005.

Huaping:1991:TSS

[848] Liu Huaping and Ma Zhien. The threshold of survival for system of two species in a polluted environment. Journal of Mathematical Biology, 30(1):49–61, October 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168006.

Hader:1991:STF

[849] Donat-P. Häder and Kurt Vogel. Simultaneous tracking of flagellates in real time by image analysis. Journal of Mathematical Biology, 30 (1):63–72, October 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168007.

Bomze:1991:CEM

[850] Immanuel M. Bomze. Cross entropy minimization in uninvadable states of complex populations. Journal of Mathematical Biology, 30 (1):73–87, October 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168008.

Cressman:1991:SSE

[851] R. Cressman and A. T. Dash. Strong stability and evolutionarily stable strategies with two types of players. Journal of Mathematical Biology, 30(1):89–99, October 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168009.

Hall:1991:SSD

[852] A. J. Hall, G. C. Wake, and P. W. Gandar. Steady size distributions for cells in one-dimensional plant tissues. Journal of Mathematical Biology, 30(2):101–123, November 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160330.

Banks:1991:EGS

[853] H. T. Banks, L. W. Botsford, F. Kappel, and C. Wane. Estimation of growth and survival in size-structured cohort data: an application to larval striped bass (Morone saxatilis). Journal of Mathematical Biology, 30 (2):125–150, November 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160331.

Salathe:1991:NLP

[854] Eric P. Salathe and Yao-Huan Xu. Non-linear phenomena in oxygen transport to tissue. Journal of Mathematical Biology, 30(2):151– 160, November 1991. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160332.

Franke:1991:MEV

[855] John E. Franke and Abdul-Aziz Yakubu. Mutual exclusion versus coexistence for discrete competitive systems. Journal of Mathematical Biology, 30(2):161–168, November 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160333.

Nagai:1991:TWC

[856] Toshitaka Nagai and Tsutomu Ikeda. Traveling waves in a chemotactic model. Journal of Mathematical Biology, 30(2):169–184, November 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00160334.

Gregorius:1991:EAT

[857] Hans-Rolf Gregorius. Establishment of allelic two-locus polymorphisms. Journal of Mathematical Biology, 30(2):185–197, November 1991. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00160335.

Burger:1991:MCP

[858] Reinhard Burger. Moments, cumulants, and polygenic dynamics. Journal of Mathematical Biology, 30(2):199–213, November 1991. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00160336.

Mogami:1992:CBT

[859] Yoshihiro Mogami, Jörg Pernberg, and Hans Machemer. Ciliary beating in three dimensions: Steps of a quantitative description. Journal of Mathematical Biology, 30(3):215–249, January 1992. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00176150.

Gurney:1992:LLC

[860] William S. C. Gurney, Philip H. Crowley, and Roger M. Nisbet. Locking life-cycles onto seasons: Circle-map models of population dynamics and local adaptation. Journal of Mathematical Biology, 30(3): 251–279, January 1992. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00176151.

Jenkins:1992:SMP

[861] Michael J. Jenkins, James Sneyd, Scott Camazine, and J. D. Murray. On a simplified model for pattern formation in honey bee colonies. Journal of Mathematical Biology, 30(3):281–306, January 1992. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00176152.

Timm:1992:DDI

[862] Uwe Timm and Akira Okubo. Diffusion-driven instability in a predator– prey system with time-varying diffusivities. Journal of Mathematical Biology, 30(3):307–320, January 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00176153.

Oelschlager:1992:SPI

[863] Karl Oelschläger. The spread of a parasitic infection in a spatially distributed host population. Journal of Mathematical Biology, 30(4): 321–354, February 1992. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00173291.

da-fu:1992:BAN

[864] Ding da fu, John Totafurno, and L. E. H. Trainor. Bifurcation analysis of a nonlinear field model of regeneration. Journal of Mathematical Biology, 30(4):355–377, February 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00173292.

Hastings:1992:SOA

[865] Alan Hastings. Second-order approximations for selection coefficients at polygenic loci. Journal of Mathematical Biology, 30(4):379–388, February 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00173293.

Myerscough:1992:AOD

[866] M. R. Myerscough, B. F. Gray, W. L. Hogarth, and J. Norbury. An analysis of an ordinary differential equation model for a two-species predator–prey system with harvesting and stocking. Journal of Mathematical Biology, 30(4):389–411, February 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00173294.

Kot:1992:DTT

[867] Mark Kot. Discrete-time travelling waves: Ecological examples. Journal of Mathematical Biology, 30(4):413–436, February 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00173295.

Arino:1992:PRN

[868] Ovide Arino and Abdessamad Mortabit. A periodicity result for a nonlinear functional integral equation. Journal of Mathematical Biology, 30(5):437–456, April 1992. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160530.

Hutson:1992:TWD

[869] V. C. L. Hutson and G. T. Vickers. Travelling waves and dominance of ESS’s. Journal of Mathematical Biology, 30(5):457–471, April 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00160531.

Schwartz:1992:SAL

[870] Ira B. Schwartz. Small amplitude, long period outbreaks in seasonally driven epidemics. Journal of Mathematical Biology, 30(5):473–491, April 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00160532.

Hahnfeldt:1992:EDD

[871] P. Hahnfeldt, R. K. Sachse, and L. R. Hlatky. Evolution of DNA damage in irradiated cells. Journal of Mathematical Biology, 30(5):493–511, April 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00160533.

Mitnura:1992:FBT

[872] Masayasu Mitnura, Ryo Kobayashi, and Tohru Tsujikawa. Formation of band-type electric patterns in Characean cells. Journal of Mathematical Biology, 30(5):513–545, April 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160534.

Hoppe:1992:ARG

[873] Fred M. Hoppe. Asymptotic rates of growth of the extinction probability of a mutant gene. Journal of Mathematical Biology, 30(6):547–566, June 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00175605;
http://link.springer.com/article/10.1007/BF00948890.

Hoppe:1992:SPM

[874] Fred M. Hoppe. The survival probability of a mutant in a multidimensional population. Journal of Mathematical Biology, 30(6):567–575, June 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00175606;
http://link.springer.com/article/10.1007/BF00948891.

Athreya:1992:RDS

[875] K. B. Athreya. Rates of decay for the survival probability of a mutant gene. Journal of Mathematical Biology, 30(6):577–581, June 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00175607;
http://link.springer.com/article/10.1007/BF00948892.

Pollak:1992:SPS

[876] Edward Pollak. Survival probabilities for some multitype branching processes in genetics. Journal of Mathematical Biology, 30(6):583–596, June 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00175608;
http://link.springer.com/article/10.1007/BF00948893.

Stadler:1992:MAR

[877] Peter F. Stadler and Peter Schuster. Mutation in autocatalytic reaction networks. Journal of Mathematical Biology, 30(6):597–632, June 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00175609;
http://link.springer.com/article/10.1007/BF00948894.

Roth:1992:HAI

[878] Bradley J. Roth. How the anisotropy of the intracellular and extracellular conductivities influences stimulation of cardiac muscle. Journal of Mathematical Biology, 30(6):633–646, June 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00175610;   http://link.springer.com/
article/10.1007/BF00948895.

Lin:1992:TRE

[879] Xiaodong Lin and P. van den Driessche. A threshold result for an epidemiological model. Journal of Mathematical Biology, 30(6):647–654, June 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
BF00175611;   http://link.springer.com/article/10.1007/BF00948896.

Lopez-Gomez:1992:CSF

[880] J. López-Gómez and R. Pardo San Gil. Coexistence in a simple food chain with diffusion. Journal of Mathematical Biology, 30(7):655–668, August 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00173262.

Mahaffy:1992:OMR

[881] Joseph M. Mahaffy, David A. Jorgensen, and Robert L. Vanderheyden. Oscillations in a model of repression with external control. Journal of Mathematical Biology, 30(7):669–691, August 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00173263.

Mena-Lorcat:1992:DMI

[882] Jaime Mena-Lorcat and Herbert W. Hethcote. Dynamic models of infectious diseases as regulators of population sizes. Journal of Mathematical Biology, 30(7):693–716, August 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00173264.

Gao:1992:DTM

[883] Linda Q. Gao and Herbert W. Hethcote. Disease transmission models with density-dependent demographics. Journal of Mathematical Biology, 30(7):717–731, August 1992. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00173265.

Milner:1992:RCN

[884] Fabio A. Milner and Guglielno Rabbiolo. Rapidly converging numerical algorithms for models of population dynamics. Journal of Mathematical Biology, 30(7):733–753, August 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00173266.

Thieme:1992:CRP

[885] Horst R. Thieme. Convergence results and a Poincaré–Bendixson trichotomy for asymptotically autonomous differential equations. Journal of Mathematical Biology, 30(7):755–763, August 1992. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00173267.

Boucher:1992:GAT

[886] Wayne Boucher. Global analysis of a two-allele mating system. Journal of Mathematical Biology, 30(8):765–773, October 1992. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00176454.

Lasota:1992:SDR

[887] Andrzej Lasota, Michael C. Mackey, and Joanna Tyrcha. The statistical dynamics of recurrent biological events. Journal of Mathematical Biology, 30(8):775–800, October 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00176455.

Smith:1992:MSY

[888] Steven J. Smith. Maximum sustainable yield of a density dependent sexage-structured population model. Journal of Mathematical Biology, 30 (8):801–814, October 1992. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00176456.

Selgrade:1992:DBP

[889] James F. Selgrade and Gene Namkoong. Dynamical behavior for population genetics models of differential and difference equations with nonmonotone fitnesses. Journal of Mathematical Biology, 30(8):815– 826, October 1992. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00176457.

Collins:1992:SBB

[890] J. J. Collins and I. N. Stewart. Symmetry-breaking bifurcation: A possible mechanism for 2: 1 frequency-locking in animal locomotion. Journal of Mathematical Biology, 30(8):827–838, October 1992. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00176458.

Adkison:1992:PEM

[891] Milo D. Adkison. Parameter estimation for models of chaotic time series. Journal of Mathematical Biology, 30(8):839–852, October 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00176459.

Tang:1992:MMM

[892] Betty Tang and Gail S. K. Wolkowicz. Mathematical models of microbial growth and competition in the chemostat regulated by cell-bound extracellular enzymes. Journal of Mathematical Biology, 31(1):1–23, December 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00163841.

Lewis:1992:ADS

[893] M. A. Lewis and J. D. Murray. Analysis of dynamic and stationary pattern formation in the cell cortex. Journal of Mathematical Biology, 31(1):25–71, December 1992. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163842.

Artzrouni:1992:MTV

[894] Marc Artzrouni. A modeled time-varying density function for the incubation period of AIDS. Journal of Mathematical Biology, 31(1): 73–99, December 1992. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163843.

Holgate:1992:MHW

[895] P. Holgate. The multilocus Hardy–Weinberg law. Journal of Mathematical Biology, 31(1):101–106, December 1992. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00163844.

Gradl:1992:CTM

[896] H. Gradl and S. Walcher. On continuous time models in genetic and Bernstein algebras. Journal of Mathematical Biology, 31(1):107– 113, December 1992. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163845.

Notohara:1993:SML

[897] Morihiro Notohara. The strong-migration limit for the genealogical process in geographically structured populations. Journal of Mathematical Biology, 31(2):115–122, January 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00171220.

Notohara:1993:GPN

[898] Morihiro Notohara. The genealogical process of neutral genes with mutation in geographically structured populations. Journal of Mathematical Biology, 31(2):123–132, January 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00171221.

Kalachev:1993:RWS

[899] Leonid V. Kalachev. A relaxation wave solution of the FitzHugh–Nagumo equations. Journal of Mathematical Biology, 31(2):133–147, January 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00171222.

Boucher:1993:DAS

[900] Wayne Boucher. A deterministic analysis of self-incompatibility alleles. Journal of Mathematical Biology, 31(2):149–155, January 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00171223.

Barbour:1993:HPM

[901] A. D. Barbour and Maria Kafetzaki. A host-parasite model yielding heterogeneous parasite loads. Journal of Mathematical Biology, 31(2):

157–176, January 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00171224.

Zhivotovsky:1993:PLN

[902] Lev A. Zhivotovsky and Marcus W. Feldman. On the probability of loss of new mutations in the presence of linkage disequilibrium. Journal of Mathematical Biology, 31(2):177–188, January 1993. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00171225.

Blanchard:1993:SDA

[903] P. Blanchard, Ph. Combe, H. Nencka, and R. Rodriguez. Stochastic dynamical aspects of neuronal activity. Journal of Mathematical Biology, 31 (2):189–198, January 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00171226.

Braun:1993:DEM

[904] W. J. Braun and R. J. Kulperger. Differential equations for moments of an interacting particle process on a lattice. Journal of Mathematical Biology, 31(2):199–214, January 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00171227.

Ikeda:1993:IAR

[905] Tsutomu Ikeda and Masayasu Mimura. An interfacial approach to regional segregation of two competing species mediated by a predator. Journal of Mathematical Biology, 31(3):215–240, February 1993. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00166143.

Campbell:1993:HSM

[906] R. B. Campbell. Half-sib mating structures. Journal of Mathematical Biology, 31(3):241–252, February 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00166144.

Tuljapurkar:1993:ECD

[907] Shripad Tuljapurkar. Entropy and convergence in dynamics and demography. Journal of Mathematical Biology, 31(3):253–271, February 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00166145.

Chicone:1993:PIP

[908] C. C. Chicone, Z.-S. Cai, P. D. King, and E. L. Forker. A parameter identification problem arising from a model of canalicular bile formation. Journal of Mathematical Biology, 31(3):273–289, February 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00166146.

Sachs:1993:DDN

[909] R. K. Sachs, P. Chen, P. Hahnfeldt, D. Lai, and L. R. Hlatky. DNA damage in non-proliferating cells subjected to ionizing irradiation at high or low dose rates. Journal of Mathematical Biology, 31(3):291– 315, February 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00166147.

Szucs:1993:EDSa

[910] Joseph M. Szucs. Equilibria and dynamics of selection at a diallelic autosomal locus in the Nagylaki–Crow continuous model of a monoecious population. Journal of Mathematical Biology, 31(4):317–349, April 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00163920.

Holden:1993:UBO

[911] Lisa Holden and Thomas Erneux. Understanding bursting oscillations as periodic slow passages through bifurcation and limit points. Journal of Mathematical Biology, 31(4):351–365, April 1993. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00163921.

Silva:1993:EDT

[912] Jacques A. L. Silva and Thomas G. Hallam. Effects of delay, truncations and density dependence in reproduction schedules on stability of nonlinear Leslie matrix models. Journal of Mathematical Biology, 31 (4):367–395, April 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163922.

Gavrilets:1993:EEV

[913] Sergey Gavrilets. Equilibria in an epistatic viability model under arbitrary strength of selection. Journal of Mathematical Biology, 31(4): 397–410, April 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163923.

Vickers:1993:SPP

[914] G. T. Vickers, V. C. L. Hutson, and C. J. Budd. Spatial patterns in population conflicts. Journal of Mathematical Biology, 31(4):411–430, April 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00163924.

Chaplain:1993:MGS

[915] M. A. J. Chaplain and B. D. Sleeman. Modelling the growth of solid tumours and incorporating a method for their classification using nonlinear elasticity theory. Journal of Mathematical Biology, 31(5):431–473, May 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00173886.

Chauvet:1993:NLB

[916] Gilbert A. Chauvet. Non-locality in biological systems results from hierarchy. Journal of Mathematical Biology, 31(5):475–486, May 1993. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00173887.

Liu:1993:DDL

[917] Wei min Liu. Dose-dependent latent period and periodicity of infectious diseases. Journal of Mathematical Biology, 31(5):487–494, May 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00173888.

Derrick:1993:DTM

[918] W. R. Derrick and P. van den Driessche. A disease transmission model in a nonconstant population. Journal of Mathematical Biology, 31 (5):495–512, May 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00173889.

Freedman:1993:MCC

[919] H. I. Freedman and Xu Yuantong. Models of competition in the chemostat with instantaneous and delayed nutrient recycling. Journal of Mathematical Biology, 31(5):513–527, May 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00173890.

Heesterbeek:1993:SCR

[920] J. A. P. Heesterbeek and J. A. J. Metz. The saturating contact rate in marriage- and epidemic models. Journal of Mathematical Biology, 31(5):529–539, May 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00173891.

Lindstrom:1993:QAP

[921] Torsten Lindström. Qualitative analysis of a predator–prey system with limit cycles. Journal of Mathematical Biology, 31(6):541–561, July 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00161198.

Dickinson:1993:SMA

[922] Richard B. Dickinson and Robert T. Tranquillo. A stochastic model for adhesion-mediated cell random motility and haptotaxis. Journal of Mathematical Biology, 31(6):563–600, July 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00161199.

Szucs:1993:EDSb

[923] Joseph M. Szucs. Equilibria and dynamics of selection at a diallelic autosomal locus in the Nagylaki–Crow continuous model of a monoecious population with random mating. Journal of Mathematical Biology, 31 (6):601–609, July 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00161200.

Kappel:1993:MMF

[924] F. Kappel and R. O. Peer. A mathematical model for fundamental regulation processes in the cardiovascular system. Journal of Mathematical Biology, 31(6):611–631, July 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00161201.

Ruan:1993:PCZ

[925] Shigui Ruan. Persistence and coexistence in zooplankton-phytoplankton-nutrient models with instantaneous nutrient recycling. Journal of Mathematical Biology, 31(6):633–654, July 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00161202.

Britton:1993:ECS

[926] N. F. Britton and U. Timm. Effects of competition and shading in planktonic communities. Journal of Mathematical Biology, 31(7):655– 673, September 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160418.

Moll:1993:AWM

[927] Victor H. Moll and Aaron L. Fogelson. Activation waves in a model of platelet aggregation: existence of solutions and stability of travelling fronts. Journal of Mathematical Biology, 31(7):675–701, September 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00160419.

Sherratt:1993:AAE

[928] Jonathan A. Sherratt. Actin aggregation and embryonic epidermal wound healing. Journal of Mathematical Biology, 31(7):703– 716, September 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160420.

Lui:1993:DDS

[929] Roger Lui and James F. Selgrade. Density-dependent selection migration model with non-monotone fitness functions. Journal of Mathematical Biology, 31(7):717–734, September 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160421.

Adam:1993:EMV

[930] John A. Adam and Richard D. Noren. Equilibrium model of a vascularized spherical carcinoma with central necrosis- some properties of the solution. Journal of Mathematical Biology, 31(7):735–745, September 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00160422.

Zhivotovsky:1993:HSS

[931] Lev A. Zhivotovsky and Marcus W. Feldman. Heterogeneous selection in subdivided populations. Journal of Mathematical Biology, 31(7): 747–759, September 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160423.

Teramoto:1993:RDD

[932] Ei Teramoto. Random disturbance and diversity of competitive systems. Journal of Mathematical Biology, 31(7):761–769, September 1993. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00160424.

Chauvet:1993:LFT

[933] G. A. Chauvet. An n-level field theory of biological neural networks. Journal of Mathematical Biology, 31(8):771–795, October 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00168045.

Yip:1993:ESA

[934] Paul Yip and Ray Watson. Estimating selective advantage of two alleles in discrete time. Journal of Mathematical Biology, 31(8):797– 804, October 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168046.

Taib:1993:NMD

[935] Ziad Taib. A note on modelling the dynamics of budding yeast populations using branching processes. Journal of Mathematical Biology, 31 (8):805–815, October 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168047.

Sulsky:1993:NSS

[936] Deborah Sulsky. Numerical solution of structured population models. Journal of Mathematical Biology, 31(8):817–839, October 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00168048.

Nath:1993:CTC

[937] H. B. Nath and R. C. Griffiths. The coalescent in two colonies with symmetric migration. Journal of Mathematical Biology, 31(8): 841–851, October 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168049.

Behncke:1993:MMF

[938] Horst Behncke. A mathematical model for the force and energetics in competitive running. Journal of Mathematical Biology, 31(8): 853–878, October 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168050.

Ardito:1993:ESF

[939] A. Ardito, P. Ricciardi, and A. Schiaffino. The effects of seasonal fluctuations on an open access fishery problem. Journal of Mathematical Biology, 31(8):879–890, October 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00168051.

Hadeler:1993:PFM

[940] K. P. Hadeler. Pair formation models with maturation period. Journal of Mathematical Biology, 32(1):1–15, November 1993. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00160370.

Pal:1993:EBF

[941] D. S. Pal and Sushma Pal. Effects of blood flow, curved boundary and environmental conditions on temperature distribution in a two dimensional model of human skin and subcutaneous tissues. Journal of Mathematical Biology, 32(1):17–32, November 1993. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00160371.

Wagner:1993:WDB

[942] G. P. Wagner and P. Krall. What is the difference between models of error thresholds and Muller’s ratchet? Journal of Mathematical Biology, 32 (1):33–44, November 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160372.

Athreya:1993:RDS

[943] K. B. Athreya. Rates of decay for the survival probability of a mutant gene II the multitype case. Journal of Mathematical Biology, 32 (1):45–53, November 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160373.

Prior:1993:EGS

[944] T. G. Prior, W. G. S. Hines, and R. Cressman. Evolutionary games for spatially dispersed populations. Journal of Mathematical Biology, 32 (1):55–65, November 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160374.

Lu:1993:GAB

[945] Zhengyi Lu and Yasuhiro Takeuchi. Global asymptotic behavior in singlespecies discrete diffusion systems. Journal of Mathematical Biology, 32 (1):67–77, November 1993. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160375.

Solow:1993:IED

[946] Andrew R. Solow. Inferring extinction in a declining population. Journal of Mathematical Biology, 32(1):79–82, November 1993. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00160376.

Bowtell:1994:ABD

[947] Graham Bowtell and Thelma L. Williams. Anguilliform body dynamics: a continuum model for the interaction between muscle activation and body curvature. Journal of Mathematical Biology, 32(2):83–91, January 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00163025.

Gyllenberg:1994:PME

[948] Mats Gyllenberg, Göran Högnäs, and Timo Koski. Population models with environmental stochasticity. Journal of Mathematical Biology, 32 (2):93–108, January 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163026.

Kuznetsov:1994:BAP

[949] Yu. A. Kuznetsov and C. Piccardi. Bifurcation analysis of periodic SEIR and SIR epidemic models. Journal of Mathematical Biology, 32 (2):109–121, January 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163027.

Macken:1994:DBP

[950] Catherine A. Macken, Simon A. Levin, and Roland Waldstätter. The dynamics of bacteria-plasmid systems. Journal of Mathematical Biology, 32 (2):123–145, January 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163028.

Segel:1994:SBK

[951] Lee Segel and Albert Goldbeter. Scaling in biochemical kinetics: dissection of a relaxation oscillator. Journal of Mathematical Biology, 32 (2):147–160, January 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163029.

Model:1994:SFP

[952] Charles J. Model and Klaus Dietz. On some formulas in a partnership model from the perspective of a semi-Markov process. Journal of Mathematical Biology, 32(2):161–169, January 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00163030.

Glendinning:1994:ICM

[953] Paul Glendinning. Island chain models and gradient systems. Journal of Mathematical Biology, 32(2):171–178, January 1994. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00163031.

Crowe:1994:NET

[954] Kathleen M. Crowe. A nonlinear ergodic theorem for discrete systems. Journal of Mathematical Biology, 32(3):179–191, February 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00163877.

Burger:1994:MLM

[955] Reinhard Bürger and Josef Hofbauer. Mutation load and mutation– selection-balance in quantitative genetic traits. Journal of Mathematical Biology, 32(3):193–218, February 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00163878.

Kuznetsov:1994:CDM

[956] Yu. A. Kuznetsov, M. Ya. Antonovsky, V. N. Biktashev, and E. A. Aponina. A cross-diffusion model of forest boundary dynamics. Journal of Mathematical Biology, 32(3):219–232, February 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00163879.

Velasco-Hernandez:1994:MET

[957] Jorge X. Velasco-Hernandez and Ying-Hen Hsieh. Modelling the effect of treatment and behavioral change in HIV transmission dynamics. Journal of Mathematical Biology, 32(3):233–249, February 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00163880.

Sato:1994:PIH

[958] Kazunori Sato, Hirotsugu Matsuda, and Akira Sasaki. Pathogen invasion and host extinction in lattice structured populations. Journal of Mathematical Biology, 32(3):251–268, February 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00163881.

Gradl:1994:LTB

[959] H. Gradl. Long-time behavior of continuous time models in genetic algebras. Journal of Mathematical Biology, 32(3):269–274, February 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00163882.

Bell:1994:NMM

[960] Jonathan Bell and Mark H. Holmes. A note on modeling mechanochemical transduction with an application to a skin receptor. Journal of Mathematical Biology, 32(3):275–285, February 1994. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00163883.

Gavrilets:1994:MMV

[961] Sergey Gavrilets and Alan Hastings. Maintenance of multilocus variability under strong stabilizing selection. Journal of Mathematical Biology, 32(4):287–302, April 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160162.

Edelstein-Keshet:1994:SMT

[962] Leah Edelstein-Keshet. Simple models for trail-following behaviour; trunk trails versus individual foragers. Journal of Mathematical Biology, 32(4):303–328, April 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160163.

Nisbet:1994:PDC

[963] R. M. Nisbet and L. C. Onyiah. Population dynamic consequences of competition within and between age classes. Journal of Mathematical Biology, 32(4):329–344, April 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160164.

Dillon:1994:PFG

[964] R. Dillon, P. K. Maini, and H. G. Othmer. Pattern formation in generalized Turing systems. Journal of Mathematical Biology, 32(4): 345–393, April 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160165.

He:1994:PSL

[965] X. Z. He and K. Gopalsamy. Persistence, stability and level crossings in an integrodifferential system. Journal of Mathematical Biology, 32 (5):395–426, May 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160166.

Klebanoff:1994:CTS

[966] Aaron Klebanoff and Alan Hastings. Chaos in three species food chains. Journal of Mathematical Biology, 32(5):427–451, May 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00160167.

Capasso:1994:ABR

[967] V. Capasso and A. Di Liddo. Asymptotic behaviour of reaction–diffusion systems in population and epidemic models. Journal of Mathematical Biology, 32(5):453–463, May 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160168.

Burchard:1994:SDM

[968] Ahnut Burchard. Substrate degradation by a mutualistic association of two species in the chemostat. Journal of Mathematical Biology, 32 (5):465–489, May 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160169.

Sulsky:1994:NSS

[969] Deborah Sulsky. Numerical solution of structured population models. Journal of Mathematical Biology, 32(5):491–514, May 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00160170.

Steiner:1994:SLG

[970] Wilfried Steiner and Hans-Rolf Gregorius. Single-locus gametophytic incompatibility: the symmetric equilibrium is globally asymptotically stable. Journal of Mathematical Biology, 32(6):515–520, ???? 1994. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00573457.

Boulanger:1994:SPN

[971] E. N. Boulanger. Small perturbations in nonlinear age-structured population equations. Journal of Mathematical Biology, 32(6):521–533, ????

1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00573458.

Cooke:1994:SID

[972] Kenneth L. Cooke and Janos Turi. Stability, instability in delay equations modeling human respiration. Journal of Mathematical Biology, 32 (6):535–543, ???? 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00573459.

Phillipson:1994:MDA

[973] Paul E. Phillipson and Peter Schuster. Map dynamics of autocatalytic networks and the replicator equations. Journal of Mathematical Biology, 32(6):545–562, ???? 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00573460.

Meier:1994:SLS

[974] Christoph Meier, Walter Senn, Rudolf Hauser, and Manfred Zimmermanne. Strange limits of stability in host-parasitoid systems. Journal of Mathematical Biology, 32(6):563–572, ???? 1994. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00573461.

Busenberg:1994:MME

[975] Stavros Busenberg and Betty Tang. Mathematical models of the early embryonic cell cycle: the role of MPF activation and cyclin degradation. Journal of Mathematical Biology, 32(6):573–596, ???? 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00573462.

Broom:1994:SMG

[976] M. Broom, C. Cannings, and G. T. Vickers. Sequential methods for generating patterns of ESS’s. Journal of Mathematical Biology, 32(6): 597–615, ???? 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00573463.

Lubkin:1994:ORD

[977] Sharon Lubkin and Richard Rand. Oscillatory reaction–diffusion equations on rings. Journal of Mathematical Biology, 32(6):617–632, ???? 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00573464.

Akin:1994:AHW

[978] Ethan Akin and Joseph M. Szucs. Approaches to the Hardy–Weinberg manifold. Journal of Mathematical Biology, 32(7):633–643, August 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00163019.

Tegeder:1994:EPC

[979] Roland W. Tegeder. On the emergence of primitive cooperation in an environment of fluctuating resource supply. Journal of Mathematical Biology, 32(7):645–662, August 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00163020.

Barradas:1994:DAC

[980] Ignacio Barradas and Joel E. Cohen. Disturbances allow coexistence of competing species. Journal of Mathematical Biology, 32(7): 663–676, August 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163021.

Ward:1994:FCA

[981] Michael D. Ward and Daniel A. Hammer. Focal contact assembly through cytoskeletal polymerization: steady state analysis. Journal of Mathematical Biology, 32(7):677–704, August 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00163022.

Cushing:1994:DHA

[982] J. M. Cushing. The dynamics of hierarchical age-structured populations. Journal of Mathematical Biology, 32(7):705–729, August 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00163023.

Hsu:1994:GAM

[983] Sze-Bi Hsu, Paul Waltman, and Gail S. K. Wolkowicz. Global analysis of a model of plasmid-bearing, plasmid-free competition in a chemostat. Journal of Mathematical Biology, 32(7):731–742, August 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00163024.

Pascual:1994:PRP

[984] Mercedes Pascual. Periodic response to periodic forcing of the droop equations for phytoplankton growth. Journal of Mathematical Biology, 32 (8):743–759, October 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168795.

Schwegler:1994:FCC

[985] Helmut Schwegler and Michael C. Mackey. Fluctuations in circulating cell numbers following chemotherapy or bone marrow transplant. Journal of Mathematical Biology, 32(8):761–770, October 1994. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00168796.

Franke:1994:PEO

[986] John E. Franke and Abdul-Aziz Yakubu. Pioneer exclusion in a one-hump discrete pioneer-climax competitive system. Journal of Mathematical Biology, 32(8):771–787, October 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00168797.

Takada:1994:RBT

[987] Takenori Takada and Toshihiko Hara. The relationship between the transition matrix model and the diffusion model. Journal of Mathematical Biology, 32(8):789–807, October 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00168798.

Zhou:1994:PSD

[988] Jinshi Zhou and Herbert W. Hethcote. Population size dependent incidence in models for diseases without immunity. Journal of Mathematical Biology, 32(8):809–834, October 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00168799.

Shnol:1994:SRB

[989] Emmanuil E. Shnol and Alexey S. Kondrashov. Some relations between different characteristics of selection. Journal of Mathematical Biology, 32 (8):835–840, October 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168800.

Chiu:1994:ACS

[990] Chichia Chiu, Frank C. Hoppensteadt, and Willi Jäger. Analysis and computer simulation of accretion patterns in bacterial cultures. Journal of Mathematical Biology, 32(8):841–855, October 1994. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00168801.

Beltrami:1994:MRV

[991] Edward Beltrami and T. O. Carroll. Modeling the role of viral disease in recurrent phytoplankton blooms. Journal of Mathematical Biology, 32 (8):857–863, October 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168802.

Hoppensteadt:1994:RSC

[992] Frank C. Hoppensteadt and Paul M. Johnson. Response of a solutivory chain to a nutrient pulse. Journal of Mathematical Biology, 32 (8):865–867, October 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00168803.

Witelski:1994:AST

[993] Thomas P. Witelski. An asymptotic solution for traveling waves of a nonlinear-diffusion Fisher’s equation. Journal of Mathematical Biology, 33(1):1–16, November 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160171.

Pruss:1994:PAD

[994] Jan Prüss and Wilhelm Schappacher. Persistent age-distributions for a pair-formation model. Journal of Mathematical Biology, 33(1):17– 33, November 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160172.

Gyllenberg:1994:QSD

[995] Mats Gyllenberg and Dmitrii S. Silvestrov. Quasi-stationary distributions of a stochastic metapopulation model. Journal of Mathematical Biology, 33(1):35–70, November 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160173.

Av-Ron:1994:RTP

[996] Evyatar Av-Ron. The role of a transient potassium current in a bursting neuron model. Journal of Mathematical Biology, 33(1):71–87, November 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00160174.

Mackey:1994:GSD

[997] Michael C. Mackey and Ryszard Rudnicki. Global stability in a delayed partial differential equation describing cellular replication. Journal of Mathematical Biology, 33(1):89–109, November 1994. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00160175.

Stadler:1994:INM

[998] Peter F. Stadler, Peter Schuster, and Alan S. Perelson. Immune networks modeled by replicator equations. Journal of Mathematical Biology, 33(2): 111–137, December 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160176.

Grunbaum:1994:TSD

[999] Daniel Grünbaum. Translating stochastic density-dependent individual behavior with sensory constraints to an Eulerian model of animal swarming. Journal of Mathematical Biology, 33(2):139–161, December 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00160177.

Sanchez-Garduno:1994:EUS

[1000] Faustino Sánchez-Garduño and Philip K. Maini. Existence and uniqueness of a sharp travelling wave in degenerate non-linear diffusion Fisher–KPP equations. Journal of Mathematical Biology, 33(2):163– 192, December 1994. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160178.

Cruywagen:1994:TWT

[1001] G. C. Cruywagen, P. K. Maini, and J. D. Murray. Travelling waves in a tissue interaction model for skin pattern formation. Journal of Mathematical Biology, 33(2):193–210, December 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00160179.

Gonzalez-Gunman:1994:AMB

[1002] Jorge González-Gunmán and Raul Naulin. Analysis of a model of bovine brucellosis using singular perturbations. Journal of Mathematical Biology, 33(2):211–223, December 1994. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160180.

Christiansen:1995:GPH

[1003] Freddy Bugge Christiansen, Viggo Andreasen, and Ebbe Thue Poulsen. Genotypic proportions in hybrid zones. Journal of Mathematical Biology, 33(3):225–249, December 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00169562.

Beretta:1995:GSS

[1004] Edoardo Beretta and Yasuhiro Takeuchi. Global stability of an SIR epidemic model with time delays. Journal of Mathematical Biology, 33 (3):250–260, December 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00169563.

Kopell:1995:APS

[1005] Nancy Kopell and David Somers. Anti-phase solutions in relaxation oscillators coupled through excitatory interactions. Journal of Mathematical Biology, 33(3):261–280, December 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00169564.

Dalgaard:1995:TTA

[1006] P. Dalgaard. Transient transport across an inhomogeneous blood-retina barrier. Journal of Mathematical Biology, 33(3):281–294, December 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00169565.

Sherratt:1995:TBT

[1007] Jonathan A. Sherratt. Turing bifurcations with a temporally varying diffusion coefficient. Journal of Mathematical Biology, 33(3):295– 308, December 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00169566.

Baert:1995:AAP

[1008] S. M. Baert, J. Rinzel, and H. Carrillo. Analysis of an autonomous phase model for neuronal parabolic bursting. Journal of Mathematical Biology, 33(3):309–333, December 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00169567.

Calsina:1995:MPS

[1009] Àngel Calsina and Joan Saldaña. A model of physiologically structured population dynamics with a nonlinear individual growth rate. Journal of Mathematical Biology, 33(4):335–364, February 1995. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00176377.

Focardi:1995:MFO

[1010] Stefano Focardi and Paolo Marcellini. A mathematical framework for optimal foraging of herbivores. Journal of Mathematical Biology, 33(4): 365–387, February 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00176378.

Rascle:1995:FTB

[1011] M. Rascle and C. Ziti. Finite time blow-up in some models of chemotaxis. Journal of Mathematical Biology, 33(4):388–414, February 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00176379.

Heesterbeek:1995:TQH

[1012] J. A. P. Heesterbeek and M. G. Roberts. Threshold quantities for Helminth infections. Journal of Mathematical Biology, 33(4):415– 434, February 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00176380.

Ballyk:1995:ETE

[1013] Mary M. Ballyk and Gail S. K. Wolkowicz. An examination of the thresholds of enrichment: a resource-based growth model. Journal of Mathematical Biology, 33(4):435–457, February 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00176381.

Watmough:1995:ODM

[1014] James Watmough and Leah Edelstein-Keshet. A one-dimensional model of trail propagation by army ants. Journal of Mathematical Biology, 33(5):459–476, April 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163038.

Newson:1995:CMP

[1015] Roger Newson. A canonical model for production and distribution of root mass in space and time. Journal of Mathematical Biology, 33 (5):477–488, April 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163039.

Ngwa:1995:STP

[1016] G. A. Ngwa and P. K. Maini. Spatio-temporal patterns in a mechanical model for mesenchymal morphogenesis. Journal of Mathematical Biology, 33(5):489–520, April 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163040.

Kaitala:1995:PDH

[1017] Veijo Kaitala and Wayne M. Getz. Population dynamics and harvesting of semelparous species with phenotypic and genotypic variability in reproductive age. Journal of Mathematical Biology, 33(5):521–556, April 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00163041.

Burger:1995:FPA

[1018] Reinhard Bürger and Warren J. Ewens. Fixation probabilities of additive alleles in diploid populations. Journal of Mathematical Biology, 33(5):557–575, April 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00163042.

Steiner:1995:SND

[1019] Wilfried Steiner and Hans-Rolf Gregorius. A supplementary note on the dynamics of the model of gametophytic incompatibility. Journal of Mathematical Biology, 33(6):577–580, June 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF00298643.

vanHerwaarden:1995:SEM

[1020] Onno A. van Herwaarden and Johan Grasman. Stochastic epidemics: major outbreaks and the duration of the endemic period. Journal of Mathematical Biology, 33(6):581–601, June 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/BF00298644.

Joyce:1995:DRA

[1021] Paul Joyce and Simon Tavaré. The distribution of rare alleles. Journal of Mathematical Biology, 33(6):602–618, June 1995. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00298645.

Mogilner:1995:SCD

[1022] Alex Mogilner and Leah Edelstein-Keshet. Selecting a common direction. Journal of Mathematical Biology, 33(6):619–660, June 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00298646.

Altmann:1995:SIR

[1023] M. Altmann. Susceptible-infected-removed epidemic models with dynamic partnerships. Journal of Mathematical Biology, 33(6):661– 675, June 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00298647.

Higgs:1995:AMA

[1024] Paul G. Higgs and Glenn Woodcock. The accumulation of mutations in asexual populations and the structure of genealogical trees in the presence of selection. Journal of Mathematical Biology, 33(7): 677–702, August 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00184644.

Nowak:1995:ARG

[1025] Martin A. Nowak, Karl Sigmund, and Esam El-Sedy. Automata, repeated games and noise. Journal of Mathematical Biology, 33(7): 703–722, August 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00184645.

Buffoni:1995:SDC

[1026] Giuseppe Buffoni, Annalisa Griffa, Zhengyuan Li, and Piero de Mottoni. Spatially distributed communities: the resource-consumer system. Journal of Mathematical Biology, 33(7):723–743, August 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00184646.

Chaplain:1995:MAM

[1027] M. A. J. Chaplain, Susan M. Giles, B. D. Sleeman, and R. J. Jarvis. A mathematical analysis of a model for tumour angiogenesis. Journal of Mathematical Biology, 33(7):744–770, August 1995. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00184647.

Stephenson:1995:WNS

[1028] Laura E. Stephenson and David J. Wollkind. Weakly nonlinear stability analyses of one-dimensional Turing pattern formation in activatorinhibitor/immobilizer model systems. Journal of Mathematical Biology, 33(8):771–815, October 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00187282.

Ardito:1995:LFG

[1029] A. Ardito and P. Ricciardi. Lyapunov functions for a generalized Gausetype model. Journal of Mathematical Biology, 33(8):816–828, October 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00187283.

Willner:1995:SOO

[1030] Barry E. Willner, Chien-Ping Lu, and Willard L. Miranker. Selforganization of an oscillatory neural system. Journal of Mathematical Biology, 33(8):829–866, October 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00187284.

Yashin:1995:EHH

[1031] Anatoli I. Yashin, Kenneth G. Manton, Max A. Woodbury, and Eric Stallard. The effects of health histories on stochastic process models of aging and mortality. Journal of Mathematical Biology, 34(1):1–16, ???? 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00180134.

Yan:1995:FDP

[1032] Joseph G.-G. Yan. Formation and deformation of patterns through diffusion. Journal of Mathematical Biology, 34(1):17–40, ???? 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00180135.

Luo:1995:GAM

[1033] Ting-Kung Luo and Sze-Bi Hsu. Global analysis of a model of plasmidbearing, plasmid-free competition in a chemostat with inhibitions. Journal of Mathematical Biology, 34(1):41–76, ???? 1995. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00180136.

Lubkin:1995:MEB

[1034] Sharon R. Lubkin and J. D. Murray. A mechanism for early branching in lung morphogenesis. Journal of Mathematical Biology, 34(1):77–94, ???? 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00180137.

Stephan:1995:PAT

[1035] Wolfgang Stephan. Perturbation analysis of a two-locus model with directional selection and recombination. Journal of Mathematical Biology, 34(1):95–109, ???? 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00180138.

Kornet:1995:IEC

[1036] D. J. Kornet, J. A. J. Metz, and H. A. J. M. Schellinx. Internodons as equivalence classes in genealogical networks: building-blocks for a rigorous species concept. Journal of Mathematical Biology, 34 (1):110–122, ???? 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00180139.

Grist:1995:SSS

[1037] Eric P. M. Grist and William S. C. Gurney. Stage-specificity and the synchronisation of life-cycles to periodic environmental variations. Journal of Mathematical Biology, 34(2):123–147, ???? 1995. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00178770.

Ruuth:1995:IEM

[1038] Steven J. Ruuth. Implicit-explicit methods for reaction–diffusion problems in pattern formation. Journal of Mathematical Biology, 34(2): 148–176, ???? 1995. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00178771.

Hethcote:1995:SEM

[1039] Herbert W. Hethcote and P. van den Driessche. An SIS epidemic model with variable population size and a delay. Journal of Mathematical Biology, 34(2):177–194, ???? 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00178772.

Jansen:1995:EDT

[1040] V. A. A. Jansen. Effects of dispersal in a tri-trophic metapopulation model. Journal of Mathematical Biology, 34(2):195–224, ???? 1995. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00178773.

Hsu:1995:CBP

[1041] S. B. Hsu, Ting-Kung Luo, and Paul Waltman. Competition between plasmid-bearing and plasmid-free organisms in a chemostat with an inhibitor. Journal of Mathematical Biology, 34(2):225–238, ???? 1995. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00178774.

Liberman:1996:MRS

[1042] Uri Liberman and Marcus W. Feldman. On the modification of recombination with sex-dependent fitnesses and linkage. Journal of Mathematical Biology, 34(3):239–252, February 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00160495.

Mesterton-Gibbons:1996:WAO

[1043] Michael Mesterton-Gibbons. On the war of attrition and other games among kin. Journal of Mathematical Biology, 34(3):253–270, February 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF00160496.

Bardsley:1996:STF

[1044] W. G. Bardsley and E. K. Kyprianou. A statistical theory for flow cytometry profiles in terms of the binding of ligands to cell surface receptors and changes in gene expression. Journal of Mathematical Biology, 34 (3):271–296, February 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160497.

Gourley:1996:PPR

[1045] S. A. Gourley and N. F. Britton. A predator–prey reaction–diffusion system with nonlocal effects. Journal of Mathematical Biology, 34(3): 297–333, February 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160498.

Nagylaki:1996:DMM

[1046] Thomas Nagylaki. The diffusion model for migration and selection in a dioecious population. Journal of Mathematical Biology, 34(3): 334–360, February 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00160499.

Tranquillo:1996:SMR

[1047] Robert T. Tranquillo and Wolfgang Alt. Stochastic model of receptormediated cytomechanics and dynamic morphology of leukocytes. Journal of Mathematical Biology, 34(4):361–412, March 1996. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF00167941.

Castillo-Chavez:1996:ESP

[1048] Carlos Castillo-Chavez, Wenzhang Huang, and Jia Li. On the existence of stable pairing distributions. Journal of Mathematical Biology, 34(4):413–441, March 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00167942.

Franke:1996:ESA

[1049] John E. Franke and Abdul-Aziz Yakubu. Extinction of species in agestructured, discrete noncooperative systems. Journal of Mathematical Biology, 34(4):442–454, March 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00167943.

Allen:1996:DCM

[1050] Edward J. Allen, Linda J. S. Allen, and Xiaoning Gilliam. Dispersal and competition models for plants. Journal of Mathematical Biology, 34 (4):455–481, March 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00167944.

Diekmann:1996:E

[1051] Odo Diekmann, Freddy Christiansen, and Richard Law. Editorial. Journal of Mathematical Biology, 34(5–6):483, ???? 1996. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/BF02409746.

Eshel:1996:CCE

[1052] Ilan Eshel. On the changing concept of evolutionary population stability as a reflection of a changing point of view in the quantitative theory of evolution. Journal of Mathematical Biology, 34(5–6):485–510, ???? 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF02409747.

Hammerstein:1996:DAP

[1053] Peter Hammerstein. Darwinian adaptation, population genetics and the streetcar theory of evolution. Journal of Mathematical Biology, 34(5– 6):511–532, ???? 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF02409748.

Weissing:1996:GVP

[1054] Franz J. Weissing. Genetic versus phenotypic models of selection: can genetics be neglected in a long-term perspective? Journal of Mathematical Biology, 34(5–6):533–555, ???? 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF02409749.

Marrow:1996:EDP

[1055] Paul Marrow, Ulf Dieckmann, and Richard Law. Evolutionary dynamics of predator–prey systems: an ecological perspective. Journal of Mathematical Biology, 34(5–6):556–578, ???? 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/BF02409750.

Dieckmann:1996:DTC

[1056] Ulf Dieckmann and Richard Law. The dynamical theory of coevolution: a derivation from stochastic ecological processes. Journal of Mathematical Biology, 34(5–6):579–612, ???? 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF02409751.

Matessi:1996:LTE

[1057] Carlo Matessi and Cristina Di Pasquale. Long-term evolution of multilocus traits. Journal of Mathematical Biology, 34(5–6):613–653, ???? 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF02409752.

Taylor:1996:IFA

[1058] Peter D. Taylor. Inclusive fitness arguments in genetic models of behaviour. Journal of Mathematical Biology, 34(5–6):654–674, ???? 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF02409753.

Hofbauer:1996:EDB

[1059] Josef Hofbauer. Evolutionary dynamics for bimatrix games: A Hamiltonian system? Journal of Mathematical Biology, 34(5–6):675–688, ???? 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF02409754.

Morimoto:1996:DME

[1060] Hiroshi Morimoto. A diffusion model for the evolution of metalloporphyrin. Journal of Mathematical Biology, 34(7):689–706, August 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00161515.

Cressman:1996:ESS

[1061] R. Cressman, J. Hofbauer, and W. G. S. Hines. Evolutionary stability in strategic models of single-locus frequency-dependent viability selection. Journal of Mathematical Biology, 34(7):707–733, August 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00161516.

Smith:1996:DSS

[1062] H. L. Smith. A discrete, size-structured model of microbial growth and competition in the chemostat. Journal of Mathematical Biology, 34(7): 734–754, August 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00161517.

Henson:1996:HMI

[1063] Shandelle M. Henson and J. M. Cushing. Hierarchical models of intraspecific competition: scramble versus contest. Journal of Mathematical Biology, 34(7):755–772, August 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF00161518.

Liberman:1996:CPI

[1064] U. Liberman and H. Manos. C–L processes involving higher order positive (negative) interference for any two collections of disjoint chromosomal regions. Journal of Mathematical Biology, 34(7):773–788, August 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF00161519.

Merino:1996:CCT

[1065] Sandro Merino. Cyclic competition of three species in the time periodic and diffusive case. Journal of Mathematical Biology, 34(7): 789–809, August 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF00161520.

Mogilner:1996:SCD

[1066] Alexander Mogilner, Leah Edelstein-Keshet, and G. Bard Ermentrout. Selecting a common direction. Journal of Mathematical Biology, 34(8): 811–842, September 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF01834821.

Fierro:1996:LSA

[1067] Raúl Fierro. Large-sample analysis for a stochastic epidemic model and its parameter estimators. Journal of Mathematical Biology, 34(8): 843–856, September 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/BF01834822.

Gourley:1996:MSD

[1068] S. A. Gourley, N. F. Britton, M. A. J. Chaplain, and H. M. Byrne. Mechanisms for stabilisation and destabilisation of systems of reaction–diffusion equations. Journal of Mathematical Biology, 34(8):857–877, September 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF01834823.

Gatton:1996:MPT

[1069] M. Gatton, W. Hogarth, A. Saul, and P. Dayananda. A model for predicting the transmission rate of malaria from serological data. Journal of Mathematical Biology, 34(8):878–888, September 1996. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/BF01834824.

Skalak:1996:CGR

[1070] Richard Skalak, Stephen Zargaryan, Rakesh K. Jain, Paolo A. Netti, and Anne Hoger. Compatibility and the genesis of residual stress by volumetric growth. Journal of Mathematical Biology, 34(8):889–914, September 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/BF01834825.

Schinazi:1996:IPS

[1071] Rinaldo Schinazi. On an interacting particle system modeling an epidemic. Journal of Mathematical Biology, 34(8):915–925, September 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/BF01834826.

Fischbacher:1996:FMP

[1072] Urs Fischbacher. Finding the maximum a posteriori probability (MAP) in a Bayesian taxonomic key is NP-hard. Journal of Mathematical Biology, 34(8):926–936, September 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/BF01834827.

Beretta:1996:MMD

[1073] Edoardo Beretta, Fortunata Solimano, and Yasuhiro Takeuchi. A mathematical model for drug administration by using the phagocytosis of red blood cells. Journal of Mathematical Biology, 35(1):1–19, November 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/s002850050039.

Lizana:1996:MBM

[1074] Marcos Lizana and Jesus Rivero. Multiparametric bifurcations for a model in epidemiology. Journal of Mathematical Biology, 35(1):21– 36, November 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s002850050040.

Schreiber:1996:GSC

[1075] Sebastian J. Schreiber. Global stability in consumer-resource cascades. Journal of Mathematical Biology, 35(1):37–48, November 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/s002850050041.

Hillen:1996:TMC

[1076] Thomas Hillen. A Turing model with correlated random walk. Journal of Mathematical Biology, 35(1):49–72, November 1996. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/s002850050042.

Watanabe:1996:SCC

[1077] Mari Watanabe and Robert F. Gilmour, Jr. Strategy for control of complex low-dimensional dynamics in cardiac tissue. Journal of Mathematical Biology, 35(1):73–87, November 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s002850050043.

Gimelfarb:1996:SAR

[1078] A. Gimelfarb. Some additional results about polymorphisms in models of an additive quantitative trait under stabilizing selection. Journal of Mathematical Biology, 35(1):88–96, November 1996. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/s002850050044.

Schwarz:1996:MTM

[1079] Richard I. Schwarz. Modeling tendon morphogenesis in vivo based on cell density signaling in cell culture. Journal of Mathematical Biology, 35 (1):97–113, November 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s002850050045.

Soto-Trevino:1996:PBR

[1080] C. Soto-Treviño, N. Kopell, and D. Watson. Parabolic bursting revisited. Journal of Mathematical Biology, 35(1):114–128, November 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s002850050046.

vandenBosch:1996:DID

[1081] Frank van den Bosch and André M. de Roos. The dynamics of infectious diseases in orchards with roguing and replanting as control strategy. Journal of Mathematical Biology, 35(2):129–157, December 1996. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s002850050047.

Rinaldi:1996:RBS

[1082] S. Rinaldi, S. Dal Bo, and E. De Nittis. On the role of body size in a tritrophic metapopulation model. Journal of Mathematical Biology, 35(2):

158–176, December 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s002850050048.

Herrero:1996:CCK

[1083] Miguel A. Herrero and Juan J. L. Velázquez. Chemotactic collapse for the Keller–Segel model. Journal of Mathematical Biology, 35(2): 177–194, December 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s002850050049.

Wikan:1996:ORG

[1084] Arild Wikan and Einar Mjølhus. Overcompensatory recruitment and generation delay in discrete age-structured population models. Journal of Mathematical Biology, 35(2):195–239, December 1996. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/article/10.1007/s002850050050.

Cooke:1996:ASE

[1085] K. L. Cooke and P. van den Driessche. Analysis of an SEIRS epidemic model with two delays. Journal of Mathematical Biology, 35(2): 240–260, December 1996. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s002850050051.

Bonneuil:1997:VPP

[1086] Noël Bonneuil and Katharina Müllers. Viable populations in a preypredator system. Journal of Mathematical Biology, 35(3):261–293, February 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050052.pdf.

Charter:1997:HCM

[1087] Kevin Charter and Thomas Rogers. Homoclinic chaos in a model of natural selection. Journal of Mathematical Biology, 35(3):294–320, February 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050053.pdf.

Baake:1997:BHD

[1088] Ellen Baake and Thomas Wiehe. Bifurcations in haploid and diploid sequence space models. Journal of Mathematical Biology, 35(3):321– 343, February 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050054.pdf.

Satulovsky:1997:SIL

[1089] Javier E. Satulovsky and Tânia Tomé. Spatial instabilities and local oscillations in a lattice gas Lotka–Volterra model. Journal of Mathematical Biology, 35(3):344–358, February 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050055.pdf.

Glendinning:1997:MAC

[1090] Paul Glendinning and Louise P. Perry. Melnikov analysis of chaos in a simple epidemiological model. Journal of Mathematical Biology, 35(3): 359–373, February 1997. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850050056.pdf.

Behncke:1997:OMF

[1091] Horst Behncke. Optimization models for the force and energy in competitive running. Journal of Mathematical Biology, 35(4):375– 390, March 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050057.pdf.

Kirzhner:1997:MDU

[1092] Valery Kirzhner and Yuri Lyubich. Multilocus dynamics under haploid selection. Journal of Mathematical Biology, 35(4):391–408, March 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050058.pdf.

Nagylaki:1997:DMM

[1093] Thomas Nagylaki. The diffusion model for migration and selection in a plant population. Journal of Mathematical Biology, 35(4):409– 431, March 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050059.pdf.

Dodd:1997:POA

[1094] R. K. Dodd. Periodic orbits arising from Hopf bifurcations in a Volterra prey-predator model. Journal of Mathematical Biology, 35(4):432– 452, March 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050060.pdf.

Tang:1997:PDC

[1095] Betty Tang, Ann Sitomer, and Trachette Jackson. Population dynamics and competition in chemostat models with adaptive nutrient uptake. Journal of Mathematical Biology, 35(4):453–479, March 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850050061.pdf.

Ackleh:1997:MAG

[1096] Azmy S. Ackleh and Ben G. Fitzpatrick. Modeling aggregation and growth processes in an algal population model: analysis and computations. Journal of Mathematical Biology, 35(4):480–502, March 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050062.pdf.

Castillo-Chavez:1997:EFS

[1097] Carlos Castillo-Chavez, Wenzhang Huang, and Jia Li. The effects of females’ susceptibility on the coexistence of multiple pathogen strains of sexually transmitted diseases. Journal of Mathematical Biology, 35(5): 503–522, April 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050063.pdf.

Feng:1997:CEV

[1098] Zhilan Feng and Jorge X. Velasco-Hernández. Competitive exclusion in a vector-host model for the dengue fever. Journal of Mathematical Biology, 35(5):523–544, April 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050064.pdf.

Smith:1997:PFD

[1099] H. L. Smith. The periodically forced droop model for phytoplankton growth in a chemostat. Journal of Mathematical Biology, 35(5):545– 556, April 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050065.pdf.

Vellekoop:1997:UFC

[1100] M. H. Vellekoop and G. Högnäs. A unifying framework for chaos and stochastic stability in discrete population models. Journal of Mathematical Biology, 35(5):557–588, April 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050066.pdf.

Davidson:1997:TWP

[1101] F. A. Davidson, B. D. Sleeman, A. D. M. Rayner, J. W. Crawford, and K. Ritz. Travelling waves and pattern formation in a model for fungal development. Journal of Mathematical Biology, 35(5):589– 608, April 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050067.pdf.

Nauroschat:1997:TAG

[1102] Jürgen Nauroschat and Uwe an der Heiden. A theoretical approach to G-protein modulation of cellular responsiveness. Journal of Mathematical Biology, 35(5):609–627, April 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050068.pdf.

Castillo-Chavez:1997:TTC

[1103] Carlos Castillo-Chavez and Zhilan Feng. To treat or not to treat: the case of tuberculosis. Journal of Mathematical Biology, 35(6):629– 656, June 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050069.pdf.

Lemire:1997:NEO

[1104] Mathieu Lemire and Sabin Lessard. On the non-existence of an optimal migration rate. Journal of Mathematical Biology, 35(6):657– 682, June 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050070.pdf.

Feng:1997:CEP

[1105] Wei Feng. Competitive exclusion and persistence in models of resource and sexual competition. Journal of Mathematical Biology, 35(6):683– 694, June 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050071.pdf.

Witelski:1997:SMD

[1106] Thomas P. Witelski. Segregation and mixing in degenerate diffusion in population dynamics. Journal of Mathematical Biology, 35(6):695– 712, June 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050072.pdf.

Sanchez-Garduno:1997:TWP

[1107] Faustino Sánchez-Garduño and Philip K. Maini. Travelling wave phenomena in non-linear diffusion degenerate Nagumo equations. Journal of Mathematical Biology, 35(6):713–728, June 1997. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050073.pdf.

Powell:1997:CSF

[1108] James A. Powell. Conditional stability of front solutions. Journal of Mathematical Biology, 35(6):729–747, June 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850050074.pdf.

Lewis:1997:AMW

[1109] M. A. Lewis, K. A. J. White, and J. D. Murray. Analysis of a model for wolf territories. Journal of Mathematical Biology, 35(7):749– 774, August 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050075.pdf.

Kirschner:1997:OCC

[1110] Denise Kirschner, Suzanne Lenhart, and Steve Serbin. Optimal control of the chemotherapy of HIV. Journal of Mathematical Biology, 35(7): 775–792, August 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050076.pdf.

vanHerwaarden:1997:SEP

[1111] Onno A. van Herwaarden. Stochastic epidemics: the probability of extinction of an infectious disease at the end of a major outbreak. Journal of Mathematical Biology, 35(7):793–813, August 1997. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050077.pdf.

Teso:1997:SPB

[1112] A. Teso, E. Drigo Filho, and A. Agostinho Neto. Solution of the Poisson– Boltzmann equation for a system with four ionic species. Journal of Mathematical Biology, 35(7):814–824, August 1997. CODEN JMBLAJ.

ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850050078.pdf.

Andreasen:1997:DCI

[1113] Viggo Andreasen, Juan Lin, and Simon A. Levin. The dynamics of cocirculating influenza strains conferring partial cross-immunity. Journal of Mathematical Biology, 35(7):825–842, August 1997. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050079.pdf.

Gourley:1997:PDI

[1114] S. A. Gourley and M. V. Bartuccelli. Parameter domains for instability of uniform states in systems with many delays. Journal of Mathematical Biology, 35(7):843–867, August 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050080.pdf.

Skalak:1997:KSG

[1115] R. Skalak, D. A. Farrow, and A. Hoger. Kinematics of surface growth. Journal of Mathematical Biology, 35(8):869–907, September 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050081.pdf.

Huyer:1997:PCS

[1116] Waltraud Huyer. On periodic cohort solutions of a size-structured population model. Journal of Mathematical Biology, 35(8):908–934, September 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050082.pdf.

Hart:1997:SMS

[1117] Deborah R. Hart and Robert H. Gardner. A spatial model for the spread of invading organisms subject to competition. Journal of Mathematical Biology, 35(8):935–948, September 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050083.pdf.

Allegretto:1997:PSC

[1118] W. Allegretto, H. Xie, and Shixin Yang. Properties of solutions for a chemotaxis system. Journal of Mathematical Biology, 35(8):949–966, September 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050084.pdf.

Calsina:1997:ABM

[1119] Àngel Calsina and Joan Saldaña. Asymptotic behaviour of a model of hierarchically structured population dynamics. Journal of Mathematical Biology, 35(8):967–987, September 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050085.pdf.

Hoppensteadt:1997:WPM

[1120] F. C. Hoppensteadt and H. D. Mittelmann. Wave propagation in mathematical models of striated cortex. Journal of Mathematical Biology, 35(8):988–994, September 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050086.pdf.

Perlitz:1997:MVN

[1121] Michael Perlitz and Wolfgang Stephan. The mean and variance of the number of segregating sites since the last hitchhiking event. Journal of Mathematical Biology, 36(1):1–23, November 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850050087.pdf.

Kooi:1997:CDB

[1122] B. W. Kooi, M. P. Boer, and S. A. L. M. Kooijman. Complex dynamic behaviour of autonomous microbial food chains. Journal of Mathematical Biology, 36(1):24–40, November 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050088.pdf.

Ottesen:1997:MBF

[1123] Johnny T. Ottesen. Modelling of the baroreflex-feedback mechanism with time-delay. Journal of Mathematical Biology, 36(1):41–63, November 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050089.pdf.

Coward:1997:ETF

[1124] Eivind Coward. Equivalence of two Fourier methods for biological sequences. Journal of Mathematical Biology, 36(1):64–70, November 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050090.pdf.

Shukla:1997:MDC

[1125] J. B. Shukla and B. Dubey. Modelling the depletion and conservation of forestry resources: effects of population and pollution. Journal of Mathematical Biology, 36(1):71–94, November 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050091.pdf.

Chavez-Ross:1997:MCO

[1126] A. Chávez-Ross, S. Franks, H. D. Mason, K. Hardy, and J. Stark. Modelling the control of ovulation and polycystic ovary syndrome. Journal of Mathematical Biology, 36(1):95–118, November 1997. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050092.pdf.

Sturis:1997:LGB

[1127] J. Sturis and M. Brøns. Local and global bifurcations at infinity in models of glycolytic oscillations. Journal of Mathematical Biology, 36(2):119– 132, November 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050093.pdf.

Schreiber:1997:GSP

[1128] Sebastian J. Schreiber. Generalist and specialist predators that mediate permanence in ecological communities. Journal of Mathematical Biology, 36(2):133–148, November 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050094.pdf.

Collings:1997:EFR

[1129] John B. Collings. The effects of the functional response on the bifurcation behavior of a mite predator–prey interaction model. Journal of Mathematical Biology, 36(2):149–168, November 1997. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850050095.pdf.

Hanson:1997:PGR

[1130] Floyd B. Hanson and Henry C. Tuckwell. Population growth with randomly distributed jumps. Journal of Mathematical Biology, 36(2):169– 187, November 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050096.pdf.

Notohara:1997:NSS

[1131] Morihiro Notohara. The number of segregating sites in a sample of DNA sequences from a geographically structured population. Journal of Mathematical Biology, 36(2):188–200, November 1997. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050097.pdf.

Henson:1997:EPH

[1132] Shandelle M. Henson and J. M. Cushing. The effect of periodic habitat fluctuations on a nonlinear insect population model. Journal of Mathematical Biology, 36(2):201–226, November 1997. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850050098.pdf.

Dushoff:1998:BBC

[1133] Jonathan Dushoff, Wenzhang Huang, and Carlos Castillo-Chavez. Backwards bifurcations and catastrophe in simple models of fatal diseases. Journal of Mathematical Biology, 36(3):227–248, February 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850050099.pdf.

Esipov:1998:KMP

[1134] Sergei E. Esipov and J. A. Shapiro. Kinetic model of Proteus mirabilis swarm colony development. Journal of Mathematical Biology, 36(3):249– 268, February 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050100.pdf.

Bees:1998:AAO

[1135] M. A. Bees, N. A. Hill, and T. J. Pedley. Analytical approximations for the orientation distribution of small dipolar particles in steady shear flows. Journal of Mathematical Biology, 36(3):269–298, February 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050101.pdf.

Beder:1998:CSG

[1136] Jay H. Beder and Richard Gomulkiewicz. Computing the selection gradient and evolutionary response of an infinite-dimensional trait. Journal of Mathematical Biology, 36(3):299–319, February 1998. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050102.pdf.

Plahte:1998:MBD

[1137] Erik Plahte, Thomas Mestl, and Stig W. Omholt. A methodological basis for description and analysis of systems with complex switchlike interactions. Journal of Mathematical Biology, 36(4):321–348, March 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050103.pdf.

Diekmann:1998:FAG

[1138] Odo Diekmann, Mats Gyllenberg, J. A. J. Metz, and Horst R. Thieme. On the formulation and analysis of general deterministic structured population models I. Linear theory. Journal of Mathematical Biology, 36(4): 349–388, March 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050104.pdf.

Kuang:1998:GQA

[1139] Yang Kuang and Edoardo Beretta. Global qualitative analysis of a ratio-dependent predator–prey system. Journal of Mathematical Biology, 36(4):389–406, March 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050105.pdf.

Zegeling:1998:ULC

[1140] André Zegeling and Robert E. Kooij. Uniqueness of limit cycles in models for microparasitic and macroparasitic diseases. Journal of Mathematical Biology, 36(4):407–417, March 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050106.pdf.

Pugliese:1998:AMM

[1141] A. Pugliese, R. Rosà, and M. L. Damaggio. Analysis of a model for macroparasitic infection with variable aggregation and clumped infections. Journal of Mathematical Biology, 36(5):419–447, March 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050107.pdf.

Manos:1998:DCF

[1142] H. Manos and U. Liberman. Discrete chiasma formation models and their associated high order interference. Journal of Mathematical Biology, 36(5):448–468, March 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050108.pdf.

Ellner:1998:SIL

[1143] Stephen P. Ellner, Akira Sasaki, Yoshihiro Haraguchi, and Hirotsugu Matsuda. Speed of invasion in lattice population models: pairedge approximation. Journal of Mathematical Biology, 36(5):469–484, March 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050109.pdf.

Haccou:1998:ROM

[1144] Patsy Haccou and Yoh Iwasa. Robustness of optimal mixed strategies. Journal of Mathematical Biology, 36(5):485–496, March 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850050110.pdf.

Takada:1998:TIP

[1145] Takenori Takada and Hisao Nakajima. Theorems on the invasion process in stage-structured populations with density-dependent dynamics. Journal of Mathematical Biology, 36(5):497–514, March 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850050111.pdf.

Edelstein-Keshet:1998:DTB

[1146] Leah Edelstein-Keshet, James Watmough, and Daniel Grunbaum. Do travelling band solutions describe cohesive swarms? an investigation for migratory locusts. Journal of Mathematical Biology, 36(6):515– 549, July 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050112.pdf.

Doebeli:1998:SGM

[1147] Michael Doebeli and Gerdien de Jong. A simple genetic model with non-equilibrium dynamics. Journal of Mathematical Biology, 36(6):550– 556, July 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050113.pdf.

Luo:1998:TMP

[1148] Zong-Ping Luo and Kai-Nan An. A theoretical model to predict distribution of the fabric tensor and apparent density in cancellous bone. Journal of Mathematical Biology, 36(6):557–568, July 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850050114.pdf.

Wu:1998:AED

[1149] H.-Y. Wu and S. M. Baer. Analysis of an excitable dendritic spine with an activity-dependent stem conductance. Journal of Mathematical Biology, 36(6):569–592, July 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050115.pdf.

Stiefenhofer:1998:QSS

[1150] Matthias Stiefenhofer. Quasi-steady-state approximation for chemical reaction networks. Journal of Mathematical Biology, 36(6):593– 609, July 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050116.pdf.

Gueron:1998:SSD

[1151] Shay Gueron. The steady-state distributions of coagulation–fragmentation processes. Journal of Mathematical Biology, 37(1):1–27, July 1998. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s002850050117.
pdf.

vandenBerg:1998:OAB

[1152] Hugo A. van den Berg, Yuri N. Kiselev, S. A. L. M. Kooijman, and Michael V. Orlov. Optimal allocation between nutrient uptake and growth in a microbial trichome. Journal of Mathematical Biology, 37 (1):28–48, July 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050118.pdf.

Yakovlev:1998:SMB

[1153] A. Y. Yakovlev, M. Mayer-Proschel, and M. Noble. A stochastic model of brain cell differentiation in tissue culture. Journal of Mathematical Biology, 37(1):49–60, July 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050119.pdf.

Dockery:1998:ESD

[1154] Jack Dockery, Vivian Hutson, Konstantin Mischaikow, and Mark Pernarowski. The evolution of slow dispersal rates: a reaction diffusion model. Journal of Mathematical Biology, 37(1):61–83, July 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050120.pdf.

Shen:1998:CAP

[1155] Wenxian Shen and Yingfei Yi. Convergence in almost periodic Fisher and Kolmogorov models. Journal of Mathematical Biology, 37(1):84– 102, July 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050121.pdf.

Cantrell:1998:ESH

[1156] Robert Stephen Cantrell and Chris Cosner. On the effects of spatial heterogeneity on the persistence of interacting species. Journal of Mathematical Biology, 37(2):103–145, August 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050122.pdf.

Garay:1998:GRW

[1157] József Garay and M. Barnabas Garay. Genetical reachability: When does a sexual population realize all phenotypic states? Journal of Mathematical Biology, 37(2):146–154, August 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050123.pdf.

Alvarez:1998:OHS

[1158] Luis H. R. Alvarez and Larry A. Shepp. Optimal harvesting of stochastically fluctuating populations. Journal of Mathematical Biology, 37(2): 155–177, August 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050124.pdf.

Vandermeer:1998:MCP

[1159] John Vandermeer and Andrew Kaufmann. Models of coupled population oscillators using 1-D maps. Journal of Mathematical Biology, 37(2):178– 202, August 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050125.pdf.

Abbiw-Jackson:1998:GIO

[1160] Roselyn M. Abbiw-Jackson and William F. Langford. Gain-induced oscillations in blood pressure. Journal of Mathematical Biology, 37(3):203– 234, September 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050126.pdf.

Kirschner:1998:MIT

[1161] Denise Kirschner and John Carl Panetta. Modeling immunotherapy of the tumor- immune interaction. Journal of Mathematical Biology, 37(3): 235–252, September 1998. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850050127.pdf.

He:1998:GSC

[1162] Xue-Zhong He and Shigui Ruan. Global stability in chemostat-type plankton models with delayed nutrient recycling. Journal of Mathematical Biology, 37(3):253–271, September 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050128.pdf.

Roberts:1998:SPM

[1163] M. G. Roberts and J. A. P. Heesterbeek. A simple parasite model with complicated dynamics. Journal of Mathematical Biology, 37(3):272–290, September 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050129.pdf.

Bernard:1998:QBS

[1164] Olivier Bernard and Sami Souissi. Qualitative behavior of stagestructured populations: application to structural validation. Journal of Mathematical Biology, 37(4):291–308, October 1998. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050130.pdf.

Henson:1998:LMM

[1165] Shandelle M. Henson. Leslie matrix models as “stroboscopic snapshots” of McKendrick PDE models. Journal of Mathematical Biology, 37(4): 309–328, October 1998. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850050131.pdf.

Bruni:1998:ISM

[1166] C. Bruni, L. Ferrante, and G. Koch. Identifiability of a stochastic model for cell debris in flow cytometry. Journal of Mathematical Biology, 37(4):329–340, October 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050132.pdf.

Baker:1998:MAT

[1167] C. T. H. Baker, G. A. Bocharov, C. A. H. Paul, and F. A. Rihan. Modelling and analysis of time-lags in some basic patterns of cell proliferation. Journal of Mathematical Biology, 37(4):341–371, October 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850050133.pdf.

Chiu:1998:ETP

[1168] Chuang-Hsiung Chiu and Sze-Bi Hsu. Extinction of top-predator in a three-level food-chain model. Journal of Mathematical Biology, 37(4): 372–380, October 1998. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850050134.pdf.

Benson:1998:UTB

[1169] Debbie L. Benson, Philip K. Maini, and Jonathan A. Sherratt. Unravelling the Turing bifurcation using spatially varying diffusion coefficients. Journal of Mathematical Biology, 37(5):381–417, November 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050135.pdf.

Taylor:1998:SME

[1170] David Taylor and Philip Holmes. Simple models for excitable and oscillatory neural networks. Journal of Mathematical Biology, 37(5):419– 446, November 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050136.pdf.

Hutson:1998:APP

[1171] V. Hutson and K. Mischaikow. An approach to practical persistence. Journal of Mathematical Biology, 37(5):447–466, November 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050137.pdf.

Gatto:1998:ICA

[1172] M. Gatto and G. A. De Leo. Interspecific competition among macroparasites in a density-dependent host population. Journal of Mathematical Biology, 37(5):467–490, November 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050138.pdf.

Cantrell:1998:CRI

[1173] Robert Stephen Cantrell, Chris Cosner, and William F. Fagan. Competitive reversals inside ecological reserves: the role of external habitat degradation. Journal of Mathematical Biology, 37(6):491–533, December 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050139.pdf.

Wilkinson-Herbots:1998:GSD

[1174] Hilde M. Wilkinson-Herbots. Genealogy and subpopulation differentiation under various models of population structure. Journal of Mathematical Biology, 37(6):535–585, December 1998. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050140.pdf.

Chen:1999:PWS

[1175] Xinfu Chen and Stuart P. Hastings. Pulse waves for a semi-discrete Morris–Lecar type model. Journal of Mathematical Biology, 38(1):1– 20, February 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002859970001.pdf.

Deng:1999:GIP

[1176] Bo Deng. Glucose-induced period-doubling cascade in the electrical activity of pancreatic β-cells. Journal of Mathematical Biology, 38(1):21– 78, February 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050141.pdf.

Lizana:1999:SDM

[1177] Marcos Lizana and Victor Padron. A spatially discrete model for aggregating populations. Journal of Mathematical Biology, 38(1):79–102, February 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050142.pdf.

Nagylaki:1999:CMS

[1178] Thomas Nagylaki, Josef Hofbauer, and Pavol Brunovský. Convergence of multilocus systems under weak epistasis or weak selection. Journal of Mathematical Biology, 38(2):103–133, February 1999. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050143.pdf.

Bees:1999:NLB

[1179] M. A. Bees and N. A. Hill. Non-linear bioconvection in a deep suspension of gyrotactic swimming micro-organisms. Journal of Mathematical Biology, 38(2):135–168, February 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050144.pdf.

Grunbaum:1999:ADE

[1180] Daniel Grünbaum. Advection–diffusion equations for generalized tactic searching behaviors. Journal of Mathematical Biology, 38(2):169– 194, February 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050145.pdf.

Mackey:1999:NCG

[1181] Michael C. Mackey and Ryszard Rudnicki. A new criterion for the global stability of simultaneous cell replication and maturation processes. Journal of Mathematical Biology, 38(3):195–219, March 1999. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050146.pdf.

Esteva:1999:MDD

[1182] Lourdes Esteva and Cristobal Vargas. A model for dengue disease with variable human population. Journal of Mathematical Biology, 38(3):220– 240, March 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050147.pdf.

Lasota:1999:CDS

[1183] Andrzej Lasota and Michael C. Mackey. Cell division and the stability of cellular populations. Journal of Mathematical Biology, 38(3):241– 261, March 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050148.pdf.

Friedman:1999:AMM

[1184] Avner Friedman and Fernando Reitich. Analysis of a mathematical model for the growth of tumors. Journal of Mathematical Biology, 38(3): 262–284, March 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050149.pdf.

Wolkowicz:1999:GDC

[1185] Gail S. K. Wolkowicz, Huaxing Xia, and Jianhong Wu. Global dynamics of a chemostat competition model with distributed delay. Journal of Mathematical Biology, 38(4):285–316, April 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850050150.pdf.

Briggs:1999:DFM

[1186] Cheryl J. Briggs, Roger M. Nisbet, and William W. Murdoch. Delayed feedback and multiple attractors in a host-parasitoid system. Journal of Mathematical Biology, 38(4):317–345, April 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850050151.pdf.

Lika:1999:TWS

[1187] Konstadia Lika and Thomas G. Hallam. Traveling wave solutions of a nonlinear reaction–advection equation. Journal of Mathematical Biology, 38(4):346–358, April 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050152.pdf.

Tyson:1999:MAC

[1188] Rebecca Tyson, S. R. Lubkin, and J. D. Murray. Model and analysis of chemotactic bacterial patterns in a liquid medium. Journal of Mathematical Biology, 38(4):359–375, April 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050153.pdf.

Hassell:1999:MAJ

[1189] D. C. Hassell, D. J. Allwright, and A. C. Fowler. A mathematical analysis of Jones’s site model for spruce budworm infestations. Journal of Mathematical Biology, 38(5):377–421, May 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050154.pdf.

Happel:1999:ARC

[1190] Robert Happel and Peter F. Stadler. Autocatalytic replication in a CSTR and constant organization. Journal of Mathematical Biology, 38(5):422– 434, May 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050155.pdf.

Stadler:1999:RFM

[1191] Peter F. Stadler and Robert Happel. Random field models for fitness landscapes. Journal of Mathematical Biology, 38(5):435–478, May 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050156.pdf.

LoFaro:1999:TRN

[1192] Thomas LoFaro and Nancy Kopell. Timing regulation in a network reduced from voltage-gated equations to a one-dimensional map. Journal of Mathematical Biology, 38(6):479–533, June 1999. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050157.pdf.

Mogilner:1999:NLM

[1193] Alexander Mogilner and Leah Edelstein-Keshet. A non-local model for a swarm. Journal of Mathematical Biology, 38(6):534–570, June 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050158.pdf.

LoFaro:1999:AVM

[1194] Thomas LoFaro and Richard Gomulkiewicz. Adaptation versus migration in demographically unstable populations. Journal of Mathematical Biology, 38(6):571–584, June 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050159.pdf.

Epstein:1999:ISS

[1195] Marcelo Epstein and Clifton Johnston. Improved solution for solitary waves in arteries. Journal of Mathematical Biology, 39(1):1– 18, July 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050160.pdf.

Boer:1999:HHO

[1196] M. P. Boer, B. W. Kooi, and S. A. L. M. Kooijman. Homoclinic and heteroclinic orbits to a cycle in a tri-trophic food chain. Journal of Mathematical Biology, 39(1):19–38, July 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050161.pdf.

Tapaswi:1999:EEF

[1197] P. K. Tapaswi and A. Mukhopadhyay. Effects of environmental fluctuation on plankton allelopathy. Journal of Mathematical Biology, 39(1): 39–58, July 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050162.pdf.

Byrne:1999:WNA

[1198] Helen M. Byrne. A weakly nonlinear analysis of a model of avascular solid tumour growth. Journal of Mathematical Biology, 39(1): 59–89, July 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050163.pdf.

Mesterton-Gibbons:1999:SCG

[1199] Michael Mesterton-Gibbons. On sperm competition games: raffles and roles revisited. Journal of Mathematical Biology, 39(2):91–108, August 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050164.pdf.

Regalado:1999:ACM

[1200] Carlos M. Regalado and Brian D. Sleeman. Aggregation and collapse in a mechanical model of fungal tip growth. Journal of Mathematical Biology, 39(2):109–138, August 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050165.pdf.

Othmer:1999:SRC

[1201] Hans G. Othmer and Min Xie. Subharmonic resonance and chaos in forced excitable systems. Journal of Mathematical Biology, 39(2):139– 171, August 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050166.pdf.

Yannacopoulos:1999:EDV

[1202] A. N. Yannacopoulos and G. Rowlands. Effective drift velocities and effective diffusivities of swimming microorganisms in external flows. Journal of Mathematical Biology, 39(2):172–192, August 1999. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050167.pdf.

Wiegmann:1999:SSI

[1203] Daniel D. Wiegmann, Kajal Mukhopadhyay, and Leslie A. Real. Sequential search and the influence of male quality on female mating decisions. Journal of Mathematical Biology, 39(3):193–216, September 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050168.pdf.

Henson:1999:CAS

[1204] Shandelle M. Henson. A continuous, age-structured insect population model. Journal of Mathematical Biology, 39(3):217–243, September 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050169.pdf.

Percus:1999:ILD

[1205] O. E. Percus and J. K. Percus. Island length distribution in genome sequencing. Journal of Mathematical Biology, 39(3):244–268, September 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050170.pdf.

Lu:1999:PGA

[1206] Zhengyi Lu and Wendi Wang. Permanence and global attractivity for Lotka–Volterra difference systems. Journal of Mathematical Biology, 39(3):269–282, September 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050171.pdf.

Yi:1999:NFD

[1207] Tao Yi, R. Cressman, and B. Brooks. Nonlinear frequency-dependent selection at a single locus with two alleles and two phenotypes. Journal of Mathematical Biology, 39(4):283–308, October 1999. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050192.pdf.

Clancy:1999:OIE

[1208] Damian Clancy. Optimal intervention for epidemic models with general infection and removal rate functions. Journal of Mathematical Biology, 39(4):309–331, October 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050193.pdf.

Cooke:1999:IMD

[1209] K. Cooke, P. van den Driessche, and X. Zou. Interaction of maturation delay and nonlinear birth in population and epidemic models. Journal of Mathematical Biology, 39(4):332–352, October 1999. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050194.pdf.

Jackson:1999:TAC

[1210] Trachette L. Jackson, Sharon R. Lubkin, and James D. Murray. Theoretical analysis of conjugate localization in two-step cancer chemotherapy. Journal of Mathematical Biology, 39(4):353–376, October 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850050195.pdf.

Clother:1999:EAS

[1211] D. R. Clother and J. Brindley. Excitability of an age-structured plankton ecosystem. Journal of Mathematical Biology, 39(5):377–420, November 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050172.pdf.

Joyce:1999:NBA

[1212] Paul Joyce. No BLUE among phylogenetic estimators. Journal of Mathematical Biology, 39(5):421–438, November 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850050173.pdf.

King:1999:RBH

[1213] Aaron A. King and William M. Schaffer. The rainbow bridge: Hamiltonian limits and resonance in predator–prey dynamics. Journal of Mathematical Biology, 39(5):439–469, November 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850050174.pdf.

Milner:1999:PSR

[1214] F. A. Milner and A. Pugliese. Periodic solutions: a robust numerical method for an S-I-R model of epidemics. Journal of Mathematical Biology, 39(6):471–492, December 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050175.pdf.

Boukal:1999:LFL

[1215] David S. Boukal and Vlastimil Kĭvan. Lyapunov functions for Lotka– Volterra predator–prey models with optimal foraging behavior. Journal of Mathematical Biology, 39(6):493–517, December 1999. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850050009.pdf.

Barradas:1999:RCV

[1216] Ignacio Barradas and Karla Tassier. Reducing competition vs. improving resistance to disturbances in the environment. Journal of Mathematical Biology, 39(6):518–532, December 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050176.pdf.

Edwards:1999:TES

[1217] David A. Edwards, Byron Goldstein, and Donald S. Cohen. Transport effects on surface-volume biological reactions. Journal of Mathematical Biology, 39(6):533–561, December 1999. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050177.pdf.

Anonymous:2000:E

[1218] Anonymous. Editorial. Journal of Mathematical Biology, 40(1):1, January 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050001.pdf.

Anonymous:2000:ASJ

[1219] Anonymous. Aims and scope of the journal. Journal of Mathematical Biology, 40(1):2, January 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050002.pdf.

Hethcote:2000:TSE

[1220] Herbert W. Hethcote and P. van den Driessche. Two SIS epidemiologic models with delays. Journal of Mathematical Biology, 40(1):3– 26, January 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050003.pdf.

Bees:2000:ITF

[1221] M. A. Bees, P. Andresén, E. Mosekilde, and M. Givskov. The interaction of thin-film flow, bacterial swarming and cell differentiation in colonies of Serratia liquefaciens. Journal of Mathematical Biology, 40(1): 27–63, January 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050004.pdf.

Edelstein-Keshet:2000:MSP

[1222] Leah Edelstein-Keshet and G. Bard Ermentrout. Models for spatial polymerization dynamics of rod-like polymers. Journal of Mathematical Biology, 40(1):64–96, January 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050005.pdf.

Dickinson:2000:GTM

[1223] Richard B. Dickinson. A generalized transport model for biased cell migration in an anisotropic environment. Journal of Mathematical Biology, 40(2):97–135, February 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050006.pdf.

DeGaetano:2000:MMI

[1224] Andrea De Gaetano and Ovide Arino. Mathematical modelling of the intravenous glucose tolerance test. Journal of Mathematical Biology, 40(2):136–168, February 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050007.pdf.

Bressloff:2000:TWP

[1225] Paul C. Bressloff. Traveling waves and pulses in a one-dimensional network of excitable integrate-and-fire neurons. Journal of Mathematical Biology, 40(2):169–198, February 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850050008.pdf.

Lopez:2000:UPS

[1226] L. F. Lopez and F. A. B. Coutinho. On the uniqueness of the positive solution of an integral equation which appears in epidemiological models.

Journal of Mathematical Biology, 40(3):199–228, March 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850050178.pdf.

Jang:2000:DVY

[1227] S. R.-J. Jang. Dynamics of variable-yield nutrient-phytoplankton-zooplankton models with nutrient recycling and self-shading. Journal of Mathematical Biology, 40(3):229–250, March 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850050179.pdf.

Bonneuil:2000:PPT

[1228] Noël Bonneuil and Patrick Saint-Pierre. Protected polymorphism in the two-locus haploid model with unpredictable fitnesses. Journal of Mathematical Biology, 40(3):251–277, March 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050180.pdf.

Percus:2000:RBD

[1229] Jerome K. Percus, Ora E. Percus, and Alan S. Perelson. Random binding of dimers to chains. Journal of Mathematical Biology, 40(3):278– 294, March 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050181.pdf.

Redheffer:2000:GMI

[1230] Ray Redheffer. Generalized monotonicity, integral conditions and partial survival. Journal of Mathematical Biology, 40(4):295–320, April 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050182.pdf.

Feltham:2000:ASM

[1231] D. L. Feltham and M. A. J. Chaplain. Analytical solutions of a minimal model of species migration in a bounded domain. Journal of Mathematical Biology, 40(4):321–342, April 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850050183.pdf.

Herbert:2000:SHP

[1232] Julian Herbert and Valerie Isham. Stochastic host-parasite interaction models. Journal of Mathematical Biology, 40(4):343–371, April 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850050184.pdf.

Ratnanather:2000:AHC

[1233] J. Tilak Ratnanather, Aleksander S. Popel, and William E. Brownell. An analysis of the hydraulic conductivity of the extracisternal space of the cochlear outer hair cell. Journal of Mathematical Biology, 40(4): 372–382, April 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850050185.pdf.

Alvarez:2000:OIR

[1234] Luis H. R. Alvarez. On the option interpretation of rational harvesting planning. Journal of Mathematical Biology, 40(5):383–405, May 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002859900020.pdf.

Broom:2000:PES

[1235] M. Broom. Patterns of evolutionarily stable strategies: the maximal pattern conjecture revisited. Journal of Mathematical Biology, 40(5): 406–412, May 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000023.pdf.

Behncke:2000:PC

[1236] Horst Behncke. Periodical cicadas. Journal of Mathematical Biology, 40 (5):413–431, May 2000. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850000024.pdf.

Alvarez:2000:CSP

[1237] Luis H. R. Alvarez. On the comparative static properties of the expected population density in the presence of stochastic fluctuations. Journal of Mathematical Biology, 40(5):432–442, May 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850000028.pdf.

Schneider:2000:MRE

[1238] Klaus R. Schneider and Thomas Wilhelm. Model reduction by extended quasi-steady-state approximation. Journal of Mathematical Biology, 40 (5):443–450, May 2000. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850000026.pdf.

Calsina:2000:MMP

[1239] Angel Calsina, José M. Mazón, and Manuel Serra. A mathematical model for the phase of sexual reproduction in monogonont rotifers. Journal of Mathematical Biology, 40(5):451–471, May 2000. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850000021.pdf.

Jones:2000:MMS

[1240] A. F. Jones, H. M. Byrne, J. S. Gibson, and J. W. Dold. A mathematical model of the stress induced during avascular tumour growth. Journal of Mathematical Biology, 40(6):473–499, June 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850000033.pdf.

Schley:2000:LSC

[1241] D. Schley and S. A. Gourley. Linear stability criteria for population models with periodically perturbed delays. Journal of Mathematical Biology, 40(6):500–524, June 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000034.pdf.

vandenDriessche:2000:SSE

[1242] P. van den Driessche and James Watmough. A simple SIS epidemic model with a backward bifurcation. Journal of Mathematical Biology, 40(6): 525–540, June 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000032.pdf.

Rocheleau:2000:SAP

[1243] Ghislain Rocheleau and Sabin Lessard. Stability analysis of the partial selfing selection model. Journal of Mathematical Biology, 40(6): 541–574, June 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000030.pdf.

Anonymous:2000:LDN

[1244] Anonymous. Life and death near a windy oasis. Journal of Mathematical Biology, 41(1):1–23, July 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000025.pdf.

Sikder:2000:UPS

[1245] Asim Sikder. Uniform persistence for sigmoidal diet selection with keystone prey species. Journal of Mathematical Biology, 41(1):25– 44, July 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000027.pdf.

Batzel:2000:SHRa

[1246] J. J. Batzel and H. T. Tran. Stability of the human respiratory control system I. Analysis of a two-dimensional delay state-space model. Journal of Mathematical Biology, 41(1):45–79, July 2000. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850000044.pdf.

Batzel:2000:SHRb

[1247] J. J. Batzel and H. T. Tran. Stability of the human respiratory control system II. Analysis of a three-dimensional delay state-space model. Journal of Mathematical Biology, 41(1):80–102, July 2000. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850000045.pdf.

Neubert:2000:DDV

[1248] Michael G. Neubert and Hal Caswell. Density-dependent vital rates and their population dynamic consequences. Journal of Mathematical Biology, 41(2):103–121, August 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850070001.pdf.

Nagylaki:2000:GIS

[1249] Thomas Nagylaki. Geographical invariance and the strong-migration limit in subdivided populations. Journal of Mathematical Biology, 41(2): 123–142, August 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850070002.pdf.

Muller:2000:RV

[1250] Johannes Müller, Birgitt Schönfisch, and Markus Kirkilionis. Ring vaccination. Journal of Mathematical Biology, 41(2):143–171, August 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850070003.pdf.

Trapa:2000:NEE

[1251] Peter E. Trapa and Martin A. Nowak. Nash equilibria for an evolutionary language game. Journal of Mathematical Biology, 41(2):172– 188, August 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850070004.pdf.

Elston:2000:MDB

[1252] Timothy C. Elston. A macroscopic description of biomolecular transport. Journal of Mathematical Biology, 41(3):189–206, September 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850000043.pdf.

Aita:2000:AWF

[1253] Takuyo Aita and Yuzuru Husimi. Adaptive walks by the fittest among finite random mutants on a Mt. Fuji-type fitness landscape II. Effect of small non-additivity. Journal of Mathematical Biology, 41(3):207– 231, September 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000046.pdf.

Jansen:2000:LSA

[1254] Vincent A. A. Jansen and Alun L. Lloyd. Local stability analysis of spatially homogeneous solutions of multi-patch systems. Journal of Mathematical Biology, 41(3):232–252, September 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850000048.pdf.

Watkins:2000:CFT

[1255] Joseph C. Watkins. Consistency and fluctuation theorems for discrete time structured population models having demographic stochasticity. Journal of Mathematical Biology, 41(3):253–271, September 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850000041.pdf.

Gourley:2000:TFS

[1256] S. A. Gourley. Travelling front solutions of a nonlocal Fisher equation. Journal of Mathematical Biology, 41(3):272–284, September 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850000047.pdf.

Painter:2000:DAM

[1257] K. J. Painter, P. K. Maini, and H. G. Othmer. Development and applications of a model for cellular response to multiple chemotactic cues. Journal of Mathematical Biology, 41(4):285–314, October 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850000035.pdf.

Wagle:2000:SCC

[1258] Mihir A. Wagle and Robert T. Tranquillo. A self-consistent cell flux expression for simultaneous chemotaxis and contact guidance in tissues. Journal of Mathematical Biology, 41(4):315–330, October 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850000040.pdf.

Qian:2000:MAB

[1259] Hong Qian. A mathematical analysis for the Brownian dynamics of a DNA tether. Journal of Mathematical Biology, 41(4):331–340, October 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000055.pdf.

Cornell:2000:DRP

[1260] Stephen J. Cornell, Valerie S. Isham, and Bryan T. Grenfell. Drugresistant parasites and aggregated infection- early-season dynamics. Journal of Mathematical Biology, 41(4):341–360, October 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850000051.pdf.

Lika:2000:DEB

[1261] Konstadia Lika and Roger M. Nisbet. A dynamic energy budget model based on partitioning of net production. Journal of Mathematical Biology, 41(4):361–386, October 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000049.pdf.

Lewis:2000:MAS

[1262] M. A. Lewis and S. Pacala. Modeling and analysis of stochastic invasion processes. Journal of Mathematical Biology, 41(5):387–429, November 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000050.pdf.

Lewis:2000:SRN

[1263] M. A. Lewis. Spread rate for a nonlinear stochastic invasion. Journal of Mathematical Biology, 41(5):430–454, November 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850000022.pdf.

Tyson:2000:FSM

[1264] Rebecca Tyson, L. G. Stern, and Randall J. LeVeque. Fractional step methods applied to a chemotaxis model. Journal of Mathematical Biology, 41(5):455–475, November 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000038.pdf.

Krzyzanski:2000:ATN

[1265] Wojciech Krzyzanski. Asymptotics of the total net direct pharmacological effect for large drug doses. Journal of Mathematical Biology, 41(6):477–492, December 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000052.pdf.

Satnoianu:2000:TIG

[1266] Razvan A. Satnoianu, Michael Menzinger, and Philip K. Maini. Turing instabilities in general systems. Journal of Mathematical Biology, 41(6):493–512, December 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000056.pdf.

Rubin:2000:ACF

[1267] Jonathan Rubin and David Terman. Analysis of clustered firing patterns in synaptically coupled networks of oscillators. Journal of Mathematical Biology, 41(6):513–545, December 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000065.pdf.

Vielle:2000:NES

[1268] Bruno Vielle. A new explicit stability criterion for human periodic breathing. Journal of Mathematical Biology, 41(6):546–558, December 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000061.pdf.

Andersson:2000:SED

[1269] Håkan Andersson and Tom Britton. Stochastic epidemics in dynamic populations: quasi-stationarity and extinction. Journal of Mathematical Biology, 41(6):559–580, December 2000. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000060.pdf.

DiCrescenzo:2001:SPM

[1270] A. Di Crescenzo, V. Giorno, A. G. Nobile, and L. M. Ricciardi. Stochastic population models with interacting species. Journal of Mathematical Biology, 42(1):1–25, January 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/PL00000070.pdf.

Khanin:2001:PME

[1271] Kostya Khanin and Raya Khanin. A probabilistic model for the establishment of neuron polarity. Journal of Mathematical Biology, 42(1): 26–40, January 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/PL00000071.pdf.

Slade:2001:SHH

[1272] Paul F. Slade. Simulation of ‘hitch-hiking’ genealogies. Journal of Mathematical Biology, 42(1):41–70, January 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/PL00000072.pdf.

Ai:2001:PSM

[1273] Shangbing Ai. Periodic solutions in a model of competition between plasmid-bearing and plasmid-free organisms in a chemostat with an inhibitor. Journal of Mathematical Biology, 42(1):71–94, January 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/content/pdf/10.1007/PL00000073.pdf.

Cantrell:2001:BBB

[1274] Robert Stephen Cantrell, Chris Cosner, and William F. Faganantrell. Brucellosis, botflies, and brainworms: the impact of edge habitats on pathogen transmission and species extinction. Journal of Mathematical Biology, 42(2):95–119, February 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000064.pdf.

Chiu:2001:MMS

[1275] Chichia Chiu and Frank C. Hoppensteadt. Mathematical models and simulations of bacterial growth and chemotaxis in a diffusion gradient chamber. Journal of Mathematical Biology, 42(2):120–144, February 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000069.pdf.

Angerer:2001:ERL

[1276] Wolfgang P. Angerer. An explicit representation of the Luria– Delbrück distribution. Journal of Mathematical Biology, 42(2):145–174, February 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000053.pdf.

Switkes:2001:CIB

[1277] Jennifer M. Switkes and Michael E. Moody. Coevolutionary interactions between a haploid species and a diploid species. Journal of Mathematical Biology, 42(2):175–194, February 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100071.pdf.

Levine:2001:MMO

[1278] Howard A. Levine, Brian D. Sleeman, and Marit Nilsen-Hamilton. Mathematical modeling of the onset of capillary formation initiating angiogenesis. Journal of Mathematical Biology, 42(3):195–238, March 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000037.pdf.

Schreiber:2001:CPD

[1279] Sebastian J. Schreiber. Chaos and population disappearances in simple ecological models. Journal of Mathematical Biology, 42(3):239– 260, March 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000070.pdf.

Mulder:2001:DMP

[1280] B. M. Mulder and A. M. C. Emons. A dynamical model for plant cell wall architecture formation. Journal of Mathematical Biology, 42(3): 261–289, March 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000063.pdf.

Buono:2001:MCPa

[1281] Pietro-Luciano Buono and Martin Golubitsky. Models of central pattern generators for quadruped locomotion I. Primary gaits. Journal of Mathematical Biology, 42(4):291–326, April 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850000058.pdf.

Buono:2001:MCPb

[1282] Pietro-Luciano Buono. Models of central pattern generators for quadruped locomotion II. Secondary gaits. Journal of Mathematical Biology, 42(4):327–346, April 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000073.pdf.

Roose:2001:MMP

[1283] T. Roose, A. C. Fowler, and P. R. Darrah. A mathematical model of plant nutrient uptake. Journal of Mathematical Biology, 42(4):347– 360, April 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000075.pdf.

Engelborghs:2001:NBA

[1284] K. Engelborghs, V. Lemaire, J. Bélair, and D. Roose. Numerical bifurcation analysis of delay differential equations arising from physiological modeling. Journal of Mathematical Biology, 42(4):361– 385, April 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000072.pdf.

Chaplain:2001:STP

[1285] M. A. J. Chaplain, M. Ganesh, and I. G. Graham. Spatio-temporal pattern formation on spherical surfaces: numerical simulation and application to solid tumour growth. Journal of Mathematical Biology, 42(5): 387–423, May 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000067.pdf.

ElHoussif:2001:DSS

[1286] N. E. El Houssif. A deterministic size-structured population model for the worm Naidis elinguis. Journal of Mathematical Biology, 42(5): 424–438, May 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850000059.pdf.

Teng:2001:EDS

[1287] Zhidong Teng and Zhengyi Lu. The effect of dispersal on single-species nonautonomous dispersal models with delays. Journal of Mathematical Biology, 42(5):439–454, May 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000076.pdf.

Baake:2001:MRT

[1288] Ellen Baake. Mutation and recombination with tight linkage. Journal of Mathematical Biology, 42(5):455–488, May 2001. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850000077.pdf.

Hsu:2001:GAM

[1289] Sze-Bi Hsu, Tzy-Wei Hwang, and Yang Kuang. Global analysis of the Michaelis–Menten-type ratio-dependent predator–prey system. Journal of Mathematical Biology, 42(6):489–506, June 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850100079.pdf.

Zacher:2001:PSA

[1290] Rico Zacher. Persistent solutions for age-dependent pair-formation models. Journal of Mathematical Biology, 42(6):507–531, June 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850100084.pdf.

Luchsinger:2001:SMP

[1291] C. J. Luchsinger. Stochastic models of a parasitic infection, exhibiting three basic reproduction ratios. Journal of Mathematical Biology, 42(6): 532–554, June 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100082.pdf.

Luchsinger:2001:ALT

[1292] C. J. Luchsinger. Approximating the long-term behaviour of a model for parasitic infection. Journal of Mathematical Biology, 42(6):555– 581, June 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100083.pdf.

Lemire:2001:ESS

[1293] Mathieu Lemire and Sabin Lessard. Equilibrium structure and stability in a frequency-dependent, two-population diploid model. Journal of Mathematical Biology, 43(1):1–21, July 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850100080.pdf.

Boucher:2001:ASM

[1294] Kenneth Boucher, Alexander Zorin, Andrej Y. Yakovlev, Margot Mayer-Proschel, and Mark Noble. An alternative stochastic model of generation of oligodendrocytes in cell culture. Journal of Mathematical Biology, 43 (1):22–36, July 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100085.pdf.

So:2001:SPT

[1295] Joseph W.-H. So, Jianhong Wu, and Xingfu Zou. Structured population on two patches: modeling dispersal and delay. Journal of Mathematical Biology, 43(1):37–51, July 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100081.pdf.

Feng:2001:IDT

[1296] Zhilan Feng, Jason Curtis, and Dennis J. Minchella. The influence of drug treatment on the maintenance of schistosome genetic diversity. Journal of Mathematical Biology, 43(1):52–68, July 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850100092.pdf.

Hsieh:2001:EDD

[1297] Ying-Hen Hsieh and Shin-Pyng Sheu. The effect of density-dependent treatment and behavior change on the dynamics of HIV transmission. Journal of Mathematical Biology, 43(1):69–80, July 2001. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850100087.pdf.

Curtu:2001:ORN

[1298] Rodica Curtu and Bard Ermentrout. Oscillations in a refractory neural net. Journal of Mathematical Biology, 43(1):81–100, July 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850100089.pdf.

Iwasaki:2001:IBP

[1299] Katsunori Iwasaki and Yutaka Kamimura. Inverse bifurcation problem, singular Wiener–Hopf equations, and mathematical models in ecology. Journal of Mathematical Biology, 43(2):101–143, August 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850100096.pdf.

Dreyer:2001:ASA

[1300] Olaf Dreyer and Raymond Puzio. Allometric scaling in animals and plants. Journal of Mathematical Biology, 43(2):144–156, August 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850170001.pdf.

Diekmann:2001:FAG

[1301] O. Diekmann, M. Gyllenberg, H. Huang, M. Kirkilionis, J. A. J. Metz, and H. R. Thieme. On the formulation and analysis of general deterministic structured population models II. Nonlinear theory. Journal of Mathematical Biology, 43(2):157–189, August 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850170002.pdf.

Chen:2001:IGI

[1302] C. Y. Chen, H. M. Byrne, and J. R. King. The influence of growthinduced stress from the surrounding medium on the development of multicell spheroids. Journal of Mathematical Biology, 43(3):191–220, September 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100091.pdf.

Berezovskaya:2001:PAR

[1303] F. Berezovskaya, G. Karev, and R. Arditi. Parametric analysis of the ratio-dependent predator–prey model. Journal of Mathematical Biology, 43(3):221–246, September 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850000078.pdf.

Martin:2001:PPM

[1304] Annik Martin and Shigui Ruan. predator–prey models with delay and prey harvesting. Journal of Mathematical Biology, 43(3):247–267, September 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100095.pdf.

Xiao:2001:GDR

[1305] Dongmei Xiao and Shigui Ruan. Global dynamics of a ratio-dependent predator–prey system. Journal of Mathematical Biology, 43(3):268–290, September 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100097.pdf.

Sherratt:2001:NMM

[1306] Jonathan A. Sherratt and Mark A. J. Chaplain. A new mathematical model for avascular tumour growth. Journal of Mathematical Biology, 43(4):291–312, October 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100088.pdf.

Pradines:2001:CLP

[1307] Joël R. Pradines, Jeff Hasty, and Khashayar Pakdaman. Complex ligand-protein systems: a globally convergent iterative method for the n× m case. Journal of Mathematical Biology, 43(4):313–324, October 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100086.pdf.

Edelstein-Keshet:2001:MAF

[1308] Leah Edelstein-Keshet and G. Bard Ermentrout. A model for actinfilament length distribution in a lamellipod. Journal of Mathematical Biology, 43(4):325–355, October 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100102.pdf.

Ponomarev:2001:RBD

[1309] A. L. Ponomarev and R. K. Sachs. Radiation breakage of DNA: a model based on random-walk chromatin structure. Journal of Mathematical Biology, 43(4):356–376, October 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100098.pdf.

Hsu:2001:RDR

[1310] Sze-Bi Hsu, Tzy-Wei Hwang, and Yang Kuang. Rich dynamics of a ratio-dependent one-prey two-predators model. Journal of Mathematical Biology, 43(5):377–396, November 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100100.pdf.

Bahlo:2001:CTT

[1311] Melanie Bahlo and Robert C. Griffiths. Coalescence time for two genes from a subdivided population. Journal of Mathematical Biology, 43(5):

397–410, November 2001. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850100104.pdf.

Plahte:2001:PFD

[1312] Erik Plahte. Pattern formation in discrete cell lattices. Journal of Mathematical Biology, 43(5):411–445, November 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850100105.pdf.

Mohtashemi:2001:TDE

[1313] Mojdeh Mohtashemi and Richard Levins. Transient dynamics and early diagnostics in infectious disease. Journal of Mathematical Biology, 43(5): 446–470, November 2001. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850100103.pdf.

Wikan:2001:CCA

[1314] Arild Wikan. From chaos to chaos. An analysis of a discrete agestructured prey-predator model. Journal of Mathematical Biology, 43(6): 471–500, December 2001. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850100101.pdf.

Hutson:2001:EDR

[1315] V. Hutson, K. Mischaikow, and P. Polácik. The evolution of dispersal rates in a heterogeneous time-periodic environment. Journal of Mathematical Biology, 43(6):501–533, December 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850100106.pdf.

Alvarez:2001:DIS

[1316] Luis H. R. Alvarez. Does increased stochasticity speed up extinction? Journal of Mathematical Biology, 43(6):534–544, December 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850100108.pdf.

Gyllenberg:2001:FSM

[1317] Mats Gyllenberg and J. A. J. Metz. On fitness in structured metapopulations. Journal of Mathematical Biology, 43(6):545–560, December 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100113.pdf.

Huang:2001:PMR

[1318] Yunxin Huang and Odo Diekmann. Predator migration in response to prey density: What are the consequences? Journal of Mathematical Biology, 43(6):561–581, December 2001. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100107.pdf.

Stacey:2002:MCO

[1319] Mark T. Stacey, Kristina S. Mead, and Mimi A. R. Koehl. Molecule capture by olfactory antennules: Mantis shrimp. Journal of Mathematical Biology, 44(1):1–30, January 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100111.pdf.

Billings:2002:ECN

[1320] L. Billings and I. B. Schwartz. Exciting chaos with noise: unexpected dynamics in epidemic outbreaks. Journal of Mathematical Biology, 44(1): 31–48, January 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100110.pdf.

Gourley:2002:DFL

[1321] S. A. Gourley and J. W.-H. So. Dynamics of a food-limited population model incorporating nonlocal delays on a finite domain. Journal of Mathematical Biology, 44(1):49–78, January 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850100109.pdf.

Agur:2002:UPS

[1322] Zvia Agur, Yoaz Daniel, and Yuval Ginosar. The universal properties of stem cells as pinpointed by a simple discrete model. Journal of Mathematical Biology, 44(1):79–86, January 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850100115.pdf.

Tonnelier:2002:WPD

[1323] Arnaud Tonnelier. Wave propagation in discrete media. Journal of Mathematical Biology, 44(1):87–105, January 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850100114.pdf.

Crampin:2002:MDT

[1324] E. J. Crampin, E. A. Gaffney, and P. K. Maini. Mode-doubling and tripling in reaction–diffusion patterns on growing domains: A piecewise linear model. Journal of Mathematical Biology, 44(2):107–128, February 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100112.pdf.

Jang:2002:DHP

[1325] S. R.-J. Jang. Dynamics of herbivore-plant-pollinator models. Journal of Mathematical Biology, 44(2):129–149, February 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850100117.pdf.

Wang:2002:IEA

[1326] Mei-Hui Wang, Mark Kot, and Michael G. Neubert. Integrodifference equations, Allee effects, and invasions. Journal of Mathematical Biology, 44(2):150–168, February 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100116.pdf.

Gog:2002:SBA

[1327] J. R. Gog and J. Swinton. A status-based approach to multiple strain dynamics. Journal of Mathematical Biology, 44(2):169–184, February 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100120.pdf.

Tang:2002:DDB

[1328] Sanyi Tang and Lansun Chen. Density-dependent birth rate, birth pulses and their population dynamic consequences. Journal of Mathematical Biology, 44(2):185–199, February 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100121.pdf.

Jackson:2002:VTG

[1329] Trachette L. Jackson. Vascular tumor growth and treatment: Consequences of polyclonality, competition and dynamic vascular support. Journal of Mathematical Biology, 44(3):201–226, March 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s002850100118.pdf.

OCallaghan:2002:TDM

[1330] M. O’Callaghan and A. G. Murray. A tractable deterministic model with realistic latent period for an epidemic in a linear habitat. Journal of Mathematical Biology, 44(3):227–251, March 2002. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850100119.pdf.

Nagylaki:2002:WWW

[1331] Thomas Nagylaki. When and where was the most recent common ancestor? Journal of Mathematical Biology, 44(3):253–275, March 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100122.pdf.

vandenBerg:2002:OAB

[1332] Hugo A. van den Berg, Yuri N. Kiselev, and Michael V. Orlov. Optimal allocation of building blocks between nutrient uptake systems in a microbe. Journal of Mathematical Biology, 44(3):276–296, March 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100123.pdf.

Sikder:2002:LVC

[1333] Asim Sikder. A Lotka–Volterra competition model and its global convergence to a definite axial equilibrium. Journal of Mathematical Biology, 44(4):297–308, April 2002. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850100125.pdf.

Kazmierczak:2002:HSM

[1334] B. Kazmierczak and T. Lipniacki. Homoclinic solutions in mechanical systems with small dissipation. Application to DNA dynamics. Journal of Mathematical Biology, 44(4):309–329, April 2002. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850100131.pdf.

Torres:2002:RIE

[1335] M. Torres, J. L. Aragón, P. Domínguez, and D. Gil. Regularity in irregular echinoids. Journal of Mathematical Biology, 44(4):330– 340, April 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100126.pdf.

Abundo:2002:APE

[1336] Mario Abundo, Luigi Accardi, Nicola Rosato, and Lorenzo Stella. Analysing protein energy data by a stochastic model for cooperative interactions: comparison and characterization of cooperativity. Journal of Mathematical Biology, 44(4):341–359, April 2002. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850100133.pdf.

Krenz:2002:VDF

[1337] Gary S. Krenz and Christopher A. Dawson. Vessel distensibility and flow distribution in vascular trees. Journal of Mathematical Biology, 44(4): 360–374, April 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100127.pdf.

Hwang:2002:FVL

[1338] W. Hwang, P. L. Krapivsky, and S. Redner. Fitness versus longevity in age-structured population dynamics. Journal of Mathematical Biology, 44(4):375–393, April 2002. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850100128.pdf.

Cui:2002:AMM

[1339] Shangbin Cui. Analysis of a mathematical model for the growth of tumors under the action of external inhibitors. Journal of Mathematical Biology, 44(5):395–426, May 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100130.pdf.

Slade:2002:CAT

[1340] Paul F. Slade. Conditional ancestral times and related properties of an extant mutant at a locus A given the number of segregating sites observed at a linked locus B. Journal of Mathematical Biology, 44(5): 427–449, May 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100137.pdf.

Li:2002:APF

[1341] Chi-Kwong Li and Hans Schneider. Applications of Perron–Frobenius theory to population dynamics. Journal of Mathematical Biology, 44(5): 450–462, May 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100132.pdf.

Horstmann:2002:ERS

[1342] Dirk Horstmann. On the existence of radially symmetric blow-up solutions for the Keller–Segel model. Journal of Mathematical Biology, 44(5): 463–478, May 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850100134.pdf.

Chiu:2002:NAF

[1343] Chichia Chiu and Charles S. Peskin. Numerical analysis of the flux and force generated by a correlation ratchet mechanism. Journal of Mathematical Biology, 44(5):479–501, May 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850100129.pdf.

Whiteley:2002:MMO

[1344] Jonathan P. Whiteley, David J. Gavaghan, and Clive E. W. Hahn. Mathematical modelling of oxygen transport to tissue. Journal of Mathematical Biology, 44(6):503–522, June 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850200135.pdf.

Bonneuil:2002:MNG

[1345] Noël Bonneuil and Patrick Saint-Pierre. Minimal number of generations out of polymorphism in the one-locus two-allele model with unpredictable fertilities. Journal of Mathematical Biology, 44(6):523– 547, June 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850200139.pdf.

Geritz:2002:IDA

[1346] S. A. H. Geritz, M. Gyllenberg, F. J. A. Jacobs, and K. Parvinen. Invasion dynamics and attractor inheritance. Journal of Mathematical Biology, 44(6):548–560, June 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850100136.pdf.

Greenhalgh:2002:GMA

[1347] David Greenhalgh and Fraser Lewis. The general mixing of addicts and needles in a variable-infectivity needle-sharing environment. Journal of Mathematical Biology, 44(6):561–598, June 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s002850200140.pdf.

Eshel:2002:LTG

[1348] Ilan Eshel, Emilia Sansone, and Frans Jacobs. A long-term genetic model for the evolution of sexual preference: the theories of Fisher and zahavi re-examined. Journal of Mathematical Biology, 45(1):1– 21, July 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850200138.pdf.

Calsina:2002:HBS

[1349] Àngel Calsina and Jordi Ripoll. Hopf bifurcation in a structured population model for the sexual phase of monogonont rotifers. Journal of Mathematical Biology, 45(1):22–36, July 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850200147.pdf.

Biscari:2002:CEM

[1350] Paolo Biscari, Fulvio Bisi, and Riccardo Rosso. Curvature effects on membrane-mediated interactions of inclusions. Journal of Mathematical Biology, 45(1):37–56, July 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850200142.pdf.

Chapman:2002:SES

[1351] Clare L. Chapman, James J. Wright, and Paul D. Bourke. Spatial eigenmodes and synchronous oscillation: Co-incidence detection in simulated cerebral cortex. Journal of Mathematical Biology, 45(1):57– 78, July 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850200141.pdf.

Gyllenberg:2002:ESE

[1352] Mats Gyllenberg, Kalle Parvinen, and Ulf Dieckmann. Evolutionary suicide and evolution of dispersal in structured metapopulations. Journal of Mathematical Biology, 45(2):79–105, August 2002. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850200151.pdf.

Parvinen:2002:EBD

[1353] Kalle Parvinen. Evolutionary branching of dispersal strategies in structured metapopulations. Journal of Mathematical Biology, 45(2):106– 124, August 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850200150.pdf.

Breward:2002:RCC

[1354] C. J. W. Breward, H. M. Byrne, and C. E. Lewis. The role of cell-cell interactions in a two-phase model for avascular tumour growth. Journal of Mathematical Biology, 45(2):125–152, August 2002. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850200149.pdf.

Schuster:2002:RRB

[1355] S. Schuster, C. Hilgetag, J. H. Woods, and D. A. Fell. Reaction routes in biochemical reaction systems: Algebraic properties, validated calculation procedure and example from nucleotide metabolism. Journal of Mathematical Biology, 45(2):153–181, August 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850200143.pdf.

Weinberger:2002:ALD

[1356] Hans F. Weinberger, Mark A. Lewis, and Bingtuan Li. Analysis of linear determinacy for spread in cooperative models. Journal of Mathematical Biology, 45(3):183–218, September 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850200145.pdf.

Lewis:2002:SSL

[1357] Mark A. Lewis, Bingtuan Li, and Hans F. Weinberger. Spreading speed and linear determinacy for two-species competition models. Journal of Mathematical Biology, 45(3):219–233, September 2002. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850200144.pdf.

Lutscher:2002:MAM

[1358] Frithjof Lutscher. Modeling alignment and movement of animals and cells. Journal of Mathematical Biology, 45(3):234–260, September 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850200146.pdf.

Grassi:2002:DHR

[1359] Gabriele Grassi, Mario Grassi, Anne Kuhn, and Reinhard Kandolf. Determination of hammerhead ribozyme kinetic constants at high molar ratio ribozyme-substrate. Journal of Mathematical Biology, 45(3):261– 277, September 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850200152.pdf.

Olofsson:2002:ESF

[1360] Peter Olofsson and Chad A. Shaw. Exact sampling formulas for multitype Galton–Watson processes. Journal of Mathematical Biology, 45(4): 279–293, October 2002. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s002850200148.pdf.

Al-Omari:2002:MTF

[1361] J. Al-Omari and S. A. Gourley. Monotone travelling fronts in an agestructured reaction–diffusion model of a single species. Journal of Mathematical Biology, 45(4):294–312, October 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850200159.pdf.

Arino:2002:DSS

[1362] Julien Arino, Jean-Luc Gouzé, and Antoine Sciandra. A discrete, sizestructured model of phytoplankton growth in the chemostat. Journal of Mathematical Biology, 45(4):313–336, October 2002. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s002850200160.pdf.

Gaffney:2002:ISM

[1363] E. A. Gaffney, K. Pugh, P. K. Maini, and F. Arnold. Investigating a simple model of cutaneous wound healing angiogenesis. Journal of Mathematical Biology, 45(4):337–374, October 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850200161.pdf.

Bauch:2002:VOA

[1364] C. T. Bauch. A versatile ODE approximation to a network model for the spread of sexually transmitted diseases. Journal of Mathematical Biology, 45(5):375–395, November 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850200153.pdf.

Lindstrom:2002:DDF

[1365] Torsten Lindström. On the dynamics of discrete food chains: Low- and high-frequency behavior and optimality of chaos. Journal of Mathematical Biology, 45(5):396–418, November 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s002850200154.pdf.

Arrigoni:2002:LMP

[1366] F. Arrigoni and A. Pugliese. Limits of a multi-patch SIS epidemic model. Journal of Mathematical Biology, 45(5):419–440, November 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s002850200155.pdf.

Haslach:2002:NDM

[1367] Henry W. Haslach. A nonlinear dynamical mechanism for bruit generation by an intracranial saccular aneurysm. Journal of Mathematical Biology, 45(5):441–460, November 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850200156.pdf.

Stockmarr:2002:DPP

[1368] Anders Stockmarr. The distribution of particles in the plane dispersed by a simple 3-dimensional diffusion process. Journal of Mathematical Biology, 45(5):461–469, November 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s002850200157.pdf.

Cooke:2002:IMD

[1369] K. Cooke, P. van den Driessche, and X. Zou. Interaction of maturation delay and nonlinear birth in population and epidemic models. Journal of Mathematical Biology, 45(5):470, November 2002. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-002-0162-x.pdf.

Dawes:2002:OOD

[1370] J. H. P. Dawes and J. R. Gog. The onset of oscillatory dynamics in models of multiple disease strains. Journal of Mathematical Biology, 45(6):471– 510, December 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0163-9.pdf.

Weinberger:2002:SST

[1371] Hans F. Weinberger. On spreading speeds and traveling waves for growth and migration models in a periodic habitat. Journal of Mathematical Biology, 45(6):511–548, December 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-002-0169-3.pdf.

Layton:2002:NMR

[1372] Anita T. Layton and Harold E. Layton. A numerical method for renal models that represent tubules with abrupt changes in membrane properties. Journal of Mathematical Biology, 45(6):549–567, December 2002. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0166-6.pdf.

Whitaker:2003:DBB

[1373] Shree Y. Whitaker, Hien T. Tran, and Christopher J. Portier. Development of a biologically-based controlled growth and differentiation model for developmental toxicology. Journal of Mathematical Biology, 46(1): 1–16, January 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0164-8.pdf.

Hwang:2003:DEE

[1374] Tzy-Wei Hwang and Yang Kuang. Deterministic extinction effect of parasites on host populations. Journal of Mathematical Biology, 46(1): 17–30, January 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0165-7.pdf.

Esteva:2003:CDS

[1375] Lourdes Esteva and Cristobal Vargas. Coexistence of different serotypes of dengue virus. Journal of Mathematical Biology, 46(1):31–47, January 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0168-4.pdf.

DeLeenheer:2003:FCC

[1376] Patrick De Leenheer and Hal Smith. Feedback control for chemostat models. Journal of Mathematical Biology, 46(1):48–70, January 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-002-0170-x.pdf.

Lessard:2003:CFR

[1377] Sabin Lessard and Ghislain Rocheleau. Change in frequency of a rare mutant allele: A general formula and applications to partial inbreeding models. Journal of Mathematical Biology, 46(1):71–94, January 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s002850200158.pdf.

Davydova:2003:YCC

[1378] N. V. Davydova, O. Diekmann, and S. A. van Gils. Year class coexistence or competitive exclusion for strict biennials? Journal of Mathematical Biology, 46(2):95–131, February 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-002-0167-5.pdf.

Huang:2003:ETW

[1379] Jianhua Huang, Gang Lu, and Shigui Ruan. Existence of traveling wave solutions in a diffusive predator–prey model. Journal of Mathematical Biology, 46(2):132–152, February 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-002-0171-9.pdf.

Dolak:2003:CMCa

[1380] Y. Dolak and T. Hillen. Cattaneo models for chemosensitive movement. Journal of Mathematical Biology, 46(2):153–170, February 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-002-0173-7.pdf.

Schliekelman:2003:TDM

[1381] Paul Schliekelman. Transient dynamics in multilocus invasions by transgenic organisms. Journal of Mathematical Biology, 46(2):171–188, February 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0176-4.pdf.

Gaffney:2003:ISM

[1382] E. A. Gaffney, K. Pugh, P. K. Maini, and F. Arnold. Investigating a simple model of cutaneous wound healing angiogenesis. Journal of Mathematical Biology, 46(2):189, February 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-002-0182-6.pdf.

Weinberger:2003:SST

[1383] Hans F. Weinberger. On spreading speeds and traveling waves for growth and migration models in a periodic habitat. Journal of Mathematical Biology, 46(2):190, February 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-002-0189-z.pdf.

Cristini:2003:NST

[1384] Vittorio Cristini, John Lowengrub, and Qing Nie. Nonlinear simulation of tumor growth. Journal of Mathematical Biology, 46(3):191– 224, March 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0174-6.pdf.

Waxman:2003:NES

[1385] D. Waxman. Numerical and exact solutions for continuum of alleles models. Journal of Mathematical Biology, 46(3):225–240, March 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-002-0178-2.pdf.

Wiuf:2003:IPH

[1386] Carsten Wiuf. Inferring population history from genealogical trees. Journal of Mathematical Biology, 46(3):241–264, March 2003. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-002-0180-8.pdf.

Mitchener:2003:BAF

[1387] W. Garrett Mitchener. Bifurcation analysis of the fully symmetric language dynamical equation. Journal of Mathematical Biology, 46(3):265– 285, March 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0172-8.pdf.

Salathe:2003:MAO

[1388] Eric P. Salathe. Mathematical analysis of oxygen concentration in a two dimensional array of capillaries. Journal of Mathematical Biology, 46(4): 287–308, April 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0175-5.pdf.

Ottesen:2003:VPF

[1389] J. T. Ottesen. Valveless pumping in a fluid-filled closed elastic tubesystem: one-dimensional theory with experimental validation. Journal of Mathematical Biology, 46(4):309–332, April 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-002-0179-1.pdf.

Ghosal:2003:SMI

[1390] Sandip Ghosal and Shreyas Mandre. A simple model illustrating the role of turbulence on phytoplankton blooms. Journal of Mathematical Biology, 46(4):333–346, April 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-002-0184-4.pdf.

Lakin:2003:WBM

[1391] William D. Lakin, Scott A. Stevens, Bruce I. Tranmer, and Paul L. Penar. A whole-body mathematical model for intracranial pressure dynamics. Journal of Mathematical Biology, 46(4):347–383, April 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-002-0177-3.pdf.

Martcheva:2003:PAE

[1392] Maia Martcheva and Horst R. Thieme. Progression age enhanced backward bifurcation in an epidemic model with super-infection. Journal of Mathematical Biology, 46(5):385–424, May 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-002-0181-7.pdf.

Culshaw:2003:MMC

[1393] Rebecca V. Culshaw, Shigui Ruan, and Glenn Webb. A mathematical model of cell-to-cell spread of HIV-1 that includes a time delay. Journal of Mathematical Biology, 46(5):425–444, May 2003. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-002-0191-5.pdf.

Eshel:2003:EDS

[1394] Ilan Eshel and Emilia Sansone. Evolutionary and dynamic stability in continuous population games. Journal of Mathematical Biology, 46(5): 445–459, May 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0194-2.pdf.

Dolak:2003:CMCb

[1395] Y. Dolak and T. Hillen. Cattaneo models for chemosensitive movement. Journal of Mathematical Biology, 46(5):460, May 2003. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0222-x.pdf.

Dolak:2003:CMCc

[1396] Y. Dolak and T. Hillen. Cattaneo models for chemosensitive movement. Journal of Mathematical Biology, 46(5):461–478, May 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0221-y.pdf.

Basu:2003:WDS

[1397] Srabashi Basu, Debi Prosad Burma, and Probal Chaudhuri. Words in DNA sequences: some case studies based on their frequency statistics. Journal of Mathematical Biology, 46(6):479–503, June 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-002-0185-3.pdf.

Andreasen:2003:DAI

[1398] Viggo Andreasen. Dynamics of annual influenza a epidemics with immuno-selection. Journal of Mathematical Biology, 46(6):504–536, June 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0186-2.pdf.

Geigant:2003:BAO

[1399] Edith Geigant and Michael Stoll. Bifurcation analysis of an orientational aggregation model. Journal of Mathematical Biology, 46(6):537– 563, June 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0187-1.pdf.

Margheri:2003:SEP

[1400] A. Margheri and C. Rebelo. Some examples of persistence in epidemiological models. Journal of Mathematical Biology, 46(6):564– 570, June 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0193-3.pdf.

Hernandez:2003:VOP

[1401] Maria-Josefina Hernandez and Ignacio Barradas. Variation in the outcome of population interactions: bifurcations and catastrophes. Journal of Mathematical Biology, 46(6):571–594, June 2003. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-002-0192-4.pdf.

Casasnovas:2003:ARR

[1402] Jaume Casasnovas, Joe Miro-Julia, and Francesc Rosselló. On the algebraic representation of RNA secondary structures with G · U pairs. Journal of Mathematical Biology, 47(1):1–22, July 2003. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-002-0188-0.pdf.

Ward:2003:EDQ

[1403] John P. Ward, John R. King, Adrian J. Koerber, Julie M. Croft, R. Elizabeth Sockett, and Paul Williams. Early development and quorum sensing in bacterial biofilms. Journal of Mathematical Biology, 47(1): 23–55, July 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-002-0190-6.pdf.

Schreiber:2003:ERU

[1404] Sebastian J. Schreiber and Glory A. Tobiason. The evolution of resource use. Journal of Mathematical Biology, 47(1):56–78, July 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0195-9.pdf.

Whiteley:2003:MMP

[1405] Jonathan P. Whiteley, David J. Gavaghan, and Clive E. W. Hahn. Mathematical modelling of pulmonary gas transport. Journal of Mathematical Biology, 47(1):79–99, July 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0196-8.pdf.

Ostby:2003:MMR

[1406] Ivar Østby, Leiv S. Rusten, Gunnar Kvalheim, and Per Grøttum. A mathematical model for reconstitution of granulopoiesis after high dose chemotherapy with autologous stem cell transplantation. Journal of Mathematical Biology, 47(2):101–136, August 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-003-0198-6.pdf.

Jones:2003:FMS

[1407] Don Jones, Hristo V. Kojouharov, Dung Le, and Hal Smith. The Freter model: A simple model of biofilm formation. Journal of Mathematical Biology, 47(2):137–152, August 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0202-1.pdf.

Ackleh:2003:CEC

[1408] Azmy S. Ackleh and Linda J. S. Allen. Competitive exclusion and coexistence for pathogens in an epidemic model with variable population size. Journal of Mathematical Biology, 47(2):153–168, August 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0207-9.pdf.

Stiedl:2003:FDC

[1409] O. Stiedl and M. Meyer. Fractal dynamics in circadian cardiac time series of corticotropin-releasing factor receptor subtype-2 deficient mice. Journal of Mathematical Biology, 47(2):169–197, August 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0197-7.pdf.

Vargas:2003:NAG

[1410] J. A. Vargas and R. F. del Castillo. Nuclear androdioecy and gynodioecy. Journal of Mathematical Biology, 47(3):199–221, August 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0200-3.pdf.

Jacobs:2003:CAC

[1411] F. J. A. Jacobs and J. A. J. Metz. On the concept of attractor for community-dynamical processes i: the case of unstructured populations. Journal of Mathematical Biology, 47(3):222–234, August 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0204-z.pdf.

Gyllenberg:2003:CAC

[1412] Mats Gyllenberg, F. J. A. Jacobs, and J. A. J. Metz. On the concept of attractor for community-dynamical processes II: the case of structured populations. Journal of Mathematical Biology, 47(3):235– 248, August 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0213-y.pdf.

Timofeeva:2003:WBP

[1413] Y. Timofeeva and S. Coombes. Wave bifurcation and propagation failure in a model of Ca^{2+} release. Journal of Mathematical Biology, 47(3):249– 269, August 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0205-y.pdf.

Villasana:2003:DDE

[1414] Minaya Villasana and Ami Radunskaya. A delay differential equation model for tumor growth. Journal of Mathematical Biology, 47(3):270– 294, August 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0211-0.pdf.

Basse:2003:MMA

[1415] Britta Basse, Bruce C. Baguley, Elaine S. Marshall, Wayne R. Joseph, Bruce van Brunt, Graeme Wake, and David J. N. Wall. A mathematical model for analysis of the cell cycle in cell lines derived from human tumors. Journal of Mathematical Biology, 47(4):295–312, September 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0203-0.pdf.

Ma:2003:EMQ

[1416] Chang-Xing Ma, George Casella, Ramon C. Littell, André I. Khuri, and Rongling Wu. Exponential mapping of quantitative trait loci governing allometric relationships in organisms. Journal of Mathematical Biology, 47(4):313–324, September 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0212-z.pdf.

Baird:2003:NAM

[1417] Mark E. Baird. Numerical approximations of the mean absorption crosssection of a variety of randomly oriented microalgal shapes. Journal of Mathematical Biology, 47(4):325–336, September 2003. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0215-9.pdf.

Mohle:2003:CPD

[1418] Martin Möhle and Serik Sagitov. Coalescent patterns in diploid exchangeable population models. Journal of Mathematical Biology, 47(4): 337–352, September 2003. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-003-0218-6.pdf.

Mogilner:2003:MIP

[1419] A. Mogilner, L. Edelstein-Keshet, L. Bent, and A. Spiros. Mutual interactions, potentials, and individual distance in a social aggregation. Journal of Mathematical Biology, 47(4):353–389, September 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0209-7.pdf.

Olofsson:2003:ESF

[1420] Peter Olofsson and Chad A. Shaw. Exact sampling formulas for multitype Galton–Watson processes. Journal of Mathematical Biology, 47(4): 390, September 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0233-7.pdf.

Friedman:2003:AMV

[1421] Avner Friedman and Youshan Tao. Analysis of a model of a virus that replicates selectively in tumor cells. Journal of Mathematical Biology, 47(5):391–423, November 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0199-5.pdf.

Franks:2003:MEG

[1422] S. J. Franks, H. M. Byrne, J. R. King, J. C. E. Underwood, and C. E. Lewis. Modelling the early growth of ductal carcinoma in situ of the breast. Journal of Mathematical Biology, 47(5):424–452, November 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0214-x.pdf.

Maia:2003:ASE

[1423] Leonardo P. Maia, Daniela F. Botelho, and José F. Fontanari. Analytical solution of the evolution dynamics on a multiplicative-fitness landscape. Journal of Mathematical Biology, 47(5):453–456, November 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-003-0208-8.pdf.

Garay:2003:WDV

[1424] József Garay. When does the variance of replicator fitness decrease? Journal of Mathematical Biology, 47(5):457–464, November 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-003-0219-5.pdf.

Cressman:2003:ESS

[1425] Ross Cressman, József Garay, and Zoltán Varga. Evolutionarily stable sets in the single-locus frequency-dependent model of natural selection. Journal of Mathematical Biology, 47(5):465–482, November 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-003-0217-7.pdf.

Hutson:2003:ED

[1426] V. Hutson, S. Martinez, K. Mischaikow, and G. T. Vickers. The evolution of dispersal. Journal of Mathematical Biology, 47(6):483–517, December 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0210-1.pdf.

Chen:2003:CBE

[1427] Kevin C. Chen, Roseanne M. Ford, and Peter T. Cummings. Cell balance equation for chemotactic bacteria with a biphasic tumbling frequency. Journal of Mathematical Biology, 47(6):518–546, December 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-003-0216-8.pdf.

Li:2003:CPS

[1428] Jia Li, Zhien Ma, Steve P. Blythe, and Carlos Castillo-Chavez. Coexistence of pathogens in sexually-transmitted disease models. Journal of Mathematical Biology, 47(6):547–568, December 2003. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0235-5.pdf.

Dercole:2003:RBE

[1429] Fabio Dercole. Remarks on branching-extinction evolutionary cycles. Journal of Mathematical Biology, 47(6):569–580, December 2003. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-003-0236-4.pdf.

deKoeijer:2004:QBC

[1430] Aline de Koeijer, Hans Heesterbeek, Bram Schreuder, Radulf Oberthür, John Wilesmith, Herman van Roermund, and Mart de Jong. Quantifying BSE control by calculating the basic reproduction ratio R_{0} for the infection among cattle. Journal of Mathematical Biology, 48(1):1– 22, January 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0206-x.pdf.

Ionides:2004:SMC

[1431] Edward L. Ionides, Kathy S. Fang, R. Rivkah Isseroff, and George F. Oster. Stochastic models for cell motion and taxis. Journal of Mathematical Biology, 48(1):23–37, January 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-003-0220-z.pdf.

Brette:2004:DOD

[1432] Romain Brette. Dynamics of one-dimensional spiking neuron models. Journal of Mathematical Biology, 48(1):38–56, January 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0223-9.pdf.

Kon:2004:PDT

[1433] Ryusuke Kon. Permanence of discrete-time Kolmogorov systems for two species and saturated fixed points. Journal of Mathematical Biology, 48(1):57–81, January 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0224-8.pdf.

Stoleriu:2004:QSS

[1434] I. Stoleriu, F. A. Davidson, and J. L. Liu. Quasi-steady state assumptions for non-isolated enzyme-catalysed reactions. Journal of Mathematical Biology, 48(1):82–104, January 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0225-7.pdf.

Smith:2004:MAD

[1435] J. D. H. Smith. A macroscopic approach to demography. Journal of Mathematical Biology, 48(1):105–118, January 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-003-0231-9.pdf.

Norman:2004:RNV

[1436] Rachel Norman, David Ross, M. Karen Laurenson, and Peter J. Hudson. The role of non-viraemic transmission on the persistence and dynamics of a tick borne virus — louping ill in red grouse (Lagopus lagopus scoticus) and mountain hares (Lepus timidus). Journal of Mathematical Biology, 48(2):119–134, February 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-002-0183-5.pdf.

Calsina:2004:SMR

[1437] Àngel Calsina and Sílvia Cuadrado. Small mutation rate and evolutionarily stable strategies in infinite dimensional adaptive dynamics. Journal of Mathematical Biology, 48(2):135–159, February 2004. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0226-6.pdf.

Song:2004:MNR

[1438] Yun S. Song and Jotun Hein. On the minimum number of recombination events in the evolutionary history of DNA sequences. Journal of Mathematical Biology, 48(2):160–186, February 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-003-0227-5.pdf.

Cantrell:2004:DRD

[1439] R. S. Cantrell and C. Cosner. Deriving reaction–diffusion models in ecology from interacting particle systems. Journal of Mathematical Biology, 48(2):187–217, February 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0229-3.pdf.

Gammack:2004:MRM

[1440] D. Gammack, C. R. Doering, and D. E. Kirschner. Macrophage response to mycobacteriumtuberculosis infection. Journal of Mathematical Biology, 48(2):218–242, February 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0232-8.pdf.

Osan:2004:MSW

[1441] Remus Osan, Rodica Curtu, Jonathan Rubin, and Bard Ermentrout. Multiple-spike waves in a one-dimensional integrate-and-fire neural network. Journal of Mathematical Biology, 48(3):243–274, March 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0228-4.pdf.

Lessard:2004:TLA

[1442] Sabin Lessard and John Wakeley. The two-locus ancestral graph in a subdivided population: convergence as the number of demes grows in the island model. Journal of Mathematical Biology, 48(3):275– 292, March 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0230-x.pdf.

Lutscher:2004:SEM

[1443] Frithjof Lutscher and Mark A. Lewis. Spatially-explicit matrix models. Journal of Mathematical Biology, 48(3):293–324, March 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0234-6.pdf.

Luther:2004:RCE

[1444] R. M. Luther and M. Broom. Rapid convergence to an equilibrium state in kleptoparasitic populations. Journal of Mathematical Biology, 48(3): 325–339, March 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0237-3.pdf.

Courteau:2004:SRE

[1445] Josiane Courteau and Sabin Lessard. Sex ratio evolution through group selection using diffusion approximation. Journal of Mathematical Biology, 48(3):340–356, March 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0240-8.pdf.

Tang:2004:ESH

[1446] Sanyi Tang and Lansun Chen. The effect of seasonal harvesting on stagestructured population models. Journal of Mathematical Biology, 48(4): 357–374, April 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0243-5.pdf.

Gaffney:2004:AMM

[1447] E. A. Gaffney. The application of mathematical modelling to aspects of adjuvant chemotherapy scheduling. Journal of Mathematical Biology, 48 (4):375–422, April 2004. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-003-0246-2.pdf.

Knolle:2004:DBP

[1448] Helmut Knolle. A discrete branching process model for the spread of HIV via steady sexual partnerships. Journal of Mathematical Biology, 48(4):

423–443, April 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0241-7.pdf.

Webb:2004:OPS

[1449] Steven D. Webb and Markus R. Owen. Oscillations and patterns in spatially discrete models for developmental intercellular signalling. Journal of Mathematical Biology, 48(4):444–476, April 2004. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-003-0247-1.pdf.

Ambrosi:2004:RSG

[1450] D. Ambrosi and F. Mollica. The role of stress in the growth of a multicell spheroid. Journal of Mathematical Biology, 48(5):477– 499, May 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0238-2.pdf.

Pollak:2004:MPR

[1451] Edward Pollak. Malthusian parameters, reproductive values and change under selection in self fertilizing age-structured populations. Journal of Mathematical Biology, 48(5):500–514, May 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-003-0242-6.pdf.

Kon:2004:PSS

[1452] Ryusuke Kon, Yasuhisa Saito, and Yasuhiro Takeuchi. Permanence of single-species stage-structured models. Journal of Mathematical Biology, 48(5):515–528, May 2004. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-003-0239-1.pdf.

Broom:2004:FMA

[1453] M. Broom and G. D. Ruxton. A framework for modelling and analysing conspecific brood parasitism. Journal of Mathematical Biology, 48(5): 529–544, May 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0244-4.pdf.

Culshaw:2004:OHT

[1454] Rebecca V. Culshaw, Shigui Ruan, and Raymond J. Spiteri. Optimal HIV treatment by maximising immune response. Journal of Mathematical Biology, 48(5):545–562, May 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0245-3.pdf.

Ahlbrandt:2004:MEP

[1455] Calvin Ahlbrandt, Gary Benson, and William Casey. Minimal entropy probability paths between genome families. Journal of Mathematical Biology, 48(5):563–590, May 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0248-0.pdf.

Tornoe:2004:GBP

[1456] Christoffer Wenzel Tornøe, Judith L. Jacobsen, and Henrik Madsen. Grey-box pharmacokinetic/pharmacodynamic modelling of a euglycaemic clamp study. Journal of Mathematical Biology, 48(6):591– 604, June 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0257-z.pdf.

Ledder:2004:DEB

[1457] Glenn Ledder, J. David Logan, and Anthony Joern. Dynamic energy budget models with size-dependent hazard rates. Journal of Mathematical Biology, 48(6):605–622, June 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0263-1.pdf.

Gunther:2004:HLD

[1458] Michael Günther, Valentin Keppler, André Seyfarth, and Reinhard Blickhan. Human leg design: optimal axial alignment under constraints. Journal of Mathematical Biology, 48(6):623–646, June 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-004-0269-3.pdf.

Greenman:2004:LAS

[1459] J. V. Greenman and T. G. Benton. Large amplification in stagestructured models: Arnol’d tongues revisited. Journal of Mathematical Biology, 48(6):647–671, June 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-004-0264-8.pdf.

Barnes:2004:MIE

[1460] B. Barnes, G. Farquhar, and K. Gan. Modelling the isotope enrichment of leaf water. Journal of Mathematical Biology, 48(6):672– 702, June 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0260-4.pdf.

Singh:2004:EVRa

[1461] Pawan P. Singh, Dirk E. Maier, John H. Cushman, Kamyar Haghighi, and Carlos Corvalan. Effect of viscoelastic relaxation on moisture transport in foods. Part I: Solution of general transport equation. Journal of Mathematical Biology, 49(1):1–19, July 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-003-0249-z.pdf.

Singh:2004:EVRb

[1462] Pawan P. Singh, Dirk E. Maier, John H. Cushman, and Osvaldo H. Campanella. Effect of viscoelastic relaxation on moisture transport in foods. Part II: Sorption and drying of soybeans. Journal of Mathematical Biology, 49(1):20–34, July 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0250-6.pdf.

Wikan:2004:DCH

[1463] Arild Wikan. Dynamical consequences of harvest in discrete agestructured population models. Journal of Mathematical Biology, 49(1): 35–55, July 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0251-5.pdf.

Cummings:2004:EUS

[1464] L. J. Cummings, S. L. Waters, J. A. D. Wattis, and S. J. Graham. The effect of ureteric stents on urine flow: Reflux. Journal of Mathematical Biology, 49(1):56–82, July 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0252-4.pdf.

Pugliese:2004:TMI

[1465] Andrea Pugliese and Lorenza Tonetto. Thresholds for macroparasite infections. Journal of Mathematical Biology, 49(1):83–110, July 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-004-0266-6.pdf.

Mantzaris:2004:MMT

[1466] Nikos V. Mantzaris, Steve Webb, and Hans G. Othmer. Mathematical modeling of tumor-induced angiogenesis. Journal of Mathematical Biology, 49(2):111–187, August 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0262-2.pdf.

Gourley:2004:SSP

[1467] Stephen A. Gourley and Yang Kuang. A stage structured predator– prey model and its dependence on maturation delay and death rate. Journal of Mathematical Biology, 49(2):188–200, August 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-004-0278-2.pdf.

Redner:2004:UCD

[1468] Oliver Redner and Michael Baake. Unequal crossover dynamics in discrete and continuous time. Journal of Mathematical Biology, 49(2):201– 226, August 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0273-7.pdf.

Kooi:2004:CSF

[1469] B. W. Kooi, L. D. J. Kuijper, and S. A. L. M. Kooijman. Consequences of symbiosis for food web dynamics. Journal of Mathematical Biology, 49(3):227–271, September 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0256-0.pdf.

Edwards:2004:RMR

[1470] David A. Edwards. Refining the measurement of rate constants in the BIAcore. Journal of Mathematical Biology, 49(3):272–292, September 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0270-x.pdf.

Hofbauer:2004:IBT

[1471] Franz Hofbauer, Josef Hofbauer, Peter Raith, and Thomas Steinberger. Intermingled basins in a two species system. Journal of Mathematical Biology, 49(3):293–309, September 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0253-3.pdf.

Choi:2004:TWS

[1472] Y. S. Choi, Juliet Lee, and Roger Lui. Traveling wave solutions for a onedimensional crawling nematode sperm cell model. Journal of Mathematical Biology, 49(3):310–328, September 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-003-0255-1.pdf.

Basse:2004:MCD

[1473] Britta Basse, Bruce C. Baguley, Elaine S. Marshall, Wayne R. Joseph, Bruce van Brunt, Graeme Wake, and David J. N. Wall. Modelling cell death in human tumour cell lines exposed to the anticancer drug paclitaxel. Journal of Mathematical Biology, 49(4):329– 357, October 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0254-2.pdf.

Iron:2004:SAT

[1474] David Iron, Juncheng Wei, and Matthias Winter. Stability analysis of Turing patterns generated by the Schnakenberg model. Journal of Mathematical Biology, 49(4):358–390, October 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-003-0258-y.pdf.

Vincent:2004:ENU

[1475] Andrew D. Vincent and Mary R. Myerscough. The effect of a nonuniform turning kernel on ant trail morphology. Journal of Mathematical Biology, 49(4):391–432, October 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-003-0201-2.pdf.

Margaria:2004:SIA

[1476] Gabriella Margaria, Eva Riccomagno, and Lisa J. White. Structural identifiability analysis of some highly structured families of statespace models using differential algebra. Journal of Mathematical Biology, 49(5):433– 454, November 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-003-0261-3.pdf.

Morgenthaler:2004:MCF

[1477] Stephan Morgenthaler, Pablo Herrero, and William G. Thilly. Multistage carcinogenesis and the fraction at risk. Journal of Mathematical Biology, 49(5):455–467, November 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-004-0271-9.pdf.

Barbour:2004:CSM

[1478] A. D. Barbour and A. Pugliese. Convergence of a structured metapopulation model to Levins’s model. Journal of Mathematical Biology, 49(5): 468–500, November 2004. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-004-0272-8.pdf.

Ostby:2004:SMH

[1479] Ivar Østby and Ragnar Winther. Stability of a model of human granulopoiesis using continuous maturation. Journal of Mathematical Biology, 49(5):501–536, November 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-004-0274-6.pdf.

MacArthur:2004:RSG

[1480] Ben D. MacArthur and Colin P. Please. Residual stress generation and necrosis formation in multi-cell tumour spheroids. Journal of Mathematical Biology, 49(6):537–552, December 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-004-0265-7.pdf.

Trapman:2004:BMS

[1481] Pieter Trapman, Ronald Meester, and Hans Heesterbeek. A branching model for the spread of infectious animal diseases in varying environments. Journal of Mathematical Biology, 49(6):553–576, December 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0267-5.pdf.

Gedeon:2004:PLI

[1482] Tomás Gedeon and Matt Holzer. Phase locking in integrate-and-fire models with refractory periods and modulation. Journal of Mathematical Biology, 49(6):577–603, December 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-004-0268-4.pdf.

Byrne:2004:NIK

[1483] Helen M. Byrne and Markus R. Owen. A new interpretation of the Keller–Segel model based on multiphase modelling. Journal of Mathematical Biology, 49(6):604–626, December 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-004-0276-4.pdf.

Enciso:2004:SMT

[1484] German Enciso and Eduardo D. Sontag. On the stability of a model of testosterone dynamics. Journal of Mathematical Biology, 49(6):627– 634, December 2004. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0291-5.pdf.

Mjolhus:2005:SSP

[1485] E. Mjølhus, A. Wikan, and T. Solberg. On synchronization in semelparous populations. Journal of Mathematical Biology, 50(1):1–21, January 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0275-5.pdf.

Beamish:2005:CDO

[1486] Dan Beamish, Christopher Peskun, and Jianhong Wu. Critical delay for overshooting in planned arm movements with delayed feedback. Journal of Mathematical Biology, 50(1):22–48, January 2005. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-004-0277-3.pdf.

Morale:2005:IPS

[1487] Daniela Morale, Vincenzo Capasso, and Karl Oelschläger. An interacting particle system modelling aggregation behavior: from individuals to populations. Journal of Mathematical Biology, 50(1):49–66, January 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0279-1.pdf.

Geritz:2005:RID

[1488] Stefan A. H. Geritz. Resident-invader dynamics and the coexistence of similar strategies. Journal of Mathematical Biology, 50(1):67–82, January 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0280-8.pdf.

Baake:2005:AMP

[1489] Ellen Baake, Michael Baake, Anton Bovier, and Markus Klein. An asymptotic maximum principle for essentially linear evolution models. Journal of Mathematical Biology, 50(1):83–114, January 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-004-0281-7.pdf.

Stoleriu:2005:EPI

[1490] I. Stoleriu, F. A. Davidson, and J. L. Liu. Effects of periodic input on the quasi-steady state assumptions for enzyme-catalysed reactions. Journal of Mathematical Biology, 50(2):115–132, February 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-004-0282-6.pdf.

Gyllenberg:2005:ICI

[1491] Mats Gyllenberg and Géza Meszéna. On the impossibility of coexistence of infinitely many strategies. Journal of Mathematical Biology, 50(2):133– 160, February 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0283-5.pdf.

Carrillo:2005:SED

[1492] C. Carrillo and P. Fife. Spatial effects in discrete generation population models. Journal of Mathematical Biology, 50(2):161–188, February 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0284-4.pdf.

Filbet:2005:DHM

[1493] Francis Filbet, Philippe Laurençot, and Benoît Perthame. Derivation of hyperbolic models for chemosensitive movement. Journal of Mathematical Biology, 50(2):189–207, February 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-004-0286-2.pdf.

Hossjer:2005:IEN

[1494] Ola Hössjer. Information and effective number of meioses in linkage analysis. Journal of Mathematical Biology, 50(2):208–232, February 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-004-0289-z.pdf.

deGunst:2005:MSP

[1495] M. C. M. de Gunst and J. G. Schouten. Model selection and parameter estimation for ion channel recordings with an application to the K^{+} outward-rectifier in barley leaf. Journal of Mathematical Biology, 50(3): 233–256, March 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0287-1.pdf.

Tang:2005:SDI

[1496] Sanyi Tang and Robert A. Cheke. State-dependent impulsive models of integrated pest management (IPM) strategies and their dynamic consequences. Journal of Mathematical Biology, 50(3):257– 292, March 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0290-6.pdf.

Batzel:2005:CRC

[1497] Jerry J. Batzel, Franz Kappel, and Susanne Timischl-Teschl. A cardiovascular-respiratory control system model including state delay with application to congestive heart failure in humans. Journal of Mathematical Biology, 50(3):293–335, March 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-004-0293-3.pdf.

Liu:2005:AMD

[1498] Tian Liu, Wei Zhao, Lili Tian, and Rongling Wu. An algorithm for molecular dissection of tumor progression. Journal of Mathematical Biology, 50(3):336–354, March 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-004-0297-z.pdf.

Burger:2005:MAI

[1499] Reinhard Bürger. A multilocus analysis of intraspecific competition and stabilizing selection on a quantitative trait. Journal of Mathematical Biology, 50(4):355–396, April 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-004-0294-2.pdf.

Scholz:2005:MHG

[1500] M. Scholz, C. Engel, and M. Loeffler. Modelling human granulopoiesis under poly-chemotherapy with G-CSF support. Journal of Mathematical Biology, 50(4):397–439, April 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-004-0295-1.pdf.

Koerber:2005:DSM

[1501] A. J. Koerber, J. R. King, and P. Williams. Deterministic and stochastic modelling of endosome escape by Staphylococcus aureus: “quorum” sensing by a single bacterium. Journal of Mathematical Biology, 50(4):

440–488, April 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0296-0.pdf.

Xia:2005:TOI

[1502] Huaxing Xia, Gail S. K. Wolkowicz, and Lin Wang. Transient oscillations induced by delayed growth response in the chemostat. Journal of Mathematical Biology, 50(5):489–530, April 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-004-0311-5.pdf.

Abu-Raddad:2005:CSE

[1503] L. J. Abu-Raddad and N. M. Ferguson. Characterizing the symmetric equilibrium of multi-strain host-pathogen systems in the presence of cross immunity. Journal of Mathematical Biology, 50(5):531– 558, April 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0292-4.pdf.

Schofield:2005:DHS

[1504] Peter G. Schofield, Mark A. J. Chaplain, and Stephen F. Hubbard. Dynamic heterogeneous spatio-temporal pattern formation in hostparasitoid systems with synchronised generations. Journal of Mathematical Biology, 50(5):559–583, April 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-004-0298-y.pdf.

Castrejon-Pita:2005:FDB

[1505] A. A. Castrejón-Pita, A. Sarmiento-Galán, J. R. Castrejón-Pita, and R. Castrejón-García. Fractal dimension in butterflies’ wings: a novel approach to understanding wing patterns? Journal of Mathematical Biology, 50(5):584–594, April 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-004-0302-6.pdf.

Vielle:2005:MAM

[1506] B. Vielle. Mathematical analysis of Mayer waves. Journal of Mathematical Biology, 50(5):595–606, April 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-004-0305-3.pdf.

Banks:2005:PSM

[1507] H. T. Banks and D. M. Bortz. A parameter sensitivity methodology in the context of HIV delay equation models. Journal of Mathematical Biology, 50(6):607–625, June 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-004-0299-x.pdf.

Hyman:2005:DSE

[1508] James M. Hyman and Jia Li. Differential susceptibility epidemic models. Journal of Mathematical Biology, 50(6):626–644, June 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-004-0301-7.pdf.

Joyce:2005:MIP

[1509] Paul Joyce, Zaid Abdo, José M. Ponciano, Leen De Gelder, Larry J. Forney, and Eva M. Top. Modeling the impact of periodic bottlenecks, unidirectional mutation, and observational error in experimental evolution. Journal of Mathematical Biology, 50(6):645–662, June 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-004-0300-8.pdf.

Xu:2005:HSE

[1510] Cailin Xu, Mark S. Boyce, and Daryl J. Daley. Harvesting in seasonal environments. Journal of Mathematical Biology, 50(6):663– 682, June 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0303-5.pdf.

Pedersen:2005:WSD

[1511] Morten Gram Pedersen. Wave speeds of density dependent Nagumo diffusion equations — inspired by oscillating gap-junction conductance in the islets of Langerhans. Journal of Mathematical Biology, 50(6): 683–698, June 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0304-4.pdf.

Tang:2005:HBR

[1512] Yilei Tang and Weinian Zhang. Heteroclinic bifurcation in a ratiodependent predator–prey system. Journal of Mathematical Biology, 50 (6):699–712, June 2005. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-004-0307-1.pdf.

Buffoni:2005:MPP

[1513] G. Buffoni, M. P. Cassinari, M. Groppi, and M. Serluca. Modelling of predator–prey trophic interactions. Part I: two trophic levels. Journal of Mathematical Biology, 50(6):713–732, June 2005. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-004-0312-4.pdf.

Merkin:2005:SM

[1514] J. H. Merkin and B. D. Sleeman. On the spread of morphogens. Journal of Mathematical Biology, 51(1):1–17, July 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-004-0308-0.pdf.

Sumner:2005:EIP

[1515] J. G. Sumner and P. D. Jarvis. Entanglement invariants and phylogenetic branching. Journal of Mathematical Biology, 51(1):18–36, July 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0309-z.pdf.

Tao:2005:CDB

[1516] Youshan Tao and Qian Guo. The competitive dynamics between tumor cells, a replication-competent virus and an immune response. Journal of Mathematical Biology, 51(1):37–74, July 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-004-0310-6.pdf.

Berestycki:2005:APF

[1517] Henri Berestycki, François Hamel, and Lionel Roques. Analysis of the periodically fragmented environment model: I — species persistence. Journal of Mathematical Biology, 51(1):75–113, July 2005. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-004-0313-3.pdf.

Maia:2005:DWM

[1518] Leonardo P. Maia. The dynamical way to mutation–selection balance of an infinite population evolving on a truncated fitness landscape. Journal of Mathematical Biology, 51(1):114–122, July 2005. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0327-5.pdf.

Veliov:2005:EPH

[1519] Vladimir M. Veliov. On the effect of population heterogeneity on dynamics of epidemic diseases. Journal of Mathematical Biology, 51(2): 123–143, August 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0288-0.pdf.

Hornaes:2005:CPC

[1520] Hans Petter Hornæs, Jan Henrik Wold, and Ivar Farup. Colorimetry and prime colours- a theorem. Journal of Mathematical Biology, 51(2): 144–156, August 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0306-2.pdf.

Rubel:2005:AVG

[1521] J. Rübel and J. Starke. Analysis and visualization of growth-related and treatment-induced craniofacial changes. Journal of Mathematical Biology, 51(2):157–170, August 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-005-0314-x.pdf.

Baroni:2005:BNH

[1522] Mihaela Baroni, Stefan Grünewald, Vincent Moulton, and Charles Semple. Bounding the number of hybridisation events for a consistent evolutionary history. Journal of Mathematical Biology, 51(2):171–182, August 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0315-9.pdf.

Sherratt:2005:AVS

[1523] Jonathan A. Sherratt. An analysis of vegetation stripe formation in semi–arid landscapes. Journal of Mathematical Biology, 51(2):183– 197, August 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0319-5.pdf.

Mazza:2005:SSN

[1524] Christian Mazza. Strand separation in negatively supercoiled DNA. Journal of Mathematical Biology, 51(2):198–216, August 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0320-z.pdf.

Friedman:2005:MIT

[1525] Avner Friedman and Gheorghe Craciun. A model of intracellular transport of particles in an axon. Journal of Mathematical Biology, 51(2): 217–246, August 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-004-0285-3.pdf.

Murase:2005:SAP

[1526] Akiko Murase, Toru Sasaki, and Tsuyoshi Kajiwara. Stability analysis of pathogen-immune interaction dynamics. Journal of Mathematical Biology, 51(3):247–267, September 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-005-0321-y.pdf.

Nagylaki:2005:SMP

[1527] Thomas Nagylaki. A stochastic model for a progressive chronic disease. Journal of Mathematical Biology, 51(3):268–280, September 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-005-0323-9.pdf.

Cohen:2005:ESM

[1528] Elisheva Cohen and David A. Kessler. Equilibrium state of molecular breeding. Journal of Mathematical Biology, 51(3):281–301, September 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-005-0324-8.pdf.

Hilpert:2005:LBM

[1529] Markus Hilpert. Lattice-Boltzmann model for bacterial chemotaxis. Journal of Mathematical Biology, 51(3):302–332, September 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-005-0318-6.pdf.

Dercole:2005:STC

[1530] F. Dercole, K. Niklas, and R. Rand. Self-thinning and community persistence in a simple size-structured dynamical model of plant growth. Journal of Mathematical Biology, 51(3):333–354, September 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0322-x.pdf.

Goriely:2005:GIC

[1531] Alain Goriely, György Károlyi, and Michael Tabor. Growth induced curve dynamics for filamentary micro-organisms. Journal of Mathematical Biology, 51(3):355–366, September 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-005-0328-4.pdf.

Allegretto:2005:PSM

[1532] Walter Allegretto, Chiara Mocenni, and Antonio Vicino. Periodic solutions in modelling lagoon ecological interactions. Journal of Mathematical Biology, 51(4):367–388, October 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-005-0325-7.pdf.

Golinski:2005:ECF

[1533] Michael Golinski, Ernest Barany, and Mary Ballyk. Ecological conditions that favor the evolution of intermediate-virulence in an environmentally transmitted parasite. Journal of Mathematical Biology, 51(4):389– 402, October 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0326-6.pdf.

Yin:2005:GMF

[1534] Yajun Yin, Jie Yin, and Dong Ni. General mathematical frame for open or closed biomembranes (part i): Equilibrium theory and geometrically constraint equation. Journal of Mathematical Biology, 51(4):403– 413, October 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0330-x.pdf.

Gomes:2005:DBE

[1535] M. G. M. Gomes, A. Margheri, G. F. Medley, and C. Rebelo. Dynamical behaviour of epidemiological models with sub-optimal immunity and nonlinear incidence. Journal of Mathematical Biology, 51(4):414– 430, October 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0331-9.pdf.

Sanchez:2005:GAG

[1536] R. Sanchez, E. Morgado, and R. Grau. Gene algebra from a genetic code algebraic structure. Journal of Mathematical Biology, 51(4):431– 457, October 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0332-8.pdf.

Ballyk:2005:GAC

[1537] Mary M. Ballyk, C. Connell McCluskey, and Gail S. K. Wolkowicz. Global analysis of competition for perfectly substitutable resources with linear response. Journal of Mathematical Biology, 51(4):458–490, October 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0333-7.pdf.

Jung:2005:MMC

[1538] Andreas Jung, Rupert Faltermeier, Ralf Rothoerl, and Alexander Brawanski. A mathematical model of cerebral circulation and oxygen supply. Journal of Mathematical Biology, 51(5):491–507, November 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-005-0343-5.pdf.

Fowler:2005:DRM

[1539] A. C. Fowler and M. J. McGuinness. A delay recruitment model of the cardiovascular control system. Journal of Mathematical Biology, 51(5): 508–526, November 2005. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-005-0339-1.pdf.

Codling:2005:CSS

[1540] E. A. Codling and N. A. Hill. Calculating spatial statistics for velocity jump processes with experimentally observed reorientation parameters. Journal of Mathematical Biology, 51(5):527–556, November 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-005-0317-7.pdf.

Anguige:2005:MAA

[1541] K. Anguige, J. R. King, and J. P. Ward. Modelling antibiotic- and anti-quorum sensing treatment of a spatially-structured Pseudomonas aeruginosa population. Journal of Mathematical Biology, 51(5):557–594, November 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0316-8.pdf.

Dolak:2005:KMC

[1542] Y. Dolak and C. Schmeiser. Kinetic models for chemotaxis: Hydrodynamic limits and spatio-temporal mechanisms. Journal of Mathematical Biology, 51(6):595–615, December 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-005-0334-6.pdf.

Foppa:2005:BRN

[1543] Ivo M. Foppa. The basic reproductive number of tick-borne encephalitis virus. Journal of Mathematical Biology, 51(6):616–628, December 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0337-3.pdf.

Hethcote:2005:SCP

[1544] Herbert W. Hethcote, Wendi Wang, and Yi Li. Species coexistence and periodicity in host–host-pathogen models. Journal of Mathematical Biology, 51(6):629–660, December 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-005-0335-5.pdf.

Akman:2005:EMI

[1545] O. E. Akman, D. S. Broomhead, R. V. Abadi, and R. A. Clement. Eye movement instabilities and nystagmus can be predicted by a nonlinear dynamics model of the saccadic system. Journal of Mathematical Biology, 51(6):661–694, December 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-005-0336-4.pdf.

Getto:2005:DAC

[1546] Ph. Getto, O. Diekmann, and A. M. de Roos. On the (dis) advantages of cannibalism. Journal of Mathematical Biology, 51(6):695–712, December 2005. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0342-6.pdf.

Parvinen:2006:FVA

[1547] Kalle Parvinen, Ulf Dieckmann, and Mikko Heino. Function-valued adaptive dynamics and the calculus of variations. Journal of Mathematical Biology, 52(1):1–26, January 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-005-0329-3.pdf.

Casey:2006:PLM

[1548] Richard Casey, Hidde de Jong, and Jean-Luc Gouzé. Piecewise-linear models of genetic regulatory networks: Equilibria and their stability. Journal of Mathematical Biology, 52(1):27–56, January 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0338-2.pdf.

Garcia:2006:TTG

[1549] M. V. P. Garcia, C. Possani, R. Ranvaud, and F. A. Tal. Three theorems on gradient fields potentially useful in homing pigeon navigation. Journal of Mathematical Biology, 52(1):57–69, January 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0340-8.pdf.

Seno:2006:GSD

[1550] H. Seno. Group size determined by fusion and fission a mathematical modelling with inclusive fitness. Journal of Mathematical Biology, 52(1): 70–92, January 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0341-7.pdf.

Ayati:2006:SPM

[1551] Bruce P. Ayati. A structured-population model of Proteus mirabilis swarm-colony development. Journal of Mathematical Biology, 52(1):93– 114, January 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0345-3.pdf.

Jones:2006:LNP

[1552] S. R. Jones and N. Kopell. Local network parameters can affect internetwork phase lags in central pattern generators. Journal of Mathematical Biology, 52(1):115–140, January 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-005-0348-0.pdf.

Moehlis:2006:CRH

[1553] Jeff Moehlis. Canards for a reduction of the Hodgkin–Huxley equations. Journal of Mathematical Biology, 52(2):141–153, February 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0347-1.pdf.

Nikolaou:2006:NMA

[1554] Michael Nikolaou and Vincent H. Tam. A new modeling approach to the effect of antimicrobial agents on heterogeneous microbial populations. Journal of Mathematical Biology, 52(2):154–182, February 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0350-6.pdf.

deCastro:2006:SAS

[1555] Manuela L. de Castro, Jacques A. L. Silva, and Dagoberto A. R. Justo. Stability in an age-structured metapopulation model. Journal of Mathematical Biology, 52(2):183–208, February 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-005-0352-4.pdf.

Medvedev:2006:KPD

[1556] E. S. Medvedev and A. A. Stuchebrukhov. Kinetics of proton diffusion in the regimes of fast and slow exchange between the membrane surface and the bulk solution. Journal of Mathematical Biology, 52(2):209– 234, February 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0354-2.pdf.

Yoshiyama:2006:CSV

[1557] Kohei Yoshiyama and Hisao Nakajima. Catastrophic shifts in vertical distributions of phytoplankton the existence of a bifurcation set. Journal of Mathematical Biology, 52(2):235–276, February 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0349-z.pdf.

Qian:2006:ACA

[1558] H. Qian. An asymptotic comparative analysis of the thermodynamics of non–covalent association. Journal of Mathematical Biology, 52(3): 277–289, March 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0353-3.pdf.

Alexander:2006:MEB

[1559] M. E. Alexander, S. M. Moghadas, P. Rohani, and A. R. Summers. Modelling the effect of a booster vaccination on disease epidemiology. Journal of Mathematical Biology, 52(3):290–306, March 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0356-0.pdf.

Anderson:2006:FAC

[1560] James W. Anderson, Keith R. Fox, and Graham A. Niblo. A fast algorithm for the construction of universal footprinting templates in DNA. Journal of Mathematical Biology, 52(3):307–342, March 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0357-z.pdf.

Guedon:2006:AGN

[1561] Yann Guédon, Yves d’Aubenton Carafa, and Claude Thermes. Analysing grouping of nucleotides in DNA sequences using lumped processes constructed from Markov chains. Journal of Mathematical Biology, 52(3): 343–372, March 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0358-y.pdf.

Farcot:2006:GPC

[1562] Etienne Farcot. Geometric properties of a class of piecewise affine biological network models. Journal of Mathematical Biology, 52(3):373– 418, March 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0360-4.pdf.

DeLeenheer:2006:GSC

[1563] Patrick De Leenheer, Simon A. Levin, Eduardo D. Sontag, and Christopher A. Klausmeier. Global stability in a chemostat with multiple nutrients. Journal of Mathematical Biology, 52(4):419–438, April 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-005-0344-4.pdf.

Hopfner:2006:SMF

[1564] Reinhard Höpfner and Klaus Brodda. A stochastic model and a functional central limit theorem for information processing in large systems of neurons. Journal of Mathematical Biology, 52(4):439– 457, April 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0361-3.pdf.

Baker:2006:MIC

[1565] R. E. Baker, S. Schnell, and P. K. Maini. A mathematical investigation of a clock and wavefront model for somitogenesis. Journal of Mathematical Biology, 52(4):458–482, April 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-005-0362-2.pdf.

Schneider:2006:MMA

[1566] Kristan A. Schneider. A multilocus-multiallele analysis of frequencydependent selection induced by intraspecific competition. Journal of Mathematical Biology, 52(4):483–523, April 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-005-0355-1.pdf.

Coutinho:2006:DTP

[1567] R. Coutinho, B. Fernandez, R. Lima, and A. Meyroneinc. Discrete time piecewise affine models of genetic regulatory networks. Journal of Mathematical Biology, 52(4):524–570, April 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-005-0359-x.pdf.

Lemon:2006:MME

[1568] Greg Lemon, John R. King, Helen M. Byrne, Oliver E. Jensen, and Kevin M. Shakesheff. Mathematical modelling of engineered tissue growth using a multiphase porous flow mixture theory. Journal of Mathematical Biology, 52(5):571–594, May 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-005-0363-1.pdf.

Eshel:2006:GBE

[1569] Ilan Eshel, Emilia Sansone, and Avner Shaked. Gregarious behaviour of evasive prey. Journal of Mathematical Biology, 52(5):595– 612, May 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0364-0.pdf.

Huber:2006:PNM

[1570] K. T. Huber and V. Moulton. Phylogenetic networks from multilabelled trees. Journal of Mathematical Biology, 52(5):613–632, May 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-005-0365-z.pdf.

DeLara:2006:MWD

[1571] Michel De Lara. Mum, why do you keep on growing?Impacts of environmental variability on optimal growth and reproduction allocation strategies of annual plants. Journal of Mathematical Biology, 52(5):

633–666, May 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0366-y.pdf.

Imhof:2006:EGD

[1572] Lorens A. Imhof and Martin A. Nowak. Evolutionary game dynamics in a Wright–Fisher process. Journal of Mathematical Biology, 52(5): 667–681, May 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0369-8.pdf.

Litcanu:2006:SPA

[1573] Gabriela Litcanu and Juan J. L. Velázquez. Singular perturbation analysis of cAMP signalling in Dictyostelium discoideum aggregates. Journal of Mathematical Biology, 52(5):682–718, May 2006. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-005-0370-2.pdf.

Schreiber:2006:HPD

[1574] Sebastian J. Schreiber. Host-parasitoid dynamics of a generalized Thompson model. Journal of Mathematical Biology, 52(6):719–732, June 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0346-2.pdf.

Chow:2006:SMT

[1575] K. W. Chow and C. C. Mak. A simple model for the two dimensional blood flow in the collapse of veins. Journal of Mathematical Biology, 52(6):733–744, June 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-005-0351-5.pdf.

Baer:2006:MLC

[1576] Steven M. Baer, Bingtuan Li, and Hal L. Smith. Multiple limit cycles in the standard model of three species competition for three essential resources. Journal of Mathematical Biology, 52(6):745–760, June 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-005-0367-x.pdf.

Ross:2006:SMM

[1577] J. V. Ross. A stochastic metapopulation model accounting for habitat dynamics. Journal of Mathematical Biology, 52(6):788–806, June 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-006-0372-8.pdf.

Shi:2006:PRD

[1578] Junping Shi and Ratnasingham Shivaji. Persistence in reaction diffusion models with weak Allee effect. Journal of Mathematical Biology, 52(6): 807–829, June 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0373-7.pdf.

Meyer:2006:FRE

[1579] M. Meyer and O. Stiedl. Fractal rigidity by enhanced sympatho-vagal antagonism in heartbeat interval dynamics elicited by central application of corticotropin-releasing factor in mice. Journal of Mathematical Biology, 52(6):830–874, June 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0375-5.pdf.

Miller:2006:FLE

[1580] J. R. Miller, M. C. Pugh, and M. B. Hamilton. A finite locus effect diffusion model for the evolution of a quantitative trait. Journal of Mathematical Biology, 52(6):??, June 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-005-0368-9.pdf.

Thygesen:2006:DTS

[1581] Uffe Høgsbro Thygesen and Thomas Kiørboe. Diffusive transport in Stokeslet flow and its application to plankton ecology. Journal of Mathematical Biology, 53(1):1–14, July 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0374-6.pdf.

Grafen:2006:TFR

[1582] Alan Grafen. A theory of Fisher’s reproductive value. Journal of Mathematical Biology, 53(1):15–60, July 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0376-4.pdf.

Sharkey:2006:PLA

[1583] Kieran J. Sharkey, Carmen Fernandez, Kenton L. Morgan, Edmund Peeler, Mark Thrush, James F. Turnbull, and Roger G. Bowers. Pairlevel approximations to the spatio-temporal dynamics of epidemics on asymmetric contact networks. Journal of Mathematical Biology, 53(1): 61–85, July 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0377-3.pdf.

Hogea:2006:SCT

[1584] Cosmina S. Hogea, Bruce T. Murray, and James A. Sethian. Simulating complex tumor dynamics from avascular to vascular growth using a general level-set method. Journal of Mathematical Biology, 53(1): 86–134, July 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0378-2.pdf.

Clote:2006:STW

[1585] Peter Clote and Jürg Straubhaar. Symmetric time warping, Boltzmann pair probabilities and functional genomics. Journal of Mathematical Biology, 53(1):135–161, July 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0379-1.pdf.

Friedrich:2006:MMH

[1586] Benjamin Friedrich. A mesoscopic model for helical bacterial flagella. Journal of Mathematical Biology, 53(1):162–178, July 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0380-8.pdf.

Guo:2006:ABS

[1587] Jong-Shenq Guo and Je-Chiang Tsai. The asymptotic behavior of solutions of the buffered bistable system. Journal of Mathematical Biology, 53(1):179–213, July 2006. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-006-0381-7.pdf.

Malik:2006:RBM

[1588] Tufail Malik and Hal Smith. A resource-based model of microbial quiescence. Journal of Mathematical Biology, 53(2):231–252, August 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0003-4.pdf.

Chen:2006:SIE

[1589] Frederick H. Chen. A susceptible-infected epidemic model with voluntary vaccinations. Journal of Mathematical Biology, 53(2):253–272, August 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0006-1.pdf.

Kobayashi:2006:MMR

[1590] Ryo Kobayashi, Atsushi Tero, and Toshiyuki Nakagaki. Mathematical model for rhythmic protoplasmic movement in the true slime mold. Journal of Mathematical Biology, 53(2):273–286, August 2006. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0007-0.pdf.

Whitman:2006:ATI

[1591] Alan M. Whitman and Hashem Ashrafiuon. Asymptotic theory of an infectious disease model. Journal of Mathematical Biology, 53(2):287– 304, August 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0009-y.pdf.

Morozov:2006:PAP

[1592] Andrew Y. Morozov and Bai-Lian Li. Parametric analysis of a predator– prey system stabilized by a top predator. Journal of Mathematical Biology, 53(2):305–335, August 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0008-z.pdf.

OHely:2006:DAA

[1593] Martin O’Hely. A diffusion approach to approximating preservation probabilities for gene duplicates. Journal of Mathematical Biology, 53(2): ??, August 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0001-6.pdf.

Alt:2006:EES

[1594] Wolfgang Alt. ESMTB — European Society for Mathematical and Theoretical Biology. Journal of Mathematical Biology, 53(3):337–339, September 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0030-1.pdf.

Llabres:2006:ACT

[1595] Mercè Llabrés, Jairo Rocha, Francesc Rosselló, and Gabriel Valiente. On the ancestral compatibility of two phylogenetic trees with nested taxa. Journal of Mathematical Biology, 53(3):340–364, September 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-006-0011-4.pdf.

El-Morshedy:2006:GAF

[1596] Hassan A. El-Morshedy and Eduardo Liz. Globally attracting fixed points in higher order discrete population models. Journal of Mathematical Biology, 53(3):365–384, September 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-006-0014-1.pdf.

Zhang:2006:MMN

[1597] Wensheng Zhang and Aurélie Edwards. Mathematical model of nitric oxide convection and diffusion in a renal medullary vas rectum. Journal of Mathematical Biology, 53(3):385–420, September 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0018-x.pdf.

Bacaer:2006:ETV

[1598] Nicolas Bacaër and Souad Guernaoui. The epidemic threshold of vectorborne diseases with seasonality. Journal of Mathematical Biology, 53(3): 421–436, September 2006. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-006-0015-0.pdf.

Belik:2006:SMP

[1599] Pavel Belík, P. W. A. Dayananda, John T. Kemper, and Mikhail M. Shvartsman. A stochastic model for PSA levels: behavior of solutions and population statistics. Journal of Mathematical Biology, 53(3):437– 463, September 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0016-z.pdf.

Sparenberg:2006:NGB

[1600] J. A. Sparenberg. Note on the glide of a bird with wings bent downwards. Journal of Mathematical Biology, 53(3):464–473, September 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-006-0019-9.pdf.

Pinto:2006:CPG

[1601] Carla M. A. Pinto and Martin Golubitsky. Central pattern generators for bipedal locomotion. Journal of Mathematical Biology, 53(3):474– 489, September 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0021-2.pdf.

Sumner:2006:EIP

[1602] J. G. Sumner and P. D. Jarvis. Entanglement invariants and phylogenetic branching. Journal of Mathematical Biology, 53(3):490, September 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0010-5.pdf.

Hillen:2006:P

[1603] Thomas Hillen, Frithjof Lutscher, and Johannes Müller. Preface. Journal of Mathematical Biology, 53(4):491–495, October 2006. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0025-y.pdf.

Levin:2006:KHB

[1604] Simon Levin. On Karl Hadeler becoming 70. Journal of Mathematical Biology, 53(4):496–498, October 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0026-x.pdf.

Colijn:2006:HFS

[1605] Caroline Colijn, A. C. Fowler, and Michael C. Mackey. High frequency spikes in long period blood cell oscillations. Journal of Mathematical Biology, 53(4):499–519, October 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0027-9.pdf.

Cushing:2006:JAM

[1606] J. M. Cushing. A juvenile-adult model with periodic vital rates. Journal of Mathematical Biology, 53(4):520–539, October 2006. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0382-6.pdf.

Esteva:2006:MPP

[1607] Lourdes Esteva, Gerardo Rivas, and Hyun Mo Yang. Modelling parasitism and predation of mosquitoes by water mites. Journal of Mathematical Biology, 53(4):540–555, October 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-006-0020-3.pdf.

Hesseler:2006:CCR

[1608] Julia Heßeler, Julia K. Schmidt, Udo Reichl, and Dietrich Flockerzi. Coexistence in the chemostat as a result of metabolic by-products. Journal of Mathematical Biology, 53(4):556–584, October 2006. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0012-3.pdf.

Hillen:2006:MMM

[1609] Thomas Hillen. M^{5} mesoscopic and macroscopic models for mesenchymal motion. Journal of Mathematical Biology, 53(4):585–616, October 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0017-y.pdf.

Iida:2006:DCD

[1610] Masato Iida, Masayasu Mimura, and Hirokazu Ninomiya. Diffusion, cross-diffusion and competitive interaction. Journal of Mathematical Biology, 53(4):617–641, October 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0013-2.pdf.

Martcheva:2006:KDE

[1611] Maia Martcheva, Horst R. Thieme, and Thanate Dhirasakdanon. Kolmogorov’s differential equations and positive semigroups on first moment sequence spaces. Journal of Mathematical Biology, 53(4):642– 671, October 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0002-5.pdf.

Muller:2006:CCC

[1612] Johannes Müller, Christina Kuttler, Burkard A. Hense, Michael Rothballer, and Anton Hartmann. Cell-cell communication by quorum sensing and dimension-reduction. Journal of Mathematical Biology, 53(4):672– 702, October 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0024-z.pdf.

Safan:2006:MER

[1613] Muntaser Safan, Hans Heesterbeek, and Klaus Dietz. The minimum effort required to eradicate infections in models with backward bifurcation. Journal of Mathematical Biology, 53(4):703–718, October 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0028-8.pdf.

Shim:2006:ASE

[1614] E. Shim, Z. Feng, M. Martcheva, and C. Castillo-Chavez. An agestructured epidemic model of rotavirus with vaccination. Journal of Mathematical Biology, 53(4):719–746, October 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-006-0023-0.pdf.

Smith:2006:DDM

[1615] H. L. Smith. The discrete dynamics of monotonically decomposable maps. Journal of Mathematical Biology, 53(4):??, October 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0004-3.pdf.

Rudnicki:2006:MEP

[1616] Ryszard Rudnicki, Jerzy Tiuryn, and Damian Wójtowicz. A model for the evolution of paralog families in genomes. Journal of Mathematical Biology, 53(5):759–770, November 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0040-z.pdf.

Picchini:2006:MEH

[1617] Umberto Picchini, Susanne Ditlevsen, and Andrea De Gaetano. Modeling the euglycemic hyperinsulinemic clamp by stochastic differential equations. Journal of Mathematical Biology, 53(5):771–796, November 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-006-0032-z.pdf.

Alford:2006:RWS

[1618] John G. Alford and Giles Auchmuty. Rotating wave solutions of the FitzHugh–Nagumo equations. Journal of Mathematical Biology, 53(5): 797–819, November 2006. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-006-0022-1.pdf.

Wiuf:2006:CEP

[1619] Carsten Wiuf. Consistency of estimators of population scaled parameters using composite likelihood. Journal of Mathematical Biology, 53(5):821– 841, November 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0031-0.pdf.

McMillen:2006:ERM

[1620] T. McMillen and P. Holmes. An elastic rod model for anguilliform swimming. Journal of Mathematical Biology, 53(5):843–886, November 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0036-8.pdf.

Deutsch:2006:EES

[1621] Andreas Deutsch. ESMTB-European society for mathematical and theoretical biology. Journal of Mathematical Biology, 53(5):887–888, November 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0041-y.pdf.

Grover:2006:DA

[1622] N. B. Grover. Diffusion with attrition. Journal of Mathematical Biology, 53(6):889–903, December 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0029-7.pdf.

Muller:2006:GMR

[1623] Stefan Müller, Josef Hofbauer, Lukas Endler, Christoph Flamm, Stefanie Widder, and Peter Schuster. A generalized model of the repressilator. Journal of Mathematical Biology, 53(6):905–937, December 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0035-9.pdf.

Preziosi:2006:HMM

[1624] Luigi Preziosi. Hybrid and multiscale modelling. Journal of Mathematical Biology, 53(6):977–978, December 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0042-x.pdf.

Wood:2006:LSN

[1625] B. P. Wood and J. R. Miller. Linked selected and neutral loci in heterogeneous environments. Journal of Mathematical Biology, 53(6):??, December 2006. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0038-6.pdf.

Jahnke:2007:SCM

[1626] Tobias Jahnke and Wilhelm Huisinga. Solving the chemical master equation for monomolecular reaction systems analytically. Journal of Mathematical Biology, 54(1):1–26, January 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-006-0034-x.pdf.

Song:2007:DSH

[1627] Baojun Song and Diana M. Thomas. Dynamics of starvation in humans. Journal of Mathematical Biology, 54(1):27–43, January 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0037-7.pdf.

Pietruszka:2007:APG

[1628] M. Pietruszka and S. Lewicka. Anisotropic plant growth due to phototropism. Journal of Mathematical Biology, 54(1):45–55, January 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0045-7.pdf.

Luzyanina:2007:CAC

[1629] Tatyana Luzyanina, Sonja Mrusek, John T. Edwards, Dirk Roose, Stephan Ehl, and Gennady Bocharov. Computational analysis of CFSE proliferation assay. Journal of Mathematical Biology, 54(1):57– 89, January 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0046-6.pdf.

Lewicka:2007:APC

[1630] S. Lewicka and M. Pietruszka. Anisotropic plant cell elongation due to ortho-gravitropism. Journal of Mathematical Biology, 54(1):91–100, January 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0049-3.pdf.

Westerhoff:2007:MTB

[1631] Hans V. Westerhoff. Mathematical and theoretical biology for systems biology, and then … vice versa. Journal of Mathematical Biology, 54(1):147–150, January 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0043-9.pdf.

Inaba:2007:ASH

[1632] Hisashi Inaba. Age-structured homogeneous epidemic systems with application to the MSEIR epidemic model. Journal of Mathematical Biology, 54(1):??, January 2007. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-006-0033-y.pdf.

Broom:2007:EKS

[1633] M. Broom and J. Rychtár. The evolution of a kleptoparasitic system under adaptive dynamics. Journal of Mathematical Biology, 54(2):151– 177, February 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0005-2.pdf.

Iber:2007:MSS

[1634] Dagmar Iber and Giorgio Gaglia. The mechanism of sudden stripe formation during dorso-ventral patterning in Drosophila. Journal of Mathematical Biology, 54(2):179–198, February 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-006-0047-5.pdf.

Sun:2007:ADW

[1635] L. Sun and M. Wang. An algorithm for a decomposition of weighted digraphs: with applications to life cycle analysis in ecology. Journal of Mathematical Biology, 54(2):199–226, February 2007. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0048-4.pdf.

Ellner:2007:SSP

[1636] Stephen P. Ellner and Mark Rees. Stochastic stable population growth in integral projection models: theory and application. Journal of Mathematical Biology, 54(2):227–256, February 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-006-0044-8.pdf.

Baake:2007:MSA

[1637] Ellen Baake and Hans-Otto Georgii. Mutation, selection, and ancestry in branching models: a variational approach. Journal of Mathematical Biology, 54(2):257–303, February 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0039-5.pdf.

Coombes:2007:MN

[1638] S. Coombes. Mathematical neuroscience. Journal of Mathematical Biology, 54(2):305–307, February 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0066-2.pdf.

Gourley:2007:EVB

[1639] Stephen A. Gourley, Rongsong Liu, and Jianhong Wu. Eradicating vector-borne diseases via age-structured culling. Journal of Mathematical Biology, 54(3):309–335, March 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0050-x.pdf.

Hiebeler:2007:CPF

[1640] David E. Hiebeler. Competing populations on fragmented landscapes with spatially structured heterogeneities: improved landscape generation and mixed dispersal strategies. Journal of Mathematical Biology, 54(3): 337–356, March 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0054-6.pdf.

Adams:2007:ACC

[1641] Malcolm R. Adams and Andrew T. Sornborger. Analysis of a certain class of replicator equations. Journal of Mathematical Biology, 54(3): 357–384, March 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0055-5.pdf.

Krakovska:2007:CVB

[1642] O. Krakovska and L. M. Wahl. Costs versus benefits: best possible and best practical treatment regimens for HIV. Journal of Mathematical Biology, 54(3):385–406, March 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0059-1.pdf.

Froda:2007:PPP

[1643] Sorana Froda and Sévérien Nkurunziza. Prediction of predator–prey populations modelled by perturbed ODEs. Journal of Mathematical Biology, 54(3):407–451, March 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0051-9.pdf.

Giesl:2007:LFB

[1644] Peter Giesl and Heiko Wagner. Lyapunov function and the basin of attraction for a single-joint muscle-skeletal model. Journal of Mathematical Biology, 54(4):453–464, April 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-006-0052-8.pdf.

Krivan:2007:LFP

[1645] Vlastimil Krivan and Ivo Vrkoc. A Lyapunov function for piecewiseindependent differential equations: stability of the ideal free distribution in two patch environments. Journal of Mathematical Biology, 54(4): 465–488, April 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0053-7.pdf.

Calsina:2007:ASE

[1646] Àngel Calsina and Sílvia Cuadrado. Asymptotic stability of equilibria of selection–mutation equations. Journal of Mathematical Biology, 54(4): 489–511, April 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0056-4.pdf.

Tsai:2007:BBU

[1647] Je-Chiang Tsai and James Sneyd. Are buffers boring? uniqueness and asymptotical stability of traveling wave fronts in the buffered bistable system. Journal of Mathematical Biology, 54(4):513–553, April 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0057-3.pdf.

Buffoni:2007:SPD

[1648] Giuseppe Buffoni and Sara Pasquali. Structured population dynamics: continuous size and discontinuous stage structures. Journal of Mathematical Biology, 54(4):555–595, April 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-006-0058-2.pdf.

Baker:2007:MMC

[1649] R. E. Baker and P. K. Maini. A mechanism for morphogen-controlled domain growth. Journal of Mathematical Biology, 54(5):597–622, May 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-006-0060-8.pdf.

Buffoni:2007:MPP

[1650] Giuseppe Buffoni, Maria Paola Cassinari, and Maria Groppi. Modelling of predator–prey trophic interactions. Part II: Three trophic levels. Journal of Mathematical Biology, 54(5):623–644, May 2007. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0061-7.pdf.

Panaretos:2007:POB

[1651] Victor M. Panaretos. Partially observed branching processes for stochastic epidemics. Journal of Mathematical Biology, 54(5):645– 668, May 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0062-6.pdf.

Mukandavire:2007:SSH

[1652] Z. Mukandavire and W. Garira. Sex-structured HIV/AIDS model to analyse the effects of condom use with application to Zimbabwe. Journal of Mathematical Biology, 54(5):669–699, May 2007. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-006-0063-5.pdf.

Breda:2007:SAA

[1653] Dimitri Breda, Caterina Cusulin, Mimmo Iannelli, Stefano Maset, and Rossana Vermiglio. Stability analysis of age-structured population equations by pseudospectral differencing methods. Journal of Mathematical Biology, 54(5):701–720, May 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-006-0064-4.pdf.

Lessard:2007:PFS

[1654] Sabin Lessard and Véronique Ladret. The probability of fixation of a single mutant in an exchangeable selection model. Journal of Mathematical Biology, 54(5):721–744, May 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0069-7.pdf.

Varea:2007:SBB

[1655] C. Varea, D. Hernández, and R. A. Barrio. Soliton behaviour in a bistable reaction diffusion model. Journal of Mathematical Biology, 54(6):797– 813, June 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0071-0.pdf.

Neumann:2007:CMR

[1656] Gunter Neumann and Stefan Schuster. Continuous model for the rock– scissors–paper game between bacteriocin producing bacteria. Journal of Mathematical Biology, 54(6):815–846, June 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-006-0065-3.pdf.

Erban:2007:TEA

[1657] Radek Erban and Hans G. Othmer. Taxis equations for amoeboid cells. Journal of Mathematical Biology, 54(6):847–885, June 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0070-1.pdf.

Herrero:2007:RMB

[1658] Miguel A. Herrero. On the role of mathematics in biology. Journal of Mathematical Biology, 54(6):887–889, June 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-007-0095-5.pdf.

Nagylaki:2007:EML

[1659] Thomas Nagylaki and Yuan Lou. Evolution at a multiallelic locus under migration and uniform selection. Journal of Mathematical Biology, 54 (6):??, June 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0077-7.pdf.

Wang:2007:EPM

[1660] Guanyu Wang. Estimation of the proliferation and maturation functions in a physiologically structured model of thymocyte development. Journal of Mathematical Biology, 54(6):??, June 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-006-0068-0.pdf.

Yamada:2007:DRO

[1661] H. Yamada, T. Nakagaki, R. E. Baker, and P. K. Maini. Dispersion relation in oscillatory reaction–diffusion systems with self-consistent flow in true slime mold. Journal of Mathematical Biology, 54(6): ??, June 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-006-0067-1.pdf.

Shaw:2007:UDR

[1662] Leah B. Shaw, Lora Billings, and Ira B. Schwartz. Using dimension reduction to improve outbreak predictability of multistrain diseases. Journal of Mathematical Biology, 55(1):1–19, July 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0074-x.pdf.

Maquet:2007:GMC

[1663] J. Maquet, C. Letellier, and Luis A. Aguirre. Global models from the Canadian lynx cycles as a direct evidence for chaos in real ecosystems. Journal of Mathematical Biology, 55(1):21–39, July 2007. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0075-9.pdf.

Locsei:2007:PDI

[1664] J. T. Locsei. Persistence of direction increases the drift velocity of run and tumble chemotaxis. Journal of Mathematical Biology, 55(1): 41–60, July 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0080-z.pdf.

Mincheva:2007:GTMa

[1665] Maya Mincheva and Marc R. Roussel. Graph-theoretic methods for the analysis of chemical and biochemical networks. I. Multistability and oscillations in ordinary differential equation models. Journal of Mathematical Biology, 55(1):61–86, July 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0099-1.pdf.

Mincheva:2007:GTMb

[1666] Maya Mincheva and Marc R. Roussel. Graph-theoretic methods for the analysis of chemical and biochemical networks. II. Oscillations in networks with delays. Journal of Mathematical Biology, 55(1):87– 104, July 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0098-2.pdf.

Nevai:2007:PIC

[1667] Andrew L. Nevai and Richard R. Vance. Plant interspecies competition for sunlight: a mathematical model of canopy partitioning. Journal of Mathematical Biology, 55(1):105–145, July 2007. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0073-y.pdf.

Roques:2007:SPD

[1668] Lionel Roques and Radu S. Stoica. Species persistence decreases with habitat fragmentation: an analysis in periodic stochastic environments. Journal of Mathematical Biology, 55(2):189–205, August 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0076-8.pdf.

Weinberger:2007:ASS

[1669] Hans F. Weinberger, Mark A. Lewis, and Bingtuan Li. Anomalous spreading speeds of cooperative recursion systems. Journal of Mathematical Biology, 55(2):207–222, August 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0078-6.pdf.

Boys:2007:BIS

[1670] Richard J. Boys and Philip R. Giles. Bayesian inference for stochastic epidemic models with time-inhomogeneous removal rates. Journal of Mathematical Biology, 55(2):223–247, August 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0081-y.pdf.

Gander:2007:SGE

[1671] Martin J. Gander, Christian Mazza, and Hansklaus Rummler. Stochastic gene expression in switching environments. Journal of Mathematical Biology, 55(2):249–269, August 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0083-9.pdf.

Edwards:2007:PPL

[1672] R. Edwards, P. van den Driessche, and Lin Wang. Periodicity in piecewise-linear switching networks with delay. Journal of Mathematical Biology, 55(2):271–298, August 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0084-8.pdf.

Champagnat:2007:IAE

[1673] Nicolas Champagnat and Sylvie Méléard. Invasion and adaptive evolution for individual-based spatially structured populations. Journal of Mathematical Biology, 55(2):??, August 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0072-z.pdf.

Bold:2007:EFA

[1674] Katherine A. Bold, Yu Zou, Ioannis G. Kevrekidis, and Michael A. Henson. An equation-free approach to analyzing heterogeneous cell population dynamics. Journal of Mathematical Biology, 55(3):331–352, September 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0086-6.pdf.

Melodelima:2007:MAP

[1675] Christelle Melodelima, Christian Gautier, and Didier Piau. A Markovian approach for the prediction of mouse isochores. Journal of Mathematical Biology, 55(3):353–364, September 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0087-5.pdf.

Pearce:2007:CIS

[1676] Ian G. Pearce, Mark A. J. Chaplain, Pietà G. Schofield, Alexander R. A. Anderson, and Stephen F. Hubbard. Chemotaxis-induced spatiotemporal heterogeneity in multi-species host-parasitoid systems. Journal of Mathematical Biology, 55(3):365–388, September 2007. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0088-4.pdf.

Keeling:2007:RTT

[1677] Stephen L. Keeling, Roland Bammer, and Rudolf Stollberger. Revision of the theory of tracer transport and the convolution model of dynamic contrast enhanced magnetic resonance imaging. Journal of Mathematical Biology, 55(3):389–411, September 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0089-3.pdf.

Innocentini:2007:MSG

[1678] G. C. P. Innocentini and J. E. M. Hornos. Modeling stochastic gene expression under repression. Journal of Mathematical Biology, 55(3):413– 431, September 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0090-x.pdf.

Cain:2007:CSR

[1679] John W. Cain. Criterion for stable reentry in a ring of cardiac tissue. Journal of Mathematical Biology, 55(3):433–448, September 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-007-0100-z.pdf.

Pfaffelhuber:2007:AGP

[1680] P. Pfaffelhuber and A. Studeny. Approximating genealogies for partially linked neutral loci under a selective sweep. Journal of Mathematical Biology, 55(3):??, September 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0085-7.pdf.

Lemon:2007:TWB

[1681] G. Lemon and J. R. King. Travelling-wave behaviour in a multiphase model of a population of cells in an artificial scaffold. Journal of Mathematical Biology, 55(4):449–480, October 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0091-9.pdf.

Li:2007:GDM

[1682] Bingtuan Li and Hal L. Smith. Global dynamics of microbial competition for two resources with internal storage. Journal of Mathematical Biology, 55(4):481–515, October 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0092-8.pdf.

Edwards:2007:SHE

[1683] David A. Edwards. Steric hindrance effects on surface reactions: applications to BIAcore. Journal of Mathematical Biology, 55(4):517– 539, October 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0093-7.pdf.

Jones:2007:ERP

[1684] Laura E. Jones and Stephen P. Ellner. Effects of rapid prey evolution on predator–prey cycles. Journal of Mathematical Biology, 55(4):541– 573, October 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0094-6.pdf.

Trewenack:2007:DST

[1685] Abbey J. Trewenack, Kerry A. Landman, and Ben D. Bell. Dispersal and settling of translocated populations: a general study and a New Zealand amphibian case study. Journal of Mathematical Biology, 55(4):575– 604, October 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0096-4.pdf.

Kon:2007:SCO

[1686] Ryusuke Kon and Yoh Iwasa. Single-class orbits in nonlinear Leslie matrix models for semelparous populations. Journal of Mathematical Biology, 55(5–6):781–802, November 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0111-9.pdf.

Villacorta:2007:MFD

[1687] José A. Villacorta, Jorge Castro, Pilar Negredo, and Carlos Avendaño. Mathematical foundations of the dendritic growth models. Journal of Mathematical Biology, 55(5–6):817–859, November 2007. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0113-7.pdf.

Beder:2007:OSF

[1688] Jay H. Beder and Richard Gomulkiewicz. Optimizing selection for function-valued traits. Journal of Mathematical Biology, 55(5–6):??, November 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0114-6.pdf.

Boczko:2007:DSR

[1689] Erik Boczko, Tomás Gedeon, and Konstantin Mischaikow. Dynamics of a simple regulatory switch. Journal of Mathematical Biology, 55(5–6): ??, November 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0102-x.pdf.

English:2007:RMA

[1690] Anthony E. English, Conrad P. Plaut, and Alan B. Moy. A Riemannian manifold analysis of endothelial cell monolayer impedance parameter precision. Journal of Mathematical Biology, 55(5–6):??, November 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0103-9.pdf.

French:2007:PAS

[1691] Donald A. French and David A. Edwards. Perturbation approximation of solutions of a nonlinear inverse problem arising in olfaction experimentation. Journal of Mathematical Biology, 55(5–6):??, November 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0104-8.pdf.

Rikvold:2007:SOC

[1692] Per Arne Rikvold. Self-optimization, community stability, and fluctuations in two individual-based models of biological coevolution. Journal of Mathematical Biology, 55(5–6):??, November 2007. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0101-y.pdf.

Roberts:2007:MCE

[1693] M. G. Roberts and J. A. P. Heesterbeek. Model-consistent estimation of the basic reproduction number from the incidence of an emerging infection. Journal of Mathematical Biology, 55(5–6):??, November 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0112-8.pdf.

Smallbone:2007:QMC

[1694] Kieran Smallbone, David J. Gavaghan, Philip K. Maini, and J. Michael Brady. Quiescence as a mechanism for cyclical hypoxia and acidosis. Journal of Mathematical Biology, 55(5–6):??, November 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0105-7.pdf.

Tsai:2007:TWD

[1695] Je-Chiang Tsai and James Sneyd. Traveling waves in the discrete fast buffered bistable system. Journal of Mathematical Biology, 55(5–6):??, November 2007. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0097-3.pdf.

Anonymous:2008:JSI

[1696] Anonymous. JMB special issues on computational biology. Journal of Mathematical Biology, 56(1–2):1, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0144-0.pdf.

Clote:2008:ISI

[1697] Peter Clote. Introduction to special issue on RNA. Journal of Mathematical Biology, 56(1–2):3–13, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0130-6.pdf.

Machado-Lima:2008:CMN

[1698] Ariane Machado-Lima, Hernando A. del Portillo, and Alan Mitchell Durham. Computational methods in noncoding RNA research. Journal of Mathematical Biology, 56(1–2):15–49, January 2008. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0122-6.pdf.

Lladser:2008:MPM

[1699] Manuel E. Lladser, M. D. Betterton, and Rob Knight. Multiple pattern matching: a Markov chain approach. Journal of Mathematical Biology, 56(1–2):51–92, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0109-3.pdf.

Chan:2008:BEF

[1700] Chi Yu Chan and Ye Ding. Boltzmann ensemble features of RNA secondary structures: a comparative analysis of biological RNA sequences and random shuffles. Journal of Mathematical Biology, 56(1–2):93– 105, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0129-z.pdf.

Ponty:2008:ESR

[1701] Yann Ponty. Efficient sampling of RNA secondary structures from the Boltzmann ensemble of low-energy. Journal of Mathematical Biology, 56(1–2):107–127, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0137-z.pdf.

Bompfunewerer:2008:VRF

[1702] Athanasius F. Bompfünewerer, Rolf Backofen, Stephan H. Bernhart, Jana Hertel, Ivo L. Hofacker, Peter F. Stadler, and Sebastian Will. Variations on RNA folding and alignment: lessons from Benasque. Journal of Mathematical Biology, 56(1–2):129–144, January 2008. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0107-5.pdf.

Zhao:2008:RIP

[1703] Jizhen Zhao, Russell L. Malmberg, and Liming Cai. Rapid ab initio prediction of RNA pseudoknots via graph tree decomposition. Journal of Mathematical Biology, 56(1–2):145–159, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0124-4.pdf.

Metzler:2008:PRS

[1704] Dirk Metzler and Markus E. Nebel. Predicting RNA secondary structures with pseudoknots by MCMC sampling. Journal of Mathematical Biology, 56(1–2):161–181, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0106-6.pdf.

Tjaden:2008:PSN

[1705] Brian Tjaden. Prediction of small, noncoding RNAs in bacteria using heterogeneous data. Journal of Mathematical Biology, 56(1–2):183– 200, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0079-5.pdf.

Sato:2008:PPS

[1706] Kengo Sato, Kensuke Morita, and Yasubumi Sakakibara. PSSMTS: position specific scoring matrices on tree structures. Journal of Mathematical Biology, 56(1–2):201–214, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0108-4.pdf.

Sarver:2008:FFL

[1707] Michael Sarver, Craig L. Zirbel, Jesse Stombaugh, Ali Mokdad, and Neocles B. Leontis. FR3D: finding local and composite recurrent structural motifs in RNA 3D structures. Journal of Mathematical Biology, 56(1–2):215–252, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0110-x.pdf.

Wang:2008:RFK

[1708] Xueyi Wang, Gary Kapral, Laura Murray, David Richardson, Jane Richardson, and Jack Snoeyink. RNABC: forward kinematics to reduce all-atom steric clashes in RNA backbone. Journal of Mathematical Biology, 56(1–2):253–278, January 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0082-x.pdf.

Aase:2008:SAO

[1709] Sven Ole Aase and Peter Ruoff. Semi-algebraic optimization of temperature compensation in a general switch-type negative feedback model of circadian clocks. Journal of Mathematical Biology, 56(3):279–292, March 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0115-5.pdf.

Volz:2008:SDR

[1710] Erik Volz. SIR dynamics in random networks with heterogeneous connectivity. Journal of Mathematical Biology, 56(3):293–310, March 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-007-0116-4.pdf. See note [1924].

Cessac:2008:DTN

[1711] B. Cessac. A discrete time neural network model with spiking neurons. Journal of Mathematical Biology, 56(3):311–345, March 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0117-3.pdf.

Wang:2008:MSM

[1712] Xiaoqiang Wang and Qiang Du. Modelling and simulations of multicomponent lipid membranes and open membranes via diffuse interface approaches. Journal of Mathematical Biology, 56(3):347–371, March 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0118-2.pdf.

Jiang:2008:CCD

[1713] Jifa Jiang and Caichun Chai. The complete classification for dynamics in a homosexually-transmitted disease model. Journal of Mathematical Biology, 56(3):373–390, March 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0119-1.pdf.

Minin:2008:CLT

[1714] Vladimir N. Minin and Marc A. Suchard. Counting labeled transitions in continuous-time Markov models of evolution. Journal of Mathematical Biology, 56(3):391–412, March 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0120-8.pdf.

Chen:2008:EDP

[1715] Christina T. L. Chen, Quo-Shin Chi, and Stanley A. Sawyer. Effects of dominance on the probability of fixation of a mutant allele. Journal of Mathematical Biology, 56(3):413–434, March 2008. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0121-7.pdf.

Thul:2008:BTM

[1716] R. Thul, G. D. Smith, and S. Coombes. A bidomain threshold model of propagating calcium waves. Journal of Mathematical Biology, 56(4): 435–463, April 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0123-5.pdf.

Grunewald:2008:EPT

[1717] Stefan Grünewald, Katharina T. Huber, Vincent Moulton, and Charles Semple. Encoding phylogenetic trees in terms of weighted quartets. Journal of Mathematical Biology, 56(4):465–477, April 2008. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0125-3.pdf.

Haake:2008:SVP

[1718] Claus-Jochen Haake, Akemi Kashiwada, and Francis Edward Su. The Shapley value of phylogenetic trees. Journal of Mathematical Biology, 56 (4):479–497, April 2008. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-007-0126-2.pdf.

Lutscher:2008:DDD

[1719] Frithjof Lutscher. Density-dependent dispersal in integrodifference equations. Journal of Mathematical Biology, 56(4):499–524, April 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-007-0127-1.pdf.

Aldous:2008:SMP

[1720] David Aldous, Maxim Krikun, and Lea Popovic. Stochastic models for phylogenetic trees on higher-order taxa. Journal of Mathematical Biology, 56(4):525–557, April 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0128-0.pdf.

Bielecki:2008:MNF

[1721] Andrzej Bielecki and Piotr Kalita. Model of neurotransmitter fast transport in axon terminal of presynaptic neuron. Journal of Mathematical Biology, 56(4):559–576, April 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0131-5.pdf.

Martcheva:2008:KDE

[1722] Maia Martcheva, Horst R. Thieme, and Thanate Dhirasakdanon. Kolmogorov’s differential equations and positive semigroups on first moment sequence spaces. Journal of Mathematical Biology, 56(4):577– 578, April 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0154-y.pdf.

Kettemann:2008:DAS

[1723] Anita Kettemann and Maria Neuss-Radu. Derivation and analysis of a system modeling the chemotactic movement of hematopoietic stem cells. Journal of Mathematical Biology, 56(5):579–610, May 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0132-4.pdf.

Chen:2008:EEE

[1724] Chu-Chih Chen, Meng-Chiuan Shih, Kuen-Yuh Wu, and Pranab K. Sen. Exterior exposure estimation using a one-compartment toxicokinetic model with blood sample measurements. Journal of Mathematical Biology, 56(5):611–633, May 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0133-3.pdf.

Boldin:2008:SCI

[1725] Barbara Boldin and Odo Diekmann. Superinfections can induce evolutionarily stable coexistence of pathogens. Journal of Mathematical Biology, 56(5):635–672, May 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0135-1.pdf.

Durinx:2008:ADP

[1726] Michel Durinx, J. A. J. (Hans) Metz, and Géza Meszéna. Adaptive dynamics for physiologically structured population models. Journal of Mathematical Biology, 56(5):673–742, May 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-007-0134-2.pdf.

Stone:2008:SMH

[1727] Patricia Stone, Hilde Wilkinson-Herbots, and Valerie Isham. A stochastic model for head lice infections. Journal of Mathematical Biology, 56(6): 743–763, June 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0136-0.pdf.

Collet:2008:AVO

[1728] P. Collet and S. Martínez. Asymptotic velocity of one dimensional diffusions with periodic drift. Journal of Mathematical Biology, 56(6): 765–792, June 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0138-y.pdf.

Hogea:2008:IDP

[1729] Cosmina Hogea, Christos Davatzikos, and George Biros. An image-driven parameter estimation problem for a reaction–diffusion glioma growth model with mass effects. Journal of Mathematical Biology, 56(6):793– 825, June 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0139-x.pdf.

Olinky:2008:SDT

[1730] Ronen Olinky, Amit Huppert, and Lewi Stone. Seasonal dynamics and thresholds governing recurrent epidemics. Journal of Mathematical Biology, 56(6):827–839, June 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-007-0140-4.pdf.

Chen:2008:ACT

[1731] Zhi-Min Chen and W. G. Price. An analysis of the coexistence of two host species with a shared pathogen. Journal of Mathematical Biology, 56(6): 841–859, June 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0141-3.pdf.

Subramanian:2008:DEV

[1732] Vijay G. Subramanian, Ken R. Duffy, Marian L. Turner, and Philip D. Hodgkin. Determining the expected variability of immune responses using the cyton model. Journal of Mathematical Biology, 56(6):861– 892, June 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0142-2.pdf.

Brook:2008:MCI

[1733] B. S. Brook and S. L. Waters. Mathematical challenges in integrative physiology. Journal of Mathematical Biology, 56(6):893– 896, June 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0159-1.pdf.

Franceschetti:2008:TBS

[1734] Andrea Franceschetti and Andrea Pugliese. Threshold behaviour of a SIR epidemic model with age structure and immigration. Journal of Mathematical Biology, 57(1):1–27, July 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0143-1.pdf.

Wang:2008:NSS

[1735] Liming Wang and Eduardo D. Sontag. On the number of steady states in a multiple futile cycle. Journal of Mathematical Biology, 57(1): 29–52, July 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0145-z.pdf.

Wei:2008:SMS

[1736] Juncheng Wei and Matthias Winter. Stationary multiple spots for reaction–diffusion systems. Journal of Mathematical Biology, 57(1): 53–89, July 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0146-y.pdf.

Brikci:2008:ACS

[1737] Fadia Bekkal Brikci, Jean Clairambault, Benjamin Ribba, and Benoît Perthame. An age-and-cyclin-structured cell population model for healthy and tumoral tissues. Journal of Mathematical Biology, 57(1): 91–110, July 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0147-x.pdf.

Dudkowska:2008:SHM

[1738] Aleksandra Dudkowska and Danuta Makowiec. Seidel-herzel model of human baroreflex in cardiorespiratory system with stochastic delays. Journal of Mathematical Biology, 57(1):111–137, July 2008. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0148-9.pdf.

Barnett:2008:ASS

[1739] Alex H. Barnett and Paul R. Moorcroft. Analytic steady-state space use patterns and rapid computations in mechanistic home range analysis. Journal of Mathematical Biology, 57(1):139–159, July 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0149-8.pdf.

Mouser:2008:MPC

[1740] Christina Mouser, Farzan Nadim, and Amitabha Bose. Maintaining phase of the crustacean tri-phasic pyloric rhythm. Journal of Mathematical Biology, 57(2):161–181, August 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-007-0150-2.pdf.

Futakata:2008:FAR

[1741] Y. Futakata and T. Iwasaki. Formal analysis of resonance entrainment by central pattern generator. Journal of Mathematical Biology, 57(2): 183–207, August 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0151-1.pdf.

Adams:2008:GDD

[1742] Michael J. Adams. Graph decompositions for demographic loop analysis. Journal of Mathematical Biology, 57(2):209–221, August 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-007-0152-0.pdf.

Ruan:2008:CHB

[1743] Shigui Ruan, Yilei Tang, and Weinian Zhang. Computing the heteroclinic bifurcation curves in predator–prey systems with ratio-dependent functional response. Journal of Mathematical Biology, 57(2):223–241, August 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-007-0153-z.pdf.

VanCleve:2008:SLP

[1744] Jeremy Van Cleve and Marcus W. Feldman. Stable long-period cycling and complex dynamics in a single-locus fertility model with genomic imprinting. Journal of Mathematical Biology, 57(2):243–264, August 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0156-4.pdf.

Copene:2008:ECC

[1745] Elizabeth D. Copene and James P. Keener. Ephaptic coupling of cardiac cells through the junctional electric potential. Journal of Mathematical Biology, 57(2):265–284, August 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0157-3.pdf.

Gyllenberg:2008:ECD

[1746] Mats Gyllenberg, Éva Kisdi, and Margarete Utz. Evolution of conditiondependent dispersal under kin competition. Journal of Mathematical Biology, 57(2):285–307, August 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0158-2.pdf.

Smith:2008:DDM

[1747] H. L. Smith. The discrete dynamics of monotonically decomposable maps. Journal of Mathematical Biology, 57(2):309–310, August 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-008-0155-5.pdf.

Sharkey:2008:DEM

[1748] Kieran J. Sharkey. Deterministic epidemiological models at the individual level. Journal of Mathematical Biology, 57(3):311–331, September 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-008-0161-7.pdf.

Weller:2008:PDN

[1749] Frédéric Frank Weller. Platelet deposition in non-parallel flow. Journal of Mathematical Biology, 57(3):333–359, September 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0163-5.pdf.

Chen:2008:ECD

[1750] Xinfu Chen, Richard Hambrock, and Yuan Lou. Evolution of conditional dispersal: a reaction–diffusion–advection model. Journal of Mathematical Biology, 57(3):361–386, September 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-008-0166-2.pdf.

Weinberger:2008:SSS

[1751] Hans F. Weinberger, Kohkichi Kawasaki, and Nanako Shigesada. Spreading speeds of spatially periodic integro-difference models for populations with nonmonotone recruitment functions. Journal of Mathematical Biology, 57(3):387–411, September 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0168-0.pdf.

Braverman:2008:CVP

[1752] Elena Braverman and Reneeta Mamdani. Continuous versus pulse harvesting for population models in constant and variable environment. Journal of Mathematical Biology, 57(3):413–434, September 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-008-0169-z.pdf.

Birkner:2008:CLC

[1753] Matthias Birkner and Jochen Blath. Computing likelihoods for coalescents with multiple collisions in the infinitely many sites model. Journal of Mathematical Biology, 57(3):435–465, September 2008. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0170-6.pdf.

Pindera:2008:CEM

[1754] Maciej Z. Pindera, Hui Ding, and Zhijian Chen. Convected element method for simulation of angiogenesis. Journal of Mathematical Biology, 57(4):467–495, October 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0171-5.pdf.

Dong:2008:MCS

[1755] Fang Dong and William F. Langford. Models of cheyne–Stokes respiration with cardiovascular pathologies. Journal of Mathematical Biology, 57(4):497–519, October 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0173-3.pdf.

Ladau:2008:RHT

[1756] Joshua Ladau and Steven J. Schwager. Robust hypothesis tests for independence in community assembly. Journal of Mathematical Biology, 57(4):537–555, October 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0176-0.pdf.

Artiles:2008:PSI

[1757] W. Artiles, P. G. S. Carvalho, and R. A. Kraenkel. Patch-size and isolation effects in the Fisher–Kolmogorov equation. Journal of Mathematical Biology, 57(4):??, October 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0174-2.pdf.

Bacaer:2008:MJE

[1758] Nicolas Bacaër, Rachid Ouifki, Carel Pretorius, Robin Wood, and Brian Williams. Modeling the joint epidemics of TB and HIV in a south African township. Journal of Mathematical Biology, 57(4): ??, October 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0177-z.pdf.

Frazer:2008:SLI

[1759] L. Neil Frazer. Sea-lice infection models for fishes. Journal of Mathematical Biology, 57(4):??, October 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0181-3.pdf.

Schlicht:2008:DSP

[1760] Robert Schlicht and Gerhard Winkler. A delay stochastic process with applications in molecular biology. Journal of Mathematical Biology, 57 (5):613–648, November 2008. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-008-0178-y.pdf.

Ryan:2008:MCT

[1761] Paul Ryan, Rabia Djellouli, and Randy Cohen. Modeling capsule tissue growth around disk-shaped implants: a numerical and in vivo study. Journal of Mathematical Biology, 57(5):675–695, November 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-008-0184-0.pdf.

Hadeler:2008:MDS

[1762] K. P. Hadeler and H. R. Thieme. Monotone dependence of the spectral bound on the transition rates in linear compartment models. Journal of Mathematical Biology, 57(5):697–712, November 2008. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0185-z.pdf.

Gernhard:2008:SPG

[1763] Tanja Gernhard, Klaas Hartmann, and Mike Steel. Stochastic properties of generalised Yule models, with biodiversity applications. Journal of Mathematical Biology, 57(5):713–735, November 2008. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0186-y.pdf.

Bacaer:2008:RET

[1764] Nicolas Bacaër and Xamxinur Abdurahman. Resonance of the epidemic threshold in a periodic environment. Journal of Mathematical Biology, 57 (5):??, November 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0183-1.pdf.

Gallegos:2008:MMC

[1765] Angela Gallegos, Tenecia Plummer, David Uminsky, Cinthia Vega, Clare Wickman, and Michael Zawoiski. A mathematical model of a crocodilian population using delay-differential equations. Journal of Mathematical Biology, 57(5):??, November 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0187-x.pdf.

Franke:2008:DIM

[1766] John E. Franke and Abdul-Aziz Yakubu. Disease-induced mortality in density-dependent discrete-time S-I-S epidemic models. Journal of Mathematical Biology, 57(6):755–790, December 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-008-0188-9.pdf.

Bennett:2008:BMC

[1767] Rachel Bennett and Roger G. Bowers. A baseline model for the coevolution of hosts and pathogens. Journal of Mathematical Biology, 57 (6):791–809, December 2008. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-008-0189-8.pdf.

Zint:2008:HCU

[1768] Natali Zint, Ellen Baake, and Frank den Hollander. How T-cells use large deviations to recognize foreign antigens. Journal of Mathematical Biology, 57(6):841–861, December 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0191-1.pdf.

Hofbauer:2008:QPL

[1769] Josef Hofbauer, Ryusuke Kon, and Yasuhisa Saito. Qualitative permanence of Lotka–Volterra equations. Journal of Mathematical Biology, 57(6):863–881, December 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0192-0.pdf.

Marshall:2008:BPE

[1770] John M. Marshall. A branching process for the early spread of a transposable element in a diploid population. Journal of Mathematical Biology, 57 (6):??, December 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0190-2.pdf.

Pickard:2008:MSA

[1771] William F. Pickard. Modelling the swelling assay for aquaporin expression. Journal of Mathematical Biology, 57(6):??, December 2008. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-008-0196-9.pdf.

Capasso:2009:SMT

[1772] Vincenzo Capasso and Daniela Morale. Stochastic modelling of tumourinduced angiogenesis. Journal of Mathematical Biology, 58(1–2):219– 233, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0193-z.pdf.

Foley:2009:DHD

[1773] Catherine Foley and Michael C. Mackey. Dynamic hematological disease: a review. Journal of Mathematical Biology, 58(1–2):285–322, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0165-3.pdf.

Alt:2009:CCB

[1774] Wolfgang Alt, Andreas Deutsch, and Luigi Preziosi. Computational cell biology: Second theme issue on “Computational biology”. Journal of Mathematical Biology, 58(1–2):??, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-008-0207-x.pdf.

Ambrosi:2009:TPT

[1775] D. Ambrosi, A. Duperray, V. Peschetola, and C. Verdier. Traction patterns of tumor cells. Journal of Mathematical Biology, 58(1–2): ??, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0167-1.pdf.

Dickinson:2009:MAP

[1776] Richard B. Dickinson. Models for actin polymerization motors. Journal of Mathematical Biology, 58(1–2):??, January 2009. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0200-4.pdf.

Galle:2009:SCT

[1777] J. Galle, M. Hoffmann, and G. Aust. From single cells to tissue architecture — a bottom-up approach to modelling the spatio-temporal organisation of complex multi-cellular systems. Journal of Mathematical Biology, 58(1–2):??, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0172-4.pdf.

Hillen:2009:UGP

[1778] T. Hillen and K. J. Painter. A user’s guide to PDE models for chemotaxis. Journal of Mathematical Biology, 58(1–2):??, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0201-3.pdf.

Kalaidzidis:2009:MOT

[1779] Yannis Kalaidzidis. Multiple objects tracking in fluorescence microscopy. Journal of Mathematical Biology, 58(1–2):??, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0180-4.pdf.

Kuusela:2009:CMC

[1780] Esa Kuusela and Wolfgang Alt. Continuum model of cell adhesion and migration. Journal of Mathematical Biology, 58(1–2):??, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0179-x.pdf.

Mogilner:2009:MCM

[1781] Alex Mogilner. Mathematics of cell motility: have we got its number? Journal of Mathematical Biology, 58(1–2):??, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0182-2.pdf.

Stelling:2009:SCC

[1782] Jörg Stelling and Boris N. Kholodenko. Signaling cascades as cellular devices for spatial computations. Journal of Mathematical Biology, 58(1– 2):??, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0162-6.pdf.

Verdier:2009:MCI

[1783] Claude Verdier, Cécile Couzon, Alain Duperray, and Pushpendra Singh. Modeling cell interactions under flow. Journal of Mathematical Biology, 58(1–2):??, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-008-0164-4.pdf.

Westerhoff:2009:SBT

[1784] Hans V. Westerhoff, Alexey Kolodkin, Riaan Conradie, Stephen J. Wilkinson, Frank J. Bruggeman, Klaas Krab, Jan H. van Schuppen, Hanna Hardin, Barbara M. Bakker, Martijn J. Moné, Katja N. Rybakova, Marco Eijken, Hans J. P. van Leeuwen, and Jacky L. Snoep. Systems biology towards life in silico: mathematics of the control of living cells. Journal of Mathematical Biology, 58(1–2):??, January 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0160-8.pdf.

Li:2009:ETW

[1785] Bingtuan Li, Mark A. Lewis, and Hans F. Weinberger. Existence of traveling waves for integral recursions with nonmonotone growth functions. Journal of Mathematical Biology, 58(3):323–338, March 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0175-1.pdf.

Allen:2009:SPD

[1786] L. J. S. Allen, Y. Lou, and A. L. Nevai. Spatial patterns in a discretetime SIS patch model. Journal of Mathematical Biology, 58(3):339– 375, March 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0194-y.pdf.

Kooijman:2009:WEC

[1787] S. A. L. M. Kooijman. What the egg can tell about its hen: Embryonic development on the basis of dynamic energy budgets. Journal of Mathematical Biology, 58(3):377–394, March 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-008-0195-x.pdf.

Gautrais:2009:AFM

[1788] Jacques Gautrais, Christian Jost, Marc Soria, Alexandre Campo, Sébastien Motsch, Richard Fournier, Stéphane Blanco, and Guy Theraulaz. Analyzing fish movement as a persistent turning walker. Journal of Mathematical Biology, 58(3):429–445, March 2009. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0198-7.pdf.

Keener:2009:IMR

[1789] James P. Keener. Invariant manifold reductions for Markovian ion channel dynamics. Journal of Mathematical Biology, 58(3):447–457, March 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0199-6.pdf.

Kuwamura:2009:MMP

[1790] Masataka Kuwamura, Takefumi Nakazawa, and Toshiyuki Ogawa. A minimum model of prey-predator system with dormancy of predators and the paradox of enrichment. Journal of Mathematical Biology, 58(3): 459–479, March 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0203-1.pdf.

Anguige:2009:ODM

[1791] K. Anguige and C. Schmeiser. A one-dimensional model of cell diffusion and aggregation, incorporating volume filling and cell-to-cell adhesion. Journal of Mathematical Biology, 58(3):??, March 2009. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0197-8.pdf.

Zhang:2009:MAB

[1792] Le Zhang, Zhihui Wang, Jonathan A. Sagotsky, and Thomas S. Deisboeck. Multiscale agent-based cancer modeling. Journal of Mathematical Biology, 58(4–5):545–559, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0211-1.pdf.

Macklin:2009:MMN

[1793] Paul Macklin, Steven McDougall, Alexander R. A. Anderson, Mark A. J. Chaplain, Vittorio Cristini, and John Lowengrub. Multiscale modelling and nonlinear simulation of vascular tumour growth. Journal of Mathematical Biology, 58(4–5):765–798, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-008-0216-9.pdf.

Anderson:2009:MDI

[1794] Alexander R. A. Anderson, Katarzyna A. Rejniak, Philip Gerlee, and Vito Quaranta. Microenvironment driven invasion: a multiscale multimodel investigation. Journal of Mathematical Biology, 58(4–5): ??, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0210-2.pdf.

Byrne:2009:IBC

[1795] Helen Byrne and Dirk Drasdo. Individual-based and continuum models of growing cell populations: a comparison. Journal of Mathematical Biology, 58(4–5):??, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0212-0.pdf.

Chaplain:2009:P

[1796] Mark Chaplain. Preface. Journal of Mathematical Biology, 58(4–5): ??, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0241-8.pdf.

Cristini:2009:NSS

[1797] Vittorio Cristini, Xiangrong Li, John S. Lowengrub, and Steven M. Wise. Nonlinear simulations of solid tumor growth using a mixture model: invasion and branching. Journal of Mathematical Biology, 58 (4–5):??, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0215-x.pdf.

ORourke:2009:LQT

[1798] S. F. C. O’Rourke, H. McAneney, and T. Hillen. Linear quadratic and tumour control probability modelling in external beam radiotherapy. Journal of Mathematical Biology, 58(4–5):??, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0222-y.pdf.

Owen:2009:AVR

[1799] Markus R. Owen, Tomás Alarcón, Philip K. Maini, and Helen M. Byrne. Angiogenesis and vascular remodelling in normal and cancerous tissues. Journal of Mathematical Biology, 58(4–5):??, April 2009. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0213-z.pdf.

Painter:2009:MCM

[1800] K. J. Painter. Modelling cell migration strategies in the extracellular matrix. Journal of Mathematical Biology, 58(4–5):??, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0217-8.pdf.

Preziosi:2009:MMT

[1801] Luigi Preziosi and Andrea Tosin. Multiphase modelling of tumour growth and extracellular matrix interaction: mathematical tools and applications. Journal of Mathematical Biology, 58(4–5):??, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0218-7.pdf.

Rockne:2009:MMB

[1802] R. Rockne, E. C. Alvord, Jr., J. K. Rockhill, and K. R. Swanson. A mathematical model for brain tumor response to radiation therapy. Journal of Mathematical Biology, 58(4–5):??, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-008-0219-6.pdf.

Sinek:2009:PDP

[1803] John P. Sinek, Sandeep Sanga, Xiaoming Zheng, Hermann B. Frieboes, Mauro Ferrari, and Vittorio Cristini. Predicting drug pharmacokinetics and effect in vascularized tumors using computer simulation. Journal of Mathematical Biology, 58(4–5):??, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-008-0214-y.pdf.

Szymanska:2009:MMI

[1804] Zuzanna Szymańska, Jakub Urbański, and Anna Marciniak-Czochra. Mathematical modelling of the influence of heat shock proteins on cancer invasion of tissue. Journal of Mathematical Biology, 58(4–5): ??, April 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0220-0.pdf.

Pearson:2009:MMC

[1805] T. Pearson, J. A. D. Wattis, B. O’Malley, L. Pickersgill, H. Blackburn, K. G. Jackson, and H. M. Byrne. Mathematical modelling of competitive LDL/VLDL binding and uptake by hepatocytes. Journal of Mathematical Biology, 58(6):845–880, June 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-008-0205-z.pdf.

Blyuss:2009:SBM

[1806] Konstantin B. Blyuss and Sunetra Gupta. Stability and bifurcations in a model of antigenic variation in malaria. Journal of Mathematical Biology, 58(6):??, June 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0204-0.pdf.

Burger:2009:MSSa

[1807] Reinhard Bürger. Multilocus selection in subdivided populations II. Maintenance of polymorphism under weak or strong migration. Journal of Mathematical Biology, 58(6):??, June 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-008-0237-4.pdf.

Burger:2009:MSSb

[1808] Reinhard Bürger. Multilocus selection in subdivided populations I. Convergence properties for weak or strong migration. Journal of Mathematical Biology, 58(6):??, June 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0236-5.pdf.

Meleard:2009:TSS

[1809] Sylvie Méléard and Viet Chi Tran. Trait substitution sequence process and canonical equation for age-structured populations. Journal of Mathematical Biology, 58(6):??, June 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0202-2.pdf.

Eftimie:2009:WNA

[1810] R. Eftimie, G. de Vries, and M. A. Lewis. Weakly nonlinear analysis of a hyperbolic model for animal group formation. Journal of Mathematical Biology, 59(1):37–74, July 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0209-8.pdf.

Kolokolnikov:2009:PDP

[1811] Theodore Kolokolnikov, Chunhua Ou, and Yuan Yuan. Phytoplankton depth profiles and their transitions near the critical sinking velocity. Journal of Mathematical Biology, 59(1):105–122, July 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0221-z.pdf.

Capasso:2009:MSG

[1812] Vincenzo Capasso. Multiple scales and geometric structures: additional sources of randomness. Journal of Mathematical Biology, 59(1): 143–146, July 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0272-9.pdf.

Cushing:2009:TSS

[1813] J. M. Cushing. Three stage semelparous Leslie models. Journal of Mathematical Biology, 59(1):??, July 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0208-9.pdf.

Gadgil:2009:WCP

[1814] Siddhartha Gadgil. Watson–Crick pairing, the Heisenberg group and Milnor invariants. Journal of Mathematical Biology, 59(1):??, July 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-008-0223-x.pdf.

Greenhalgh:2009:BBE

[1815] David Greenhalgh and Martin Griffiths. Backward bifurcation, equilibrium and stability phenomena in a three-stage extended BRSV epidemic model. Journal of Mathematical Biology, 59(1):??, July 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0206-y.pdf.

Fiedler:2009:NPC

[1816] Bernold Fiedler and Sze-Bi Hsu. Non-periodicity in chemostat equations: a multi-dimensional negative Bendixson–Dulac criterion. Journal of Mathematical Biology, 59(2):233–253, August 2009. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0229-4.pdf.

Duffy:2009:ICB

[1817] Ken R. Duffy and Vijay G. Subramanian. On the impact of correlation between collaterally consanguineous cells on lymphocyte population dynamics. Journal of Mathematical Biology, 59(2):255–285, August 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0231-x.pdf.

Bates:2009:GPD

[1818] P. W. Bates, Zhan Chen, Yuhui Sun, Guo-Wei Wei, and Shan Zhao. Geometric and potential driving formation and evolution of biomolecular surfaces. Journal of Mathematical Biology, 59(2):??, August 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0226-7.pdf.

Nykamp:2009:SDC

[1819] Duane Q. Nykamp. A stimulus-dependent connectivity analysis of neuronal networks. Journal of Mathematical Biology, 59(2):??, August 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0224-9.pdf.

Xiong:2009:SME

[1820] Xiaoping Xiong, James M. Boyett, Robert G. Webster, and Juergen Stech. A stochastic model for estimation of mutation rates in multiplereplication proliferation processes. Journal of Mathematical Biology, 59 (2):??, August 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0225-8.pdf.

Keef:2009:AEI

[1821] T. Keef and R. Twarock. Affine extensions of the icosahedral group with applications to the three-dimensional organisation of simple viruses. Journal of Mathematical Biology, 59(3):287–313, September 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-008-0228-5.pdf.

Chaves:2009:GTP

[1822] Madalena Chaves, Anirvan Sengupta, and Eduardo D. Sontag. Geometry and topology of parameter space: investigating measures of robustness in regulatory networks. Journal of Mathematical Biology, 59(3):315– 358, September 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0230-y.pdf.

Zhu:2009:STN

[1823] Junmei Zhu and Christoph von der Malsburg. Steps toward numerical mode analysis of organizing systems. Journal of Mathematical Biology, 59(3):359–376, September 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0233-8.pdf.

Ermentrout:2009:SSA

[1824] Bard Ermentrout and Stuart Hastings. Steady-state analysis of a continuum model for super-infection. Journal of Mathematical Biology, 59(3): 415–438, September 2009. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-008-0232-9.pdf.

Reiter:2009:EDG

[1825] Philipp Reiter, Dieter Felix, Heiko von der Mosel, and Wolfgang Alt. Energetics and dynamics of global integrals modeling interaction between stiff filaments. Journal of Mathematical Biology, 59(3): ??, September 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0227-6.pdf.

Frenod:2009:EMV

[1826] Emmanuel Frénod and Olivier Sire. An explanatory model to validate the way water activity rules periodic terrace generation in Proteus mirabilis swarm. Journal of Mathematical Biology, 59(4):439– 466, October 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0235-6.pdf.

Domijan:2009:BOC

[1827] Mirela Domijan and Markus Kirkilionis. Bistability and oscillations in chemical reaction networks. Journal of Mathematical Biology, 59(4):467– 501, October 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0234-7.pdf.

de-Camino-Beck:2009:GTM

[1828] Tomás de Camino-Beck, Mark A. Lewis, and P. van den Driessche. A graph-theoretic method for the basic reproduction number in continuous time epidemiological models. Journal of Mathematical Biology, 59(4):503–516, October 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0240-9.pdf.

Gerberry:2009:SMC

[1829] David J. Gerberry and Fabio A. Milner. An SEIQR model for childhood diseases. Journal of Mathematical Biology, 59(4):535–561, October 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0239-2.pdf.

DeLeenheer:2009:FAT

[1830] Patrick De Leenheer and N. G. Cogan. Failure of antibiotic treatment in microbial populations. Journal of Mathematical Biology, 59(4):563– 579, October 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0243-6.pdf.

Lalam:2009:QAP

[1831] Nadia Lalam. A quantitative approach for polymerase chain reactions based on a hidden Markov model. Journal of Mathematical Biology, 59 (4):??, October 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0238-3.pdf.

Javierre:2009:MAP

[1832] E. Javierre, F. J. Vermolen, C. Vuik, and S. van der Zwaag. A mathematical analysis of physiological and morphological aspects of wound closure. Journal of Mathematical Biology, 59(5):605–630, November 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-008-0242-7.pdf.

Broom:2009:GTM

[1833] Mark Broom and Jan Rychtár. A game theoretical model of kleptoparasitism with incomplete information. Journal of Mathematical Biology, 59(5):631–649, November 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0247-2.pdf.

Ediev:2009:DRV

[1834] Dalkhat M. Ediev. On the definition of the reproductive value: response to the discussion by Bacaër and abdurahman. Journal of Mathematical Biology, 59(5):651–657, November 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0246-3.pdf.

Lessard:2009:DAO

[1835] Sabin Lessard. Diffusion approximations for one-locus multi-allele kin selection, mutation and random drift in group-structured populations: a unifying approach to selection models in population genetics. Journal of Mathematical Biology, 59(5):659–696, November 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0248-1.pdf.

Smith:2009:CVR

[1836] Robert J. Smith and B. D. Aggarwala. Can the viral reservoir of latently infected CD4^{+} T cells be eradicated with antiretroviral HIV drugs? Journal of Mathematical Biology, 59(5):697–715, November 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-008-0245-4.pdf.

Geli:2009:MMP

[1837] Patricia Geli. Modeling the mechanism of postantibiotic effect and determining implications for dosing regimens. Journal of Mathematical Biology, 59(5):??, November 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0249-8.pdf.

Luzyanina:2009:DPI

[1838] Tatyana Luzyanina, Dirk Roose, and Gennady Bocharov. Distributed parameter identification for a label-structured cell population dynamics model using CFSE histogram time-series data. Journal of Mathematical Biology, 59(5):??, November 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-008-0244-5.pdf.

Linninger:2009:MMB

[1839] Andreas A. Linninger, Michalis Xenos, Brian Sweetman, Sukruti Ponkshe, Xiaodong Guo, and Richard Penn. A mathematical model of blood, cerebrospinal fluid and brain dynamics. Journal of Mathematical Biology, 59(6):729–759, December 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-009-0250-2.pdf.

Band:2009:MCA

[1840] L. R. Band, L. J. Cummings, S. L. Waters, and J. A. D. Wattis. Modelling crystal aggregation and deposition in the catheterised lower urinary tract. Journal of Mathematical Biology, 59(6):809–840, December 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0253-z.pdf.

Taylor:2009:SEM

[1841] Michael L. Taylor and Thomas W. Carr. An SIR epidemic model with partial temporary immunity modeled with delay. Journal of Mathematical Biology, 59(6):841–880, December 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-009-0256-9.pdf.

Langlands:2009:FCE

[1842] T. A. M. Langlands, B. I. Henry, and S. L. Wearne. Fractional cable equation models for anomalous electrodiffusion in nerve cells: infinite domain solutions. Journal of Mathematical Biology, 59(6):??, December 2009. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0251-1.pdf.

Levine:2010:DRC

[1843] Howard A. Levine, Maria P. McGee, and Susana Serna. Diffusion and reaction in the cell glycocalyx and the extracellular matrix. Journal of Mathematical Biology, 60(1):1–26, January 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-009-0254-y.pdf.

Tang:2010:MFT

[1844] Moxun Tang. The mean frequency of transcriptional bursting and its variation in single cells. Journal of Mathematical Biology, 60(1):27– 58, January 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0258-7.pdf.

Hasik:2010:PPS

[1845] Karel Hasík. On a predator–prey system of Gause type. Journal of Mathematical Biology, 60(1):59–74, January 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-009-0257-8.pdf.

Pozrikidis:2010:NSB

[1846] C. Pozrikidis. Numerical simulation of blood and interstitial flow through a solid tumor. Journal of Mathematical Biology, 60(1):75– 94, January 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0259-6.pdf.

Liu:2010:MTC

[1847] Liang Liu, Lili Yu, and Dennis K. Pearl. Maximum tree: a consistent estimator of the species tree. Journal of Mathematical Biology, 60(1): 95–106, January 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0260-0.pdf.

Karev:2010:MTS

[1848] Georgiy P. Karev. On mathematical theory of selection: continuous time population dynamics. Journal of Mathematical Biology, 60(1):107– 129, January 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0252-0.pdf.

Mirams:2010:MTA

[1849] Gary R. Mirams, Helen M. Byrne, and John R. King. A multiple timescale analysis of a mathematical model of the Wnt/β-catenin signalling pathway. Journal of Mathematical Biology, 60(1):131–160, January 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0262-y.pdf.

Pedersen:2010:ITS

[1850] Morten Gram Pedersen and Alberto Maria Bersani. Introducing total substrates simplifies theoretical analysis at non-negligible enzyme concentrations: pseudo first-order kinetics and the loss of zeroorder ultrasensitivity. Journal of Mathematical Biology, 60(2):267–283, February 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0267-6.pdf.

Calsina:2010:EAD

[1851] Àngel Calsina and Jordi Ripoll. Evolution of age-dependent sex-reversal under adaptive dynamics. Journal of Mathematical Biology, 60(2):??, February 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0261-z.pdf.

Dong:2010:IFL

[1852] Chao-Yi Dong, Tae-Woong Yoon, Declan G. Bates, and Kwang-Hyun Cho. Identification of feedback loops embedded in cellular circuits by investigating non-causal impulse response components. Journal of Mathematical Biology, 60(2):??, February 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-009-0263-x.pdf.

Mynard:2010:PS

[1853] Frédéric Mynard and Gavin J. Seal. Phenotype spaces. Journal of Mathematical Biology, 60(2):??, February 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-009-0265-8.pdf.

Watkins:2010:CTE

[1854] Joseph C. Watkins. Convergence time to the Ewens sampling formula in the infinite alleles Moran model. Journal of Mathematical Biology, 60 (2):??, February 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0255-x.pdf.

Zeiser:2010:AGN

[1855] Stefan Zeiser, Uwe Franz, and Volkmar Liebscher. Autocatalytic genetic networks modeled by piecewise-deterministic Markov processes. Journal of Mathematical Biology, 60(2):??, February 2010. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0264-9.pdf.

Hek:2010:GSP

[1856] Geertje Hek. Geometric singular perturbation theory in biological practice. Journal of Mathematical Biology, 60(3):347–386, March 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-009-0266-7.pdf.

Durand:2010:PAG

[1857] Eric Durand and Olivier François. Probabilistic analysis of a genealogical model of animal group patterns. Journal of Mathematical Biology, 60 (3):??, March 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0270-y.pdf.

Lambert:2010:PGE

[1858] Amaury Lambert. Population genetics, ecology and the size of populations. Journal of Mathematical Biology, 60(3):??, March 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0286-3.pdf.

Lee:2010:MTS

[1859] Chang Hyeong Lee and Hans G. Othmer. A multi-time-scale analysis of chemical reaction networks: I. Deterministic systems. Journal of Mathematical Biology, 60(3):??, March 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-009-0269-4.pdf.

Woodroffe:2010:MAG

[1860] P. J. Woodroffe, L. J. Bridge, J. R. King, C. Y. Chen, and S. J. Hill. Modelling of the activation of G-protein coupled receptors: drug free constitutive receptor activity. Journal of Mathematical Biology, 60(3): ??, March 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0268-5.pdf.

Keener:2010:ERM

[1861] James P. Keener. Exact reductions of Markovian dynamics for ion channels with a single permissive state. Journal of Mathematical Biology, 60 (4):473–479, April 2010. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-009-0271-x.pdf.

Chang:2010:LTP

[1862] Huilan Chang and Michael Fuchs. Limit theorems for patterns in phylogenetic trees. Journal of Mathematical Biology, 60(4):481–512, April 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0275-6.pdf.

Lu:2010:SPP

[1863] Linghong Lu and Roderick Edwards. Structural principles for periodic orbits in glass networks. Journal of Mathematical Biology, 60(4):513– 541, April 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0273-8.pdf.

Wiegmann:2010:MAM

[1864] Daniel D. Wiegmann, Kelly L. Weinersmith, and Steven M. Seubert. Multi-attribute mate choice decisions and uncertainty in the decision process: a generalized sequential search strategy. Journal of Mathematical Biology, 60(4):543–572, April 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0274-7.pdf.

Hews:2010:RDH

[1865] Sarah Hews, Steffen Eikenberry, John D. Nagy, and Yang Kuang. Rich dynamics of a hepatitis B viral infection model with logistic hepatocyte growth. Journal of Mathematical Biology, 60(4):573– 590, April 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0278-3.pdf.

Srinivasu:2010:TOC

[1866] P. D. N. Srinivasu and B. S. R. V. Prasad. Time optimal control of an additional food provided predator–prey system with applications to pest management and biological conservation. Journal of Mathematical Biology, 60(4):591–613, April 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0279-2.pdf.

Rempe:2010:MMS

[1867] Michael J. Rempe, Janet Best, and David Terman. A mathematical model of the sleep/wake cycle. Journal of Mathematical Biology, 60(5): 615–644, May 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0276-5.pdf.

Li:2010:DEM

[1868] Jing Li and Xingfu Zou. Dynamics of an epidemic model with non-local infections for diseases with latency over a patchy environment. Journal of Mathematical Biology, 60(5):645–686, May 2010. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0280-9.pdf.

McWalter:2010:RRI

[1869] Thomas Andrew McWalter and Alex Welte. Relating recent infection prevalence to incidence with a sub-population of assay non-progressors. Journal of Mathematical Biology, 60(5):687–710, May 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0282-7.pdf.

Smith:2010:ATP

[1870] Amber M. Smith, Frederick R. Adler, and Alan S. Perelson. An accurate two-phase approximate solution to an acute viral infection model. Journal of Mathematical Biology, 60(5):711–726, May 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0281-8.pdf.

vonWangenheim:2010:SCR

[1871] Ute von Wangenheim, Ellen Baake, and Michael Baake. Single-crossover recombination in discrete time. Journal of Mathematical Biology, 60(5): 727–760, May 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0277-4.pdf.

Jagers:2010:PSP

[1872] Peter Jagers. A plea for stochastic population dynamics. Journal of Mathematical Biology, 60(5):761–764, May 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-010-0335-y.pdf.

Kelkel:2010:SRD

[1873] Jan Kelkel and Christina Surulescu. On a stochastic reaction–diffusion system modeling pattern formation on seashells. Journal of Mathematical Biology, 60(6):765–796, June 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0284-5.pdf.

Cattiaux:2010:CWC

[1874] Patrick Cattiaux and Sylvie Méléard. Competitive or weak cooperative stochastic Lotka–Volterra systems conditioned on non-extinction. Journal of Mathematical Biology, 60(6):797–829, June 2010. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0285-4.pdf.

Buffoni:2010:IBM

[1875] Giuseppe Buffoni and Sara Pasquali. Individual-based models for stage structured populations: formulation of “no regression” development equations. Journal of Mathematical Biology, 60(6):831–848, June 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0287-2.pdf.

Tang:2010:ADF

[1876] Fusheng Tang and Weijiu Liu. An age-dependent feedback control model of calcium dynamics in yeast cells. Journal of Mathematical Biology, 60(6):849–879, June 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0289-0.pdf.

Friedman:2010:CCC

[1877] Avner Friedman, Bei Hu, and Chiu-Yen Kao. Cell cycle control at the first restriction point and its effect on tissue growth. Journal of Mathematical Biology, 60(6):881–907, June 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-009-0290-7.pdf.

Luo:2010:DEM

[1878] Jiangtao Luo, William W. Hager, and Rongling Wu. A differential equation model for functional mapping of a virus-cell dynamic system. Journal of Mathematical Biology, 61(1):1–15, July 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-009-0288-1.pdf.

Jabbari:2010:MMA

[1879] Sara Jabbari, John R. King, Adrian J. Koerber, and Paul Williams. Mathematical modelling of the agr operon in Staphylococcus aureus. Journal of Mathematical Biology, 61(1):17–54, July 2010. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0291-6.pdf.

Angerer:2010:PMD

[1880] Wolfgang P. Angerer. Proliferation model dependence in fluctuation analysis: the neutral case. Journal of Mathematical Biology, 61(1): 55–93, July 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0294-3.pdf.

Schneider:2010:MPF

[1881] Kristan Alexander Schneider. Maximization principles for frequencydependent selection II: the one-locus multiallele case. Journal of Mathematical Biology, 61(1):95–132, July 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-009-0292-5.pdf.

Madzvamuse:2010:SAN

[1882] Anotida Madzvamuse, Eamonn A. Gaffney, and Philip K. Maini. Stability analysis of non-autonomous reaction–diffusion systems: the effects of growing domains. Journal of Mathematical Biology, 61(1):133– 164, July 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0293-4.pdf.

Kisdi:2010:ADF

[1883] Éva Kisdi and Stefan A. H. Geritz. Adaptive dynamics: a framework to model evolution in the ecological theatre. Journal of Mathematical Biology, 61(1):165–169, July 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0300-9.pdf.

Palumbo:2010:IPM

[1884] Pasquale Palumbo and Andrea De Gaetano. An islet population model of the endocrine pancreas. Journal of Mathematical Biology, 61(2):171– 205, August 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0297-0.pdf.

Uriu:2010:SOS

[1885] Koichiro Uriu, Yoshihiro Morishita, and Yoh Iwasa. Synchronized oscillation of the segmentation clock gene in vertebrate development. Journal of Mathematical Biology, 61(2):207–229, August 2010. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0296-1.pdf.

Smiley:2010:GED

[1886] Michael W. Smiley and Stephen R. Proulx. Gene expression dynamics in randomly varying environments. Journal of Mathematical Biology, 61(2): 231–251, August 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0298-z.pdf.

Cardona:2010:NDR

[1887] Gabriel Cardona, Mercè Llabrés, Francesc Rosselló, and Gabriel Valiente. Nodal distances for rooted phylogenetic trees. Journal of Mathematical Biology, 61(2):253–276, August 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0295-2.pdf.

Diekmann:2010:DRL

[1888] Odo Diekmann, Mats Gyllenberg, J. A. J. Metz, Shinji Nakaoka, and Andre M. de Roos. Daphnia revisited: local stability and bifurcation theory for physiologically structured population models explained by way of an example. Journal of Mathematical Biology, 61(2):277– 318, August 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0299-y.pdf. See erratum [2662].

Srinivasu:2010:ETO

[1889] P. D. N. Srinivasu and B. S. R. V. Prasad. Erratum to: Time optimal control of an additional food provided predator–prey system with applications to pest management and biological conservation. Journal of Mathematical Biology, 61(2):319–321, August 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-009-0301-8.pdf.

Vanpeteghem:2010:AAS

[1890] Dimitri Vanpeteghem and Bart Haegeman. An analytical approach to spatio-temporal dynamics of neutral community models. Journal of Mathematical Biology, 61(3):323–357, September 2010. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0303-6.pdf.

Iizuka:2010:EPS

[1891] Masaru Iizuka. Effective population size of a population with stochastically varying size. Journal of Mathematical Biology, 61(3):359–375, September 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0304-5.pdf.

Balasuriya:2010:WRD

[1892] Sanjeeva Balasuriya and Georg A. Gottwald. Wavespeed in reaction– diffusion systems, with applications to chemotaxis and population pressure. Journal of Mathematical Biology, 61(3):377–399, September 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-009-0305-4.pdf.

Kim:2010:TEC

[1893] Yangjin Kim, Julie Wallace, Fu Li, Michael Ostrowski, and Avner Friedman. Transformed epithelial cells and fibroblasts/myofibroblasts interaction in breast tumor: a mathematical model and experiments. Journal of Mathematical Biology, 61(3):401–421, September 2010. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0307-2.pdf.

Terekhov:2010:AAP

[1894] Alexander V. Terekhov, Yakov B. Pesin, Xun Niu, Mark L. Latash, and Vladimir M. Zatsiorsky. An analytical approach to the problem of inverse optimization with additive objective functions: an application to human prehension. Journal of Mathematical Biology, 61(3):423– 453, September 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0306-3.pdf.

Bobrowski:2010:ABM

[1895] Adam Bobrowski, Tomasz Wojdyla, and Marek Kimmel. Asymptotic behavior of a Moran model with mutations, drift and recombination among multiple loci. Journal of Mathematical Biology, 61(3):455–473, September 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0308-1.pdf.

DeLeenheer:2010:SAR

[1896] Patrick De Leenheer, Jack Dockery, Tomás Gedeon, and Sergei S. Pilyugin. Senescence and antibiotic resistance in an age-structured population model. Journal of Mathematical Biology, 61(4):475–499, October 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0302-7.pdf.

Small:2010:UBP

[1897] Michael Small, Hugh P. C. Robinson, Ingo C. Kleppe, and Chi Kong Tse. Uncovering bifurcation patterns in cortical synapses. Journal of Mathematical Biology, 61(4):501–526, October 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-009-0312-5.pdf.

Clancy:2010:EWI

[1898] Damian Clancy and Sang Taphou Mendy. The effect of waning immunity on long-term behaviour of stochastic models for the spread of infection. Journal of Mathematical Biology, 61(4):527–544, October 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0313-4.pdf.

Fetecau:2010:INH

[1899] Razvan C. Fetecau and Raluca Eftimie. An investigation of a nonlocal hyperbolic model for self-organization of biological groups. Journal of Mathematical Biology, 61(4):545–579, October 2010. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0311-6.pdf.

Angeli:2010:GTC

[1900] David Angeli, Patrick De Leenheer, and Eduardo Sontag. Graphtheoretic characterizations of monotonicity of chemical networks in reaction coordinates. Journal of Mathematical Biology, 61(4):581–616, October 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0309-0.pdf.

Groh:2010:SMB

[1901] Andreas Groh and Alfred K. Louis. Stochastic modelling of biased cell migration and collagen matrix modification. Journal of Mathematical Biology, 61(5):617–647, November 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0314-3.pdf.

Bergdorf:2010:LPM

[1902] Michael Bergdorf, Ivo F. Sbalzarini, and Petros Koumoutsakos. A Lagrangian particle method for reaction–diffusion systems on deforming surfaces. Journal of Mathematical Biology, 61(5):649–663, November 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-009-0315-2.pdf.

Hamelin:2010:DGT

[1903] Frédéric M. Hamelin and Mark A. Lewis. A differential game theoretical analysis of mechanistic models for territoriality. Journal of Mathematical Biology, 61(5):665–694, November 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0316-1.pdf.

Notohara:2010:ACL

[1904] Morihiro Notohara. An application of the central limit theorem to coalescence times in the structured coalescent model with strong migration. Journal of Mathematical Biology, 61(5):695–714, November 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-009-0318-z.pdf.

Linz:2010:ARR

[1905] Simone Linz, Charles Semple, and Tanja Stadler. Analyzing and reconstructing reticulation networks under timing constraints. Journal of Mathematical Biology, 61(5):715–737, November 2010. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0319-y.pdf.

Lui:2010:TWS

[1906] Roger Lui and Zhi An Wang. Traveling wave solutions from microscopic to macroscopic chemotaxis models. Journal of Mathematical Biology, 61(5):739–761, November 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0317-0.pdf.

Britton:2010:TES

[1907] Tom Britton and Peter Neal. The time to extinction for a stochastic SIShousehold-epidemic model. Journal of Mathematical Biology, 61(6):763– 779, December 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0320-5.pdf.

Li:2010:GKM

[1908] Yongfeng Li and Jeyaraman Srividhya. Goldbeter–Koshland model for open signaling cascades: a mathematical study. Journal of Mathematical Biology, 61(6):781–803, December 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-009-0322-3.pdf.

Weller:2010:FBP

[1909] Frédéric Frank Weller. A free boundary problem modeling thrombus growth. Journal of Mathematical Biology, 61(6):805–818, December 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0324-1.pdf.

Leocard:2010:EAR

[1910] Stephanie Leocard and Etienne Pardoux. Evolution of the ancestral recombination graph along the genome in case of selective sweep. Journal of Mathematical Biology, 61(6):819–841, December 2010. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-009-0321-4.pdf.

Terry:2010:ICS

[1911] Alan J. Terry. Impulsive culling of a structured population on two patches. Journal of Mathematical Biology, 61(6):843–875, December 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0325-0.pdf.

Stadler:2010:CVF

[1912] Bärbel M. R. Stadler and Peter F. Stadler. Combinatorial vector fields and the valley structure of fitness landscapes. Journal of Mathematical Biology, 61(6):877–898, December 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0326-z.pdf.

Alt:2010:MSD

[1913] Wolfgang Alt. Model-supported data analysis: some biological principles and examples. Journal of Mathematical Biology, 61(6):899–903, December 2010. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-009-0310-7.pdf.

Kazmierczak:2011:CWF

[1914] Bogdan Kaźmierczak and Zbigniew Peradzyński. Calcium waves with fast buffers and mechanical effects. Journal of Mathematical Biology, 62 (1):1–38, January 2011. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-009-0323-2.pdf.

Bonzi:2011:SDS

[1915] B. Bonzi, A. A. Fall, A. Iggidr, and G. Sallet. Stability of differential susceptibility and infectivity epidemic models. Journal of Mathematical Biology, 62(1):39–64, January 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0327-y.pdf.

RoyChoudhury:2011:CLB

[1916] Arindam RoyChoudhury. Composite likelihood-based inferences on genetic data from dependent loci. Journal of Mathematical Biology, 62(1): 65–80, January 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0329-9.pdf.

Ballyk:2011:CRB

[1917] Mary M. Ballyk and Gail S. K. Wolkowicz. Classical and resource-based competition: a unifying graphical approach. Journal of Mathematical Biology, 62(1):81–109, January 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0328-x.pdf.

Gandolfi:2011:ASM

[1918] Alberto Gandolfi, Mimmo Iannelli, and Gabriela Marinoschi. An agestructured model of epidermis growth. Journal of Mathematical Biology, 62(1):111–141, January 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0330-3.pdf.

Lindquist:2011:EDN

[1919] Jennifer Lindquist, Junling Ma, P. van den Driessche, and Frederick H. Willeboordse. Effective degree network disease models. Journal of Mathematical Biology, 62(2):143–164, February 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-010-0331-2.pdf.

Lin:2011:SST

[1920] Guo Lin, Wan-Tong Li, and Shigui Ruan. Spreading speeds and traveling waves in competitive recursion systems. Journal of Mathematical Biology, 62(2):165–201, February 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0334-z.pdf.

Leviyang:2011:DOG

[1921] Sivan Leviyang. The distribution of F_{st} and other genetic statistics for a class of population structure models. Journal of Mathematical Biology, 62(2):203–289, February 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0333-0.pdf.

Wang:2011:PPS

[1922] Jinfeng Wang, Junping Shi, and Junjie Wei. predator–prey system with strong Allee effect in prey. Journal of Mathematical Biology, 62(3):291– 331, March 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0332-1.pdf.

Andersson:2011:SSE

[1923] Patrik Andersson and David Lindenstrand. A stochastic SIS epidemic with demography: initial stages and time to extinction. Journal of Mathematical Biology, 62(3):333–348, March 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-010-0336-x.pdf.

Miller:2011:NPE

[1924] Joel C. Miller. A note on a paper by Erik Volz: SIR dynamics in random networks. Journal of Mathematical Biology, 62(3):349– 358, March 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0337-9.pdf. See [1710].

Gyllenberg:2011:NSC

[1925] Mats Gyllenberg and Robert Service. Necessary and sufficient conditions for the existence of an optimisation principle in evolution. Journal of Mathematical Biology, 62(3):359–369, March 2011. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0340-1.pdf.

Blackstone:2011:MCI

[1926] T. Blackstone, R. Scharein, B. Borgo, R. Varela, Y. Diao, and J. Arsuaga. Modeling of chromosome intermingling by partially overlapping uniform random polygons. Journal of Mathematical Biology, 62(3):371– 389, March 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0338-8.pdf.

Souplet:2011:ISD

[1927] Philippe Souplet and Michael Winkler. The influence of space dimension on the large-time behavior in a reaction–diffusion system modeling diallelic selection. Journal of Mathematical Biology, 62(3):391– 421, March 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0339-7.pdf.

Hansen:2011:OCE

[1928] Elsa Hansen and Troy Day. Optimal control of epidemics with limited resources. Journal of Mathematical Biology, 62(3):423–451, March 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0341-0.pdf.

Edwards:2011:CDS

[1929] Roderick Edwards, Sehjeong Kim, and P. van den Driessche. Control design for sustained oscillation in a two-gene regulatory network. Journal of Mathematical Biology, 62(4):453–478, April 2011. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0343-y.pdf.

Simon:2011:EEM

[1930] Péter L. Simon, Michael Taylor, and Istvan Z. Kiss. Exact epidemic models on graphs using graph-automorphism driven lumping. Journal of Mathematical Biology, 62(4):479–508, April 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-010-0344-x.pdf.

McCluskey:2011:ACM

[1931] C. Connell McCluskey and David J. D. Earn. Attractivity of coherent manifolds in metapopulation models. Journal of Mathematical Biology, 62(4):509–541, April 2011. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-010-0342-z.pdf.

Lou:2011:RDM

[1932] Yijun Lou and Xiao-Qiang Zhao. A reaction–diffusion malaria model with incubation period in the vector population. Journal of Mathematical Biology, 62(4):543–568, April 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0346-8.pdf.

Cristiani:2011:EAI

[1933] Emiliano Cristiani, Paolo Frasca, and Benedetto Piccoli. Effects of anisotropic interactions on the structure of animal groups. Journal of Mathematical Biology, 62(4):569–588, April 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-010-0347-7.pdf.

Previte:2011:ERF

[1934] Joseph P. Previte, Natalie Sheils, Kathleen A. Hoffman, Tim Kiemel, and Eric D. Tytell. Entrainment ranges of forced phase oscillators. Journal of Mathematical Biology, 62(4):589–603, April 2011. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0348-6.pdf.

Raghib:2011:MME

[1935] Michael Raghib, Nicholas A. Hill, and Ulf Dieckmann. A multiscale maximum entropy moment closure for locally regulated space-time point process models of population dynamics. Journal of Mathematical Biology, 62(5):605–653, May 2011. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-010-0345-9.pdf.

Schreiber:2011:PFE

[1936] Sebastian J. Schreiber, Michel Benaïm, and Kolawolé A. S. Atchadé. Persistence in fluctuating environments. Journal of Mathematical Biology, 62(5):655–683, May 2011. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-010-0349-5.pdf.

Marucci:2011:DIV

[1937] Lucia Marucci, Stefania Santini, Mario di Bernardo, and Diego di Bernardo. Derivation, identification and validation of a computational model of a novel synthetic regulatory network in yeast. Journal of Mathematical Biology, 62(5):685–706, May 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-010-0350-z.pdf.

Gyrya:2011:ESV

[1938] V. Gyrya, K. Lipnikov, I. S. Aranson, and L. Berlyand. Effective shear viscosity and dynamics of suspensions of micro-swimmers from small to moderate concentrations. Journal of Mathematical Biology, 62(5): 707–740, May 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0351-y.pdf.

Bacaer:2011:GSB

[1939] Nicolas Bacaër and El Hadi Ait Dads. Genealogy with seasonality, the basic reproduction number, and the influenza pandemic. Journal of Mathematical Biology, 62(5):741–762, May 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-010-0354-8.pdf.

Park:2011:BTL

[1940] Su-Chan Park and Joachim Krug. Bistability in two-locus models with selection, mutation, and recombination. Journal of Mathematical Biology, 62(5):763–788, May 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0352-x.pdf.

Schneider:2011:AHE

[1941] Kristan A. Schneider and Yuseob Kim. Approximations for the hitchhiking effect caused by the evolution of antimalarial-drug resistance. Journal of Mathematical Biology, 62(6):789–832, June 2011. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0353-9.pdf.

Allman:2011:IRS

[1942] Elizabeth S. Allman, James H. Degnan, and John A. Rhodes. Identifying the rooted species tree from the distribution of unrooted gene trees under the coalescent. Journal of Mathematical Biology, 62(6): 833–862, June 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0355-7.pdf.

Cessac:2011:DTN

[1943] B. Cessac. A discrete time neural network model with spiking neurons: II: Dynamics with noise. Journal of Mathematical Biology, 62(6): 863–900, June 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0358-4.pdf.

Bowers:2011:DEO

[1944] Roger G. Bowers. On the determination of evolutionary outcomes directly from the population dynamics of the resident. Journal of Mathematical Biology, 62(6):901–924, June 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-010-0356-6.pdf.

Kang:2011:EED

[1945] Yun Kang and Nicolas Lanchier. Expansion or extinction: deterministic and stochastic two-patch models with Allee effects. Journal of Mathematical Biology, 62(6):925–973, June 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-010-0359-3.pdf.

Ainseba:2011:GDH

[1946] Bedr’Eddine Aïnseba and Chahrazed Benosman. Global dynamics of hematopoietic stem cells and differentiated cells in a chronic myeloid leukemia model. Journal of Mathematical Biology, 62(6):975– 997, June 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0360-x.pdf.

Biler:2011:LMS

[1947] Piotr Biler, Lucilla Corrias, and Jean Dolbeault. Large mass self-similar solutions of the parabolic-parabolic Keller–Segel model of chemotaxis. Journal of Mathematical Biology, 63(1):1–32, July 2011. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0357-5.pdf.

Jones:2011:CMT

[1948] Graham Jones. Calculations for multi-type age-dependent binary branching processes. Journal of Mathematical Biology, 63(1):33– 56, July 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0362-8.pdf.

Lambert:2011:SAD

[1949] Amaury Lambert. Species abundance distributions in neutral models with immigration or mutation and general lifetimes. Journal of Mathematical Biology, 63(1):57–72, July 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0361-9.pdf.

Kwok:2011:PGL

[1950] Rex Bing Hung Kwok. Phylogeny, genealogy and the Linnaean hierarchy: a logical analysis. Journal of Mathematical Biology, 63(1):73– 108, July 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0364-6.pdf.

Forsyth:2011:PGG

[1951] Peter A. I. Forsyth and Christoph Hauert. Public goods games with reward in finite populations. Journal of Mathematical Biology, 63(1):

109–123, July 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0363-7.pdf.

Huang:2011:GAD

[1952] Gang Huang and Yasuhiro Takeuchi. Global analysis on delay epidemiological dynamic models with nonlinear incidence. Journal of Mathematical Biology, 63(1):125–139, July 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0368-2.pdf.

Andasari:2011:MMC

[1953] Vivi Andasari, Alf Gerisch, Georgios Lolas, Andrew P. South, and Mark A. J. Chaplain. Mathematical modeling of cancer cell invasion of tissue: biological insight from mathematical analysis and computational simulation. Journal of Mathematical Biology, 63(1):141– 171, July 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0369-1.pdf.

Mente:2011:PEN

[1954] Carsten Mente, Ina Prade, Lutz Brusch, Georg Breier, and Andreas Deutsch. Parameter estimation with a novel gradient-based optimization method for biological lattice-gas cellular automaton models. Journal of Mathematical Biology, 63(1):173–200, July 2011. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0366-4.pdf.

Delgado-Eckert:2011:MHR

[1955] Edgar Delgado-Eckert and Michael Shapiro. A model of host response to a multi-stage pathogen. Journal of Mathematical Biology, 63(2):201– 227, August 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0365-5.pdf.

Lee:2011:MFV

[1956] K.-C. Lee, A. Gopinathan, and J. M. Schwarz. Modeling the formation of in vitro filopodia. Journal of Mathematical Biology, 63(2):229– 261, August 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0371-7.pdf.

Brannstrom:2011:CFG

[1957] Åke Brännström, Thilo Gross, Bernd Blasius, and Ulf Dieckmann. Consequences of fluctuating group size for the evolution of cooperation. Journal of Mathematical Biology, 63(2):263–281, August 2011. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0367-3.pdf.

Sarkar:2011:DRD

[1958] Ram Rup Sarkar, R. Maithreye, and Somdatta Sinha. Design of regulation and dynamics in simple biochemical pathways. Journal of Mathematical Biology, 63(2):283–307, August 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-010-0375-3.pdf.

Ball:2011:HEM

[1959] Frank Ball, Tom Britton, and David Sirl. Household epidemic models with varying infection response. Journal of Mathematical Biology, 63(2): 309–337, August 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0372-6.pdf.

Benson:2011:GMD

[1960] James D. Benson, Carmen C. Chicone, and John K. Critser. A general model for the dynamics of cell volume, global stability, and optimal control. Journal of Mathematical Biology, 63(2):339–359, August 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0374-4.pdf.

Kisdi:2011:EBM

[1961] Éva Kisdi and Tadeas Priklopil. Evolutionary branching of a magic trait. Journal of Mathematical Biology, 63(2):361–397, August 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0377-1.pdf.

Ewing:2011:SSR

[1962] G. Ewing, J. Hermisson, P. Pfaffelhuber, and J. Rudolf. Selective sweeps for recessive alleles and for other modes of dominance. Journal of Mathematical Biology, 63(3):399–431, September 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-010-0382-4.pdf.

Baxendale:2011:SOD

[1963] Peter H. Baxendale and Priscilla E. Greenwood. Sustained oscillations for density dependent Markov processes. Journal of Mathematical Biology, 63(3):433–457, September 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0376-2.pdf.

Yarahmadian:2011:ESS

[1964] Shantia Yarahmadian, Blake Barker, Kevin Zumbrun, and Sidney L. Shaw. Existence and stability of steady states of a reaction convection diffusion equation modeling microtubule formation. Journal of Mathematical Biology, 63(3):459–492, September 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-010-0379-z.pdf.

Jabin:2011:SDC

[1965] Pierre-Emmanuel Jabin and Gaël Raoul. On selection dynamics for competitive interactions. Journal of Mathematical Biology, 63(3):493–517, September 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0370-8.pdf.

Caceres:2011:RLM

[1966] Manuel O. Cáceres and Iris Cáceres-Saez. Random Leslie matrices in population dynamics. Journal of Mathematical Biology, 63(3):519–556, September 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0378-0.pdf.

Su:2011:PSB

[1967] Ying Su, Shigui Ruan, and Junjie Wei. Periodicity and synchronization in blood-stage malaria infection. Journal of Mathematical Biology, 63(3):557–574, September 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0381-5.pdf.

Brannstrom:2011:RCF

[1968] Åke Brännström, Linus Carlsson, and Axel G. Rossberg. Rigorous conditions for food-web intervality in high-dimensional trophic niche spaces. Journal of Mathematical Biology, 63(3):575–592, September 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-010-0383-3.pdf.

Richard:2011:SPN

[1969] Adrien Richard and Jean-Paul Comet. Stable periodicity and negative circuits in differential systems. Journal of Mathematical Biology, 63(3):593–600, September 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0388-y.pdf.

Sasaki:2011:ODA

[1970] Akira Sasaki and Ulf Dieckmann. Oligomorphic dynamics for analyzing the quantitative genetics of adaptive speciation. Journal of Mathematical Biology, 63(4):601–635, October 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0380-6.pdf.

Bailly-Bechet:2011:EMN

[1971] Marc Bailly-Bechet, Arndt Benecke, Wolf Dietrich Hardt, Valentina Lanza, Alexander Sturm, and Riccardo Zecchina. An externally modulated, noise-driven switch for the regulation of SPI1 in Salmonella enterica serovar typhimurium. Journal of Mathematical Biology, 63(4):637– 662, October 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0385-1.pdf.

Vinther:2011:MMH

[1972] Frank Vinther, Morten Andersen, and Johnny T. Ottesen. The minimal model of the hypothalamic-pituitary-adrenal axis. Journal of Mathematical Biology, 63(4):663–690, October 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-010-0384-2.pdf.

Pellis:2011:EGR

[1973] Lorenzo Pellis, Neil M. Ferguson, and Christophe Fraser. Epidemic growth rate and household reproduction number in communities of households, schools and workplaces. Journal of Mathematical Biology, 63(4):691–734, October 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0386-0.pdf.

Quint:2011:OOB

[1974] D. A. Quint and J. M. Schwarz. Optimal orientation in branched cytoskeletal networks. Journal of Mathematical Biology, 63(4):735–755, October 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0389-x.pdf.

Grote:2011:DFO

[1975] Marcus J. Grote, Viviana Palumberi, Barbara Wagner, Andrea Barbero, and Ivan Martin. Dynamic formation of oriented patches in chondrocyte cell cultures. Journal of Mathematical Biology, 63(4):757– 777, October 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0390-4.pdf.

Datta:2011:SAP

[1976] Samik Datta, Gustav W. Delius, Richard Law, and Michael J. Plank. A stability analysis of the power-law steady state of marine size spectra. Journal of Mathematical Biology, 63(4):779–799, October 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0387-z.pdf.

Carrillo:2011:DMF

[1977] José Antonio Carrillo, Stéphane Cordier, and Simona Mancini. A decision-making Fokker–Planck model in computational neuroscience. Journal of Mathematical Biology, 63(5):801–830, November 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-010-0391-3.pdf.

Levy:2011:MCS

[1978] Chris Levy and David Iron. Model of cell signal transduction in a threedimensional domain. Journal of Mathematical Biology, 63(5):831–854, November 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0394-0.pdf.

Efimov:2011:PRC

[1979] D. V. Efimov. Phase resetting control based on direct phase response curve. Journal of Mathematical Biology, 63(5):855–879, November 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-010-0396-y.pdf.

Alshamary:2011:GOP

[1980] Bader Alshamary and Ovidiu Calin. Geometric optimization for preypredator strategies. Journal of Mathematical Biology, 63(5):881–900, November 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0393-1.pdf.

Li:2011:GAS

[1981] Xiong Li, Hao Wang, and Yang Kuang. Global analysis of a stoichiometric producer-grazer model with Holling type functional responses. Journal of Mathematical Biology, 63(5):901–932, November 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0392-2.pdf.

Yousfi:2011:MAI

[1982] Noura Yousfi, Khalid Hattaf, and Abdessamad Tridane. Modeling the adaptive immune response in HBV infection. Journal of Mathematical Biology, 63(5):933–957, November 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0397-x.pdf.

King:2011:MMB

[1983] Julian King, Karl Unterkofler, Gerald Teschl, Susanne Teschl, Helin Koc, Hartmann Hinterhuber, and Anton Amann. A mathematical model for breath gas analysis of volatile organic compounds with special emphasis on acetone. Journal of Mathematical Biology, 63(5):959–999, November 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0398-9.pdf.

Hu:2011:TAF

[1984] Jifeng Hu and Hans G. Othmer. A theoretical analysis of filament length fluctuations in actin and other polymers. Journal of Mathematical Biology, 63(6):1001–1049, December 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-010-0400-6.pdf.

Buckwar:2011:ESH

[1985] Evelyn Buckwar and Martin G. Riedler. An exact stochastic hybrid model of excitable membranes including spatio-temporal evolution. Journal of Mathematical Biology, 63(6):1051–1093, December 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-010-0395-z.pdf.

Barreira:2011:SFE

[1986] R. Barreira, C. M. Elliott, and A. Madzvamuse. The surface finite element method for pattern formation on evolving biological surfaces. Journal of Mathematical Biology, 63(6):1095–1119, December 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0401-0.pdf.

Seubert:2011:VMQ

[1987] Steven M. Seubert, Gordon A. Wade, and Daniel D. Wiegmann. The variability of male quality and female mate choice decisions: secondorder stochastic dominance and the behavior of searchers under a sequential search strategy. Journal of Mathematical Biology, 63(6):1121– 1138, December 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0399-8.pdf.

Chen:2011:DGB

[1988] Zhan Chen, Nathan A. Baker, and G. W. Wei. Differential geometry based solvation model II: Lagrangian formulation. Journal of Mathematical Biology, 63(6):1139–1200, December 2011. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-011-0402-z.pdf.

Blossey:2012:SPM

[1989] Ralf Blossey, Jean-François Bodart, Anne Devys, Thierry Goudon, and Pauline Lafitte. Signal propagation of the MAPK cascade in xenopus oocytes: role of bistability and ultrasensitivity for a mixed problem. Journal of Mathematical Biology, 64(1–2):1–39, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0403-y.pdf.

Venkataraman:2012:GES

[1990] Chandrasekhar Venkataraman, Omar Lakkis, and Anotida Madzvamuse. Global existence for semilinear reaction–diffusion systems on evolving domains. Journal of Mathematical Biology, 64(1–2):41–67, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0404-x.pdf.

Bordewich:2012:BNR

[1991] Magnus Bordewich and Charles Semple. Budgeted nature reserve selection with diversity feature loss and arbitrary split systems. Journal of Mathematical Biology, 64(1–2):69–85, January 2012. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0405-9.pdf.

Reddy:2012:RRB

[1992] Shashir B. Reddy and Noah A. Rosenberg. Refining the relationship between homozygosity and the frequency of the most frequent allele. Journal of Mathematical Biology, 64(1–2):87–108, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0406-8.pdf.

Hsu:2012:LVC

[1993] Sze-Bi Hsu and Xiao-Qiang Zhao. A Lotka–Volterra competition model with seasonal succession. Journal of Mathematical Biology, 64(1–2):109– 130, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0408-6.pdf.

Taati:2012:IDB

[1994] Siamak Taati, Enrico Formenti, Jean-Paul Comet, and Gilles Bernot. On the impact of the distance between two genes on their interaction curve. Journal of Mathematical Biology, 64(1–2):131–147, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-010-0373-5.pdf.

Bryant:2012:LBS

[1995] David Bryant and Steffen Klaere. The link between segregation and phylogenetic diversity. Journal of Mathematical Biology, 64(1–2):149– 162, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0409-5.pdf.

Billiard:2012:GSM

[1996] Sylvain Billiard and Viet Chi Tran. A general stochastic model for sporophytic self-incompatibility. Journal of Mathematical Biology, 64(1–2): 163–210, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-011-0410-z.pdf.

Wei:2012:FDS

[1997] Juncheng Wei and Matthias Winter. Flow-distributed spikes for Schnakenberg kinetics. Journal of Mathematical Biology, 64(1–2):211– 254, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0412-x.pdf.

Gomez-Corral:2012:ETS

[1998] A. Gómez-Corral and M. López García. Extinction times and size of the surviving species in a two-species competition process. Journal of Mathematical Biology, 64(1–2):255–289, January 2012. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0414-8.pdf.

Pokalyuk:2012:ERM

[1999] Cornelia Pokalyuk. The effect of recurrent mutation on the linkage disequilibrium under a selective sweep. Journal of Mathematical Biology, 64(1–2):291–317, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0411-y.pdf.

Gutierrez:2012:ATC

[2000] Juan B. Gutierrez, Monica K. Hurdal, Rana D. Parshad, and John L. Teem. Analysis of the Trojan Y chromosome model for eradication of invasive species in a dendritic riverine system. Journal of Mathematical Biology, 64(1–2):319–340, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0413-9.pdf.

Sieber:2012:HEP

[2001] Michael Sieber and Frank M. Hilker. The hydra effect in predator– prey models. Journal of Mathematical Biology, 64(1–2):341–360, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0416-6.pdf.

Hacker:2012:MMM

[2002] Anita Häcker. A mathematical model for mesenchymal and chemosensitive cell dynamics. Journal of Mathematical Biology, 64(1–2):361– 401, January 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0415-7.pdf.

Bacaer:2012:MKM

[2003] Nicolas Bacaër. The model of Kermack and McKendrick for the plague epidemic in Bombay and the type reproduction number with seasonality.

Journal of Mathematical Biology, 64(3):403–422, February 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0417-5.pdf.

Arino:2012:MMM

[2004] Julien Arino, Arnaud Ducrot, and Pascal Zongo. A metapopulation model for malaria with transmission-blocking partial immunity in hosts. Journal of Mathematical Biology, 64(3):423–448, February 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0418-4.pdf.

Duff:2012:MMS

[2005] Campbell Duff, Kate Smith-Miles, Leo Lopes, and Tianhai Tian. Mathematical modelling of stem cell differentiation: the PU.1-GATA-1 interaction. Journal of Mathematical Biology, 64(3):449–468, February 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-011-0419-3.pdf.

Sun:2012:MGT

[2006] Qiwen Sun, Moxun Tang, and Jianshe Yu. Modulation of gene transcription noise by competing transcription factors. Journal of Mathematical Biology, 64(3):469–494, February 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0420-x.pdf.

Wu:2012:EGT

[2007] Yu Wu, Mingjun Zhang, Jing Wu, Xiaopeng Zhao, and Lijin Xia. Evolutionary game theoretic strategy for optimal drug delivery to influence selection pressure in treatment of HIV-1. Journal of Mathematical Biology, 64(3):495–512, February 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0422-8.pdf.

Oelz:2012:SLF

[2008] Dietmar Oelz and Christian Schmeiser. Simulation of lamellipodial fragments. Journal of Mathematical Biology, 64(3):513–528, February 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-011-0421-9.pdf.

Li:2012:CRR

[2009] Thomas J. X. Li and Christian M. Reidys. Combinatorics of RNA– RNA interaction. Journal of Mathematical Biology, 64(3):529–556, February 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0423-7.pdf.

Ledzewicz:2012:ORC

[2010] Urszula Ledzewicz, Mohammad Naghnaeian, and Heinz Schättler. Optimal response to chemotherapy for a mathematical model of tumorimmune dynamics. Journal of Mathematical Biology, 64(3):557–577, February 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0424-6.pdf.

ODea:2012:CLP

[2011] R. D. O’Dea and J. R. King. Continuum limits of pattern formation in hexagonal-cell monolayers. Journal of Mathematical Biology, 64(3):579– 610, February 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0427-3.pdf.

Hadeler:2012:FSI

[2012] Karl Hadeler and Alan Hastings. Forward to special issue. Journal of Mathematical Biology, 64(4):611–612, March 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-011-0473-x.pdf.

Hadeler:2012:PF

[2013] K. P. Hadeler. Pair formation. Journal of Mathematical Biology, 64(4): 613–645, March 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0454-0.pdf.

Giuggioli:2012:LAM

[2014] Luca Giuggioli and Frederic Bartumeus. Linking animal movement to site fidelity. Journal of Mathematical Biology, 64(4):647–656, March 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0431-7.pdf.

Durrett:2012:EDD

[2015] Rick Durrett and Daniel Remenik. Evolution of dispersal distance. Journal of Mathematical Biology, 64(4):657–666, March 2012. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0444-2.pdf.

Langebrake:2012:DMM

[2016] Jessica Langebrake, Louise Riotte-Lambert, Craig W. Osenberg, and Patrick De Leenheer. Differential movement and movement bias models for marine protected areas. Journal of Mathematical Biology, 64(4): 667–696, March 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0407-7.pdf.

Jiang:2012:NNM

[2017] Xingpeng Jiang, Joshua S. Weitz, and Jonathan Dushoff. A nonnegative matrix factorization framework for identifying modular patterns in metagenomic profile data. Journal of Mathematical Biology, 64(4): 697–711, March 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0428-2.pdf.

Grover:2012:CBM

[2018] James P. Grover, Sze-Bi Hsu, and Feng-Bin Wang. Competition between microorganisms for a single limiting resource with cell quota structure and spatial variation. Journal of Mathematical Biology, 64(5):713– 743, April 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0426-4.pdf.

Indelicato:2012:CAS

[2019] Giuliana Indelicato, Paolo Cermelli, David G. Salthouse, Simone Racca, Giovanni Zanzotto, and Reidun Twarock. A crystallographic approach to structural transitions in icosahedral viruses. Journal of Mathematical Biology, 64(5):745–773, April 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0425-5.pdf.

McVinish:2012:LBM

[2020] R. McVinish and P. K. Pollett. The limiting behaviour of a mainlandisland metapopulation. Journal of Mathematical Biology, 64(5):775– 801, April 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0429-1.pdf.

Wu:2012:HSS

[2021] Bin Wu, Chaitanya S. Gokhale, Long Wang, and Arne Traulsen. How small are small mutation rates? Journal of Mathematical Biology, 64(5): 803–827, April 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0430-8.pdf.

Bokes:2012:EAD

[2022] Pavol Bokes, John R. King, Andrew T. A. Wood, and Matthew Loose. Exact and approximate distributions of protein and mRNA levels in the low-copy regime of gene expression. Journal of Mathematical Biology, 64 (5):829–854, April 2012. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-011-0433-5.pdf.

Wang:2012:PON

[2023] Yuanshi Wang, Hong Wu, and Shigui Ruan. Periodic orbits near heteroclinic cycles in a cyclic replicator system. Journal of Mathematical Biology, 64(5):855–872, April 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0435-3.pdf.

Mori:2012:EMP

[2024] Yoichiro Mori. Erratum to: Mathematical properties of pump-leak models of cell volume control and electrolyte balance. Journal of Mathematical Biology, 64(5):873–916, April 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-011-0499-0.pdf.

Curuksu:2012:ACS

[2025] Jeremy Curuksu. Adaptive conformational sampling based on replicas. Journal of Mathematical Biology, 64(6):917–931, May 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0432-6.pdf.

Rebelo:2012:PSF

[2026] Carlota Rebelo, Alessandro Margheri, and Nicolas Bacaër. Persistence in seasonally forced epidemiological models. Journal of Mathematical Biology, 64(6):933–949, May 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0440-6.pdf.

Smith:2012:PBP

[2027] Hal L. Smith and Horst R. Thieme. Persistence of bacteria and phages in a chemostat. Journal of Mathematical Biology, 64(6):951– 979, May 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0434-4.pdf.

Nabi:2012:TOC

[2028] Ali Nabi and Jeff Moehlis. Time optimal control of spiking neurons. Journal of Mathematical Biology, 64(6):981–1004, May 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0441-5.pdf.

Li:2012:JEM

[2029] Michael Y. Li and Hongying Shu. Joint effects of mitosis and intracellular delay on viral dynamics: two-parameter bifurcation analysis. Journal of Mathematical Biology, 64(6):1005–1020, May 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0436-2.pdf.

Taylor:2012:MPE

[2030] Michael Taylor, Péter L. Simon, Darren M. Green, Thomas House, and Istvan Z. Kiss. From Markovian to pairwise epidemic models and the performance of moment closure approximations. Journal of Mathematical Biology, 64(6):1021–1042, May 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0443-3.pdf.

Northcott:2012:CPV

[2031] Katherine Northcott, Mudassar Imran, and Gail S. K. Wolkowicz. Competition in the presence of a virus in an aquatic system: an SIS model in the chemostat. Journal of Mathematical Biology, 64(6):1043– 1086, May 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0439-z.pdf.

Diao:2012:EDT

[2032] Y. Diao, K. Hinson, R. Kaplan, M. Vazquez, and J. Arsuaga. The effects of density on the topological structure of the mitochondrial DNA from trypanosomes. Journal of Mathematical Biology, 64(6):1087– 1108, May 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0438-0.pdf.

Hautphenne:2012:MBT

[2033] Sophie Hautphenne and Guy Latouche. The Markovian binary tree applied to demography. Journal of Mathematical Biology, 64(7):1109– 1135, June 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0437-1.pdf.

Priklopil:2012:IBU

[2034] Tadeas Priklopil. On invasion boundaries and the unprotected coexistence of two strategies. Journal of Mathematical Biology, 64(7):1137– 1156, June 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0448-y.pdf.

Stroyan:2012:VDM

[2035] Keith Stroyan and Mark Nawrot. Visual depth from motion parallax and eye pursuit. Journal of Mathematical Biology, 64(7):1157– 1188, June 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0445-1.pdf.

Mirrahimi:2012:EST

[2036] Sepideh Mirrahimi, Benoît Perthame, and Joe Yuichiro Wakano. Evolution of species trait through resource competition. Journal of Mathematical Biology, 64(7):1189–1223, June 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-011-0447-z.pdf.

Weng:2012:PST

[2037] J. F. Weng, I. Mareels, and D. A. Thomas. Probability Steiner trees and maximum parsimony in phylogenetic analysis. Journal of Mathematical Biology, 64(7):1225–1251, June 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0442-4.pdf.

Hodge:2012:CMN

[2038] Neil Hodge and Panayiotis Papadopoulos. Continuum modeling and numerical simulation of cell motility. Journal of Mathematical Biology, 64 (7):1253–1279, June 2012. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-011-0446-0.pdf.

Berec:2012:DIS

[2039] Ludek Berec and Daniel Maxin. Double impact of sterilizing pathogens: added value of increased life expectancy on pest control effectiveness. Journal of Mathematical Biology, 64(7):1281–1311, June 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0449-x.pdf.

Rosenbaum:2012:FVA

[2040] Robert Rosenbaum, Fabien Marpeau, Jianfu Ma, Aditya Barua, and Kresimir Josić. Finite volume and asymptotic methods for stochastic neuron models with correlated inputs. Journal of Mathematical Biology, 65(1):1–34, July 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0451-3.pdf.

Eftimie:2012:HKM

[2041] Raluca Eftimie. Hyperbolic and kinetic models for self-organized biological aggregations and movement: a brief review. Journal of Mathematical Biology, 65(1):35–75, July 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0452-2.pdf.

Dress:2012:LPT

[2042] Andreas W. M. Dress, Katharina T. Huber, and Mike Steel. ‘Lassoing’ a phylogenetic tree i: basic properties, shellings, and covers. Journal of Mathematical Biology, 65(1):77–105, July 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-011-0450-4.pdf.

Conradi:2012:MMA

[2043] Carsten Conradi and Dietrich Flockerzi. Multistationarity in mass action networks with applications to ERK activation. Journal of Mathematical Biology, 65(1):107–156, July 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0453-1.pdf.

Gambette:2012:EPN

[2044] Philippe Gambette and Katharina T. Huber. On encodings of phylogenetic networks of bounded level. Journal of Mathematical Biology, 65(1):

157–180, July 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0456-y.pdf.

Lang:2012:STP

[2045] John Lang and Michael Y. Li. Stable and transient periodic oscillations in a mathematical model for CTL response to HTLV-I infection. Journal of Mathematical Biology, 65(1):181–199, July 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0455-z.pdf.

Yang:2012:EHR

[2046] Yiding Yang, Zhilan Feng, Dashun Xu, Gregory J. Sandland, and Dennis J. Minchella. Evolution of host resistance to parasite infection in the snail-schistosome-human system. Journal of Mathematical Biology, 65(2):201–236, August 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0457-x.pdf.

Katriel:2012:SEP

[2047] Guy Katriel. The size of epidemics in populations with heterogeneous susceptibility. Journal of Mathematical Biology, 65(2):237–262, August 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0460-2.pdf.

Terry:2012:MMP

[2048] Emmanuelle Terry, Jacqueline Marvel, Christophe Arpin, Olivier Gandrillon, and Fabien Crauste. Mathematical model of the primary CD8 T cell immune response: stability analysis of a nonlinear agestructured system. Journal of Mathematical Biology, 65(2):263–291, August 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0459-8.pdf.

Fischer:2012:NHM

[2049] Mareike Fischer. Non-hereditary maximum parsimony trees. Journal of Mathematical Biology, 65(2):293–308, August 2012. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0458-9.pdf.

Inaba:2012:NPB

[2050] Hisashi Inaba. On a new perspective of the basic reproduction number in heterogeneous environments. Journal of Mathematical Biology, 65(2): 309–348, August 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0463-z.pdf.

ElKhatib:2012:RDM

[2051] N. El Khatib, S. Genieys, B. Kazmierczak, and V. Volpert. Reaction– diffusion model of atherosclerosis development. Journal of Mathematical Biology, 65(2):349–374, August 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0461-1.pdf.

Domijan:2012:IGS

[2052] Mirela Domijan and Elisabeth Pécou. The interaction graph structure of mass-action reaction networks. Journal of Mathematical Biology, 65(2): 375–402, August 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0462-0.pdf.

Jin:2012:SIP

[2053] Yu Jin and Mark A. Lewis. Seasonal influences on population spread and persistence in streams: spreading speeds. Journal of Mathematical Biology, 65(3):403–439, September 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0465-x.pdf.

Smith:2012:ICS

[2054] Aaron M. Smith, Ruth E. Baker, David Kay, and Philip K. Maini. Incorporating chemical signalling factors into cell-based models of growing epithelial tissues. Journal of Mathematical Biology, 65(3):441–463, September 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0464-y.pdf.

Lee:2012:BOS

[2055] Keum W. Lee and Sahjendra N. Singh. Bifurcation of orbits and synchrony in inferior olive neurons. Journal of Mathematical Biology, 65(3): 465–491, September 2012. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-011-0466-9.pdf.

Bokes:2012:MSM

[2056] Pavol Bokes, John R. King, Andrew T. A. Wood, and Matthew Loose. Multiscale stochastic modelling of gene expression. Journal of Mathematical Biology, 65(3):493–520, September 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-011-0468-7.pdf.

Liu:2012:MDW

[2057] Rongsong Liu, Stephen A. Gourley, Donald L. DeAngelis, and John P. Bryant. Modeling the dynamics of woody plant-herbivore interactions with age-dependent toxicity. Journal of Mathematical Biology, 65(3):521– 552, September 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0470-0.pdf.

Crawford:2012:TPG

[2058] Forrest W. Crawford and Marc A. Suchard. Transition probabilities for general birth-death processes with applications in ecology, genetics, and evolution. Journal of Mathematical Biology, 65(3):553–580, September 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0471-z.pdf.

Clote:2012:EDB

[2059] Peter Clote, Yann Ponty, and Jean-Marc Steyaert. Expected distance between terminal nucleotides of RNA secondary structures. Journal of Mathematical Biology, 65(3):581–599, September 2012. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0467-8.pdf.

Bacaer:2012:BID

[2060] Nicolas Bacaër and El Hadi Ait Dads. On the biological interpretation of a definition for the parameter R_{0} in periodic population models. Journal of Mathematical Biology, 65(4):601–621, October 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0479-4.pdf.

Lou:2012:TVD

[2061] Jie Lou, Yijun Lou, and Jianhong Wu. Threshold virus dynamics with impulsive antiretroviral drug effects. Journal of Mathematical Biology, 65(4):623–652, October 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0474-9.pdf.

Li:2012:IBM

[2062] Yibao Li, Ana Yun, and Junseok Kim. An immersed boundary method for simulating a single axisymmetric cell growth and division. Journal of Mathematical Biology, 65(4):653–675, October 2012. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0476-7.pdf.

Jourdain:2012:LFE

[2063] Benjamin Jourdain, Sylvie Méléard, and Wojbor A. Woyczynski. Lévy flights in evolutionary ecology. Journal of Mathematical Biology, 65(4): 677–707, October 2012. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-011-0478-5.pdf.

Li:2012:GMO

[2064] Yongfeng Li. A generic model for open signaling cascades with forward activation. Journal of Mathematical Biology, 65(4):709–742, October 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0480-y.pdf.

Band:2012:MMA

[2065] L. R. Band and J. R. King. Multiscale modelling of auxin transport in the plant-root elongation zone. Journal of Mathematical Biology, 65(4):743– 785, October 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0472-y.pdf.

Zhao:2012:GDR

[2066] Xiao-Qiang Zhao. Global dynamics of a reaction and diffusion model for Lyme disease. Journal of Mathematical Biology, 65(4):787–808, October 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0482-9.pdf.

Dumont:2012:MSS

[2067] Y. Dumont and J. M. Tchuenche. Mathematical studies on the sterile insect technique for the Chikungunya disease and Aedes albopictus. Journal of Mathematical Biology, 65(5):809–854, November 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0477-6.pdf.

Li:2012:TWS

[2068] Bingtuan Li. Traveling wave solutions in a plant population model with a seed bank. Journal of Mathematical Biology, 65(5):855–873, November 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0481-x.pdf.

Mori:2012:MPP

[2069] Yoichiro Mori. Mathematical properties of pump-leak models of cell volume control and electrolyte balance. Journal of Mathematical Biology, 65(5):875–918, November 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0483-8.pdf.

Ramirez:2012:PPU

[2070] Jorge M. Ramirez. Population persistence under advection–diffusion in river networks. Journal of Mathematical Biology, 65(5):919–942, November 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0485-6.pdf.

Cantrell:2012:ESI

[2071] Robert Stephen Cantrell, Chris Cosner, and Yuan Lou. Evolutionary stability of ideal free dispersal strategies in patchy environments. Journal of Mathematical Biology, 65(5):943–965, November 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0486-5.pdf.

Vermolen:2012:FEM

[2072] F. J. Vermolen and E. Javierre. A finite-element model for healing of cutaneous wounds combining contraction, angiogenesis and closure. Journal of Mathematical Biology, 65(5):967–996, November 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-011-0487-4.pdf.

Liz:2012:HEB

[2073] Eduardo Liz and Alfonso Ruiz-Herrera. The hydra effect, bubbles, and chaos in a simple discrete population model with constant effort harvesting. Journal of Mathematical Biology, 65(5):997–1016, November 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0489-2.pdf.

Kang:2012:NMC

[2074] Hye-Won Kang, Likun Zheng, and Hans G. Othmer. A new method for choosing the computational cell in stochastic reaction–diffusion systems. Journal of Mathematical Biology, 65(6–7):1017–1099, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0469-6.pdf.

Alzahrani:2012:RIB

[2075] Ebraheem O. Alzahrani, Fordyce A. Davidson, and Niall Dodds. Reversing invasion in bistable systems. Journal of Mathematical Biology, 65(6–7):1101–1124, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0490-9.pdf.

Gourley:2012:DEM

[2076] Stephen A. Gourley and Shigui Ruan. A delay equation model for oviposition habitat selection by mosquitoes. Journal of Mathematical Biology, 65(6–7):1125–1148, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0491-8.pdf.

Thomas:2012:GSM

[2077] Peter J. Thomas and Jack D. Cowan. Generalized spin models for coupled cortical feature maps obtained by coarse graining correlation based synaptic learning rules. Journal of Mathematical Biology, 65(6–7):1149– 1186, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0484-7.pdf.

Mishra:2012:KTU

[2078] Rama Mishra and Shantha Bhushan. Knot theory in understanding proteins. Journal of Mathematical Biology, 65(6–7):1187–1213, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0488-3.pdf.

Ratz:2012:TIM

[2079] Andreas Rätz and Matthias Röger. Turing instabilities in a mathematical model for signaling networks. Journal of Mathematical Biology, 65(6–7):1215–1244, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0495-4.pdf.

Hanes:2012:MRV

[2080] Douglas A. Hanes. Mathematical requirements of visual-vestibular integration. Journal of Mathematical Biology, 65(6–7):1245–1266, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0494-5.pdf.

Moulton:2012:BEC

[2081] Vincent Moulton and Mike Steel. The ‘butterfly effect’ in Cayley graphs with applications to genomics. Journal of Mathematical Biology, 65(6–7): 1267–1284, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-011-0498-1.pdf.

Leviyang:2012:ASP

[2082] Sivan Leviyang. Analysis of a stochastic predator–prey model with applications to intrahost HIV genetic diversity. Journal of Mathematical Biology, 65(6–7):1285–1336, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-011-0497-2.pdf.

Clote:2012:PNR

[2083] Peter Clote, Stefan Dobrev, Ivan Dotu, Evangelos Kranakis, Danny Krizanc, and Jorge Urrutia. On the page number of RNA secondary structures with pseudoknots. Journal of Mathematical Biology, 65(6–7): 1337–1357, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-011-0493-6.pdf.

Henkel:2012:MSP

[2084] A. Henkel, J. Müller, and C. Pötzsche. Modeling the spread of phytophthora. Journal of Mathematical Biology, 65(6–7):1359–1385, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0492-7.pdf.

Guo:2012:TDI

[2085] Zhiming Guo, Feng-Bin Wang, and Xingfu Zou. Threshold dynamics of an infective disease model with a fixed latent period and nonlocal infections. Journal of Mathematical Biology, 65(6–7):1387–1410, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0500-y.pdf.

Geritz:2012:MEW

[2086] Stefan A. H. Geritz and Éva Kisdi. Mathematical ecology: why mechanistic models? Journal of Mathematical Biology, 65(6–7):1411–1415, December 2012. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0496-3.pdf.

Tang:2013:TCI

[2087] Sanyi Tang, Juhua Liang, Yuanshun Tan, and Robert A. Cheke. Threshold conditions for integrated pest management models with pesticides that have residual effects. Journal of Mathematical Biology, 66(1–2): 1–35, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-011-0501-x.pdf.

Thatte:2013:RPS

[2088] Bhalchandra D. Thatte. Reconstructing pedigrees: some identifiability questions for a recombination–mutation model. Journal of Mathematical Biology, 66(1–2):37–74, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0503-8.pdf.

Ma:2013:EDH

[2089] Junling Ma, P. van den Driessche, and Frederick H. Willeboordse. Effective degree household network disease model. Journal of Mathematical Biology, 66(1–2):75–94, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-011-0502-9.pdf.

Bichsel:2013:EFE

[2090] Manuel Bichsel, A. D. Barbour, and Andreas Wagner. Estimating the fitness effect of an insertion sequence. Journal of Mathematical Biology, 66(1–2):95–114, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0504-2.pdf.

Blyuss:2013:ESD

[2091] Konstantin B. Blyuss. The effects of symmetry on the dynamics of antigenic variation. Journal of Mathematical Biology, 66(1–2):115–137, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0508-y.pdf.

Lin:2013:LPT

[2092] Kevin K. Lin, Kyle C. A. Wedgwood, Stephen Coombes, and Lai-Sang Young. Limitations of perturbative techniques in the analysis of rhythms and oscillations. Journal of Mathematical Biology, 66(1–2):139– 161, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0506-0.pdf.

Tosin:2013:IBV

[2093] Andrea Tosin. Initial/ boundary-value problems of tumor growth within a host tissue. Journal of Mathematical Biology, 66(1–2):163– 202, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0505-1.pdf.

Durzinsky:2013:REP

[2094] Markus Durzinsky, Wolfgang Marwan, and Annegret Wagler. Reconstruction of extended Petri nets from time-series data by using logical control functions. Journal of Mathematical Biology, 66(1–2):203– 223, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0511-3.pdf.

Rueffler:2013:WLC

[2095] Claus Rueffler, Johan A. J. Metz, and Tom J. M. Van Dooren. What life cycle graphs can tell about the evolution of life histories. Journal of Mathematical Biology, 66(1–2):225–279, January 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0509-x.pdf.

Feliu:2013:VEP

[2096] Elisenda Feliu and Carsten Wiuf. Variable elimination in posttranslational modification reaction networks with mass-action kinetics. Journal of Mathematical Biology, 66(1–2):281–310, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-012-0510-4.pdf.

Bertuzzi:2013:OSC

[2097] A. Bertuzzi, C. Bruni, F. Papa, and C. Sinisgalli. Optimal solution for a cancer radiotherapy problem. Journal of Mathematical Biology, 66(1–2): 311–349, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-012-0512-2.pdf.

Prokharau:2013:SAP

[2098] Pavel Prokharau and Fred Vermolen. Stability analysis for a peri-implant osseointegration model. Journal of Mathematical Biology, 66(1–2):351– 382, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0513-1.pdf.

Ding:2013:ESS

[2099] Zhanwen Ding, Shuxun Wang, and Honglin Yang. Evolutionarily stable strategy and invader strategy in matrix games. Journal of Mathematical Biology, 66(1–2):383–397, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0516-y.pdf.

Hellmuth:2013:ORS

[2100] Marc Hellmuth, Maribel Hernandez-Rosales, Katharina T. Huber, Vincent Moulton, Peter F. Stadler, and Nicolas Wieseke. Orthology relations, symbolic ultrametrics, and cographs. Journal of Mathematical Biology, 66(1–2):399–420, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0525-x.pdf.

Ratz:2013:ETI

[2101] Andreas Rätz and Matthias Röger. Erratum to: Turing instabilities in a mathematical model for signaling networks. Journal of Mathematical Biology, 66(1–2):421–422, January 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0630-x.pdf.

Evans:2013:SPG

[2102] Steven N. Evans, Peter L. Ralph, Sebastian J. Schreiber, and Arnab Sen. Stochastic population growth in spatially heterogeneous environments. Journal of Mathematical Biology, 66(3):423–476, February 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0514-0.pdf.

Martini:2013:MVD

[2103] Johannes W. R. Martini and G. Matthias Ullmann. A mathematical view on the decoupled sites representation. Journal of Mathematical Biology, 66(3):477–503, February 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0517-x.pdf.

Depperschmidt:2013:BRP

[2104] A. Depperschmidt, N. Ketterer, and P. Pfaffelhuber. A Brownian ratchet for protein translocation including dissociation of ratcheting sites. Journal of Mathematical Biology, 66(3):505–534, February 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0519-8.pdf.

Breban:2013:REP

[2105] Romulus Breban. Role of environmental persistence in pathogen transmission: a mathematical modeling approach. Journal of Mathematical Biology, 66(3):535–546, February 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0520-2.pdf.

George:2013:MMN

[2106] Uduak Z. George, Angélique Stéphanou, and Anotida Madzvamuse. Mathematical modelling and numerical simulations of actin dynamics in the eukaryotic cell. Journal of Mathematical Biology, 66(3):547– 593, February 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0521-1.pdf.

Serna:2013:AUB

[2107] Susana Serna, Jasmine A. Nirody, and Miklós Z. Rácz. Analysis of unstable behavior in a mathematical model for erythropoiesis. Journal of Mathematical Biology, 66(3):595–625, February 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0524-y.pdf.

Bertuzzi:2013:EOS

[2108] A. Bertuzzi, C. Bruni, F. Papa, and C. Sinisgalli. Erratum to: Optimal solution for a cancer radiotherapy problem. Journal of Mathematical Biology, 66(3):627–630, February 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0637-3.pdf.

Gyllenberg:2013:HOD

[2109] Mats Gyllenberg. In honour of odo diekmann on the occasion of his 65th birthday. Journal of Mathematical Biology, 66(4–5):631– 634, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0641-2.pdf.

Metz:2013:CIE

[2110] J. A. J. Metz. On the concept of individual in ecology and evolution. Journal of Mathematical Biology, 66(4–5):635–647, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0610-1.pdf.

Hadeler:2013:QEH

[2111] K. P. Hadeler. Quiescence, excitability, and heterogeneity in ecological models. Journal of Mathematical Biology, 66(4–5):649–684, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0590-1.pdf.

Shaw:2013:EIB

[2112] Allison K. Shaw and Simon A. Levin. The evolution of intermittent breeding. Journal of Mathematical Biology, 66(4–5):685–703, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0603-0.pdf.

Geritz:2013:GDP

[2113] S. A. H. Geritz and M. Gyllenberg. Group defence and the predator’s functional response. Journal of Mathematical Biology, 66(4–5): 705–717, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0617-7.pdf.

Benson:2013:CMS

[2114] Neil Benson, Piet H. van der Graaf, and Lambertus A. Peletier. Crossmembrane signal transduction of receptor tyrosine kinases (RTKs): from systems biology to systems pharmacology. Journal of Mathematical Biology, 66(4–5):719–742, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0620-z.pdf.

Berestycki:2013:ILF

[2115] Henri Berestycki, Jean-Michel Roquejoffre, and Luca Rossi. The influence of a line with fast diffusion on Fisher–KPP propagation. Journal of Mathematical Biology, 66(4–5):743–766, March 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0604-z.pdf.

Stankova:2013:IPD

[2116] Katerina Stanková, Alessandro Abate, and Maurice W. Sabelis. Irreversible prey diapause as an optimal strategy of a physiologically extended Lotka–Volterra model. Journal of Mathematical Biology, 66(4–5): 767–794, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0599-5.pdf.

Kooijman:2013:YEP

[2117] S. A. L. M. Kooijman. Yolky eggs prepare for metabolic acceleration. Journal of Mathematical Biology, 66(4–5):795–805, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0597-7.pdf.

Nakaoka:2013:SSS

[2118] Shinji Nakaoka and Kazuyuki Aihara. Stochastic simulation of structured skin cell population dynamics. Journal of Mathematical Biology, 66(4–5): 807–835, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0618-6.pdf.

vanGils:2013:LBN

[2119] S. A. van Gils, S. G. Janssens, Yu. A. Kuznetsov, and S. Visser. On local bifurcations in neural field models with transmission delays. Journal of Mathematical Biology, 66(4–5):837–887, March 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0598-6.pdf.

DeRoos:2013:OSA

[2120] André M. De Roos, Johan A. J. Metz, and Lennart Persson. Ontogenetic symmetry and asymmetry in energetics. Journal of Mathematical Biology, 66(4–5):889–914, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0583-0.pdf.

Metz:2013:DIB

[2121] J. A. J. Metz and Viet Chi Tran. Daphnias: from the individual based model to the large population equation. Journal of Mathematical Biology, 66(4–5):915–933, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0619-5.pdf.

Corson:2013:TSO

[2122] S. Corson, D. Greenhalgh, and S. J. Hutchinson. A time since onset of injection model for hepatitis C spread amongst injecting drug users. Journal of Mathematical Biology, 66(4–5):935–978, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0577-y.pdf.

Ball:2013:NTC

[2123] Frank Ball, Tom Britton, and David Sirl. A network with tunable clustering, degree correlation and degree distribution, and an epidemic thereon. Journal of Mathematical Biology, 66(4–5):979–1019, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0609-7.pdf.

Kisdi:2013:CMS

[2124] Éva Kisdi and Barbara Boldin. A construction method to study the role of incidence in the adaptive dynamics of pathogens with direct and environmental transmission. Journal of Mathematical Biology, 66(4–5):1021– 1044, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0563-4.pdf.

Roberts:2013:CNG

[2125] M. G. Roberts and J. A. P. Heesterbeek. Characterizing the nextgeneration matrix and basic reproduction number in ecological epidemiology. Journal of Mathematical Biology, 66(4–5):1045–1064, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-012-0602-1.pdf.

Inaba:2013:DCT

[2126] Hisashi Inaba. On the definition and the computation of the typereproduction number T for structured populations in heterogeneous environments. Journal of Mathematical Biology, 66(4–5):1065–1097, March 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0522-0.pdf.

Rueffler:2013:NSC

[2127] Claus Rueffler and Johan A. J. Metz. Necessary and sufficient conditions for R_{0} to be a sum of contributions of fertility loops. Journal of Mathematical Biology, 66(4–5):1099–1122, March 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0575-0.pdf.

Borovkov:2013:HHD

[2128] K. Borovkov, R. Day, and T. Rice. High host density favors greater virulence: a model of parasite-host dynamics based on multi-type branching processes. Journal of Mathematical Biology, 66(6):1123– 1153, May 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0526-9.pdf.

Wang:2013:DPP

[2129] Yuanshi Wang. Dynamics of plant-pollinator-robber systems. Journal of Mathematical Biology, 66(6):1155–1177, May 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-012-0527-8.pdf.

Schuff:2013:MTMa

[2130] M. M. Schuff, J. P. Gore, and E. A. Nauman. A mixture theory model of fluid and solute transport in the microvasculature of normal and malignant tissues. I. Theory. Journal of Mathematical Biology, 66(6):1179– 1207, May 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0528-7.pdf.

Fuertinger:2013:MEA

[2131] Doris H. Fuertinger, Franz Kappel, Stephan Thijssen, Nathan W. Levin, and Peter Kotanko. A model of erythropoiesis in adults with sufficient iron availability. Journal of Mathematical Biology, 66(6):1209– 1240, May 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0530-0.pdf.

Wang:2013:STL

[2132] Xuefeng Wang and Qian Xu. Spiky and transition layer steady states of chemotaxis systems via global bifurcation and Helly’s compactness theorem. Journal of Mathematical Biology, 66(6):1241– 1266, May 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0533-x.pdf.

Boissard:2013:TFB

[2133] Emmanuel Boissard, Pierre Degond, and Sebastien Motsch. Trail formation based on directed pheromone deposition. Journal of Mathematical Biology, 66(6):1267–1301, May 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0529-6.pdf.

Faye:2013:LSU

[2134] Grégory Faye, James Rankin, and Pascal Chossat. Localized states in an unbounded neural field equation with smooth firing rate function: a multi-parameter analysis. Journal of Mathematical Biology, 66(6):1303– 1338, May 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0532-y.pdf.

Wang:2013:EIE

[2135] Mei-Hui Wang, Mark Kot, and Michael G. Neubert. Erratum to: Integrodifference equations, Allee effects, and invasions. Journal of Mathematical Biology, 66(6):1339–1340, May 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0643-0.pdf.

Bearon:2013:HSC

[2136] R. N. Bearon. Helical swimming can provide robust upwards transport for gravitactic single-cell algae; a mechanistic model. Journal of Mathematical Biology, 66(7):1341–1359, June 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0531-z.pdf.

Malerba:2013:PRR

[2137] Paola Malerba and Nancy Kopell. Phase resetting reduces theta-gamma rhythmic interaction to a one-dimensional map. Journal of Mathematical Biology, 66(7):1361–1386, June 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0534-9.pdf.

Clarelli:2013:FDM

[2138] F. Clarelli, C. Di Russo, R. Natalini, and M. Ribot. A fluid dynamics model of the growth of phototrophic biofilms. Journal of Mathematical Biology, 66(7):1387–1408, June 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0538-5.pdf.

DeMatteis:2013:RSC

[2139] Giovanni De Matteis, Alex Graudenzi, and Marco Antoniotti. A review of spatial computational models for multi-cellular systems, with regard to intestinal crypts and colorectal cancer development. Journal of Mathematical Biology, 66(7):1409–1462, June 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0539-4.pdf.

Roberts:2013:EMU

[2140] M. G. Roberts. Epidemic models with uncertainty in the reproduction number. Journal of Mathematical Biology, 66(7):1463–1474, June 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0540-y.pdf.

Robinson:2013:NFT

[2141] P. A. Robinson. Neural field theory with variance dynamics. Journal of Mathematical Biology, 66(7):1475–1497, June 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0541-x.pdf.

Bressloff:2013:PCT

[2142] Paul C. Bressloff. Propagation of CaMKII translocation waves in heterogeneous spiny dendrites. Journal of Mathematical Biology, 66(7):1499– 1525, June 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0542-9.pdf.

Reluga:2013:GAD

[2143] Timothy C. Reluga and Jing Li. Games of age-dependent prevention of chronic infections by social distancing. Journal of Mathematical Biology, 66(7):1527–1553, June 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0543-8.pdf.

Clairambault:2013:CTH

[2144] Jean Clairambault. Can theorems help treat cancer? Journal of Mathematical Biology, 66(7):1555–1558, June 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0518-9.pdf.

Lassas:2013:FIP

[2145] Matti Lassas and Samuli Siltanen. Foreword: inverse problems in biology. Journal of Mathematical Biology, 67(1):1, July 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0700-8.pdf.

Raol:2013:IPN

[2146] Jay Raol and Steven J. Cox. Inverse problems in neuronal calcium signaling. Journal of Mathematical Biology, 67(1):3–23, July 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0507-z.pdf.

Capasso:2013:SIP

[2147] V. Capasso, H. E. Kunze, D. La Torre, and E. R. Vrscay. Solving inverse problems for biological models using the collage method for differential equations. Journal of Mathematical Biology, 67(1):25– 38, July 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0515-z.pdf.

Olufsen:2013:PAP

[2148] Mette S. Olufsen and Johnny T. Ottesen. A practical approach to parameter estimation applied to model predicting heart rate regulation. Journal of Mathematical Biology, 67(1):39–68, July 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0535-8.pdf.

Doumic:2013:EDR

[2149] M. Doumic and Léon M. Tine. Estimating the division rate for the growth-fragmentation equation. Journal of Mathematical Biology, 67(1): 69–103, July 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0553-6.pdf.

Anderssen:2013:IRM

[2150] Robert S. Anderssen and Christopher A. Helliwell. Information recovery in molecular biology: causal modelling of regulated promoter switching experiments. Journal of Mathematical Biology, 67(1):105– 122, July 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0536-7.pdf.

Avdonin:2013:DDP

[2151] Sergei Avdonin and Jonathan Bell. Determining a distributed parameter in a neural cable model via a boundary control method. Journal of Mathematical Biology, 67(1):123–141, July 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0537-6.pdf.

Lu:2013:IPB

[2152] James Lu, Elias August, and Heinz Koeppl. Inverse problems from biomedicine. Journal of Mathematical Biology, 67(1):143–168, July 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-012-0523-z.pdf.

Doebeli:2013:SCG

[2153] Michael Doebeli and Iaroslav Ispolatov. Symmetric competition as a general model for single-species adaptive dynamics. Journal of Mathematical Biology, 67(2):169–184, August 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0547-4.pdf.

Cai:2013:MBE

[2154] Linlin Cai, Guoting Chen, and Dongmei Xiao. Multiparametric bifurcations of an epidemiological model with strong Allee effect. Journal of Mathematical Biology, 67(2):185–215, August 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0546-5.pdf.

Koch:2013:ERR

[2155] Dean Koch, Reinhard Illner, and Junling Ma. Edge removal in random contact networks and the basic reproduction number. Journal of Mathematical Biology, 67(2):217–238, August 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0545-6.pdf.

Ditlevsen:2013:MLN

[2156] Susanne Ditlevsen and Priscilla Greenwood. The Morris–Lecar neuron model embeds a leaky integrate-and-fire model. Journal of Mathematical Biology, 67(2):239–259, August 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0552-7.pdf.

Kuang:2013:MNP

[2157] Juhong Kuang, Moxun Tang, and Jianshe Yu. The mean and noise of protein numbers in stochastic gene expression. Journal of Mathematical Biology, 67(2):261–291, August 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0551-8.pdf.

ORegan:2013:CSP

[2158] Suzanne M. O’Regan, Thomas C. Kelly, Andrei Korobeinikov, Michael J. A. O’Callaghan, Alexei V. Pokrovskii, and Dmitrii Rachinskii. Chaos in a seasonally perturbed SIR model: avian influenza in a seabird colony as a paradigm. Journal of Mathematical Biology, 67(2):293– 327, August 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0550-9.pdf.

Cressman:2013:TPP

[2159] Ross Cressman and Vlastimil Krivan. Two-patch population models with adaptive dispersal: the effects of varying dispersal speeds. Journal of Mathematical Biology, 67(2):329–358, August 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0548-3.pdf.

Nagaiah:2013:OCA

[2160] Chamakuri Nagaiah, Karl Kunisch, and Gernot Plank. Optimal control approach to termination of re-entry waves in cardiac electrophysiology. Journal of Mathematical Biology, 67(2):359–388, August 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0557-2.pdf.

Kong:2013:LDS

[2161] Yong Kong. Length distribution of sequencing by synthesis: fixed flow cycle model. Journal of Mathematical Biology, 67(2):389–410, August 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0556-3.pdf.

Juher:2013:OAS

[2162] David Juher, Jordi Ripoll, and Joan Saldaña. Outbreak analysis of an SIS epidemic model with rewiring. Journal of Mathematical Biology, 67(2): 411–432, August 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0555-4.pdf.

Shao:2013:SDS

[2163] Jianfeng Shao, Xiaohua Yan, and Shuo Shao. SNR of DNA sequences mapped by general affine transformations of the indicator sequences. Journal of Mathematical Biology, 67(2):433–451, August 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0564-3.pdf.

Dumont:2013:PDM

[2164] Grégory Dumont and Jacques Henry. Population density models of integrate-and-fire neurons with jumps: well-posedness. Journal of Mathematical Biology, 67(3):453–481, September 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-012-0554-5.pdf.

Mummert:2013:SRP

[2165] Anna Mummert. Studying the recovery procedure for the time-dependent transmission rate(s) in epidemic models. Journal of Mathematical Biology, 67(3):483–507, September 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0558-1.pdf.

Parvinen:2013:FVA

[2166] Kalle Parvinen, Mikko Heino, and Ulf Dieckmann. Function-valued adaptive dynamics and optimal control theory. Journal of Mathematical Biology, 67(3):509–533, September 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0549-2.pdf.

Xiao:2013:EDR

[2167] MingQing Xiao, John D. Reeve, Dashun Xu, and James T. Cronin. Estimation of the diffusion rate and crossing probability for biased edge movement between two different types of habitat. Journal of Mathematical Biology, 67(3):535–567, September 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0561-6.pdf.

Collet:2013:RMS

[2168] Pierre Collet, Sylvie Méléard, and Johan A. J. Metz. A rigorous model study of the adaptive dynamics of Mendelian diploids. Journal of Mathematical Biology, 67(3):569–607, September 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-012-0562-5.pdf.

Catsigeras:2013:IFN

[2169] Eleonora Catsigeras and Pierre Guiraud. Integrate and fire neural networks, piecewise contractive maps and limit cycles. Journal of Mathematical Biology, 67(3):609–655, September 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0560-7.pdf.

Diego:2013:MCB

[2170] David Diego, Gabriel F. Calvo, and Víctor M. Pérez-García. Modeling the connection between primary and metastatic tumors. Journal of Mathematical Biology, 67(3):657–692, September 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0565-2.pdf.

McVinish:2013:LBS

[2171] R. McVinish and P. K. Pollett. The limiting behaviour of a stochastic patch occupancy model. Journal of Mathematical Biology, 67(3):693– 716, September 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0568-z.pdf.

Saisho:2013:MMG

[2172] Yasumasa Saisho and Atsushi Ito. Mathematical models of the generation of radiation-induced DNA double-strand breaks. Journal of Mathematical Biology, 67(3):717–736, September 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0567-0.pdf.

Weinberger:2013:ERL

[2173] Hans F. Weinberger. Erratum to: The retreat of the less fit allele in a population-controlled model for population genetics. Journal of Mathematical Biology, 67(3):737, September 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0676-4.pdf.

Mailier:2013:SIM

[2174] Johan Mailier, Marcel Remy, and Alain Vande Wouwer. Stoichiometric identification with maximum likelihood principal component analysis. Journal of Mathematical Biology, 67(4):739–765, October 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0559-0.pdf.

Mossel:2013:IIN

[2175] Elchanan Mossel and Sebastien Roch. Identifiability and inference of nonparametric rates-across-sites models on large-scale phylogenies. Journal of Mathematical Biology, 67(4):767–797, October 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0571-4.pdf.

Artalejo:2013:SEM

[2176] J. R. Artalejo, A. Economou, and M. J. Lopez-Herrero. Stochastic epidemic models with random environment: quasi-stationarity, extinction and final size. Journal of Mathematical Biology, 67(4):799– 831, October 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0570-5.pdf.

Djellouli:2013:TSF

[2177] R. Djellouli, S. Mahserejian, A. Mokrane, M. Moussaoui, and T. M. Laleg-Kirati. Theoretical study of the fibrous capsule tissue growth around a disk-shaped implant. Journal of Mathematical Biology, 67(4): 833–867, October 2013. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-012-0566-1.pdf.

Miller:2013:MHE

[2178] Joel C. Miller and Erik M. Volz. Model hierarchies in edge-based compartmental modeling for infectious disease spread. Journal of Mathematical Biology, 67(4):869–899, October 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0572-3.pdf.

Wang:2013:VDM

[2179] Yan Wang, Yicang Zhou, Fred Brauer, and Jane M. Heffernan. Viral dynamics model with CTL immune response incorporating antiretroviral therapy. Journal of Mathematical Biology, 67(4):901–934, October 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0580-3.pdf.

Argasinski:2013:ETE

[2180] K. Argasinski and M. Broom. Ecological theatre and the evolutionary game: how environmental and demographic factors determine payoffs in evolutionary games. Journal of Mathematical Biology, 67(4):935– 962, October 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0573-2.pdf.

Clancy:2013:EPH

[2181] Damian Clancy and Christopher J. Pearce. The effect of population heterogeneities upon spread of infection. Journal of Mathematical Biology, 67(4):963–987, October 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0578-x.pdf.

Desroches:2013:ICE

[2182] M. Desroches, M. Krupa, and S. Rodrigues. Inflection, canards and excitability threshold in neuronal models. Journal of Mathematical Biology, 67(4):989–1017, October 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0576-z.pdf.

Kooijman:2013:EYE

[2183] S. A. L. M. Kooijman. Erratum to: Yolky eggs prepare for metabolic acceleration. Journal of Mathematical Biology, 67(4):1019–1021, October 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0697-z.pdf.

Muller:2013:ADC

[2184] Johannes Müller and Hannes Uecker. Approximating the dynamics of communicating cells in a diffusive medium by ODEs-homogenization with localization. Journal of Mathematical Biology, 67(5):1023–1065, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0569-y.pdf.

Shuai:2013:ETR

[2185] Zhisheng Shuai, J. A. P. Heesterbeek, and P. van den Driessche. Extending the type reproduction number to infectious disease control targeting contacts between types. Journal of Mathematical Biology, 67(5):1067– 1082, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0579-9.pdf.

Safan:2013:VBC

[2186] Muntaser Safan, Mirjam Kretzschmar, and Karl P. Hadeler. Vaccination based control of infections in SIRS models with reinfection: special reference to pertussis. Journal of Mathematical Biology, 67(5):1083– 1110, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0582-1.pdf.

Kitayimbwa:2013:RBM

[2187] John M. Kitayimbwa, Joseph Y. T. Mugisha, and Roberto A. Saenz. The role of backward mutations on the within-host dynamics of HIV-1. Journal of Mathematical Biology, 67(5):1111–1139, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0581-2.pdf.

Wendl:2013:CTM

[2188] Michael C. Wendl, Karthik Kota, George M. Weinstock, and Makedonka Mitreva. Coverage theories for metagenomic DNA sequencing based on a generalization of Stevens’ theorem. Journal of Mathematical Biology, 67(5):1141–1161, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0586-x.pdf.

Hartmann:2013:ETP

[2189] Klaas Hartmann. The equivalence of two phylogenetic biodiversity measures: the Shapley value and fair proportion index. Journal of Mathematical Biology, 67(5):1163–1170, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0585-y.pdf.

Claus:2013:SAS

[2190] J. Claus, E. Friedmann, U. Klingmüller, R. Rannacher, and T. Szekeres. Spatial aspects in the SMAD signaling pathway. Journal of Mathematical Biology, 67(5):1171–1197, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0574-1.pdf.

ODea:2013:IBT

[2191] R. D. O’Dea, J. M. Osborne, A. J. El Haj, H. M. Byrne, and S. L. Waters. The interplay between tissue growth and scaffold degradation in engineered tissue constructs. Journal of Mathematical Biology, 67(5):1199– 1225, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0587-9.pdf.

Kang:2013:DIP

[2192] Yun Kang and Lauren Wedekin. Dynamics of a intraguild predation model with generalist or specialist predator. Journal of Mathematical Biology, 67(5):1227–1259, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0584-z.pdf.

Andersen:2013:TCE

[2193] J. E. Andersen, R. C. Penner, C. M. Reidys, and M. S. Waterman. Topological classification and enumeration of RNA structures by genus. Journal of Mathematical Biology, 67(5):1261–1278, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0594-x.pdf.

Draelants:2013:NBA

[2194] Delphine Draelants, Jan Broeckhove, Gerrit T. S. Beemster, and Wim Vanroose. Numerical bifurcation analysis of the pattern formation in a cell based auxin transport model. Journal of Mathematical Biology, 67(5):1279–1305, November 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0588-8.pdf.

Schuff:2013:MTMb

[2195] M. M. Schuff, J. P. Gore, and E. A. Nauman. A mixture theory model of fluid and solute transport in the microvasculature of normal and malignant tissues. II: Factor sensitivity analysis, calibration, and validation. Journal of Mathematical Biology, 67(6–7):1307–1337, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0544-7.pdf.

Rempala:2013:MDO

[2196] Grzegorz A. Rempala and Michal Seweryn. Methods for diversity and overlap analysis in T-cell receptor populations. Journal of Mathematical Biology, 67(6–7):1339–1368, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0589-7.pdf.

Achdou:2013:QMA

[2197] Yves Achdou, Bruno Franchi, Norina Marcello, and Maria Carla Tesi. A qualitative model for aggregation and diffusion of β-amyloid in Alzheimer’s disease. Journal of Mathematical Biology, 67(6–7):1369– 1392, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0591-0.pdf.

Middleton:2013:WPS

[2198] A. M. Middleton, J. R. King, and M. Loose. Wave pinning and spatial patterning in a mathematical model of Antivin/ lefty-nodal signalling. Journal of Mathematical Biology, 67(6–7):1393–1424, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0592-z.pdf.

Dunia:2013:MMV

[2199] Ricardo Dunia and Roger Bonnecaze. Mathematical modeling of viral infection dynamics in spherical organs. Journal of Mathematical Biology, 67(6–7):1425–1455, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0593-y.pdf.

Hawkins-Daarud:2013:BCV

[2200] Andrea Hawkins-Daarud, Serge Prudhomme, Kristoffer G. van der Zee, and J. Tinsley Oden. Bayesian calibration, validation, and uncertainty quantification of diffuse interface models of tumor growth. Journal of Mathematical Biology, 67(6–7):1457–1485, December 2013. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0595-9.pdf.

Setchi:2013:MMF

[2201] Adriana Setchi, A. Jonathan Mestel, Jennifer H. Siggers, Kim H. Parker, Ming Wang Tan, and Kangwen Wong. Mathematical model of flow through the patent ductus arteriosus. Journal of Mathematical Biology, 67(6–7):1487–1506, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0596-8.pdf.

Chan:2013:BSC

[2202] Bernard S. Chan and Pei Yu. Bifurcation, stability, and cluster formation of multi-strain infection models. Journal of Mathematical Biology, 67(6–7):1507–1532, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0600-3.pdf.

Cortez:2013:WDP

[2203] Michael H. Cortez. When does pathogen evolution maximize the basic reproductive number in well-mixed host-pathogen systems? Journal of Mathematical Biology, 67(6–7):1533–1585, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0601-2.pdf.

Tsai:2013:DCB

[2204] Je-Chiang Tsai. Do calcium buffers always slow down the propagation of calcium waves? Journal of Mathematical Biology, 67(6–7):1587–1632, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0605-y.pdf.

Franco:2013:SPM

[2205] Elisa Franco and Franco Blanchini. Structural properties of the MAPK pathway topologies in PC12 cells. Journal of Mathematical Biology, 67(6–7):1633–1668, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0606-x.pdf.

Bressloff:2013:DNV

[2206] Paul C. Bressloff and Yi Ming Lai. Dispersal and noise: Various modes of synchrony in ecological oscillators. Journal of Mathematical Biology, 67(6–7):1669–1690, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0607-9.pdf.

Levy:2013:DST

[2207] Chris Levy and David Iron. Dynamics and stability of a threedimensional model of cell signal transduction. Journal of Mathematical Biology, 67(6–7):1691–1728, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-012-0608-8.pdf.

Bacaer:2013:BRN

[2208] Nicolas Bacaër and Mohamed Khaladi. On the basic reproduction number in a random environment. Journal of Mathematical Biology, 67(6–7):1729–1739, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0611-0.pdf.

Taylor:2013:SFM

[2209] Rachel A. Taylor, Jonathan A. Sherratt, and Andrew White. Seasonal forcing and multi-year cycles in interacting populations: lessons from a predator–prey model. Journal of Mathematical Biology, 67(6–7):1741– 1764, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0612-z.pdf.

Saez:2013:MMC

[2210] Pablo Sáez, Estefanía Peña, Miguel Ángel Martínez, and Ellen Kuhl. Mathematical modeling of collagen turnover in biological tissue. Journal of Mathematical Biology, 67(6–7):1765–1793, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0613-y.pdf.

Berthoumieux:2013:IMN

[2211] Sara Berthoumieux, Matteo Brilli, Daniel Kahn, Hidde de Jong, and Eugenio Cinquemani. On the identifiability of metabolic network models. Journal of Mathematical Biology, 67(6–7):1795–1832, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0614-x.pdf.

Cardona:2013:EFV

[2212] Gabriel Cardona, Arnau Mir, and Francesc Rosselló. Exact formulas for the variance of several balance indices under the Yule model. Journal of Mathematical Biology, 67(6–7):1833–1846, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0615-9.pdf.

Mikhal:2013:DAV

[2213] Julia Mikhal and Bernard J. Geurts. Development and application of a volume penalization immersed boundary method for the computation of blood flow and shear stresses in cerebral vessels and aneurysms. Journal of Mathematical Biology, 67(6–7):1847–1875, December 2013. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0627-5.pdf.

Wei:2014:SCS

[2214] Juncheng Wei and Matthias Winter. Stability of cluster solutions in a cooperative consumer chain model. Journal of Mathematical Biology, 68(1– 2):1–39, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0616-8.pdf.

Hartung:2014:PNI

[2215] Niklas Hartung. Parameter non-identifiability of the Gyllenberg– Webb ODE model. Journal of Mathematical Biology, 68(1–2):41–55, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0724-0.pdf.

Zheng:2014:VMB

[2216] Xiaoming Zheng and Chunjing Xie. A viscoelastic model of blood capillary extension and regression: derivation, analysis, and simulation. Journal of Mathematical Biology, 68(1–2):57–80, January 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0624-8.pdf.

Moulton:2014:SGK

[2217] Derek E. Moulton and Alain Goriely. Surface growth kinematics via local curve evolution. Journal of Mathematical Biology, 68(1–2):81– 108, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0625-7.pdf.

Allen:2014:MSC

[2218] Benjamin Allen and Corina E. Tarnita. Measures of success in a class of evolutionary models with fixed population size and structure. Journal of Mathematical Biology, 68(1–2):109–143, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0622-x.pdf.

Barbi:2014:DTC

[2219] Maria Barbi, Julien Mozziconacci, Hua Wong, and Jean-Marc Victor. DNA topology in chromosomes: a quantitative survey and its physiological implications. Journal of Mathematical Biology, 68(1–2):145– 179, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0621-y.pdf.

Veloz:2014:RNE

[2220] Tomas Veloz, Pablo Razeto-Barry, Peter Dittrich, and Alejandro Fajardo. Reaction networks and evolutionary game theory. Journal of Mathematical Biology, 68(1–2):181–206, January 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0626-6.pdf.

Miller:2014:RLS

[2221] Judith R. Miller and Huihui Zeng. Range limits in spatially explicit models of quantitative traits. Journal of Mathematical Biology, 68(1–2):207– 234, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0628-4.pdf.

Kavian:2014:CEM

[2222] Otared Kavian, Michael Leguèbe, Clair Poignard, and Lisl Weynans. “Classical” electropermeabilization modeling at the cell scale. Journal of Mathematical Biology, 68(1–2):235–265, January 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0629-3.pdf.

Sakamoto:2014:MAP

[2223] Y. Sakamoto, S. Prudhomme, and M. H. Zaman. Modeling of adhesion, protrusion, and contraction coordination for cell migration simulations. Journal of Mathematical Biology, 68(1–2):267–302, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0634-6.pdf.

Schwemmer:2014:RPL

[2224] Michael A. Schwemmer and Timothy J. Lewis. The robustness of phaselocking in neurons with dendro-dendritic electrical coupling. Journal of Mathematical Biology, 68(1–2):303–340, January 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0635-5.pdf.

Fusy:2014:CLO

[2225] Éric Fusy and Peter Clote. Combinatorics of locally optimal RNA secondary structures. Journal of Mathematical Biology, 68(1–2):341– 375, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0631-9.pdf.

Castella:2014:CPG

[2226] François Castella and Sten Madec. Coexistence phenomena and global bifurcation structure in a chemostat-like model with species-dependent diffusion rates. Journal of Mathematical Biology, 68(1–2):377–415, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0633-7.pdf.

Wang:2014:DAI

[2227] Chuncheng Wang, Stephen A. Gourley, and Rongsong Liu. Delayed action insecticides and their role in mosquito and malaria control. Journal of Mathematical Biology, 68(1–2):417–451, January 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-012-0638-2.pdf.

Ackleh:2014:RUP

[2228] Azmy S. Ackleh and Paul L. Salceanu. Robust uniform persistence and competitive exclusion in a nonautonomous multi-strain SIR epidemic model with disease-induced mortality. Journal of Mathematical Biology, 68(1–2):453–475, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0636-4.pdf.

Shu:2014:STO

[2229] Hongying Shu, Lin Wang, and James Watmough. Sustained and transient oscillations and chaos induced by delayed antiviral immune response in an immunosuppressive infection model. Journal of Mathematical Biology, 68(1–2):477–503, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0639-1.pdf.

Planque:2014:MBS

[2230] Robert Planqué, Nicholas F. Britton, and Hans Slabbekoorn. On the maintenance of bird song dialects. Journal of Mathematical Biology, 68(1–2):505–531, January 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0632-8.pdf.

Bacaer:2014:PEP

[2231] Nicolas Bacaër and El Hadi Ait Dads. On the probability of extinction in a periodic environment. Journal of Mathematical Biology, 68(3):533– 548, February 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-012-0623-9.pdf.

Cobbold:2014:MOT

[2232] Christina A. Cobbold and Frithjof Lutscher. Mean occupancy time: linking mechanistic movement models, population dynamics and landscape ecology to population persistence. Journal of Mathematical Biology, 68(3):549–579, February 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0642-1.pdf.

Riaz:2014:DSP

[2233] Syed Shahed Riaz and Michael C. Mackey. Dynamic spatial pattern formation in the sea urchin embryo. Journal of Mathematical Biology, 68(3):581–608, February 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-012-0640-8.pdf.

Lopez-Caamal:2014:ESC

[2234] Fernando López-Caamal, Richard H. Middleton, and Heinrich J. Huber. Equilibria and stability of a class of positive feedback loops. Journal of Mathematical Biology, 68(3):609–645, February 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0644-z.pdf.

Seguin:2014:MDC

[2235] Brian Seguin and Eliot Fried. Microphysical derivation of the canham– Helfrich free-energy density. Journal of Mathematical Biology, 68(3):647– 665, February 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0647-9.pdf.

Davis:2014:TFM

[2236] Lisa Davis, Tomás Gedeon, Jakub Gedeon, and Jennifer Thorenson. A traffic flow model for bio-polymerization processes. Journal of Mathematical Biology, 68(3):667–700, February 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0651-0.pdf.

Augeraud-Veron:2014:SDS

[2237] E. Augeraud-Véron and N. Sari. Seasonal dynamics in an SIR epidemic system. Journal of Mathematical Biology, 68(3):701–725, February 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0645-y.pdf.

Huillet:2014:PGA

[2238] Thierry E. Huillet. Pareto genealogies arising from a Poisson branching evolution model with selection. Journal of Mathematical Biology, 68(3): 727–761, February 2014. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-013-0649-7.pdf.

Onyango:2014:DOV

[2239] Nelson Owuor Onyango and Johannes Müller. Determination of optimal vaccination strategies using an orbital stability threshold from periodically driven systems. Journal of Mathematical Biology, 68(3):763– 784, February 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0648-8.pdf.

Lemon:2014:MBC

[2240] Greg Lemon, Ylva Gustafsson, Johannes C. Haag, Mei L. Lim, Sebastian Sjöqvist, Fatemeh Ajalloueian, Philipp Jungebluth, and Paolo Macchiarini. Modelling biological cell attachment and growth on adherent surfaces. Journal of Mathematical Biology, 68(4):785–813, March 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0653-y.pdf.

Kloosterman:2014:CNM

[2241] Matt Kloosterman, Sue Ann Campbell, and Francis J. Poulin. A closed NPZ model with delayed nutrient recycling. Journal of Mathematical Biology, 68(4):815–850, March 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0646-x.pdf.

Lam:2014:ECD

[2242] King-Yeung Lam and Yuan Lou. Evolution of conditional dispersal: evolutionarily stable strategies in spatial models. Journal of Mathematical Biology, 68(4):851–877, March 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0650-1.pdf.

Feller:2014:PFA

[2243] C. Feller, J. P. Gabriel, C. Mazza, and F. Yerly. Pattern formation in auxin flux. Journal of Mathematical Biology, 68(4):879–909, March 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0655-9.pdf.

Zhang:2014:MMP

[2244] Zhigang Zhang, Jeffrey J. Morgan, and Paul A. Lindahl. Mathematical model for positioning the FtsZ contractile ring in Escherichia coli. Journal of Mathematical Biology, 68(4):911–930, March 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0652-z.pdf.

Mirrahimi:2014:DCR

[2245] Sepideh Mirrahimi, Benoît Perthame, and Joe Yuichiro Wakano. Direct competition results from strong competition for limited resource. Journal of Mathematical Biology, 68(4):931–949, March 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0659-5.pdf.

Svensson:2014:WWI

[2246] Åke Svensson. Who was the infector-probabilities in the presence of variability in latent and infectious times. Journal of Mathematical Biology, 68(4):951–967, March 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0658-6.pdf.

Magori:2014:ECA

[2247] Krisztian Magori and Andrew W. Park. The evolutionary consequences of alternative types of imperfect vaccines. Journal of Mathematical Biology, 68(4):969–987, March 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0654-x.pdf.

Vernerey:2014:MMC

[2248] Franck J. Vernerey and Mehdi Farsad. A mathematical model of the coupled mechanisms of cell adhesion, contraction and spreading. Journal of Mathematical Biology, 68(4):989–1022, March 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0656-8.pdf.

Ryu:2014:TFF

[2249] Hwayeon Ryu and Anita T. Layton. Tubular fluid flow and distal NaCl delivery mediated by tubuloglomerular feedback in the rat kidney. Journal of Mathematical Biology, 68(4):1023–1049, March 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0667-5.pdf.

Yvinec:2014:ARM

[2250] Romain Yvinec, Changjing Zhuge, Jinzhi Lei, and Michael C. Mackey. Adiabatic reduction of a model of stochastic gene expression with jump Markov process. Journal of Mathematical Biology, 68(5):1051– 1070, April 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0661-y.pdf.

Paniello:2014:MDG

[2251] Irene Paniello. Marginal distributions of genetic coalgebras. Journal of Mathematical Biology, 68(5):1071–1087, April 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0663-9.pdf.

Chalub:2014:FDW

[2252] Fabio A. C. C. Chalub and Max O. Souza. The frequency-dependent Wright–Fisher model: diffusive and non-diffusive approximations. Journal of Mathematical Biology, 68(5):1089–1133, April 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0657-7.pdf.

Akerman:2014:CGF

[2253] Ada Akerman and Reinhard Bürger. The consequences of gene flow for local adaptation and differentiation: a two-locus two-deme model. Journal of Mathematical Biology, 68(5):1135–1198, April 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0660-z.pdf.

McGillen:2014:GRD

[2254] Jessica B. McGillen, Eamonn A. Gaffney, Natasha K. Martin, and Philip K. Maini. A general reaction–diffusion model of acidity in cancer invasion. Journal of Mathematical Biology, 68(5):1199–1224, April 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0665-7.pdf.

Biktashev:2014:SMM

[2255] Vadim N. Biktashev. A simple mathematical model of gradual Darwinian evolution: emergence of a Gaussian trait distribution in adaptation along a fitness gradient. Journal of Mathematical Biology, 68(5): 1225–1248, April 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0669-3.pdf.

Meijer:2014:TWN

[2256] Hil G. E. Meijer and Stephen Coombes. Travelling waves in a neural field model with refractoriness. Journal of Mathematical Biology, 68(5): 1249–1268, April 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0670-x.pdf.

Souza:2014:MAV

[2257] Max O. Souza. Multiscale analysis for a vector-borne epidemic model. Journal of Mathematical Biology, 68(5):1269–1293, April 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0666-6.pdf.

Weinberger:2014:RLF

[2258] Hans F. Weinberger. The retreat of the less fit allele in a populationcontrolled model for population genetics. Journal of Mathematical Biology, 68(5):1295–1316, April 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0673-7.pdf.

Wang:2014:MRA

[2259] Wendi Wang and Xingfu Zou. Modeling the role of altruism of antibiotic-resistant bacteria. Journal of Mathematical Biology, 68(6): 1317–1339, May 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0668-4.pdf.

Kim:2014:GAS

[2260] Yong-Jung Kim, Ohsang Kwon, and Fang Li. Global asymptotic stability and the ideal free distribution in a starvation driven diffusion. Journal of Mathematical Biology, 68(6):1341–1370, May 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0674-6.pdf.

Baake:2014:SCR

[2261] Ellen Baake and Ute von Wangenheim. Single-crossover recombination and ancestral recombination trees. Journal of Mathematical Biology, 68(6):1371–1402, May 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0662-x.pdf.

Dagbovie:2014:ASD

[2262] Ayawoa S. Dagbovie and Jonathan A. Sherratt. Absolute stability and dynamical stabilisation in predator–prey systems. Journal of Mathematical Biology, 68(6):1403–1421, May 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0672-8.pdf.

Andersen:2014:ECI

[2263] Lars Nørvang Andersen, Thomas Mailund, and Asger Hobolth. Efficient computation in the IM model. Journal of Mathematical Biology, 68(6): 1423–1451, May 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0671-9.pdf.

Aston:2014:MAR

[2264] Philip J. Aston, Gianne Derks, Balaji M. Agoram, and Piet H. van der Graaf. A mathematical analysis of rebound in a target-mediated drug disposition model: I. Without feedback. Journal of Mathematical Biology, 68(6):1453–1478, May 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0675-5.pdf.

Wang:2014:SMM

[2265] Chuncheng Wang, Rongsong Liu, Junping Shi, and Carlos Martinez del Rio. Spatiotemporal mutualistic model of mistletoes and birds. Journal of Mathematical Biology, 68(6):1479–1520, May 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0664-8.pdf.

Hirt:2014:MRT

[2266] Bartholomäus V. Hirt, Jonathan A. D. Wattis, and Simon P. Preston. Modelling the regulation of telomere length: the effects of telomerase and G-Quadruplex stabilising drugs. Journal of Mathematical Biology, 68(6): 1521–1552, May 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0678-2.pdf.

Abdelrazec:2014:TDW

[2267] Ahmed Abdelrazec, Suzanne Lenhart, and Huaiping Zhu. Transmission dynamics of West Nile virus in mosquitoes and corvids and non-corvids. Journal of Mathematical Biology, 68(6):1553–1582, May 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0677-3.pdf.

Graham:2014:DSE

[2268] Matthew Graham and Thomas House. Dynamics of stochastic epidemics on heterogeneous networks. Journal of Mathematical Biology, 68(7): 1583–1605, June 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0679-1.pdf. See erratum [2533].

Stinner:2014:CET

[2269] C. Stinner, J. I. Tello, and M. Winkler. Competitive exclusion in a two-species chemotaxis model. Journal of Mathematical Biology, 68(7): 1607–1626, June 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0681-7.pdf.

Fernandez:2014:TTC

[2270] Bastien Fernandez and Lev S. Tsimring. Typical trajectories of coupled degrade-and-fire oscillators: from dispersed populations to massive clustering. Journal of Mathematical Biology, 68(7):1627–1652, June 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0680-8.pdf.

Oelz:2014:VTP

[2271] Dietmar Oelz. A viscous two-phase model for contractile actomyosin bundles. Journal of Mathematical Biology, 68(7):1653–1676, June 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-013-0682-6.pdf.

Vorotnikov:2014:AAB

[2272] Dmitry Vorotnikov. Analytical aspects of the Brownian motor effect in randomly flashing ratchets. Journal of Mathematical Biology, 68(7): 1677–1705, June 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0684-4.pdf.

vanIersel:2014:TET

[2273] Leo van Iersel and Vincent Moulton. Trinets encode tree-child and level-2 phylogenetic networks. Journal of Mathematical Biology, 68(7):1707– 1729, June 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0683-5.pdf.

Wang:2014:ATC

[2274] Xueying Wang, Jay R. Walton, Rana D. Parshad, Katie Storey, and May Boggess. Analysis of the Trojan Y-Chromosome eradication strategy for an invasive species. Journal of Mathematical Biology, 68(7):1731– 1756, June 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0687-1.pdf.

Kugler:2014:IAJ

[2275] Philipp Kügler and Wei Yang. Identification of alterations in the Jacobian of biochemical reaction networks from steady state covariance data at two conditions. Journal of Mathematical Biology, 68(7):1757– 1783, June 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0685-3.pdf.

Hidalgo:2014:NAS

[2276] A. Hidalgo, L. Tello, and E. F. Toro. Numerical and analytical study of an atherosclerosis inflammatory disease model. Journal of Mathematical Biology, 68(7):1785–1814, June 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0688-0.pdf.

Haegeman:2014:PCP

[2277] Bart Haegeman, Tewfik Sari, and Rampal S. Etienne. Predicting coexistence of plants subject to a tolerance-competition trade-off. Journal of Mathematical Biology, 68(7):1815–1847, June 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0692-4.pdf.

Eager:2014:GAS

[2278] Eric Alan Eager, Richard Rebarber, and Brigitte Tenhumberg. Global asymptotic stability of plant-seed bank models. Journal of Mathematical Biology, 69(1):1–37, July 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0689-z.pdf.

Kong:2014:DPS

[2279] Yong Kong. Distributions of positive signals in pyrosequencing. Journal of Mathematical Biology, 69(1):39–54, July 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-013-0691-5.pdf.

Pauleve:2014:DPD

[2280] Loïc Paulevé, Gheorghe Craciun, and Heinz Koeppl. Dynamical properties of discrete reaction networks. Journal of Mathematical Biology, 69 (1):55–72, July 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0686-2.pdf.

Bacaer:2014:LBD

[2281] Nicolas Bacaër and Abdelkarim Ed-Darraz. On linear birth-and-death processes in a random environment. Journal of Mathematical Biology, 69 (1):73–90, July 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0696-0.pdf.

Marth:2014:SNC

[2282] Wieland Marth and Axel Voigt. Signaling networks and cell motility: a computational approach using a phase field description. Journal of Mathematical Biology, 69(1):91–112, July 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0704-4.pdf.

Xiao:2014:TDV

[2283] Yanyu Xiao and Xingfu Zou. Transmission dynamics for vector-borne diseases in a patchy environment. Journal of Mathematical Biology, 69(1): 113–146, July 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0695-1.pdf.

Lladser:2014:AST

[2284] Manuel E. Lladser and Stephen R. Chestnut. Approximation of sojourntimes via maximal couplings: motif frequency distributions. Journal of Mathematical Biology, 69(1):147–182, July 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-013-0690-6.pdf.

Taylor:2014:IHE

[2285] Timothy J. Taylor and Istvan Z. Kiss. Interdependency and hierarchy of exact and approximate epidemic models on networks. Journal of Mathematical Biology, 69(1):183–211, July 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0699-x.pdf.

Hwang:2014:IRC

[2286] Yongyun Hwang, Praveen Kumar, and Abdul I. Barakat. Intracellular regulation of cell signaling cascades: how location makes a difference. Journal of Mathematical Biology, 69(1):213–242, July 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0701-7.pdf.

Egri-Nagy:2014:GTM

[2287] Attila Egri-Nagy, Volker Gebhardt, Mark M. Tanaka, and Andrew R. Francis. Group-theoretic models of the inversion process in bacterial genomes. Journal of Mathematical Biology, 69(1):243–265, July 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-013-0702-6.pdf.

Strickland:2014:MPP

[2288] Christopher Strickland, Gerhard Dangelmayr, and Patrick D. Shipman. Modeling the presence probability of invasive plant species with nonlocal dispersal. Journal of Mathematical Biology, 69(2):267–294, August 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0693-3.pdf.

Batty:2014:FMT

[2289] Charles J. K. Batty, Paul Crewe, Alan Grafen, and Richard Gratwick. Foundations of a mathematical theory of Darwinism. Journal of Mathematical Biology, 69(2):295–334, August 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0706-2.pdf.

Poignard:2014:ICG

[2290] Camille Poignard. Inducing chaos in a gene regulatory network by coupling an oscillating dynamics with a hysteresis-type one. Journal of Mathematical Biology, 69(2):335–368, August 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0703-5.pdf.

Garlick:2014:HSD

[2291] Martha J. Garlick, James A. Powell, Mevin B. Hooten, and Leslie R. MacFarlane. Homogenization, sex, and differential motility predict spread of chronic wasting disease in mule deer in southern Utah. Journal of Mathematical Biology, 69(2):369–399, August 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0709-z.pdf.

Sarhad:2014:PPR

[2292] Jonathan Sarhad, Robert Carlson, and Kurt E. Anderson. Population persistence in river networks. Journal of Mathematical Biology, 69(2): 401–448, August 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0710-6.pdf.

Dumett:2014:PIC

[2293] Miguel A. Dumett and James P. Keener. The pyrite iron cycle catalyzed by acidithiobacillus ferrooxidans. Journal of Mathematical Biology, 69(2):449–467, August 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0708-0.pdf.

Ghandi:2014:RMF

[2294] Mahmoud Ghandi, Morteza Mohammad-Noori, and Michael A. Beer. Robust k-mer frequency estimation using gapped k-mers. Journal of Mathematical Biology, 69(2):469–500, August 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-013-0705-3.pdf.

Wang:2014:SMT

[2295] Xueying Wang, Raju Gautam, Pablo J. Pinedo, Linda J. S. Allen, and Renata Ivanek. A stochastic model for transmission, extinction and outbreak of escherichia coli O157:H7 in cattle as affected by ambient temperature and cleaning practices. Journal of Mathematical Biology, 69(2):

501–532, August 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0707-1.pdf.

Ducrot:2014:CPT

[2296] Arnaud Ducrot and Thomas Giletti. Convergence to a pulsating travelling wave for an epidemic reaction–diffusion system with non-diffusive susceptible population. Journal of Mathematical Biology, 69(3):533–552, September 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0713-3.pdf.

Garcia-Penarrubia:2014:MMC

[2297] Pilar García-Peñarrubia, Juan J. Gálvez, and Jesús Gálvez. Mathematical modelling and computational study of two-dimensional and threedimensional dynamics of receptor-ligand interactions in signalling response mechanisms. Journal of Mathematical Biology, 69(3):553–582, September 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0712-4.pdf.

Musgrave:2014:IEPa

[2298] Jeffrey Musgrave and Frithjof Lutscher. Integrodifference equations in patchy landscapes. Journal of Mathematical Biology, 69(3):583–615, September 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0714-2.pdf.

Musgrave:2014:IEPb

[2299] Jeffrey Musgrave and Frithjof Lutscher. Integrodifference equations in patchy landscapes. Journal of Mathematical Biology, 69(3):617–658, September 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0715-1.pdf.

An:2014:QMP

[2300] Le Thi Thanh An and Willi Jäger. A quantitative model of population dynamics in malaria with drug treatment. Journal of Mathematical Biology, 69(3):659–685, September 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0716-0.pdf.

Hasenauer:2014:MCM

[2301] J. Hasenauer, V. Wolf, A. Kazeroonian, and F. J. Theis. Method of conditional moments (MCM) for the chemical master equation. Journal of Mathematical Biology, 69(3):687–735, September 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0711-5.pdf.

Gamado:2014:MUR

[2302] Kokouvi M. Gamado, George Streftaris, and Stan Zachary. Modelling under-reporting in epidemics. Journal of Mathematical Biology, 69(3): 737–765, September 2014. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-013-0717-z.pdf.

Ganguly:2014:MCA

[2303] Arnab Ganguly, Tatjana Petrov, and Heinz Koeppl. Markov chain aggregation and its applications to combinatorial reaction networks. Journal of Mathematical Biology, 69(3):767–797, September 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0738-7.pdf.

Bocker:2014:CGR

[2304] Sebastian Böcker and Stephan Wagner. Counting glycans revisited. Journal of Mathematical Biology, 69(4):799–816, October 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0721-3.pdf.

Tao:2014:PDE

[2305] Youshan Tao, Qian Guo, and Kazuyuki Aihara. A partial differential equation model and its reduction to an ordinary differential equation model for prostate tumor growth under intermittent hormone therapy. Journal of Mathematical Biology, 69(4):817–838, October 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0718-y.pdf.

Al-Husari:2014:CAM

[2306] Maymona Al-Husari, Craig Murdoch, and Steven D. Webb. A cellular automaton model examining the effects of oxygen, hydrogen ions and lactate on early tumour growth. Journal of Mathematical Biology, 69(4):839–873, October 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0719-x.pdf.

Yuan:2014:TDS

[2307] Yuan Yuan and Jacques Bélair. Threshold dynamics in an SEIRS model with latency and temporary immunity. Journal of Mathematical Biology, 69(4):875–904, October 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0720-4.pdf.

Boldin:2014:ECE

[2308] Barbara Boldin and Odo Diekmann. An extension of the classification of evolutionarily singular strategies in adaptive dynamics. Journal of Mathematical Biology, 69(4):905–940, October 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0725-z.pdf.

Newby:2014:MBM

[2309] Jay Newby and Jon Chapman. Metastable behavior in Markov processes with internal states. Journal of Mathematical Biology, 69(4):941– 976, October 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0723-1.pdf.

Ryerson:2014:UIM

[2310] Shane Ryerson and Germán A. Enciso. Ultrasensitivity in independent multisite systems. Journal of Mathematical Biology, 69(4):977– 999, October 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0727-x.pdf.

Bolzoni:2014:RWW

[2311] Luca Bolzoni, Valentina Tessoni, Maria Groppi, and Giulio A. De Leo. React or wait: which optimal culling strategy to control infectious diseases in wildlife. Journal of Mathematical Biology, 69(4):1001–1025, October 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0726-y.pdf.

Barbarossa:2014:SDN

[2312] M. V. Barbarossa, K. P. Hadeler, and C. Kuttler. State-dependent neutral delay equations from population dynamics. Journal of Mathematical Biology, 69(4):1027–1056, October 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0821-8.pdf.

Hossjer:2014:QEV

[2313] Ola Hössjer and Nils Ryman. Quasi equilibrium, variance effective size and fixation index for populations with substructure. Journal of Mathematical Biology, 69(5):1057–1128, November 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-013-0728-9.pdf.

Bundschuh:2014:UAP

[2314] Ralf Bundschuh. Unified approach to partition functions of RNA secondary structures. Journal of Mathematical Biology, 69(5):1129–1150, November 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0737-8.pdf.

Muller:2014:FMM

[2315] Arne C. Müller and Alexander Bockmayr. Flux modules in metabolic networks. Journal of Mathematical Biology, 69(5):1151–1179, November 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-013-0731-1.pdf.

Korobenko:2014:ESC

[2316] L. Korobenko and E. Braverman. On evolutionary stability of carrying capacity driven dispersal in competition with regularly diffusing populations. Journal of Mathematical Biology, 69(5):1181–1206, November 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0729-8.pdf.

Helal:2014:ADA

[2317] Mohamed Helal, Erwan Hingant, Laurent Pujo-Menjouet, and Glenn F. Webb. Alzheimer’s disease: analysis of a mathematical model incorporating the role of prions. Journal of Mathematical Biology, 69(5):1207– 1235, November 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0732-0.pdf.

Calcagno:2014:HOF

[2318] Vincent Calcagno, Ludovic Mailleret, Éric Wajnberg, and Frédéric Grognard. How optimal foragers should respond to habitat changes: a reanalysis of the marginal value theorem. Journal of Mathematical Biology, 69(5):1237–1265, November 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0734-y.pdf.

Roth:2014:PFE

[2319] Gregory Roth and Sebastian J. Schreiber. Persistence in fluctuating environments for interacting structured populations. Journal of Mathematical Biology, 69(5):1267–1317, November 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0739-6.pdf.

Lou:2014:EDO

[2320] Yuan Lou and Frithjof Lutscher. Evolution of dispersal in open advective environments. Journal of Mathematical Biology, 69(6–7):1319– 1342, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0730-2.pdf.

Cosner:2014:WPQ

[2321] Chris Cosner and Michael Winkler. Well-posedness and qualitative properties of a dynamical model for the ideal free distribution. Journal of Mathematical Biology, 69(6–7):1343–1382, December 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0733-z.pdf.

Klika:2014:CMB

[2322] Václav Klika, Maria Angelés Pérez, José Manuel García-Aznar, Frantisek Marsík, and Manuel Doblaré. A coupled mechano-biochemical model for bone adaptation. Journal of Mathematical Biology, 69(6–7):1383– 1429, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0736-9.pdf.

Blyuss:2014:ASM

[2323] Konstantin B. Blyuss. Analysis of symmetries in models of multistrain infections. Journal of Mathematical Biology, 69(6–7):1431–1459, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0740-0.pdf.

Abou-Jaoude:2014:LBT

[2324] Wassim Abou-Jaoudé, Madalena Chaves, and Jean-Luc Gouzé. Links between topology of the transition graph and limit cycles in a twodimensional piecewise affine biological model. Journal of Mathematical Biology, 69(6–7):1461–1495, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-013-0735-x.pdf.

Krivan:2014:ATI

[2325] Vlastimil Krivan. The Allee-type ideal free distribution. Journal of Mathematical Biology, 69(6–7):1497–1513, December 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0742-y.pdf.

Prokopiou:2014:MAM

[2326] Sotiris A. Prokopiou, Helen M. Byrne, Mike R. Jeffrey, Robert S. Robinson, George E. Mann, and Markus R. Owen. Mathematical analysis of a model for the growth of the bovine corpus luteum. Journal of Mathematical Biology, 69(6–7):1515–1546, December 2014. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-013-0722-2.pdf.

Luzyanina:2014:MMC

[2327] Tatyana Luzyanina, Jovana Cupovic, Burkhard Ludewig, and Gennady Bocharov. Mathematical models for CFSE labelled lymphocyte dynamics: asymmetry and time-lag in division. Journal of Mathematical Biology, 69(6–7):1547–1583, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0741-z.pdf.

Cantini:2014:SAS

[2328] Laura Cantini, Claudia Cianci, Duccio Fanelli, Emma Massi, Luigi Barletti, and Malbor Asllani. Stochastic amplification of spatial modes in a system with one diffusing species. Journal of Mathematical Biology, 69(6–7):1585–1608, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0743-x.pdf.

Massarelli:2014:EPP

[2329] Nicole Massarelli, Kathleen Hoffman, and Joseph P. Previte. Effect of parity on productivity and sustainability of Lotka–Volterra food chains. Journal of Mathematical Biology, 69(6–7):1609–1626, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0746-7.pdf.

Noel:2014:SDC

[2330] Pierre-André Noël, Antoine Allard, Laurent Hébert-Dufresne, Vincent Marceau, and Louis J. Dubé. Spreading dynamics on complex networks: a general stochastic approach. Journal of Mathematical Biology, 69(6–7):1627–1660, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0744-9.pdf.

Ironi:2014:APQ

[2331] Liliana Ironi and Ettore Lanzarone. Assigning probabilities to qualitative dynamics of gene regulatory networks. Journal of Mathematical Biology, 69(6–7):1661–1692, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0765-z.pdf.

Francis:2014:AVB

[2332] Andrew R. Francis. An algebraic view of bacterial genome evolution. Journal of Mathematical Biology, 69(6–7):1693–1718, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0747-6.pdf.

Bernhard:2014:SSG

[2333] Pierre Bernhard and Frédéric M. Hamelin. Simple signaling games of sexual selection (Grafen’s revisited). Journal of Mathematical Biology, 69(6–7):1719–1742, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0745-8.pdf.

Poznanovic:2014:ADM

[2334] Svetlana Poznanović and Christine E. Heitsch. Asymptotic distribution of motifs in a stochastic context-free grammar model of RNA folding. Journal of Mathematical Biology, 69(6–7):1743–1772, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0750-y.pdf.

Gujarati:2014:VAD

[2335] Tanvi P. Gujarati and G. Ambika. Virus antibody dynamics in primary and secondary dengue infections. Journal of Mathematical Biology, 69(6–7):1773–1800, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0749-4.pdf.

Barros:2014:EDH

[2336] Manuel Barros and Angel Ferrández. On the energy density of helical proteins. Journal of Mathematical Biology, 69(6–7):1801–1813, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-013-0752-9.pdf.

Adamson:2014:DDS

[2337] M. W. Adamson and A. Yu. Morozov. Defining and detecting structural sensitivity in biological models: developing a new framework. Journal of Mathematical Biology, 69(6–7):1815–1848, December 2014. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0753-3.pdf.

Xue:2015:MEB

[2338] Chuan Xue. Macroscopic equations for bacterial chemotaxis: integration of detailed biochemistry of cell signaling. Journal of Mathematical Biology, 70(1–2):1–44, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0748-5.pdf.

Koslicki:2015:CSD

[2339] David Koslicki and Daniel J. Thompson. Coding sequence density estimation via topological pressure. Journal of Mathematical Biology, 70(1–2): 45–69, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0754-2.pdf.

Eymard:2015:RSO

[2340] N. Eymard, N. Bessonov, O. Gandrillon, M. J. Koury, and V. Volpert. The role of spatial organization of cells in erythropoiesis. Journal of Mathematical Biology, 70(1–2):71–97, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-014-0758-y.pdf.

Mayer:2015:SMC

[2341] Hannah Mayer and Anton Bovier. Stochastic modelling of T-cell activation. Journal of Mathematical Biology, 70(1–2):99–132, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0759-x.pdf.

Kimpton:2015:PMC

[2342] L. S. Kimpton, J. P. Whiteley, S. L. Waters, and J. M. Oliver. On a poroviscoelastic model for cell crawling. Journal of Mathematical Biology, 70(1–2):133–171, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0755-1.pdf.

Senter:2015:RFP

[2343] Evan Senter, Ivan Dotu, and Peter Clote. RNA folding pathways and kinetics using 2D energy landscapes. Journal of Mathematical Biology, 70(1–2):173–196, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0760-4.pdf.

Molina:2015:MMB

[2344] Manuel Molina, Manuel Mota, and Alfonso Ramos. Mathematical modeling in biological populations through branching processes. Application to salmonid populations. Journal of Mathematical Biology, 70(1–2):197– 212, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0762-2.pdf.

Kassa:2015:ISP

[2345] Semu Mitiku Kassa and Aziz Ouhinou. The impact of self-protective measures in the optimal interventions for controlling infectious diseases of human population. Journal of Mathematical Biology, 70(1–2):213– 236, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0761-3.pdf.

George:2015:MVA

[2346] Stuart George, Jamie M. Foster, and Giles Richardson. Modelling in vivo action potential propagation along a giant axon. Journal of Mathematical Biology, 70(1–2):237–263, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-013-0751-x.pdf.

Li:2015:SFS

[2347] Bingtuan Li, William F. Fagan, and Kimberly I. Meyer. Success, failure, and spreading speeds for invasions on spatial gradients. Journal of Mathematical Biology, 70(1–2):265–287, January 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0766-y.pdf.

Coron:2015:OAP

[2348] Jean-Michel Coron, Pierre Gabriel, and Peipei Shang. Optimization of an amplification protocol for misfolded proteins by using relaxed control. Journal of Mathematical Biology, 70(1–2):289–327, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0768-9.pdf.

Schertzer:2015:ISD

[2349] E. Schertzer, A. C. Staver, and S. A. Levin. Implications of the spatial dynamics of fire spread for the bistability of savanna and forest. Journal of Mathematical Biology, 70(1–2):329–341, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0757-z.pdf.

Qesmi:2015:IEM

[2350] Redouane Qesmi, Jane M. Heffernan, and Jianhong Wu. An immunoepidemiological model with threshold delay: a study of the effects of multiple exposures to a pathogen. Journal of Mathematical Biology, 70(1–2): 343–366, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-014-0764-0.pdf.

Lambert:2015:RTL

[2351] Amaury Lambert, Hélène Morlon, and Rampal S. Etienne. The reconstructed tree in the lineage-based model of protracted speciation. Journal of Mathematical Biology, 70(1–2):367–397, January 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0767-x.pdf.

Gandolfi:2015:EDH

[2352] Alberto Gandolfi, Andrea Pugliese, and Carmela Sinisgalli. Epidemic dynamics and host immune response: a nested approach. Journal of Mathematical Biology, 70(3):399–435, February 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0769-8.pdf.

Kiss:2015:EDR

[2353] Istvan Z. Kiss, Charles G. Morris, Fanni Sélley, Péter L. Simon, and Robert R. Wilkinson. Exact deterministic representation of Markovian SIR epidemics on networks with and without loops. Journal of Mathematical Biology, 70(3):437–464, February 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0772-0.pdf.

Nakajima:2015:EDF

[2354] Yohei Nakajima and Naoki Masuda. Evolutionary dynamics in finite populations with zealots. Journal of Mathematical Biology, 70(3):465– 484, February 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0770-2.pdf.

Spill:2015:MCM

[2355] F. Spill, P. Guerrero, T. Alarcon, P. K. Maini, and H. M. Byrne. Mesoscopic and continuum modelling of angiogenesis. Journal of Mathematical Biology, 70(3):485–532, February 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0771-1.pdf.

Beyer:2015:MSC

[2356] Robert Beyer, Octave Etard, Paul-Henry Cournède, and Pascal Laurent-Gengoux. Modeling spatial competition for light in plant populations with the porous medium equation. Journal of Mathematical Biology, 70 (3):533–547, February 2015. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-014-0763-1.pdf.

Jacobsen:2015:IMP

[2357] Jon Jacobsen, Yu Jin, and Mark A. Lewis. Integrodifference models for persistence in temporally varying river environments. Journal of Mathematical Biology, 70(3):549–590, February 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0774-y.pdf.

West:2015:MSC

[2358] Simon West, Lloyd J. Bridge, Michael R. H. White, Pawel Paszek, and Vadim N. Biktashev. A method of ‘speed coefficients’ for biochemical model reduction applied to the NF-κ B system. Journal of Mathematical Biology, 70(3):591–620, February 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0775-x.pdf.

Guy:2015:AEM

[2359] Romain Guy, Catherine Larédo, and Elisabeta Vergu. Approximation of epidemic models by diffusion processes and their statistical inference.

Journal of Mathematical Biology, 70(3):621–646, February 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0777-8.pdf.

Lin:2015:DSEa

[2360] Yen Ting Lin, Hyejin Kim, and Charles R. Doering. Demographic stochasticity and evolution of dispersion I. Spatially homogeneous environments. Journal of Mathematical Biology, 70(3):647–678, February 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0776-9.pdf.

Lin:2015:DSEb

[2361] Yen Ting Lin, Hyejin Kim, and Charles R. Doering. Demographic stochasticity and evolution of dispersion II: Spatially inhomogeneous environments. Journal of Mathematical Biology, 70(3):679–707, February 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0756-0.pdf.

Madzvamuse:2015:CDD

[2362] Anotida Madzvamuse, Hussaini S. Ndakwo, and Raquel Barreira. Cross-diffusion-driven instability for reaction–diffusion systems: analysis and simulations. Journal of Mathematical Biology, 70(4):709– 743, March 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0779-6.pdf.

Caiazzo:2015:CHS

[2363] Alfonso Caiazzo, Gino Montecinos, Lucas O. Müller, E. Mark Haacke, and Eleuterio F. Toro. Computational haemodynamics in stenotic internal jugular veins. Journal of Mathematical Biology, 70(4):745– 772, March 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0778-7.pdf.

Jabbari:2015:MMR

[2364] Sara Jabbari, Stephen T. Cartman, and John R. King. Mathematical modelling reveals properties of TcdC required for it to be a negative regulator of toxin production in clostridium difficile. Journal of Mathematical Biology, 70(4):773–804, March 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0780-0.pdf.

Holcman:2015:PTR

[2365] D. Holcman, K. Dao Duc, A. Jones, H. Byrne, and K. Burrage. Posttranscriptional regulation in the nucleus and cytoplasm: study of mean time to threshold (MTT) and narrow escape problem. Journal of Mathematical Biology, 70(4):805–828, March 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0782-y.pdf.

Fontbona:2015:NLL

[2366] Joaquin Fontbona and Sylvie Méléard. Non local Lotka–Volterra system with cross-diffusion in an heterogeneous medium. Journal of Mathematical Biology, 70(4):829–854, March 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0781-z.pdf.

Fernandez-Sanchez:2015:LMM

[2367] Jesús Fernández-Sánchez, Jeremy G. Sumner, Peter D. Jarvis, and Michael D. Woodhams. Lie Markov models with purine/pyrimidine symmetry. Journal of Mathematical Biology, 70(4):855–891, March 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0773-z.pdf.

Wang:2015:FPC

[2368] Yuandi Wang, Mingya Dou, and Zhigang Zhou. The fencing problem and coleochaete cell division. Journal of Mathematical Biology, 70(4): 893–912, March 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0784-9.pdf.

Sargsyan:2015:AFG

[2369] Ori Sargsyan. An analytical framework in the general coalescent tree setting for analyzing polymorphisms created by two mutations. Journal of Mathematical Biology, 70(4):913–956, March 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0785-8.pdf.

Richard:2015:ESP

[2370] Adrien Richard and Jean-Paul Comet. Erratum to: Stable periodicity and negative circuits in differential systems. Journal of Mathematical Biology, 70(4):957–958, March 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0855-y.pdf.

Grognard:2015:CBP

[2371] Frédéric Grognard, Pierre Masci, Eric Benoît, and Olivier Bernard. Competition between phytoplankton and bacteria: exclusion and coexistence. Journal of Mathematical Biology, 70(5):959–1006, April 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0783-x.pdf.

House:2015:PMF

[2372] Thomas House. For principled model fitting in mathematical biology. Journal of Mathematical Biology, 70(5):1007–1013, April 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0787-6.pdf.

Guiver:2015:ICP

[2373] Chris Guiver, Hartmut Logemann, Richard Rebarber, Adam Bill, Brigitte Tenhumberg, Dave Hodgson, and Stuart Townley. Integral control for population management. Journal of Mathematical Biology, 70(5): 1015–1063, April 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0789-4.pdf.

Tien:2015:DIC

[2374] Joseph H. Tien, Zhisheng Shuai, Marisa C. Eisenberg, and P. van den Driessche. Disease invasion on community networks with environmental pathogen movement. Journal of Mathematical Biology, 70(5):1065– 1092, April 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0791-x.pdf.

Kisdi:2015:CMT

[2375] Éva Kisdi. Construction of multiple trade-offs to obtain arbitrary singularities of adaptive dynamics. Journal of Mathematical Biology, 70(5): 1093–1117, April 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0788-5.pdf.

Donovan:2015:SPF

[2376] Graham M. Donovan and Thibaut Kritter. Spatial pattern formation in the lung. Journal of Mathematical Biology, 70(5):1119– 1149, April 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0792-9.pdf.

Breitsch:2015:CCD

[2377] Nathan Breitsch, Gregory Moses, Erik Boczko, and Todd Young. Cell cycle dynamics: clustering is universal in negative feedback systems. Journal of Mathematical Biology, 70(5):1151–1175, April 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0786-7.pdf.

Rodriguez:2015:IDM

[2378] Nancy Rodríguez. On an integro-differential model for pest control in a heterogeneous environment. Journal of Mathematical Biology, 70(5): 1177–1206, April 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0793-8.pdf.

Sainudiin:2015:FBR

[2379] Raazesh Sainudiin, Tanja Stadler, and Amandine Véber. Finding the best resolution for the Kingman–Tajima coalescent: theory and applications. Journal of Mathematical Biology, 70(6):1207–1247, May 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0796-5.pdf.

Cangelosi:2015:NSA

[2380] Richard A. Cangelosi, David J. Wollkind, Bonni J. Kealy-Dichone, and Inthira Chaiya. Nonlinear stability analyses of Turing patterns for a mussel-algae model. Journal of Mathematical Biology, 70(6):1249– 1294, May 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0794-7.pdf.

Rudnicki:2015:MPE

[2381] Ryszard Rudnicki and Pawel Zwoleński. Model of phenotypic evolution in hermaphroditic populations. Journal of Mathematical Biology, 70(6): 1295–1321, May 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0798-3.pdf.

VanOosterwyck:2015:CMM

[2382] Hans Van Oosterwyck. Computational mechanobiology: may the force be with you. Journal of Mathematical Biology, 70(6):1323– 1326, May 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0795-6.pdf.

Martinez:2015:CAD

[2383] Luis Martínez, Martin Milanic, Leire Legarreta, Paul Medvedev, Iker Malaina, and Ildefonso M. de la Fuente. A combinatorial approach to the design of vaccines. Journal of Mathematical Biology, 70(6):1327– 1358, May 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0797-4.pdf.

Yoon:2015:BCG

[2384] Changwook Yoon and Yong-Jung Kim. Bacterial chemotaxis without gradient-sensing. Journal of Mathematical Biology, 70(6):1359– 1380, May 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0790-y.pdf.

Maxin:2015:MBH

[2385] Daniel Maxin, Ludek Berec, Adrienna Bingham, Denali Molitor, and Julie Pattyson. Is more better? higher sterilization of infected hosts need not result in reduced pest population size. Journal of Mathematical Biology, 70(6):1381–1409, May 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0800-0.pdf.

Nakata:2015:GAS

[2386] Yukihiko Nakata and Gergely Röst. Global analysis for spread of infectious diseases via transportation networks. Journal of Mathematical Biology, 70(6):1411–1456, May 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0801-z.pdf.

Aviziotis:2015:CGA

[2387] I. G. Aviziotis, M. E. Kavousanakis, I. A. Bitsanis, and A. G. Boudouvis. Coarse-grained analysis of stochastically simulated cell populations with a positive feedback genetic network architecture. Journal of Mathematical Biology, 70(7):1457–1484, June 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0799-2.pdf.

Hyrien:2015:SMS

[2388] O. Hyrien, S. A. Peslak, N. M. Yanev, and J. Palis. Stochastic modeling of stress erythropoiesis using a two-type age-dependent branching process with immigration. Journal of Mathematical Biology, 70(7):1485– 1521, June 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0803-x.pdf.

Uecker:2015:AGI

[2389] Hildegard Uecker, Derek Setter, and Joachim Hermisson. Adaptive gene introgression after secondary contact. Journal of Mathematical Biology, 70(7):1523–1580, June 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0802-y.pdf.

Ngonghala:2015:POB

[2390] Calistus N. Ngonghala, Miranda I. Teboh-Ewungkem, and Gideon A. Ngwa. Persistent oscillations and backward bifurcation in a malaria model with varying human and mosquito populations: implications for control. Journal of Mathematical Biology, 70(7):1581–1622, June 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0804-9.pdf.

Fimmel:2015:CCS

[2391] Elena Fimmel, Simone Giannerini, Diego Luis Gonzalez, and Lutz Strüngmann. Circular codes, symmetries and transformations. Journal of Mathematical Biology, 70(7):1623–1644, June 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0806-7.pdf.

Severino:2015:BLS

[2392] Gerardo Severino and Daniel M. Tartakovsky. A boundary-layer solution for flow at the soil-root interface. Journal of Mathematical Biology, 70(7): 1645–1668, June 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0813-8.pdf.

Millan:2015:IDR

[2393] Mercedes Pérez Millán and Alicia Dickenstein. Implicit dose-response curves. Journal of Mathematical Biology, 70(7):1669–1684, June 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0809-4.pdf.

Hoze:2015:KAF

[2394] Nathanael Hoze and David Holcman. Kinetics of aggregation with a finite number of particles and application to viral capsid assembly. Journal of Mathematical Biology, 70(7):1685–1705, June 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0819-2.pdf.

Leung:2015:SID

[2395] Ka Yin Leung, Mirjam Kretzschmar, and Odo Diekmann. SI infection on a dynamic partnership network: characterization of R_{0}. Journal of Mathematical Biology, 71(1):1–56, July 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0808-5.pdf.

Torres:2015:POM

[2396] Pedro J. Torres. Periodic oscillations of a model for membrane permeability with fluctuating environmental conditions. Journal of Mathematical Biology, 71(1):57–68, July 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0815-6.pdf.

Rapaport:2015:GDB

[2397] Alain Rapaport, Ihab Haidar, and Jérôme Harmand. Global dynamics of the buffered chemostat for a general class of response functions. Journal of Mathematical Biology, 71(1):69–98, July 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0814-7.pdf.

Zhu:2015:CCC

[2398] Sha Zhu, Cuong Than, and Taoyang Wu. Clades and clans: a comparison study of two evolutionary models. Journal of Mathematical Biology, 71 (1):99–124, July 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0817-4.pdf.

Kuwamura:2015:TIP

[2399] Masataka Kuwamura. Turing instabilities in prey-predator systems with dormancy of predators. Journal of Mathematical Biology, 71(1):125– 149, July 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0816-5.pdf.

Jarrett:2015:GSA

[2400] Angela M. Jarrett, Yaning Liu, N. G. Cogan, and M. Yousuff Hussaini. Global sensitivity analysis used to interpret biological experimental results. Journal of Mathematical Biology, 71(1):151–170, July 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0818-3.pdf.

DiCostanzo:2015:HMM

[2401] E. Di Costanzo, R. Natalini, and L. Preziosi. A hybrid mathematical model for self-organizing cell migration in the zebrafish lateral line. Journal of Mathematical Biology, 71(1):171–214, July 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0812-9.pdf.

Browne:2015:RPV

[2402] Cameron J. Browne, Robert J. Smith, and Lydia Bourouiba. From regional pulse vaccination to global disease eradication: insights from a mathematical model of poliomyelitis. Journal of Mathematical Biology, 71(1):215–253, July 2015. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-014-0810-y.pdf.

Shuai:2015:EET

[2403] Zhisheng Shuai, J. A. P. Heesterbeek, and P. van den Driessche. Erratum to: Extending the type reproduction number to infectious disease control targeting contacts between types. Journal of Mathematical Biology, 71(1):255–257, July 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0858-3.pdf.

Faugeras:2015:SNF

[2404] O. Faugeras and J. Inglis. Stochastic neural field equations: a rigorous footing. Journal of Mathematical Biology, 71(2):259–300, August 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0807-6.pdf.

Pribylova:2015:PIS

[2405] Lenka Pribylová and Ludek Berec. Predator interference and stability of predator–prey dynamics. Journal of Mathematical Biology, 71(2):301– 323, August 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0820-9.pdf.

Evans:2015:PPE

[2406] Steven N. Evans, Alexandru Hening, and Sebastian J. Schreiber. Protected polymorphisms and evolutionary stability of patch-selection strategies in stochastic environments. Journal of Mathematical Biology, 71(2):325–359, August 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0824-5.pdf.

Bruni:2015:OWS

[2407] C. Bruni, F. Conte, F. Papa, and C. Sinisgalli. Optimal weekly scheduling in fractionated radiotherapy: effect of an upper bound on the dose fraction size. Journal of Mathematical Biology, 71(2):361– 398, August 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0805-8.pdf.

Croft:2015:PIP

[2408] Wayne Croft, Charles M. Elliott, Graham Ladds, Björn Stinner, Chandrasekhar Venkataraman, and Cathryn Weston. Parameter identification problems in the modelling of cell motility. Journal of Mathematical Biology, 71(2):399–436, August 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0823-6.pdf.

Antunes:2015:QGE

[2409] Duarte Antunes and Abhyudai Singh. Quantifying gene expression variability arising from randomness in cell division times. Journal of Mathematical Biology, 71(2):437–463, August 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0811-x.pdf.

Roques:2015:ECN

[2410] L. Roques, Y. Hosono, O. Bonnefon, and T. Boivin. The effect of competition on the neutral intraspecific diversity of invasive species. Journal of Mathematical Biology, 71(2):465–489, August 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0825-4.pdf.

Bacaer:2015:SSE

[2411] Nicolas Bacaër. On the stochastic SIS epidemic model in a periodic environment. Journal of Mathematical Biology, 71(2):491–511, August 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0828-1.pdf.

Ciupe:2015:MMM

[2412] Stanca M. Ciupe. Mathematical model of multivalent virus-antibody complex formation in humans following acute and chronic HIV infections. Journal of Mathematical Biology, 71(3):513–532, September 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0826-3.pdf.

Zumbrum:2015:CMO

[2413] Matthew E. Zumbrum and David A. Edwards. Conformal mapping in optical biosensor applications. Journal of Mathematical Biology, 71(3):533– 550, September 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0827-2.pdf.

Engwer:2015:GFW

[2414] Christian Engwer, Thomas Hillen, Markus Knappitsch, and Christina Surulescu. Glioma follow white matter tracts: a multiscale DTI-based model. Journal of Mathematical Biology, 71(3):551–582, September 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0822-7.pdf.

Rao:2015:GCE

[2415] Arni S. R. Srinivasa Rao and James R. Carey. Generalization of Carey’s equality and a theorem on stationary population. Journal of Mathematical Biology, 71(3):583–594, September 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0831-6.pdf.

Hossjer:2015:EES

[2416] Ola Hössjer. On the eigenvalue effective size of structured populations. Journal of Mathematical Biology, 71(3):595–646, September 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0832-5.pdf.

Winstanley:2015:CCB

[2417] H. F. Winstanley, M. Chapwanya, A. C. Fowler, and S. B. G. O’Brien. A 2D channel-clogging biofilm model. Journal of Mathematical Biology, 71(3):647–668, September 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0833-4.pdf.

Seki:2015:GMC

[2418] Motohide Seki, Tomohiko Yoshida, and Takenori Takada. A general method for calculating the optimal leaf longevity from the viewpoint of carbon economy. Journal of Mathematical Biology, 71(3):669–690, September 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0830-7.pdf.

Dyson:2015:IVE

[2419] Louise Dyson and Ruth E. Baker. The importance of volume exclusion in modelling cellular migration. Journal of Mathematical Biology, 71(3):691–711, September 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0829-0.pdf.

Barbour:2015:IPB

[2420] A. D. Barbour and M. J. Luczak. Individual and patch behaviour in structured metapopulation models. Journal of Mathematical Biology, 71(3):713–733, September 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0834-3.pdf.

Amaris:2015:FTR

[2421] Armando J. R. Amaris and Murray P. Cox. A flexible theoretical representation for the temporal dynamics of structured populations as paths on polytope complexes. Journal of Mathematical Biology, 71(3):735– 766, September 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0841-4.pdf.

Bilge:2015:UEM

[2422] Ayse Humeyra Bilge, Funda Samanlioglu, and Onder Ergonul. On the uniqueness of epidemic models fitting a normalized curve of removed individuals. Journal of Mathematical Biology, 71(4):767–794, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0838-z.pdf.

Claus:2015:GHB

[2423] Juliane Claus, Mariya Ptashnyk, Ansgar Bohmann, and Andrés Chavarría-Krauser. Global Hopf bifurcation in the ZIP regulatory system. Journal of Mathematical Biology, 71(4):795–816, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0836-1.pdf.

Salau:2015:BAS

[2424] Kehinde R. Salau and Eli P. Fenichel. Bioeconomic analysis supports the endangered species act. Journal of Mathematical Biology, 71(4):817– 846, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0840-5.pdf.

Buono:2015:SPF

[2425] Pietro-Luciano Buono and Raluca Eftimie. Symmetries and pattern formation in hyperbolic versus parabolic models of self-organised aggregation. Journal of Mathematical Biology, 71(4):847–881, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0842-3.pdf.

Sun:2015:SCP

[2426] Zheng Sun and Natalia L. Komarova. Stochastic control of proliferation and differentiation in stem cell dynamics. Journal of Mathematical Biology, 71(4):883–901, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0835-2.pdf.

Oddsdottir:2015:DGR

[2427] Hildur Æsa Oddsdóttir, Erika Hagrot, Véronique Chotteau, and Anders Forsgren. On dynamically generating relevant elementary flux modes in a metabolic network using optimization. Journal of Mathematical Biology, 71(4):903–920, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0844-1.pdf.

Sturrock:2015:MFA

[2428] M. Sturrock, P. J. Murray, A. Matzavinos, and M. A. J. Chaplain. Mean field analysis of a spatial stochastic model of a gene regulatory network. Journal of Mathematical Biology, 71(4):921–959, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0837-0.pdf.

Li:2015:MDD

[2429] Meili Li, Junling Ma, and P. van den Driessche. Model for disease dynamics of a waterborne pathogen on a random network. Journal of Mathematical Biology, 71(4):961–977, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-014-0839-y.pdf.

Broom:2015:EFC

[2430] Mark Broom, Michal Johanis, and Jan Rychtár. The effect of fight cost structure on fighting behaviour. Journal of Mathematical Biology, 71(4):979–996, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0848-x.pdf.

Harris:2015:RQD

[2431] Terence Harris and Anna Q. Cai. Red Queen dynamics in specific predator–prey systems. Journal of Mathematical Biology, 71(4):997– 1016, October 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0850-3.pdf.

Fan:2015:DDS

[2432] Guihong Fan, Horst R. Thieme, and Huaiping Zhu. Delay differential systems for tick population dynamics. Journal of Mathematical Biology, 71(5):1017–1048, November 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0845-0.pdf.

Colombi:2015:DCB

[2433] Annachiara Colombi, Marco Scianna, and Andrea Tosin. Differentiated cell behavior: a multiscale approach using measure theory. Journal of Mathematical Biology, 71(5):1049–1079, November 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0846-z.pdf.

Zienkiewicz:2015:DDS

[2434] Adam Zienkiewicz, David A. W. Barton, Maurizio Porfiri, and Mario di Bernardo. Data-driven stochastic modelling of zebrafish locomotion. Journal of Mathematical Biology, 71(5):1081–1105, November 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-014-0843-2.pdf.

Capone:2015:IDS

[2435] Florinda Capone, Valentina De Cataldis, and Roberta De Luca. Influence of diffusion on the stability of equilibria in a reaction–diffusion system modeling cholera dynamic. Journal of Mathematical Biology, 71(5):1107– 1131, November 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0849-9.pdf.

Fuchs:2015:ESV

[2436] Michael Fuchs and Emma Yu Jin. Equality of Shapley value and fair proportion index in phylogenetic trees. Journal of Mathematical Biology, 71(5):1133–1147, November 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-014-0853-0.pdf.

Bhatia:2015:ADC

[2437] Sangeeta Bhatia, Attila Egri-Nagy, and Andrew R. Francis. Algebraic double cut and join. Journal of Mathematical Biology, 71(5):1149–1178, November 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0852-1.pdf.

Chan:2015:ESG

[2438] Yao ban Chan, Vincent Ranwez, and Céline Scornavacca. Exploring the space of gene/species reconciliations with transfers. Journal of Mathematical Biology, 71(5):1179–1209, November 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-014-0851-2.pdf.

Billiard:2015:SDA

[2439] Sylvain Billiard, Régis Ferrière, Sylvie Méléard, and Viet Chi Tran. Stochastic dynamics of adaptive trait and neutral marker driven by ecoevolutionary feedbacks. Journal of Mathematical Biology, 71(5):1211– 1242, November 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-014-0847-y.pdf.

Arino:2015:EIM

[2440] Julien Arino and Stéphanie Portet. Epidemiological implications of mobility between a large urban centre and smaller satellite cities. Journal of Mathematical Biology, 71(5):1243–1265, November 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-014-0854-z.pdf.

Capone:2015:EID

[2441] Florinda Capone, Valentina De Cataldis, and Roberta De Luca. Erratum to: Influence of diffusion on the stability of equilibria in a reaction– diffusion system modeling cholera dynamic. Journal of Mathematical Biology, 71(5):1267–1268, November 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0915-y.pdf.

Shu:2015:DIS

[2442] Hongying Shu, Xi Hu, Lin Wang, and James Watmough. Delay induced stability switch, multitype bistability and chaos in an intraguild predation model. Journal of Mathematical Biology, 71(6–7):1269–1298, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0857-4.pdf.

Monk:2015:ECO

[2443] Travis Monk, Michael G. Paulin, and Peter Green. Ecological constraints on the origin of neurones. Journal of Mathematical Biology, 71(6–7):1299–1324, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0862-7.pdf.

Terry:2015:PPM

[2444] Alan J. Terry. predator–prey models with component Allee effect for predator reproduction. Journal of Mathematical Biology, 71(6–7):1325– 1352, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0856-5.pdf.

Eren:2015:SMP

[2445] Ezgi C. Eren, Ram Dixit, and Natarajan Gautam. Stochastic models for plant microtubule self-organization and structure. Journal of Mathematical Biology, 71(6–7):1353–1385, December 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0860-9.pdf.

Privault:2015:CFM

[2446] Nicolas Privault and Stéphane Guindon. Closed form modeling of evolutionary rates by exponential Brownian functionals. Journal of Mathematical Biology, 71(6–7):1387–1409, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-015-0863-6.pdf.

Goudenege:2015:WFM

[2447] Ludovic Goudenège and Pierre-André Zitt. A Wright–Fisher model with indirect selection. Journal of Mathematical Biology, 71(6–7):1411–1450, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0859-2.pdf.

Chalmers:2015:BDM

[2448] Alexander D. Chalmers, Anna Cohen, Christina A. Bursill, and Mary R. Myerscough. Bifurcation and dynamics in a mathematical model of early atherosclerosis. Journal of Mathematical Biology, 71(6–7):1451– 1480, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0864-5.pdf.

Barbour:2015:CDS

[2449] A. D. Barbour, R. McVinish, and P. K. Pollett. Connecting deterministic and stochastic metapopulation models. Journal of Mathematical Biology, 71(6–7):1481–1504, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0865-4.pdf.

Kebir:2015:UHS

[2450] Amira Kebir, Nina H. Fefferman, and Slimane Ben Miled. Understanding hermaphrodite species through game theory. Journal of Mathematical Biology, 71(6–7):1505–1524, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0866-3.pdf.

Qureshi:2015:CSP

[2451] M. Umar Qureshi and N. A. Hill. A computational study of pressure wave reflections in the pulmonary arteries. Journal of Mathematical Biology, 71(6–7):1525–1549, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0867-2.pdf.

Broom:2015:SDM

[2452] Mark Broom, Charlotte Lafaye, Karan Pattni, and Jan Rychtár. A study of the dynamics of multi-player games on small networks using territorial interactions. Journal of Mathematical Biology, 71(6–7):1551– 1574, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0868-1.pdf.

Kooi:2015:MDT

[2453] B. W. Kooi. Modelling the dynamics of traits involved in fightingpredators-prey system. Journal of Mathematical Biology, 71(6–7):1575– 1605, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0869-0.pdf.

Coudiere:2015:TLM

[2454] Yves Coudière, Jacques Henry, and Simon Labarthe. A two layers monodomain model of cardiac electrophysiology of the atria. Journal of Mathematical Biology, 71(6–7):1607–1641, December 2015. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0861-8.pdf.

Margheri:2015:CBV

[2455] Alessandro Margheri, Carlota Rebelo, and M. Gabriela M. Gomes. On the correlation between variance in individual susceptibilities and infection prevalence in populations. Journal of Mathematical Biology, 71(6–7):1643–1661, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0870-7.pdf.

Gaubert:2015:DLM

[2456] Stéphane Gaubert and Thomas Lepoutre. Discrete limit and monotonicity properties of the Floquet eigenvalue in an age structured cell division cycle model. Journal of Mathematical Biology, 71(6–7):1663– 1703, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0874-3.pdf.

Ball:2015:EWH

[2457] Frank Ball and Laurence Shaw. Estimating the within-household infection rate in emerging SIR epidemics among a community of households. Journal of Mathematical Biology, 71(6–7):1705–1735, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0872-5.pdf.

Barbarossa:2015:IEP

[2458] M. V. Barbarossa and G. Röst. Immuno-epidemiology of a population structured by immune status: a mathematical study of waning immunity and immune system boosting. Journal of Mathematical Biology, 71(6–7):1737–1770, December 2015. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0880-5.pdf.

Kucharski:2016:CDP

[2459] Adam J. Kucharski, Viggo Andreasen, and Julia R. Gog. Capturing the dynamics of pathogens with many strains. Journal of Mathematical Biology, 72(1–2):1–24, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0873-4.pdf.

Potts:2016:TPF

[2460] Jonathan R. Potts and Mark A. Lewis. Territorial pattern formation in the absence of an attractive potential. Journal of Mathematical Biology, 72(1–2):25–46, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0881-4.pdf.

Greenman:2016:MES

[2461] C. D. Greenman, S. L. Cooke, J. Marshall, M. R. Stratton, and P. J. Campbell. Modeling the evolution space of breakage fusion bridge cycles with a stochastic folding process. Journal of Mathematical Biology, 72(1–2):47–86, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0875-2.pdf.

Popovic:2016:GAF

[2462] Nikola Popović, Carsten Marr, and Peter S. Swain. A geometric analysis of fast-slow models for stochastic gene expression. Journal of Mathematical Biology, 72(1–2):87–122, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0876-1.pdf.

Guerrero:2016:ILI

[2463] Pilar Guerrero, Helen M. Byrne, Philip K. Maini, and Tomás Alarcón. From invasion to latency: intracellular noise and cell motility as key controls of the competition between resource-limited cellular populations. Journal of Mathematical Biology, 72(1–2):123–156, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-015-0883-2.pdf.

Albert:2016:PDM

[2464] Jaroslav Albert and Marianne Rooman. Probability distributions for multimeric systems. Journal of Mathematical Biology, 72(1–2):157– 169, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0877-0.pdf.

Coron:2016:SFS

[2465] Camille Coron. Slow-fast stochastic diffusion dynamics and quasistationarity for diploid populations with varying size. Journal of Mathematical Biology, 72(1–2):171–202, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-015-0878-z.pdf.

McAvoy:2016:SCS

[2466] Alex McAvoy and Christoph Hauert. Structure coefficients and strategy selection in multiplayer games. Journal of Mathematical Biology, 72(1–2):203–238, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0882-3.pdf.

DeAngelis:2016:DSH

[2467] Donald L. DeAngelis, Wei-Ming Ni, and Bo Zhang. Dispersal and spatial heterogeneity: single species. Journal of Mathematical Biology, 72(1–2):239–254, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0879-y.pdf.

Ritchie:2016:BCM

[2468] Martin Ritchie, Luc Berthouze, and Istvan Z. Kiss. Beyond clustering: mean-field dynamics on networks with arbitrary subgraph composition. Journal of Mathematical Biology, 72(1–2):255–281, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0884-1.pdf.

Chalub:2016:FLP

[2469] Fabio A. C. C. Chalub and Max O. Souza. Fixation in large populations: a continuous view of a discrete problem. Journal of Mathematical Biology, 72(1–2):283–330, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0889-9.pdf.

Cha:2016:SMR

[2470] Ji Hwan Cha and Maxim Finkelstein. On some mortality rate processes and mortality deceleration with age. Journal of Mathematical Biology, 72(1–2):331–342, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0885-0.pdf.

Xiao:2016:CTI

[2471] Yanyu Xiao, Fred Brauer, and Seyed M. Moghadas. Can treatment increase the epidemic size? Journal of Mathematical Biology, 72(1–2):343– 361, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0887-y.pdf.

Sainudiin:2016:ARD

[2472] R. Sainudiin, B. Thatte, and A. Véber. Ancestries of a recombining diploid population. Journal of Mathematical Biology, 72(1–2):363– 408, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0886-z.pdf.

Gholami:2016:IPF

[2473] Amir Gholami, Andreas Mang, and George Biros. An inverse problem formulation for parameter estimation of a reaction–diffusion model of low grade gliomas. Journal of Mathematical Biology, 72(1–2):409– 433, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0888-x.pdf.

Pfab:2016:RPP

[2474] Ferdinand Pfab, Wilfried Gabriel, and Margarete Utz. Reversible phenotypic plasticity with continuous adaptation. Journal of Mathematical Biology, 72(1–2):435–466, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0890-3.pdf.

Johnston:2016:CAP

[2475] Matthew D. Johnston, Casian Pantea, and Pete Donnell. A computational approach to persistence, permanence, and endotacticity of biochemical reaction systems. Journal of Mathematical Biology, 72(1–2): 467–498, January 2016. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-015-0892-1.pdf.

Gou:2016:HMS

[2476] Kun Gou and Thomas J. Pence. Hyperelastic modeling of swelling in fibrous soft tissue with application to tracheal angioedema. Journal of Mathematical Biology, 72(1–2):499–526, January 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0893-0.pdf.

Jin:2016:RSS

[2477] Emma Yu Jin and Markus E. Nebel. RNA secondary structures in a polymer-zeta model how foldings should be shaped for sparsification to establish a linear speedup. Journal of Mathematical Biology, 72(3):527– 571, February 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0894-z.pdf.

Costa:2016:SEE

[2478] Manon Costa, Céline Hauzy, Nicolas Loeuille, and Sylvie Méléard. Stochastic eco-evolutionary model of a prey-predator community. Journal of Mathematical Biology, 72(3):573–622, February 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0895-y.pdf.

Ibanez-Marcelo:2016:SEE

[2479] Esther Ibáñez-Marcelo and Tomás Alarcón. Surviving evolutionary escape on complex genotype-phenotype networks. Journal of Mathematical Biology, 72(3):623–647, February 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0896-x.pdf.

Tosenberger:2016:MPF

[2480] A. Tosenberger, F. Ataullakhanov, N. Bessonov, M. Panteleev, A. Tokarev, and V. Volpert. Modelling of platelet-fibrin clot formation in flow with a DPD–PDE method. Journal of Mathematical Biology, 72(3):649–681, February 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0891-2.pdf.

Ruiz-Herrera:2016:ADE

[2481] Alfonso Ruiz-Herrera. Analysis of dispersal effects in metapopulation models. Journal of Mathematical Biology, 72(3):683–698, February 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-015-0897-9.pdf.

Huber:2016:SPN

[2482] Katharina T. Huber, Simone Linz, Vincent Moulton, and Taoyang Wu. Spaces of phylogenetic networks from generalized nearest-neighbor interchange operations. Journal of Mathematical Biology, 72(3):699–725, February 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0899-7.pdf.

Chartrand:2016:SGU

[2483] Thomas Chartrand, Gin McCollum, Douglas A. Hanes, and Richard D. Boyle. Symmetries of a generic utricular projection: neural connectivity and the distribution of utricular information. Journal of Mathematical Biology, 72(3):727–753, February 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0900-5.pdf.

Peng:2016:NPR

[2484] Rui Peng and Xiao-Qiang Zhao. A nonlocal and periodic reaction– diffusion–advection model of a single phytoplankton species. Journal of Mathematical Biology, 72(3):755–791, February 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0904-1.pdf.

Corander:2016:TMG

[2485] Jukka Corander, Odo Diekmann, and Timo Koski. A tribute to mats gyllenberg, on the occasion of his 60th birthday. Journal of Mathematical Biology, 72(4):793–795, March 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-0965-9.pdf.

Hamza:2016:EPI

[2486] Kais Hamza, Peter Jagers, and Fima C. Klebaner. On the establishment, persistence, and inevitable extinction of populations. Journal of Mathematical Biology, 72(4):797–820, March 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0903-2.pdf.

Jin:2016:PVE

[2487] Wen Jin, Hal L. Smith, and Horst R. Thieme. Persistence versus extinction for a class of discrete-time structured population models. Journal of Mathematical Biology, 72(4):821–850, March 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0898-8.pdf.

Vasilyeva:2016:ASP

[2488] Olga Vasilyeva, Frithjof Lutscher, and Mark Lewis. Analysis of spread and persistence for stream insects with winged adult stages. Journal of Mathematical Biology, 72(4):851–875, March 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-015-0932-x.pdf.

Diekmann:2016:CEU

[2489] Odo Diekmann, Philipp Getto, and Yukihiko Nakata. On the characteristic equation λ = α_{1} + (α_{2} + α_{3}λ)e^{−λ} and its use in the context of a cell population model. Journal of Mathematical Biology, 72(4):877– 908, March 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0918-8.pdf.

Nakaoka:2016:DHI

[2490] Shinji Nakaoka, Shingo Iwami, and Kei Sato. Dynamics of HIV infection in lymphoid tissue network. Journal of Mathematical Biology, 72(4): 909–938, March 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0940-x.pdf.

Jiang:2016:HCC

[2491] Jifa Jiang, Lei Niu, and Yi Wang. On heteroclinic cycles of competitive maps via carrying simplices. Journal of Mathematical Biology, 72(4):

939–972, March 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0920-1.pdf.

Yang:2016:MMS

[2492] Yifan Yang, Willi Jäger, Maria Neuss-Radu, and Thomas Richter. Mathematical modeling and simulation of the evolution of plaques in blood vessels. Journal of Mathematical Biology, 72(4):973–996, March 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0934-8.pdf.

Pacheco:2016:EDC

[2493] Jorge M. Pacheco, Francisco C. Santos, and Simon A. Levin. Evolutionary dynamics of collective index insurance. Journal of Mathematical Biology, 72(4):997–1010, March 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0939-3.pdf.

Metz:2016:FDA

[2494] Johan A. J. Metz and Stefan A. H. Geritz. Frequency dependence 3.0: an attempt at codifying the evolutionary ecology perspective. Journal of Mathematical Biology, 72(4):1011–1037, March 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0956-2.pdf.

Kisdi:2016:ADS

[2495] Éva Kisdi and Stefan A. H. Geritz. Adaptive dynamics of saturated polymorphisms. Journal of Mathematical Biology, 72(4):1039– 1079, March 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0948-2.pdf. See correction [3024].

Geritz:2016:MIN

[2496] Stefan A. H. Geritz, Johan A. J. Metz, and Claus Rueffler. Mutual invadability near evolutionarily singular strategies for multivariate traits, with special reference to the strongly convergence stable case. Journal of Mathematical Biology, 72(4):1081–1099, March 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0944-6.pdf.

Boldin:2016:EST

[2497] Barbara Boldin and Éva Kisdi. Evolutionary suicide through a non-catastrophic bifurcation: adaptive dynamics of pathogens with frequency-dependent transmission. Journal of Mathematical Biology, 72 (4):1101–1124, March 2016. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-015-0945-5.pdf.

Metz:2016:CEA

[2498] Johan A. Jacob Metz, Katerina Stanková, and Jacob Johansson. The canonical equation of adaptive dynamics for life histories: from fitnessreturns to selection gradients and Pontryagin’s maximum principle. Journal of Mathematical Biology, 72(4):1125–1152, March 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0938-4.pdf.

Szabo-Solticzky:2016:OED

[2499] András Szabó-Solticzky, Luc Berthouze, Istvan Z. Kiss, and Péter L. Simon. Oscillating epidemics in a dynamic network model: stochastic and mean-field analysis. Journal of Mathematical Biology, 72(5):1153– 1176, April 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0902-3.pdf.

Ouboter:2016:SSE

[2500] Tanneke Ouboter, Ronald Meester, and Pieter Trapman. Stochastic SIR epidemics in a population with households and schools. Journal of Mathematical Biology, 72(5):1177–1193, April 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0901-4.pdf.

Romanescu:2016:MTS

[2501] Razvan Romanescu and Rob Deardon. Modeling two strains of disease via aggregate-level infectivity curves. Journal of Mathematical Biology, 72(5):1195–1224, April 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0910-3.pdf.

Shi:2016:MOM

[2502] Chunmei Shi, Lingling Zhao, Junjie Wang, Chiping Zhang, Xiaohong Su, and Peijun Ma. Micro-object motion tracking based on the probability hypothesis density particle tracker. Journal of Mathematical Biology, 72(5):1225–1254, April 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0909-9.pdf.

Schattler:2016:DPM

[2503] Heinz Schättler, Urszula Ledzewicz, and Behrooz Amini. Dynamical properties of a minimally parameterized mathematical model for metronomic chemotherapy. Journal of Mathematical Biology, 72(5):1255– 1280, April 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0907-y.pdf.

Culos:2016:USP

[2504] G. J. Culos, D. D. Olesky, and P. van den Driessche. Using sign patterns to detect the possibility of periodicity in biological systems. Journal of Mathematical Biology, 72(5):1281–1300, April 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0906-z.pdf.

Badri:2016:ORD

[2505] H. Badri, K. Pitter, E. C. Holland, F. Michor, and K. Leder. Optimization of radiation dosing schedules for proneural glioblastoma. Journal of Mathematical Biology, 72(5):1301–1336, April 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0908-x.pdf.

Kosiuk:2016:GAG

[2506] Ilona Kosiuk and Peter Szmolyan. Geometric analysis of the Goldbeter minimal model for the embryonic cell cycle. Journal of Mathematical Biology, 72(5):1337–1368, April 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0905-0.pdf.

Durrett:2016:SMM

[2507] R. Durrett, J. Foo, and K. Leder. Spatial Moran models, II: cancer initiation in spatially structured tissue. Journal of Mathematical Biology, 72(5):1369–1400, April 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0912-1.pdf.

Hsu:2016:TMS

[2508] Sze-Bi Hsu and Ya-Tang Yang. Theory of a microfluidic serial dilution bioreactor for growth of planktonic and biofilm populations. Journal of Mathematical Biology, 72(5):1401–1427, April 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0913-0.pdf.

Du:2016:PES

[2509] Zengji Du and Rui Peng. A priori L^{∞} estimates for solutions of a class of reaction–diffusion systems. Journal of Mathematical Biology, 72(6): 1429–1439, May 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0914-z.pdf.

Bendahmane:2016:TPD

[2510] Mostafa Bendahmane, Ricardo Ruiz-Baier, and Canrong Tian. Turing pattern dynamics and adaptive discretization for a super-diffusive Lotka–Volterra model. Journal of Mathematical Biology, 72(6):1441– 1465, May 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0917-9.pdf.

Guiver:2016:RSP

[2511] Chris Guiver, Markus Mueller, Dave Hodgson, and Stuart Townley. Robust set-point regulation for ecological models with multiple management goals. Journal of Mathematical Biology, 72(6):1467– 1529, May 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0919-7.pdf.

Silva:2016:CFH

[2512] Jacques A. L. Silva. Cluster formation in a heterogeneous metapopulation model. Journal of Mathematical Biology, 72(6):1531–1553, May 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0916-x.pdf.

Davis:2016:MMD

[2513] Jason K. Davis and Suzanne S. Sindi. A mathematical model of the dynamics of prion aggregates with chaperone-mediated fragmentation. Journal of Mathematical Biology, 72(6):1555–1578, May 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0921-0.pdf.

Ryan:2016:MCD

[2514] Shawn D. Ryan. A model for collective dynamics in ant raids. Journal of Mathematical Biology, 72(6):1579–1606, May 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0929-5.pdf.

Bernuau:2016:SIA

[2515] Emmanuel Bernuau, Denis Efimov, and Wilfrid Perruquetti. Scale invariance analysis for genetic networks applying homogeneity. Journal of Mathematical Biology, 72(6):1607–1632, May 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-015-0923-y.pdf.

Hanin:2016:UMM

[2516] Leonid Hanin, Karen Seidel, and Dietrich Stoevesandt. A ”universal” model of metastatic cancer, its parametric forms and their identification: what can be learned from site-specific volumes of metastases. Journal of Mathematical Biology, 72(6):1633–1662, May 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0928-6.pdf.

Wolkowicz:2016:ELG

[2517] Gail S. K. Wolkowicz and Yuan Yuan. Effect of light on the growth of non-nitrogen-fixing and nitrogen-fixing phytoplankton in an aquatic system. Journal of Mathematical Biology, 72(6):1663– 1692, May 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0924-x.pdf.

Berestycki:2016:PCP

[2518] Henri Berestycki, Jérôme Coville, and Hoang-Hung Vo. Persistence criteria for populations with non-local dispersion. Journal of Mathematical Biology, 72(7):1693–1745, June 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0911-2.pdf.

Martini:2016:CBM

[2519] Johannes W. R. Martini, Luis Diambra, and Michael Habeck. Cooperative binding: a multiple personality. Journal of Mathematical Biology, 72(7):1747–1774, June 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0922-z.pdf.

Dyson:2016:IIE

[2520] R. J. Dyson, J. E. F. Green, J. P. Whiteley, and H. M. Byrne. An investigation of the influence of extracellular matrix anisotropy and cell– matrix interactions on tissue architecture. Journal of Mathematical Biology, 72(7):1775–1809, June 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0927-7.pdf.

Ranwez:2016:IGD

[2521] Vincent Ranwez, Celine Scornavacca, Jean-Philippe Doyon, and Vincent Berry. Inferring gene duplications, transfers and losses can be done in a discrete framework. Journal of Mathematical Biology, 72(7):1811– 1844, June 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0930-z.pdf.

Nie:2016:ACW

[2522] Hua Nie, Sze-Bi Hsu, and J. P. Grover. Algal competition in a water column with excessive dioxide in the atmosphere. Journal of Mathematical Biology, 72(7):1845–1892, June 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0926-8.pdf.

Wang:2016:CAA

[2523] Qixuan Wang and Hans G. Othmer. Computational analysis of amoeboid swimming at low Reynolds number. Journal of Mathematical Biology, 72(7):1893–1926, June 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0925-9.pdf.

Giordano:2016:CSI

[2524] Giulia Giordano, Christian Cuba Samaniego, Elisa Franco, and Franco Blanchini. Computing the structural influence matrix for biological systems. Journal of Mathematical Biology, 72(7):1927–1958, June 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-015-0933-9.pdf.

Roberts:2016:ROD

[2525] Paul A. Roberts, Eamonn A. Gaffney, Philip J. Luthert, Alexander J. E. Foss, and Helen M. Byrne. Retinal oxygen distribution and the role of neuroglobin. Journal of Mathematical Biology, 73(1):1– 38, July 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0931-y.pdf.

Borri:2016:GSP

[2526] Alessandro Borri, Simona Panunzi, and Andrea De Gaetano. A glycemiastructured population model. Journal of Mathematical Biology, 73(1): 39–62, July 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0935-7.pdf.

Dawes:2016:LPF

[2527] J. H. P. Dawes and J. L. M. Williams. Localised pattern formation in a model for dryland vegetation. Journal of Mathematical Biology, 73 (1):63–90, July 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0937-5.pdf.

Economou:2016:DSE

[2528] Antonis Economou and Maria Jesus Lopez-Herrero. The deterministic SIS epidemic model in a Markovian random environment. Journal of Mathematical Biology, 73(1):91–121, July 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0943-7.pdf.

Drmota:2016:SAE

[2529] Michael Drmota, Michael Fuchs, and Yi-Wen Lee. Stochastic analysis of the extra clustering model for animal grouping. Journal of Mathematical Biology, 73(1):123–159, July 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0941-9.pdf.

Esser:2016:PDL

[2530] Mareike Esser, Sebastian Probst, and Ellen Baake. Partitioning, duality, and linkage disequilibria in the Moran model with recombination. Journal of Mathematical Biology, 73(1):161–197, July 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0936-6.pdf.

Sherratt:2016:WDC

[2531] Jonathan A. Sherratt. When does colonisation of a semi-arid hillslope generate vegetation patterns? Journal of Mathematical Biology, 73(1):

199–226, July 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0942-8.pdf.

Tilles:2016:HAM

[2532] Paulo F. C. Tilles and Sergei V. Petrovskii. How animals move along? Exactly solvable model of superdiffusive spread resulting from animal’s decision making. Journal of Mathematical Biology, 73(1):227– 255, July 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0947-3.pdf.

Graham:2016:EDS

[2533] Matthew Graham and Thomas House. Erratum to: Dynamics of stochastic epidemics on heterogeneous networks. Journal of Mathematical Biology, 73(1):257–258, July 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-1004-6.pdf. See [2268].

Jarvis:2016:MGS

[2534] Peter D. Jarvis and Jeremy G. Sumner. Matrix group structure and Markov invariants in the strand symmetric phylogenetic substitution model. Journal of Mathematical Biology, 73(2):259–282, August 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0951-7.pdf.

Bordewich:2016:DPN

[2535] Magnus Bordewich and Charles Semple. Determining phylogenetic networks from inter-taxa distances. Journal of Mathematical Biology, 73(2): 283–303, August 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0950-8.pdf.

Demasse:2016:OCA

[2536] Ramses Djidjou Demasse, Jean-Jules Tewa, Samuel Bowong, and Yves Emvudu. Optimal control for an age-structured model for the transmission of hepatitis B. Journal of Mathematical Biology, 73(2):305– 333, August 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0952-6.pdf.

Vikulova:2016:MEF

[2537] Nathalie A. Vikulova, Leonid B. Katsnelson, Alexander G. Kursanov, Olga Solovyova, and Vladimir S. Markhasin. Mechano-electric feedback in one-dimensional model of myocardium. Journal of Mathematical Biology, 73(2):335–366, August 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0953-5.pdf.

Mackey:2016:LDB

[2538] Michael C. Mackey and Marta Tyran-Kamińska. The limiting dynamics of a bistable molecular switch with and without noise. Journal of Mathematical Biology, 73(2):367–395, August 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0949-1.pdf.

Roques:2016:UGD

[2539] L. Roques, E. Walker, P. Franck, S. Soubeyrand, and E. K. Klein. Using genetic data to estimate diffusion rates in heterogeneous landscapes. Journal of Mathematical Biology, 73(2):397–422, August 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0954-4.pdf.

Yang:2016:SSE

[2540] J. Yang, I. Kupka, Z. Schuss, and D. Holcman. Search for a small egg by spermatozoa in restricted geometries. Journal of Mathematical Biology, 73(2):423–446, August 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0955-3.pdf.

Forcey:2016:FBM

[2541] Stefan Forcey, Logan Keefe, and William Sands. Facets of the balanced minimal evolution polytope. Journal of Mathematical Biology, 73(2): 447–468, August 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0957-1.pdf.

Coloma:2016:PRB

[2542] M. Coloma, J. D. Schaffer, R. O. Carare, P. R. Chiarot, and P. Huang. Pulsations with reflected boundary waves: a hydrodynamic reverse transport mechanism for perivascular drainage in the brain. Journal of Mathematical Biology, 73(2):469–490, August 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-015-0960-6.pdf.

Weber:2016:IAG

[2543] Tom S. Weber, Leïla Perié, and Ken R. Duffy. Inferring average generation via division-linked labeling. Journal of Mathematical Biology, 73(2): 491–523, August 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0963-3.pdf.

Wang:2016:STF

[2544] Xiaohui Wang and Martin Golubitsky. Singularity theory of fitness functions under dimorphism equivalence. Journal of Mathematical Biology, 73(3):525–573, September 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-015-0958-0.pdf.

Subbey:2016:ESI

[2545] Sam Subbey, Benjamin Planque, and Ulf Lindstrøm. Exploring stochasticity and imprecise knowledge based on linear inequality constraints. Journal of Mathematical Biology, 73(3):575–595, September 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-015-0959-z.pdf.

Liu:2016:AST

[2546] Meng Liu and Chuanzhi Bai. Analysis of a stochastic tri-trophic foodchain model with harvesting. Journal of Mathematical Biology, 73(3): 597–625, September 2016. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-016-0970-z.pdf.

Delaporte:2016:MPS

[2547] Cécile Delaporte, Guillaume Achaz, and Amaury Lambert. Mutational pattern of a sample from a critical branching population. Journal of Mathematical Biology, 73(3):627–664, September 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-015-0964-2.pdf.

Yuan:2016:DIR

[2548] Sanling Yuan, P. van den Driessche, Frederick H. Willeboordse, Zhisheng Shuai, and Junling Ma. Disease invasion risk in a growing population. Journal of Mathematical Biology, 73(3):665–681, September 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-015-0962-4.pdf.

Bengfort:2016:FPL

[2549] Michael Bengfort, Horst Malchow, and Frank M. Hilker. The Fokker– Planck law of diffusion and pattern formation in heterogeneous environments. Journal of Mathematical Biology, 73(3):683–704, September 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-016-0966-8.pdf.

Montenegro-Johnson:2016:TDF

[2550] T. D. Montenegro-Johnson, D. I. Baker, D. J. Smith, and S. S. Lopes. Three-dimensional flow in Kupffer’s Vesicle. Journal of Mathematical Biology, 73(3):705–725, September 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-0967-7.pdf.

Thalhofer:2016:SMC

[2551] V. Thalhofer, M. Annunziato, and A. Borzì. Stochastic modelling and control of antibiotic subtilin production. Journal of Mathematical Biology, 73(3):727–749, September 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-0968-6.pdf.

Chong:2016:AOF

[2552] Nyuk Sian Chong, Benoit Dionne, and Robert Smith. An avian-only Filippov model incorporating culling of both susceptible and infected birds in combating avian influenza. Journal of Mathematical Biology, 73(3):751–784, September 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-0971-y.pdf. See erratum [2553].

Chong:2016:EAO

[2553] Nyuk Sian Chong, Benoit Dionne, and Robert Smith. Erratum to: An avian-only Filippov model incorporating culling of both susceptible and infected birds in combating avian influenza. Journal of Mathematical Biology, 73(3):785–786, September 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-1003-7.pdf. See [2552].

Yan:2016:EPM

[2554] Ada W. C. Yan, Pengxing Cao, and James M. McCaw. On the extinction probability in models of within-host infection: the role of latency and immunity. Journal of Mathematical Biology, 73(4):787–813, October 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-015-0961-5.pdf.

Schlagel:2016:FAR

[2555] Ulrike E. Schlägel and Mark A. Lewis. A framework for analyzing the robustness of movement models to variable step discretization. Journal of Mathematical Biology, 73(4):815–845, October 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-0969-5.pdf.

Bacaer:2016:MSS

[2556] Nicolas Bacaër. Le modèle stochastique SIS pour une épidémie dans un environnement aléatoire. (French) [The SIS stochastic model for an epidemic in a random environment]. Journal of Mathematical Biology, 73(4):847–866, October 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-0974-8.pdf.

Forys:2016:ADS

[2557] U. Foryś, M. Bodnar, and Y. Kogan. Asymptotic dynamics of some t-periodic one-dimensional model with application to prostate cancer immunotherapy. Journal of Mathematical Biology, 73(4):867–883, October 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-0978-4.pdf.

Rozins:2016:DEL

[2558] Carly Rozins and Troy Day. Disease eradication on large industrial farms. Journal of Mathematical Biology, 73(4):885–902, October 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-0973-9.pdf.

Parvinen:2016:FMB

[2559] Kalle Parvinen and Anne Seppänen. On fitness in metapopulations that are both size- and stage-structured. Journal of Mathematical Biology, 73(4):903–917, October 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-0975-7.pdf.

Sanchez-Taltavull:2016:SME

[2560] Daniel Sánchez-Taltavull, Arturo Vieiro, and Tomás Alarcón. Stochastic modelling of the eradication of the HIV-1 infection by stimulation of latently infected cells in patients under highly active antiretroviral therapy. Journal of Mathematical Biology, 73(4):919–946, October 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-0977-5.pdf.

Zhang:2016:BBT

[2561] Wenjing Zhang, Lindi M. Wahl, and Pei Yu. Backward bifurcations, turning points and rich dynamics in simple disease models. Journal of Mathematical Biology, 73(4):947–976, October 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-016-0976-6.pdf.

Xue:2016:MFM

[2562] Chuan Xue and Xige Yang. Moment-flux models for bacterial chemotaxis in large signal gradients. Journal of Mathematical Biology, 73(4):977– 1000, October 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-0981-9.pdf.

Busse:2016:MCN

[2563] J.-E. Busse, P. Gwiazda, and A. Marciniak-Czochra. Mass concentration in a nonlocal model of clonal selection. Journal of Mathematical Biology, 73(4):1001–1033, October 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-0979-3.pdf.

Sircar:2016:SDS

[2564] Sarthok Sircar and Anthony J. Roberts. Surface deformation and shear flow in ligand mediated cell adhesion. Journal of Mathematical Biology, 73(4):1035–1052, October 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-0983-7.pdf.

Yin:2016:PPS

[2565] Changchuan Yin and Jiasong Wang. Periodic power spectrum with applications in detection of latent periodicities in DNA sequences. Journal of Mathematical Biology, 73(5):1053–1079, November 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-0982-8.pdf.

Kan:2016:MTS

[2566] Xingye Kan, Chang Hyeong Lee, and Hans G. Othmer. A multi-time-scale analysis of chemical reaction networks: II. Stochastic systems. Journal of Mathematical Biology, 73(5):1081–1129, November 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-0980-x.pdf.

McCleney:2016:ESN

[2567] Zachary T. McCleney and Zachary P. Kilpatrick. Entrainment in up and down states of neural populations: non-smooth and stochastic models. Journal of Mathematical Biology, 73(5):1131–1160, November 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-016-0984-6.pdf.

Perthame:2016:DBR

[2568] Benoît Perthame, Min Tang, and Nicolas Vauchelet. Derivation of the bacterial run-and-tumble kinetic equation from a model with biochemical pathway. Journal of Mathematical Biology, 73(5):1161–1178, November 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-0985-5.pdf.

Wang:2016:MFE

[2569] Xiaoying Wang, Liana Zanette, and Xingfu Zou. Modelling the fear effect in predator–prey interactions. Journal of Mathematical Biology, 73(5):1179–1204, November 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-0989-1.pdf.

Menichini:2016:MMT

[2570] Claudia Menichini and Xiao Yun Xu. Mathematical modeling of thrombus formation in idealized models of aortic dissection: initial findings and potential applications. Journal of Mathematical Biology, 73(5):1205– 1226, November 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-0986-4.pdf.

Yang:2016:EIA

[2571] Junyuan Yang, Yuming Chen, and Fei Xu. Effect of infection age on an SIS epidemic model on complex networks. Journal of Mathematical Biology, 73(5):1227–1249, November 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-016-0991-7.pdf.

Popovic:2016:TIY

[2572] Lea Popovic and Mariolys Rivas. Topology and inference for Yule trees with multiple states. Journal of Mathematical Biology, 73(5):1251–1291, November 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-0992-6.pdf.

Gu:2016:TPF

[2573] Rui Gu, Xiaoqiang Wang, and Max Gunzburger. A two phase field model for tracking vesicle–vesicle adhesion. Journal of Mathematical Biology, 73(5):1293–1319, November 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-016-0994-4.pdf.

Ermentrout:2016:RWS

[2574] B. Ermentrout and B. I. S. van der Ventel. Rotating waves in simple scalar excitable media: approximations and numerical solutions. Journal of Mathematical Biology, 73(6–7):1321–1351, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-0988-2.pdf.

Boelle:2016:RAL

[2575] Pierre-Yves Boëlle and Guy Thomas. Resistance to antibiotics: limit theorems for a stochastic SIS model structured by level of resistance. Journal of Mathematical Biology, 73(6–7):1353–1378, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-0996-2.pdf.

Rasmussen:2016:TTM

[2576] Martin Rasmussen, Alan Hastings, Matthew J. Smith, Folashade B. Agusto, Benito M. Chen-Charpentier, Forrest M. Hoffman, Jiang Jiang, Katherine E. O. Todd-Brown, Ying Wang, Ying-Ping Wang, and Yiqi Luo. Transit times and mean ages for nonautonomous and autonomous compartmental systems. Journal of Mathematical Biology, 73(6–7):1379– 1398, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-0990-8.pdf.

Abeliovich:2016:HCS

[2577] Hagai Abeliovich. On Hill coefficients and subunit interaction energies. Journal of Mathematical Biology, 73(6–7):1399–1411, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1001-9.pdf.

Dumont:2016:NTN

[2578] G. Dumont, J. Henry, and C. O. Tarniceriu. Noisy threshold in neuronal models: connections with the noisy leaky integrate-and-fire model. Journal of Mathematical Biology, 73(6–7):1413–1436, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1002-8.pdf.

Doutor:2016:OVS

[2579] Paulo Doutor, Paula Rodrigues, Maria do Céu Soares, and Fabio A. C. C. Chalub. Optimal vaccination strategies and rational behaviour in seasonal epidemics. Journal of Mathematical Biology, 73(6–7):1437– 1465, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-0997-1.pdf.

Chung:2016:DTS

[2580] K. W. Chung and Roger Lui. Dynamics of two-strain influenza model with cross-immunity and no quarantine class. Journal of Mathematical Biology, 73(6–7):1467–1489, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-016-1000-x.pdf.

Belavkin:2016:MFL

[2581] Roman V. Belavkin, Alastair Channon, Elizabeth Aston, John Aston, Rok Krasovec, and Christopher G. Knight. Monotonicity of fitness landscapes and mutation rate control. Journal of Mathematical Biology, 73(6–7):1491–1524, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-0995-3.pdf.

Zhao:2016:ZVL

[2582] Songnian Zhao, Yan Kuang, Chih-Hang Wu, David Ben-Arieh, Marcelo Ramalho-Ortigao, and Kaiming Bi. Zoonotic visceral leishmaniasis transmission: modeling, backward bifurcation, and optimal control. Journal of Mathematical Biology, 73(6–7):1525–1560, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-0999-z.pdf.

Peng:2016:PID

[2583] Xiao-Long Peng, Xin-Jian Xu, Michael Small, Xinchu Fu, and Zhen Jin. Prevention of infectious diseases by public vaccination and individual protection. Journal of Mathematical Biology, 73(6–7):1561–1594, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1007-3.pdf.

Gandolfi:2016:SSE

[2584] Alberto Gandolfi, Mimmo Iannelli, and Gabriela Marinoschi. The steady state of epidermis: mathematical modeling and numerical simulations. Journal of Mathematical Biology, 73(6–7):1595–1626, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1006-4.pdf.

Bigan:2016:NSC

[2585] Erwan Bigan, Loïc Paulevé, Jean-Marc Steyaert, and Stéphane Douady. Necessary and sufficient conditions for protocell growth. Journal of Mathematical Biology, 73(6–7):1627–1664, December 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-0998-0.pdf.

Rodriguez:2016:ECP

[2586] Irene Núñez Rodríguez and Armando G. M. Neves. Evolution of cooperation in a particular case of the infinitely repeated prisoner’s dilemma with three strategies. Journal of Mathematical Biology, 73(6–7):1665– 1690, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1009-1.pdf.

Schlagel:2016:RMM

[2587] Ulrike E. Schlägel and Mark A. Lewis. Robustness of movement models: can models bridge the gap between temporal scales of data sets and behavioural processes? Journal of Mathematical Biology, 73(6–7):1691– 1726, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1005-5.pdf.

Duong:2016:AED

[2588] Manh Hong Duong and The Anh Han. Analysis of the expected density of internal equilibria in random evolutionary multi-player multistrategy games. Journal of Mathematical Biology, 73(6–7):1727–1760, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1010-8.pdf.

Huber:2016:FUP

[2589] Katharina T. Huber, Vincent Moulton, Mike Steel, and Taoyang Wu. Folding and unfolding phylogenetic trees and networks. Journal of Mathematical Biology, 73(6–7):1761–1780, December 2016. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-0993-5.pdf.

Campillo:2016:LBD

[2590] Fabien Campillo, Nicolas Champagnat, and Coralie Fritsch. Links between deterministic and stochastic approaches for invasion in growth– fragmentation–death models. Journal of Mathematical Biology, 73(6–7): 1781–1821, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00285-016-1012-6.pdf.

Calvetti:2016:UQF

[2591] Daniela Calvetti, Yougan Cheng, and Erkki Somersalo. Uncertainty quantification in flux balance analysis of spatially lumped and distributed models of neuron–astrocyte metabolism. Journal of Mathematical Biology, 73(6–7):1823–1849, December 2016. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-1011-7.pdf.

Hirsch:2017:MMM

[2592] S. Hirsch, A. Manhart, and C. Schmeiser. Mathematical modeling of Myosin induced bistability of Lamellipodial fragments. Journal of Mathematical Biology, 74(1–2):1–22, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-016-1008-2;
 http://link.springer.com/content/pdf/10.1007/s00285-016-1008-
2.pdf.

Enciso:2017:ESS

[2593] German A. Enciso and Shane Ryerson. The effect of site-to-site variability in ultrasensitive dose responses. Journal of Mathematical Biology, 74 (1–2):23–41, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-016-1013-5;   http://link.springer.com/content/
pdf/10.1007/s00285-016-1013-5.pdf.

Berman:2017:PRI

[2594] Gennady P. Berman, Alexander I. Nesterov, Shmuel Gurvitz, and Richard T. Sayre. Possible role of interference, protein noise, and sink effects in nonphotochemical quenching in photosynthetic complexes. Journal of Mathematical Biology, 74(1–2):43–76, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1016-2;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1016-2.pdf.

Lladser:2017:RPI

[2595] Manuel E. Lladser, Joseph G. Azofeifa, Mary A. Allen, and Robin D. Dowell. RNA Pol II transcription model and interpretation of GROseq data. Journal of Mathematical Biology, 74(1–2):77–97, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1014-4;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1014-4.pdf.

Layer:2017:PTE

[2596] Mark Layer and John A. Rhodes. Phylogenetic trees and Euclidean embeddings. Journal of Mathematical Biology, 74(1–2):99–111, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1018-0;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1018-0.pdf.

Massaccesi:2017:NDS

[2597] Annalisa Massaccesi and Enrico Valdinoci. Is a nonlocal diffusion strategy convenient for biological populations in competition? Journal of Mathematical Biology, 74(1–2):113–147, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1019-z;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1019-z.pdf.

Paniello:2017:EOG

[2598] Irene Paniello. On evolution operators of genetic coalgebras. Journal of Mathematical Biology, 74(1–2):149–168, January 2017. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-016-1025-1;
 http://link.springer.com/content/pdf/10.1007/s00285-016-1025-
1.pdf.

Manhart:2017:EDE

[2599] Angelika Manhart and Christian Schmeiser. Existence of and decay to equilibrium of the filament end density along the leading edge of the lamellipodium. Journal of Mathematical Biology, 74(1–2):169–193, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1027-z;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1027-z.pdf.

Saez:2017:GRR

[2600] Meritxell Sáez, Carsten Wiuf, and Elisenda Feliu. Graphical reduction of reaction networks by linear elimination of species. Journal of Mathematical Biology, 74(1–2):195–237, January 2017. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-016-1028-y;
 http://link.springer.com/content/pdf/10.1007/s00285-016-1028-
y.pdf.

Bordewich:2017:FPT

[2601] Magnus Bordewich, Celine Scornavacca, Nihan Tokac, and Mathias Weller. On the fixed parameter tractability of agreement-based phylogenetic distances. Journal of Mathematical Biology, 74(1–2):239–257, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1023-3;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1023-3.pdf.

Banks:2017:NSE

[2602] Harvey Thomas Banks, Marie Doumic, and Carola Kruse. A numerical scheme for the early steps of nucleation–aggregation models. Journal of Mathematical Biology, 74(1–2):259–287, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1026-0;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1026-0.pdf.

Fung:2017:SAD

[2603] Tak Fung, James P. O’Dwyer, and Ryan A. Chisholm. Speciesabundance distributions under colored environmental noise. Journal of Mathematical Biology, 74(1–2):289–311, January 2017. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-016-1022-4;
 http://link.springer.com/content/pdf/10.1007/s00285-016-1022-
4.pdf.

Ge:2017:BEE

[2604] Peinan Ge, William J. Bottega, Jonathan L. Prenner, and Howard F. Fine. On the behavior of an eye encircled by a scleral buckle. Journal of Mathematical Biology, 74(1–2):313–332, January 2017. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-016-1015-3;
 http://link.springer.com/content/pdf/10.1007/s00285-016-1015-
3.pdf.

Lee:2017:NAP

[2605] S. Seirin Lee, S. Tashiro, A. Awazu, and R. Kobayashi. A new application of the phase-field method for understanding the mechanisms of nuclear architecture reorganization. Journal of Mathematical Biology, 74(1–2): 333–354, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-016-1031-3;   http://link.springer.com/content/
pdf/10.1007/s00285-016-1031-3.pdf.

Ane:2017:PTC

[2606] Cécile Ané, Lam Si Tung Ho, and Sebastien Roch. Phase transition on the convergence rate of parameter estimation under an Ornstein– Uhlenbeck diffusion on a tree. Journal of Mathematical Biology, 74(1–2): 355–385, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-016-1029-x;   http://link.springer.com/content/
pdf/10.1007/s00285-016-1029-x.pdf.

Golubitsky:2017:HSN

[2607] Martin Golubitsky and Ian Stewart. Homeostasis, singularities, and networks. Journal of Mathematical Biology, 74(1–2):387–407, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1024-2;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1024-2.pdf.

Nev:2017:VIS

[2608] Olga A. Nev and Hugo A. van den Berg. Variable-internal-stores models of microbial growth and metabolism with dynamic allocation of cellular resources. Journal of Mathematical Biology, 74(1–2):409–445, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1030-4;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1030-4.pdf. See erratum [2612].

Jones:2017:AIS

[2609] Graham Jones. Algorithmic improvements to species delimitation and phylogeny estimation under the multispecies coalescent. Journal of Mathematical Biology, 74(1–2):447–467, January 2017. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-016-1034-0;
 http://link.springer.com/content/pdf/10.1007/s00285-016-1034-
0.pdf.

Lahrouz:2017:ESP

[2610] Aadil Lahrouz, Adel Settati, and Abdelhadi Akharif. Effects of stochastic perturbation on the SIS epidemic system. Journal of Mathematical Biology, 74(1–2):469–498, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-016-1033-1;
 http://link.springer.com/content/pdf/10.1007/s00285-016-1033-
1.pdf.

Molina:2017:ESC

[2611] Chai Molina and David J. D. Earn. Evolutionary stability in continuous nonlinear public goods games. Journal of Mathematical Biology, 74(1– 2):499–529, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-016-1017-1;   http://link.springer.com/content/
pdf/10.1007/s00285-016-1017-1.pdf.

Nev:2017:EVI

[2612] Olga A. Nev and Hugo A. van den Berg. Erratum to: Variable–Internal-Stores models of microbial growth and metabolism with dynamic allocation of cellular resources. Journal of Mathematical Biology, 74(1– 2):531, January 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-016-1044-y;   http://link.springer.com/content/
pdf/10.1007/s00285-016-1044-y.pdf. See [2608].

Groh:2017:NRF

[2613] Andreas Groh, Holger Kohr, and Alfred K. Louis. Numerical rate function determination in partial differential equations modeling cell population dynamics. Journal of Mathematical Biology, 74(3):533–565, February 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1032-2.pdf.

Bonifaci:2017:RMF

[2614] Vincenzo Bonifaci. A revised model of fluid transport optimization in Physarum polycephalum. Journal of Mathematical Biology, 74(3):567– 581, February 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1036-y.pdf.

Marciniak-Czochra:2017:ITP

[2615] Anna Marciniak-Czochra, Grzegorz Karch, and Kanako Suzuki. Instability of Turing patterns in reaction–diffusion-ODE systems. Journal of Mathematical Biology, 74(3):583–618, February 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1035-z.pdf.

Leung:2017:DCB

[2616] Ka Yin Leung and Odo Diekmann. Dangerous connections: on binding site models of infectious disease dynamics. Journal of Mathematical Biology, 74(3):619–671, February 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-1037-x.pdf.

Artalejo:2017:SDS

[2617] J. R. Artalejo, A. Gómez-Corral, M. López-García, and C. Molina-París. Stochastic descriptors to study the fate and potential of naive T cell clonotypes in the periphery. Journal of Mathematical Biology, 74(3):673– 708, February 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1020-6.pdf.

Rao:2017:GSC

[2618] Shodhan Rao. Global stability of a class of futile cycles. Journal of Mathematical Biology, 74(3):709–726, February 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00285-016-1039-8.pdf.

Dallon:2017:CTM

[2619] J. C. Dallon, Lynnae C. Despain, Emily J. Evans, Christopher P. Grant, and W. V. Smith. A continuous-time model of centrally coordinated motion with random switching. Journal of Mathematical Biology, 74(3):727– 753, February 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1040-2.pdf.

Reimer:2017:CDS

[2620] Jody R. Reimer, Michael B. Bonsall, and Philip K. Maini. The critical domain size of stochastic population models. Journal of Mathematical Biology, 74(3):755–782, February 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-1021-5.pdf.

Colombi:2017:CMS

[2621] A. Colombi, M. Scianna, and L. Preziosi. Coherent modelling switch between pointwise and distributed representations of cell aggregates. Journal of Mathematical Biology, 74(4):783–808, March 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1042-0.pdf.

Fermin:2017:VSA

[2622] Lisandro J. Fermín and Jacques Lévy-Véhel. Variability and singularity arising from poor compliance in a pharmacokinetic model II: the multi-oral case. Journal of Mathematical Biology, 74(4):809– 841, March 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1041-1.pdf.

Janson:2017:NCS

[2623] Svante Janson, Malwina Luczak, Peter Windridge, and Thomas House. Near-critical SIR epidemic on a random graph with given degrees. Journal of Mathematical Biology, 74(4):843–886, March 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1043-z.pdf.

deFreitas:2017:ICP

[2624] Michael Marcondes de Freitas, Elisenda Feliu, and Carsten Wiuf. Intermediates, catalysts, persistence, and boundary steady states. Journal of Mathematical Biology, 74(4):887–932, March 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1046-9.pdf.

Milisic:2017:MML

[2625] Vuk Milisic and Gilles Wainrib. Mathematical modeling of lymphocytes selection in the germinal center. Journal of Mathematical Biology, 74(4): 933–979, March 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1038-9.pdf.

Sarhad:2017:GIP

[2626] Jonathan Sarhad, Scott Manifold, and Kurt E. Anderson. Geometric indicators of population persistence in branching continuousspace networks. Journal of Mathematical Biology, 74(4):981–1009, March 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1045-x.pdf.

Liu:2017:HDH

[2627] Sensen Liu and ShiNung Ching. Homeostatic dynamics, hysteresis and synchronization in a low-dimensional model of burst suppression. Journal of Mathematical Biology, 74(4):1011–1035, March 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1048-7.pdf.

Avesani:2017:ASP

[2628] Diego Avesani, Michael Dumbser, Gabriele Chiogna, and Alberto Bellin. An alternative smooth particle hydrodynamics formulation to simulate chemotaxis in porous media. Journal of Mathematical Biology, 74(5): 1037–1058, April 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1049-6.pdf.

Iniguez:2017:SPF

[2629] Abdon Iniguez and Jun Allard. Spatial pattern formation in microtubule post-translational modifications and the tight localization of motor-driven cargo. Journal of Mathematical Biology, 74(5):1059– 1080, April 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1053-x.pdf.

Tsachev:2017:SME

[2630] Tsvetomir Tsachev, Vladimir M. Veliov, and Andreas Widder. Setmembership estimations for the evolution of infectious diseases in heterogeneous populations. Journal of Mathematical Biology, 74(5):1081– 1106, April 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1050-0.pdf.

Casanellas:2017:PML

[2631] Marta Casanellas and Mike Steel. Phylogenetic mixtures and linear invariants for equal input models. Journal of Mathematical Biology, 74(5): 1107–1138, April 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1055-8.pdf.

Mackey:2017:ROD

[2632] Michael C. Mackey, Marta Tyran-Kamińska, and Hans-Otto Walther. Response of an oscillatory differential delay equation to a single stimulus. Journal of Mathematical Biology, 74(5):1139–1196, April 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1051-z.pdf.

DeVille:2017:FSE

[2633] Lee DeVille and Meghan Galiardi. Finite-size effects and switching times for Moran process with mutation. Journal of Mathematical Biology, 74(5):1197–1222, April 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-1056-7.pdf.

Jiang:2017:ECT

[2634] Jifa Jiang and Lei Niu. On the equivalent classification of threedimensional competitive Leslie/Gower models via the boundary dynamics on the carrying simplex. Journal of Mathematical Biology, 74(5): 1223–1261, April 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1052-y.pdf.

Sun:2017:ESA

[2635] Mengfeng Sun, Haifeng Zhang, Huiyan Kang, Guanghu Zhu, and Xinchu Fu. Epidemic spreading on adaptively weighted scale-free networks. Journal of Mathematical Biology, 74(5):1263–1298, April 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1057-6.pdf.

Foxall:2017:SER

[2636] Eric Foxall and Nicolas Lanchier. Survival and extinction results for a patch model with sexual reproduction. Journal of Mathematical Biology, 74(6):1299–1349, May 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-1058-5.pdf.

Abdelrazec:2017:MAR

[2637] Ahmed Abdelrazec and Abba B. Gumel. Mathematical assessment of the role of temperature and rainfall on mosquito population dynamics. Journal of Mathematical Biology, 74(6):1351–1395, May 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1054-9.pdf.

Cai:2017:GPV

[2638] Li-Ming Cai, Xue-Zhi Li, Bin Fang, and Shigui Ruan. Global properties of vector-host disease models with time delays. Journal of Mathematical Biology, 74(6):1397–1423, May 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-1047-8.pdf.

Yatat:2017:IMF

[2639] V. Yatat, P. Couteron, J. J. Tewa, S. Bowong, and Y. Dumont. An impulsive modelling framework of fire occurrence in a size-structured model of tree-grass interactions for savanna ecosystems. Journal of Mathematical Biology, 74(6):1425–1482, May 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-1060-y.pdf.

Bokes:2017:GEN

[2640] Pavol Bokes and Abhyudai Singh. Gene expression noise is affected differentially by feedback in burst frequency and burst size. Journal of Mathematical Biology, 74(6):1483–1509, May 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-016-1059-4.pdf.

DOnofrio:2017:TBF

[2641] Giuseppe D’Onofrio and Enrica Pirozzi. Two-boundary first exit time of Gauss–Markov processes for stochastic modeling of acto-myosin dynamics. Journal of Mathematical Biology, 74(6):1511–1531, May 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-016-1061-x.pdf.

Pimenov:2017:MAB

[2642] Alexander Pimenov, Thomas C. Kelly, Andrei Korobeinikov, Michael J. O’Callaghan, and Dmitrii Rachinskii. Memory and adaptive behavior in population dynamics: anti-predator behavior as a case study. Journal of Mathematical Biology, 74(6):1533–1559, May 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1065-6.pdf.

Bajeux:2017:ABW

[2643] Nicolas Bajeux, Frédéric Grognard, and Ludovic Mailleret. Augmentative biocontrol when natural enemies are subject to Allee effects. Journal of Mathematical Biology, 74(7):1561–1587, June 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1063-8.pdf.

Sardanyes:2017:ATT

[2644] Josep Sardanyés, Regina Martínez, Carles Simó, and Ricard Solé. Abrupt transitions to tumor extinction: a phenotypic quasispecies model. Journal of Mathematical Biology, 74(7):1589–1609, June 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1062-9.pdf.

Hart:2017:EBT

[2645] Andrew Hart and Servet Martínez. An entropy-based technique for classifying bacterial chromosomes according to synonymous codon usage. Journal of Mathematical Biology, 74(7):1611–1625, June 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1067-4.pdf.

Zhang:2017:RPW

[2646] Calvin Zhang and Timothy J. Lewis. Robust phase-waves in chains of half-center oscillators. Journal of Mathematical Biology, 74(7):1627– 1656, June 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1066-5.pdf.

Mancini:2017:PMC

[2647] Simona Mancini, René-Marc Mège, Benoit Sarels, and Pierre-Olivier Strale. A phenomenological model of cell–cell adhesion mediated by cadherins. Journal of Mathematical Biology, 74(7):1657–1678, June 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-016-1072-7.pdf.

Stefan:2017:CCD

[2648] Melanie I. Stefan. Cooperativity: a competition of definitions. Journal of Mathematical Biology, 74(7):1679–1681, June 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1069-2.pdf.

Gamado:2017:EUR

[2649] Kokouvi Gamado, George Streftaris, and Stan Zachary. Estimation of under-reporting in epidemics using approximations. Journal of Mathematical Biology, 74(7):1683–1707, June 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-016-1064-7.pdf.

Paeng:2017:CDS

[2650] Seong-Hun Paeng and Jonggul Lee. Continuous and discrete SIR-models with spatial distributions. Journal of Mathematical Biology, 74(7):1709– 1727, June 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1071-8.pdf.

Gambette:2017:CRL

[2651] Philippe Gambette, K. T. Huber, and S. Kelk. On the challenge of reconstructing level-1 phylogenetic networks from triplets and clusters. Journal of Mathematical Biology, 74(7):1729–1751, June 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00285-016-1068-3.pdf.

Rowthorn:2017:OTI

[2652] Robert Rowthorn and Selma Walther. The optimal treatment of an infectious disease with two strains. Journal of Mathematical Biology, 74(7): 1753–1791, June 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1074-5.pdf.

Li:2017:STR

[2653] Thomas J. X. Li and Christian M. Reidys. Statistics of topological RNA structures. Journal of Mathematical Biology, 74(7):1793– 1821, June 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1078-1.pdf.

Koppenol:2017:MMS

[2654] Daniël C. Koppenol, Fred J. Vermolen, Gabriela V. Koppenol-Gonzalez, Frank B. Niessen, Paul P. M. van Zuijlen, and Kees Vuik. A mathematical model for the simulation of the contraction of burns. Journal of Mathematical Biology, 75(1):1–31, July 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-016-1075-4.pdf.

Aston:2017:MAR

[2655] Philip J. Aston, Gianne Derks, Balaji M. Agoram, and Piet H. van der Graaf. A mathematical analysis of rebound in a target-mediated drug disposition model: II. With feedback. Journal of Mathematical Biology, 75 (1):33–84, July 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1073-6.pdf.

Bacaer:2017:PLN

[2656] Nicolas Bacaër. Sur les processus linéaires de naissance et de mort souscritiques dans un environnement aléatoire. (French) [On subcritical linear birth and death processes in a random environment]. Journal of Mathematical Biology, 75(1):85–108, July 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-016-1079-0.pdf.

Rus:2017:AST

[2657] Jernej Rus. Antiparallel d-stable traces and a stronger version of ore problem. Journal of Mathematical Biology, 75(1):109–127, July 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00285-016-1077-2.pdf.

Wang:2017:THU

[2658] Lei Wang and Xiao-Song Yang. Topological horseshoe and its uniform hyperbolicity in the HP model. Journal of Mathematical Biology, 75(1): 129–143, July 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1076-3.pdf.

Neukirch:2017:SRA

[2659] Rebecca Neukirch and Anton Bovier. Survival of a recessive allele in a Mendelian diploid model. Journal of Mathematical Biology, 75(1):

145–198, July 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-016-1081-6.pdf.

Hellmuth:2017:MXC

[2660] Marc Hellmuth, Peter F. Stadler, and Nicolas Wieseke. The mathematics of xenology: di-cographs, symbolic ultrametrics, 2-structures and tree-representable systems of binary relations. Journal of Mathematical Biology, 75(1):199–237, July 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00285-016-1084-3.pdf.

He:2017:PZD

[2661] Xiao He and Sining Zheng. Protection zone in a diffusive predator–prey model with Beddington–DeAngelis functional response. Journal of Mathematical Biology, 75(1):239–257, July 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00285-016-1082-5.pdf.

Diekmann:2017:EVD

[2662] Odo Diekmann, Mats Gyllenberg, J. A. J. Metz, Shinji Nakaoka, and André M. de Roos. Erratum to: Daphnia revisited: local stability and bifurcation theory for physiologically structured population models explained by way of an example. Journal of Mathematical Biology, 75(1): 259–261, July 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00285-017-1148-z.pdf. See [1888].

Gallinato:2017:FBP

[2663] Olivier Gallinato, Masahito Ohta, Clair Poignard, and Takashi Suzuki. Free boundary problem for cell protrusion formations: theoretical and numerical aspects. Journal of Mathematical Biology, 75(2):263–307, August 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1080-7;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1080-7.pdf.

Morsky:2017:HRE

[2664] Bryce Morsky, Ross Cressman, and C. T. Bauch. Homophilic replicator equations. Journal of Mathematical Biology, 75(2):309–325, August 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-

016-1083-4;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1083-4.pdf.

Armbruster:2017:EPC

[2665] Benjamin Armbruster and Ekkehard Beck. Elementary proof of convergence to the mean-field model for the SIR process. Journal of Mathematical Biology, 75(2):327–339, August 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-016-1086-1;
 http://link.springer.com/content/pdf/10.1007/s00285-016-1086-
1.pdf.

Olobatuyi:2017:RDM

[2666] Oluwole Olobatuyi, Gerda de Vries, and Thomas Hillen. A reaction– diffusion model for radiation-induced bystander effects. Journal of Mathematical Biology, 75(2):341–372, August 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-016-1090-5;
 http://link.springer.com/content/pdf/10.1007/s00285-016-1090-
5.pdf.

Lombardo:2017:DPM

[2667] M. C. Lombardo, R. Barresi, E. Bilotta, F. Gargano, P. Pantano, and M. Sammartino. Demyelination patterns in a mathematical model of multiple sclerosis. Journal of Mathematical Biology, 75(2):373–417, August 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1087-0;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1087-0.pdf.

Clamer:2017:HCM

[2668] Valentina Clamer, Andrea Pugliese, Davide Liessi, and Dimitri Breda. Host coexistence in a model for two host-one parasitoid interactions. Journal of Mathematical Biology, 75(2):419–441, August 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1088-z;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1088-z.pdf.

Promislow:2017:EBG

[2669] Keith Promislow and Qiliang Wu. Existence, bifurcation, and geometric evolution of quasi-bilayers in the multicomponent functionalized Cahn– Hilliard equation. Journal of Mathematical Biology, 75(2):443–489, August 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1089-y;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1089-y.pdf.

Cushing:2017:BTE

[2670] J. M. Cushing, F. Martins, A. A. Pinto, and Amy Veprauskas. A bifurcation theorem for evolutionary matrix models with multiple traits. Journal of Mathematical Biology, 75(2):491–520, August 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
016-1091-4;   http://link.springer.com/content/pdf/10.1007/
s00285-016-1091-4.pdf.

Rebuli:2017:HMC

[2671] Nicolas P. Rebuli, N. G. Bean, and J. V. Ross. Hybrid Markov chain models of S-I-R disease dynamics. Journal of Mathematical Biology, 75(3): 521–541, September 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-016-1085-2;   http://link.springer.com/content/
pdf/10.1007/s00285-016-1085-2.pdf.

Madec:2017:BIG

[2672] Sten Madec, Jérôme Casas, Guy Barles, and Christelle Suppo. Bistability induced by generalist natural enemies can reverse pest invasions. Journal of Mathematical Biology, 75(3):543–575, September 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1093-x;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1093-x.pdf.

Ball:2017:HNE

[2673] Frank Ball and Thomas House. Heterogeneous network epidemics: real-time growth, variance and extinction of infection. Journal of Mathematical Biology, 75(3):577–619, September 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-016-1092-3;
 http://link.springer.com/content/pdf/10.1007/s00285-016-1092-
3.pdf.

Farkas:2017:MWI

[2674] József Z. Farkas, Stephen A. Gourley, Rongsong Liu, and Abdul-Aziz Yakubu. Modelling Wolbachia infection in a sex-structured mosquito population carrying West Nile virus. Journal of Mathematical Biology, 75(3):621–647, September 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1096-7;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1096-
7.pdf.

Zhou:2017:DTM

[2675] Ying Zhou and William F. Fagan. A discrete-time model for population persistence in habitats with time-varying sizes. Journal of Mathematical Biology, 75(3):649–704, September 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1095-8;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1095-
8.pdf.

Iron:2017:MCS

[2676] D. Iron and J. Rumsey. A model of cell surface receptor aggregation. Journal of Mathematical Biology, 75(3):705–731, September 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1094-9;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1094-9.pdf.

Besse:2017:LTT

[2677] Apollos Besse, Thomas Lepoutre, and Samuel Bernard. Long-term treatment effects in chronic myeloid leukemia. Journal of Mathematical Biology, 75(3):733–758, September 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1098-5;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1098-
5.pdf.

Lanzarone:2017:BEA

[2678] E. Lanzarone, S. Pasquali, G. Gilioli, and E. Marchesini. A Bayesian estimation approach for the mortality in a stage-structured demographic model. Journal of Mathematical Biology, 75(3):759–779, September 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1099-4;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1099-4.pdf.

Anonymous:2017:M

[2679] Anonymous. In memoriam: [Karl Peter Hadeler (16.10.1936– 3.2.2017)]. Journal of Mathematical Biology, 75(4):781, October 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1110-0;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1110-0.pdf.

Thieme:2017:HEP

[2680] Horst R. Thieme. From homogeneous eigenvalue problems to twosex population dynamics. Journal of Mathematical Biology, 75(4): 783–804, October 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1114-9;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1114-9.pdf.

Cloez:2017:GAC

[2681] Bertrand Cloez and Coralie Fritsch. Gaussian approximations for chemostat models in finite and infinite dimensions. Journal of Mathematical Biology, 75(4):805–843, October 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1097-6;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1097-
6.pdf.

Ostergaard:2017:OSC

[2682] Jacob Østergaard, Anders Rahbek, and Susanne Ditlevsen. Oscillating systems with cointegrated phase processes. Journal of Mathematical Biology, 75(4):845–883, October 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1100-2;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1100-
2.pdf.

Avitable:2017:MCS

[2683] Daniele Avitable and Kyle C. A. Wedgwood. Macroscopic coherent structures in a stochastic neural network: from interface dynamics to coarse-grained bifurcation analysis. Journal of Mathematical Biology, 75 (4):885–928, October 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-016-1070-9;   http://link.springer.com/content/
pdf/10.1007/s00285-016-1070-9.pdf.

Drabek:2017:CTW

[2684] Pavel Drábek and Peter Takác. Convergence to travelling waves in Fisher’s population genetics model with a non-Lipschitzian reaction term. Journal of Mathematical Biology, 75(4):929–972, October 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1103-z;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1103-z.pdf.

Ji:2017:QFA

[2685] Hao Ji, Hans-Georg Müller, Nikos T. Papadopoulos, and James R. Carey. Quantifying functionals of age distributions in the wild by solving an operator equation. Journal of Mathematical Biology, 75(4):973–984, October 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1105-x;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1105-x.pdf.

Baker:2017:MMC

[2686] Michelle Baker, Bindi S. Brook, and Markus R. Owen. Mathematical modelling of cytokines, MMPs and fibronectin fragments in osteoarthritic cartilage. Journal of Mathematical Biology, 75(4):985–1024, October 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1104-y;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1104-y.pdf.

Browne:2017:GPN

[2687] Cameron Browne. Global properties of nested network model with application to multi-epitope HIV/CTL dynamics. Journal of Mathematical Biology, 75(4):1025–1046, October 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1102-0;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1102-
0.pdf.

Scheel:2017:WSC

[2688] Arnd Scheel and Angela Stevens. Wavenumber selection in coupled transport equations. Journal of Mathematical Biology, 75(5):1047– 1073, November 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1107-8;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1107-8.pdf.

Nava-Sedeno:2017:ECA

[2689] J. M. Nava-Sedeño, H. Hatzikirou, F. Peruani, and A. Deutsch. Extracting cellular automaton rules from physical Langevin equation models for single and collective cell migration. Journal of Mathematical Biology, 75(5):1075–1100, November 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1106-9;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1106-
9.pdf.

Lee:2017:MMA

[2690] Seongwon Lee, Se woong Kim, Youngmin Oh, and Hyung Ju Hwang. Mathematical modeling and its analysis for instability of the immune system induced by chemotaxis. Journal of Mathematical Biology, 75(5): 1101–1131, November 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1108-7;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1108-7.pdf.

Meunier:2017:TQR

[2691] F. Meunier, V. Couvreur, X. Draye, J. Vanderborght, and M. Javaux. Towards quantitative root hydraulic phenotyping: novel mathematical functions to calculate plant-scale hydraulic parameters from root system functional and structural traits. Journal of Mathematical Biology, 75(5): 1133–1170, November 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1111-z;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1111-z.pdf.

Emerick:2017:KMS

[2692] Brooks Emerick, Gilberto Schleiniger, and Bruce M. Boman. A kinetic model to study the regulation of β-catenin, APC, and Axin in the human colonic crypt. Journal of Mathematical Biology, 75(5):1171– 1202, November 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1112-y;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1112-y.pdf.

Ninomiya:2017:RDN

[2693] Hirokazu Ninomiya, Yoshitaro Tanaka, and Hiroko Yamamoto. Reaction, diffusion and non-local interaction. Journal of Mathematical Biology, 75(5):1203–1233, November 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1113-x;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1113-
x.pdf.

Hajnova:2017:TPB

[2694] Veronika Hajnová and Lenka Pribylová. Two-parameter bifurcations in LPA model. Journal of Mathematical Biology, 75(5):1235–1251, November 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1115-8;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1115-8.pdf.

Dessalles:2017:SAA

[2695] Renaud Dessalles, Vincent Fromion, and Philippe Robert. A stochastic analysis of autoregulation of gene expression. Journal of Mathematical Biology, 75(5):1253–1283, November 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1116-7;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1116-
7.pdf.

Sample:2017:LWS

[2696] Christine Sample and Benjamin Allen. The limits of weak selection and large population size in evolutionary game theory. Journal of Mathematical Biology, 75(5):1285–1317, November 2017. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1119-4;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1119-
4.pdf.

Hautphenne:2017:HOB

[2697] Sophie Hautphenne, Melanie Massaro, and Peter Taylor. How old is this bird? The age distribution under some phase sampling schemes. Journal of Mathematical Biology, 75(6–7):1319–1347, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1121-x;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1121-x.pdf.

Tournier:2017:ORA

[2698] Laurent Tournier, Anne Goelzer, and Vincent Fromion. Optimal resource allocation enables mathematical exploration of microbial metabolic configurations. Journal of Mathematical Biology, 75(6–7):1349–1380, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1118-5;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1118-5.pdf.

Lin:2017:SSM

[2699] Zhigui Lin and Huaiping Zhu. Spatial spreading model and dynamics of West Nile virus in birds and mosquitoes with free boundary. Journal of Mathematical Biology, 75(6–7):1381–1409, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1124-7;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1124-7.pdf.

Langlois:2017:NPD

[2700] Gabriel P. Langlois, Morgan Craig, Antony R. Humphries, Michael C. Mackey, Joseph M. Mahaffy, Jacques Bélair, Thibault Moulin, Sean R. Sinclair, and Liangliang Wang. Normal and pathological dynamics of platelets in humans. Journal of Mathematical Biology, 75(6–7):1411– 1462, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1125-6;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1125-6.pdf.

Cen:2017:BAG

[2701] Xiuli Cen, Zhilan Feng, Yiqiang Zheng, and Yulin Zhao. Bifurcation analysis and global dynamics of a mathematical model of antibiotic resistance in hospitals. Journal of Mathematical Biology, 75(6–7):1463– 1485, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1128-3;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1128-3.pdf.

Zhao:2017:DFB

[2702] Xiao Zhao, Stephan Noack, Wolfgang Wiechert, and Eric von Lieres. Dynamic flux balance analysis with nonlinear objective function. Journal of Mathematical Biology, 75(6–7):1487–1515, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1127-4;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1127-4.pdf.

Domschke:2017:SMC

[2703] Pia Domschke, Dumitru Trucu, Alf Gerisch, and Mark A. J. Chaplain. Structured models of cell migration incorporating molecular binding processes. Journal of Mathematical Biology, 75(6–7):1517–1561, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1120-y;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1120-y.pdf.

Wilkinson:2017:RBM

[2704] Robert R. Wilkinson, Frank G. Ball, and Kieran J. Sharkey. The relationships between message passing, pairwise, Kermack– McKendrick and stochastic SIR epidemic models. Journal of Mathematical Biology, 75(6–7):1563–1590, December 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1123-8;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1123-
8.pdf.

Magpantay:2017:VIH

[2705] F. M. G. Magpantay. Vaccine impact in homogeneous and age-structured models. Journal of Mathematical Biology, 75(6–7):1591–1617, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1126-5;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1126-5.pdf.

Sumner:2017:DSP

[2706] Jeremy G. Sumner, Amelia Taylor, Barbara R. Holland, and Peter D. Jarvis. Developing a statistically powerful measure for quartet tree inference using phylogenetic identities and Markov invariants. Journal of Mathematical Biology, 75(6–7):1619–1654, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1129-2;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1129-2.pdf.

Barrientos:2017:CDS

[2707] Pablo G. Barrientos, J. Ángel Rodríguez, and Alfonso Ruiz-Herrera. Chaotic dynamics in the seasonally forced SIR epidemic model. Journal of Mathematical Biology, 75(6–7):1655–1668, December 2017. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1130-9;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1130-
9.pdf.

Cardona:2017:RLN

[2708] Gabriel Cardona and Joan Carles Pons. Reconstruction of LGT networks from tri-LGT-nets. Journal of Mathematical Biology, 75(6–7): 1669–1692, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1131-8;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1131-8.pdf.

Stephenson:2017:OCV

[2709] Brittany Stephenson, Cristina Lanzas, Suzanne Lenhart, and Judy Day. Optimal control of vaccination rate in an epidemiological model of Clostridium difficile transmission. Journal of Mathematical Biology, 75(6–7):1693–1713, December 2017. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1133-6;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1133-
6.pdf.

Rozikov:2017:THD

[2710] U. A. Rozikov. Tree-hierarchy of DNA and distribution of Holliday junctions. Journal of Mathematical Biology, 75(6–7):1715–1733, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1136-3;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1136-3.pdf.

Chalub:2017:SEF

[2711] Fabio A. C. C. Chalub and Max O. Souza. On the stochastic evolution of finite populations. Journal of Mathematical Biology, 75(6–7): 1735–1774, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1135-4;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1135-4.pdf.

Hsu:2017:SSG

[2712] Sze-Bi Hsu, King-Yeung Lam, and Feng-Bin Wang. Single species growth consuming inorganic carbon with internal storage in a poorly mixed habitat. Journal of Mathematical Biology, 75(6–7):1775–1825, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1134-5;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1134-5.pdf.

Huber:2017:MTC

[2713] K. T. Huber, V. Moulton, and M. Steel. Minimum triplet covers of binary phylogenetic X-trees. Journal of Mathematical Biology, 75(6–7): 1827–1840, December 2017. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1117-6;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1117-6.pdf.

Chumley:2018:MTB

[2714] Timothy Chumley, Ozgur Aydogmus, Anastasios Matzavinos, and Alexander Roitershtein. Moran-type bounds for the fixation probability in a frequency-dependent Wright–Fisher model. Journal of Mathematical Biology, 76(1–2):1–35, January 2018. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1137-2;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1137-
2.pdf.

Wilson:2018:GAB

[2715] Dan Wilson and Bard Ermentrout. Greater accuracy and broadened applicability of phase reduction using isostable coordinates. Journal of Mathematical Biology, 76(1–2):37–66, January 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1141-6;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1141-
6.pdf.

Cuesta:2018:SSP

[2716] José A. Cuesta, Gustav W. Delius, and Richard Law. Sheldon spectrum and the plankton paradox: two sides of the same coin — a trait-based plankton size-spectrum model. Journal of Mathematical Biology, 76 (1–2):67–96, January 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1132-7;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1132-7.pdf.

Chen:2018:KMM

[2717] Kuan-Wei Chen, Kang-Ling Liao, and Chih-Wen Shih. The kinetics in mathematical models on segmentation clock genes in zebrafish. Journal of Mathematical Biology, 76(1–2):97–150, January 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1138-1;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1138-1.pdf.

Pontz:2018:EDT

[2718] Martin Pontz, Josef Hofbauer, and Reinhard Bürger. Evolutionary dynamics in the two-locus two-allele model with weak selection. Journal of Mathematical Biology, 76(1–2):151–203, January 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1140-7;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1140-7.pdf.

Motsch:2018:SRR

[2719] Sebastien Motsch and Diane Peurichard. From short-range repulsion to Hele–Shaw problem in a model of tumor growth. Journal of Mathematical Biology, 76(1–2):205–234, January 2018. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1143-4;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1143-
4.pdf.

Zheng:2018:WSD

[2720] Bo Zheng, Moxun Tang, Jianshe Yu, and Junxiong Qiu. Wolbachia spreading dynamics in mosquitoes with imperfect maternal transmission. Journal of Mathematical Biology, 76(1–2):235–263, January 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1142-5;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1142-5.pdf.

Lashari:2018:BPA

[2721] Abid Ali Lashari and Pieter Trapman. Branching process approach for epidemics in dynamic partnership network. Journal of Mathematical Biology, 76(1–2):265–294, January 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1147-0;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1147-
0.pdf.

Dawson:2018:MMS

[2722] Donald A. Dawson. Multilevel mutation–selection systems and setvalued duals. Journal of Mathematical Biology, 76(1–2):295–378, January 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1145-2;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1145-2.pdf.

Abboubakar:2018:BTO

[2723] Hamadjam Abboubakar, Jean Claude Kamgang, Leontine Nkague Nkamba, and Daniel Tieudjo. Bifurcation thresholds and optimal control in transmission dynamics of arboviral diseases. Journal of Mathematical Biology, 76(1–2):379–427, January 2018. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1146-1;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1146-
1.pdf.

Buttenschon:2018:SJD

[2724] Andreas Buttenschön, Thomas Hillen, Alf Gerisch, and Kevin J. Painter. A space-jump derivation for non-local models of cell-cell adhesion and non-local chemotaxis. Journal of Mathematical Biology, 76(1–2):429– 456, January 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1144-3;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1144-3.pdf.

Broom:2018:EFC

[2725] Mark Broom, Michal Johanis, and Jan Rychtár. The effect of fight cost structure on fighting behaviour involving simultaneous decisions and variable investment levels. Journal of Mathematical Biology, 76(1– 2):457–482, January 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1149-y;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1149-y.pdf.

Ball:2018:EVS

[2726] Frank Ball and David Sirl. Evaluation of vaccination strategies for SIR epidemics on random networks incorporating household structure. Journal of Mathematical Biology, 76(1–2):483–530, January 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-

017-1139-0;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1139-0.pdf.

Hilhorst:2018:DTF

[2727] Danielle Hilhorst, Yong-Jung Kim, Dohyun Kwon, and Thanh Nam Nguyen. Dispersal towards food: the singular limit of an Allen–Cahn equation. Journal of Mathematical Biology, 76(3):531–565, February 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1150-5;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1150-5.pdf.

Renault:2018:MTS

[2728] Vincent Renault, Michèle Thieullen, and Emmanuel Trélat. Minimal time spiking in various ChR2-controlled neuron models. Journal of Mathematical Biology, 76(3):567–608, February 2018. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1101-1;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1101-
1.pdf.

Kong:2018:SOM

[2729] Jude D. Kong, Paul Salceanu, and Hao Wang. A stoichiometric organic matter decomposition model in a chemostat culture. Journal of Mathematical Biology, 76(3):609–644, February 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1152-3;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1152-3.pdf.

Molina:2018:SFP

[2730] Chai Molina and David J. D. Earn. On selection in finite populations. Journal of Mathematical Biology, 76(3):645–678, February 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1151-4;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1151-4.pdf.

Chigansky:2018:WCO

[2731] Pavel Chigansky, Peter Jagers, and Fima C. Klebaner. What can be observed in real time PCR and when does it show? Journal of Mathematical Biology, 76(3):679–695, February 2018. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1154-1;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1154-
1.pdf.

Hening:2018:SPG

[2732] Alexandru Hening, Dang H. Nguyen, and George Yin. Stochastic population growth in spatially heterogeneous environments: the densitydependent case. Journal of Mathematical Biology, 76(3):697–754, February 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1153-2;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1153-2.pdf.

Sherborne:2018:MFM

[2733] Neil Sherborne, Joel C. Miller, Konstantin B. Blyuss, and Istvan Z. Kiss. Mean-field models for non-Markovian epidemics on networks. Journal of Mathematical Biology, 76(3):755–778, February 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1155-0;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1155-0.pdf.

Yan:2018:ASM

[2734] Ke Yan, Ho-Lun Cheng, Zhiwei Ji, Xin Zhang, and Huijuan Lu. Accelerating smooth molecular surface calculation. Journal of Mathematical Biology, 76(3):779–793, February 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1156-z;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1156-
z.pdf.

Evans:2018:RHO

[2735] Ryan M. Evans and David A. Edwards. Receptor heterogeneity in optical biosensors. Journal of Mathematical Biology, 76(4):795–816, March 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1158-x;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1158-x.pdf.

Garcia-Nieto:2018:PME

[2736] P. J. García-Nieto, E. García-Gonzalo, J. R. Alonso Fernández, and C. Díaz Muñiz. Predictive modelling of eutrophication in the Pozón de la Dolores lake (Northern Spain) by using an evolutionary support vector machines approach. Journal of Mathematical Biology, 76(4):817–840, March 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1161-2;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1161-2.pdf.

Bao:2018:FBM

[2737] Wendi Bao, Yihong Du, Zhigui Lin, and Huaiping Zhu. Free boundary models for mosquito range movement driven by climate warming. Journal of Mathematical Biology, 76(4):841–875, March 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1159-9;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1159-9.pdf.

Koslicki:2018:EPI

[2738] David Koslicki and Mark Novak. Exact probabilities for the indeterminacy of complex networks as perceived through press perturbations. Journal of Mathematical Biology, 76(4):877–909, March 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1163-0;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1163-0.pdf.

Ho:2018:BBD

[2739] Lam Si Tung Ho, Jason Xu, Forrest W. Crawford, Vladimir N. Minin, and Marc A. Suchard. Birth/birth-death processes and their computable transition probabilities with biological applications. Journal of Mathematical Biology, 76(4):911–944, March 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1160-3;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1160-3.pdf.

Mattei:2018:CDA

[2740] M. R. Mattei, L. Frunzo, B. D’Acunto, Y. Pechaud, F. Pirozzi, and G. Esposito. Continuum and discrete approach in modeling biofilm development and structure: a review. Journal of Mathematical Biology, 76 (4):945–1003, March 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1165-y;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1165-y.pdf.

Christen:2018:MSE

[2741] Alejandra Christen, M. Angélica Maulén-Yañez, Eduardo González-Olivares, and Michel Curé. Modeling a SI epidemic with stochastic transmission: hyperbolic incidence rate. Journal of Mathematical Biology, 76 (4):1005–1026, March 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1162-1;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1162-1.pdf.

Just:2018:OEM

[2742] Winfried Just, Joan Saldaña, and Ying Xin. Oscillations in epidemic models with spread of awareness. Journal of Mathematical Biology, 76 (4):1027–1057, March 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1166-x;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1166-x.pdf.

Crewe:2018:DFU

[2743] Paul Crewe, Richard Gratwick, and Alan Grafen. Defining fitness in an uncertain world. Journal of Mathematical Biology, 76(5):1059–1099, April 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1164-z;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1164-z.pdf.

Gavryushkin:2018:CDT

[2744] Alex Gavryushkin, Chris Whidden, and Frederick A. Matsen IV. The combinatorics of discrete time-trees: theory and open problems. Journal of Mathematical Biology, 76(5):1101–1121, April 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1167-9;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1167-9.pdf.

Manda:2018:MCW

[2745] Edna Chilenje Manda and Faraimunashe Chirove. Modelling coupled within host and population dynamics of R_{5} and X_{4} HIV infection. Journal of Mathematical Biology, 76(5):1123–1158, April 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1170-1;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1170-1.pdf.

Zhang:2018:MMA

[2746] Jimin Zhang, Junping Shi, and Xiaoyuan Chang. A mathematical model of algae growth in a pelagic-benthic coupled shallow aquatic ecosystem. Journal of Mathematical Biology, 76(5):1159–1193, April 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1168-8;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1168-8.pdf.

Clote:2018:RFK

[2747] Peter Clote and Amir H. Bayegan. RNA folding kinetics using Monte Carlo and Gillespie algorithms. Journal of Mathematical Biology, 76 (5):1195–1227, April 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1169-7;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1169-7.pdf.

Francis:2018:BPN

[2748] Andrew Francis, Katharina T. Huber, Vincent Moulton, and Taoyang Wu. Bounds for phylogenetic network space metrics. Journal of Mathematical Biology, 76(5):1229–1248, April 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1171-0;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1171-
0.pdf.

Song:2018:GHB

[2749] Pengfei Song and Yanni Xiao. Global Hopf bifurcation of a delayed equation describing the lag effect of media impact on the spread of infectious disease. Journal of Mathematical Biology, 76(5):1249–1267, April 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1173-y;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1173-y.pdf.

Bliman:2018:ESI

[2750] Pierre-Alexandre Bliman, M. Soledad Aronna, Flávio C. Coelho, and Moacyr A. H. B. da Silva. Ensuring successful introduction of Wolbachia in natural populations of Aedes aegypti by means of feedback control. Journal of Mathematical Biology, 76(5):1269–1300, April 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-

017-1174-x;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1174-x.pdf.

Fan:2018:GSM

[2751] Dejun Fan, Pengmiao Hao, and Dongyan Sun. Global stability of multi-group viral models with general incidence functions. Journal of Mathematical Biology, 76(5):1301–1326, April 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1178-6;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1178-
6.pdf.

Cael:2018:LCP

[2752] B. B. Cael and Courtenay Strong. A Laplacian characterization of phytoplankton shape. Journal of Mathematical Biology, 76(6):1327–1338, May 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1176-8;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1176-8.pdf.

Huang:2018:GSE

[2753] Yu-Jhe Huang, Jonq Juang, Yu-Hao Liang, and Hsin-Yu Wang. Global stability for epidemic models on multiplex networks. Journal of Mathematical Biology, 76(6):1339–1356, May 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1179-5;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1179-
5.pdf.

Pellacci:2018:BDS

[2754] Benedetta Pellacci and Gianmaria Verzini. Best dispersal strategies in spatially heterogeneous environments: optimization of the principal eigenvalue for indefinite fractional Neumann problems. Journal of Mathematical Biology, 76(6):1357–1386, May 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1180-z;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1180-
z.pdf.

Yang:2018:EIC

[2755] Zhichun Yang, Chuangxia Huang, and Xingfu Zou. Effect of impulsive controls in a model system for age-structured population over a patchy environment. Journal of Mathematical Biology, 76(6):1387–1419, May 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1172-z;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1172-z.pdf.

Coron:2018:SMS

[2756] Camille Coron, Manon Costa, Hélène Leman, and Charline Smadi. A stochastic model for speciation by mating preferences. Journal of Mathematical Biology, 76(6):1421–1463, May 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1175-9;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1175-
9.pdf.

Hadjichrysanthou:2018:MKN

[2757] Christoforos Hadjichrysanthou, Mark Broom, and Jan Rychtár. Models of kleptoparasitism on networks: the effect of population structure on food stealing behaviour. Journal of Mathematical Biology, 76(6): 1465–1488, May 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1177-7;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1177-7.pdf.

Nadin:2018:HBF

[2758] Grégoire Nadin, Martin Strugarek, and Nicolas Vauchelet. Hindrances to bistable front propagation: application to Wolbachia invasion. Journal of Mathematical Biology, 76(6):1489–1533, May 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1181-y;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1181-y.pdf.

Johnston:2018:CEE

[2759] Matthew D. Johnston, David F. Anderson, Gheorghe Craciun, and Robert Brijder. Conditions for extinction events in chemical reaction networks with discrete state spaces. Journal of Mathematical Biology, 76 (6):1535–1558, May 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1182-x;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1182-x.pdf.

Damiano:2018:TAT

[2760] David B. Damiano and Melissa R. McGuirl. A topological analysis of targeted In-111 uptake in SPECT images of murine tumors. Journal of Mathematical Biology, 76(6):1559–1587, May 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1184-8;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1184-8.pdf.

Basener:2018:FTN

[2761] William F. Basener and John C. Sanford. The fundamental theorem of natural selection with mutations. Journal of Mathematical Biology, 76 (7):1589–1622, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1190-x;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1190-x.pdf. See reply [3619].

Hellmuth:2018:IPT

[2762] Marc Hellmuth, Maribel Hernandez-Rosales, Yangjing Long, and Peter F. Stadler. Inferring phylogenetic trees from the knowledge of rare evolutionary events. Journal of Mathematical Biology, 76(7):1623–1653, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1194-6;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1194-6.pdf.

Sovrano:2018:NAC

[2763] Elisa Sovrano. A negative answer to a conjecture arising in the study of selection–migration models in population genetics. Journal of Mathematical Biology, 76(7):1655–1672, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1185-7;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1185-
7.pdf.

Lambert:2018:BIA

[2764] Ben Lambert, Adam L. MacLean, Alexander G. Fletcher, Alexander N. Combes, Melissa H. Little, and Helen M. Byrne. Bayesian inference of agent-based models: a tool for studying kidney branching morphogenesis. Journal of Mathematical Biology, 76(7):1673–1697, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-

018-1208-z;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1208-z.pdf.

Wang:2018:AMM

[2765] Qixuan Wang and Hans G. Othmer. Analysis of a model microswimmer with applications to blebbing cells and mini-robots. Journal of Mathematical Biology, 76(7):1699–1763, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1225-y;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1225-
y.pdf.

Girel:2018:ESP

[2766] Simon Girel and Fabien Crauste. Existence and stability of periodic solutions of an impulsive differential equation and application to CD8 T-cell differentiation. Journal of Mathematical Biology, 76(7):1765–1795, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1220-3;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1220-3.pdf.

Emerick:2018:MSM

[2767] Brooks Emerick, Gilberto Schleiniger, and Bruce M. Boman. Multiscale modeling of APC and β-catenin regulation in the human colonic crypt. Journal of Mathematical Biology, 76(7):1797–1830, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1204-8;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1204-8.pdf.

Ren:2018:RDW

[2768] Xinzhi Ren, Yanni Tian, Lili Liu, and Xianning Liu. A reaction– diffusion within-host HIV model with cell-to-cell transmission. Journal of Mathematical Biology, 76(7):1831–1872, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1202-x;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1202-
x.pdf.

Seo:2018:SDG

[2769] Gunog Seo and Gail S. K. Wolkowicz. Sensitivity of the dynamics of the general Rosenzweig–MacArthur model to the mathematical form of the functional response: a bifurcation theory approach. Journal of Mathematical Biology, 76(7):1873–1906, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1201-y;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1201-y.pdf.

Campo-Duarte:2018:OCA

[2770] Doris E. Campo-Duarte, Olga Vasilieva, Daiver Cardona-Salgado, and Mikhail Svinin. Optimal control approach for establishing wMelPop Wolbachia infection among wild Aedes aegypti populations. Journal of Mathematical Biology, 76(7):1907–1950, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1213-2;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1213-2.pdf.

Garay:2018:ERE

[2771] József Garay, Ross Cressman, Tamás F. Móri, and Tamás Varga. The ESS and replicator equation in matrix games under time constraints. Journal of Mathematical Biology, 76(7):1951–1973, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1207-0;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1207-0.pdf.

Li:2018:MIS

[2772] Bingtuan Li. Multiple invasion speeds in a two-species integro-difference competition model. Journal of Mathematical Biology, 76(7):1975–2009, June 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1200-z;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1200-z.pdf.

Wang:2018:CBM

[2773] Xiunan Wang and Xiao-Qiang Zhao. A climate-based malaria model with the use of bed nets. Journal of Mathematical Biology, 77(1):1–25, July 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1183-9;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1183-9.pdf.

Yurk:2018:HAI

[2774] Brian P. Yurk. Homogenization analysis of invasion dynamics in heterogeneous landscapes with differential bias and motility. Journal of Mathematical Biology, 77(1):27–54, July 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1186-6;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1186-
6.pdf.

Dowdall:2018:IPP

[2775] James Dowdall, Victor LeBlanc, and Frithjof Lutscher. Invasion pinning in a periodically fragmented habitat. Journal of Mathematical Biology, 77(1):55–78, July 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1188-4;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1188-4.pdf.

Patel:2018:RPE

[2776] Swati Patel and Sebastian J. Schreiber. Robust permanence for ecological equations with internal and external feedbacks. Journal of Mathematical Biology, 77(1):79–105, July 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1187-5;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1187-
5.pdf.

Bichara:2018:MPM

[2777] Derdei Bichara and Abderrahman Iggidr. Multi-patch and multi-group epidemic models: a new framework. Journal of Mathematical Biology, 77(1):107–134, July 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1191-9;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1191-9.pdf.

Hening:2018:SLV

[2778] Alexandru Hening and Dang H. Nguyen. Stochastic Lotka–Volterra food chains. Journal of Mathematical Biology, 77(1):135–163, July 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1192-8;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1192-8.pdf.

Dalwadi:2018:MTA

[2779] Mohit P. Dalwadi, John R. King, and Nigel P. Minton. Multitimescale analysis of a metabolic network in synthetic biology: a kinetic model for 3-hydroxypropionic acid production via betaalanine. Journal of Mathematical Biology, 77(1):165–199, July 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1189-3;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1189-3.pdf.

Bai:2018:RDM

[2780] Zhenguo Bai, Rui Peng, and Xiao-Qiang Zhao. A reaction–diffusion malaria model with seasonality and incubation period. Journal of Mathematical Biology, 77(1):201–228, July 2018. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1193-7;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1193-
7.pdf.

Lessard:2018:FDG

[2781] Sabin Lessard and Cíntia Dalila Soares. Frequency-dependent growth in class-structured populations: continuous dynamics in the limit of weak selection. Journal of Mathematical Biology, 77(1):229–259, July 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1195-5;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1195-5.pdf.

Veerman:2018:TDP

[2782] Frits Veerman, Carsten Marr, and Nikola Popović. Time-dependent propagators for stochastic models of gene expression: an analytical method. Journal of Mathematical Biology, 77(2):261–312, August 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1196-4;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1196-4.pdf.

Prohaska:2018:EGC

[2783] Sonja J. Prohaska, Sarah J. Berkemer, Fabian Gärtner, Thomas Gatter, Nancy Retzlaff, Christian Höner zu Siederdissen, and Peter F. Stadler. Expansion of gene clusters, circular orders, and the shortest Hamiltonian path problem. Journal of Mathematical Biology, 77(2):313–341, August 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
017-1197-3;   http://link.springer.com/content/pdf/10.1007/
s00285-017-1197-3.pdf.

Al-Darabsah:2018:PDT

[2784] Isam Al-Darabsah and Yuan Yuan. A periodic disease transmission model with asymptomatic carriage and latency periods. Journal of Mathematical Biology, 77(2):343–376, August 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1199-1;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1199-
1.pdf.

Reis:2018:GSC

[2785] Matthias Reis, Justus A. Kromer, and Edda Klipp. General solution of the chemical master equation and modality of marginal distributions for hierarchic first-order reaction networks. Journal of Mathematical Biology, 77(2):377–419, August 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1205-2;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1205-2.pdf.

Rohan:2018:MCE

[2786] Eduard Rohan, Vladimír Lukes, and Alena Jonásová. Modeling of the contrast-enhanced perfusion test in liver based on the multicompartment flow in porous media. Journal of Mathematical Biology, 77 (2):421–454, August 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1209-y;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1209-y.pdf.

Brightwell:2018:ETS

[2787] Graham Brightwell, Thomas House, and Malwina Luczak. Extinction times in the subcritical stochastic SIS logistic epidemic. Journal of Mathematical Biology, 77(2):455–493, August 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1210-5;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1210-
5.pdf.

Peckham:2018:CTE

[2788] Scott D. Peckham, Edward C. Waymire, and Patrick De Leenheer. Critical thresholds for eventual extinction in randomly disturbed population growth models. Journal of Mathematical Biology, 77(2):495–525, August 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1217-y;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1217-y.pdf.

Bordewich:2018:ICD

[2789] Magnus Bordewich, Ina Maria Deutschmann, Mareike Fischer, Elisa Kasbohm, Charles Semple, and Mike Steel. On the information content of discrete phylogenetic characters. Journal of Mathematical Biology, 77(3): 527–544, September 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-017-1198-2;   http://link.springer.com/content/
pdf/10.1007/s00285-017-1198-2.pdf.

Clancy:2018:PTS

[2790] Damian Clancy. Persistence time of SIS infections in heterogeneous populations and networks. Journal of Mathematical Biology, 77(3): 545–570, September 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1222-1;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1222-1.pdf.

Bordewich:2018:RNN

[2791] Magnus Bordewich, Katharina T. Huber, Vincent Moulton, and Charles Semple. Recovering normal networks from shortest inter-taxa distance information. Journal of Mathematical Biology, 77(3):571–594, September 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1218-x;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1218-x.pdf.

Wu:2018:GSS

[2792] Hao Wu, Marco Avila Ponce de León, and Hans G. Othmer. Getting in shape and swimming: the role of cortical forces and membrane heterogeneity in eukaryotic cells. Journal of Mathematical Biology, 77(3): 595–626, September 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1223-0;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1223-0.pdf.

Kleshnina:2018:EGU

[2793] Maria Kleshnina, Jerzy A. Filar, Vladimir Ejov, and Jody C. McKerral. Evolutionary games under incompetence. Journal of Mathematical Biology, 77(3):627–646, September 2018. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1221-2;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1221-
2.pdf.

Mushanyu:2018:API

[2794] J. Mushanyu, F. Nyabadza, G. Muchatibaya, P. Mafuta, and G. Nhawu. Assessing the potential impact of limited public health resources on the spread and control of typhoid. Journal of Mathematical Biology, 77(3): 647–670, September 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1219-9;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1219-9.pdf.

Pham:2018:NST

[2795] Kara Pham, Emma Turian, Kai Liu, Shuwang Li, and John Lowengrub. Nonlinear studies of tumor morphological stability using a two-fluid flow model. Journal of Mathematical Biology, 77(3):671–709, September 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1212-3;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1212-3.pdf.

Aw:2018:BMG

[2796] Alan J. Aw and Noah A. Rosenberg. Bounding measures of genetic similarity and diversity using majorization. Journal of Mathematical Biology, 77(3):711–737, September 2018. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1226-x;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1226-
x.pdf.

Eigentler:2018:AMB

[2797] Lukas Eigentler and Jonathan A. Sherratt. Analysis of a model for banded vegetation patterns in semi-arid environments with nonlocal dispersal. Journal of Mathematical Biology, 77(3):739–763, September 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-

018-1233-y;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1233-y.pdf.

Barbour:2018:LAM

[2798] A. D. Barbour, R. McVinish, and P. K. Pollett. Local approximation of a metapopulation’s equilibrium. Journal of Mathematical Biology, 77(3): 765–793, September 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1231-0;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1231-0.pdf.

Baake:2018:PVD

[2799] Ellen Baake, Fernando Cordero, and Sebastian Hummel. A probabilistic view on the deterministic mutation–selection equation: dynamics, equilibria, and ancestry via individual lines of descent. Journal of Mathematical Biology, 77(3):795–820, September 2018. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1228-8;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1228-
8.pdf.

Michalaki:2018:AAT

[2800] Lida I. Michalaki and Dimitris A. Goussis. Asymptotic analysis of a TMDD model: when a reaction contributes to the destruction of its product. Journal of Mathematical Biology, 77(3):821–855, September 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1234-x;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1234-x.pdf.

Eikenberry:2018:MMC

[2801] Steffen E. Eikenberry and Abba B. Gumel. Mathematical modeling of climate change and malaria transmission dynamics: a historical review. Journal of Mathematical Biology, 77(4):857–933, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1229-7;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1229-7.pdf.

McClelland:2018:EEL

[2802] Jason McClelland and David Koslicki. EMDUniFrac: exact linear time computation of the UniFrac metric and identification of differentially abundant organisms. Journal of Mathematical Biology, 77(4):

935–949, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1235-9;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1235-9.pdf.

Lienkaemper:2018:GPF

[2803] Caitlin Lienkaemper, Lisa Lamberti, James Drain, Niko Beerenwinkel, and Alex Gavryushkin. The geometry of partial fitness orders and an efficient method for detecting genetic interactions. Journal of Mathematical Biology, 77(4):951–970, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1237-7;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1237-7.pdf.

Bovier:2018:RRA

[2804] Anton Bovier, Loren Coquille, and Rebecca Neukirch. The recovery of a recessive allele in a Mendelian diploid model. Journal of Mathematical Biology, 77(4):971–1033, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1240-z;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1240-
z.pdf.

Almocera:2018:MMW

[2805] Alexis Erich S. Almocera, Van Kinh Nguyen, and Esteban A. Hernandez-Vargas. Multiscale model within-host and between-host for viral infectious diseases. Journal of Mathematical Biology, 77(4):1035–1057, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1241-y;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1241-y.pdf.

Garnaev:2018:BAR

[2806] Andrey Garnaev and Wade Trappe. A bargaining approach for resolving the tradeoff between beneficial and harmful drug responses. Journal of Mathematical Biology, 77(4):1059–1072, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1242-x;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1242-x.pdf.

Colin:2018:EES

[2807] Thierry Colin, Guillaume Dechristé, Jérôme Fehrenbach, Ludivine Guillaume, Valérie Lobjois, and Clair Poignard. Experimental estimation of stored stress within spherical microtissues. Journal of Mathematical Biology, 77(4):1073–1092, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1243-9;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1243-
9.pdf.

Wang:2018:SMI

[2808] Lulu Wang, Zhen Jin, and Hao Wang. A switching model for the impact of toxins on the spread of infectious diseases. Journal of Mathematical Biology, 77(4):1093–1115, October 2018. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1245-7;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1245-
7.pdf.

Li:2018:SDR

[2809] Jinxian Li, Jing Wang, and Zhen Jin. SIR dynamics in random networks with communities. Journal of Mathematical Biology, 77(4):1117– 1151, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1247-5;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1247-5.pdf.

Czuppon:2018:LNA

[2810] Peter Czuppon and Peter Pfaffelhuber. Limits of noise for autoregulated gene expression. Journal of Mathematical Biology, 77(4):1153–1191, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1248-4;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1248-4.pdf.

Parsons:2018:IPH

[2811] Todd L. Parsons. Invasion probabilities, hitting times, and some fluctuation theory for the stochastic logistic process. Journal of Mathematical Biology, 77(4):1193–1231, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1250-x;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1250-
x.pdf.

Czuppon:2018:FPP

[2812] Peter Czuppon and Arne Traulsen. Fixation probabilities in populations under demographic fluctuations. Journal of Mathematical Biology, 77(4): 1233–1277, October 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1251-9;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1251-9.pdf.

Shchur:2018:NST

[2813] Vladimir Shchur and Rasmus Nielsen. On the number of siblings and p-th cousins in a large population sample. Journal of Mathematical Biology, 77(5):1279–1298, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1252-8;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1252-
8.pdf.

Inaba:2018:ASE

[2814] Hisashi Inaba, Ryohei Saito, and Nicolas Bacaër. An age-structured epidemic model for the demographic transition. Journal of Mathematical Biology, 77(5):1299–1339, November 2018. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1253-7;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1253-
7.pdf.

Berrios:2018:SAB

[2815] Soledad Berríos, Julio López Fenner, and Aude Maignan. Simulating aggregates of bivalents in 2n = 40 mouse meiotic spermatocytes through inhomogeneous site percolation processes. Journal of Mathematical Biology, 77(5):1341–1362, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1254-6;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1254-
6.pdf.

Shimadzu:2018:SRR

[2816] Hideyasu Shimadzu. On species richness and rarefaction: size- and coverage-based techniques quantify different characteristics of richness change in biodiversity. Journal of Mathematical Biology, 77(5):1363– 1381, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1255-5;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1255-5.pdf.

Contento:2018:EIC

[2817] Lorenzo Contento, Danielle Hilhorst, and Masayasu Mimura. Ecological invasion in competition–diffusion systems when the exotic species is either very strong or very weak. Journal of Mathematical Biology, 77(5): 1383–1405, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1256-4;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1256-4.pdf.

Ma:2018:CBP

[2818] Jie Ma and James P. Keener. The computation of biomarkers in pharmacokinetics with the aid of singular perturbation methods. Journal of Mathematical Biology, 77(5):1407–1430, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1257-3;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1257-3.pdf.

Kopp:2018:PLP

[2819] Michael Kopp, Elma Nassar, and Etienne Pardoux. Phenotypic lag and population extinction in the moving-optimum model: insights from a small-jumps limit. Journal of Mathematical Biology, 77(5):1431– 1458, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1258-2;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1258-2.pdf.

Geiss:2018:RGT

[2820] Manuela Geiß, John Anders, Peter F. Stadler, Nicolas Wieseke, and Marc Hellmuth. Reconstructing gene trees from Fitch’s xenology relation. Journal of Mathematical Biology, 77(5):1459–1491, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1260-8;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1260-8.pdf.

Oba:2018:WDI

[2821] Takuji Oba and Jun Kigami. Why does invasion imply substitution? Beyond the paradigm of invasion fitness. Journal of Mathematical Biology, 77(5):1493–1532, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1261-7;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1261-
7.pdf.

Moore:2018:OCT

[2822] Helen Moore, Lewis Strauss, and Urszula Ledzewicz. Optimization of combination therapy for chronic myeloid leukemia with dosing constraints. Journal of Mathematical Biology, 77(5):1533–1561, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1262-6;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1262-6.pdf.

Jaramillo:2018:HCS

[2823] J. M. Jaramillo, Junling Ma, P. van den Driessche, and Sanling Yuan. Host contact structure is important for the recurrence of Influenza A. Journal of Mathematical Biology, 77(5):1563–1588, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1263-5;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1263-5.pdf.

Zheng:2018:MMA

[2824] Xiaoming Zheng and Mohye Sweidan. A mathematical model of angiogenesis and tumor growth: analysis and application in anti-angiogenesis therapy. Journal of Mathematical Biology, 77(5):1589–1622, November 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1264-4;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1264-4.pdf.

Diekmann:2018:KPH

[2825] Odo Diekmann, Klaus Dietz, Thomas Hillen, and Horst Thieme. Karl-Peter Hadeler: His legacy in mathematical biology. Journal of Mathematical Biology, 77(6–7):1623–1627, December 2018. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1259-1;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1259-
1.pdf.

Magal:2018:PIP

[2826] Pierre Magal and Glenn Webb. The parameter identification problem for SIR epidemic models: identifying unreported cases. Journal of Mathematical Biology, 77(6–7):1629–1648, December 2018. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-017-1203-9;
 http://link.springer.com/content/pdf/10.1007/s00285-017-1203-
9.pdf.

Lewis:2018:IEP

[2827] Mark A. Lewis, Nathan G. Marculis, and Zhongwei Shen. Integrodifference equations in the presence of climate change: persistence criterion, travelling waves and inside dynamics. Journal of Mathematical Biology, 77(6–7):1649–1687, December 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1206-1;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1206-
1.pdf.

Cushing:2018:PMM

[2828] J. M. Cushing and Shandelle M. Henson. Periodic matrix models for seasonal dynamics of structured populations with application to a seabird population. Journal of Mathematical Biology, 77(6–7):1689– 1720, December 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1211-4;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1211-4.pdf.

Pillay:2018:IEP

[2829] Samara Pillay, Helen M. Byrne, and Philip K. Maini. The impact of exclusion processes on angiogenesis models. Journal of Mathematical Biology, 77(6–7):1721–1759, December 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1214-1;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1214-
1.pdf.

Alfaro:2018:NSV

[2830] Matthieu Alfaro, Hirofumi Izuhara, and Masayasu Mimura. On a nonlocal system for vegetation in drylands. Journal of Mathematical Biology, 77(6–7):1761–1793, December 2018. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1215-0;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1215-
0.pdf.

Poghotanyan:2018:CMP

[2831] Gayane Poghotanyan, Zhilan Feng, John W. Glasser, and Andrew N. Hill. Constrained minimization problems for the reproduction number in meta-population models. Journal of Mathematical Biology, 77(6–7): 1795–1831, December 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1216-z;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1216-z.pdf.

Browne:2018:DVI

[2832] Cameron J. Browne and Hal L. Smith. Dynamics of virus and immune response in multi-epitope network. Journal of Mathematical Biology, 77(6–7):1833–1870, December 2018. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1224-z;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1224-
z.pdf.

Zhao:2018:TWS

[2833] Lin Zhao, Zhi-Cheng Wang, and Shigui Ruan. Traveling wave solutions in a two-group SIR epidemic model with constant recruitment. Journal of Mathematical Biology, 77(6–7):1871–1915, December 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1227-9;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1227-9.pdf.

Saad-Roy:2018:EST

[2834] C. M. Saad-Roy, Junling Ma, and P. van den Driessche. The effect of sexual transmission on Zika virus dynamics. Journal of Mathematical Biology, 77(6–7):1917–1941, December 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1230-1;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1230-
1.pdf.

Geritz:2018:ACB

[2835] S. Geritz, M. Gyllenberg, and J. Toivonen. Adaptive correlations between seed size and germination time. Journal of Mathematical Biology, 77(6–7):1943–1968, December 2018. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1232-z;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1232-
z.pdf.

Olobatuyi:2018:EGC

[2836] Oluwole Olobatuyi, Gerda de Vries, and Thomas Hillen. Effects of G2-checkpoint dynamics on low-dose hyper-radiosensitivity. Journal of Mathematical Biology, 77(6–7):1969–1997, December 2018. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1236-8;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1236-
8.pdf.

Rutter:2018:EIH

[2837] E. M. Rutter, H. T. Banks, and K. B. Flores. Estimating intratumoral heterogeneity from spatiotemporal data. Journal of Mathematical Biology, 77(6–7):1999–2022, December 2018. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1238-6;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1238-
6.pdf.

Diekmann:2018:WBD

[2838] O. Diekmann, W. F. de Graaf, M. E. E. Kretzschmar, and P. F. M. Teunis. Waning and boosting: on the dynamics of immune status. Journal of Mathematical Biology, 77(6–7):2023–2048, December 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1239-5;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1239-5.pdf.

MacDonald:2018:IBH

[2839] Jane S. MacDonald and Frithjof Lutscher. Individual behavior at habitat edges may help populations persist in moving habitats. Journal of Mathematical Biology, 77(6–7):2049–2077, December 2018. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1244-8;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1244-8.pdf.

Pugliese:2018:SPM

[2840] Andrea Pugliese and Fabio Milner. A structured population model with diffusion in structure space. Journal of Mathematical Biology, 77(6–7): 2079–2102, December 2018. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1246-6;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1246-6.pdf.

Farrell:2018:DFI

[2841] Alex P. Farrell, James P. Collins, Amy L. Greer, and Horst R. Thieme. Do fatal infectious diseases eradicate host species? Journal of Mathematical Biology, 77(6–7):2103–2164, December 2018. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1249-3;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1249-
3.pdf.

Bergeron:2019:GRP

[2842] François Bergeron and Christophe Reutenauer. Golden ratio and phyllotaxis, a clear mathematical link. Journal of Mathematical Biology, 78 (1–2):1–19, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1265-3;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1265-3.pdf.

Dione:2019:SAP

[2843] Ibrahima Dione, Nicolas Doyon, and Jean Deteix. Sensitivity analysis of the Poisson Nernst–Planck equations: a finite element approximation for the sensitive analysis of an electrodiffusion model. Journal of Mathematical Biology, 78(1–2):21–56, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1266-2;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1266-2.pdf.

Helal:2019:SAS

[2844] Mohammed Helal, Angélique Igel-Egalon, Abdelkader Lakmeche, Pauline Mazzocco, Angélique Perrillat-Mercerot, Laurent Pujo-Menjouet, Human Rezaei, and Léon M. Tine. Stability analysis of a steady state of a model describing Alzheimer’s disease and interactions with prion proteins. Journal of Mathematical Biology, 78(1–2):57–81, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1267-1;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1267-1.pdf.

Li:2019:HWD

[2845] Yao Li, Logan Chariker, and Lai-Sang Young. How well do reduced models capture the dynamics in models of interacting neurons? Journal of Mathematical Biology, 78(1–2):83–115, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1268-0;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1268-0.pdf.

Hoppe:2019:SNC

[2846] Anica Hoppe, Sonja Türpitz, and Mike Steel. Species notions that combine phylogenetic trees and phenotypic partitions. Journal of Mathematical Biology, 78(1–2):117–134, January 2019. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1269-z;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1269-
z.pdf.

Wang:2019:EWS

[2847] Zhuoqun Wang and Rick Durrett. Extrapolating weak selection in evolutionary games. Journal of Mathematical Biology, 78(1–2):135–154, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1270-6;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1270-6.pdf.

Disanto:2019:ECC

[2848] Filippo Disanto and Noah A. Rosenberg. Enumeration of compact coalescent histories for matching gene trees and species trees. Journal of Mathematical Biology, 78(1–2):155–188, January 2019. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1271-5;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1271-
5.pdf.

Grusea:2019:CTT

[2849] Simona Grusea, Willy Rodríguez, Didier Pinchon, Lounès Chikhi, Simon Boitard, and Olivier Mazet. Coalescence times for three genes provide sufficient information to distinguish population structure from population size changes. Journal of Mathematical Biology, 78(1–2): 189–224, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1272-4;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1272-4.pdf.

Iggidr:2019:SES

[2850] A. Iggidr and M. O. Souza. State estimators for some epidemiological systems. Journal of Mathematical Biology, 78(1–2):225–256, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1273-3;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1273-3.pdf.

Miller:2019:IWP

[2851] Judith R. Miller. Invasion waves and pinning in the Kirkpatrick– Barton model of evolutionary range dynamics. Journal of Mathematical Biology, 78(1–2):257–292, January 2019. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1274-2;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1274-
2.pdf.

Hening:2019:AHP

[2852] Alexandru Hening, Dang H. Nguyen, Sergiu C. Ungureanu, and Tak Kwong Wong. Asymptotic harvesting of populations in random environments. Journal of Mathematical Biology, 78(1–2):293–329, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1275-1;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1275-1.pdf.

Duong:2019:DNI

[2853] Manh Hong Duong, Hoang Minh Tran, and The Anh Han. On the distribution of the number of internal equilibria in random evolutionary games. Journal of Mathematical Biology, 78(1–2):331–371, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1276-0;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1276-0.pdf.

Canizo:2019:EEG

[2854] José A. Cañizo, José A. Carrillo, and Manuel Pájaro. Exponential equilibration of genetic circuits using entropy methods. Journal of Mathematical Biology, 78(1–2):373–411, January 2019. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1277-z;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1277-
z.pdf.

Kruff:2019:RMS

[2855] Niclas Kruff, Christian Lax, Volkmar Liebscher, and Sebastian Walcher. The Rosenzweig–MacArthur system via reduction of an individual based model. Journal of Mathematical Biology, 78(1–2):413–439, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1278-y;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1278-y.pdf.

Ge:2019:IAP

[2856] Li Ge, Jiaguo Liu, Yusen Zhang, and Matthias Dehmer. Identifying anticancer peptides by using a generalized chaos game representation. Journal of Mathematical Biology, 78(1–2):441–463, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1279-x;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1279-x.pdf.

Lemarre:2019:GMM

[2857] Paul Lemarre, Laurent Pujo-Menjouet, and Suzanne S. Sindi. Generalizing a mathematical model of prion aggregation allows strain coexistence and co-stability by including a novel misfolded species. Journal of Mathematical Biology, 78(1–2):465–495, January 2019. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1280-4;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1280-
4.pdf.

Camacho:2019:BMT

[2858] Ariel Camacho and Silvia Jerez. Bone metastasis treatment modeling via optimal control. Journal of Mathematical Biology, 78(1–2):497–526, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-

018-1281-3;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1281-3.pdf.

Kelk:2019:FMP

[2859] Steven Kelk, Fabio Pardi, Celine Scornavacca, and Leo van Iersel. Finding a most parsimonious or likely tree in a network with respect to an alignment. Journal of Mathematical Biology, 78(1–2):527–547, January 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1282-2;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1282-2.pdf.

Coron:2019:IDE

[2860] Camille Coron, Sylvie Méléard, and Denis Villemonais. Impact of demography on extinction/fixation events. Journal of Mathematical Biology, 78 (3):549–577, February 2019. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1283-1;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1283-1.pdf.

Lange:2019:FCC

[2861] Alexander Lange, Robert Schwieger, Julia Plöntzke, Stefan Schäfer, and Susanna Röblitz. Follicular competition in cows: the selection of dominant follicles as a synergistic effect. Journal of Mathematical Biology, 78 (3):579–606, February 2019. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1284-0;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1284-0.pdf.

Roberts:2019:SIM

[2862] M. G. Roberts, R. I. Hickson, J. M. McCaw, and L. Talarmain. A simple influenza model with complicated dynamics. Journal of Mathematical Biology, 78(3):607–624, February 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1285-z;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1285-
z.pdf.

Oliveri:2019:RPC

[2863] Hadrien Oliveri, Jan Traas, Christophe Godin, and Olivier Ali. Regulation of plant cell wall stiffness by mechanical stress: a mesoscale physical model. Journal of Mathematical Biology, 78(3):625–653, February 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1286-y;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1286-y.pdf.

Manhart:2019:CPW

[2864] Angelika Manhart. Counter-propagating wave patterns in a swarm model with memory. Journal of Mathematical Biology, 78(3):655–682, February 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1287-x;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1287-x.pdf.

Wang:2019:GDM

[2865] Yuanshi Wang, Hong Wu, and Donald L. DeAngelis. Global dynamics of a mutualism-competition model with one resource and multiple consumers. Journal of Mathematical Biology, 78(3):683–710, February 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1288-9;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1288-9.pdf.

Wang:2019:TPP

[2866] Xiaoying Wang and Frithjof Lutscher. Turing patterns in a predator– prey model with seasonality. Journal of Mathematical Biology, 78(3): 711–737, February 2019. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1289-8;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1289-8.pdf.

Shan:2019:DIE

[2867] Chunhua Shan and Qihua Huang. Direct and indirect effects of toxins on competition dynamics of species in an aquatic environment. Journal of Mathematical Biology, 78(3):739–766, February 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1290-2;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1290-2.pdf.

Luo:2019:MPB

[2868] Qingyang Luo. A minimally parametrized branching process explaining plateau phase of qPCR amplification. Journal of Mathematical Biology, 78(3):767–776, February 2019. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1291-1;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1291-
1.pdf.

Almet:2019:PBB

[2869] Axel A. Almet, Helen M. Byrne, Philip K. Maini, and Derek E. Moulton. Post-buckling behaviour of a growing elastic rod. Journal of Mathematical Biology, 78(3):777–814, February 2019. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1292-0;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1292-
0.pdf.

Bennett:2019:LSP

[2870] Jamie J. R. Bennett and Jonathan A. Sherratt. Large scale patterns in mussel beds: stripes or spots? Journal of Mathematical Biology, 78 (3):815–835, February 2019. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1293-z;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1293-z.pdf.

Semenov:2019:GQM

[2871] Yuri S. Semenov and Artem S. Novozhilov. Generalized quasispecies model on finite metric spaces: isometry groups and spectral properties of evolutionary matrices. Journal of Mathematical Biology, 78(3): 837–878, February 2019. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1294-y;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1294-y.pdf.

Ren:2019:DMB

[2872] Yue Ren, Johannes W. R. Martini, and Jacinta Torres. Decoupled molecules with binding polynomials of bidegree (n, 2). Journal of Mathematical Biology, 78(4):879–898, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1295-x;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1295-
x.pdf.

Pons:2019:TBN

[2873] Joan Carles Pons, Charles Semple, and Mike Steel. Tree-based networks: characterisations, metrics, and support trees. Journal of Mathematical Biology, 78(4):899–918, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1296-9;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1296-
9.pdf.

Gray:2019:CTT

[2874] Catheryn W. Gray and Adelle C. F. Coster. Crosstalk in transition: the translocation of Akt. Journal of Mathematical Biology, 78(4):919–942, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1297-8;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1297-8.pdf.

Amorim:2019:ANM

[2875] Paulo Amorim, Thierry Goudon, and Fernando Peruani. An ant navigation model based on Weber’s law. Journal of Mathematical Biology, 78 (4):943–984, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1298-7;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1298-7.pdf.

Yegorov:2019:OCB

[2876] Ivan Yegorov, Francis Mairet, Hidde de Jong, and Jean-Luc Gouzé. Optimal control of bacterial growth for the maximization of metabolite production. Journal of Mathematical Biology, 78(4):985–1032, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1299-6;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1299-6.pdf.

deSouza:2019:FPM

[2877] Evandro P. de Souza, Eliza M. Ferreira, and Armando G. M. Neves. Fixation probabilities for the Moran process in evolutionary games with two strategies: graph shapes and large population asymptotics. Journal of Mathematical Biology, 78(4):1033–1065, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1300-4;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1300-4.pdf.

Bayen:2019:SSO

[2878] Térence Bayen and Pedro Gajardo. On the steady state optimization of the biogas production in a two-stage anaerobic digestion model. Journal of Mathematical Biology, 78(4):1067–1087, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1301-3;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1301-3.pdf.

Buonomo:2019:OTP

[2879] Bruno Buonomo, Piero Manfredi, and Alberto d’Onofrio. Optimal timeprofiles of public health intervention to shape voluntary vaccination for childhood diseases. Journal of Mathematical Biology, 78(4):1089–1113, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1303-1;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1303-1.pdf.

Giaroli:2019:RMC

[2880] Magalí Giaroli, Frédéric Bihan, and Alicia Dickenstein. Regions of multistationarity in cascades of Goldbeter–Koshland loops. Journal of Mathematical Biology, 78(4):1115–1145, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1304-0;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1304-
0.pdf.

Allen:2019:MFN

[2881] Benjamin Allen and Alex McAvoy. A mathematical formalism for natural selection with arbitrary spatial and genetic structure. Journal of Mathematical Biology, 78(4):1147–1210, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1305-z;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1305-
z.pdf.

Burden:2019:SDS

[2882] Conrad J. Burden and Robert C. Griffiths. The stationary distribution of a sample from the Wright–Fisher diffusion model with general small mutation rates. Journal of Mathematical Biology, 78(4):1211–1224, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1306-y;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1306-y.pdf.

Rozikov:2019:DTD

[2883] U. A. Rozikov and M. V. Velasco. A discrete-time dynamical system and an evolution algebra of mosquito population. Journal of Mathematical Biology, 78(4):1225–1244, March 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1307-x;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1307-
x.pdf.

Moraki:2019:SMC

[2884] E. Moraki, R. Grima, and K. J. Painter. A stochastic model of corneal epithelium maintenance and recovery following perturbation. Journal of Mathematical Biology, 78(5):1245–1276, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1308-9;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1308-9.pdf.

Li:2019:IDM

[2885] Beibei Li, Steven M. Roper, Lei Wang, Xiaoyu Luo, and N. A. Hill. An incremental deformation model of arterial dissection. Journal of Mathematical Biology, 78(5):1277–1298, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1309-8;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1309-
8.pdf.

Stadler:2019:ESA

[2886] Eva Stadler. Eigensolutions and spectral analysis of a model for vertical gene transfer of plasmids. Journal of Mathematical Biology, 78 (5):1299–1330, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1310-2;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1310-2.pdf.

Jennings:2019:HTK

[2887] Rachel Jennings, Yang Kuang, Horst R. Thieme, Jianhong Wu, and Xiaotian Wu. How ticks keep ticking in the adversity of host immune reactions. Journal of Mathematical Biology, 78(5):1331–1364, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1311-1;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1311-1.pdf.

Barany:2019:ISS

[2888] Balázs Bárány, Gregory Moses, and Todd Young. Instability of the steady state solution in cell cycle population structure models with feedback. Journal of Mathematical Biology, 78(5):1365–1387, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1312-0;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1312-0.pdf.

Wang:2019:UNS

[2889] Aili Wang, Yanni Xiao, and Robert Smith. Using non-smooth models to determine thresholds for microbial pest management. Journal of Mathematical Biology, 78(5):1389–1424, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1313-z;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1313-
z.pdf.

Befekadu:2019:OCD

[2890] Getachew K. Befekadu and Quanyan Zhu. Optimal control of diffusion processes pertaining to an opioid epidemic dynamical model with random perturbations. Journal of Mathematical Biology, 78(5):1425–1438, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1314-y;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1314-y.pdf.

Morais:2019:CPD

[2891] J. V. Morais, A. L. Custódio, and G. M. Marques. Calibration of parameters in Dynamic Energy Budget models using Direct–Search methods. Journal of Mathematical Biology, 78(5):1439–1458, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1315-x;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1315-x.pdf.

Tefagh:2019:QFC

[2892] Mojtaba Tefagh and Stephen P. Boyd. Quantitative flux coupling analysis. Journal of Mathematical Biology, 78(5):1459–1484, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1316-9;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1316-9.pdf.

Alqawasmeh:2019:PSS

[2893] Yousef Alqawasmeh and Frithjof Lutscher. Persistence and spread of stage-structured populations in heterogeneous landscapes. Journal of Mathematical Biology, 78(5):1485–1527, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1317-8;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1317-
8.pdf.

Kesmia:2019:NDT

[2894] M. Kesmia, S. Boughaba, and S. Jacquir. Nonlinear dynamics of two-dimensional cardiac action potential duration mapping model with memory. Journal of Mathematical Biology, 78(5):1529–1552, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1318-7;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1318-7.pdf.

Li:2019:PSE

[2895] Fuxiang Li and Xiao-Qiang Zhao. A periodic SEIRS epidemic model with a time-dependent latent period. Journal of Mathematical Biology, 78(5):1553–1579, April 2019. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL http://link.springer.com/article/10.
1007/s00285-018-1319-6;   http://link.springer.com/content/
pdf/10.1007/s00285-018-1319-6.pdf.

Guan:2019:CNS

[2896] Li Guan, Dong Li, Ke Wang, and Kun Zhao. On a class of nonlocal SIR models. Journal of Mathematical Biology, 78(6):1581–1604, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1320-0;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1320-0.pdf.

He:2019:DCR

[2897] Xiaoqing He, King-Yeung Lam, Yuan Lou, and Wei-Ming Ni. Dynamics of a consumer-resource reaction–diffusion model. Journal of Mathematical Biology, 78(6):1605–1636, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1321-z;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1321-
z.pdf.

Souza:2019:ROD

[2898] Daniel C. De Souza and Michael C. Mackey. Response of an oscillatory differential delay equation to a periodic stimulus. Journal of Mathematical Biology, 78(6):1637–1679, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

http://link.springer.com/article/10.1007/s00285-018-1322-y;
 http://link.springer.com/content/pdf/10.1007/s00285-018-1322-
y.pdf.

Plaza:2019:DBN

[2899] Ramón G. Plaza. Derivation of a bacterial nutrient-taxis system with doubly degenerate cross-diffusion as the parabolic limit of a velocityjump process. Journal of Mathematical Biology, 78(6):1681–1711, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1323-x;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1323-x.pdf.

Meehan:2019:GSP

[2900] Michael T. Meehan, Daniel G. Cocks, Johannes Müller, and Emma S. McBryde. Global stability properties of a class of renewal epidemic models. Journal of Mathematical Biology, 78(6):1713–1725, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-018-01324-1;   http://link.springer.com/content/pdf/
10.1007/s00285-018-01324-1.pdf.

Speed:2019:GFM

[2901] Maria Simonsen Speed, David Joseph Balding, and Asger Hobolth. A general framework for moment-based analysis of genetic data. Journal of Mathematical Biology, 78(6):1727–1769, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/s00285-018-01325-0;   http://link.
springer.com/content/pdf/10.1007/s00285-018-01325-0.pdf.

Bossy:2019:SSM

[2902] Mireille Bossy, Joaquín Fontbona, and Héctor Olivero. Synchronization of stochastic mean field networks of Hodgkin–Huxley neurons with noisy channels. Journal of Mathematical Biology, 78(6):1771–1820, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01326-7;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01326-7.pdf.

Kura:2019:OCD

[2903] Klodeta Kura, Doran Khamis, Claire El Mouden, and Michael B. Bonsall. Optimal control for disease vector management in SIT models: an integrodifference equation approach. Journal of Mathematical Biology, 78(6):1821–1839, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01327-6;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01327-6.pdf.

Alsmeyer:2019:EPD

[2904] Gerold Alsmeyer and Kilian Raschel. The extinction problem for a distylous plant population with sporophytic self-incompatibility. Journal of Mathematical Biology, 78(6):1841–1874, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/s00285-019-01328-5;   http://link.
springer.com/content/pdf/10.1007/s00285-019-01328-5.pdf.

Ball:2019:SSN

[2905] Frank Ball, Tom Britton, Ka Yin Leung, and David Sirl. A stochastic SIR network epidemic model with preventive dropping of edges. Journal of Mathematical Biology, 78(6):1875–1951, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01329-4;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01329-4.pdf.

Herbst:2019:QAA

[2906] Lina Herbst, Heyang Li, and Mike Steel. Quantifying the accuracy of ancestral state prediction in a phylogenetic tree under maximum parsimony. Journal of Mathematical Biology, 78(6):1953–1979, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01330-x;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01330-x.pdf.

Hasic:2019:GTS

[2907] Damir Hasić and Eric Tannier. Gene tree species tree reconciliation with gene conversion. Journal of Mathematical Biology, 78(6):1981– 2014, May 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01331-w;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01331-w.pdf.

Geiss:2019:BMG

[2908] Manuela Geiß, Edgar Chávez, Marcos González Laffitte, Alitzel López Sánchez, Bärbel M. R. Stadler, Dulce I. Valdivia, Marc Hellmuth, Maribel Hernández Rosales, and Peter F. Stadler. Best match graphs. Journal of Mathematical Biology, 78(7):2015–2057, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01332-9;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01332-9.pdf.

Kaouri:2019:SMM

[2909] K. Kaouri, P. K. Maini, P. A. Skourides, N. Christodoulou, and S. J. Chapman. A simple mechanochemical model for calcium signalling in embryonic epithelial cells. Journal of Mathematical Biology, 78(7):2059–2092, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01333-8;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01333-8.pdf.

Wang:2019:PEP

[2910] Yan Wang, Junping Shi, and Jinfeng Wang. Persistence and extinction of population in reaction–diffusion–advection model with strong Allee effect growth. Journal of Mathematical Biology, 78(7):2093–2140, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01334-7;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01334-7.pdf.

Choi:2019:IPE

[2911] Wonhyung Choi, Seunghyeon Baek, and Inkyung Ahn. Intraguild predation with evolutionary dispersal in a spatially heterogeneous environment. Journal of Mathematical Biology, 78(7):2141–2169, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01336-5;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01336-5.pdf.

During:2019:AIM

[2912] Bertram Düring, Carsten Gottschlich, Stephan Huckemann, Lisa Maria Kreusser, and Carola-Bibiane Schönlieb. An anisotropic interaction model for simulating fingerprints. Journal of Mathematical Biology, 78(7):2171–2206, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01338-3;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01338-3.pdf.

Oelker:2019:SAG

[2913] A. Oelker, T. Horger, and C. Kuttler. From Staphylococcus aureus gene regulation to its pattern formation. Journal of Mathematical Biology, 78(7):2207–2234, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01340-9;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01340-9.pdf.

Duarte:2019:CAE

[2914] Jorge Duarte, Cristina Januário, Nuno Martins, Svitlana Rogovchenko, and Yuriy Rogovchenko. Chaos analysis and explicit series solutions to the seasonally forced SIR epidemic model. Journal of Mathematical Biology, 78(7):2235–2258, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01342-7;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01342-7.pdf.

Lou:2019:MDM

[2915] Yijun Lou, Kaihui Liu, Daihai He, Daozhou Gao, and Shigui Ruan. Modelling diapause in mosquito population growth. Journal of Mathematical Biology, 78(7):2259–2288, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01343-6;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01343-6.pdf.

Malik:2019:MMC

[2916] Adam A. Malik and Philip Gerlee. Mathematical modelling of cell migration: stiffness dependent jump rates result in durotaxis. Journal of Mathematical Biology, 78(7):2289–2315, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01344-5;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01344-5.pdf.

Lewis:2019:GTT

[2917] Mark A. Lewis, Zhisheng Shuai, and P. van den Driessche. A general theory for target reproduction numbers with applications to ecology and epidemiology. Journal of Mathematical Biology, 78(7):2317–2339, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01345-4;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01345-4.pdf.

Tyler:2019:RSI

[2918] Jonathan Tyler, Anne Shiu, and Jay Walton. Revisiting a synthetic intracellular regulatory network that exhibits oscillations. Journal of Mathematical Biology, 78(7):2341–2368, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/s00285-019-01346-3;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01346-3.pdf.

Fransson:2019:SEV

[2919] Carolina Fransson and Pieter Trapman. SIR epidemics and vaccination on random graphs with clustering. Journal of Mathematical Biology, 78(7):2369–2398, June 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01347-2;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01347-2.pdf.

Tillquist:2019:LDR

[2920] Richard C. Tillquist and Manuel E. Lladser. Low-dimensional representation of genomic sequences. Journal of Mathematical Biology, 79(1): 1–29, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01348-1;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01348-1.pdf.

Montagnon:2019:SSM

[2921] Pierre Montagnon. A stochastic SIR model on a graph with epidemiological and population dynamics occurring over the same time scale. Journal of Mathematical Biology, 79(1):31–62, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01349-0;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01349-0.pdf.

Vasilyeva:2019:PDR

[2922] Olga Vasilyeva. Population dynamics in river networks: analysis of steady states. Journal of Mathematical Biology, 79(1):63–100, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/

s00285-019-01350-7;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01350-7.pdf.

Cooney:2019:RDM

[2923] Daniel B. Cooney. The replicator dynamics for multilevel selection in evolutionary games. Journal of Mathematical Biology, 79(1):101–154, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01352-5;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01352-5.pdf.

Cartailler:2019:SSV

[2924] J. Cartailler and D. Holcman. Steady-state voltage distribution in threedimensional cusp-shaped funnels modeled by PNP. Journal of Mathematical Biology, 79(1):155–185, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01353-4;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01353-4.pdf.

Kristiansen:2019:GSP

[2925] Kristian Uldall Kristiansen. Geometric singular perturbation analysis of a dynamical target mediated drug disposition model. Journal of Mathematical Biology, 79(1):187–222, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/s00285-019-01354-3;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01354-3.pdf.

Esposito:2019:OJS

[2926] Flavia Esposito, Nicolas Gillis, and Nicoletta Del Buono. Orthogonal joint sparse NMF for microarray data analysis. Journal of Mathematical Biology, 79(1):223–247, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01355-2;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01355-2.pdf.

Ruschel:2019:SDD

[2927] Stefan Ruschel, Tiago Pereira, Serhiy Yanchuk, and Lai-Sang Young. An SIQ delay differential equations model for disease control via isolation. Journal of Mathematical Biology, 79(1):249–279, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01356-1;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01356-1.pdf.

Getto:2019:SAS

[2928] Philipp Getto, Mats Gyllenberg, Yukihiko Nakata, and Francesca Scarabel. Stability analysis of a state-dependent delay differential equation for cell maturation: analytical and numerical methods. Journal of Mathematical Biology, 79(1):281–328, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/s00285-019-01357-0;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01357-0.pdf.

Kato:2019:SBA

[2929] Hayato Kato and Takenori Takada. Stability and bifurcation analysis of a ratio-dependent community dynamics model on Batesian mimicry. Journal of Mathematical Biology, 79(1):329–368, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01359-y;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01359-y.pdf.

Blath:2019:SPS

[2930] Jochen Blath, Eugenio Buzzoni, Adrián González Casanova, and Maite Wilke Berenguer. Structural properties of the seed bank and the two island diffusion. Journal of Mathematical Biology, 79(1):369–392, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01360-5;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01360-5.pdf.

Benaim:2019:PES

[2931] Michel Benaïm and Sebastian J. Schreiber. Persistence and extinction for stochastic ecological models with internal and external variables. Journal of Mathematical Biology, 79(1):393–431, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01361-4;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01361-4.pdf.

Wang:2019:SSW

[2932] Zhiguo Wang, Hua Nie, and Yihong Du. Spreading speed for a West Nile virus model with free boundary. Journal of Mathematical Biology, 79(2):433–466, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01363-2;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01363-2.pdf.

Gonzalez-Parra:2019:RVT

[2933] Gilberto González-Parra and Hana M. Dobrovolny. The rate of viral transfer between upper and lower respiratory tracts determines RSV illness duration. Journal of Mathematical Biology, 79(2):467–483, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01364-1;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01364-1.pdf.

Felipe:2019:DAB

[2934] Miraine Dávila Felipe, Jean-Baka Domelevo Entfellner, Frédéric Lemoine, Jakub Truszkowski, and Olivier Gascuel. Distribution and asymptotic behavior of the phylogenetic transfer distance. Journal of Mathematical Biology, 79(2):485–508, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/s00285-019-01365-0;   http://link.
springer.com/content/pdf/10.1007/s00285-019-01365-0.pdf.

Yamakou:2019:SFN

[2935] Marius E. Yamakou, Tat Dat Tran, Luu Hoang Duc, and Jürgen Jost. The stochastic Fitzhugh–Nagumo neuron model in the excitable regime embeds a leaky integrate-and-fire model. Journal of Mathematical Biology, 79(2):509–532, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01366-z;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01366-z.pdf.

Hening:2019:HIS

[2936] Alexandru Hening, Ky Quan Tran, Tien Trong Phan, and George Yin. Harvesting of interacting stochastic populations. Journal of Mathematical Biology, 79(2):533–570, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01368-x;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01368-x.pdf.

Allard:2019:BST

[2937] Jun Allard, Marie Doumic, Alex Mogilner, and Dietmar Oelz. Bidirectional sliding of two parallel microtubules generated by multiple identical motors. Journal of Mathematical Biology, 79(2):571–594, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01369-w;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01369-w.pdf.

Ward:2019:MMC

[2938] J. P. Ward, S. J. Franks, M. J. Tindall, J. R. King, A. Curtis, and G. S. Evans. Mathematical modelling of contact dermatitis from nickel and chromium. Journal of Mathematical Biology, 79(2):595–630, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01371-2;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01371-2.pdf.

Hazra:2019:SGP

[2939] Tania Hazra, Sheik Ahmed Ullah, Siwen Wang, Emil Alexov, and Shan Zhao. A super-Gaussian Poisson–Boltzmann model for electrostatic free energy calculation: smooth dielectric distribution for protein cavities and in both water and vacuum states. Journal of Mathematical Biology, 79(2):631–672, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01372-1;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01372-1.pdf.

Meli:2019:SPP

[2940] Gianfelice Meli, Tom S. Weber, and Ken R. Duffy. Sample path properties of the average generation of a Bellman–Harris process. Journal of Mathematical Biology, 79(2):673–704, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01373-0;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01373-0.pdf.

Mummert:2019:PIS

[2941] Anna Mummert and Olusegun M. Otunuga. Parameter identification for a stochastic SEIRS epidemic model: case study influenza. Journal of Mathematical Biology, 79(2):705–729, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01374-z;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01374-z.pdf.

Inaba:2019:BRN

[2942] Hisashi Inaba. The basic reproduction number R_{0} in time-heterogeneous environments. Journal of Mathematical Biology, 79(2):731–764, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01375-y;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01375-y.pdf.

Abboud:2019:DLI

[2943] Candy Abboud, Olivier Bonnefon, Eric Parent, and Samuel Soubeyrand. Dating and localizing an invasion from post-introduction data and a coupled reaction–diffusion–absorption model. Journal of Mathematical Biology, 79(2):765–789, July 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01376-x;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01376-x.pdf.

Li:2019:BSR

[2944] Thomas J. X. Li, Christie S. Burris, and Christian M. Reidys. The block spectrum of RNA pseudoknot structures. Journal of Mathematical Biology, 79(3):791–822, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01379-8;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01379-8.pdf.

Barnard:2019:ETP

[2945] Rosanna C. Barnard, Luc Berthouze, Péter L. Simon, and István Z. Kiss. Epidemic threshold in pairwise models for clustered networks: closures and fast correlations. Journal of Mathematical Biology, 79(3):823–860, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01380-1;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01380-1.pdf.

Eble:2019:CPF

[2946] Holger Eble, Michael Joswig, Lisa Lamberti, and William B. Ludington. Cluster partitions and fitness landscapes of the Drosophila fly microbiome. Journal of Mathematical Biology, 79(3):861–899, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01381-0;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01381-0.pdf.

Meleard:2019:BDM

[2947] Sylvie Méléard, Michael Rera, and Tristan Roget. A birth-death model of ageing: from individual-based dynamics to evolutive differential inclusions. Journal of Mathematical Biology, 79(3):901–939, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01382-z;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01382-z.pdf.

Subramanian:2019:SGG

[2948] Shashank Subramanian, Amir Gholami, and George Biros. Simulation of glioblastoma growth using a 3D multispecies tumor model with mass effect. Journal of Mathematical Biology, 79(3):941–967, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01383-y;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01383-y.pdf.

Hellmuth:2019:ACF

[2949] Marc Hellmuth and Carsten R. Seemann. Alternative characterizations of Fitch’s xenology relation. Journal of Mathematical Biology, 79(3):969–986, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01384-x;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01384-x.pdf.

Williams:2019:CDD

[2950] Nakeya D. Williams, Renee Brady, Steven Gilmore, Pierre Gremaud, Hien T. Tran, Johnny T. Ottesen, Jesper Mehlsen, and Mette S. Olufsen. Cardiovascular dynamics during head-up tilt assessed via pulsatile and non-pulsatile models. Journal of Mathematical Biology, 79(3):987–1014, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01386-9;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01386-9.pdf.

Almarashi:2019:EII

[2951] Reem M. Almarashi and C. Connell McCluskey. The effect of immigration of infectives on disease-free equilibria. Journal of Mathematical Biology, 79(3):1015–1028, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01387-8;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01387-8.pdf.

Franco:2019:BPS

[2952] D. Franco, C. Guiver, H. Logemann, and J. Perán. Boundedness, persistence and stability for classes of forced difference equations arising in population ecology. Journal of Mathematical Biology, 79(3):1029–1076, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01388-7;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01388-7.pdf.

Miyaoka:2019:OCV

[2953] Tiago Yuzo Miyaoka, Suzanne Lenhart, and João F. C. A. Meyer. Optimal control of vaccination in a vector-borne reaction–diffusion model applied to Zika virus. Journal of Mathematical Biology, 79(3):1077–1104, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01390-z;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01390-z.pdf.

Coronado:2019:BIP

[2954] Tomás M. Coronado, Arnau Mir, Francesc Rosselló, and Gabriel Valiente. A balance index for phylogenetic trees based on rooted quartets. Journal of Mathematical Biology, 79(3):1105–1148, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01377-w;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01377-w.pdf.

Calamoneri:2019:CDT

[2955] T. Calamoneri, A. Monti, and B. Sinaimeri. Co-divergence and tree topology. Journal of Mathematical Biology, 79(3):1149–1167, August 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01385-w;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01385-w.pdf.

Li:2019:SNF

[2956] Yao Li and Hui Xu. Stochastic neural field model: multiple firing events and correlations. Journal of Mathematical Biology, 79(4): 1169–1204, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01389-6;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01389-6.pdf.

Disanto:2019:UTM

[2957] Filippo Disanto, Pasquale Miglionico, and Guido Narduzzi. On the unranked topology of maximally probable ranked gene tree topologies. Journal of Mathematical Biology, 79(4):1205–1225, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01392-x;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01392-x.pdf.

Kuehn:2019:GFF

[2958] Christian Kuehn and Jonas M. Tölle. A gradient flow formulation for the stochastic Amari neural field model. Journal of Mathematical Biology, 79(4):1227–1252, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01393-w;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01393-w.pdf.

Anderson:2019:DBE

[2959] David F. Anderson and Daniele Cappelletti. Discrepancies between extinction events and boundary equilibria in reaction networks. Journal of Mathematical Biology, 79(4):1253–1277, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01394-9;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01394-9.pdf.

Li:2019:DAP

[2960] Huicong Li and Rui Peng. Dynamics and asymptotic profiles of endemic equilibrium for SIS epidemic patch models. Journal of Mathematical Biology, 79(4):1279–1317, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01395-8;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01395-8.pdf.

Tian:2019:SDR

[2961] Chenwei Tian, Qingyan Shi, Xinping Cui, Jingzhe Guo, Zhenbiao Yang, and Junping Shi. Spatiotemporal dynamics of a reaction-diffusion model of pollen tube tip growth. Journal of Mathematical Biology, 79(4):1319–1355, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01396-7;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01396-7.pdf.

Espejo:2019:SBP

[2962] Elio Espejo, Karina Vilches, and Carlos Conca. A simultaneous blowup problem arising in tumor modeling. Journal of Mathematical Biology, 79(4):1357–1399, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01397-6;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01397-6.pdf.

Zadorin:2019:NSC

[2963] A. S. Zadorin and Y. Rondelez. Natural selection in compartmentalized environment with reshuffling. Journal of Mathematical Biology, 79(4):1401–1454, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01399-4;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01399-4.pdf.

Salako:2019:CCS

[2964] Rachidi B. Salako, Wenxian Shen, and Shuwen Xue. Can chemotaxis speed up or slow down the spatial spreading in parabolic-elliptic Keller– Segel systems with logistic source? Journal of Mathematical Biology, 79(4):1455–1490, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01400-0;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01400-0.pdf.

Xin:2019:OMI

[2965] Ying Xin, David Gerberry, and Winfried Just. Open-minded imitation can achieve near-optimal vaccination coverage. Journal of Mathematical Biology, 79(4):1491–1514, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/s00285-019-01401-z;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01401-z.pdf.

Obatake:2019:OBM

[2966] Nida Obatake, Anne Shiu, Xiaoxian Tang, and Angélica Torres. Oscillations and bistability in a model of ERK regulation. Journal of Mathematical Biology, 79(4):1515–1549, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01402-y;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01402-y.pdf.

Pargaei:2019:MSC

[2967] Meena Pargaei, B. V. Rathish Kumar, and Luca F. Pavarino. Modeling and simulation of cardiac electric activity in a human cardiac tissue with multiple ischemic zones. Journal of Mathematical Biology, 79(4):1551–1586, September 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01403-x;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01403-x.pdf.

Lorenzi:2019:SPM

[2968] Tommaso Lorenzi, Anna Marciniak-Czochra, and Thomas Stiehl. A structured population model of clonal selection in acute leukemias with multiple maturation stages. Journal of Mathematical Biology, 79(5):1587–1621, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01404-w;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01404-w.pdf.

Erdos:2019:APN

[2969] Péter L. Erdős, Leo van Iersel, and Mark Jones. Not all phylogenetic networks are leaf-reconstructible. Journal of Mathematical Biology, 79(5):1623–1638, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01405-9;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01405-9.pdf.

Ajazi:2019:NRT

[2970] Fioralba Ajazi, Valérie Chavez-Demoulin, and Tatyana Turova. Networks of random trees as a model of neuronal connectivity. Journal of Mathematical Biology, 79(5):1639–1663, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/s00285-019-01406-8;   http://link.
springer.com/content/pdf/10.1007/s00285-019-01406-8.pdf.

Etchegaray:2019:SMC

[2971] Christèle Etchegaray and Nicolas Meunier. A stochastic model for cell adhesion to the vascular wall. Journal of Mathematical Biology, 79(5):1665–1697, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01407-7;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01407-7.pdf.

Kraut:2019:ADA

[2972] Anna Kraut and Anton Bovier. From adaptive dynamics to adaptive walks. Journal of Mathematical Biology, 79(5):1699–1747, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01408-6;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01408-6.pdf.

Rohl:2019:FMM

[2973] Annika Röhl and Alexander Bockmayr. Finding MEMo: minimum sets of elementary flux modes. Journal of Mathematical Biology, 79 (5):1749–1777, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01409-5;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01409-5.pdf.

Sunkara:2019:AEC

[2974] Vikram Sunkara. Algebraic expressions of conditional expectations in gene regulatory networks. Journal of Mathematical Biology, 79 (5):1779–1829, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01410-y;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01410-y.pdf.

Hurtado:2019:GLC

[2975] Paul J. Hurtado and Adam S. Kirosingh. Generalizations of the ‘linear chain trick’: incorporating more flexible dwell time distributions into mean field ODE models. Journal of Mathematical Biology, 79 (5):1831–1883, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01412-w;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01412-w.pdf.

Hellmuth:2019:REL

[2976] Marc Hellmuth, Katharina T. Huber, and Vincent Moulton. Reconciling event-labeled gene trees with MUL-trees and species networks. Journal of Mathematical Biology, 79(5):1885–1925, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01414-8;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01414-8.pdf.

Hilker:2019:PTH

[2977] Frank M. Hilker and Eduardo Liz. Proportional threshold harvesting in discrete-time population models. Journal of Mathematical Biology, 79(5):1927–1951, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01415-7;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01415-7.pdf.

Buttenschon:2019:CRW

[2978] Andreas Buttenschön and Leah Edelstein-Keshet. Correlated random walks inside a cell: actin branching and microtubule dynamics. Journal of Mathematical Biology, 79(5):1953–1972, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01416-6;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01416-6.pdf.

Dragicevic:2019:CRC

[2979] Arnaud Z. Dragicevic. Conditional rehabilitation of cooperation under strategic uncertainty. Journal of Mathematical Biology, 79(5):1973–2003, October 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01417-5;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01417-5.pdf.

Ramos:2019:KMC

[2980] M. P. Machado Ramos, C. Ribeiro, and A. J. Soares. A kinetic model of T cell autoreactivity in autoimmune diseases. Journal of Mathematical Biology, 79(6–7):2005–2031, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/s00285-019-01418-4;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01418-4.pdf.

Kunisch:2019:ILE

[2981] Karl Kunisch, Aurel Neic, Gernot Plank, and Philip Trautmann. Inverse localization of earliest cardiac activation sites from activation maps based on the viscous eikonal equation. Journal of Mathematical Biology, 79(6–7):2033–2068, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01419-3;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01419-3.pdf.

Soares:2019:DAA

[2982] Cíntia Dalila Soares and Sabin Lessard. Diffusion approximation for an age-class-structured population under viability and fertility selection with application to fixation probability of an advantageous mutant. Journal of Mathematical Biology, 79(6–7):2069–2110, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01420-w;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01420-w.pdf.

Laurent-Gengoux:2019:MAM

[2983] Pascal Laurent-Gengoux, Valérie Petit, and Lionel Larue. Modeling and analysis of melanoblast motion. Journal of Mathematical Biology, 79(6–7):2111–2132, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01422-8;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01422-8.pdf.

Mata:2019:RFA

[2984] May Anne Mata, Rebecca C. Tyson, and Priscilla Greenwood. Random fluctuations around a stable limit cycle in a stochastic system with parametric forcing. Journal of Mathematical Biology, 79(6–7): 2133–2155, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01423-7;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01423-7.pdf.

Wang:2019:DBS

[2985] Zhen Wang, Gergely Röst, and Seyed M. Moghadas. Delay in booster schedule as a control parameter in vaccination dynamics. Journal of Mathematical Biology, 79(6–7):2157–2182, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01424-6;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01424-6.pdf.

Asfaw:2019:SPH

[2986] Manalebish Debalike Asfaw, Semu Mitiku Kassa, and Edward M. Lungu. Stochastic plant-herbivore interaction model with Allee effect. Journal of Mathematical Biology, 79(6–7):2183–2209, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01425-5;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01425-5.pdf.

Albert:2019:PIA

[2987] Jaroslav Albert. Path integral approach to generating functions for multistep post-transcription and post-translation processes and arbitrary initial conditions. Journal of Mathematical Biology, 79(6–7):

2211–2236, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01426-4;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01426-4.pdf.

Perepelitsa:2019:ALL

[2988] Misha Perepelitsa. Adaptive learning in large populations. Journal of Mathematical Biology, 79(6–7):2237–2253, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://

link.springer.com/article/10.1007/s00285-019-01427-3;   http:
//link.springer.com/content/pdf/10.1007/s00285-019-01427-3.
pdf.

Gao:2019:HFP

[2989] Daozhou Gao, P. van den Driessche, and Chris Cosner. Habitat fragmentation promotes malaria persistence. Journal of Mathematical Biology, 79(6–7):2255–2280, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01428-2;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01428-2.pdf.

Rogg:2019:OED

[2990] S. Rogg, D. H. Fuertinger, S. Volkwein, F. Kappel, and P. Kotanko. Optimal EPO dosing in hemodialysis patients using a non-linear model predictive control approach. Journal of Mathematical Biology, 79(6– 7):2281–2313, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01429-1;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01429-1.pdf.

Burden:2019:TDS

[2991] Conrad J. Burden and Robert C. Griffiths. The transition distribution of a sample from a Wright–Fisher diffusion with general small mutation rates. Journal of Mathematical Biology, 79(6–7):2315–2342, December 2019. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01430-8;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01430-8.pdf.

Diekmann:2020:SIJ

[2992] Odo Diekmann, Sergey Gavrilets, Mats Gyllenberg, Simon Levin, and Mark Lewis. Special issue of the Journal of Mathematical Biology to honor Alan Hastings’ 65th birthday. Journal of Mathematical Biology, 80(1–2):1–2, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01472-3;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01472-3.pdf.

Cantrell:2020:EDS

[2993] Robert Stephen Cantrell, Chris Cosner, Mark A. Lewis, and Yuan Lou. Evolution of dispersal in spatial population models with multiple timescales. Journal of Mathematical Biology, 80(1–2):3–37, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00285-
018-1302-2;   http://link.springer.com/content/pdf/10.1007/
s00285-018-1302-2.pdf.

Poggiale:2020:APP

[2994] Jean-Christophe Poggiale, Clément Aldebert, Benjamin Girardot, and Bob W. Kooi. Analysis of a predator–prey model with specific time scales: a geometrical approach proving the occurrence of canard solutions. Journal of Mathematical Biology, 80(1–2):39–60, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01337-4;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01337-4.pdf.

Maciel:2020:ESM

[2995] Gabriel Maciel, Chris Cosner, Robert Stephen Cantrell, and Frithjof Lutscher. Evolutionarily stable movement strategies in reaction-diffusion models with edge behavior. Journal of Mathematical Biology, 80(1–2):61– 92, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01339-2;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01339-2.pdf.

Labrum:2020:MAE

[2996] Matthew J. Labrum and Richard Gomulkiewicz. Metapopulation Allee effects, habitat destruction, and extinction in metacommunities. Journal of Mathematical Biology, 80(1–2):93–110, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01341-8;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01341-8.pdf.

Sandhu:2020:MEV

[2997] Simran K. Sandhu, Andrew Yu. Morozov, and József Z. Farkas. Modelling evolution of virulence in populations with a distributed parasite load. Journal of Mathematical Biology, 80(1–2):111–141, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01351-6;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01351-6.pdf.

Boettiger:2020:BST

[2998] Carl Boettiger and Ryan Batt. Bifurcation or state tipping: assessing transition type in a model trophic cascade. Journal of Mathematical Biology, 80(1–2):143–155, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01358-z;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01358-z.pdf.

Marculis:2020:IDS

[2999] Nathan G. Marculis, Jimmy Garnier, Roger Lui, and Mark A. Lewis. Inside dynamics for stage-structured integrodifference equations. Journal of Mathematical Biology, 80(1–2):157–187, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01378-9;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01378-9.pdf.

Diekmann:2020:MPS

[3000] Odo Diekmann, Mats Gyllenberg, and Johan A. J. Metz. On models of physiologically structured populations and their reduction to ordinary differential equations. Journal of Mathematical Biology, 80 (1–2):189–204, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01431-7;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01431-7.pdf.

Diekmann:2020:FDS

[3001] Odo Diekmann, Mats Gyllenberg, and Johan A. J. Metz. Finite dimensional state representation of physiologically structured populations. Journal of Mathematical Biology, 80(1–2):205–273, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01454-0;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01454-0.pdf.

Hillen:2020:MPP

[3002] Thomas Hillen, Kevin J. Painter, Magdalena A. Stolarska, and Chuan Xue. Multiscale phenomena and patterns in biological systems: special issue in honour of Hans Othmer. Journal of Mathematical Biology, 80(1–2):275–281, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01473-2;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01473-2.pdf.

Diekmann:2020:WTI

[3003] Odo Diekmann and Robert Planqué. The winner takes it all: how semelparous insects can become periodical. Journal of Mathematical Biology, 80(1–2):283–301, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01362-3;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01362-3.pdf.

Contento:2020:CPF

[3004] Lorenzo Contento and Masayasu Mimura. Complex pattern formation driven by the interaction of stable fronts in a competitiondiffusion system. Journal of Mathematical Biology, 80(1–2):303–342, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01370-3;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01370-3.pdf.

Chaplain:2020:BGB

[3005] Mark A. J. Chaplain, Tommaso Lorenzi, and Fiona R. Macfarlane. Bridging the gap between individual-based and continuum models of growing cell populations. Journal of Mathematical Biology, 80(1–2):343–371, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01391-y;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01391-y.pdf.

Loy:2020:KMN

[3006] Nadia Loy and Luigi Preziosi. Kinetic models with non-local sensing determining cell polarization and speed according to independent cues. Journal of Mathematical Biology, 80(1–2):373–421, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01411-x;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01411-x.pdf.

Colombi:2020:MCR

[3007] A. Colombi, M. Scianna, K. J. Painter, and L. Preziosi. Modelling chaseand-run migration in heterogeneous populations. Journal of Mathematical Biology, 80(1–2):423–456, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/s00285-019-01421-9;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01421-9.pdf.

Erban:2020:CGM

[3008] Radek Erban. Coarse-graining molecular dynamics: stochastic models with non-Gaussian force distributions. Journal of Mathematical Biology, 80(1–2):457–479, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01433-5;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01433-5.pdf.

Giniunaite:2020:MCC

[3009] Rasa Giniūnaitė, Ruth E. Baker, Paul M. Kulesa, and Philip K. Maini. Modelling collective cell migration: neural crest as a model paradigm. Journal of Mathematical Biology, 80(1–2):481–504, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01436-2;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01436-2.pdf.

Li:2020:EBM

[3010] Linlin Li, Xu Wang, Mary C. Mullins, and David M. Umulis. Evaluation of BMP-mediated patterning in a 3D mathematical model of the zebrafish blastula embryo. Journal of Mathematical Biology, 80 (1–2):505–520, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01449-x;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01449-x.pdf.

Hao:2020:SPF

[3011] Wenrui Hao and Chuan Xue. Spatial pattern formation in reaction– diffusion models: a computational approach. Journal of Mathematical Biology, 80(1–2):521–543, January 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01462-0;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01462-0.pdf.

Vermolen:2020:UQS

[3012] Fred Vermolen and Ilkka Pölönen. Uncertainty quantification on a spatial Markov-chain model for the progression of skin cancer. Journal of Mathematical Biology, 80(3):545–573, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01367-y;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01367-y.pdf.

Knauer:2020:OWB

[3013] Franziska Knauer, Thomas Stiehl, and Anna Marciniak-Czochra. Oscillations in a white blood cell production model with multiple differentiation stages. Journal of Mathematical Biology, 80(3):575–600, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01432-6;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01432-6.pdf.

Kuznetsov:2020:IST

[3014] Maxim Kuznetsov and Andrey Kolobov. Investigation of solid tumor progression with account of proliferation/migration dichotomy via Darwinian mathematical model. Journal of Mathematical Biology, 80(3):601–626, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01434-4;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01434-4.pdf.

Fuchs:2020:CBS

[3015] Michael Fuchs and Ariel R. Paningbatan. Correlation between Shapley values of rooted phylogenetic trees under the beta-splitting model. Journal of Mathematical Biology, 80(3):627–653, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01435-3;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01435-3.pdf.

Zaytseva:2020:MPF

[3016] Sofya Zaytseva, Junping Shi, and Leah B. Shaw. Model of pattern formation in marsh ecosystems with nonlocal interactions. Journal of Mathematical Biology, 80(3):655–686, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/s00285-019-01437-1;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01437-1.pdf.

Wicke:2020:CPP

[3017] Kristina Wicke and Mike Steel. Combinatorial properties of phylogenetic diversity indices. Journal of Mathematical Biology, 80(3):687–715, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01438-0;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01438-0.pdf.

Stahn:2020:BSV

[3018] Hubert Stahn. Biodiversity, Shapley value and phylogenetic trees: some remarks. Journal of Mathematical Biology, 80(3):717–741, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01439-z;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01439-z.pdf.

Varga:2020:EEM

[3019] Tamás Varga, Tamás F. Móri, and József Garay. The ESS for evolutionary matrix games under time constraints and its relationship with the asymptotically stable rest point of the replicator dynamics. Journal of Mathematical Biology, 80(3):743–774, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01440-6;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01440-6.pdf.

Ardaseva:2020:EDC

[3020] Aleksandra Ardaševa, Robert A. Gatenby, Alexander R. A. Anderson, Helen M. Byrne, Philip K. Maini, and Tommaso Lorenzi. Evolutionary dynamics of competing phenotype-structured populations in periodically fluctuating environments. Journal of Mathematical Biology, 80(3):775–807, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01441-5;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01441-5.pdf.

Qiu:2020:MDG

[3021] Zhipeng Qiu, Xuerui Wei, Chunhua Shan, and Huaiping Zhu. Monotone dynamics and global behaviors of a West Nile virus model with mosquito demographics. Journal of Mathematical Biology, 80(3):809–834, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01442-4;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01442-4.pdf.

Wang:2020:APR

[3022] Bin-Guo Wang, Lizhong Qiang, and Zhi-Cheng Wang. An almost periodic Ross–Macdonald model with structured vector population in a patchy environment. Journal of Mathematical Biology, 80(3):835–863, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01443-3;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01443-3.pdf.

Geiss:2020:RBM

[3023] Manuela Geiß, Peter F. Stadler, and Marc Hellmuth. Reciprocal best match graphs. Journal of Mathematical Biology, 80(3):865–953, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01444-2;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01444-2.pdf.

Kisdi:2020:CAD

[3024] Éva Kisdi and Stefan A. H. Geritz. Correction to: Adaptive dynamics of saturated polymorphisms. Journal of Mathematical Biology, 80(3):955–957, February 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01398-5;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01398-5.pdf. See [2495].

Chen:2020:LTG

[3025] Xian Chen and Chen Jia. Limit theorems for generalized densitydependent Markov chains and bursty stochastic gene regulatory networks. Journal of Mathematical Biology, 80(4):959–994, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01445-1;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01445-1.pdf.

Casanellas:2020:ERI

[3026] Marta Casanellas, Jesús Fernández-Sánchez, and Jordi Roca-Lacostena. Embeddability and rate identifiability of Kimura 2-parameter matrices. Journal of Mathematical Biology, 80(4):995–1019, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01446-0;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01446-0.pdf.

Bacaer:2020:DMP

[3027] Nicolas Bacaër. Deux modèles de population dans un environnement périodique lent ou rapide. (French) [Two population models in a slow or fast periodic environment]. Journal of Mathematical Biology, 80 (4):1021–1037, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01447-z;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01447-z.pdf.

Wattis:2020:MMT

[3028] Jonathan A. D. Wattis, Qi Qi, and Helen M. Byrne. Mathematical modelling of telomere length dynamics. Journal of Mathematical Biology, 80(4):1039–1076, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01448-y;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01448-y.pdf.

Donde:2020:UPP

[3029] Tobia Dondè. Uniform persistence in a prey–predator model with a diseased predator. Journal of Mathematical Biology, 80(4):1077–1093, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01451-3;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01451-3.pdf.

Bai:2020:BRR

[3030] Zhenguo Bai and Xiao-Qiang Zhao. Basic reproduction ratios for periodic and time-delayed compartmental models with impulses. Journal of Mathematical Biology, 80(4):1095–1117, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/s00285-019-01452-2;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01452-2.pdf.

Ho:2020:CML

[3031] Lam Si Tung Ho, Vu Dinh, Frederick A. Matsen IV, and Marc A. Suchard. On the convergence of the maximum likelihood estimator for the transition rate under a 2-state symmetric model. Journal of Mathematical Biology, 80(4):1119–1138, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.

com/article/10.1007/s00285-019-01453-1;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01453-1.pdf.

Brandle:2020:STA

[3032] Stefanie Brändle, Syn Schmitt, and Matthias A. Müller. A systemstheoretic analysis of low-level human motor control: application to a single-joint arm model. Journal of Mathematical Biology, 80(4):1139– 1158, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01455-z;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01455-z.pdf.

Feliu:2020:RAM

[3033] Elisenda Feliu. On the role of algebra in models in molecular biology. Journal of Mathematical Biology, 80(4):1159–1161, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01456-y;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01456-y.pdf.

Golubitsky:2020:IHT

[3034] Martin Golubitsky and Yangyang Wang. Infinitesimal homeostasis in three-node input-output networks. Journal of Mathematical Biology, 80(4):1163–1185, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01457-x;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01457-x.pdf.

Kon:2020:BCN

[3035] Ryusuke Kon. Bifurcations of cycles in nonlinear semelparous Leslie matrix models. Journal of Mathematical Biology, 80(4):1187–1207, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01459-9;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01459-9.pdf.

Chladna:2020:GDS

[3036] Zuzana Chladná, Jana Kopfová, Dmitrii Rachinskii, and Samiha C. Rouf. Global dynamics of SIR model with switched transmission rate. Journal of Mathematical Biology, 80(4):1209–1233, March 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-019-01460-2;   http://link.springer.com/content/pdf/
10.1007/s00285-019-01460-2.pdf.

Durell:2020:LPN

[3037] Cassandra Durell and Stefan Forcey. Level-1 phylogenetic networks and their balanced minimum evolution polytopes. Journal of Mathematical Biology, 80(5):1235–1263, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01458-w;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01458-w.pdf.

Hendriksen:2020:POC

[3038] Michael Hendriksen and Andrew Francis. A partial order and clustersimilarity metric on rooted phylogenetic trees. Journal of Mathematical Biology, 80(5):1265–1290, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01461-1;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01461-1.pdf.

Clote:2020:RNS

[3039] P. Clote. Are RNA networks scale-free? Journal of Mathematical Biology, 80(5):1291–1321, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01463-z;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01463-z.pdf.

Hening:2020:CEP

[3040] Alexandru Hening and Dang H. Nguyen. The competitive exclusion principle in stochastic environments. Journal of Mathematical Biology, 80(5):1323–1351, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01464-y;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01464-y.pdf.

Chauve:2020:CSG

[3041] Cedric Chauve, Yann Ponty, and Michael Wallner. Counting and sampling gene family evolutionary histories in the duplication-loss and duplication-loss-transfer models. Journal of Mathematical Biology, 80(5):1353–1388, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-019-01465-x;   http://link.springer.
com/content/pdf/10.1007/s00285-019-01465-x.pdf.

Gai:2020:LOR

[3042] Chunyi Gai, David Iron, and Theodore Kolokolnikov. Localized outbreaks in an S-I-R model with diffusion. Journal of Mathematical Biology, 80(5):1389–1411, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01466-1;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01466-1.pdf.

Alhasanat:2020:CPW

[3043] Ahmad Alhasanat and Chunhua Ou. On the conjecture for the pushed wavefront to the diffusive Lotka–Volterra competition model. Journal of Mathematical Biology, 80(5):1413–1422, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/s00285-020-01467-0;   http://link.
springer.com/content/pdf/10.1007/s00285-020-01467-0.pdf.

Bo:2020:EIV

[3044] Wei-Jian Bo, Guo Lin, and Shigui Ruan. The effect of initial values on extinction or persistence in degenerate diffusion competition systems. Journal of Mathematical Biology, 80(5):1423–1458, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01468-z;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01468-z.pdf.

Geiss:2020:BMG

[3045] Manuela Geiß, Marcos E. González Laffitte, Alitzel López Sánchez, Dulce I. Valdivia, Marc Hellmuth, Maribel Hernández Rosales, and Peter F. Stadler. Best match graphs and reconciliation of gene trees with species trees. Journal of Mathematical Biology, 80(5):1459–1495, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01469-y;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01469-y.pdf.

Freund:2020:CMP

[3046] Fabian Freund. Cannings models, population size changes and multiplemerger coalescents. Journal of Mathematical Biology, 80(5):1497–1521, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01470-5;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01470-5.pdf.

Berry:2020:UEE

[3047] Eric Berry, Bree Cummins, Robert R. Nerem, Lauren M. Smith, Steven B. Haase, and Tomas Gedeon. Using extremal events to characterize noisy time series. Journal of Mathematical Biology, 80(5):1523– 1557, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01471-4;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01471-4.pdf.

Du:2020:FKE

[3048] Yihong Du, Bendong Lou, Rui Peng, and Maolin Zhou. The Fisher–KPP equation over simple graphs: varied persistence states in river networks. Journal of Mathematical Biology, 80(5):1559–1616, April 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01474-1;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01474-1.pdf.

Rahman:2020:DUC

[3049] Bootan Rahman, Yuliya N. Kyrychko, and Konstantin B. Blyuss. Dynamics of unidirectionally-coupled ring neural network with discrete and distributed delays. Journal of Mathematical Biology, 80(6):1617– 1653, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01475-0;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01475-0.pdf.

Popovic:2020:DDC

[3050] Lea Popovic and Liam Peuckert. Diffusion dynamics on the coexistence subspace in a stochastic evolutionary game. Journal of Mathematical Biology, 80(6):1655–1682, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01476-z;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01476-z.pdf.

Gross:2020:JDR

[3051] Elizabeth Gross, Heather Harrington, Nicolette Meshkat, and Anne Shiu. Joining and decomposing reaction networks. Journal of Mathematical Biology, 80(6):1683–1731, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01477-y;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01477-y.pdf.

Baker:2020:EM

[3052] Jeremy Baker, Pavel Chigansky, Peter Jagers, and Fima C. Klebaner. On the establishment of a mutant. Journal of Mathematical Biology, 80(6):1733–1757, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01478-x;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01478-x.pdf.

Loy:2020:MPL

[3053] Nadia Loy and Luigi Preziosi. Modelling physical limits of migration by a kinetic model with non-local sensing. Journal of Mathematical Biology, 80(6):1759–1801, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01479-w;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01479-w.pdf.

Gulbudak:2020:ISA

[3054] Hayriye Gulbudak and Cameron J. Browne. Infection severity across scales in multi-strain immuno-epidemiological Dengue model structured by host antibody level. Journal of Mathematical Biology, 80(6):1803– 1843, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01480-3;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01480-3.pdf.

Finkelstein:2020:SSP

[3055] Maxim Finkelstein and James Vaupel. On some stochastic properties of lives lived and left in non-stationary populations. Journal of Mathematical Biology, 80(6):1845–1856, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01481-2;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01481-2.pdf.

Friedman:2020:AMM

[3056] Avner Friedman and King-Yeung Lam. Analysis of a mathematical model of rheumatoid arthritis. Journal of Mathematical Biology, 80(6):1857–1883, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01482-1;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01482-1.pdf.

Seirin-Lee:2020:TSF

[3057] S. Seirin-Lee, T. Sukekawa, T. Nakahara, H. Ishii, and S.-I. Ei. Transitions to slow or fast diffusions provide a general property for inphase or anti-phase polarity in a cell. Journal of Mathematical Biology, 80(6):1885–1917, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01484-z;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01484-z.pdf.

Anderson:2020:TDP

[3058] David F. Anderson, David Schnoerr, and Chaojie Yuan. Time-dependent product-form Poisson distributions for reaction networks with higher order complexes. Journal of Mathematical Biology, 80(6):1919–1951, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01485-y;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01485-y.pdf.

Ito:2020:DOC

[3059] Ryo Ito. Determining the optimal coefficient of the spatially periodic Fisher–KPP equation that minimizes the spreading speed. Journal of Mathematical Biology, 80(6):1953–1970, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/s00285-020-01486-x;   http://link.
springer.com/content/pdf/10.1007/s00285-020-01486-x.pdf.

Francis:2020:MFP

[3060] Andrew R. Francis and Henry P. Wynn. A mean first passage time genome rearrangement distance. Journal of Mathematical Biology, 80(6):1971–1992, May 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01487-w;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01487-w.pdf.

Coronado:2020:MVC

[3061] Tomás M. Coronado, Mareike Fischer, Lina Herbst, Francesc Rosselló, and Kristina Wicke. On the minimum value of the Colless index and the bifurcating trees that achieve it. Journal of Mathematical Biology, 80(7):1993–2054, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01488-9;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01488-9.pdf.

Shaffery:2020:NSR

[3062] Peter Shaffery, Bret D. Elderd, and Vanja Dukic. A note on species richness and the variance of epidemic severity. Journal of Mathematical Biology, 80(7):2055–2074, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01489-8;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01489-8.pdf.

Ersoz:2020:CIC

[3063] Elif Köksal Ersöz, Mathieu Desroches, Antoni Guillamon, John Rinzel, and Joël Tabak. Canard-induced complex oscillations in an excitatory network. Journal of Mathematical Biology, 80(7):2075–2107, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01490-1;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01490-1.pdf.

Seno:2020:SME

[3064] Hiromi Seno. An SIS model for the epidemic dynamics with two phases of the human day-to-day activity. Journal of Mathematical Biology, 80(7):2109–2140, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01491-0;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01491-0.pdf.

Ito:2020:LVA

[3065] Hiroshi C. Ito, Ulf Dieckmann, and Johan A. J. Metz. Lotka–Volterra approximations for evolutionary trait-substitution processes. Journal of Mathematical Biology, 80(7):2141–2226, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/s00285-020-01493-y;   http://link.
springer.com/content/pdf/10.1007/s00285-020-01493-y.pdf.

Flegg:2020:PEP

[3066] Mark B. Flegg, Mario A. Muñoz, Kate Smith-Miles, Wai Shan Yuen, Jennifer A. Flegg, and John G. Carroll. Parameter estimation for a point-source diffusion–decay morphogen model. Journal of Mathematical Biology, 80(7):2227–2255, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01494-x;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01494-x.pdf.

Fu:2020:CCR

[3067] Xiaoming Fu, Quentin Griette, and Pierre Magal. A cell-cell repulsion model on a hyperbolic Keller–Segel equation. Journal of Mathematical Biology, 80(7):2257–2300, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01495-w;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01495-w.pdf.

Lawley:2020:DEF

[3068] Sean D. Lawley. Distribution of extreme first passage times of diffusion. Journal of Mathematical Biology, 80(7):2301–2325, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01496-9;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01496-9.pdf.

Chen:2020:APS

[3069] Shanshan Chen, Junping Shi, Zhisheng Shuai, and Yixiang Wu. Asymptotic profiles of the steady states for an SIS epidemic patch model with asymmetric connectivity matrix. Journal of Mathematical Biology, 80(7):2327–2361, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01497-8;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01497-8.pdf.

Magal:2020:SMH

[3070] P. Magal, G. F. Webb, and Yixiang Wu. A spatial model of honey bee colony collapse due to pesticide contamination of foraging bees. Journal of Mathematical Biology, 80(7):2363–2393, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL http://link.springer.com/article/10.1007/
s00285-020-01498-7;   http://link.springer.com/content/pdf/
10.1007/s00285-020-01498-7.pdf.

Tourigny:2020:DMR

[3071] David S. Tourigny. Dynamic metabolic resource allocation based on the maximum entropy principle. Journal of Mathematical Biology, 80(7):2395–2430, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01499-6;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01499-6.pdf.

Berardo:2020:IBD

[3072] Cecilia Berardo, Stefan Geritz, Mats Gyllenberg, and Gaël Raoul. Interactions between different predator–prey states: a method for the derivation of the functional and numerical response. Journal of Mathematical Biology, 80(7):2431–2468, June 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.springer.com/

article/10.1007/s00285-020-01500-2;   http://link.springer.
com/content/pdf/10.1007/s00285-020-01500-2.pdf.

Akin:2020:CDM

[3073] Elvan Akin, Gülsah Yeni, and Alan S. Perelson. Continuous and discrete modeling of HIV-1 decline on therapy. Journal of Mathematical Biology, 81(1):1–24, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01492-z.

Wilson:2020:OOL

[3074] Dan Wilson. Optimal open-loop desynchronization of neural oscillator populations. Journal of Mathematical Biology, 81(1):25–64, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01501-1.

Hening:2020:HSS

[3075] Alexandru Hening and Ky Quan Tran. Harvesting and seeding of stochastic populations: analysis and numerical approximation. Journal of Mathematical Biology, 81(1):65–112, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-020-01502-0.

Enahoro:2020:LLI

[3076] Iboi Enahoro, Steffen Eikenberry, Abba B. Gumel, Silvie Huijben, and Krijn Paaijmans. Long-lasting insecticidal nets and the quest for malaria eradication: a mathematical modeling approach. Journal of Mathematical Biology, 81(1):113–158, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-020-01503-z.

Dalziel:2020:GAP

[3077] Benjamin D. Dalziel, Enrique Thomann, Jan Medlock, and Patrick De Leenheer. Global analysis of a predator–prey model with variable predator search rate. Journal of Mathematical Biology, 81(1):159–183, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01504-y.

Vigeland:2020:RCP

[3078] Magnus Dehli Vigeland. Relatedness coefficients in pedigrees with inbred founders. Journal of Mathematical Biology, 81(1):185–207, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01505-x.

Ishaku:2020:AOC

[3079] A. Ishaku, A. M. Gazali, S. A. Abdullahi, and N. Hussaini. Analysis and optimal control of an HIV model based on CD4 count. Journal of Mathematical Biology, 81(1):209–241, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-020-01508-8.

Zhang:2020:MDW

[3080] Xianghong Zhang, Qiyong Liu, and Huaiping Zhu. Modeling and dynamics of Wolbachia-infected male releases and mating competition on mosquito control. Journal of Mathematical Biology, 81(1):243–276, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01509-7.

Ferreira:2020:FPM

[3081] Eliza M. Ferreira and Armando G. M. Neves. Fixation probabilities for the Moran process with three or more strategies: general and coupling results. Journal of Mathematical Biology, 81(1):277–314, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01510-0.

Wang:2020:PAT

[3082] Yuanshi Wang, Hong Wu, Yiyang He, Zhihui Wang, and Kun Hu. Population abundance of two-patch competitive systems with asymmetric dispersal. Journal of Mathematical Biology, 81(1):315–341, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01511-z.

Bokes:2020:MDS

[3083] Pavol Bokes, Alessandro Borri, Pasquale Palumbo, and Abhyudai Singh. Mixture distributions in a stochastic gene expression model with delayed feedback: a WKB approximation approach. Journal of Mathematical Biology, 81(1):343–367, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-020-01512-y.

Guo:2020:MRM

[3084] Ting Guo, Zhipeng Qiu, and Libin Rong. Modeling the role of macrophages in HIV persistence during antiretroviral therapy. Journal of Mathematical Biology, 81(1):369–402, July 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-020-01513-x.

Guo:2020:ECC

[3085] Qian Guo, Xiaoqing He, and Wei-Ming Ni. On the effects of carrying capacity and intrinsic growth rate on single and multiple species in spatially heterogeneous environments. Journal of Mathematical Biology, 81 (2):403–433, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01507-9.

Chen:2020:TTP

[3086] Can Chen, Pengde Wang, and Litao Zhang. A two-thresholds policy for a Filippov model in combating influenza. Journal of Mathematical Biology, 81(2):435–461, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01514-w.

Tonello:2020:BAL

[3087] Elisa Tonello and Heike Siebert. Boolean analysis of lateral inhibition. Journal of Mathematical Biology, 81(2):463–486, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-020-01515-9.

Allen:2020:RVS

[3088] E. J. Allen, L. J. S. Allen, and H. L. Smith. On real-valued SDE and nonnegative-valued SDE population models with demographic variability. Journal of Mathematical Biology, 81(2):487–515, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01516-8.

Boahen:2020:MDS

[3089] Frank Boahen and Nicolas Doyon. Modelling dendritic spines with the finite element method, investigating the impact of geometry on electric and calcic responses. Journal of Mathematical Biology, 81(2): 517–547, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01517-7.

Shore:2020:IMC

[3090] Julia A. Shore, Jeremy G. Sumner, and Barbara R. Holland. The impracticalities of multiplicatively-closed codon models: a retreat to linear alternatives. Journal of Mathematical Biology, 81(2):549–573, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01519-5.

Gackou:2020:QAD

[3091] Gorgui Gackou, Arnaud Guillin, and Arnaud Personne. Quantitative approximation of the discrete Moran process by a Wright–Fisher diffusion. Journal of Mathematical Biology, 81(2):575–602, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-020-01520-y.

Schmiester:2020:PMM

[3092] Leonard Schmiester, Daniel Weindl, and Jan Hasenauer. Parameterization of mechanistic models from qualitative data using an efficient optimal scaling approach. Journal of Mathematical Biology, 81(2): 603–623, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01522-w.

Wang:2020:TRN

[3093] Xueying Wang and Xiao-Qiang Zhao. Target reproduction numbers for reaction–diffusion population models. Journal of Mathematical Biology, 81(2):625–647, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01523-9.

Dalwadi:2020:USP

[3094] Mohit P. Dalwadi, Diego Orol, Frederik Walter, Nigel P. Minton, John R. King, and Katalin Kovács. Using singular perturbation theory to determine kinetic parameters in a non-standard coupled enzyme assay. Journal of Mathematical Biology, 81(2):649–690, August 2020. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-020-01524-8.

Ahrabi:2020:MPM

[3095] Sima Sarv Ahrabi and Alireza Momenzadeh. Metaheuristics and Pontryagin’s minimum principle for optimal therapeutic protocols in cancer immunotherapy: a case study and methods comparison. Journal of Mathematical Biology, 81(2):691–723, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-020-01525-7.

Watson:2020:MMG

[3096] Michael G. Watson, Helen M. Byrne, Charlie Macaskill, and Mary R. Myerscough. A multiphase model of growth factor-regulated atherosclerotic cap formation. Journal of Mathematical Biology, 81(2):725–767, August 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01526-6.

Krishnan:2020:MIM

[3097] Jeyashree Krishnan, Reza Torabi, Andreas Schuppert, and Edoardo Di Napoli. A modified Ising model of Barabási–Albert network with genetype spins. Journal of Mathematical Biology, 81(3):769–798, September 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01518-6.

Rhodes:2020:IIM

[3098] Adam Rhodes and Thomas Hillen. Implications of immune-mediated metastatic growth on metastatic dormancy, blow-up, early detection, and treatment. Journal of Mathematical Biology, 81(3):799–843, September 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01521-x.

Jagers:2020:PES

[3099] Peter Jagers and Sergei Zuyev. Populations in environments with a soft carrying capacity are eventually extinct. Journal of Mathematical Biology, 81(3):845–851, September 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-020-01527-5.

Thong:2020:LLR

[3100] David Thong, George Streftaris, and Gavin J. Gibson. Latent likelihood ratio tests for assessing spatial kernels in epidemic models. Journal of Mathematical Biology, 81(3):853–873, September 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-020-01529-3.

Eigentler:2020:IMV

[3101] Lukas Eigentler and Jonathan A. Sherratt. An integrodifference model for vegetation patterns in semi-arid environments with seasonality. Journal of Mathematical Biology, 81(3):875–904, September 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-020-01530-w.

Diekmann:2020:CFD

[3102] Odo Diekmann, Mats Gyllenberg, and Johan A. J. Metz. Correction to: Finite dimensional state representation of physiologically structured populations. Journal of Mathematical Biology, 81(3):905–906, September 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL https://link.springer.com/article/10.
1007/s00285-020-01506-w.

Cai:2020:RID

[3103] Yuhua Cai and Stefan A. H. Geritz. Resident-invader dynamics of similar strategies in fluctuating environments. Journal of Mathematical Biology, 81(4–5):907–959, November 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-020-01532-8.

Linz:2020:CTF

[3104] Simone Linz and Charles Semple. Caterpillars on three and four leaves are sufficient to reconstruct binary normal networks. Journal of Mathematical Biology, 81(4–5):961–980, November 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-020-01533-7.

Ei:2020:CMS

[3105] Shin-Ichiro Ei, Hiroshi Ishii, Makoto Sato, Yoshitaro Tanaka, Miaoxing Wang, and Tetsuo Yasugi. A continuation method for spatially discretized models with nonlocal interactions conserving size and shape of cells and lattices. Journal of Mathematical Biology, 81(4–5):981– 1028, November 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01534-6. See correction [3176].

Manuel:2020:SST

[3106] Cassius Manuel and Arndt von Haeseler. Structure of the space of taboo-free sequences. Journal of Mathematical Biology, 81(4–5):1029– 1057, November 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01535-5.

Berestycki:2020:IRP

[3107] Henri Berestycki, Romain Ducasse, and Luca Rossi. Influence of a road on a population in an ecological niche facing climate change. Journal of Mathematical Biology, 81(4–5):1059–1097, November 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-020-01537-3.

Li:2020:FRM

[3108] Jingwei Li, Hao Ge, and Yunxin Zhang. Fluctuating-rate model with multiple gene states. Journal of Mathematical Biology, 81(4–5):1099– 1141, November 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01538-2.

Oliveira:2020:EES

[3109] Gonçalo Oliveira. Early epidemic spread, percolation and Covid-19. Journal of Mathematical Biology, 81(4–5):1143–1168, November 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01539-1.

Kaihnsa:2020:CAI

[3110] Nidhi Kaihnsa, Yue Ren, Mohab Safey El Din, and Johannes W. R. Martini. Cooperativity, absolute interaction, and algebraic optimization. Journal of Mathematical Biology, 81(4–5):1169–1191, November 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01540-8.

Li:2020:CBZ

[3111] Bingtuan Li, Minghua Zhang, and Bradley Coffman. Can a barrier zone stop invasion of a population? Journal of Mathematical Biology, 81(4–5): 1193–1216, November 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01541-7.

Boca:2020:HAP

[3112] Simina M. Boca, Lucy Huang, and Noah A. Rosenberg. On the heterozygosity of an admixed population. Journal of Mathematical Biology, 81(6–7):1217–1250, December 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-020-01531-9.

Eckardt:2020:NLM

[3113] Maria Eckardt, Kevin J. Painter, Christina Surulescu, and Anna Zhigun. Nonlocal and local models for taxis in cell migration: a rigorous limit procedure. Journal of Mathematical Biology, 81(6–7):1251– 1298, December 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01536-4.

Prasse:2020:TDS

[3114] Bastian Prasse and Piet Van Mieghem. Time-dependent solution of the NIMFA equations around the epidemic threshold. Journal of Mathematical Biology, 81(6–7):1299–1355, December 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-020-01542-6.

Bouvel:2020:CPN

[3115] Mathilde Bouvel, Philippe Gambette, and Marefatollah Mansouri. Counting phylogenetic networks of level 1 and 2. Journal of Mathematical Biology, 81(6–7):1357–1395, December 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-020-01543-5.

Edwards:2020:MME

[3116] David A. Edwards, Brooks Emerick, Anna Georgieva Kondic, Kristian Kiradjiev, Christopher Raymond, and Maxim Zyskin. Mathematical models for the effect of anti-vascular endothelial growth factor on visual acuity. Journal of Mathematical Biology, 81(6–7):1397– 1428, December 2020. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-020-01544-4.

Nie:2020:IAP

[3117] Hua Nie, Biao Wang, and Jianhua Wu. Invasion analysis on a predator– prey system in open advective environments. Journal of Mathematical Biology, 81(6–7):1429–1463, December 2020. CODEN JMBLAJ.

ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-020-01545-3.

Tsai:2020:PDT

[3118] Je-Chiang Tsai and Yu-Yu Weng. Propagation direction of traveling waves for a class of bistable epidemic models. Journal of Mathematical Biology, 81(6–7):1465–1493, December 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-020-01546-2.

Li:2020:TWS

[3119] Yifei Li, Peter van Heijster, Robert Marangell, and Matthew J. Simpson. Travelling wave solutions in a negative nonlinear diffusion–reaction model. Journal of Mathematical Biology, 81(6–7):1495–1522, December 2020. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-020-01547-1.

Brechmann:2021:USM

[3120] Pia Brechmann and Alan D. Rendall. Unbounded solutions of models for glycolysis. Journal of Mathematical Biology, 82(1–2):??, January 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01560-y.

Nagahara:2021:MTP

[3121] Kentaro Nagahara, Yuan Lou, and Eiji Yanagida. Maximizing the total population with logistic growth in a patchy environment. Journal of Mathematical Biology, 82(1–2):??, January 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01565-7.

Lopez:2021:DCI

[3122] Víctor Jiménez López and Eduardo Liz. Destabilization and chaos induced by harvesting: insights from one-dimensional discrete-time models. Journal of Mathematical Biology, 82(1–2):??, January 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01557-7.

VanGorder:2021:TCP

[3123] Robert A. Van Gorder, Václav Klika, and Andrew L. Krause. Turing conditions for pattern forming systems on evolving manifolds. Journal of Mathematical Biology, 82(1–2):??, January 2021. CODEN JMBLAJ.

ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01552-y.

Tourigny:2021:CMR

[3124] David S. Tourigny. Cooperative metabolic resource allocation in spatially-structured systems. Journal of Mathematical Biology, 82(1– 2):??, January 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01558-6.

Favero:2021:DPC

[3125] Martina Favero, Henrik Hult, and Timo Koski. A dual process for the coupled Wright–Fisher diffusion. Journal of Mathematical Biology, 82 (1–2):??, January 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01555-9.

Wang:2021:LVC

[3126] Zhi-An Wang and Jiao Xu. On the Lotka–Volterra competition system with dynamical resources and density-dependent diffusion. Journal of Mathematical Biology, 82(1–2):??, January 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01562-w.

Collienne:2021:CNN

[3127] Lena Collienne and Alex Gavryushkin. Computing nearest neighbour interchange distances between ranked phylogenetic trees. Journal of Mathematical Biology, 82(1–2):??, January 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01567-5.

Kurasov:2021:ASS

[3128] Pavel Kurasov, Delio Mugnolo, and Verena Wolf. Analytic solutions for stochastic hybrid models of gene regulatory networks. Journal of Mathematical Biology, 82(1–2):??, January 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01549-7.

Engwer:2021:EEG

[3129] Christian Engwer and Michael Wenske. Estimating the extent of glioblastoma invasion. Journal of Mathematical Biology, 82(1–2):??, January 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01563-9.

Robin:2021:RMS

[3130] François Robin, Anne Van Gorp, and Amandine Véber. The role of mode switching in a population of actin polymers with constraints. Journal of Mathematical Biology, 82(3):??, February 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01551-z.

Clement:2021:SNM

[3131] Frédérique Clément, Frédérique Robin, and Romain Yvinec. Stochastic nonlinear model for somatic cell population dynamics during ovarian follicle activation. Journal of Mathematical Biology, 82(3):??, February 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01561-x.

Martinez:2021:GDC

[3132] Carlos Martínez and Jean-Luc Gouzé. Global dynamics of the chemostat with overflow metabolism. Journal of Mathematical Biology, 82 (3):??, February 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01566-6.

McAvoy:2021:FPE

[3133] Alex McAvoy and Benjamin Allen. Fixation probabilities in evolutionary dynamics under weak selection. Journal of Mathematical Biology, 82(3):??, February 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01568-4.

Drozhzhin:2021:FOE

[3134] Sergei Drozhzhin, Tatiana Yakushkina, and Alexander S. Bratus. Fitness optimization and evolution of permanent replicator systems. Journal of Mathematical Biology, 82(3):??, February 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01548-8.

Djidjou-Demasse:2021:WHB

[3135] Ramsès Djidjou-Demasse, Samuel Alizon, and Mircea T. Sofonea. Within-host bacterial growth dynamics with both mutation and horizontal gene transfer. Journal of Mathematical Biology, 82(3):??, February 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01571-9.

Salagnac:2021:CSS

[3136] Olivier Salagnac and John Wakeley. The consequences of switching strategies in a two-player iterated survival game. Journal of Mathematical Biology, 82(3):??, February 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01569-3.

Ding:2021:TWN

[3137] Yujie Ding and Bard Ermentrout. Traveling waves in non-local pulsecoupled networks. Journal of Mathematical Biology, 82(3):??, February 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01572-8.

Garba:2021:IGP

[3138] M. K. Garba, T. M. W. Nye, J. Lueg, and S. F. Huckemann. Information geometry for phylogenetic trees. Journal of Mathematical Biology, 82(3):??, February 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01553-x.

Schaller:2021:CCI

[3139] David Schaller, Manuela Geiß, Peter F. Stadler, and Marc Hellmuth. Complete characterization of incorrect orthology assignments in best match graphs. Journal of Mathematical Biology, 82(3):??, February 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/

s00285-021-01564-8. See correction [3166].

Martinson:2021:CAC

[3140] W. Duncan Martinson, Hirokazu Ninomiya, Helen M. Byrne, and Philip K. Maini. Comparative analysis of continuum angiogenesis models. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01570-w.

Kaj:2021:ADD

[3141] Ingemar Kaj, Sylvain Glémin, Daniah Tahir, and Martin Lascoux. Analysis of diversity-dependent species evolution using concepts in population genetics. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01559-5.

Lu:2021:HMC

[3142] Hannah Lu, Kimoon Um, and Daniel M. Tartakovsky. Hybrid models of chemotaxis with application to leukocyte migration. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01581-7.

Shi:2021:ATS

[3143] Yangyang Shi and Hongyong Zhao. Analysis of a two-strain malaria transmission model with spatial heterogeneity and vector-bias. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01577-3.

Borri:2021:QER

[3144] Alessandro Borri and Andrea De Gaetano. A quasi-equilibrium reduced model of pancreatic insulin secretion. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01575-5.

Fu:2021:TWT

[3145] Sheng-Chen Fu, Masayasu Mimura, and Je-Chiang Tsai. Traveling waves for a three-component reaction–diffusion model of farmers and huntergatherers in the Neolithic transition. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01585-3.

Zhang:2021:DPT

[3146] Xue Zhang, Bei Sun, and Yijun Lou. Dynamics of a periodic tick-borne disease model with co-feeding and multiple patches. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01582-6.

Liu:2021:SSS

[3147] Yue Liu, Elisabeth G. Rens, and Leah Edelstein-Keshet. Spots, stripes, and spiral waves in models for static and motile cells. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01550-0.

Oelz:2021:QSS

[3148] Dietmar B. Oelz. Quasi-steady-state reduction of a model for cytoplasmic transport of secretory vesicles in stimulated chromaffin cells. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01583-5.

Zhang:2021:GTD

[3149] Hui Zhang. A game-theoretical dynamic imitation model on networks. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-021-01573-7.

Youlden:2021:TDA

[3150] George H. Youlden, Vito Ricci, Xuan Wang-Kan, Laura J. V. Piddock, Sara Jabbari, and John R. King. Time dependent asymptotic analysis of the gene regulatory network of the AcrAB-TolC efflux pump system in gram-negative bacteria. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01576-4.

Conradi:2021:DER

[3151] Carsten Conradi, Nida Obatake, Anne Shiu, and Xiaoxian Tang. Dynamics of ERK regulation in the processive limit. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01574-6.

Shi:2021:SMD

[3152] Qingyan Shi, Junping Shi, and Hao Wang. Spatial movement with distributed memory. Journal of Mathematical Biology, 82(4):??, March 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01588-0.

Vittadello:2021:NMM

[3153] Sean T. Vittadello, Scott W. McCue, Gency Gunasingh, Nikolas K. Haass, and Matthew J. Simpson. A novel mathematical model of heterogeneous cell proliferation. Journal of Mathematical Biology, 82(5):

??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL https://link.springer.com/article/10.
1007/s00285-021-01580-8.

Pybus:2021:RBM

[3154] Hannah J. Pybus, Amanda L. Tatler, Lowell T. Edgar, Reuben D. O’Dea, and Bindi S. Brook. Reduced biomechanical models for precisioncut lung-slice stretching experiments. Journal of Mathematical Biology, 82(5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01578-2.

Bouin:2021:EAD

[3155] Emeric Bouin, Guillaume Legendre, Yuan Lou, and Nichole Slover. Evolution of anisotropic diffusion in two-dimensional heterogeneous environments. Journal of Mathematical Biology, 82(5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01579-1.

Pijpers:2021:NPM

[3156] Frank P. Pijpers. A non-parametric method for determining epidemiological reproduction numbers. Journal of Mathematical Biology, 82 (5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL https://link.springer.com/article/10.
1007/s00285-021-01590-6.

Izuhara:2021:FSF

[3157] Hirofumi Izuhara, Harunori Monobe, and Chang-Hong Wu. The formation of spreading front: the singular limit of three-component reaction–diffusion models. Journal of Mathematical Biology, 82(5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01591-5.

Andrade-Restrepo:2021:RDM

[3158] Martin Andrade-Restrepo, Ionel Sorin Ciuperca, Paul Lemarre, Laurent Pujo-Menjouet, and Léon Matar Tine. A reaction–diffusion model of spatial propagation of Aβ oligomers in early stage Alzheimer’s disease. Journal of Mathematical Biology, 82(5):??, April 2021. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-021-01593-3.

Huber:2021:RHN

[3159] Katharina T. Huber, Simone Linz, and Vincent Moulton. The rigid hybrid number for two phylogenetic trees. Journal of Mathematical Biology, 82(5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01594-2.

Alberti:2021:SMR

[3160] F. Alberti, E. Baake, I. Letter, and S. Martínez. Solving the migrationrecombination equation from a genealogical point of view. Journal of Mathematical Biology, 82(5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01584-4.

Feng:2021:DTA

[3161] Tao Feng, Daniel Charbonneau, Zhipeng Qiu, and Yun Kang. Dynamics of task allocation in social insect colonies: scaling effects of colony size versus work activities. Journal of Mathematical Biology, 82(5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01589-z.

Overton:2021:EBH

[3162] Christopher E. Overton and Kieran J. Sharkey. Evolutionary bethedging in structured populations. Journal of Mathematical Biology, 82(5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01597-z.

Argasinski:2021:TRD

[3163] K. Argasinski and M. Broom. Towards a replicator dynamics model of age structured populations. Journal of Mathematical Biology, 82 (5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL https://link.springer.com/article/10.
1007/s00285-021-01592-4.

Guo:2021:AAS

[3164] Zhong-Kai Guo, Hong Xiang, and Hai-Feng Huo. Analysis of an age-structured tuberculosis model with treatment and relapse. Journal of Mathematical Biology, 82(5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01595-1.

Gomez-Corral:2021:TDV

[3165] A. Gómez-Corral, M. López-García, and M. T. Rodríguez-Bernal. On time-discretized versions of the stochastic SIS epidemic model: a comparative analysis. Journal of Mathematical Biology, 82(5):??, April 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01598-y.

Schaller:2021:CBM

[3166] David Schaller, Manuela Geiß, Edgar Chávez, Marcos González Laffitte, Alitzel López Sánchez, Bärbel M. R. Stadler, Dulce I. Valdivia, Marc Hellmuth, Maribel Hernández Rosales, and Peter F. Stadler. Corrigendum to “Best match graphs”. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01601-6. See [3139].

Allen:2021:SMI

[3167] Linda J. S. Allen and Xueying Wang. Stochastic models of infectious diseases in a periodic environment with application to cholera epidemics. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01603-4.

Kumar:2021:MMG

[3168] Pawan Kumar, Jing Li, and Christina Surulescu. Multiscale modeling of glioma pseudopalisades: contributions from the tumor microenvironment. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01599-x.

Reichenbach:2021:SPN

[3169] Matt Reichenbach, Richard Rebarber, and Brigitte Tenhumberg. Spectral properties of a non-compact operator in ecology. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01600-7.

Guo:2021:DNH

[3170] Ting Guo, Zhipeng Qiu, Mingwang Shen, and Libin Rong. Dynamics of a new HIV model with the activation status of infected cells. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ.

ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01604-3.

Besse:2021:DES

[3171] Christophe Besse and Grégory Faye. Dynamics of epidemic spreading on connected graphs. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01602-5.

Blath:2021:STS

[3172] Jochen Blath, Eugenio Buzzoni, Adrián González Casanova, and Maite Wilke Berenguer. Separation of timescales for the seed bank diffusion and its jump-diffusion limit. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01596-0.

Schecter:2021:GSP

[3173] Stephen Schecter. Geometric singular perturbation theory analysis of an epidemic model with spontaneous human behavioral change. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01605-2.

Sommariva:2021:GLF

[3174] Sara Sommariva, Giacomo Caviglia, and Michele Piana. Gain and loss of function mutations in biological chemical reaction networks: a mathematical model with application to colorectal cancer cells. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01607-0.

Hening:2021:GTC

[3175] Alexandru Hening, Dang H. Nguyen, and Peter Chesson. A general theory of coexistence and extinction for stochastic ecological communities. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01606-1.

Ei:2021:CCM

[3176] Shin-Ichiro Ei, Hiroshi Ishii, Makoto Sato, Yoshitaro Tanaka, Miaoxing Wang, and Tetsuo Yasugi. Correction to: A continuation method for spatially discretized models with nonlocal interactions conserving size and shape of cells and lattices. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01610-5. See [3105].

Bonnet:2021:LFM

[3177] Céline Bonnet and Sylvie Méléard. Large fluctuations in multi-scale modeling for rest hematopoiesis. Journal of Mathematical Biology, 82(6):??, May 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01611-4.

Albi:2021:CUD

[3178] Giacomo Albi, Lorenzo Pareschi, and Mattia Zanella. Control with uncertain data of socially structured compartmental epidemic models. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01617-y.

Feng:2021:SFL

[3179] Yuanyuan Feng, Gautam Iyer, and Lei Li. Scheduling fixed length quarantines to minimize the total number of fatalities during an epidemic. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-021-01615-0.

FitzGerald:2021:SSS

[3180] Cody E. FitzGerald and James P. Keener. A systematic search for switchlike behavior in type II toxin-antitoxin systems. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01608-z.

Gulbudak:2021:DMP

[3181] Hayriye Gulbudak, Paul L. Salceanu, and Gail S. K. Wolkowicz. A delay model for persistent viral infections in replicating cells. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01612-3.

Hening:2021:SDP

[3182] Alexandru Hening and Yao Li. Stationary distributions of persistent ecological systems. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01613-2.

Hoessly:2021:SDD

[3183] Linard Hoessly. Stationary distributions via decomposition of stochastic reaction networks. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01620-3.

Jarvis:2021:CMG

[3184] Peter D. Jarvis and Jeremy G. Sumner. Correction to: Matrix group structure and Markov invariants in the strand symmetric phylogenetic substitution model. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01625-y.

Morita:2021:LTB

[3185] Yoshihisa Morita and Sungrim Seirin-Lee. Long time behavior and stable patterns in high-dimensional polarity models of asymmetric cell division. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01619-w.

Richter:2021:SAT

[3186] Hendrik Richter. Spectral analysis of transient amplifiers for death-birth updating constructed from regular graphs. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01609-y.

Tricot:2021:MVD

[3187] A. Tricot, I. M. Sokolov, and D. Holcman. Modeling the voltage distribution in a non-locally but globally electroneutral confined electrolyte medium: applications for nanophysiology. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01618-x.

Wang:2021:SIH

[3188] Yangyang Wang, Zhengyuan Huang, and Martin Golubitsky. The structure of infinitesimal homeostasis in input-output networks. Journal of Mathematical Biology, 82(7):??, June 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01614-1.

Aftalion:2021:PMC

[3189] Amandine Aftalion and Emmanuel Trélat. Pace and motor control optimization for a runner. Journal of Mathematical Biology, 83(1): ??, July 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01632-z.

Dimarco:2021:KME

[3190] G. Dimarco, B. Perthame, and M. Zanella. Kinetic models for epidemic dynamics with social heterogeneity. Journal of Mathematical Biology, 83(1):??, July 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01630-1.

Friedman:2021:AMM

[3191] Avner Friedman and King-Yeung Lam. Analysis of a mathematical model of immune response to fungal infection. Journal of Mathematical Biology, 83(1):??, July 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01633-y.

Hurry:2021:MIB

[3192] Christian John Hurry, Alexander Mozeika, and Alessia Annibale. Modelling the interplay between the CD4^{+} /CD8^{+} T-cell ratio and the expression of MHC-I in tumours. Journal of Mathematical Biology, 83(1):??, July 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01622-1.

Jagers:2021:APE

[3193] Peter Jagers and Sergei Zuyev. Amendment to: populations in environments with a soft carrying capacity are eventually extinct. Journal of Mathematical Biology, 83(1):??, July 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01624-z.

Kang:2021:MAA

[3194] Hao Kang and Shigui Ruan. Mathematical analysis on an age-structured SIS epidemic model with nonlocal diffusion. Journal of Mathematical Biology, 83(1):??, July 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01634-x.

Ketcheson:2021:OCS

[3195] David I. Ketcheson. Optimal control of an SIR epidemic through finitetime non-pharmaceutical intervention. Journal of Mathematical Biology, 83(1):??, July 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01628-9.

Schaller:2021:IIH

[3196] David Schaller, Manuel Lafond, and Marc Hellmuth. Indirect identification of horizontal gene transfer. Journal of Mathematical Biology, 83(1):??, July 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01631-0.

Scheel:2021:SGS

[3197] Arnd Scheel, Angela Stevens, and Christoph Tenbrock. Signaling gradients in surface dynamics as basis for planarian regeneration. Journal of Mathematical Biology, 83(1):??, July 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01627-w.

Williamson:2021:PEF

[3198] Daniel E. Williamson, Erik Sahai, and John R. King. Parameter estimation in fluorescence recovery after photobleaching: quantitative analysis of protein binding reactions and diffusion. Journal of Mathematical Biology, 83(1):??, July 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01616-z.

Bidari:2021:HGS

[3199] Subekshya Bidari and Zachary P Kilpatrick. Hive geometry shapes the recruitment rate of honeybee colonies. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01644-9.

Blath:2021:BPM

[3200] Jochen Blath, Felix Hermann, and Martin Slowik. A branching process model for dormancy and seed banks in randomly fluctuating environments. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01639-6.

Croix:2021:BID

[3201] Jean-Charles Croix, Nicolas Durrande, and Mauricio A. Alvarez. Bayesian inversion of a diffusion model with application to biology. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01621-2.

Hamel:2021:AHE

[3202] François Hamel, Florian Lavigne, and Lionel Roques. Adaptation in a heterogeneous environment I: persistence versus extinction. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01637-8.

Hong:2021:EMS

[3203] Won Eui Hong and Robert L. Pego. Exclusion and multiplicity for stable communities in Lotka–Volterra systems. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01638-7.

Kopfova:2021:DSM

[3204] Jana Kopfová, Petra Nábelková, and Samiha C. Rouf. Dynamics of SIR model with vaccination and heterogeneous behavioral response of individuals modeled by the Preisach operator. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01629-8.

Molina:2021:IES

[3205] Chai Molina and David J. D. Earn. On inferring evolutionary stability in finite populations using infinite population models. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01636-9.

Nadin:2021:TMO

[3206] Grégoire Nadin, Eric Ogier-Denis, and Hatem Zaag. A Turing mechanism in order to explain the patchy nature of Crohn’s disease. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01635-w.

Phan:2021:DSM

[3207] Tuan Anh Phan, Hai Dang Nguyen, and Jianjun Paul Tian. Deterministic and stochastic modeling for PDGF-driven gliomas reveals a classification of gliomas. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01647-6.

Thieme:2021:DTP

[3208] Horst R. Thieme. Discrete-time population dynamics of spatially distributed semelparous two-sex populations. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01649-4.

Wang:2021:DVM

[3209] Feng-Bin Wang and Chang-Yuan Cheng. A diffusive virus model with a fixed intracellular delay and combined drug treatments. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01646-7.

Zhang:2021:MAG

[3210] Jimin Zhang, Junping Shi, and Xiaoyuan Chang. A model of algal growth depending on nutrients and inorganic carbon in a poorly mixed water column. Journal of Mathematical Biology, 83(2):??, August 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01640-z.

Aggarwal:2021:PIE

[3211] Manu Aggarwal, N. G. Cogan, and Owen L. Lewis. Physiological insights into electrodiffusive maintenance of gastric mucus through sensitivity analysis and simulations. Journal of Mathematical Biology, 83 (3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01643-w.

Almeida:2021:TIS

[3212] Luis Almeida, Gissell Estrada-Rodriguez, and Francois Vallette. Treatment-induced shrinking of tumour aggregates: a nonlinear volumefilling chemotactic approach. Journal of Mathematical Biology, 83(3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01642-x.

Ardiyansyah:2021:MSM

[3213] Muhammad Ardiyansyah, Dimitra Kosta, and Kaie Kubjas. The modelspecific Markov embedding problem for symmetric group-based models. Journal of Mathematical Biology, 83(3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01656-5.

Egberts:2021:SOD

[3214] Ginger Egberts, Fred Vermolen, and Paul van Zuijlen. Stability of a onedimensional morphoelastic model for post-burn contraction. Journal of Mathematical Biology, 83(3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01648-5.

Gross:2021:DLP

[3215] Elizabeth Gross, Leo van Iersel, and Yukihiro Murakami. Distinguishing level-1 phylogenetic networks on the basis of data generated by Markov processes. Journal of Mathematical Biology, 83(3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01653-8.

Hara:2021:WMD

[3216] Akane Hara and Akiko Satake. Why meals during resting time cause fat accumulation in mammals? Mathematical modeling of circadian regulation on glucose metabolism. Journal of Mathematical Biology, 83 (3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01645-8.

Khan:2021:IBU

[3217] Kamruzzaman Khan, Shuang Liu, and Yihong Du. Invasive behaviour under competition via a free boundary model: a numerical approach. Journal of Mathematical Biology, 83(3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01641-y.

Semple:2021:TEO

[3218] Charles Semple and Gerry Toft. Trinets encode orchard phylogenetic networks. Journal of Mathematical Biology, 83(3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01654-7.

Streipert:2021:ADP

[3219] Sabrina H. Streipert and Gail S. K. Wolkowicz. An alternative delayed population growth difference equation model. Journal of Mathematical Biology, 83(3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01652-9.

Xue:2021:IBM

[3220] Xiaoru Xue and Min Tang. Individual based models exhibiting Lévyflight type movement induced by intracellular noise. Journal of Mathematical Biology, 83(3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01651-w.

Zhang:2021:ITP

[3221] Wenjing Zhang, Leif Ellingson, and Jane Heffernan. An investigation of tuberculosis progression revealing the role of macrophages apoptosis via sensitivity and bifurcation analysis. Journal of Mathematical Biology, 83(3):??, September 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01655-6.

Bolzoni:2021:OCS

[3222] Luca Bolzoni, Rossella Della Marca, and Maria Groppi. On the optimal control of SIR model with Erlang-distributed infectious period: isolation strategies. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01668-1.

Calleri:2021:CTS

[3223] Fabiana Calleri, Giovanni Nastasi, and Vittorio Romano. Continuoustime stochastic processes for the spread of COVID-19 disease simulated via a Monte Carlo approach and comparison with deterministic models. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01657-4.

Castillo-Laborde:2021:AET

[3224] Carla Castillo-Laborde, Taco de Wolff, and Héctor Ramírez C. Assessment of event-triggered policies of nonpharmaceutical interventions based on epidemiological indicators. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01669-0.

Choi:2021:AJD

[3225] Kwok Pui Choi, Gursharn Kaur, and Taoyang Wu. On asymptotic joint distributions of cherries and pitchforks for random phylogenetic trees. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-021-01667-2.

Cristiani:2021:ALA

[3226] Emiliano Cristiani, Marta Menci, and Léonard Brafman. An all-leader agent-based model for turning and flocking birds. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01675-2.

Geng:2021:CCS

[3227] Yunfeng Geng and Frithjof Lutscher. Competitive coexistence of seasonal breeders. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01665-4.

Jardon-Kojakhmetov:2021:GAS

[3228] Hildeberto Jardón-Kojakhmetov, Christian Kuehn, and Mattia Sensi. A geometric analysis of the SIRS epidemiological model on a homogeneous network. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01664-5.

Jerez:2021:ART

[3229] Silvia Jerez, Emilene Pliego, and Anna K. Miller. Antigen receptor therapy in bone metastasis via optimal control for different human life stages. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-021-01673-4.

Neto:2021:DMB

[3230] Joana Pinheiro Neto, Irina Alho, and Susana Vinga. Dynamic modeling of bone remodeling, osteolytic metastasis and PK/PD therapy: introducing variable order derivatives as a simplification technique. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01666-3.

Papanicolaou:2021:BSS

[3231] Vassilis G. Papanicolaou. A binary search scheme for determining all contaminated specimens. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01663-6.

Shu:2021:TDN

[3232] Hongying Shu, Zongwei Ma, and Xiang-Sheng Wang. Threshold dynamics of a nonlocal and delayed cholera model in a spatially heterogeneous environment. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01672-5.

Zaidi:2021:OBO

[3233] Faheem Zaidi, Alona Ben-Tal, and Mick Roberts. Is our breathing optimal? Solving a piecewise linear model with constraints. Journal of Mathematical Biology, 83(4):??, October 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01661-8.

Albert:2021:DRP

[3234] Jaroslav Albert. Dimensionality reduction via path integration for computing mRNA distributions. Journal of Mathematical Biology, 83 (5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01683-2.

Bedekar:2021:RDM

[3235] Prajakta Bedekar, Ilya Timofeyev, and Misha Perepelitsa. Reactiondiffusion models for morphological patterning of hESCs. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01674-3.

Bienvenu:2021:RSS

[3236] François Bienvenu, Gabriel Cardona, and Celine Scornavacca. Revisiting Shao and Sokal’s B_{2} index of phylogenetic balance. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01662-7.

Brocchieri:2021:EDD

[3237] Elisabetta Brocchieri, Lucilla Corrias, and Yong-Jung Kim. Evolution of dietary diversity and a starvation driven cross-diffusion system as its singular limit. Journal of Mathematical Biology, 83(5): ??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01679-y.

Capera-Aragones:2021:DEM

[3238] Pau Capera-Aragones, Eric Foxall, and Rebecca C. Tyson. Differential equation model for central-place foragers with memory: implications for bumble bee crop pollination. Journal of Mathematical Biology, 83 (5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01676-1.

Caputo:2021:ATN

[3239] Jean-Guy Caputo, Valérie Girardin, and Quentin Noguès. Analysis of trophic networks: an optimisation approach. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01682-3.

Collienne:2021:DCT

[3240] Lena Collienne, Kieran Elmes, and Alex Gavryushkin. Discrete coalescent trees. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01685-0.

Davies:2021:ICD

[3241] Colleen M. Davies and Hao Wang. Incorporating carbon dioxide into a stoichiometric producer-grazer model. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01658-3.

Lefevre:2021:CBE

[3242] Claude Lefèvre, Philippe Picard, and Sergey Utev. A chain binomial epidemic with asymptomatics motivated by COVID-19 modelling. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01680-5.

Maiorana:2021:EIM

[3243] Andrea Maiorana, Marco Meneghelli, and Mario Resnati. Effectiveness of isolation measures with app support to contain COVID-19 epidemics: a parametric approach. Journal of Mathematical Biology, 83 (5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01660-9.

Miles:2021:ESJ

[3244] Jason G. Miles and Nicholas A. Battista. Exploring the sensitivity in jellyfish locomotion under variations in scale, frequency, and duty cycle. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01678-z.

Vastola:2021:SCM

[3245] John J. Vastola. Solving the chemical master equation for monomolecular reaction systems and beyond: a Doi–Peliti path integral view. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01670-7.

Ventre:2021:RSM

[3246] Elias Ventre, Thibault Espinasse, and Olivier Gandrillon. Reduction of a stochastic model of gene expression: Lagrangian dynamics gives access to basins of attraction as cell types and metastabilty. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01684-1.

Xu:2021:UKM

[3247] Yian Xu, Masoud Asadi-Zeydabadi, and Orrin Shindell. Universality in kinetic models of circadian rhythms in Arabidopsis thaliana. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01677-0.

Yoon:2021:TMC

[3248] Nara Yoon, Nikhil Krishnan, and Jacob Scott. Theoretical modeling of collaterally sensitive drug cycles: shaping heterogeneity to allow adaptive therapy. Journal of Mathematical Biology, 83(5):??, November 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01671-6.

Boenkost:2021:HFC

[3249] Florin Boenkost, Adrián González Casanova, and Anton Wakolbinger. Haldane’s formula in Cannings models: the case of moderately strong selection. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01698-9.

Cantin:2021:MMF

[3250] Guillaume Cantin, Arnaud Ducrot, and Beatriz M. Funatsu. Mathematical modeling of forest ecosystems by a reaction-diffusion-advection system: impacts of climate change and deforestation. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01696-x.

Cloez:2021:LTB

[3251] Bertrand Cloez, Benoîte de Saporta, and Tristan Roget. Long-time behavior and Darwinian optimality for an asymmetric size-structured branching process. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01695-y.

Dawi:2021:SBD

[3252] Malik A. Dawi and Jose J. Muñoz. Stability bounds of a delay viscoelastic rheological model with substrate friction. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01699-8.

Elbasha:2021:VHI

[3253] Elamin H. Elbasha and Abba B. Gumel. Vaccination and herd immunity thresholds in heterogeneous populations. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01686-z.

Girardin:2021:DFC

[3254] Léo Girardin and Florence Débarre. Demographic feedbacks can hamper the spatial spread of a gene drive. Journal of Mathematical Biology, 83 (6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01702-2.

Guilberteau:2021:MBB

[3255] Jules Guilberteau, Camille Pouchol, and Nastassia Pouradier Duteil. Monostability and bistability of biological switches. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01687-y.

Hobolth:2021:MPT

[3256] Asger Hobolth, Mogens Bladt, and Lars Nørvang Andersen. Multivariate phase-type theory for the site frequency spectrum. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01689-w.

Kozlov:2021:GSA

[3257] Vladimir Kozlov, Sonja Radosavljevic, and Uno Wennergren. Global stability of an age-structured population model on several temporally variable patches. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01701-3.

Le:2021:IID

[3258] Adam Le, Aaron A. King, and Pejman Rohani. The impact of infectionderived immunity on disease dynamics. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01681-4.

Lorenzi:2021:EDS

[3259] Tommaso Lorenzi, Andrea Pugliese, and Agnese Zardini. Evolutionary dynamics in an SI epidemic model with phenotype-structured susceptible compartment. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01703-1.

Lotstedt:2021:DCM

[3260] Per Lötstedt. Derivation of continuum models from discrete models of mechanical forces in cell populations. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01697-w.

Ma:2021:SEE

[3261] Zhihui Ma, Shuyan Han, and Shenghua Li. A stochastic ecoepidemiological system with patchy structure and transport-related infection. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

https://link.springer.com/article/10.1007/s00285-021-01688-
x.

Martinez:2021:MII

[3262] Ricardo Martínez and Joaquín Sánchez-Soriano. Mathematical indices for the influence of risk factors on the lethality of a disease. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01700-4.

Pasquini:2021:CHM

[3263] Mirko Pasquini and David Angeli. On convergence for hybrid models of gene regulatory networks under polytopic uncertainties: a Lyapunov approach. Journal of Mathematical Biology, 83(6–7):??, December 2021. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01690-3.

Baccelli:2022:CFE

[3264] François Baccelli and Nithin Ramesan. A computational framework for evaluating the role of mobility on the propagation of epidemics on point processes. Journal of Mathematical Biology, 84(1–2):??, January 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-021-01692-1.

Busse:2022:LAS

[3265] Jan-Erik Busse, Sílvia Cuadrado, and Anna Marciniak-Czochra. Local asymptotic stability of a system of integro-differential equations describing clonal evolution of a self-renewing cell population under mutation. Journal of Mathematical Biology, 84(1–2):??, January 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01708-w.

Delacour:2022:MMP

[3266] Julia Delacour, Marie Doumic, and Gabriele Zaffagnini. A mathematical model of p62-ubiquitin aggregates in autophagy. Journal of Mathematical Biology, 84(1–2):??, January 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01659-2.

Fai:2022:GAS

[3267] Thomas G. Fai and Youngmin Park. Global asymptotic stability of the active disassembly model of flagellar length control. Journal of Mathematical Biology, 84(1–2):??, January 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01709-9.

Gedeon:2022:ODS

[3268] Tomás Gedeon, Antony R. Humphries, and Zhao Wang. Operon dynamics with state dependent transcription and/or translation delays. Journal of Mathematical Biology, 84(1–2):??, January 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01693-0.

DeLuca:2022:VCP

[3269] L. De Luca, A. Ninno, and M. Ponsiglione. Vectorial crystallization problems and collective behavior. Journal of Mathematical Biology, 84 (1–2):??, January 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-021-01704-0.

Renardy:2022:SIA

[3270] Marissa Renardy, Denise Kirschner, and Marisa Eisenberg. Structural identifiability analysis of age-structured PDE epidemic models. Journal of Mathematical Biology, 84(1–2):??, January 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01711-1.

Ridgway:2022:QSI

[3271] Wesley Ridgway, Michael J. Ward, and Brian T. Wetton. Quorumsensing induced transitions between bistable steady-states for a cell-bulk ODE–PDE model with Lux intracellular kinetics. Journal of Mathematical Biology, 84(1–2):??, January 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-021-01705-z.

Sen:2022:BAP

[3272] Deeptajyoti Sen, Saktipada Ghorai, and Andrew Morozov. Bifurcation analysis of the predator–prey model with the Allee effect in the predator. Journal of Mathematical Biology, 84(1–2):??, January 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-021-01707-x.

Zhang:2022:DST

[3273] Juping Zhang, Wenhui Hao, and Zhen Jin. The dynamics of sexually transmitted diseases with men who have sex with men. Journal of Mathematical Biology, 84(1–2):??, January 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01694-z.

Alfaro:2022:STD

[3274] Matthieu Alfaro, Quentin Griette, and Benoît Sarels. The spatiotemporal dynamics of interacting genetic incompatibilities. Part I: the case of stacked underdominant clines. Journal of Mathematical Biology, 84(3):??, February 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01722-6.

Azhdari:2022:EVT

[3275] Ebrahim Azhdari and Jahed Naghipoor. The effect of viscoelasticity of the tissue on the magneto-responsive drug delivery system. Journal of Mathematical Biology, 84(3):??, February 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01710-2.

Chappelle:2022:OTT

[3276] George Chappelle, Alan Hastings, and Martin Rasmussen. Occupancy times for time-dependent stage-structured models. Journal of Mathematical Biology, 84(3):??, February 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01713-7.

Dekens:2022:EDC

[3277] Léonard Dekens. Evolutionary dynamics of complex traits in sexual populations in a heterogeneous environment: how normal? Journal of Mathematical Biology, 84(3):??, February 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01712-0.

Pan:2022:PDA

[3278] Yingli Pan. Propagation dynamics for an age-structured population model in time-space periodic habitat. Journal of Mathematical Biology, 84(3):??, February 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01721-7.

Pargaei:2022:CEM

[3279] Meena Pargaei, B. V. Rathish Kumar, and Simone Scacchi. Cardiac electro-mechanical activity in a deforming human cardiac tissue: modeling, existence-uniqueness, finite element computation and application to multiple ischemic disease. Journal of Mathematical Biology, 84(3):??, February 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01717-3.

Salmaniw:2022:GDD

[3280] Yurij Salmaniw, Zhongwei Shen, and Hao Wang. Global dynamics of a diffusive competition model with habitat degradation. Journal of Mathematical Biology, 84(3):??, February 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01720-8.

Videla:2022:SSP

[3281] Leonardo Videla. Strong stochastic persistence of some Lévy-driven Lotka–Volterra systems. Journal of Mathematical Biology, 84(3):??, February 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01714-6.

Wang:2022:SPD

[3282] Cuihua Wang, Sanling Yuan, and Hao Wang. Spatiotemporal patterns of a diffusive prey-predator model with spatial memory and pregnancy period in an intimidatory environment. Journal of Mathematical Biology, 84(3):??, February 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01716-4.

Yu:2022:DSM

[3283] Jianshe Yu and Jia Li. A delay suppression model with sterile mosquitoes release period equal to wild larvae maturation period. Journal of Mathematical Biology, 84(3):??, February 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01718-2.

Ho:2022:ASR

[3284] Lam Si Tung Ho and Edward Susko. Ancestral state reconstruction with large numbers of sequences and edge-length estimation. Journal of Mathematical Biology, 84(4):??, March 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01715-5.

Huang:2022:CIH

[3285] Zhengyuan Huang and Martin Golubitsky. Classification of infinitesimal homeostasis in four-node input-output networks. Journal of Mathematical Biology, 84(4):??, March 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01727-1.

Kuwamura:2022:OBS

[3286] Masataka Kuwamura, Hirofumi Izuhara, and Shin ichiro Ei. Oscillations and bifurcation structure of reaction-diffusion model for cell polarity formation. Journal of Mathematical Biology, 84(4):??, March 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01723-5.

Lang:2022:UMM

[3287] John C. Lang. Use of mathematical modelling to assess respiratory syncytial virus epidemiology and interventions: a literature review. Journal of Mathematical Biology, 84(4):??, March 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-021-01706-y.

Roy:2022:FPF

[3288] Souvik Roy, Zui Pan, and Suvra Pal. A Fokker–Planck feedback control framework for optimal personalized therapies in colon cancerinduced angiogenesis. Journal of Mathematical Biology, 84(4):??, March 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01725-3.

Vaghi:2022:MSM

[3289] Cristina Vaghi, Raphaëlle Fanciullino, and Clair Poignard. Macro-scale models for fluid flow in tumour tissues: impact of microstructure properties. Journal of Mathematical Biology, 84(4):??, March 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01719-1.

Yu:2022:HCL

[3290] Zhuojun Yu and Peter J. Thomas. A homeostasis criterion for limit cycle systems based on infinitesimal shape response curves. Journal of Mathematical Biology, 84(4):??, March 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01724-4.

Alfaro:2022:MSH

[3291] Matthieu Alfaro, Thomas Giletti, and Hyowon Seo. On the modelling of spatially heterogeneous nonlocal diffusion: deciding factors and preferential position of individuals. Journal of Mathematical Biology, 84(5):??, April 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01738-y.

Allman:2022:ISN

[3292] Elizabeth S. Allman, Hector Baños, and John A. Rhodes. Identifiability of species network topologies from genomic sequences using the logDet distance. Journal of Mathematical Biology, 84(5):??, April 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01734-2.

Cheek:2022:CTM

[3293] David Cheek. The coalescent tree of a Markov branching process with generalised logistic growth. Journal of Mathematical Biology, 84(5):??, April 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01735-1.

Dasarathy:2022:SFT

[3294] Gautam Dasarathy, Elchanan Mossel, and Sebastien Roch. A stochastic Farris transform for genetic data under the multispecies coalescent with applications to data requirements. Journal of Mathematical Biology, 84(5):??, April 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01731-5.

Hamed:2022:PFM

[3295] Mohammad Abu Hamed and Alexander A. Nepomnyashchy. Phase field model for cell spreading dynamics. Journal of Mathematical Biology, 84(5):??, April 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01732-4.

Hubert:2022:ILT

[3296] Emma Hubert, Thibaut Mastrolia, and Xavier Warin. Incentives, lockdown, and testing: from Thucydides’ analysis to the COVID-19 pandemic. Journal of Mathematical Biology, 84(5):??, April 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01736-0.

Lin:2022:PGC

[3297] Chiu-Ju Lin, Ting-Hao Hsu, and Gail S. K. Wolkowicz. Population growth and competition models with decay and competition consistent delay. Journal of Mathematical Biology, 84(5):??, April 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01741-3.

Reoch:2022:MMC

[3298] J. R. Reoch, Y. M. Stokes, and J. E. F. Green. A mathematical model for cell-induced gel contraction incorporating osmotic effects. Journal of Mathematical Biology, 84(5):??, April 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01730-6.

Wang:2022:RAD

[3299] Xueying Wang, Ruiwen Wu, and Xiao-Qiang Zhao. A reactionadvection-diffusion model of cholera epidemics with seasonality and human behavior change. Journal of Mathematical Biology, 84(5):??, April 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01733-3.

Alonzo:2022:SBG

[3300] Flavien Alonzo, Aurelien A. Serandour, and Mazen Saad. Simulating the behaviour of glioblastoma multiforme based on patient MRI during treatments. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01747-x.

Coron:2022:PBM

[3301] Camille Coron and Yves Le Jan. Pedigree in the biparental Moran model. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01752-0.

Demers:2022:IVE

[3302] Jeffery Demers, Suzanne L. Robertson, and William F. Fagan. Implicit versus explicit vector management strategies in models for vectorborne disease epidemiology. Journal of Mathematical Biology, 84(6): ??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01742-2.

Gao:2022:IVH

[3303] Shasha Gao, Maia Martcheva, and Libin Rong. The impact of vaccination on human papillomavirus infection with disassortative geographical mixing: a two-patch modeling study. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01745-z.

Garnier:2022:LHT

[3304] Jimmy Garnier and Pierre Lafontaine. Life history traits and dispersal shape neutral genetic diversity in metapopulations. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01749-9.

Hening:2022:ERS

[3305] Alexandru Hening, Ky Quan Tran, and Sergiu C. Ungureanu. The effects of random and seasonal environmental fluctuations on optimal harvesting and stocking. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01750-2.

Jing:2022:SDS

[3306] Xiaojie Jing, Guirong Liu, and Zhen Jin. Stochastic dynamics of an SIS epidemic on networks. Journal of Mathematical Biology, 84(6): ??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01754-y.

Kong:2022:CEP

[3307] Sungsik Kong, Joan Carles Pons, and Kristina Wicke. Classes of explicit phylogenetic networks and their biological and mathematical significance. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01746-y.

Kreten:2022:TWF

[3308] Florian Kreten. Traveling waves of an FKPP-type model for selforganized growth. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

https://link.springer.com/article/10.1007/s00285-022-01753-
z.

Lawley:2022:SFD

[3309] Sean D. Lawley, H. Frederik Nijhout, and Michael C. Reed. Spiracular fluttering decouples oxygen uptake and water loss: a stochastic PDE model of respiratory water loss in insects. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01740-4.

Lou:2022:GDG

[3310] Yuan Lou and Hua Nie. Global dynamics of a generalist predator–prey model in open advective environments. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01756-w.

Mirani:2022:PLR

[3311] Arjun Mirani and Alex McAvoy. Payoff landscapes and the robustness of selfish optimization in iterated games. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01758-8.

Ndongo:2022:CBM

[3312] Mamadou Sadio Ndongo, Papa Ibrahima Ndiaye, and Mohamed Aziz Darghouth. A climate-based model for tick life cycle: positive semigroup theory on Cauchy problem approach. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01755-x.

Palacios:2022:EBT

[3313] Julia A. Palacios, Anand Bhaskar, and Noah A. Rosenberg. Enumeration of binary trees compatible with a perfect phylogeny. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01748-w.

Terauds:2022:NAA

[3314] Venta Terauds and Jeremy Sumner. A new algebraic approach to genome rearrangement models. Journal of Mathematical Biology, 84 (6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL https://link.springer.com/article/10.
1007/s00285-022-01744-0.

Zhang:2022:GAR

[3315] Ran Zhang and Jinliang Wang. On the global attractivity for a reactiondiffusion malaria model with incubation period in the vector population. Journal of Mathematical Biology, 84(6):??, May 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01751-1.

Pieschner:2022:IAM

[3316] Susanne Pieschner, Jan Hasenauer, and Christiane Fuchs. Identifiability analysis for models of the translation kinetics after mRNA transfection. Journal of Mathematical Biology, 84(7):??, June 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01739-x.

Ji:2022:SDI

[3317] Juping Ji, Genghong Lin, Lin Wang, and Ali Mai. Spatiotemporal dynamics induced by intraguild predator diffusion in an intraguild predation model. Journal of Mathematical Biology, 85(1):??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01772-w.

Montalban:2022:HIU

[3318] Antonio Montalbán, Rodrigo M. Corder, and M. Gabriela M. Gomes. Herd immunity under individual variation and reinfection. Journal of Mathematical Biology, 85(1):??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01771-x.

Eilertsen:2022:SEK

[3319] Justin Eilertsen, Kashvi Srivastava, and Santiago Schnell. Stochastic enzyme kinetics and the quasi-steady-state reductions: Application of the slow scale linear noise approximation à la Fenichel. Journal of Mathematical Biology, 85(1):??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01768-6.

Tam:2022:ABF

[3320] Alexander K. Y. Tam, Alex Mogilner, and Dietmar B. Oelz. F -actin bending facilitates net actomyosin contraction by inhibiting expansion with plus-end-located myosin motors. Journal of Mathematical Biology, 85(1):??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01737-z.

Zheng:2022:TGT

[3321] Xiaoming Zheng, Kun Zhao, Trachette Jackson, and John Lowengrub. Tumor growth towards lower extracellular matrix conductivity regions under Darcy’s law and steady morphology. Journal of Mathematical Biology, 85(1):??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01759-7.

Cantrell:2022:IFD

[3322] Robert Stephen Cantrell, Chris Cosner, and Ying Zhou. Ideal free dispersal in integrodifference models. Journal of Mathematical Biology, 85(1):??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01743-1.

Dragicevic:2022:SCE

[3323] Arnaud Z. Dragicevic and Anjula Gurtoo. Stochastic control of ecological networks. Journal of Mathematical Biology, 85(1):??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01777-5.

Young:2022:FSL

[3324] Glenn Young and Andrew Belmonte. Fixation in the stochastic Lotka– Volterra model with small fitness trade-offs. Journal of Mathematical Biology, 85(1):??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01774-8.

Blazej:2022:SGC

[3325] Paweł Błażej, Dariusz R. Kowalski, Dorota Mackiewicz, Małgorzata Wnetrzak, Daniyah A. Aloqalaa, and Paweł Mackiewicz. The structure of the genetic code as an optimal graph clustering problem. Journal of Mathematical Biology, 85(1):??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01778-4.

Alharbi:2022:HIV

[3326] Mohammed H. Alharbi and Christopher M. Kribs. How influenza vaccination and virus interference may impact combined influenza– coronavirus disease burden. Journal of Mathematical Biology, 85(1): ??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01767-7.

Manson:2022:COM

[3327] Kerry Manson, Charles Semple, and Mike Steel. Counting and optimising maximum phylogenetic diversity sets. Journal of Mathematical Biology, 85(1):??, July 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01779-3.

Cooney:2022:LTB

[3328] Daniel B. Cooney and Yoichiro Mori. Long-time behavior of a PDE replicator equation for multilevel selection in group-structured populations. Journal of Mathematical Biology, 85(2):??, August 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01776-6.

Scaramangas:2022:ASP

[3329] Alan Scaramangas and Mark Broom. Aposematic signalling in preypredator systems: determining evolutionary stability when prey populations consist of a single species. Journal of Mathematical Biology, 85(2):??, August 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01762-y.

Dean:2022:NDG

[3330] Justin Dean and Ayalvadi Ganesh. Noise dissipation in gene regulatory networks via second order statistics of networks of infinite server queues. Journal of Mathematical Biology, 85(2):??, August 2022. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-022-01781-9.

Yang:2022:CDH

[3331] Yantao Yang, Chaojing Ma, and Jian Zu. Coevolutionary dynamics of host-pathogen interaction with density-dependent mortality. Journal of Mathematical Biology, 85(2):??, August 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01782-8.

Degond:2022:MSV

[3332] Pierre Degond, Sophie Hecht, Michèle Romanos, and Ariane Trescases. Multi-species viscous models for tissue growth: incompressible limit and qualitative behaviour. Journal of Mathematical Biology, 85(2):??, August 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01784-6.

Song:2022:ADE

[3333] Pengfei Song and Yanni Xiao. Analysis of a diffusive epidemic system with spatial heterogeneity and lag effect of media impact. Journal of Mathematical Biology, 85(2):??, August 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01780-w.

Zheng:2022:IRP

[3334] Bo Zheng. Impact of releasing period and magnitude on mosquito population in a sterile release model with delay. Journal of Mathematical Biology, 85(2):??, August 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01785-5.

Wu:2022:ESS

[3335] Bin Wu. Evolutionary stability is sensitive on the conflict between reproduction and survival: proofs. Journal of Mathematical Biology, 85(2):??, August 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01775-7.

Emary:2022:SIT

[3336] Clive Emary and Anne-Kathleen Malchow. Stability-instability transition in tripartite merged ecological networks. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01783-7.

Lan:2022:GDA

[3337] Yuqiong Lan, Yanqiu Li, and Dongmei Zheng. Global dynamics of an age-dependent multiscale hepatitis C virus model. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01773-9.

Chan:2022:LSA

[3338] Yao ban Chan, Qiuyi Li, and Celine Scornavacca. The large-sample asymptotic behaviour of quartet-based summary methods for species tree inference. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01786-4.

Pan:2022:SMN

[3339] Qin Pan, Jicai Huang, and Hao Wang. An SIRS model with nonmonotone incidence and saturated treatment in a changing environment. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01787-3. See correction

[3412].

Katriel:2022:DIG

[3340] Guy Katriel. Dispersal-induced growth in a time-periodic environment. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01791-7.

Peng:2022:UBA

[3341] Q. Peng and F. J. Vermolen. Upscaling between an agent-based model (smoothed particle approach) and a continuum-based model for skin contractions. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01770-y.

Unterberger:2022:SDA

[3342] Jérémie Unterberger and Philippe Nghe. Stoechiometric and dynamical autocatalysis for diluted chemical reaction networks. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01798-0.

Kroumi:2022:AAC

[3343] Dhaker Kroumi and Sabin Lessard. Average abundancy of cooperation in multi-player games with random payoffs. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01789-1.

Elias:2022:AMC

[3344] Ján Eliaš, Hirofumi Izuhara, Masayasu Mimura, and Bao Q. Tang. An aggregation model of cockroaches with fast-or-slow motion dichotomy. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01797-1.

Lewkiewicz:2022:TSP

[3345] Stephanie M. Lewkiewicz, Sebastiano De Bona, Matthew R. Helmus, and Benjamin Seibold. Temperature sensitivity of pest reproductive numbers in age-structured PDE models, with a focus on the invasive spotted lanternfly. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01800-9.

Huber:2022:HNP

[3346] K. T. Huber and L. J. Maher. The hybrid number of a ploidy profile. Journal of Mathematical Biology, 85(3):??, September 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01792-6.

Ballesteros:2022:MNS

[3347] Miguel Ballesteros and Guillermo Garro. A model and a numerical scheme for the description of distribution and abundance of individuals. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01788-2.

Harris:2022:MMS

[3348] P. J. Harris and B. E. J. Bodmann. A mathematical model for simulating the spread of a disease through a country divided into geographical regions with different population densities. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01803-6.

Dipierro:2022:EFL

[3349] Serena Dipierro, Giovanni Giacomin, and Enrico Valdinoci. Efficiency functionals for the Lévy flight foraging hypothesis. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01808-1.

Khan:2022:QST

[3350] Muhammad Salman Khan, Maria Samreen, Muhammad Ozair, Takasar Hussain, E. M. Elsayed, and J. F. Gómez-Aguilar. On the qualitative study of a two-trophic plant-herbivore model. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01809-0.

Blanchini:2022:VRR

[3351] Franco Blanchini, Patrizio Colaneri, Giulia Giordano, and Irene Zorzan. Vertex results for the robust analysis of uncertain biochemical systems. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-022-01799-z.

Melikechi:2022:LEP

[3352] Omar Melikechi, Alexander L. Young, Tao Tang, Trevor Bowman, David Dunson, and James Johndrow. Limits of epidemic prediction using SIR models. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01804-5.

Eck:2022:RSE

[3353] Daniel J. Eck, Olga Morozova, and Forrest W. Crawford. Randomization for the susceptibility effect of an infectious disease intervention. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01801-8.

Sadhu:2022:AOR

[3354] Susmita Sadhu. Analysis of the onset of a regime shift and detecting early warning signs of major population changes in a two-trophic three-species predator–prey model with long-term transients. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01805-4.

Holzer:2022:LFD

[3355] Matt Holzer, Zachary Richey, Wyatt Rush, and Samuel Schmidgall. Locked fronts in a discrete time discrete space population model. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01802-7.

Narci:2022:IGS

[3356] Romain Narci, Maud Delattre, Catherine Larédo, and Elisabeta Vergu. Inference in Gaussian state-space models with mixed effects for multiple epidemic dynamics. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01806-3.

Bakker:2022:SRA

[3357] Bente Hilde Bakker, Timothy E. Faver, Hermen Jan Hupkes, Roeland M. H. Merks, and Jelle van der Voort. Scaling relations for auxin waves. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-022-01793-5.

Carrillo:2022:NDB

[3358] José A. Carrillo, Helge Holden, and Susanne Solem. Noise-driven bifurcations in a neural field system modelling networks of grid cells. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01811-6.

Foutel-Rodier:2022:IBE

[3359] Félix Foutel-Rodier, François Blanquart, Philibert Courau, Peter Czuppon, Jean-Jil Duchamps, Jasmine Gamblin, Élise Kerdoncuff, Rob Kulathinal, Léo Régnier, Laura Vuduc, Amaury Lambert, and Emmanuel Schertzer. From individual-based epidemic models to McKendrick-von Foerster PDEs: a guide to modeling and inferring COVID-19 dynamics. Journal of Mathematical Biology, 85(4):??, October 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01794-4.

Nouaoura:2022:ODT

[3360] Sarra Nouaoura, Radhouane Fekih-Salem, Nahla Abdellatif, and Tewfik Sari. Operating diagrams for a three-tiered microbial food web in the chemostat. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01812-5.

Basser-Ravitz:2022:CAN

[3361] Etan Basser-Ravitz, Arman Darbar, and Julia Chifman. Cyclic attractors of nonexpanding q-ary networks. Journal of Mathematical Biology, 85 (5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01796-2.

Friedman:2022:CMN

[3362] Avner Friedman, Wenrui Hao, and King-Yeung Lam. A cancer model with nonlocal free boundary dynamics. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01813-4.

Kao:2022:MTP

[3363] Chiu-Yen Kao and Seyyed Abbas Mohammadi. Maximal total population of species in a diffusive logistic model. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01817-0.

Vittadello:2022:GTA

[3364] Sean T. Vittadello and Michael P. H. Stumpf. A group theoretic approach to model comparison with simplicial representations. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ.

ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01807-2.

Li:2022:SBP

[3365] Shuping Li, Xiaorong Zhao, and Ruixia Zhang. Site-bond percolation model of epidemic spreading with vaccination in complex networks. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-022-01816-1.

Sun:2022:SDV

[3366] Gui-Quan Sun, Li-Feng Hou, Li Li, Zhen Jin, and Hao Wang. Spatial dynamics of a vegetation model with uptake-diffusion feedback in an arid environment. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL https://link.springer.com/article/10.
1007/s00285-022-01825-0.

Meaney:2022:TOR

[3367] Cameron Meaney, Mohammad Kohandel, and Arian Novruzi. Temporal optimization of radiation therapy to heterogeneous tumour populations and cancer stem cells. Journal of Mathematical Biology, 85 (5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01819-y.

Frolich:2022:PGI

[3368] Emil F. Frølich and Uffe H. Thygesen. Population games with instantaneous behavior and the Rosenzweig–MacArthur model. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01821-4.

Joshi:2022:FSD

[3369] Badal Joshi and Gheorghe Craciun. Foundations of static and dynamic absolute concentration robustness. Journal of Mathematical Biology, 85 (5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01823-2.

Hofbauer:2022:PIG

[3370] Josef Hofbauer and Sebastian J. Schreiber. Permanence via invasion graphs: incorporating community assembly into modern coexistence theory. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01815-2.

Mazari:2022:SEO

[3371] Idriss Mazari and Domènec Ruiz-Balet. Spatial ecology, optimal control and game theoretical fishing problems. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01829-w.

Giunta:2022:DME

[3372] Valeria Giunta, Thomas Hillen, Mark A. Lewis, and Jonathan R. Potts. Detecting minimum energy states and multi-stability in nonlocal advection-diffusion models for interacting species. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01824-1.

Campo:2022:GTM

[3373] Vince N. Campo, John Lawrence Palacios, Hideo Nagahashi, Hyunju Oh, Jan Rychtář, and Dewey Taylor. A game-theoretic model of rabies in domestic dogs with multiple voluntary preventive measures. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01826-z.

Huang:2022:SBL

[3374] Yaodan Huang and Bei Hu. Symmetry-breaking longitude bifurcations for a free boundary problem modeling small plaques in three dimensions. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01827-y.

Li:2022:WSC

[3375] Bingtuan Li and Garrett Otto. Wave speed and critical patch size for integro-difference equations with a strong Allee effect. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01814-3.

Schramm:2022:HRR

[3376] Wolfgang Schramm. The (human) respiratory rate at rest. Journal of Mathematical Biology, 85(5):??, November 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01790-8.

Zhang:2022:BAS

[3377] Hua Zhang and Junjie Wei. Bifurcation analysis for a single population model with advection. Journal of Mathematical Biology, 85(6– 7):??, December 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01818-z.

Kuehn:2022:ITP

[3378] Christian Kuehn and Jan Mölter. The influence of a transport process on the epidemic threshold. Journal of Mathematical Biology, 85 (6–7):??, December 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01810-7.

Ball:2022:EMS

[3379] Frank Ball and Peter Neal. An epidemic model with short-lived mixing groups. Journal of Mathematical Biology, 85(6–7):??, December 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01822-3.

Alvarez:2022:ESC

[3380] Frank Ernesto Alvarez, José Antonio Carrillo, and Jean Clairambault. Evolution of a structured cell population endowed with plasticity of traits under constraints on and between the traits. Journal of Mathematical Biology, 85(6–7):??, December 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01820-5.

Resende:2022:MSA

[3381] Anna Claudia M. Resende, Ernesto A. B. F. Lima, Regina C. Almeida, Matthew T. McKenna, and Thomas E. Yankeelov. Model selection for assessing the effects of doxorubicin on triple-negative breast cancer cell lines. Journal of Mathematical Biology, 85(6–7):??, December 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01828-x.

Akjouj:2022:FSL

[3382] Imane Akjouj and Jamal Najim. Feasibility of sparse large Lotka– Volterra ecosystems. Journal of Mathematical Biology, 85(6–7):??, December 2022. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01830-3.

Craciun:2022:HIR

[3383] Gheorghe Craciun and Abhishek Deshpande. Homeostasis and injectivity: a reaction network perspective. Journal of Mathematical Biology, 85 (6–7):??, December 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01795-3.

Gutierrez:2022:PEA

[3384] R. Gutiérrez, F. Córdova-Lepe, F. N. Moreno-Gómez, and N. A. Velásquez. Plastic energy allocation toward life-history functions in a consumer-resource interaction. Journal of Mathematical Biology, 85(6– 7):??, December 2022. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01834-z.

Goh:2022:TRA

[3385] Gary Goh, Michael Fuchs, and Louxin Zhang. Two results about the Sackin and Colless indices for phylogenetic trees and their shapes. Journal of Mathematical Biology, 85(6–7):??, December 2022. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-022-01831-2.

Stotsky:2023:EIF

[3386] Jay A. Stotsky and Hans G. Othmer. The effects of internal forces and membrane heterogeneity on three-dimensional cell shapes. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01836-x.

Zhang:2023:MMA

[3387] Haifeng Zhang, Meirong Zhang, and Jinzhi Lei. A mathematical model with aberrant growth correction in tissue homeostasis and tumor cell growth. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01837-w.

Rad:2023:SES

[3388] Mohsen Kamelian Rad, Mohammad Ali Ahmadi-Pajouh, and Mehrdad Saviz. Selective electrical stimulation of low versus high diameter myelinated fibers and its application in pain relief: a modeling study. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01833-0.

Coron:2023:OPP

[3389] Camille Coron, Manon Costa, Hélène Leman, Violaine Llaurens, and Charline Smadi. Origin and persistence of polymorphism in loci targeted by disassortative preference: a general model. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01832-1.

Salako:2023:IPS

[3390] Rachidi B. Salako. Impact of population size and movement on the persistence of a two-strain infectious disease. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01842-z.

Kong:2023:ESS

[3391] Fanze Kong, Juncheng Wei, and Liangshun Xu. The existence and stability of spikes in the one-dimensional Keller–Segel model with logistic growth. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01840-1.

Georg:2023:LRT

[3392] Peter Georg, Lars Grasedyck, Maren Klever, Rudolf Schill, Rainer Spang, and Tilo Wettig. Low-rank tensor methods for Markov chains with applications to tumor progression models. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01846-9.

Cerpa:2023:PAS

[3393] Eduardo Cerpa, Matías Courdurier, Esteban Hernández, Leonel E. Medina, and Esteban Paduro. A partially averaged system to model neuron responses to interferential current stimulation. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01839-8.

Browne:2023:VID

[3394] Cameron J. Browne and Fadoua Yahia. Virus-immune dynamics determined by prey-predator interaction network and epistasis in viral fitness landscape. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01843-y.

Allman:2023:TBS

[3395] Elizabeth S. Allman, Hector Baños, Jonathan D. Mitchell, and John A. Rhodes. The tree of blobs of a species network: identifiability under the coalescent. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01838-9.

Porter:2023:TMS

[3396] Thomas K. Porter, Michael N. Heinz, Daniel James Lundberg, Allan M. Brooks, Tedrick Thomas Salim Lew, Kevin S. Silmore, Volodymyr B. Koman, Mervin Chun-Yi Ang, Duc Thinh Khong, Gajendra Pratap Singh, James W. Swan, Rajani Sarojam, Nam-Hai Chua, and Michael S. Strano. A theory of mechanical stress-induced H_{2}O_{2} signaling waveforms in Planta. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01835-y.

Xu:2023:ILG

[3397] Jingcheng Xu and Cécile Ané. Identifiability of local and global features of phylogenetic networks from average distances. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01847-8.

Wicke:2023:EDB

[3398] Kristina Wicke, Mareike Fischer, and Laura Kubatko. Effects of discordance between species and gene trees on phylogenetic diversity conservation. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01845-w.

Kim:2023:RST

[3399] Yangjin Kim, Junho Lee, Chaeyoung Lee, and Sean Lawler. Role of senescent tumor cells in building a cytokine shield in the tumor microenvironment: mathematical modeling. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01850-z.

Eigentler:2023:LRS

[3400] L. Eigentler and J. A. Sherratt. Long-range seed dispersal enables almost stationary patterns in a model for dryland vegetation. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01852-x.

Altamirano:2023:PLN

[3401] Maximiliano Altamirano, Roberto Cortez, Matthieu Jonckheere, and Lasse Leskelä. Persistence in a large network of sparsely interacting neurons. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01844-x.

Boahen:2023:SER

[3402] Frank Boahen, Nicolas Doyon, and Jean Deteix. Sensitivity of the electrical response of a node of Ranvier model to alterations of the myelin sheath geometry. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01841-0.

Remesan:2023:TPM

[3403] Gopikrishnan C. Remesan, Jennifer A. Flegg, and Helen M. Byrne. Twophase model of compressive stress induced on a surrounding hyperelastic medium by an expanding tumour. Journal of Mathematical Biology, 86(1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01851-y.

Lu:2023:BAF

[3404] Min-Jhe Lu, Wenrui Hao, Bei Hu, and Shuwang Li. Bifurcation analysis of a free boundary model of vascular tumor growth with a necrotic core and chemotaxis. Journal of Mathematical Biology, 86 (1):??, January 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01862-9.

Al-Darabsah:2023:SHM

[3405] Isam Al-Darabsah, Kang-Ling Liao, and Stéphanie Portet. A simple inhost model for COVID-19 with treatments: model prediction and calibration. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01849-6.

Ghosh:2023:ADI

[3406] Samiran Ghosh, Vitaly Volpert, and Malay Banerjee. An agedependent immuno-epidemiological model with distributed recovery and death rates. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01855-8.

Hall:2023:ESB

[3407] Cameron Luke Hall and Bram Alexander Siebert. Exact solutions and bounds for network SIR and SEIR models using a rooted-tree approximation. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-022-01854-9.

Tang:2023:PDR

[3408] De Tang and Zhi-An Wang. Population dynamics with resourcedependent dispersal: single- and two-species models. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01856-7.

Muller:2023:CTS

[3409] Johannes Müller and Volker Hösel. Contact tracing & super-spreaders in the branching-process model. Journal of Mathematical Biology, 86 (2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01857-6.

Pu:2023:WNV

[3410] Liqiong Pu, Zhigui Lin, and Yuan Lou. A West Nile virus nonlocal model with free boundaries and seasonal succession. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01860-x.

Delmas:2023:OVV

[3411] Jean-François Delmas, Dylan Dronnier, and Pierre-André Zitt. Optimal vaccination: various (counter) intuitive examples. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-022-01858-5.

Pan:2023:CSM

[3412] Qin Pan, Jicai Huang, and Hao Wang. Correction: An SIRS model with nonmonotone incidence and saturated treatment in a changing environment. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01853-w. See [3339].

Plesa:2023:SAH

[3413] Tomislav Plesa. Stochastic approximations of higher-molecular by bimolecular reactions. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-022-01848-7.

Diekmann:2023:SPI

[3414] Odo Diekmann and Hisashi Inaba. A systematic procedure for incorporating separable static heterogeneity into compartmental epidemic models. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01865-0.

Barendregt:2023:HCE

[3415] Nicholas W. Barendregt and Peter J. Thomas. Heteroclinic cycling and extinction in May–Leonard models with demographic stochasticity. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01859-4.

Ji:2023:SEC

[3416] Juping Ji, Russell Milne, and Hao Wang. Stoichiometry and environmental change drive dynamical complexity and unpredictable switches in an intraguild predation model. Journal of Mathematical Biology, 86(2):??, February 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01866-z.

Zhao:2023:BRR

[3417] Hongyong Zhao, Kai Wang, and Hao Wang. Basic reproduction ratio of a mosquito-borne disease in heterogeneous environment. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01867-y.

Othmer:2023:ISC

[3418] Hans Othmer, Yuan Lou, Philip Maini, and Urszula Ledzewicz. Introduction to the special collection in honor of Avner Friedman. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01864-7.

Barrett:2023:ATG

[3419] Christopher Barrett, Andrei Bura, Qijun He, Fenix Huang, and Christian Reidys. The arithmetic topology of genetic alignments. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01868-x.

Wang:2023:HSM

[3420] Yan Wang, Jun Liu, Xinhong Zhang, and Jane M. Heffernan. An HIV stochastic model with cell-to-cell infection, B-cell immune response and distributed delay. Journal of Mathematical Biology, 86 (3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-022-01863-8.

Yin:2023:OCD

[3421] Shuangshuang Yin, Jianhong Wu, and Pengfei Song. Optimal control by deep learning techniques and its applications on epidemic models. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01873-0.

Deng:2023:VDI

[3422] Jiawei Deng, Hongying Shu, Lin Wang, and Xiang-Sheng Wang. Viral dynamics with immune responses: effects of distributed delays and Filippov antiretroviral therapy. Journal of Mathematical Biology, 86 (3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01869-w.

Ma:2023:MTH

[3423] Shizhao Ma, Jinzhi Lei, and Xiulan Lai. Modeling tumour heterogeneity of PD-L1 expression in tumour progression and adaptive therapy. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01872-1.

Villar-Sepulveda:2023:GCT

[3424] Edgardo Villar-Sepúlveda and Alan R. Champneys. General conditions for Turing and wave instabilities in reaction–diffusion systems. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01870-3. See correction

[3442].

Ban:2023:MAT

[3425] Jung-Chao Ban, Jyy-I Hong, and Yu-Liang Wu. Mathematical analysis of topological and random m-order spread models. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01874-z.

Cantrell:2023:EDR

[3426] R. S. Cantrell, C. Cosner, and R. B. Salako. Effects of dispersal rates in a two-stage reaction-diffusion system. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01878-9.

Zhang:2023:CAD

[3427] Xue Zhang and Jianhong Wu. A coupled algebraic-delay differential system modeling tick-host interactive behavioural dynamics and multistability. Journal of Mathematical Biology, 86(3):??, March 2023. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

https://link.springer.com/article/10.1007/s00285-023-01879-
8.

Lunz:2023:OCB

[3428] Davin Lunz, J. Frédéric Bonnans, and Jakob Ruess. Optimal control of bioproduction in the presence of population heterogeneity. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01876-x.

Snyman:2023:PRF

[3429] Jandre Snyman, Colin Fox, and David Bryant. Parsimony and the rank of a flattening matrix. Journal of Mathematical Biology, 86(3): ??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-

1416 (electronic). URL https://link.springer.com/article/10.
1007/s00285-023-01875-y.

Luckhaus:2023:FBP

[3430] Stephan Luckhaus and Angela Stevens. A free boundary problemin time-for the spread of Covid-19. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01881-0.

Schmiegelt:2023:APC

[3431] Benjamin Schmiegelt and Joachim Krug. Accessibility percolation on Cartesian power graphs. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01882-z.

Fierro:2023:CDM

[3432] Raúl Fierro. Cumulative damage for multi-type epidemics and an application to infectious diseases. Journal of Mathematical Biology, 86(3):??, March 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01880-1.

Ulloa:2023:TGR

[3433] J. Ignacio Fierro Ulloa, Liu-Di Lu, and Olivier Bernard. Theoretical growth rate of microalgae under high/low-flashing light. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01871-2.

Yoon:2023:SBA

[3434] Hong-Gyu Yoon and Pilwon Kim. STDP-based associative memory formation and retrieval. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01883-y.

Huynh:2023:IDD

[3435] Linh Huynh, Jacob G. Scott, and Peter J. Thomas. Inferring densitydependent population dynamics mechanisms through rate disambiguation for logistic birth-death processes. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01877-w.

Zhang:2023:MPS

[3436] Zhaowang Zhang, Lijie Chang, Qiming Huang, Rong Yan, and Bo Zheng. A mosquito population suppression model with a saturated Wolbachia release strategy in seasonal succession. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01888-7.

Lu:2023:RPB

[3437] Min Lu, Daozhou Gao, Jicai Huang, and Hao Wang. Relative prevalencebased dispersal in an epidemic patch model. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01887-8.

Hill:2023:IID

[3438] Andrew N. Hill, John W. Glasser, and Zhilan Feng. Implications for infectious disease models of heterogeneous mixing on control thresholds. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-023-01886-9.

Heggerud:2023:NDL

[3439] Christopher M. Heggerud, King-Yeung Lam, and Hao Wang. Niche differentiation in the light spectrum promotes coexistence of phytoplankton species: a spatial modelling approach. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01890-z.

Fok:2023:SSR

[3440] Pak-Wing Fok. Shear stress regulation in cylindrical arteries through medial growth and nitric oxide release. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01894-9.

Melo:2023:FSP

[3441] Alison M. V. D. L. Melo and Matheus C. Santos. Final size and partial distance estimate for a two-group SEIRD model. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01892-x.

Villar-Sepulveda:2023:CGC

[3442] Edgardo Villar-Sepúlveda and Alan R. Champneys. Correction to: General conditions for Turing and wave instabilities in reaction–diffusion systems. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01884-x. See

[3424].

Smith:2023:IRA

[3443] Delaney Smith and Anita Layton. The intrarenal renin-angiotensin system in hypertension: insights from mathematical modelling. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01891-y.

Egberts:2023:STD

[3444] Ginger Egberts, Fred Vermolen, and Paul van Zuijlen. Stability of a twodimensional biomorphoelastic model for post-burn contraction. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01893-w.

Berestycki:2023:EMH

[3445] Henri Berestycki, Benoît Desjardins, Joshua S. Weitz, and Jean-Marc Oury. Epidemic modeling with heterogeneity and social diffusion. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ.

ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-022-01861-w.

DellaMarca:2023:SMV

[3446] Rossella Della Marca, Nadia Loy, and Andrea Tosin. An SIR model with viral load-dependent transmission. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01901-z.

Crona:2023:GFL

[3447] Kristina Crona, Joachim Krug, and Malvika Srivastava. Geometry of fitness landscapes: peaks, shapes and universal positive epistasis. Journal of Mathematical Biology, 86(4):??, April 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01889-6.

Pell:2023:EVV

[3448] Bruce Pell, Samantha Brozak, Tin Phan, Fuqing Wu, and Yang Kuang. The emergence of a virus variant: dynamics of a competition model with cross-immunity time-delay validated by wastewater surveillance data for COVID-19. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01900-0.

Lindwall:2023:FPI

[3449] Gustav Lindwall and Philip Gerlee. Fast and precise inference on diffusivity in interacting particle systems. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01902-y.

Zhang:2023:PMS

[3450] Hua Zhang, Hao Wang, and Junjie Wei. Perceptive movement of susceptible individuals with memory. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01904-w.

Webb:2023:BAC

[3451] Glenn Webb and Xinyue Evelyn Zhao. Bifurcation analysis of critical values for wound closure outcomes in wound healing experiments. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ.

ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01896-7.

Kenne:2023:NMB

[3452] Cyrille Kenne, Gisèle Mophou, and Pascal Zongo. A nested model with boosting and waning of immunity from Tilapia Lake Virus infection with distributed resistance to pathogens carrier-state. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01906-8.

Hamis:2023:SCM

[3453] Sara Hamis, Panu Somervuo, J. Arvid Ågren, Dagim Shiferaw Tadele, Juha Kesseli, Jacob G. Scott, Matti Nykter, Philip Gerlee, Dmitri Finkelshtein, and Otso Ovaskainen. Spatial cumulant models enable spatially informed treatment strategies and analysis of local interactions in cancer systems. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01903-x.

Ardiyansyah:2023:ECM

[3454] Muhammad Ardiyansyah, Dimitra Kosta, and Jordi Roca-Lacostena. Embeddability of centrosymmetric matrices capturing the double-helix structure in natural and synthetic DNA. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01895-8.

Cheek:2023:ARB

[3455] David Cheek and Samuel G. G. Johnston. Ancestral reproductive bias in branching processes. Journal of Mathematical Biology, 86(5): ??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01907-7.

Wang:2023:OPP

[3456] Hao Wang and Yurij Salmaniw. Open problems in PDE models for knowledge-based animal movement via nonlocal perception and cognitive mapping. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01905-9.

Gomez:2023:FPS

[3457] Daniel Gomez, King-Yeung Lam, and Yoichiro Mori. Front propagation in the shadow wave-pinning model. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01908-6.

Drosinou:2023:SPM

[3458] Ourania Drosinou, Christos V. Nikolopoulos, Anastasios Matzavinos, and Nikos I. Kavallaris. A stochastic parabolic model of MEMS driven by fractional Brownian motion. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01897-6.

Gurski:2023:SHT

[3459] Katharine Gurski and Kathleen Hoffman. Staged HIV transmission and treatment in a dynamic model with long-term partnerships. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01885-w.

Saldana:2023:OVA

[3460] Fernando Saldaña, Vanessa Steindorf, Akhil Kumar Srivastav, Nico Stollenwerk, and Maíra Aguiar. Optimal vaccine allocation for the control of sexually transmitted infections. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01910-y.

Loutchko:2023:ACS

[3461] Dimitri Loutchko. An algebraic characterization of self-generating chemical reaction networks using semigroup models. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01899-4.

Huang:2023:FBP

[3462] Yu-Jhe Huang, Jonq Juang, Tai-Yi Kuo, and Yu-Hao Liang. Forwardbackward and period doubling bifurcations in a discrete epidemic model with vaccination and limited medical resources. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01911-x.

Loutchko:2023:SMB

[3463] Dimitri Loutchko. Semigroup models for biochemical reaction networks. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01898-5.

Fralix:2023:MAS

[3464] Brian Fralix, Mark Holmes, and Andreas Löpker. A Markovian arrival stream approach to stochastic gene expression in cells. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01913-9.

Liu:2023:GSL

[3465] Xiaogang Liu, Yuming Chen, Xiaomin Li, and Jianquan Li. Global stability of latency-age/stage-structured epidemic models with differential infectivity. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01918-4.

Xu:2023:IQM

[3466] Wanxiao Xu, Hongying Shu, Lin Wang, Xiang-Sheng Wang, and James Watmough. The importance of quarantine: modelling the COVID-19 testing process. Journal of Mathematical Biology, 86(5): ??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01916-6.

Bai:2023:AIM

[3467] Fan Bai. An age-of-infection model with both symptomatic and asymptomatic infections. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01920-w.

Arachchilage:2023:EHE

[3468] Kalpana Hanthanan Arachchilage, Mohammed Y. Hussaini, N. G. Cogan, and Michael H. Cortez. Exploring how ecological and epidemiological processes shape multi-host disease dynamics using global sensitivity analysis. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01912-w.

Alebraheem:2023:NMM

[3469] Jawdat Alebraheem and Yahya Abu-Hassan. A novel mechanism measurement of predator interference in predator–prey models. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01914-8.

Kribs:2023:ITD

[3470] Christopher Kribs and David Greenhalgh. Impact of tetravalent dengue vaccination with screening, ADE, and altered infectivity on singleserotype dengue and Zika transmission. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01915-7.

Moyles:2023:DSI

[3471] Iain R. Moyles, Chapin S. Korosec, and Jane M. Heffernan. Determination of significant immunological timescales from mRNA-LNP-based vaccines in humans. Journal of Mathematical Biology, 86(5): ??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01919-3.

Wang:2023:IAG

[3472] Yue Wang and Siqi He. Inference on autoregulation in gene expression with variance-to-mean ratio. Journal of Mathematical Biology, 86(5):??, May 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01924-6.

Vu:2023:WCW

[3473] Nhat L. Vu, Thanh P. Nguyen, Binh T. Nguyen, Vu Dinh, and Lam Si Tung Ho. When can we reconstruct the ancestral state? Beyond Brownian motion. Journal of Mathematical Biology, 86(6):??, June 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01922-8.

Buckwar:2023:SHM

[3474] Evelyn Buckwar, Martina Conte, and Amira Meddah. A stochastic hierarchical model for low grade glioma evolution. Journal of Mathematical Biology, 86(6):??, June 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01909-5.

Lawley:2023:SFP

[3475] Sean D. Lawley and Joshua Johnson. Slowest first passage times, redundancy, and menopause timing. Journal of Mathematical Biology, 86(6):??, June 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01921-9.

Siewe:2023:HDT

[3476] Nourridine Siewe and Abdul-Aziz Yakubu. Hybrid discrete-time-continuous-time models and a SARS CoV-2 mystery: Sub-Saharan Africa’s low SARS CoV-2 disease burden. Journal of Mathematical Biology, 86(6):??, June 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01923-7.

Lemaire:2023:IMT

[3477] Louisiane Lemaire, Mathieu Desroches, Martin Krupa, and Massimo Mantegazza. Idealized multiple-timescale model of cortical spreading depolarization initiation and pre-epileptic hyperexcitability caused by Na_{V} 1.1/SCN1A mutations. Journal of Mathematical Biology, 86(6): ??, June 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01917-5.

Cao:2023:TSL

[3478] Hui Cao, Baojun Song, and Yicang Zhou. Treatment strategies for the latent tuberculosis infections. Journal of Mathematical Biology, 86(6):??, June 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01927-3.

Zhang:2023:BDG

[3479] Yuyue Zhang, Jicai Huang, and Hao Wang. Bifurcations driven by generalist and specialist predation: mathematical interpretation of Fennoscandia phenomenon. Journal of Mathematical Biology, 86(6): ??, June 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01929-1.

Watanabe:2023:EEN

[3480] Junya Watanabe. Exact expressions and numerical evaluation of average evolvability measures for characterizing and comparing G matrices. Journal of Mathematical Biology, 86(6):??, June 2023. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-023-01930-8.

Valega-Mackenzie:2023:RAP

[3481] Wencel Valega-Mackenzie, Jason Bintz, and Suzanne Lenhart. Resource allocation in a PDE ecosystem model. Journal of Mathematical Biology, 86(6):??, June 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01932-6.

Tang:2023:PFM

[3482] Xiaoxia Tang, Shuwang Li, John S. Lowengrub, and Steven M. Wise. Phase field modeling and computation of vesicle growth or shrinkage. Journal of Mathematical Biology, 86(6):??, June 2023. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-023-01928-2.

Griffiths:2023:ERR

[3483] Robert C. Griffiths and Paul A. Jenkins. An estimator for the recombination rate from a continuously observed diffusion of haplotype frequencies. Journal of Mathematical Biology, 86(6):??, June 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01931-7.

Elbetch:2023:NDM

[3484] Bilel Elbetch and Ali Moussaoui. Nonlinear diffusion in multi-patch logistic model. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01936-2.

Bansaye:2023:DTI

[3485] Vincent Bansaye and Bertrand Cloez. From the distributions of times of interactions to preys and predators dynamical systems. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01925-5.

Richter:2023:SDG

[3486] Hendrik Richter. Spectral dynamics of guided edge removals and identifying transient amplifiers for death-birth updating. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01937-1.

Hiremath:2023:DDM

[3487] Sandesh Athni Hiremath and Christina Surulescu. Data driven modeling of pseudopalisade pattern formation. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01933-5.

Xue:2023:AIS

[3488] Ling Xue, Xiulei Jin, and Huaiping Zhu. Assessing the impact of serostatus-dependent immunization on mitigating the spread of dengue virus. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01944-2.

Ribeiro:2023:CAP

[3489] Henrique A. L. Ribeiro, Yogesh Scindia, Borna Mehrad, and Reinhard Laubenbacher. COVID-19-associated pulmonary aspergillosis in immunocompetent patients: a virtual patient cohort study. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01940-6.

Zitouni:2023:SIC

[3490] Nour El Houda Zitouni, Mohamed Dellal, and Mustapha Lakrib. Substrate inhibition can produce coexistence and limit cycles in the chemostat model with allelopathy. Journal of Mathematical Biology, 87(1): ??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01943-3.

Katsaounis:2023:SDE

[3491] Dimitrios Katsaounis, Mark A. J. Chaplain, and Nikolaos Sfakianakis. Stochastic differential equation modelling of cancer cell migration and tissue invasion. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01934-4.

Zavaleta-Viveros:2023:MKM

[3492] José Alfredo Zavaleta-Viveros, Porfirio Toledo, Martha Lorena Avendaño-Garrido, Jesús Enrique Escalante-Martínez, María-Leonor López-Meraz, and Karen Paola Ramos-Riera. A modification to the Kuramoto model to simulate epileptic seizures as synchronization. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01938-0.

Poloni:2023:IME

[3493] Silas Poloni and Frithjof Lutscher. Integrodifference models for evolutionary processes in biological invasions. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01947-z.

Andersen:2023:ASP

[3494] M. Andersen, K. O. Bangsgaard, J. G. Heaf, and J. T. Ottesen. Analytical solution of phosphate kinetics for hemodialysis. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01942-4.

Meng:2023:PTE

[3495] Lingqi Meng and Naoki Masuda. Perturbation theory for evolution of cooperation on networks. Journal of Mathematical Biology, 87(1): ??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01941-5.

Clenet:2023:ESS

[3496] Maxime Clenet, François Massol, and Jamal Najim. Equilibrium and surviving species in a large Lotka–Volterra system of differential equations. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01939-z.

Sun:2023:CDS

[3497] Mengfeng Sun and Xinchu Fu. Competitive dual-strain SIS epidemiological models with awareness programs in heterogeneous networks: two modeling approaches. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01945-1.

Lee:2023:LBG

[3498] Seungjoon Lee, Yorgos M. Psarellis, Constantinos I. Siettos, and Ioannis G. Kevrekidis. Learning black- and gray-box chemotactic PDEs/ closures from agent based Monte Carlo simulation data. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01946-0.

Ducasse:2023:PPM

[3499] Romain Ducasse and Samuel Nordmann. Propagation properties in a multi-species SIR reaction–diffusion system. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01950-4.

Ruan:2023:IBT

[3500] Shigui Ruan and Dongmei Xiao. Imperfect and Bogdanov–Takens bifurcations in biological models: from harvesting of species to isolation of infectives. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01951-3.

Jaramillo-Aguayo:2023:PFM

[3501] Pedro Jaramillo-Aguayo, Annabelle Collin, and Clair Poignard. Phasefield model of bilipid membrane electroporation. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01956-y.

Chen:2023:ISP

[3502] Jun Chen, Jordy O. Rodriguez Rincon, Gloria DeGrandi-Hoffman, Jennifer Fewell, Jon Harrison, and Yun Kang. Impacts of seasonality and parasitism on honey bee population dynamics. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01952-2.

Lou:2023:HMD

[3503] Yuan Lou, Rachidi B. Salako, and Pengfei Song. Human mobility and disease prevalence. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01953-1.

Dey:2023:ADS

[3504] Subrata Dey, S. Ghorai, and Malay Banerjee. Analytical detection of stationary and dynamic patterns in a prey-predator model with reproductive Allee effect in prey growth. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01957-x.

Wang:2023:PGV

[3505] Cuihua Wang, Hao Wang, and Sanling Yuan. Precipitation governing vegetation patterns in an arid or semi-arid environment. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01954-0.

Martinez:2023:OPP

[3506] Carlos Martínez, Eugenio Cinquemani, Hidde de Jong, and Jean-Luc Gouzé. Optimal protein production by a synthetic microbial consortium: coexistence, distribution of labor, and syntrophy. Journal of Mathematical Biology, 87(1):??, July 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01935-3.

Vo:2023:AMM

[3507] MyVan Vo, Zhilan Feng, John W. Glasser, Kristie E. N. Clarke, and Jefferson N. Jones. Analysis of metapopulation models of the transmission of SARS–CoV-2 in the United States. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01948-y.

Zabelkin:2023:TBE

[3508] Alexey Zabelkin, Pavel Avdeyev, and Nikita Alexeev. TruEst: a better estimator of evolutionary distance under the INFER model. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ.

ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01955-z.

Pearson:2023:LIM

[3509] Antony Pearson and Manuel E. Lladser. On latent idealized models in symbolic datasets: unveiling signals in noisy sequencing data. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01961-1.

Patsatzis:2023:ACV

[3510] Dimitris G. Patsatzis and Dimitris A. Goussis. Algorithmic criteria for the validity of quasi-steady state and partial equilibrium models: the Michaelis–Menten reaction mechanism. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01962-0.

Mohammed-Awel:2023:CIR

[3511] Jemal Mohammed-Awel and Abba B. Gumel. Can insecticide resistance increase malaria transmission? A genetics-epidemiology mathematical modeling approach. Journal of Mathematical Biology, 87 (2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01949-x.

He:2023:HIS

[3512] Runzi He, Xiaofeng Luo, Joshua Kiddy K. Asamoah, Yongxin Zhang, Yihong Li, Zhen Jin, and Gui-Quan Sun. A hierarchical intervention scheme based on epidemic severity in a community network. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01964-y.

Klay:2023:PPF

[3513] Léna Kläy, Léo Girardin, Vincent Calvez, and Florence Débarre. Pulled, pushed or failed: the demographic impact of a gene drive can change the nature of its spatial spread. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01926-4.

Salazar:2023:IEP

[3514] Fabián Mauricio Vélez Salazar and Iván David Patiño Arcila. Influence of electric pulse characteristics on the cellular internalization of chemotherapeutic drugs and cell survival fraction in electroporated and vasoconstricted cancer tissues using boundary element techniques. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01963-z.

Djidjou-Demasse:2023:GBT

[3515] Ramsès Djidjou-Demasse, Ibou Goudiaby, and Ousmane Seydi. Growth bound and threshold dynamic for nonautonomous nondensely defined evolution problems. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01966-w.

Nareddy:2023:MPP

[3516] Vahini Reddy Nareddy, Jonathan Machta, Karen Abbott, Shadisadat Esmaeili, and Alan Hastings. Modeling and prediction of phase shifts in noisy two-cycle oscillations. Journal of Mathematical Biology, 87 (2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01960-2.

Clark:2023:AMI

[3517] Chad Clark, Julius Jonušas, James D. Mitchell, and Andrew Francis. An algebraic model for inversion and deletion in bacterial genome rearrangement. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01965-x.

Pakkanen:2023:UIP

[3518] Mikko S. Pakkanen, Xenia Miscouridou, Matthew J. Penn, Charles Whittaker, Tresnia Berah, Swapnil Mishra, Thomas A. Mellan, and Samir Bhatt. Unifying incidence and prevalence under a time-varying general branching process. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01958-w.

Kiss:2023:NSC

[3519] István Z. Kiss, Eben Kenah, and Grzegorz A. Rempała. Necessary and sufficient conditions for exact closures of epidemic equations on configuration model networks. Journal of Mathematical Biology, 87 (2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01967-9.

Kribs:2023:CFB

[3520] Christopher Kribs, Shigui Ruan, and Zhilan Feng. A celebration of Fred Brauer’s legacy in mathematical biology. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01971-z.

Hamelin:2023:SSI

[3521] Frédéric M. Hamelin, Frank M. Hilker, and Yves Dumont. Spatial spread of infectious diseases with conditional vector preferences. Journal of Mathematical Biology, 87(2):??, August 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01972-y.

He:2023:DCR

[3522] Xiaoqing He, Wei-Ming Ni, and Haoyi Wang. Dynamics of consumerresource reaction-diffusion models: single and multiple consumer species. Journal of Mathematical Biology, 87(3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01970-0.

Burie:2023:ABE

[3523] Jean-Baptiste Burie, Arnaud Ducrot, and Quentin Griette. Asymptotic behavior of an epidemic model with infinitely many variants. Journal of Mathematical Biology, 87(3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01975-9.

Wang:2023:SDM

[3524] Lei Wang, Zhidong Teng, Xi Huo, Kai Wang, and Xiaomei Feng. A stochastic dynamical model for nosocomial infections with co-circulation of sensitive and resistant bacterial strains. Journal of Mathematical Biology, 87(3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01968-8.

Goel:2023:SEM

[3525] Shashank Goel, Sumit Kaur Bhatia, Jai Prakash Tripathi, Sarita Bugalia, Mansi Rana, and Vijay Pal Bajiya. SIRC epidemic model with crossimmunity and multiple time delays. Journal of Mathematical Biology, 87(3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01974-w.

Sinzger-DAngelo:2023:BBL

[3526] Mark Sinzger-D’Angelo, Sofia Startceva, and Heinz Koeppl. Bye bye, linearity, bye: quantification of the mean for linear CRNs in a random environment. Journal of Mathematical Biology, 87(3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01973-x.

Dimaschko:2023:SBC

[3527] J. Dimaschko, V. Shlyakhover, and M. Iabluchanskyi. Strong biological correlations as a cause of autonomous oscillations in epidemics. Journal of Mathematical Biology, 87(3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01976-8.

Macdonald:2023:FHH

[3528] J. C. Macdonald and H. Gulbudak. Forward hysteresis and Hopf bifurcation in an Npzd model with application to harmful algal blooms. Journal of Mathematical Biology, 87(3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01969-7.

Kang:2023:ABB

[3529] Hao Kang. Asymptotic behavior of the basic reproduction number in an age-structured SIS epidemic patch model. Journal of Mathematical Biology, 87(3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01979-5.

Doublet:2023:SPI

[3530] Violette Doublet, Lionel Roques, Etienne K. Klein, François Lefèvre, and Thomas Boivin. Seed predation-induced Allee effects, seed dispersal and masting jointly drive the diversity of seed sources during population expansion. Journal of Mathematical Biology, 87(3): ??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01981-x.

Le:2023:QND

[3531] Thi Minh Thao Le, Erida Gjini, and Sten Madec. Quasi-neutral dynamics in a coinfection system with N strains and asymmetries along multiple traits. Journal of Mathematical Biology, 87(3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01977-7.

Bertsch:2023:RBT

[3532] Michiel Bertsch, Bruno Franchi, Maria Carla Tesi, and Veronica Tora. The role of Aβ and Tau proteins in Alzheimer’s disease: a mathematical model on graphs. Journal of Mathematical Biology, 87(3): ??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01985-7.

Wieder:2023:GEF

[3533] Frederik Wieder, Martin Henk, and Alexander Bockmayr. On the geometry of elementary flux modes. Journal of Mathematical Biology, 87 (3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01982-w.

Wright:2023:MRR

[3534] Justin Wright, Kelly Buch, Ursula K. Beattie, Brenna M. G. Gormally, L. Michael Romero, and Nina Fefferman. A mathematical representation of the reactive scope model. Journal of Mathematical Biology, 87 (3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01983-9.

Walker:2023:WPS

[3535] Christoph Walker. Well-posedness and stability analysis of an epidemic model with infection age and spatial diffusion. Journal of Mathematical Biology, 87(3):??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01980-y.

Overwater:2023:ELF

[3536] Marcus Overwater, Daniel Pelletier, and Mike Steel. The expected loss of feature diversity (versus phylogenetic diversity) following rapid extinction at the present. Journal of Mathematical Biology, 87(3): ??, September 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01988-4.

Schardt:2023:ARC

[3537] Simon Schardt and Sabine C. Fischer. Adjusting the range of cell-cell communication enables fine-tuning of cell fate patterns from checkerboard to engulfing. Journal of Mathematical Biology, 87(4):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01959-9.

Xu:2023:MET

[3538] Jingjing Xu, Zhen Wang, and Seyed M. Moghadas. Modelling the effect of travel-related policies on disease control in a meta-population structure. Journal of Mathematical Biology, 87(4):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01990-w.

Bovier:2023:SIA

[3539] Anton Bovier and Lisa Hartung. The speed of invasion in an advancing population. Journal of Mathematical Biology, 87(4):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01989-3.

Gilbertson:2023:BPI

[3540] Nora M. Gilbertson and Mark Kot. Block-pulse integrodifference equations. Journal of Mathematical Biology, 87(4):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01986-6.

Tripathi:2023:CEP

[3541] Jai Prakash Tripathi, Deepak Tripathi, Swarnendu Mandal, and Manish Dev Shrimali. Cannibalistic enemy-pest model: effect of additional food and harvesting. Journal of Mathematical Biology, 87(4):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01991-9.

Li:2023:CMA

[3542] Aaron Li, Danika Kibby, and Jasmine Foo. A comparison of mutation and amplification-driven resistance mechanisms and their impacts on tumor recurrence. Journal of Mathematical Biology, 87(4):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01992-8.

Grumbach:2023:EDA

[3543] Carolin Grumbach, Femke N. Reurik, Juan Segura, Daniel Franco, and Frank M. Hilker. The effect of dispersal on asymptotic total population size in discrete- and continuous-time two-patch models. Journal of Mathematical Biology, 87(4):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01984-8.

Elgart:2023:PAA

[3544] Shoshana Elgart. A perturbative approach to the analysis of manycompartment models characterized by the presence of waning immunity. Journal of Mathematical Biology, 87(4):??, October 2023. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-023-01994-6.

Nguyen:2023:IRD

[3545] Tung D. Nguyen, Yixiang Wu, Tingting Tang, Amy Veprauskas, Ying Zhou, Behzad Djafari Rouhani, and Zhisheng Shuai. Impact of resource distributions on the competition of species in stream environment. Journal of Mathematical Biology, 87(4):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01978-6.

Roitershtein:2023:ESE

[3546] Alexander Roitershtein, Reza Rastegar, Robert S. Chapkin, and Ivan Ivanov. Extinction scenarios in evolutionary processes: a multinomial Wright–Fisher approach. Journal of Mathematical Biology, 87 (4):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-01993-7.

Bhaumik:2023:FPE

[3547] Jnanajyoti Bhaumik and Naoki Masuda. Fixation probability in evolutionary dynamics on switching temporal networks. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-01987-5.

Fromion:2023:SMR

[3548] Vincent Fromion, Philippe Robert, and Jana Zaherddine. Stochastic models of regulation of transcription in biological cells. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-01998-2.

Guo:2023:GDT

[3549] Zhong-Kai Guo, Hai-Feng Huo, Hong Xiang, and Qiu-Yan Ren. Global dynamics of a tuberculosis model with age-dependent latency and time delays in treatment. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01999-1.

Guillen-Gonzalez:2023:FPT

[3550] F. Guillén-González, E. Sevillano-Castellano, and A. Suárez. Fitting parameters and therapies of ODE tumor models with senescence and immune system. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-02000-9.

Alfaro:2023:AHE

[3551] Matthieu Alfaro, François Hamel, Florian Patout, and Lionel Roques. Adaptation in a heterogeneous environment II: to be three or not to be. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01996-4.

Abtout:2023:ADL

[3552] Annia Abtout and Jürgen Reingruber. Analysis of dim-light responses in rod and cone photoreceptors with altered calcium kinetics. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02005-4.

Linz:2023:ESS

[3553] Simone Linz and Kristina Wicke. Exploring spaces of semi-directed level-1 networks. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-02004-5.

Yang:2023:RDB

[3554] Junyuan Yang, Peiqi Jia, Jin Wang, and Zhen Jin. Rich dynamics of a bidirectionally linked immuno-epidemiological model for cholera. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02009-0.

Zhang:2023:TDA

[3555] Tailei Zhang and Xiao-Qiang Zhao. Threshold dynamics of an almost periodic vector-borne disease model. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02002-7.

Rao:2023:DAG

[3556] Feng Rao, Zhongliang Zhang, and Jiaxu Li. Dynamical analysis of a glucose-insulin regulatory system with insulin-degrading enzyme and multiple delays. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-02003-6.

Howerton:2023:EGS

[3557] Emily Howerton, Kyle Dahlin, Christina J. Edholm, Lindsey Fox, Margaret Reynolds, Brandon Hollingsworth, George Lytle, Melody Walker, Julie Blackwood, and Suzanne Lenhart. The effect of governance structures on optimal control of two-patch epidemic models. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02001-8.

Dumm:2023:REE

[3558] Will Dumm, Mary Barker, William Howard-Snyder, William S. DeWitt III, and Frederick A. Matsen IV. Representing and extending ensembles of parsimonious evolutionary histories with a directed acyclic graph. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-02006-3.

Morrison:2023:MBS

[3559] Maike L. Morrison and Noah A. Rosenberg. Mathematical bounds on Shannon entropy given the abundance of the i-th most abundant taxon.

Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-01997-3.

Duong:2023:COH

[3560] M. H. Duong, C. M. Durbac, and T. A. Han. Cost optimisation of hybrid institutional incentives for promoting cooperation in finite populations. Journal of Mathematical Biology, 87(5):??, November 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-02011-6.

Mann-Manyombe:2023:LWB

[3561] Martin L. Mann-Manyombe, Abdoulaye Mendy, Ousmane Seydi, and Ramsès Djidjou-Demasse. Linking within- and between-host scales for understanding the evolutionary dynamics of quantitative antimicrobial resistance. Journal of Mathematical Biology, 87(6):??, December 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-02008-1.

Chowell:2023:SIA

[3562] Gerardo Chowell, Sushma Dahal, Yuganthi R. Liyanage, Amna Tariq, and Necibe Tuncer. Structural identifiability analysis of epidemic models based on differential equations: a tutorial-based primer. Journal of Mathematical Biology, 87(6):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02007-2.

Yerlanov:2023:EPS

[3563] Madi Yerlanov, Piyush Agarwal, Caroline Colijn, and Jessica E. Stockdale. Effective population size in simple infectious disease models. Journal of Mathematical Biology, 87(6):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02016-1.

Peng:2023:NSP

[3564] Rui Peng, Zhi-An Wang, Guanghui Zhang, and Maolin Zhou. Novel spatial profiles of some diffusive SIS epidemic models. Journal of Mathematical Biology, 87(6):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-02013-4.

Cammarota:2023:LVM

[3565] Denise Cammarota, Noemi Zeraick Monteiro, Rafael Menezes, Hugo Fort, and Angel M. Segura. Lotka–Volterra model with Allee effect: equilibria, coexistence and size scaling of maximum and minimum abundance. Journal of Mathematical Biology, 87(6):??, October 2023. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-023-02012-5.

Ball:2023:IHS

[3566] Frank Ball, Liam Critcher, Peter Neal, and David Sirl. The impact of household structure on disease-induced herd immunity. Journal of Mathematical Biology, 87(6):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-02010-7.

Dallaston:2023:ECC

[3567] Michael C. Dallaston, Geneva Birtles, Robyn P. Araujo, and Adrianne L. Jenner. The effect of chemotaxis on T-cell regulatory dynamics. Journal of Mathematical Biology, 87(6):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02017-0.

Phan:2023:SDH

[3568] Tuan Anh Phan, Farhana Sarower, Jinqiao Duan, and Jianjun Paul Tian. Stochastic dynamics of human papillomavirus delineates cervical cancer progression. Journal of Mathematical Biology, 87(6):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-02018-z.

Ding:2023:IPS

[3569] Ming-Hui Ding, Hongyu Liu, and Guang-Hui Zheng. On inverse problems for several coupled PDE systems arising in mathematical biology. Journal of Mathematical Biology, 87(6):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02021-4.

Navas-Zuloaga:2023:MFA

[3570] M. Gabriela Navas-Zuloaga, Kaitlin M. Baudier, Jennifer H. Fewell, Noam Ben-Asher, Theodore P. Pavlic, and Yun Kang. A modeling framework for adaptive collective defense: crisis response in socialinsect colonies. Journal of Mathematical Biology, 87(6):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-01995-5.

Olofsson:2023:MMA

[3571] Peter Olofsson, Logan Chipkin, Ryan C. Daileda, and Ricardo B. R. Azevedo. Mutational meltdown in asexual populations doomed to extinction. Journal of Mathematical Biology, 87(6):??, October 2023. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-02019-y.

Rahman:2024:MMA

[3572] Nizhum Rahman and Dietmar B. Oelz. A mathematical model for axonal transport of large cargo vesicles. Journal of Mathematical Biology, 88(1):??, January 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02022-3.

Lan:2024:TBE

[3573] Guijie Lan, Sanling Yuan, and Baojun Song. Threshold behavior and exponential ergodicity of an SIR epidemic model: the impact of random jamming and hospital capacity. Journal of Mathematical Biology, 88(1):??, January 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02024-1.

Bichara:2024:CDS

[3574] Derdei M. Bichara. Characterization of differential susceptibility and differential infectivity epidemic models. Journal of Mathematical Biology, 88(1):??, January 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02023-2.

Hillen:2024:MMD

[3575] Thomas Hillen, Nadia Loy, Kevin J. Painter, and Ryan Thiessen. Modelling microtube driven invasion of glioma. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02025-0.

Chang:2024:TIL

[3576] Lili Chang, Xinyu Wang, Guiquan Sun, Zhen Wang, and Zhen Jin. A time independent least squares algorithm for parameter identification of Turing patterns in reaction–diffusion systems. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02026-z.

Yeni:2024:TST

[3577] Gülsah Yeni, Elvan Akin, and Naveen K. Vaidya. Time scale theory on stability of explicit and implicit discrete epidemic models: applications to swine flu outbreak. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02015-2.

Mehra:2024:SHR

[3578] Somya Mehra, James M. McCaw, and Peter G. Taylor. Superinfection and the hypnozoite reservoir for Plasmodium vivax: a general framework. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-02014-3.

Shabbir:2024:EDP

[3579] Muhammad Sajjad Shabbir, Qamar Din, Manuel De la Sen, and J. F. Gómez-Aguilar. Exploring dynamics of plant-herbivore interactions: bifurcation analysis and chaos control with Holling type-II functional response. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-02020-5.

Magal:2024:RHM

[3580] Pierre Magal. A return-to-home model with commuting people and workers. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-02028-x.

Anderson:2024:GSP

[3581] Hannah G. Anderson, Gregory P. Takacs, Duane C. Harris, Yang Kuang, Jeffrey K. Harrison, and Tracy L. Stepien. Global stability and parameter analysis reinforce therapeutic targets of PD-L1-PD-1 and MDSCs for glioblastoma. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02027-y.

El-Hachem:2024:CTS

[3582] Maud El-Hachem and Nicholas J. Beeton. Coexistence in two-species competition with delayed maturation. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02031-2.

Pires:2024:ABM

[3583] Rosângela A. Pires and Armando G. M. Neves. Asymptotic behavior of mean fixation times in the Moran process with frequency-independent fitnesses. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-023-02032-1.

Ai:2024:OVA

[3584] Mingxia Ai and Wendi Wang. Optimal vaccination ages for emerging infectious diseases under limited vaccine supply. Journal of Mathematical Biology, 88(1):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02030-3.

Roy:2024:ROC

[3585] Souvik Roy, Zui Pan, Naif Abu Qarnayn, Mesfer Alajmi, Ali Alatawi, Asma Alghamdi, Ibrahem Alshaoosh, Zahra Asiri, Berlinda Batista, Shreshtha Chaturvedi, Olusola Dehinsilu, Hussein Edduweh, Rodina El-Adawy, Emran Hossen, Bardia Mojra, and Jashmon Rana. A robust optimal control framework for controlling aberrant RTK signaling pathways in esophageal cancer. Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02033-0.

Eisenbaum:2024:SLS

[3586] Nathalie Eisenbaum and Nicolas Meunier. A stochastic lipid structured model for macrophage dynamics in atherosclerotic plaques. Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02029-w.

Meyer:2024:LST

[3587] Léo Meyer, Magali Ribot, and Romain Yvinec. A Lifshitz–Slyozov type model for adipocyte size dynamics: limit from Becker–Döring system and numerical simulation. Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02036-x.

Holland:2024:DBM

[3588] Barbara Holland, Katharina T. Huber, and Vincent Moulton. A distancebased model for convergent evolution. Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02038-9.

Otunuga:2024:TGP

[3589] Olusegun Michael Otunuga. Tumor growth and population modeling in a toxicant-stressed random environment. Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-023-02035-y.

Benaim:2024:WCP

[3590] Michel Benaim, Claude Lobry, Tewfik Sari, and Edouard Strickler. When can a population spreading across sink habitats persist? Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-023-02039-8.

Cronin:2024:PIP

[3591] James T. Cronin, Jerome Goddard II, Amila Muthunayake, Juan Quiroa, and Ratnasingham Shivaji. Predator-induced prey dispersal can cause hump-shaped density–area relationships in prey populations. Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02040-1.

Cordoni:2024:SMV

[3592] Francesco G. Cordoni. A spatial measure-valued model for radiationinduced DNA damage kinetics and repair under protracted irradiation condition. Journal of Mathematical Biology, 88(2):??, ???? 2024. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL

https://link.springer.com/article/10.1007/s00285-024-02046-

Gao:2024:VBD

[3593] Daozhou Gao and Linlin Cao. Vector-borne disease models with Lagrangian approach. Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-02044-x.

Borgqvist:2024:TPF

[3594] Johannes G. Borgqvist, Philip Gerlee, and Carl Lundholm. Turing pattern formation on the sphere is robust to the removal of a hole. Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02034-z.

Martignoni:2024:MLE

[3595] Maria M. Martignoni, Rebecca C. Tyson, Oren Kolodny, and Jimmy Garnier. Mutualism at the leading edge: insights into the eco-evolutionary dynamics of host-symbiont communities during range expansion. Journal of Mathematical Biology, 88(2):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02037-w.

Albani:2024:STE

[3596] Vinicius V. L. Albani and Jorge P. Zubelli. Stochastic transmission in epidemiological models. Journal of Mathematical Biology, 88(3):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-023-02042-z.

Kanzler:2024:SMC

[3597] L. Kanzler, B. Perthame, and B. Sarels. Structured model conserving biomass for the size-spectrum evolution in aquatic ecosystems. Journal of Mathematical Biology, 88(3):??, ???? 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02043-y.

daSilva:2024:AVS

[3598] Poly H. da Silva, Arash Jamshidpey, and Simon Tavaré. Another view of sequential sampling in the birth process with immigration. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-023-02041-0.

Wang:2024:CAB

[3599] Xiaoyu Wang, Adrianne L. Jenner, Robert Salomone, David J. Warne, and Christopher Drovandi. Calibration of agent based models for monophasic and biphasic tumour growth using approximate Bayesian computation. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-024-02045-4.

Ane:2024:ANU

[3600] Cécile Ané, John Fogg, Elizabeth S. Allman, Hector Baños, and John A. Rhodes. Anomalous networks under the multispecies coalescent: theory and prevalence. Journal of Mathematical Biology, 88 (3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-024-02050-7.

Djuikem:2024:IOC

[3601] Clotilde Djuikem, Frédéric Grognard, and Suzanne Touzeau. Impact of ontogenic changes on the dynamics of a fungal crop disease model motivated by coffee leaf rust. Journal of Mathematical Biology, 88 (3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-024-02053-4.

Liu:2024:CEM

[3602] Guodong Liu, Hao Wang, and Xiaoyan Zhang. On cognitive epidemic models: spatial segregation versus nonpharmaceutical interventions. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-024-02049-0.

Freingruber:2024:BRW

[3603] Viktoria Freingruber, Kevin J. Painter, Mariya Ptashnyk, and Linus J. Schumacher. A biased random walk approach for modeling the collective chemotaxis of neural crest cells. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-024-02047-2.

Wang:2024:RNR

[3604] Jing Wang, Hongyong Zhao, and Hao Wang. The role of natural recovery category in malaria dynamics under saturated treatment. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-024-02051-6.

Clarke:2024:TSM

[3605] Cameron Clarke and Stephen Pankavich. Three-stage modeling of HIV infection and implications for antiretroviral therapy. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-024-02056-1.

Li:2024:SPI

[3606] Bingtuan Li and Garrett Otto. Spread and persistence for integrodifference equations with shifting habitat and strong Allee effect. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-024-02048-1.

Joshi:2024:BEP

[3607] Badal Joshi and Tung D. Nguyen. Bifunctional enzyme provides absolute concentration robustness in multisite covalent modification networks. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-024-02060-5.

Pandey:2024:ACS

[3608] Sanjay Kumar Pandey and Ankit Prajapati. An analytical and comparative study of swallowing in a tumor-infected oesophagus: a mathematical model. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-024-02054-3.

Li:2024:GDT

[3609] Zhimin Li and Xiao-Qiang Zhao. Global dynamics of a time-delayed nonlocal reaction–diffusion model of within-host viral infections. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-024-02052-5.

Toth:2024:ISP

[3610] Kaitlyn Toth and Dan Wilson. The influence of synaptic plasticity on critical coupling estimates for neural populations. Journal of Mathematical Biology, 88(3):??, March 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-024-02061-4.

Bordewich:2024:QDB

[3611] Magnus Bordewich and Charles Semple. Quantifying the difference between phylogenetic diversity and diversity indices. Journal of Mathematical Biology, 88(4):??, April 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.springer.

com/article/10.1007/s00285-024-02059-y.

Alvarez:2024:TCM

[3612] Frank Ernesto Alvarez and Yannick Viossat. Tumor containment: a more general mathematical analysis. Journal of Mathematical Biology, 88(4):??, April 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-024-02062-3.

Pan:2024:EWK

[3613] Qin Pan, Min Lu, Jicai Huang, and Shigui Ruan. Effects of whaling and krill fishing on the whale–krill predation dynamics: bifurcations in a harvested predator–prey model with Holling type I functional response. Journal of Mathematical Biology, 88(4):??, April 2024. CODEN JM-BLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://

link.springer.com/article/10.1007/s00285-024-02063-2.

Ni:2024:NMI

[3614] Zhuxin Ni and Qihua Huang. A nonautonomous model for the interaction between a size-structured consumer and an unstructured resource. Journal of Mathematical Biology, 88(5):??, May 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-024-02071-2.

Hayes:2024:EVG

[3615] Wayne B. Hayes. Exact p-values for global network alignments via combinatorial analysis of shared GO terms. Journal of Mathematical Biology, 88(5):??, May 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-024-02058-z.

Han:2024:FES

[3616] Zhimin Han, Yi Wang, Shan Gao, Guiquan Sun, and Hao Wang. Final epidemic size of a two-community SIR model with asymmetric coupling. Journal of Mathematical Biology, 88(5):??, May 2024. CODEN JMBLAJ.

ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-024-02073-0.

Jing:2024:GAA

[3617] Shuanglin Jing, Ling Xue, Hao Wang, and Zhihang Peng. Global analysis of an age-structured tuberculosis model with an application to Jiangsu, China. Journal of Mathematical Biology, 88(5):??, May 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https:

//link.springer.com/article/10.1007/s00285-024-02066-z.

Hindle:2024:ESC

[3618] Ivy J. Hindle, Lawrence K. Forbes, Stephen J. Walters, and Scott Carver. The effects of spatially-constrained treatment regions upon a model of wombat mange. Journal of Mathematical Biology, 88(5): ??, May 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-024-02078-9.

Hancock:2024:BFR

[3619] Zachary B. Hancock and Daniel Stern Cardinale. Back to the fundamentals: a reply to Basener and Sanford 2018. Journal of Mathematical Biology, 88(5):??, May 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-024-02077-w. See [2761].

Zhigun:2024:MNL

[3620] Anna Zhigun and Mabel Lizzy Rajendran. Modelling non-local cellcell adhesion: a multiscale approach. Journal of Mathematical Biology, 88(5):??, May 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-024-02079-8.

Wu:2024:HBA

[3621] San-Xing Wu, Zhi-Cheng Wang, and Shigui Ruan. Hopf bifurcation in an age-structured predator-prey system with Beddington–DeAngelis functional response and constant harvesting. Journal of Mathematical Biology, 88(5):??, May 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-024-02070-3.

Mielke:2024:MED

[3622] Adam Mielke, Mads Peter Sørensen, and John Wyller. Memory effects in disease modelling through kernel estimates with oscillatory time history.

Journal of Mathematical Biology, 88(5):??, May 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL https://link.

springer.com/article/10.1007/s00285-024-02080-1.

Strube:2024:III

[3623] Laura F. Strube, Shoshana Elgart, and Lauren M. Childs. Infectioninduced increases to population size during cycles in a discrete-time epidemic model. Journal of Mathematical Biology, 88(6):60:1–60:35, June 2024. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416

(electronic). URL https://link.springer.com/article/10.1007/
s00285-024-02074-z.

Cheng:2024:MIP

[3624] Tianyu Cheng and Xingfu Zou. Modelling the impact of precaution on disease dynamics and its evolution. Journal of Mathematical Biology, 89(1):??, July 2024. CODEN JMBLAJ. ISSN 0303-6812 (print),

1432-1416 (electronic). URL https://link.springer.com/article/
10.1007/s00285-024-02100-0.